A Bibliography of Publications in IEEE Computer Graphics and Applications

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

27 March 2024 Version 1.135

Title word cross-reference

+ [AJD13, CTW^{+}10]. $100 [ZMR^{+}88]. 111 [Bli07a]. 1¯1 [Bli06b]. 2 [AP11, BBT^{+}13, CCF97, CMO^{+}99, GWF98, MD00f, MKFN87, SMWS88, SSOC01, UIL^{+}06]. 25 [Ang17]. 2 × 2 [Bli96a, Bli02b]. 3 [Abo99a, ASW93, AP10, Ano15-33, AA09, BRL^{+}17, BC87, BBF^{+}98, BGSC85, Ben99, BH16, BA13, Bli88c, BFB^{+}98, BCF^{+}08, BBT^{+}13, BKSS07, Bux05, CD12, CCG^{+}04, CPCS08, Cam17, CCF97, CRB^{+}02, CBB^{+}14, CT09, COK97, CMO^{+}99, DB03, Del99b, DGA^{+}97, DNS^{+}15, Die00, DY09, DWL^{+}92, EHBPG04, Enc14a, FYZ85, FSK07, FD03, FB09, FHKH06, FZ03, FTAT00, GWF98, Ger91, Gir87, GBT^{+}02, GKB89, HPG08, HdlRL^{+}13, HWY16, HTZ^{+}19, Han01, HGMT17, HF07, Hay95b, HJF06, IKMF12, IG98, JWL^{+}18, JAC^{+}08, KK02, KS07, KTA^{+}02, KB88, KP14, KSRS13, Kon12, KZC^{+}11, Kul09, KPL17, KDMH17, LaV08, LFV^{+}97, Law99, Lea96a, Lea96d, LJB^{+}13, LYO02, LSQ^{+}12, Mac97a, Mac13, MD99f, MD99e, MD99d, MD00f, MKFN87, MSF07, MGO^{+}16, Mas98, MVWB05, McG93, MFM^{+}16, MSH^{+}08, Mil03, Mun98, NTH02, NINT89]. 3 [NE93, NM08, NF91, NLvLM09, dSNRdLdSN16, Nur08, PW97, PSX13, PLS^{+}16, PJC88, Paq96, PT90, PTR^{+}98, Pie06, PVV^{+}03, PG93, PKJ^{+}07, Red01, RLOW13, RWB^{+}14, RR03, RB00, RDH^{+}02, RT07, RFN89, RBK^{+}13, RGCB07, SRS91, SPJ99, SS17, SSG95, SBMK20, SB16, SMWS88, SRL06, SRDT01, Shn03, SM04, Sim96c, Sin01c, SSOC01, SFL^{+}04, SD99, Spa16b, SQG^{+}05, SAHK13, Tau11a, TLL^{+}96, TMB^{+}13, THB^{+}90, TUH^{+}90, TH17, US17, UBS^{+}12, WOF96, WAH^{+}21, WMS92, WK94, WLL^{+}16, Wea99b, WWV^{+}10, WL88, WSKY09, Wor96b, WZQ^{+}19, YY16, YCB05, YZM^{+}11, YYOiT92, YY99, YWR09, YBL^{+}12, ZZ05, ZCS^{+}06, ZSK^{+}21, aE01, vHdP94]. 360^{∘} [DV21, MCGS21, PD21]. 4 [Bli03e, CMPA18, HLE04, KN07, MD99e, OWZ17]. 4 × 4 [YT85]. 6 [HHLL14]. = [CTW^{+}10]. ^{∘} [WW24]. _{2} [AGBC09]. C^{1} [CW87, CW88]. D [RSDP01]. G^{2} [Sch93a]. N [Gil94, HT98, Sch93a, MG04].

-D [HTZ^{+}19, SBMK20, WAH^{+}21, WZQ^{+}19]. -Dimensional [MG04]. -DOF [HHLL14]. -Draw [SRS91]. -Guided [FHH^{+}21]. -Sided [HT98, Sch93a]. -Step [Gil94]. -Year [Ang17].

1 [Arn14a, Enc23a, WS09a]. 1/2D [WW07]. 12-bit [WWC91]. 17-inch [MD99e]. 19 [Ano20u, KCB^{+}23, OBO^{+}23]. 1996 [Lea97d].

2 [Arn14b, Bro82, Enc23b, MD99f, TT84, WS09b, WW07, WWL^{+}21, YMK^{+}12]. 2-1 [WW07]. 2-1/2D [WW07]. 2-D [TT84, WWL^{+}21]. 2-Layer [AAB^{+}23]. 2.0 [MD99f, MD99e, MD99d]. 2.5 [FM22]. 2.5-D [FM22]. 2000 [MD99d, XPH05]. 2003 [SP04b]. 2007 [PGS^{+}08]. 2009 [SR10]. 2010 [Ano10b, Ano10c, DPL^{+}11]. 2011 [OKET12]. 2012 [Kas12a, SKS^{+}13]. 2013 [LZF^{+}14]. 2014 [Ano13b]. 2016 [GGS^{+}18]. 2018 [Ano18e]. 2019 [Ano19p, Ano19-39, Ano19-40]. 2020 [Ano20t, Ano20-36, Ano20-37, TPR^{+}22]. 2021 [SLC^{+}23]. 2022 [Ano21-30, FGCB23]. 2022/2023 [FS23]. 2023 [Ano22a, Ano23a]. 21st [Ann17, Ano20r, Ano22-72, Ano22-73, Ano22-74, Ano23-74, Ano23-75, Ano23-76, Ano23-77]. 240GTX [Ake89]. 25-Year [MLJ19]. 2D [WW07].

3 [Bal98c, Bar84, BU23, CCV85, CE83, DL22, DPL23, FZY84, FPT^{+}82, FPI84, HU83, Ike84, LSFW83, LCG83, MTT83, OATO85, RSDP01, aKTRL^{+}21, TPA24, TT84, Zha22b, ZLK^{+}22b]. 3-D [Bal98c, Bar84, BU23, CCV85, CE83, DL22, DPL23, FZY84, FPT^{+}82, FPI84, HU83, Ike84, LSFW83, LCG83, MTT83, OATO85, aKTRL^{+}21, TPA24, TT84, Zha22b, ZLK^{+}22b]. 3.0 [MD99d]. 31 [Ano15c]. 3D [PM86]. 3DUI [FKK^{+}10].

4 [Lea96a, NCO03]. 4.0

[PZD^{+}18, Spa16b, Sto15, ZPL^{+}23]. 450-Year-Old [Abo99b]. 4D [Ake89]. 4D/240GTX [Ake89].

5 [MD99e].

6 [SLSP22]. 6-D [SLSP22]. 64-Bit [GKB89].

7 [MD99f]. 777 [DSW07].

8 [Bar97a]. 8-Bit [Bar97a]. ’85 [MTT85a, KR86]. 86 [Gre86a]. 88 [Ano88b, Pea88].

94 [BDDH95]. 95 [Hay96d, IS95]. 96

[Hay96a, Lea97c]. 97

[Bal98b, GH97, JLC98, Lea97d].

A-Changing [Rib98]. A-Desk [STJ^{+}20]. Aboard [Tho02a]. above [WB14]. Abstract [Rou09, SR16, Sin02b, Sin05d, Sin10a, Sin15a, Sin15b, Zha07]. Abstraction [BWD^{+}19, Sin16e]. Abstracts [Hay96b, Sin08b]. Academic [PGV20]. Accelerate [IO07]. Accelerated [FTI86, RNNTD14, SGG^{+}22]. Accelerating [MSU^{+}16, RAPL98, ZPL^{+}23]. Acceleration [JAPF07]. Accelerator [HG86, MD99d, MQR99, RLV^{+}16]. Accelerators [JAC^{+}08, Pin91]. Access [AS96, HBW^{+}24, RD90]. Accessible [SKJ^{+}13]. Accidents [NdSM^{+}16, Sin22d]. According [Pen10]. Account [Hec86a]. Accumulation [HW95]. Accuracy [BWP^{+}16, WF93]. Accurate [ANS15, BS96, LKV90, LTG92, LX04, SK97, SPA^{+}16a, Tur88, Gre91a]. Achieving [ARL^{+}10]. ACM [Ano88a, Ano19a, KWJ23]. ACM/IEEE [Ano19a]. Acoustic [LSB21, MOD00, NTM^{+}01, Sch91]. Acoustical [TCE^{+}02]. Acquire [PCJ^{+}16]. Acquires [Nea89a]. Acquiring [AXP07, NK14]. Acquisition [GJ13, dSNRdLdSN16, SKW^{+}12]. Across [FKC^{+}05, MCGS21, Sin01d]. Act [Hay96a]. Action [MP89, NTS^{+}07, VT24]. Action- [NTS^{+}07]. Actions [EBT98]. Active [AXP07, FH84b, GBT^{+}02, JRM17, OZH00, WZQ^{+}19, KÇC^{+}01]. Active-Matrix [FH84b]. Activities [EGR94, KSK21, Kah94]. Activity [BNA^{+}22, BD21, KFBI22, PD09, SAO^{+}95]. Activity-Driven [PD09]. Actor [ARL^{+}10]. Actors [BRB17, MTT87b, VMJT96, MTT91]. actually [GP91a]. Ad [Ano18d, Ano18f, Ano18q, Ano18n, Ano18z, Ano18x, Ano18-27, Ano18-28, Ano18-36, Ano18-35, Ano18-37, Ano18-38, Ano18-43, Ano18-41, Ano18-50]. Adaptation [JOG^{+}17]. Adaptive [BM03, BB00, BBB08, FR00, HW90, KS98, KS97, LGM13]. Adapts [Lea96a]. Adding [FPT^{+}82, Wil13, MSED23, Swi10, WKB11, Won08, Zac84]. Additions [Tau10b]. address [Ano96h, Ano97g]. addressability [Kla93]. Addressed [Ma23]. Addressing [LKV90]. Adds [MD99d, Wri98, Whi96a]. Adjustable [LJ15]. Adjustment [FM90]. Administration [DBV^{+}08]. Adolescents [RRV^{+}15]. Adopts [Lea96a]. Adorning [JW01]. Ads [Ano18o]. Adult [PPA14].

Adults [LTD16, RMR^{+}19]. Advanced [Ano15b, BDM02, CB86, HW82b, Kas10a, Kas10b, Kas11a, Kas11b, RDRS10, SAB^{+}23, Sto15, VPVG21]. Advances [ABB^{+}01, CGJ^{+}17, Kri06, LaV07, Lea96e, Ovi03, Yan85, Bal98c]. Advancing [Agr21a, Law99, Wor95]. Adventure [Abo99a]. Adventures [Ros21]. Adverbs [RCB98]. Adversarial [CCT^{+}20, LWZY19]. Advertisement [Ano14s, Ano14q, Ano14r, Ano14-29, Ano14-31, Ano15c, Ano15e, Ano15t, Ano15u, Ano15y, Ano15x, Ano15-35, Ano15-34, Ano15-36, Ano15-37, Ano15-38, Ano15-43, Ano15-44, Ano16d, Ano16g, Ano16r, Ano16s, Ano16t, Ano16u, Ano16v, Ano16y, Ano16x, Ano16z, Ano16-27, Ano16-28, Ano16-29, Ano16-36, Ano16-38, Ano16-42, Ano16-43, Ano16-39, Ano17a, Ano17c, Ano17p, Ano17-27, Ano14c, Ano14t, Ano14-27, Ano14-28, Ano14-30, Ano14-32, Ano14-39, Ano15d, Ano15l, Ano15w, Ano15-33, Ano15-39]. Advertiser [Ano95a, Ano95b, Ano95c, Ano95d, Ano96b, Ano96c, Ano96d, Ano96e, Ano96f, Ano96g, Ano97b, Ano97c, Ano97d, Ano97e, Ano97f, Ano98a, Ano98b, Ano98c, Ano98d, Ano98e]. Advertiser/Product [Ano95a, Ano95b, Ano95c, Ano95d, Ano96b, Ano96c, Ano96d, Ano96e, Ano96f, Ano96g, Ano97b, Ano97c, Ano97d, Ano97e, Ano97f, Ano98a, Ano98b, Ano98c, Ano98d, Ano98e]. Adviser [FPH^{+}06]. AEA [MD99e]. Aerial [FZ03, WPL22]. Aerospace [RBW05]. Aesthetic [DPL23, OF14, OC14, ODW15, RR95, WO13, Wor96a]. Aesthetically [PKM18]. Aesthetics [DL22, EH16, GGB23, GW08, MIA^{+}08, RFS08, ZNMT15, WJ93]. Affect [KKG^{+}13, Whi95d]. Affecticon [YL11]. Affective [MSU^{+}21]. Affectively [ALD^{+}06]. Affects [DPA^{+}11]. Affine [CT09]. Affine-Transformation-Invariant [CT09]. Africa [LK08]. After [BB95]. Again [Lea96c, Lea96d]. Against [Sin11a]. Age [Ano16a]. Agenda [IIH^{+}13, TC06, XAJK^{+}15]. Agent [CEMD22, CKJC12, GD13]. Agent-Based [CKJC12, GD13]. Ages [MMGG12]. Agglomerative [XJ94]. Aggregated [LHFL20]. Aggregation [XZZL19]. AGP [JM95]. Agriculture [PCJ^{+}16]. Ahead [Add95, BAB^{+}23b, WCG^{+}11]. AI [BM24, BABB^{+}21, EAM22, EKS22, PJC24, VT24]. Aid [Han00a, SHMN09, UTMI11]. Aided [BC87, BDM02, BH83, CF93, CLC^{+}96, DP82, HLC11, Hur82, KSM85, Lic84, Myk88, NTS^{+}07, Nie92, RGA^{+}83, SM86, SG86a, Bar81b]. Air [AHDG00, Bro89, GSFT16, RGG19, SK05, ZYL^{+}17, Sim95d]. Air-Quality [ZYL^{+}17]. Air-Traffic [AHDG00]. Airborne [CHC^{+}97, Cro96]. Aircraft [BSA^{+}04, DFP^{+}11, Tur02a, Bro89, Sim94c]. Akira [Nea89b]. Alchemical [Sin18a]. Alert [CAG^{+}06]. Alerts [FAL^{+}06, SZZ^{+}18]. Alfred [Shn20]. Algebra [Bli04b, DM02b, MD02e, SZMJ12, ZZ05, GJ93]. Algebraic [Fre84, SG86a, Sed90b, Tau94, Sed90a]. Algorithm [BB00, BB84, Cha88, DW87, Ell94b, Gal89, GG83, HB97, KKGH90, Lor20, McL83, MW82, Nap05, PO84, PB85, RR85, SM86, Sta17, Van84, YT84, AM91, Bli93a, Car87, PB87, Rou87, Sug92, Wan93]. Algorithmic [Sin21f]. Algorithms [BFP86, Cro87a, FM85, FIB^{+}20, Gil94, HHS^{+}92, IFH^{+}03, LKPM19, Lue01, Mam89, MD95b, Neu94, OKM84, OKKP13, Pan90, SR93, SL00, THB^{+}90, UF11, WD15b, WD15c, WPF90, Wri90, XZWH14, YKFT84, SW88]. Alias [Hay95b]. aliasing [Bli89c]. Alice [UVa95]. Alignment [RMP^{+}16]. Alis [Kub84]. Alive [Sin10b]. Alliances [Sin21e]. Alpha [SB86a, Bli96b]. Alternate [Lec17, Smi87]. Alternating [HM85]. Alternating-Pair [HM85]. Alternative [II15, Mil86]. Altia [MD99d, MD00f]. Altitude [PCJ^{+}16]. am [Bli88e]. Amazing [GP93b]. Ambient [Gol00, GSF08, SKUT^{+}10]. Ambiguities [Bre87]. America [Hay95b]. Among [HPX^{+}21, IDJW19, LWS98]. Amorphous [FH84b]. Amplification [Coh00]. Amputees [VCBS97]. Analog [Tho84]. Analyses [WSW08]. Analysis [AY16, AESF21, AVV20, AOH^{+}14, BQP^{+}19, BFR21, BHM^{+}22, BFH86, BSM^{+}15, BP06, BHS^{+}24, BG82, CMSEM08, CWB02, CTW^{+}10, Cro18, DNS^{+}15, DFL^{+}18, DL16, Dwy16, EKM^{+}22, FJKTX19, FTAT00, GGG08, GLRK06, GGS^{+}18, Gro83, GZM97, HYL^{+}22, HAH^{+}23, HT04, HLW13, Hua21, JS10, JBH^{+}10, KFBI22, Kos23, LPZ12, LHZ^{+}21, LPMAA^{+}22, LTB16, MD99e, MRB19, MTN18, Mil88b, MNC^{+}22, PA18, RFA^{+}22, RGL^{+}06, RMDT96, RLV^{+}16, SFD16, SMG^{+}21, SGG^{+}22, She06, SJB^{+}16, SJSK19, SGEF^{+}12, SW22, Swi10, TMWJK04, TA85, TLS^{+}14, UBS^{+}12, VSLG22, VFLJ02, WYG^{+}12, WBS^{+}05, Wil87b, WDK17, WHC^{+}14, Woo85, XSN^{+}20, Yan01, YBL^{+}12, ZGC14, Zha22d, FM87, GT93, LK95a, RBOV94, UO91]. Analysts [KBHP14]. Analytic [CS85, CPG^{+}16, EI17, MRB19, SP15, XAJK^{+}15, ZHZ^{+}15, JRM17]. Analytics [AAAW22, Ano14-30, Ano18c, BHPS12, BQP^{+}15, BQP^{+}19, Cha09, CZGR09, Che08, CCR10, CTW^{+}10, CP13, CL18, DFL^{+}18, ENDAD13, EE12, EBN13, End14, ECNZ15, EI17, FFN19, FDC14, GCS^{+}12, GMR^{+}23, HPX^{+}21, KCW^{+}17, KW15, KVHD17, LBDM19, LAC^{+}14, LRS12, MZC^{+}21, MMN^{+}21, MGO14, MHR^{+}09, MGA^{+}21, MLJ19, MGN^{+}11, OBO^{+}23, PGS^{+}08, PGS09, QC15, Ren15, RBD14, SBCvdS04, SMP^{+}19, SZMJ12, SLC^{+}23, TC06, TC15, TPR^{+}22, WT04, WLL^{+}09, WCG^{+}11, WSP12, WSJ^{+}12, WHC^{+}14, WKY23, WDC^{+}23, YCB^{+}22, ZY18, ZYL^{+}17]. Analyze [BKV^{+}10, Koz12]. Analyzing [AESF21, BGM^{+}07, CKJC12, EG98, Gai04, KSK21, PGM19, SRE98]. anatomic [RKRW87]. Anatomical [HBR^{+}92a, WLC^{+}12, HBR^{+}92b]. Anatomy [BWHR99, Wil97]. ancient [Cra91, YYOiT92]. Andrew [BG96, Gla02a, Gla96e, Gla96b, Gla96a, Gla96d, Gla96c, Gla97a, Gla97e, Gla97d, Gla97c, Gla97f, Gla97b, Gla98e, Gla98b, Gla98a, Gla98f, Gla98c, Gla98d, Gla99d, Gla99f, Gla99a, Gla99e, Gla99c, Gla99b, Gla00a, Gla00e, Gla00f, Gla00b, Gla00c, Gla00d, Gla01b, Gla01f, Gla01a, Gla01c, Gla01d, Gla01e, Gla02e, Gla02d, Gla02f, Gla02b, Gla02c, Gla03a, Gla03b, Gla03e, Gla03f, Gla03c, Gla03d, Gla04a, Gla04b, Gla04e, Gla04f, Gla04c, Gla04d, Gla05a, Gla05b]. Andy [De 11]. Aneurysm [MGW^{+}18]. Aneurysms [GGG08]. Angeles [Wor94]. Angels [BSHB01]. Angiograms [WOF96]. Angkor [CC20]. Angle [Bar81a, GXMW15]. Angle-Preserving [Bar81a, GXMW15]. Animal [Gir87, MR98, Wea99a, WAB^{+}13]. Animals [Wil97]. Animated [CCM15, Gir87, Kop03, LAM01, PFS03, SADC02]. Animating [FYZ85, KB96, Sin01b, MTT91]. Animation [AH95, AWM10, iAH03, Ano98g, Bad82, BL14, BHS11, BJC18, CBD^{+}93, CHP21, CE83, CMPA18, CPAV^{+}17, DiL15, EMTTT98, FSB82b, Gla88a, HFZ^{+}13, HE82, HEW17, HKSK17, JLS15, KMTM^{+}98, hKKSK17, Kla89, Kon12, LYO02, MD99f, MD99e, MD99d, MAB05, MTT83, MTT85b, MTTF85, MTT85c, NP17, NK85, PHE21, PSJ87, PJC88, Par82, RPM17, Sin01a, Sta17, SH07, SBS10, TNTM07, USW17, Wei23, Wil87b, WS90b, YM95, Zel82, Csu91, FSB82a, GS88, Bad00, DiL15]. Animations [CZP^{+}10, SC12, Ano98h, Ano98i]. Animators [SL11, Str90]. Animatronics [AJD13]. Anime [Kri06]. Animengine [NK85]. Anisotropy [IMS03]. AniViz [AWM10]. Annals [Ano21i, Ano21j, Ano22k, Ano22l, Ano22m, Ano23j, Ano23k, Ano23l, Ano23m, Ano24d, Ano24e]. Annotated [NVC92]. Annotation [BBDZ08, LHZ^{+}21, SJB^{+}16]. Announced [MD99f]. Announces [MD99f, MD99e]. Annual [Ano96a, Ano97a, Ano98f, Ano01, Ano02, Ano03a, Ano04a, Ano05a, Ano06a, Ano08a, Ano10a, Ano11a]. Anomalies [Chr83, Ala92]. Anomaly [ZHZ^{+}15]. Answer [JSDL21]. Antarctic [ASP^{+}21]. Antennas [CWGC90]. Anthropometric [Doo82]. antialiased [FG88a]. Antialiasing [Bar91, BBC^{+}87, Cro81, LJR96, Mam89, Max90, WEWL00, NINT89]. Anything [Cai20]. Aortic [GGG08]. Ape [Bar04]. Aperiodic [Gla98a]. Aperture [BKS19]. Apex [Fei85]. API [AJH^{+}12, SD99, Zhu12]. Apollo [NL84]. Apparatus [Wea99b]. Apparel [Paq96, Bre96a]. apparent [Kla93]. Appearance [OKKP13, PR12, RP95, Sto01, VB24]. Apple [CWZ99, Ano19b, Ano20a]. Apple-Sixth [Ano20a]. Apples [RSZ96]. Appliances [FJHW00, KRL97, WB00a, WB00b]. Application [Agr18a, CPG^{+}16, DNS^{+}15, Eng86, EH16, FJ87, GWA84, LS08, LSB18, MN97, RGRLFM21, RMR^{+}19, Spi86, WYM10, WCE^{+}17, WK85, RKRW87, UFK^{+}89]. Application-Driven [LSB18, WYM10]. Application-specific [Eng86]. Applications [Abo98a, Abo99b, Abo98b, AGBC09, Ano92c, Ano93f, Ano93g, Ano94c, Ano04a, AJH^{+}12, DFQ12, BFW99, CWRF05, Cha85, CP89, CDBL83, CHC^{+}97, DB96, Del97, Del98b, Del98c, Del99a, Del00b, DM02b, Ear98, EM09, FE97, GE97b, GSW24, GBT^{+}02, GPR94, Gre86b, GP93b, GM12, GDY^{+}11, HJ03, Han00a, Han00b, Han01, HU83, Hoe12, KHP^{+}11, Kri00, Kri01, Kro02, KF05, LTJ18, Lea97a, Lea97b, Lic84, ML16, MD99d, MD02e, Mil99a, Mol18b, MYK^{+}08, MP23, NL84, PJC88, Paq96, PSE^{+}09, Pot98, Pot00, Pot04, PW14, RM89, RXB09, RR96, RLG^{+}02, Rok96, Ros97, SW88, Sch08a, SMWS88, Sim95a, Sim95b, Sim95c, Sim95d, Sim96b, Sim97, Ste18, Swe01, TCB03, UMO82, WAB^{+}05, WWF00, War85, Whi96a, Whi96d, Wor94, Wri97, WDC^{+}23, Yan01, ZZ99, aE01, vdEAA^{+}23, Goe96, Rei91, Sim93a]. Applications [Ano03a, Ano21k, Ano23x, Ano23y, Ano23z, Ano23-27]. Applied [BC87, HMHM18, MC85, MD18, TF01, MNR^{+}91]. Apply [WCE^{+}17]. Applying [BK06, EH16, Rod14]. Apprentice [Sim96c]. Approach [AFYC03, BFK^{+}84, CC20, CGIB86, CM16, DJM13, DiL15, DYS^{+}07, HWP^{+}19, HPX^{+}21, Hed84, HROIK16, LBS^{+}03, MPDR14, PD09, PRN15, RPM17, SC89, SPMS04, Shi17, WS04, YS83, YCB^{+}22, Zha96b, FG88a]. Approaches [BCPS22, Gol83b, HYN03, Mil86]. Appropriate [Ano14a, Ano14b]. Approximate [FDK12]. Approximating [Ale21, PT03]. Approximation [ELK97, SF92]. Approximations [RPHW11, Tur88]. Apps [MRL13, Nav04]. Arbitrary [CW88, KSK00, lLK84, PO84, Sar90, CW87]. Arc [Gal89, GP90]. arcades [Bal98c]. Archaeological [PVV^{+}03, Sim97, VIK^{+}02, VFLJ02]. Archaeologist [PKT01]. Archaeology [CR02a, Ste91]. Archeoguide [VIK^{+}02]. Archimedean [OZH15]. Architects [NAB^{+}23]. Architectural [DD84, Mil89, RAPL98, YWR09, FJ87]. Architecture [AAB^{+}22, Ano19g, Ano22n, BHPB^{+}12, CG85, Eng86, FL81, Gol84, KB88, Kel84, KKAM84, KM88, MYPT09, MRD06, NL84, SMWS88, WMS92, Wei84, Wil86]. Architectures [CPA^{+}10, Fri20, KB92, MSU^{+}16, SB03a]. Arctic [Sin18b]. Area [Cro87a, GDY^{+}11, Pic92, Rie93, de 86]. Areas [Sar82]. Argentina [Han00b]. Arguments [SEAKC21]. ARIES [CRD^{+}18]. Arithmetic [Gol01]. Army [Del00c]. Arsenal [WF02]. Art [Ano16a, CS15a, CS15b, CF16, CHH24, CPAM08, CR02a, Cha98b, CRD^{+}18, Csu91, Dev13, Die85, Die00, Gla04c, Gla04d, Gla05a, Hag90b, Hay95a, Hay95b, kHHBEH13, JSH16, JW16, Kri06, LB20, Mar07, MR99, MS24, Mei85, MGN^{+}11, ND05, Pal85, PJC24, Pri00, RGRLFM21, RSZ20, SGE^{+}18, Sin05a, Sin08a, Sin08e, Sin09c, Sin11c, Sin13c, Sin14b, Sin15b, VCS19, WCS18, Whi95b, Wor96a, Yea97b, vOW15, Hag91a, Hag94, Mac97a]. Art-Science [CS15b]. Arthritis [ZCBM14]. Article [Ano23a]. Articles [Ano14j, Ano14m, Ano14k, Ano14l, Ano15n, Ano15o, Ano15p, Ano15t, Ano15q, Ano15u, Ano15r, Ano15s, Ano16r, Ano16s, Ano16t, Ano16u, Ano16n, Ano16o, Ano16p, Ano16q, Ano16v, Ano17p, Ano17j, Ano17k, Ano17l, Ano17m, Ano17n, Ano17o, Ano19n, Ano20j, Ano20k, Ano20l, Ano20i, Ano21y, Ano21v, Ano21w, Ano21x, Ano21u, Ano22z, Ano22-27, Ano22x, Ano22y, Ano22w, Ano22q, Ano23v, Ano23u, Ano23w, Ano23s, Ano23t, Ano24g, Ano24f, Hay96b, ISMGG22]. Articulated [BMW87, Dai89, KB82, Mac90, SS17, WH97, Wil87b]. Artifacts [SDM22, Bal98c]. Artificial [Ano23-33, CHH24, Hag91a, How00, LSVB21, MSU^{+}21, MS24, SCS22, Sin08c]. Artist [Hay97, Sin19d]. Artistic [CMPA18, KKVL05, Lai01, Mar07]. Artistry [Sin04e]. Arts [War00, FTI86]. Artwork [KJA^{+}18]. ASCI [Hee99]. ash [RG95]. Ask [SS20]. Aspect [JW01]. Aspects [BA13, CCP82, Eic96a, Eic96b]. Asperger [RMR^{+}19]. Assemblies [Arb90]. Assembling [PKT01]. Assembly [JJW^{+}99, KCMK20, RA06]. Assess [BWR24]. Assessing [PVML96, WF93, FG88a]. Assessment [Ano14b, CKPB16, EKM^{+}22, LLML16, RV82, WGZ^{+}18, WCE^{+}17, Ano14a]. Assets [Agr19]. Assignment [LLL^{+}13]. Assignments [AVV20]. Assist [AGP^{+}22, GJM18]. Assistance [BFP86, KCMK20]. Assistant [LBW^{+}17, Sim96c]. Assisted [AKME^{+}90, Bar91, CH10, CPG^{+}16, CXPG19, KTD10, KW10, KY97, LMA^{+}17, SRS83, SW22, WBS^{+}05]. Assistive [RBMS17]. Assistive-Technology [RBMS17]. Associate [Tau10a, Tau13a]. Assurance [Got82]. Asteroid [LHFL20, MPDR14]. Astronauts [NTB^{+}20]. Astronomical [BAE^{+}18, MUN^{+}10]. Astrophysical [FOI^{+}07]. Astrophysics [NPH^{+}09]. AstroSim [NPH^{+}09]. Asymmetric [SSO^{+}12]. atla [HBR^{+}92b]. Atlas [HBR^{+}92a]. Atlases [WLC^{+}12]. ATM [LFV^{+}97]. Atmosphere [GSFT16]. Atmospheric [RGG19]. Atomic [CHY99]. Attend [Bli95d]. Attention [HBB20, RS14]. Attractions [JMT^{+}01]. Attractive [HF18]. Attractors [PS88]. Attribute [Mil07]. Attributes [Mil07]. Audience [SPS99, Sch93b]. Audiences [CS17b, LCI^{+}20]. Audio [BRB17, KB09, LTB22, PHAS20, Sin01a, SHS^{+}21]. Audio-Visual [BRB17]. Audioptimization [MOD00]. Auditory [GDS^{+}18, LSV^{+}02]. Augment [Ano16b]. Augmented [AKFPW20, ABB^{+}01, BN95, BPB14, BH06, BVD19, CML^{+}23, DFP^{+}11, DJM13, FLSG98, GSN09, GSSO05, GDY^{+}11, HROIK16, HBG04, IHL18, JAM^{+}05, JBBL02, KYK^{+}24, KBHT06, Liv05, LRZ23, MSWT14, MSDS23, MSED23, MIEL00, Nav04, NTS^{+}07, NTB^{+}20, Pie06, PKN^{+}22, PXC02, RBW05, RLG^{+}02, Ros00a, SPT^{+}15, SSW^{+}07, SC21, SDPW20, VIK^{+}02, WS09a, WS09b, YNA99]. Augmented-Reality [BPB14, DJM13]. Augmenting [Bim06, JMC^{+}02, MRL13, TN10]. Australian [IS95]. Authentication [CT09]. Authoring [CRB^{+}02, IM11, KSRF11, MD99f, ZPL^{+}23].

Authors [Ano96k, BSS20]. Auto

[BB00, PHE21, BF93]. Auto-Adaptive [BB00]. Auto-Regressive [PHE21]. Automated [CH81, DLN05, Fei85, GGG08, HSK02, OATO85, Swi10, WJ93]. Automatic [ASW93, DD19, FM90, GS87b, JLW08, KH98, LHZ^{+}21, MOP89, MSF07, MGW^{+}18, MTS^{+}20, Rot03, Wil87a, YP08, ZGC14, ZLMY19]. Automatically [QNT02, Shi83, SK91]. Automating [Kru95]. Automation [May83]. Automotive [BFBK00, RBW05, SWS07, Ste07]. Autonomy [LJ15]. AutoPlant [MD99f]. Autoscan [BFB^{+}98]. Availability [KFH^{+}18, SSOC01]. Available [MD99f, MD99e]. Avatar [PMR^{+}09]. Avatars [CLP18, OBG^{+}23, PHE21, SPS99]. Average [GH86]. Averaging [CP89]. avionics [PG93]. AViSSS [EM09]. Award [Ano16e, Ano18p, Ano18q, Zha22c, Ano18d, Ano18n, Ano19a, Ano19c, Ano19d, Ano20b, Ano22a, Ano22-81, Ano23-84, Ano23-85, Ano23-86, WW95]. Awards [Ano88a, Ano16c, Ano17a, Ano19m, Ano20v, Ano21-66, Ano22b, Ano23q, Ano23r, Ano23-64, FS23, Fu24, Hay95b, Lea96a, Lea97d, Tau12c, Zha22a, Ano17-31, Ano20-51, Ano21-67, Ano21-68, Ano23-65]. Aware [Gas16, GD11, PGM19, TSLX12, TCB03, dHPP10]. Awareness [FGW00, LGZ04, PMR^{+}09, RDDT97, SLC^{+}23, SvW09]. away [Sim94a]. Axis [KPN05]. Axonometry [Kri00].

B [CP92, GS94, HN85, PT89, PT03, Rie22, RSR83a, RSR83b, SR85, SB86b, Til83a, WWL^{+}21, YSSP91]. B-Spline [HN85, PT89, RSR83a, RSR83b, SR85, SB86b, WWL^{+}21, YSSP91]. B-Splines [GS94, PT03, Rie22, Til83a, CP92]. B. [Ano19c, Ano19d, Ano20b]. Babbage [Ano18d, Ano22a]. Back [Ano16f, Bli03c, Bli07b, Enc14a, Enc14b, FGR85, LCS21, She05, Sin08d, GC91]. Back-To-Front [FGR85]. Back-to-School [Ano16f]. Backface [Bli93a]. Background [Ker13, LSFW83, Mac85]. Backpropagation [CPM^{+}18]. Bacteria [Sin23b]. Bad [WSS23]. Badges [PGV20]. Bag [Sin11d]. Baggage [TH17]. Bake [PLB^{+}01]. Bake-Off [PLB^{+}01]. Balancing [Whi94b]. Ball [LCB^{+}23]. balls [MG92]. Band [Iga19]. Bands [Sin01b]. BandWeavy [Iga19]. Barcelona [Cab89]. bare [KmKK^{+}18]. bare-hand [KmKK^{+}18]. Barium [SKS^{+}13]. Barnard [BC19]. Baroque [Sin16c]. Barriers [MRL13]. Base [UMO82]. Baseball [Gol01, LOC^{+}16]. Based [AWM10, ACL^{+}12, AS11, Ano16-43, BN95, BNA^{+}22, BSHW07, Ben99, BTM17, BF93, BMS^{+}13, BKSS07, CFF01, CRF^{+}21, CML^{+}23, CEMD22, CAP^{+}05, CM92, CMP95, CXW^{+}20, CPN09, CW96, CSI86, CKJC12, CCM15, CXPG19, De 89, DM02a, Deb02, DiS14, DNBC16, DD04, ERH^{+}11, EJ98, FH19, FH21, FPNVHH17, FJP02, FGR85, FPW^{+}00, GA97, GWW86, dOGDM^{+}19, GSF08, GD13, GHCW21, GHQ04b, HdlRL^{+}13, HFZ^{+}13, Her03, HZ11, HROIK16, HJF06, HNC^{+}04, HH01, HLW13, HKSK17, IZ07, IWF97, VCR00, JKKM^{+}03, JCM07, Jos11, KS20, KVM^{+}05, KS07, KB88, KPN05, KSD04, hKKSK17, KKAM84, KMM^{+}01, LW94, Lat98, LMWC17, LD08, LGW^{+}12, LZS04, LLZ13, LZR^{+}21, MBSD20, MSWT14, MSRMH09, Mil88b, MOD00, Mor05, NTH02, NMRI06, NL02, OF14, OC14, ODW15, PHE21, PSX13, PLS^{+}16, PPK05, PPVH23, PHAS20, PG04, RL07, RC85]. Based [RS14, RRV^{+}15, RS97, RSE99, RBE08, RGCB07, RBD14, SS17, SPMS04, SS97, SR16, SBMK20, SSB06, SWS07, SI94, SADH23, SH07, Ste84, SL00, SLSP22, SSMH03, SOM^{+}10, SSZP15, TLC02, TW88, TM88, TS13, TWB^{+}07, UCK^{+}04, Van10, VBHS11, WO13, Wei85, WWW^{+}23, WWL^{+}21, WW24, XNSW19, YST84, YK95, YL11, YSW^{+}23, ZGC14, ZHZ^{+}15, ZYJ^{+}21, ZPG^{+}21, ZSK^{+}21, Coq87, DG99, FB93, GGS^{+}18, Han00b, HJS00, HZB92, HBR^{+}92b, HBR^{+}92a, KmKK^{+}18, LK95a, MGW^{+}18, PB87, RN94, SKK91, SZ94, Sug92, WS92]. Bases [Ken81, PSJ87]. Basics [BEW91]. Basis [MBV17]. Basketball [LTB16]. BAT [WJ88]. Battle [Kri03]. BattleBridge [SBSW96]. Bay [WLVL^{+}96]. Bayou [Hay96a]. Be [Abo00c, Ano16x, Ano18s, JB02, Tac16, Sim96a]. Bead [TNTM07]. BEAMES [DCCE19]. Beaming [KND^{+}18, SSO^{+}12]. Beautification [KJR^{+}18, She22]. Beautiful [OF14, OC14, ODW15, OZH15, OWZ17, WO13]. Beauty [Cha99, Pru90, Sin07a, Sin15f]. Become [SF20]. Becomes [Sin11f]. Before [HS01]. Beginning [CMBZ00, Gol12]. Beginnings [Enc23a]. Behavior [BKM16, DEN13, KSRF11, MGA^{+}21, OKET12, PB06, SF20, SSB06, TSF^{+}98, WZM^{+}11, WAPW06, WGP97, AES93]. Behavior-Authoring [KSRF11]. Behavioral [WS90b]. Behaviors [BCI08, SQG^{+}05, WYG^{+}12]. Behind [Enc14c, OCS20, Sin06f]. Being [DeF00, SN86]. Believability [BRD17]. Believing [SH06, SP06]. Bell [Hay96c, TCE^{+}02]. Belle [DPG18]. Belle2VR [DPG18]. Benchmark [LAM01]. Benchmarks [MSC89]. Bending [KGK94]. Benefits [BBC^{+}87, Lea96d]. Bert [Ano08b]. Bertin [ZK10]. Best [Ano19-45, Dom12, Fol14, Sin12a, ZM20, Zha22a, Ano18-29]. Beta [BD85, Bar93, Gol86, GU86]. Beta-spline [BD85, GU86]. Beta-splines [Bar93, Gol86]. Beta2 [BD85]. Beta2-spline [BD85]. Better [Ano16-37, Ano16-38, Ano17c, Bro15, DeF00, Del97, FGCB23, HBB20, Shn03, Ano17b, Ano17z]. Between [KPTS89, LPMAA^{+}22, Pra84a, Sin06f, SO95, Smi87, WFOP85, BKP01, HH08, Mol18b, RAGS01, Rhy03, TYK01, YP08].

Beyond [EBN13, FZH^{+}13, Han00a, Kos23, Kra06, LS04a, NTB^{+}20, Pac04, RWB^{+}14, RR92, Ste18, Sto08, Wea98a, WWV^{+}10, ZK10, vD00, EBBS00, RRB^{+}14, WB14]. Bézier [ABB90, SF92, Pie87a]. Biases [EAM22]. Bibliography [MTT85b, MTT87a, NVC92, Bar81b]. Bibliometric [Ano14b, Ano14a]. Bicubic [PK87, PB87]. Bicycle [HS01]. Bidirectional [CPK06, RZSP04, XZZL19]. Big [Ano14-30, Ano16-28, Ano16-42, Ano16-40, Ano16-41, Ano18a, Ano19e, Ano20g, Ano21n, Ano21o, Ano21p, Ano21q, Ano22s, Ano22t, BvRS^{+}11, BCPS22, CBS^{+}22, CP13, HB16, KWJ23, KQM13, LSB18, LR14, RRB^{+}14, RBD14, SP15]. Big-Data [KQM13]. BigBarChart [PCG^{+}21]. Bilevel [WN82]. Bill [Ano20-62]. Billion [JBD^{+}21]. BIM [LHZ^{+}21]. Binary [LJB^{+}13, MPHK94, UO91]. Binary-Encoded [LJB^{+}13]. binary-swap [MPHK94]. Bindings [SG86b]. Bing [Pen10]. Bins [SCH05]. Bio [LCG^{+}02]. Bio- [LCG^{+}02]. bioelectric [MJM93]. Biological [AJH^{+}12, CCE^{+}14, MGO14, WLC^{+}12, WLVL^{+}96]. Biological-Modeling [AJH^{+}12]. Biologically [JW16]. Biology [GSSO05, KS14, PF83, SPW84]. Biomechanics [Fig02]. Biomedical [GPW^{+}19, GM12, YBL^{+}12]. Biometric [Sim94a]. Biometrics [Kro02]. Biomolecular [GGB23]. Biopsy [NCQ^{+}11]. Biosciences [ZP95]. Biplane [WOF96]. Biquadratic [Ste84]. Bisectors [EK99]. Bit [GKB89, SB86a, Zac84, WWC91, Bar97a]. Bit-Map [Zac84]. Bitcoin [KFBI22]. BitConduite [KFBI22]. bitmap [Kor87]. Bivariate [DMR83, RO86, WG95a, Rou87]. BKViz [LTB16]. Black [Lee99, ABB^{+}93]. Black-Ink [Lee99]. Blackboard [ABB^{+}93, Ano92k, Ano92l, Ano92m, Ano93v, BB95, BHMN95, CR95, CG94, Ger91, HM95, HSAKD98, Koh93, KNK99, KGM95, KF94, MJM93, Mat96, MA98, MKL97, PJJF96, PM97, RBOV94, Sch93b, Sch91, SSRH95, Ste91, SBHM97, TC96, TS97, Tre98, UVa95, VCR98, ZPA95, Zha96a, TSH96]. blend [Sch93a]. Blender [BL14]. Blending [HT98, PPK05, SHS^{+}21, SR95, YLR17, Gol83c]. Blendshape [LD08]. Blendshapes [LiA10]. Blinn [Bli92e, Bli92f, Bli87, Bli88d, Bli88e, Bli88c, Bli88b, Bli89c, Bli89e, Bli89d, Bli89b, Bli89a, Bli90b, Bli90a, Bli90c, Bli91b, Bli91a, Bli91c, Bli91d, Bli92c, Bli92a, Bli92b, Bli92d, Bli93c, Bli93b, Bli93d, Bli93a, Bli94b, Bli94c, Bli94d, Bli94a, Bli95a, Bli95b, Bli95d, Bli95e, Bli95c, Bli96a, Bli96c, Bli96b, BG96, Bli96d, Bli97b, Bli97a, Bli98a, Bli98f, Bli98e, Bli98c, Bli98d, Bli98b, Bli99c, Bli99b, Bli99a, Bli00a, Bli00c, Bli00b, Bli00d, Bli00e, Bli01a, Bli01b, Bli02a, Bli02b, Bli03e, Bli03a, Bli03b, Bli03c, Bli03d, Bli04a, Bli04b, Bli04c, Bli05b, Bli05d, Bli05a, Bli05c, Bli06c, Bli06d, Bli06a, Bli06b, Bli07a, Bli07b, Whi95d]. Bliss [Sin11f]. Blitz [Sin07b]. Blocks [Mil07]. Blood [BGSC85, CM16, QCP^{+}10]. Blt [DFL87]. Blt-stones [DFL87]. Blue [Bre90]. Bluish [VB24]. Blur [HLCC08]. Board [Ano13c, Ano18-27, Ano22-64, MD99d, Mol18c, Ano20-50, Ano21-63, Ano21-64, Ano21-65, Ano22-60, Ano22-61, Ano22-62, Ano22-63, Ano22-65, Ano23-62, ECS^{+}15, Sto07, Tau10b, Tau11d]. Boards [MD99f]. Bodies [Bar95, MH04, Wil87b]. Body [Bad82, CLG^{+}16, CMSEM08, EBT98, KK02, hKKSK17, PMR^{+}09, SHMN09, Wil82, ZGC14, KmKK^{+}18]. Boeing [Bee82, DSW07]. BoidVR [CEMD22]. Bold [Hua11, Tur03]. Bone [CF93, MSB^{+}06, YSW^{+}23]. Books [Har00]. Boolean [Arb90, PS86, YT84]. Boost [Ma07]. Booster [MD99f]. Booze [PSE^{+}09]. Border [TYK01]. Borrowed [Bre90]. Bosnia [Hay96d]. Botanical [AK84]. Both [Hag90c, Hay96a, Sin12a]. Bottleneck [PEMF92]. Bottom [MR99]. Boundaries [Sin01d]. Boundary [BW94, BDH^{+}89, KSY85, Pie87b, Woo85, Ala92, CR91, Mil93, NB94]. Bounded [KT96, PPK05, Pav85]. bounds [Cam91]. Box [FM82, LWM14, TCE^{+}02]. boxing [YSSP91]. Brachytherapy [ZYJ^{+}21]. Brad [Sin21c]. Braid [Ks22]. Brain [Add95, DPL^{+}11, EHA^{+}14, Kan09, LGM13, LZF^{+}14, PT90, PP11, Sin12b, TA85, aCJB^{+}12]. Brain-Computer [LGM13]. Branching [RY06, Sin06a]. Brave [TH01]. Braving [Tur03]. BRDF [CPK06]. BRDF-Shop [CPK06]. BRDFs [RBK^{+}13]. Break [PMQ21]. Breaking [PEMF92]. Breathing [Jac15]. Breezy [Mol18a]. Bresenham [AM91, Wri90]. bridge [PABS86]. Bridges [Mol18b]. Bridging [PvW07, vW06]. Briefs [Bal97, FJ87, Mac97a, PJC88]. Bright [Bli00a]. Brightest [Fol14]. Brightness [Sto01, WWF00]. Bring [Sch07]. Bringing [BHY18, LaV08, Lea96a, Lea97d, NF11, PTR^{+}98]. Brings [Hay96a, She05]. Broadcast [FEV^{+}01]. Broadening [LIS^{+}19, Wat06]. Broader [LCI^{+}20]. Bronze [Mac97a]. Brooks [Whi23]. Browsable [Hay96c]. Browser [JB97]. Browsers [LGZ04]. Browsing [BCI08, Whi96a]. Brush [CT04b]. Brushing [FH19, Sin04b, FH21]. BSP [GC91]. BTF [FGR85]. Bubbles [Gla00c, Gla00d]. BubbleUp [Fis21]. BUCOLIC [Mei85]. Buddy [BNA^{+}22, BG96]. Buffer [BB84, DW87, FM85, FPI84, GF86, HG86, OKM84, PEMF92, BFP86]. Buffering [RR86b]. bug [WPO96]. Build [Del00b, Hil82, Sin22b]. BuilderPro [MD99f]. Building [AL11, Bak88, BRM^{+}02, Ear98, Enc15c, Fin08, FA85, HJS00, HGK^{+}04, HW18, JOnH93, Kul09, LCC^{+}00, LHZ^{+}21, LFFB20, Mol18b, Nad99, Nea88, NAB^{+}23, PB06, Rot03, SK13, She04, Sin04c, WZM^{+}11, Whi96b, YWR09, Sim95b]. Buildings [MMGG12, Pat12]. Built [AAB^{+}22, PHAS20]. Bullet [WF02]. Bulsarapp [MNC^{+}22]. Bump [MB94]. Bunny [Cam17]. Bus [GL82]. Business [AR99, BHPS12, Bas14, FDC14, Gus83, Hop83, LWM14, Mac97a, MD99e, Mac00a, Shi83, SOSG08, VG81, Wri97, Wri98, Mac91]. Butterfly [ABS^{+}05]. Buys [Hay95b]. Buzz [Sin12c].

C [Bli00d, Bli01b]. C2 [XJW18].

C2-Weighting [XJW18]. cache [GFB87]. Caching [DDdSC11]. CAD [Ano84, Bee82, Ber99, Boh95, BA85, Bre84, Cav87, DSW07, FJ87, GSN09, GWA84, Hal87, HR98, Kas00, KBF05, Ken81, KPTS89, Laz87, Lil81, MD99f, MD99e, MD99d, MD00f, NM08, Oia82, PKJ^{+}07, RSS^{+}09, RA06, Ric85, Rie22, RSR83a, RSR83b, SW83, SK02, Smi83, Tak85, Wei84]. CAD/CAM [Ano84, Bee82, Bre84, Ken81, Lil81, MD99f, MD99e, MD99d, Ric85, RSR83a, RSR83b, SW83, Smi83, Wei84, RA06]. CADBlocks [MD99d]. Cadcentre [MD99e]. CAGD [Hof93, Nie93a, Sch93c]. Cage [KW10]. calcium [SAO^{+}95]. Calculating [Bli96c, lLK84]. Calculation [FB81]. Calculations [JF99, WNAJ92]. calculus [Gor83]. Calendar [Ano95h, Ano95i, Ano95j, Ano95k]. Caleydo [TLS^{+}14]. Calibration [Hay95b, PLS^{+}16]. California [Del00c]. Call [Ano11b, Ano14j, Ano14k, Ano14l, Ano15n, Ano15o, Ano15p, Ano15v, Ano15t, Ano15q, Ano15u, Ano15r, Ano15s, Ano16c, Ano16r, Ano16s, Ano16t, Ano16u, Ano16n, Ano16o, Ano16p, Ano16q, Ano16v, Ano17a, Ano17p, Ano17j, Ano17k, Ano17l, Ano17m, Ano17n, Ano17o, Ano18c, Ano19n, Ano19m, Ano19o, Ano20j, Ano20k, Ano20l, Ano20i, Ano20q, Ano20m, Ano20r, Ano20n, Ano20o, Ano20p, Ano20s, Ano20-35, Ano20-34, Ano20-41, Ano20-42, Ano21y, Ano21v, Ano21w, Ano21x, Ano21z, Ano21-27, Ano21-28, Ano21-29, Ano21-55, Ano21-56, Ano21-57, Ano21-58, Ano21u, Ano22z, Ano22-27, Ano22x, Ano22y, Ano22-28, Ano22-29, Ano22-47, Ano22o, Ano22-48, Ano22-49, Ano22-50, Ano22w, Ano22q, Ano23v, Ano23u, Ano23w, Ano23s, Ano23t, Ano23q, Ano23r, Ano23-28, Ano23-33]. Call [Ano23-34, Ano23-29, Ano23-30, Ano23x, Ano23-31, Ano23y, Ano23z, Ano23-32, Ano23-27, Ano24g, Ano24f, Ano24h, Ano24i, DFL87]. Calls [Ano14n, Ano14o, Ano14m]. CAM [Ano84, Bee82, Bre84, Ken81, Lil81, MD99f, MD99e, MD99d, RA06, Ric85, RSR83a, RSR83b, SW83, Smi83, Wei84]. came [Bli88a]. Camera [AXP07, BE11, GDY^{+}11, HB08, MD99d, Mor05, PBR^{+}14, PXC02, Rus11, Sch00, SMP^{+}22]. Camera-Based [Mor05]. Cameras [GLC11, MTT86, RM89]. Camp [BL14, Sim94d]. Campus [BYP21, Sou04]. Can [Abo99b, Ano17b, Ano17c, Gai17, Gla99e, JB02, MGN^{+}11, Rei23, Sch17, SF20, SL00, Zuc84]. Canal [CZP^{+}10]. Cancer [BSM^{+}15, Lea96a, SOM^{+}10, TLS^{+}14]. Cancer-Screening [SOM^{+}10]. CancerSpace [SOM^{+}10]. Cantonese [CCT^{+}20]. Cap [Tho84]. Capabilities [GKB89, Kir21]. Capability [PJC24, Tak85]. Capture [Del98c, GLC11, GM12, LD08, RPM17, SH07]. Captured [ZSK^{+}21]. Capturing [BGM19, Mas18, SPMS04]. car [Sim95a]. Card [Gla02e, Gla02d, IM14]. Card-Weaving [IM14]. Cardiac [GZM97, SADC02]. Cardiovascular [Wix83]. Cards [DJM13]. Care [Han01]. Career [Ano19-43, Ano19-41, Ano19-42, Ano20w, Ano20-54, Ano21-32, Ano21-33, Ano23-41, Ano23-42, Ano23-43, Ano23-44, Ano23-45, Ano24n, Ano24o, LWZ^{+}18, Sin10a]. Careers [HBW^{+}24]. CareerVis [LWZ^{+}18]. Caricature [LZR^{+}21]. Carl [Mac97a, vD12]. Carlo [XP05]. Carnival [BHS11]. Carol [Tau12e]. Carries [Whi95d]. Carrying [SK13]. Cartograms [CCM15, KPN05]. Cartographic [BCA^{+}20, CMSEM08, Jon89]. Cartolabe [CRF^{+}21]. Cartoon [iAH03, FM22]. Cartoons [Zha22b]. carves [Ano97i]. CASDA [MF89]. Case [AY16, Ang17, BD85, Bli06b, Bli07a, ENDAD13, Gla99a, KM00, KND^{+}18, KCMK20, LCG83, MP99, PSJ87, PS09, PCH^{+}07, SD83, UBS^{+}12, ŽLK22a, ZLK^{+}22b]. Casting [MSRMH09, RBE08]. Casual [HMC^{+}20, PKM18, ST09]. Catalog [Bli00b]. Categorization [LV23, dRvWW10]. Category [SLSP22]. Category-Level [SLSP22]. Catheterization [LGW^{+}12]. Catheterization-Training [LGW^{+}12]. Catmull [AESF21]. Causal [BWG24, Sin23c]. Causes [Cai20]. CAVE [AEFQ96, KH14, HCL^{+}20]. Caves [aE01, LKPM19]. CC [LKS96]. CCAD [Koc94]. Celebrating [KWJ23]. Cell [Saw08]. CellPAINT [GAB^{+}18]. Cells [BKt82, Saw08, SSZP15, YBL^{+}12, ASK94, BKtH82, SAO^{+}95]. Cellular [CF16, GAB^{+}18, LMP89, Car87]. Celtic [Gla99c, Gla99b, Gla00a]. Census [Bow93]. Center [Ano16x, Ano18s, Ano23-41, Ano23-42, Ano23-43, Ano23-44, Ano23-45, Del00c, JAPF07, Ano24n, Ano24o]. Centered [AAAW22, BvDE^{+}99, GHS99, Jer18a, Tal97, YCKF15]. Centers [Ano97i, BvRS^{+}11]. Centralized [OBO^{+}23]. Centric [KZC^{+}11]. Centrifugal [OKET12]. Century [Ann17, Ano22-72, Ano22-73, Ano22-74, Ano23-74, Ano23-75, Ano23-76, Ano23-77]. Ceramic [DHC^{+}22]. Ceraunoscope [Gla00e, Gla00f]. Cerebral [BGSC85, MGW^{+}18, SDT^{+}93]. Certification [Ano13a]. CFD [MD99e]. CFX [MD99e]. CFX-5 [MD99e]. CG [AMDM22, AJH^{+}12, Bux05]. CG&A [Sto07, Ano96k, Ano98f, Ano99, Ano00, Ano03b, Ano04g, Ano06b, Ano07a, Ano08c, Ano09a, Ano14n, Ano14o, Ano14j, Ano14m, Ano14k, Ano14l, Ano14-33, Ano15n, Ano15o, Ano15p, Ano15t, Ano15q, Ano15u, Ano15r, Ano15s, Ano16r, Ano16s, Ano16t, Ano16u, Ano16n, Ano16o, Ano16p, Ano16q, Ano16v, Ano17p, Ano17j, Ano17k, Ano17o, Dil04, ECS^{+}15, FS23, Hay96f, Mac97a, Möl19, O’S11, Pot12, Rho96, Tau10a, Tur02b]. CGA [Ano17l, Ano17m, Ano17n]. CGI [Mor98]. Chagas [WBA^{+}96]. Chain [Dai89]. Challenge [CN03, LHM23, Mab82, SLC^{+}23, TPR^{+}22, vL16, LWM14]. Challenges [Ahr22, Bar15, BM24, BBC^{+}20, CKP96, Cro18, Enc06, GSG^{+}21, GB16, HF07, Her96, KBF05, Li22b, Ma09, MJS20, May83, MJM^{+}06, Pac10, SGE^{+}18, Sto15, TBFGC23, WCG^{+}11, WSJ^{+}12, WDC^{+}23, vL18, OJ14]. Chameleon [Nie07]. Chance [Cha98b]. Change [Ano96h, Ano97g, BAB^{+}23b, CHW21, Gla01a, JOG^{+}17, LSVB21, Per16, Sch08b, Sin07c, WLL^{+}09, MED^{+}93, ZJH^{+}18]. Change-of-address [Ano96h, Ano97g]. Changed [ODE^{+}13]. Changes [KP12, LWY^{+}20, Möl19, WW95]. Changing [BAE^{+}18, GGB23, Pet12, Rib98, Tho02b, War00]. Chaos [Sin07a, Pic90]. Character [BRD17, KDMH17, MYT09, SL11, WC11, WTR11]. Characteristic [TF97]. Characteristics [WP87]. Characterizations [BD89]. Characterizing [CLms15, TLS^{+}14]. Characters [AMDM22, ALD^{+}06, BMH98, HZ11, KMCMT09, KP11, NP17, TWB^{+}07, WB85]. Charles [Ano22a, Ano18d, SCN90]. Chart [Zhu12]. Charting [Tau10c]. Charts [Jar83, PKM18, PBV16]. Chasing [BRM10]. Cheap [Bal97, WW95]. Cheaper [Del97]. Chemicals [Sin08f]. Cheminformatics [LCG^{+}02]. Chemistry [SPW84]. Chernobyl [Abo98a]. Chesapeake [WLVL^{+}96]. Chief [Ano94j, Ano23-63, Her96, Tau10a, Tau13a, Tau13b, Tho01, Tho02a]. Child [Wei23]. Children [BNA^{+}22, RdCFP^{+}20]. China [ZNMT15]. Chinese [CT04b, kHHBEH13, Mak87, Shi17]. Chip [CB86, GGV92, PK87, DWL^{+}92]. Chips [Mac97a]. Choice [Ano18-35, Sad16]. Choosing [RB00, Rob91]. Christoffel [ODW15]. Chromaticity [LS84]. CI [WDK17]. CIG [Ler82]. Cimlogic [MD99f]. Cinema [Agr21a, MD99e]. CinemaScience [TBO^{+}20]. Cinematic [DSP91, FKKB05]. Cinematographic [MTT86]. Cinematography [DV21, RT07]. Circle [Fre86, HB97, Wri90]. Circle-Drawing [HB97]. Circles [PT89]. Circuit [SBHM97, Tho84]. Circuits [US17]. CirculaFloor [IYFN05]. Circular [Gla98b]. Circulation [GMNF97]. CiSE [Ano16d]. Cities [RWF02, Sin03c]. City [FZ03, LTJ18, Nur08, RGCB07, Sin22b, WPL22]. Clark [AESF21]. Class [Ahr15, WD15b, WD15c, ZNMT15]. Classes [GAW86]. Classification [BRD^{+}20, BBB08, CXW^{+}20, JRM17, MSW10, MSF07, MCEF94, NRH^{+}22, Til81]. classifications [Mil93]. classified [Bar81b]. Classroom [DJP^{+}08, WKM^{+}04]. ClaytricSurface [SPMK14]. Clearing [Ano94g]. CLEVis [MZC^{+}21]. Climate [CHW21, JOG^{+}17, LTS^{+}13, LSVB21, LSC^{+}17, MFG^{+}17, RBS21, WAH^{+}21, MCW93]. Climbing [RV07]. Clinical [YMDO93]. Clipping [RR85, Bli91b, Pin91]. Clips [LYO02]. CliPSS [YMDO93]. Clock [AC05]. Close [BKS20]. Closed [SB06]. Closer [Kos22]. Cloth [ADBW96, EWS96, HEW17, NG96, YYiTI92]. Clothes [VMJT96]. Clothing [CSMT03, PCG^{+}21, TWB^{+}07]. Cloud [Ano15v, CWL^{+}10, MBSD20, PSX13, SMDBR22, SGG^{+}22, Ano14p]. Cloud-Based [MBSD20]. Clouds [Ano94g, GSFT16, Hag90c, LCW12, RGG19, SBC15, Zha22b, RG95]. Cluster [Kel84, LLR09]. Clustering [FAECOG19, FIB^{+}20, IK15, MQM22, RKJH99, WMFM13, XLJ^{+}21, XJ94, ŽLK22a]. Clustering-Driven [FAECOG19]. Clusters [HGK10, WPLM01, ZYL^{+}17]. Clutter [LSFW83]. Cluttered [AA09]. CNN [FH21]. CNS [FS86]. CNS-HLS [FS86]. CO [AGBC09, BRD17]. Co-presence [BRD17]. Coarse [CRC^{+}02]. Coast [KP12]. Cockpit [MSA^{+}94, PG93]. CoCoNutTrix [IBH^{+}09]. Code [PLM^{+}19, Sin13c, Sin16a]. Codes [NAB^{+}23]. Coding [GUFG15, LSQ^{+}12, SKK91]. Cognition [End14, ECNZ15, Jae17]. Cognitive [CMT^{+}10, PD09]. Coherence [Cro87a, GP89, SM84]. Coherent [PBR^{+}14]. Cohort [AOH^{+}14, BSM^{+}15]. Cohort-Study [AOH^{+}14]. Collaborating [LA20a]. Collaboration [AJD13, BH11, CMZ01, Dom12, FT00b, HBG04, KKVL05, LD97, MD99d, OCS20, Sin18d, VCS19, Wat06, vOW15]. Collaboration-Specific [BH11]. Collaborations [CS15b, KM13]. Collaborative [Agr09, BRG^{+}00, CRB^{+}02, CPN09, CCVA02, Enc15c, Fis02, FFA^{+}97, FLSG98, IBH^{+}09, JLJ^{+}10, JMOG99, KP18, LFV^{+}97, LJ96, MSDS23, NPH^{+}09, PW97, RDH^{+}02, SSTP18, SKS^{+}96b, SCH05, WHR19, WLL^{+}09, ZSK^{+}21, vHV09]. Collage [Fig01a]. Colleague [Ano20-61]. Collection [Cro96]. Collections [CRF^{+}21, CRD^{+}18, WSL^{+}20, YLL10, dRvWW10]. Collins [Abo98b]. Collision [Dai89, EW96, FUF06, GASF94, SGG^{+}22]. Collision-Free [Dai89, EW96]. Collisions [SGC21]. Colloidal [OOHR14]. Colonization [LCS21]. Color [Bar97a, BBK82, Bli93c, BP06, BT07, BH11, BBS09, Dia91, EH99, FZY84, FHH^{+}21, GUFG15, HML83, Hal99, Ike84, JF99, LS84, LH92, LBW^{+}17, LBSS85, Mac99a, MD99f, MD99e, Mei85, MSK04, MG88b, Mur84a, NM08, PH98, PH99, PP95, RGW05, RAGS01, RO86, RS99, RP95, SW83, SPA^{+}16a, Sto01, Sto12, TT00, War88, XJ94, YP08, ZLMY19, Dia94, FG88b, Rob88, WWC91, MD99f]. Color-Accurate [SPA^{+}16a]. Color-defective [MG88b]. Color-Map [BH11]. Color-Naming [BBK82]. Color-Shaded [SW83]. Color-Space [NM08]. Colorful [OZH15]. Coloring [PP12]. Colorization [BP06, SSC^{+}11]. Colormap [SKR18]. Colormaps [WSS23]. ColorMoves [SKR18]. Colors [Hag90a, Sin14d, Sto05, SFCD99]. ColorSketch [LBW^{+}17]. Column [Her96]. Com [Bet00]. combat [Bro89]. Combination [Ano16d]. Combinatorial [Woo85]. Combined [VCR98]. Combining [BHS11, HBB20, KSRS13, LGM13, Lec17, OZH00, PNPS03, WZM^{+}11]. Combustion [WYG^{+}12, YWG^{+}10]. Comdex [Hay96d]. Comet [Ano94d]. Comfort [AMDM22, ZR24]. Comfortable [FFN19]. Comic [JSH16]. Comics [BRCP17, SRL06]. Coming [Agr09, CS17a]. Comings [Tau11b]. Command [Sin18b]. Comments [Use93]. Committee [LRS09]. Commodity [Sto21]. Communicated [SDPW20]. Communicating [RPOC21]. Communication [AD05, CNT^{+}97, EAM22, Jae17, Lil81, Lok06, MGN^{+}11, Neu94, RB00, Sin18c, TN10, YLT18]. Communications [Agu86, OK09, Hay96a]. Communities [BvDE^{+}99, PCG^{+}21]. Community [CW09, Enc15c, GFW^{+}16, MZC^{+}21]. Compact [MD99f, MD99d, de 86]. Companies [CBS^{+}22]. Comparative [AHP^{+}10]. Compared [FH21]. Comparing [DPL23, ELK97, Hal99]. Comparison [Cro81, Eas90, HS87, LTD16, MMD^{+}19, TPR^{+}22, vL18, GG87]. Compatible [FM89a]. Compelling [HCL^{+}20, SH03]. Competing [WXK22]. Competition [JB97]. Completely [Sin23a]. Completion [MZL^{+}21, ONDA20]. Complex [BQP^{+}15, DAG95, JW16, KTA^{+}02, KMS07, MTT91, PP12, Rho97a, SSW^{+}07, Sin17a, Ste07, KS94, RN94]. Complexity [KCMK20, NAB^{+}23, SCS23, Voo89, Til81]. Component [CCB22, KRG^{+}22]. Components [HW82a, HH01, TW88]. Composers [FPJ04]. composing [Mak87]. Composite [SSH^{+}98, KGM95]. Compositing [Bli94d, NINT89, QNT02, RGL^{+}06, MPHK94]. Composition [FK87]. Composting [Bli94a]. Comprehensive [VG84]. Compressed [Van10, SR93]. Compressing [VK17]. Compression [FM89b, KB92, Lea96d, WYM10, WGZ^{+}18, Har96]. Compressive [WLH^{+}12]. Compromise [Whi95d]. COMPSAC [Ano18e, Ano19p, Ano20t, Ano21-30]. Computation [DV21, End14, ECNZ15, GWH05, Hal99, MHC98, Sar82, XNSW19, Til81]. Computational [BKLZ17, BA13, CBGH19, CP13, CN03, DM02b, DS07b, EH16, Fig01b, Gai04, GW08, HU83, Lev10, MD02e, Mas18, RWB^{+}14, SHY^{+}23, UFK^{+}89, UZM^{+}23]. computations [Woo93]. Computed [NFMD90]. Computer [AKME^{+}90, Agr18a, Ama87, AVV20, AS11, Ang17, Ano03a, Ano04a, Ano14s, Ano14q, Ano14p, Ano14r, Ano15w, Ano15y, Ano15x, Ano16e, Ano16f, Ano16y, Ano16x, Ano16z, Ano16w, Ano16-27, Ano17q, Ano18s, Ano18u, Ano18t, Ano18x, Ano18v, Ano18w, Ano18r, Ano19g, Ano19-30, Ano19-31, Ano19-32, Ano19-35, Ano19-36, Ano19-34, Ano19-33, Ano19-44, Ano19-54, Ano20-32, Ano20-30, Ano20-31, Ano20-47, Ano20-48, Ano20-43, Ano20-49, Ano20-51, Ano20-44, Ano20-38, Ano20-45, Ano20-41, Ano20-39, Ano20-50, Ano20-42, Ano20-46, Ano20-40, Ano21-39, Ano21-40, Ano21-41, Ano21-42, Ano21k, Ano21-55, Ano21-43, Ano21-66, Ano21-59, Ano21-44, Ano21-45, Ano21-56, Ano21-46, Ano21-47, Ano21-48, Ano21-49, Ano21-50, Ano21-67, Ano21-51, Ano21-52, Ano21-68, Ano21-63, Ano21-60, Ano21-64, Ano21-57, Ano21-53, Ano21-61, Ano21-58]. Computer [Ano21-54, Ano21-62, Ano21-65, Ano22-37, Ano22-39, Ano22-40, Ano22-41, Ano22-42, Ano22n, Ano22-47, Ano22-55, Ano22-51, Ano22-43, Ano22-60, Ano22o, Ano22-44, Ano22-61, Ano22-52, Ano22-48, Ano22-45, Ano22-62, Ano22-53, Ano22-49, Ano22-46, Ano22-63, Ano22-56, Ano22-54, Ano22-58, Ano22-64, Ano22-57, Ano22-50, Ano22-59, Ano22-65, Ano23-28, Ano23-29, Ano23-30, Ano23x, Ano23-31, Ano23y, Ano23z, Ano23-32, Ano23-27, Ano23-50, Ano23-63, Ano23-64, Ano23-56, Ano23-62, Ano23-51, Ano23-57, Ano23-41, Ano23-65, Ano23-52, Ano23-58, Ano23-42, Ano23-46, Ano23-59, Ano23-47, Ano23-53, Ano23-43, Ano23-48, Ano23-54, Ano23-60, Ano23-44, Ano23-49, Ano23-55, Ano23-61, Ano23-45, Ano24h, Ano24i, Ano24m, Ano24p, Ano24t, Ano24r, Ano24n, Ano24q, Ano24s, Ano24u, Ano24o, Arn14a, Arn14b, Atw84, BC87, BC05, Bar81b, Ber83]. Computer [BHS11, BT00, BB82, BBDM98, Bli98c, BDM02, BBS09, Bre96a, BH83, CDLS19, CF93, CHP21, Car83, CM86, CR02a, CLC^{+}96, CE83, DGTK07, DP82, De 11, Del02, DiS14, Die85, DDG88, Ear98, EMTTT98, Enc00, Enc06, ECS^{+}15, Enc23a, Enc23b, Fer01, Fet82, FWC84, FAB^{+}98, Fry83, FL81, Fyo97, Gar15, GB96a, Gir87, Gla96b, Gla03d, Gol83b, Gol00, Got82, GRWN87, Gre84, GM87, HW82a, HH92a, Hay96b, HJO86, Hob81, HM85, HLC11, HN85, Hur82, JCM07, KSM85, Kas11c, Ker13, Kla89, Koc94, Kon12, Kri05, Kri06, Lav83, LLML16, LGM13, LR90, LMP89, Lic84, LPMEPCM11, LSFW83, MC85, Mac99a, ML16, Mac81, Mac83, Mac85, Mac94, Mac00a, MIA^{+}08, MTT85b, MTT85c, Mar07, Mar83, Max83, Mil99b, MP23, MM14, Mye85, Myk88, NTS^{+}07, Nea86]. Computer [Nea89a, NG96, Nie92, Pal81, Pal85, PPVH23, PF83, PG04, Pot96, PFA86, Pri00, RME14, RGA^{+}83, Rho90, Rho91, Rho93, Rho97a, Rie81, RS99, SDS^{+}11, Sar90, Saw83, Sch81, SM86, SMWS88, SG11, SF20, SG86a, SRL06, She82, Sin02f, Ska86, SG86b, Sta91a, Ste18, SP06, SRS83, SDS95a, Sto21, Str90, SBE95, Stu98, SD83, Tau10c, Tho00, Tur02a, USW17, VG84, VG81, WB85, War81b, War81a, Wei23, WMB85, Whi97, XDJ23, YST84, YKR^{+}84, Yan85, Zha22b, Bro89, FG88b, GB96b, Hod92, Mac91, MG88b, PG93, RKRW87, Rho96, SW88, Sta91b, SDS95b, WFG92, Woo91a, Wor93, Ano96i, Ano97h, Ano18y, Ano18-28, Ano19-38, Ano20f, Ano22-38, KR86, Ano20-70]. Computer-Aided [BC87, BDM02, BH83, CF93, CLC^{+}96, DP82, HLC11, Hur82, KSM85, Lic84, Myk88, NTS^{+}07, Nie92, SM86, SG86a, Bar81b]. Computer-Animated [Gir87]. Computer-Assisted [AKME^{+}90, SRS83]. Computer-Generated [Gla96b, HM85, JCM07, MIA^{+}08, PF83, Yan85, WFG92]. Computer-Integrated [YKR^{+}84]. Computer-Supported [PPVH23]. Computer/Art [Pal85]. Computerized [Hue83]. Computers [CWRF05, HHLL14, HS01, MGN^{+}11, Tho84, de 86, Ano19o, Ano20q, Ano20m, Ano20n, Ano20o, Ano20p, Ano21z, Ano21-27, Ano21-28, Ano21-29, Ano22-28]. Computing [Ano87, Ano20-35, Ano20-33, Ano20-34, Ano22-30, Ano24j, Bak88, BAB^{+}23b, BvDE^{+}99, EK99, Gla01c, Gla01d, Gla01e, GP90, HJ03, Lee08, Li22b, LS04a, LCC^{+}96, Nie07, PTB^{+}15, RMG16, SZP^{+}00, Sto15, VBW94, YIÇ09, vL16, Ano15v, Ano16-29, Ano18b, Ano19f, Ano19h, Ano19k, Ano19l, Ano20j, Ano20k, Ano20l, Ano20c, Ano20h, Ano21y, Ano21a, Ano21b, Ano21c, Ano21i, Ano21j, Ano21t, Ano21r, Ano21s, Ano22z, Ano22-27, Ano22c, Ano22d, Ano22e, Ano22k, Ano22l, Ano22m, Ano22q, Ano22u, Ano22v, Ano23u, Ano23b, Ano23c, Ano23d, Ano23j, Ano23k, Ano23l, Ano23m, Ano23p, Ano24a, Ano24d, Ano24e]. ComputingEdge [Ano19q, Ano19r, Ano20d, Ano20e, Ano21d, Ano21e, Ano21f, Ano21g, Ano21h, Ano22f, Ano22g, Ano22h, Ano22i, Ano22j, Ano23e, Ano23f, Ano23g, Ano23h, Ano23i, Ano24b, Ano24c]. concept [WJ93]. Concepts [BFW99, Rou09, SCB04a]. Conceptual [JMOG99, LGK^{+}06, LO94]. Concurrent [HDKS95]. Conditioned [Mac90]. Conducting [LCB^{+}23]. Conference [Ano20r, Ano23-33, Ano23a, Kos15, Bli95d]. Conferences [Ano92d, Ano93h, Ano93i, Ano93j, Ano94g, Ano94e, Ano94f, Ano23-46, KWJ23]. Conferencing [EG98]. Confidence [WDK17]. configurable [GFB87]. Configuration [SPJ99]. Configurations [CPN09]. Configuring [PM82]. Conflicts [ZR24]. Conformal [FS90, OC14]. Confusion [Cai20]. Congress [Ano19-39, Ano19-40]. Congruent [SDM22]. Conic [HP91, Wil87c]. Conics [BS96, Far92]. Connect [Ano17-28, Ano17-29, Ano18f, Ano19s, Ano19t, Ano17-30]. Connected [Ano15-39, Ano18-42, MR87, Ano19-55]. Connecting [LPMAA^{+}22]. Connection [MD00f]. Connectionless [BS03]. Connections [LQ18, LFFB20, Whi96c]. Connectivity [COK97]. Connectome [BHAA^{+}13]. Conquering [HWK^{+}96]. Conservative [LW13]. Consider [Bli96a]. Consideration [Bli98e, Woo88]. Considerations [Cav87, MJS20, RR95, Smi84a, WXK22, BB95]. Considered [BT07]. Considers [Hay96d]. Consolidation [LCW12]. Constituents [Woo88]. Constraint [CSI86, FB93]. Constraint-Based [CSI86, FB93]. Constraints [Ahr15, BMW87, FAECOG19, HKSK17, Alp93]. Constructing [Fre86, FZ03, MW82, SK91, Wal91]. Construction [AAB^{+}22, IM14, MGP^{+}04, SKR18, She04]. Constructions [BD90]. Constructive [Men94, RR86b, SWS91, WKS86, Cam91]. Contact [CS06, Li22a, MH04, OBO^{+}23]. Contamination [GDF^{+}14]. Contemporary [RV82, Sin01a]. Content [BKSS07, DL15, HdlRL^{+}13, IM11, ITW01, JPSM13, LJB^{+}13, MSED23, TSLX12, WGZ^{+}18]. Content-Aware [TSLX12]. Content-Based [BKSS07]. Contents [Ano11c, Ano14-34, Ano14-36, Ano16-44, Ano16-45, Ano16-46, Ano16-47, Ano16-48, Ano17-33, Ano17-34, Ano17-35, Ano17-36, Ano18-44, Ano18-46, Ano18-47, Ano18-48, Ano18-49, Ano19u, Ano19-56, Ano19-57, Ano19-58, Ano19-59, Ano19-60, Ano20-64, Ano20-65, Ano20-66, Ano20-67, Ano20-68, Ano20-69, Ano21-74, Ano21-75, Ano21-76, Ano21-77, Ano21-78, Ano21-79, Ano22-75, Ano22-76, Ano22-77, Ano22-78, Ano22-79, Ano22-80, Ano23-78, Ano23-79, Ano23-80, Ano23-81, Ano23-82, Ano23-83, Ano24x, Ano24y, Ano13g, Ano14-35, Ano14-37, Ano14-38, Ano15-40, Ano15-41, Ano15-42, Ano17-32, Ano17-37, Ano18-45]. Contest [BHS^{+}24, PGS^{+}08, SMG^{+}21, Tau11b, TPR^{+}22, DPL^{+}11, FKK^{+}10, GGS^{+}18, LZF^{+}14, OKET12, SKS^{+}13]. Context [CWL^{+}10, GLRK06, GD11, KMH02, OBE^{+}84, dHPP10, JSDL21]. Context-Aware [GD11, dHPP10]. Context-Preserving [CWL^{+}10]. Contextual [BWG18]. Contiguous [CCM15]. Continent [Hag90a]. Contingency [WHC^{+}14]. Continuator [Pac04]. Continuing [Arn14b]. Continuity [BD89, BD90, Gar15, GU86]. Continuous [BD90, BW94, WN82]. Continuous-Tone [WN82]. Continuously [Mil07]. Continuum [DRHL^{+}03]. Contour [Ks22, SR85, Her86]. Contouring [DBV89]. Contours [Her24]. Contraction [Bli01b].

Control

[AMJ^{+}08, AGL87, Ber83, CEMTT98,

COK97, DLEF15, EDC96, EBN13, HPG08, JXW^{+}08, KP11, hKKSK17, LJ15, LS08, LLZ13, Mab82, MD00f, PMR^{+}09, Pie87b, PT00, PBM^{+}96, RDO^{+}96, SPMK14, SKBF12, SKS^{+}96b, Sin02d, Sin17a, VS07, Wil87a, WS90b, Zel82, BBP94, Coq87]. control-point-based [Coq87]. Controlled [IMS03, JDH^{+}89]. Controller [AP11, BDJ86, DSKH95, SHY^{+}23]. Controllers [TLT11, Jam87]. Controlling [YIK85]. Convection [ME22]. Convenient [WWV^{+}10]. converge [Sim93c]. Convergence [ARS17]. Convergent [Kel84, San01]. Converging [BHPS12]. Converging-Business-Ecosystem [BHPS12]. Conversation [WSP^{+}20]. Conversations [HZHD03]. Conversion [GPR94, GG83, MKFN87, Pav85, Vos85]. Converting [Gha12]. convexity [FM87, Rou87]. convexity-preserving [FM87, Rou87]. Conveying [ABS16]. Convolutional [CXW^{+}20, Hei18, TPA24, WZQ^{+}19]. Cooking [NOK^{+}08]. Coons [WW95]. Cooperation [Koc94]. Coordinate [Bli95b]. Coordinates [Rie81, Bli91c]. Cope [Del02, SP15]. coprocessor [Shi86]. Copyright [Whi95d]. Copyrights [VP99]. Cords [CS06]. Core [Ano15-44, Enc14d, FS01, KH98, GS12]. Cornea [SRJ99]. Cornell [Ano97i]. Corner [Bli92e, Bli92f, Bli87, Bli88d, Bli88e, Bli89c, Bli89e, Bli89d, Bli89b, Bli89a, Bli90b, Bli90a, Bli90c, Bli91b, Bli91a, Bli91c, Bli91d, Bli92c, Bli92a, Bli92b, Bli92d, Bli93c, Bli93b, Bli93d, Bli93a, Bli94b, Bli94c, Bli94d, Bli94a, Bli95a, Bli95b, Bli95d, Bli95e, Bli95c, Bli96a, Bli96c, Bli96b, Bli96d, Bli97b, Bli97a, Bli98a, Bli98f, Bli98e, Bli98c, Bli98d, Bli98b, Bli99c, Bli99b, Bli99a, Bli00a, Bli00c, Bli00b, Bli00d, Bli00e, Bli01a, Bli01b, Bli02a, Bli02b, Bli03e, Bli03a, Bli03b, Bli03c, Bli03d, Bli04a, Bli04b, Bli04c, Bli05b, Bli05d, Bli05a, Bli05c, Bli06c, Bli06d, Bli06a, Bli06b, Bli07a, Bli07b, HT98, vL16, Bli88c, Bli88b]. Corners [BG96]. Corporate [Mar84, Wri83]. Corpus [EZM^{+}97]. Correct [CPK06]. Correction [BN95, HBR^{+}92b, PP95, Bli00c]. Correlation [FAL^{+}06, WDK17]. Cortex [SMWS88]. Cortical [SM86, WVL^{+}18, ZTT99]. Cosegmentation [WZQ^{+}19]. Cosmic [SMG^{+}21, Sin06a, GR94, Sim93b]. Cosmological [HH08, SBD^{+}17, SW22]. Cost [PCJ^{+}16, SGC21, YST84]. Costs [Neu94]. Could [Whi95d]. Counterfactuals [BWG24]. Countering [CAG^{+}06]. Countries [NdSM^{+}16, RMG16]. counts [DH91]. Coupled [PIBV11, WLVL^{+}96]. Coupling [End14, ECNZ15, SCS22, YK95]. Course [Dom12, EE12, Gre84, Moh85]. Courses [AJK15, BYP21, KSK21, Ker13, ODE^{+}13, SDS^{+}11, YY16, Zhu12]. Court [LCB^{+}23, Hay96a, Whi95d]. Covariant [Bli06c]. Coven [TSF^{+}98]. Cover [Abo99a, Abo00a, Abo00b, Ano92a, Ano92b, Ano93a, Ano93b, Ano93c, Ano93d, Ano93e, Ano94a, Ano94b, Ano14e, Ano14f, Ano15g, Ano15h, Ano15j, Ano16h, Ano16i, Ano16j, Ano16k, Ano16l, Ano16m, Ano17e, Ano17f, Ano17g, Ano17h, Ano18h, Ano18i, Ano18j, Ano18k, Ano18l, Ano18m, Ano19x, Ano19y, Ano19z, Ano19-27, Ano19-28, Ano19-29, Ano20x, Ano20y, Ano20z, Ano20-27, Ano20-28, Ano20-29, Ano21-31, Ano21-34, Ano21-35, Ano21-36, Ano21-37, Ano22-31, Ano22-32, Ano22-33, Ano22-34, Ano22-35, Ano22-36, Ano23-35, Ano23-36, Ano23-37, Ano23-38, Ano23-39, Ano23-40, Ano24k, Ano24l, Bal97, Bal98a, Cha97a, Cha97b, Cha98b, Cha98a, Cha99, Del98a, Del00a, Fig00, Fig01a, FY97, Hay95a, Hay97, How00, Law99, Lea96b, Lea96c, Mus99, Sim96a, Sin01b, Sin01c, Sin02b, Sin02a, Sin02d, Sin02c, Sin02e, Sin03a]. Cover [Sin03b, Sin03d, Sin03c, Sin03e, Sin03f, Sin04a, Sin08a, Sin09a, Sin09b, Sin09c, Sin09d, Sin13a, Sin16d, Sin19d, Sin21f, Sin23b, Sin24b, TH01, Wea98a, Wea98b, Wea99b, Wea99a, Whi95b, Whi95c, Whi96b, Whi96c, Wor95, Yea97a, Ano14g, Ano14h, Ano14i, Ano15f, Ano15i, Ano15k, Ano17d, Ano17i, Ano21-38, Sin16a]. Coverage [Mil12, WEWL00, Bli96c]. Covered [Ano19-34, Ano20-43, Ano20-44, Ano20-45, Ano20-46, Ano21-59, Ano21-60, Ano21-61, Ano21-62, Ano22-55, Ano22-56, Ano22-57, Ano23-50, Ano23-51, Ano23-52, Ano23-53, Ano23-54, Ano23-55, Ano24r, Ano24s]. COVID [KCB^{+}23, OBO^{+}23]. COVID-19 [KCB^{+}23, OBO^{+}23]. CPU [GKB89]. Craft [Iga19]. Crafts [VVP^{+}11]. Craftsmanship [MCS22]. CraMs [DNS^{+}15]. Craniofacial [PVML96]. Craniometric [DNS^{+}15]. Crash [KHSE98, SWS07, Koh93, Sim95a]. Crash-Worthiness [KHSE98]. Cray [Ano17-31]. Crayons [Sin08d]. Creasing [KGK94]. Creates [Sin13e]. Creating [Abo00a, CMT^{+}10, CTM18, CPK06, CW09, FEV^{+}01, GH86, KGR^{+}18, LSV^{+}02, Mye87, SKJ^{+}13, YL09, ZED^{+}09, Sim96c]. Creation [DL15, Fei85, GS87b, KS07, PB10, Sin06d, Sin08f, Sin13a, ASW93]. Creative [Bra86, CLP^{+}14, Shu23, Sin15a]. Criteria [JKM23]. Critical [CCB22, GCS^{+}12, SRK90, TCH^{+}96]. Criticism [KDHL08]. Crop [Gla04c, Gla04d, Gla05a, MV93]. Cross [ANS15, Bli98d, Bli03e, CJS^{+}19, SK91]. Cross-Platform [CJS^{+}19]. Cross-Ratio [ANS15]. Crosscutting [SSZS21]. Crossings [Gla99f]. Crossover [Sin10d]. Crosswalk [GJM18]. Crow [Ber86]. Crowd [BKM16, CTJ^{+}17, GD11, GD13, HUS^{+}21, Hay96d, JXW^{+}08, PD09, PB06, TLC02, YIÇ09]. Crowd-Behavior [BKM16]. Crowd-Driven [HUS^{+}21]. Crowds [DPA^{+}11, Kan09, MYT09, MYPT09]. CRT [Zuc84]. Cruise [PSE^{+}09]. Cryptography [Wor96b]. Crystal [FH84b]. CS [Ano19a, Ano22a, Ano14-39, Ano15-43, Ano18g, Ano18-50, Bal97]. CSG [CH95, GMTF89, LF87, Vos85, WKS86]. CSpray [PW97]. Csuri [SCN90]. CT [FYZ85, RGA^{+}83, Wal91, ZCBM14]. CT-Aided [RGA^{+}83]. Cube [Lea96d]. Cuberille [CHRU85]. Cubes [Ano20-62, Lor20, SMMS24]. Cubic [Bli89a, Bli99b, Bli99a, Bli00b, Bli06c, Bli06a, Bli06b, Bli07a, Bli07b, FPW^{+}00, Sed90b, Bli89d, CP92, Sed90a]. Cubic-Mouse-Based [FPW^{+}00]. Cubism [CMPA18, Gla04e, Gla04f]. Cues [MCGS21]. Culling [AM04, CIF99, Bli93a]. Cultivating [SH06]. CultLab3D [Sin14a]. Cultural [Arn14a, Arn14b, BT00, BCA^{+}20, Mar07, RSZ20, Sin14a, SPM10, WSL^{+}20, ZED^{+}09]. Cultural-Heritage [SPM10]. Culture [Rus11, Cra91]. Cuneiform [AL02]. Curing [BRL^{+}17]. Current [RV83, SGE^{+}18]. Curriculum [ZD98]. Curse [Her23]. Curvature [BW94, FS89, KS01, TF97]. Curvature-Continuous [BW94]. Curvature-Dependent [KS01]. Curve [BD85, CS06, ELK97, Jos11, Til83a, Bli89a, FB93, GD93]. Curve-Based [Jos11]. Curved [ADBW96, Wee02, Wei85, GG92]. Curved-Surface [Wei85]. Curves [BD89, BD90, Bar93, Bli99b, Bli99a, Bli00b, Cha88, FS89, GP90, LMZ11, PQC^{+}12, Pie87a, PS88, RV10, SF92, Tau94, TF97, WN82, WWL^{+}21, Bli89d, Hof93]. Custom [GRWN87, PHAS20]. Custom-Built [PHAS20]. Customized [CLG^{+}16, HLC11]. Customizing [CP13]. Cut [SJB^{+}16]. Cutouts [SC12]. Cuts [MCGS21]. Cutting [BTM17, JK09]. Cuttlefish [BU23]. CVPR [Ano20u]. Cyberhuman [Fig02]. Cyberinfrastructure [JLJ^{+}10]. Cybernetic [Pac04]. Cybersecurity [Ano14-31, Ano14-32, Ano15-35, Ano15-34, Ma06]. Cybersickness [WW95]. Cyberspace [CWZ99, Pot98, Whi96b, Kri03]. Cyberworlds [Kun00]. Cycle [KM00, MD99e, PMQ21]. Cycles [MD99e]. Cyclides [DMP93]. Cycling [BM88, Woo15].

D [Abo99a, Ano18n, FYZ85, Hay95b,

MD99f, MD99e, MD00f, MKFN87, PKJ^{+}07, SD99, AP11, ASW93, AP10, Ano15-33, AA09, BRL^{+}17, BC87, Bal98c, BBF^{+}98, BGSC85, Bar84, Ben99, BH16, BA13, Bli88c, Bli03e, BFB^{+}98, BCF^{+}08, BBT^{+}13, BU23, BKSS07, Bux05, CD12, CCG^{+}04, CPCS08, Cam17, CCV85, CCF97, CRB^{+}02, CBB^{+}14, CT09, CE83, COK97, CMO^{+}99, CMPA18, DB03, Del99b, DL22, DPL23, DGA^{+}97, DNS^{+}15, Die00, DY09, DWL^{+}92, EHBPG04, Enc14a, FZY84, FSK07, FD03, FB09, FHKH06, FZ03, FPT^{+}82, FPI84, FTAT00, FM22, GWF98, Ger91, Gir87, GBT^{+}02, GKB89, HPG08, HdlRL^{+}13, HWY16, HTZ^{+}19, Han01, HGMT17, HF07, Hay95b, HU83, HJF06, HLE04, IKMF12, Ike84, IG98, JWL^{+}18, JAC^{+}08, KN07, KK02, KS07, KTA^{+}02, KB88, KP14, KSRS13, Kon12, KZC^{+}11, Kul09, KPL17, KDMH17, LaV08]. D [LFV^{+}97, Law99, Lea96a, Lea96d, LJB^{+}13, LYO02, LSQ^{+}12, LSFW83, LCG83, Mac97a, Mac13, MD99f, MD99e, MD99d, MD00f, MTT83, MKFN87, MSF07, MGO^{+}16, Mas98, MVWB05, McG93, MFM^{+}16, MSH^{+}08, Mil03, Mun98, NTH02, NINT89, NE93, NM08, NF91, NLvLM09, dSNRdLdSN16, Nur08, OATO85, OWZ17, PW97, PSX13, PLS^{+}16, PJC88, Paq96, PT90, PTR^{+}98, Pie06, PVV^{+}03, PG93, Red01, RLOW13, RWB^{+}14, RR03, RB00, RDH^{+}02, RT07, RFN89, RBK^{+}13, RGCB07, SRS91, SPJ99, SS17, SSG95, SBMK20, SB16, SMWS88, SRL06, SLC24, SRDT01, Shn03, SM04, Sim96c, Sin01c, SSOC01, SFL^{+}04, Spa16b, SQG^{+}05, SLSP22, SAHK13, aKTRL^{+}21, Tau11a, TLL^{+}96, TMB^{+}13, THB^{+}90, TUH^{+}90, TPA24, TH17, TT84, US17, UIL^{+}06, UBS^{+}12, WOF96, WAH^{+}21, WMS92, WK94, WLL^{+}16, Wea99b, WWV^{+}10, WL88, WSKY09, Wor96b, WZQ^{+}19]. D [WWL^{+}21, YY16, YCB05, YZM^{+}11, YYOiT92, YY99, YWR09, YBL^{+}12, ZZ05, ZCS^{+}06, Zha22b, ZSK^{+}21, ZLK^{+}22b, aE01, vHdP94]. D&I [Ano22-51, Ano22-52, Ano22-53, Ano22-54, Ano24q, Ano23-47, Ano23-48, Ano23-49, Ano24p]. D-Printing [RBK^{+}13]. D-Rendering [THB^{+}90]. D-Shape [KS07]. D/ [MD00f, MKFN87]. Damage [RDO^{+}96, VS07]. Dance [CWRF05, Ebe05]. Dark [Hag90a, SBD^{+}17]. Dark-Matter [SBD^{+}17]. Dart [ADBW96]. Dashboard [TBFGC23, UZM^{+}23, LOC^{+}16]. Dashboards [FT16]. Data [APB^{+}21, ABM^{+}01, Ahr15, AAW17, AGBC09, AOH^{+}14, Ano14-30, Ano16-42, Ano16-40, Ano16-41, Ano19e, Ano19-53, Ano20r, BMC84, BRCP17, BC19, BS16, BQP^{+}19, BSS20, BHAA^{+}13, BCPS22, BGM19, BHP07, BD21, CCG^{+}04, CCV85, CCE^{+}14, CEH^{+}09, CLms15, CP13, CP92, Chr83, CE98, CPG^{+}16, CCVA02, DB03, De 89, DKS^{+}05, DD19, DSW07, DP00, DCLK03, DFL^{+}18, FPNVHH17, Fis21, FM89b, FTAT00, Gai04, GV81, GDF^{+}14, GB16, GM12, Gro83, GHCW21, GDS^{+}18, HTZ^{+}19, HS87, Hed84, HKW99, HB16, HJF06, HNC^{+}04, HHH^{+}20, HHH^{+}21, HHKK17, HLE04, HW18, HMC^{+}20, IIH^{+}13, ILC10, ITSK06, IDJW19, JSH16, VCR00, Jar83, JBH^{+}10, JSDL21, KBHP14, KWK21, KCB^{+}23, KJA^{+}18, KPSN04, KQM13, Ken81, Kir21, Kla89, KSRS13, KMM^{+}01, Kos15, Kos23, Kd21, KNK99, Kro04, KPTS89, KÇC^{+}01, LOC^{+}16, LRIC15]. Data [LCI^{+}20, Lev88, Lev90, LD08, LWZ^{+}18, LK82, LRS12, LPMAA^{+}22, LPK05, LWY^{+}20, LB20, Ma00, Ma01, MP01, MRL13, MGO14, MD99e, MD00f, MHR^{+}09, MRB19, MUN^{+}10, Max90, MCS22, MBV17, Mur93, NE93, NB04, NFMD90, Nie93b, OCS20, Off20, OK09, Ozt15, PGV20, PSJ87, PJC88, PL19, Per21, PCG^{+}21, Pie87b, Pie87a, QC15, RPOC21, RFA^{+}22, RXB09, RFK^{+}13, Rei17, Rei23, RD90, RLCI24, Ric15, Rie93, RMDT96, RBS21, Ros89a, RC15, Rot03, RPM17, RM95, RSL^{+}19, RBD14, SC89, SE19, SBD^{+}17, SSB06, Sen99, She22, SP15, Sin17c, Sin19a, Sin19b, Sin21a, Sin22f, Sin23d, Sin24a, SMMS24, SH97, SS20, Ste07, TMWJK04, TBFGC23, UMO82, Use98, WYM10, Wan22, War04, Wei85, WBS^{+}05, WFOP85, Wix83, WL88, Won99, WFL^{+}00, Woo85, WKY23, YKFT84, YIK85, YNF^{+}92, YMS01, YCKF15, Zha96a]. Data [Zha21, Zha22d, ZE23, ZCBM14, ZYL^{+}17, Zhu12, ZWWS19, Ala92, BBP94, Bow93, LH92, MTS91, NB94, NFHL91, NBG^{+}94, PFH95, Pie87c, RBR^{+}93, Rob88, Rob91, RN94, RHC95, SSG95, SW88, Sch93b, Tre95, Wal91, Ano16-28, Ano18a, Ano20g, Ano21n, Ano21o, Ano21p, Ano21q, Ano22s, Ano22t]. Data-Base [UMO82]. Data-Compression [FM89b]. Data-Dependent [YMS01]. Data-Driven [BSS20, GB16, Ric15, RPM17, She22, ZE23]. Data2Vis [DD19]. Database [AL11, DR83, FDK12, KK94, KKAM84, Lea96a, WTR11]. Databases [Cat96, RLIB99, SKW^{+}12]. Databiting [RLCI24]. DataCad [MD99f]. Dataflow [Dye90, Fri20]. DataPort [Ano19-37]. Dataset [UZM^{+}23]. Datasets [BBT^{+}13, DMAS17, SBMK20, SB16]. David [CCG^{+}04, CF16, Die00]. Day [Del00b, Kop03, Whi95d]. Days [KD20, PS22]. DCT [Bli93d]. Deadline [Ano15c]. deal [Bli93d]. Dealing [Mac90]. Debriefing [GA97]. Debug [Lar10]. Debugging [CR02b, NRH^{+}22]. Decade [Die85, Ear98, LLG^{+}07, Pot04, Rho91]. Decades [Mac94]. Decay [MMGG12]. Decency [Hay96a]. Deception [CCB22]. Decide [Whi95d]. Decimation [RO96]. Decision [AZSW22, BJP23, CG85, FAZH17, Jae17, JKM23, MGP^{+}04, MHTD23, PKN^{+}22, RA93, SEAKC21, VS07, Den93]. Decision-Making [Jae17]. Decision-Support [FAZH17]. Deck [PPKM15, ZPA95]. Declarative [KSRS13, LJB^{+}13, NMRI06]. Declarative- [LJB^{+}13]. Deco [Iga11]. Decomposing [EAM22]. Decomposition [CCV85]. Deconstructed [Car16]. Deconstruction [Add95, Sin13a]. Decorations [Iga11]. Decoupled [LLZ13]. Decriminalizing [Sim94b]. Dedicated [SWDR16]. Deep [ASS^{+}21, Ano19-54, BRD^{+}20, CL18, FAECOG19, GPS^{+}21, GHCW21, HWP^{+}19, HTZ^{+}19, HPX^{+}21, KS20, KJR^{+}18, KKV^{+}02, LHFL20, LFT19, MMD^{+}19, SMDBR22, SW22, WYHS21, WFZB21, ZLMY19]. Deep-Learning [KJR^{+}18]. Deep-Learning-Assisted [SW22]. Deep-Learning-Based [GHCW21]. DeepCompare [MMD^{+}19]. DeepGD [WYHS21]. Defect [BHG23]. defective [MG88b]. Defense [CMZ01, DGTK07, ML16, ZHM^{+}03]. Deferred [KBGK17]. Defined [Cha88, HW90, Mil07, PvW07, Roc87]. Defining [BKt82, CZGR09, BKtH82]. Definition [SJSK19]. Defocus [JAPF07]. Deformable [CEO^{+}93, JK09, LMA^{+}17, MH04, SPMK14, SKP09, TW88, Whi04]. Deformation [CD12, FM22, KLR12, KY97, YCB05]. Deformations [LW94]. degenerated [Her86]. Degree [FM82]. Delaunay [Use93, FP93, FP95, Sug92]. Delivering [SADH23]. Delivers [MD99f]. Delta [HDKS95]. Delta3D [DMJ05]. demise [Cla91]. Democratizing [DL15, ZPL^{+}23]. Demonstration [MD99e, Mye87]. Demonstrational [SE19]. Demystifying [Cha09, Pan15]. Denoising [GHCW21]. Dense [AFYC03, CPCS08]. Density [SWvdWvW12, Sin23f, LGLD86]. Department [FL12, Fry83, Tau11c, Tau12c, Tau11d]. Departments [Pot12, Tau10c]. Dependent [DAG95, FM22, KIN^{+}05, KS01, NR03, STW^{+}08, YMS01]. Depicting [ND05]. Depiction [AD05]. Deploying [JKKM^{+}03]. Depolarization [Pal85]. Depression [Bli06c]. Depth [BB84, CR15, CXPG19, DNBC16, FPI84, HLCC08, HNC^{+}04, OKM84, PRdJ07, RTF^{+}05, Rok96, RR86b, WKB11, SM91]. Depth-Assisted [CXPG19]. Depth-Buffer [OKM84]. Depth-Buffering [RR86b]. Depth-of-Field [HLCC08, Rok96]. Derivative [TLT11]. Derive [Gla89, MML^{+}15]. Deriving [Gol03, HS04, SFCD99, ZZ05]. Desaparecidos [Han00b]. Description [Bar84, De 89, VG84, Wil87c, SWS91]. Descriptive [Her86]. Descriptors [MGW^{+}18, SP04a]. deserve [Ano15l]. Design [APB^{+}21, AXP07, Bar15, BQP^{+}15, DFQ12, BKLZ17, BF93, BYP21, BA13, BDM02, BCA^{+}20, Bol94, BA85, BH11, BJP23, BG82, BWHR99, BFBK00, CLG^{+}16, CM92, CRB^{+}02, CP89, CK85b, CW09, DP82, DEN13, Die11, DAG95, ESS82, GHS99, Gir87, Gla02e, Gla02d, GRWN87, HS04, HC21, HUS^{+}21, Hei94, HMRS06, HLC11, Hur82, Iga11, IM14, Iga19, II22, JJW^{+}99, Jer18a, JCM07, JKM23, KS07, KSM85, KSK21, Koc94, Koz10, LMWC17, LD97, Lic84, LMZ11, LSB18, MD99f, MSU^{+}21, Mar83, Mar84, MMMM21, May83, Mil03, MOD00, MCP^{+}06, Nie92, NZB00, OKKP13, Ovi03, PNPS03, Paq96, PRN15, PS09, PJC24, Pra84a, RR84, RDO^{+}96, SPJ99, Sad16, SFD16, Saw83, SR13, SR16, SHS^{+}21, SG86a, SI94, SB81, SRF^{+}06, Sin02e, Sin04a, Sin16d, Smi84a, SML^{+}09, TM88, Tho84, Til83b, Tre99]. Design [WXK22, WTLM15, WZL^{+}21, Wat06, Whi84, XY01, Zha21, Bar83, Bar81b, FJ87, KGM95, McG93, MF89, Nie86, Sim93b, Sim94c, WJ93, MD99d].

Design-to-Fabricate [SR13]. Designed [Sin03a]. Designer [Gol12]. Designing [AY16, AGR21b, BEW^{+}98, DKS^{+}05, FT16, GRvL^{+}23, Gus83, II10, KFH^{+}18, KKVL05, Kim84, Kul09, MR98, MJS20, Mar82, MG14, MIEL00, PCG^{+}21, PARV05, SS01, Shu23, Wan10, WSL^{+}20, WLL^{+}09, YSS^{+}16, SRS91]. Designs [SK02]. Desk [STJ^{+}20]. Desks [AMJ^{+}08]. Desktop [ADBR^{+}12, CBD^{+}93, DDG88, GP93b, LPMEPCM11, MD99f, RRB^{+}14, Sim93a]. Despite [Lea96e, KFH^{+}18, ZK10]. Detail [LZS04, RGW05]. Detailed [EHBPG04, Fin08, MSED23]. Details [Yan21, YLR17]. Detecting [JDH^{+}89, SSZP15, TMWJK04]. Detection [EWF02, FML23, FUF06, GASF94, ITSK06, KH98, RGL^{+}06, SMP^{+}19, SZZ^{+}18, ZHZ^{+}15, KmKK^{+}18]. Detections [PGM19]. Determinant [YT85]. Determination [WS90a]. Determine [Fre84]. Determining [FM82]. Deterministic [MD95b]. Detour [Add95]. DETOXER [NRH^{+}22]. Develop [Hay96c, PMQ21]. Developer [IFH^{+}03, Lue01]. Developers [ZTJ^{+}07]. Developing [CPC99, Hay96f, LZF^{+}14, Nav04, NdSM^{+}16, RMG16, Rei91, RdCFP^{+}20]. Development [CG85, Die11, Hay97, KIQ^{+}06, LPOLID18, LK08, MN97, MD00f, MKH20, Moh85, Oia82, RRF^{+}23, Rie22, SADC02, SHMN09, WAH^{+}21, WHR19, Wil86, Pap96]. Developments [GN83, HILW98, Laz87, Ped01, WHC94]. Develops [Hay95b]. Device [BSA^{+}04, MD00f, Mor05, PFA86, SB16, SQG^{+}05, War81b, War81a, Kor87]. Device-Independent [War81b]. Devices [CW93, DL15, DD04, FHKH06, HPG08, HBP^{+}07, KK02, LS08, NCO03, PARV05, WWV^{+}10, WSKY09, YIK85]. DEViSE [RXB09]. DevOps [Fis21]. Dharma [Sin21d]. Diagnostic [CDBL83, PKN^{+}22, WH99, WVL^{+}18]. diagramatic [FSB82a]. Diagrammatic [FSB82b]. Diagrams [Bli01a, Bli03c, Cho95, ITW01, LBD^{+}92, MMMM21, PvW07, TO83]. DIAL [FSB82b, FSB82a]. Dialogue [BKtH82, BKt82]. Dichromats [RGW05]. DICOMED [Hue83]. Didactic [KSK21]. Didn’t [Del00b]. Dietmar [OCS20]. Difference [HH08, Rhy03, WWW^{+}23]. Difference-Based [WWW^{+}23]. Differences [HPX^{+}21]. Different [Bli99a, Bli00b, Kos15, MHTD23, Sin23a, Bli89d]. Differentiable [Li22b]. Differential [HGMT17, SM86]. Differentiating [REF^{+}06]. DiffSeer [WWW^{+}23]. Diffusion [GHCW21, PQC^{+}12, ZGP^{+}13, CR95]. Digesting [Wea99b]. Digital [ARL^{+}10, ATS^{+}03, Ano16a, AB14a, BE11, BSS20, BRM^{+}02, BEAC^{+}18, BFW99, CFF01, CDLS19, CGMB20, Cha98a, CG84, DL15, Ear01, Enc16a, Gla02a, Gla00e, Gla00f, Gla03a, Gla03b, Gla04e, Gla04f, Gol12, HU83, IM11, Lea97a, Lea97c, LS84, Mac97a, MD99f, MD99d, MD00f, Mar07, MRD06, MTC07, Nak84, Nie07, OBO^{+}23, PMI06, Ped01, PB10, San01, SS19, Sch08a, Sch08b, SC06, Sin02a, Sin05a, Sin08d, Sin09a, Sin10b, Sin12d, Sin15b, Sin16c, Swe01, TK03, Tur02b, VP99, Wea99a, WGI15, WH99, Yan01, Zha96b, Bal98c, ARL^{+}10, PCZ^{+}09]. Digital-Content [IM11]. Digital-Image [VP99]. Digitally [Ann17, VVP^{+}11]. Digitization [KN07, SL00]. Digitize [SPM10]. Digitizer [MD99e, MD00f, TLL^{+}96, WP87, de 86]. Digitizers [MD99e, PTR^{+}98]. Digitizing [MD99d, Sin14a, ZLK^{+}22b]. Dimension [HH92b, Whi96a, MD99e]. Dimensional [CDBL83, ESS82, Gas16, Gro83, HKW99, MTT85c, MG04, Nak84, PM86, Rob87, TH84, WJ88, Wix83, XZZL19, PS86]. Dimensionality [SCS23]. Dimensions [Gol03, GZS88, Sin16b, FP95]. Dinnerware [SK02]. Dinosaur [Gla88b]. Direct [Bai01, CW93, EBN13, KP11, LiA10, NM08, SRF^{+}06, TT84, VP16, YNF^{+}92]. Direct-Touch [SRF^{+}06]. Directed [HWP^{+}19, KB82, NLvLM09]. Direction [MTT87b]. Directional [MCGS21]. Directionality [IMS03]. Directions [BCF^{+}08, ECS^{+}15, FIB^{+}20, RV83, Ros00b, SGE^{+}18]. Director [Kon12, MTTF85, SJB^{+}16, WKB11, MD99f]. Director-Oriented [MTTF85]. Dirty [Bli98b]. Disabilities [Hay96e, MFM^{+}16]. disabled [Sim94d]. Discard [Cai20]. Disciplinary [LRCT22, Sin01d]. Discipline [GH97]. Discomfort [CR15]. Discontinuity [LTG92]. Discourse [WKM^{+}04]. Discoveries [Ric15]. Discovering [CES07, DY09, Kun00, VFLJ02]. Discovery [Ano15-38, Bec97, DGTK07, PKJ^{+}07, RKJH99, Sin13d, Swe01]. Discrete [Ale21, YCK92]. Discretization [VS14]. Discriminant [Bli06a]. Discriminants [Bli00e, Bli01a, Bli02a]. Discussion [Tau11c]. Disease [WBA^{+}96]. Disentangled [SLSP22]. Disinformation [ŽLK22a]. Disk [WWL^{+}21]. Disorder [Mac09]. Disorders [HAB^{+}01]. Displacement [ABS^{+}05]. Display [Ano92e, Ano92f, Ano92g, Ano93l, Ano94h, Ano94i, Ano15b, AYS98, Bol94, BMC04, CW93, CDBL83, CG84, Cro87a, DT81, Eng86, FMM06, FGR85, FPT^{+}82, FPI84, GV81, Gol84, Gos90, GAW86, Gro83, HF18, Hag94, HCL^{+}20, HJS00, HU83, Hob81, HM85, HMRS06, IWF97, JLJ^{+}10, KIN^{+}05, Kan88, KL09, Lea96a, LGLD86, Lev88, LCC^{+}00, MD99e, Mur84b, Mye84b, Mye84c, NOK^{+}08, Nak84, NTM^{+}01, PPKM15, RR84, RCB^{+}05, RR86a, RR86b, Sad16, SPMK14, SFF^{+}00, SK86, SQG^{+}05, TT84, WAB^{+}05, WN82, YY16, ZWWS19, Zuc84, Bow93, GC91, KmKK^{+}18, KKM96, PFH95, Roc87, WB14]. Display-Rich [JLJ^{+}10]. Displaying [HML83, PJC88, PO84, Mas18]. Displays [Ano92e, Ano92f, Ano92g, Ano93k, Ano93l, Ano94h, Ano94i, Bal97, Ber83, BWP^{+}16, BB05, BBDM98, Bow93, BFBK00, CHC^{+}97, DSP^{+}14, FKKB05, FH84b, FI83, FL00, GM87, HW82a, Hag94, HCV^{+}13, JPSM13, KTA^{+}02, KKG^{+}13, KF05, Lea96d, LFOI15, MD99e, Mas18, Mor05, PMI06, RLOW13, RS14, SB97, SCH05, SSOC01, Sto01, WSK^{+}00, WLH^{+}12, Whi84, Zac84, Ano98j, BB13, LTSE92, MG88b, Rob88]. Disruptive [BAB^{+}23b]. Dissemination [BDC^{+}12, BAE^{+}18, Enc15c]. Distance [DJP^{+}08, FMM06, FUF06, Gla97a, PT92, SJSK19]. Distance-Similarity [FMM06]. Distances [LMWC17]. Distancing [CML^{+}23]. Distinguishing [Han14]. Distortion [CCF97]. Distributed [BS03, Bre84, EYY03, Kel84, LD97, LCC^{+}96, MAB05, MSA^{+}94, RDH^{+}02, Sty96]. Distributing [BBP94]. Distribution [CWB02, CPK06, FAZH17]. distributions [MTS91]. Dithering [WWC91, Bli94c]. DIVE [RBD14]. Diverse [CPA^{+}10, Enc15b, Zha22d]. Diversification [GD11]. Diversity [Gai17, LIS^{+}19, LRCT22, PRdJ07]. Dividing [NLvLM09]. Diving [Sin11b]. Division [OOHR14, OATO85, WKS86]. DiVRsify [PMQ21]. DIY [PGV20]. DLA [SW22]. DLA-VPS [SW22]. Dlabs [MD99e]. DML [MD99f]. DMorph [Gla03c]. DNA [Pot96]. Do [BG96, CGMB20, Dom00, DBDH05, EAM22, MKH20, TM05]. Doc [MD99f]. Doc-to-Net [MD99f]. Document [CRF^{+}21, EZM^{+}97]. Documentation [MPP81]. Documents [FSK07, HF07]. Does [KH14, MD99f, Rhy03, Bli92c]. DOF [HHLL14]. Dogs [Gla96e, Gla96b]. Dolby [Agr21a]. Domain [KFH^{+}18, Mil12, PP12, Sad16, TBO^{+}20, NL84]. Domains [SKS^{+}13]. Dose [ZYJ^{+}21]. Dot [Bet00, SB97]. Dots [LPMAA^{+}22]. Double [GI94, HG13, WR89]. Double- [GI94, HG13]. Double-step [WR89]. Down [Bli96d, Hay96d, Bli91b, Bli91c, Bli91d, Bli92c, Bli93b]. downers [Bli92a, Bli92b]. DPFrag [TS13]. dPVS [AM04]. Drafting [YST84, Kal87]. Drag [DiS14]. Drag-and-Drop [DiS14]. Drama [FEV^{+}01, NMRI06]. Draping [EWS96]. Draw [Whi99, Bli95a, Bli95b, Bli95e, SRS91]. Drawing [CS15a, CBB^{+}14, Eas90, HB97, II10, LBW^{+}17, OKS^{+}17, Pan90, SM04, Van84, WYHS21]. Drawings [AAB^{+}23, HE82, KSM85, WW07, YWR09]. Drawn [MIA^{+}08]. Draws [Hay96d]. Dream [CGMB20, Hay95b, Sin06e, Sin20b, MR99]. Dreamcatcher [BAAQ21]. Dreams [BAAQ21, Die00, Sin13b, TYK01]. Dreamweaver [MD99f]. DreamWorks [DiL15]. Dredging [KL18]. Dress [CLP18]. Dressing [Sin09b]. Drill [KYK^{+}24]. Driven [AGR21b, ALD^{+}06, BSS20, BHS^{+}24, FAECOG19, GB16, GHQ04a, HFZ^{+}13, HUS^{+}21, LSB18, MZC^{+}21, NTS^{+}07, PD09, Ric15, RPM17, SBRCD17, SADH23, She22, Sin23d, SOSG08, SK05, WYM10, WD15a, ZE23]. Driving [AP10, CKP96, PSE^{+}09, SAB^{+}23]. Drones [BKS19]. Drop [AKFPW20, DiS14]. Drops [SKN^{+}15]. DT2VIS [JSDL21]. Dual [TLS^{+}14]. Duck [Gla02b]. Dukane [MD99f, MD99d]. Dunhuang [HCL^{+}20]. during [KLR12, PB06, TN10]. Dust [HW95]. dv [MD99f]. dv/MockUp [MD99f]. Dynamic [AM04, ACL^{+}12, Bac16, BN95, BPB09, CR15, CdHM97, CWL^{+}10, Dai89, DDG88, EBG^{+}10, FM86, GAB^{+}18, GSG90, HLB^{+}07, JLW08, KN07, KK02, KP18, LFOI15, LX04, MBSD20, Mol18c, Rhe02, RBE08, SPMK14, SBM86, Ste18, TS13, TS22, WAH^{+}21, WWW^{+}23, Whi94b, Wil87b, XPH05]. Dynamical [OZH15, WGP97]. Dynamically [BMH98, CMBZ00, Mil12].

Dynamics [Ano94g, Bar97b, BTM17, CIF99, Gai04, KDH^{+}95a, KB96, MCP^{+}06, ND05, Sin07c, SBS10]. Dysplasia [WVL^{+}18].

E-Books [Har00]. E-mail [MD99d]. Ear [Ken20, PHAS20]. Early [BQP^{+}15, Enc23b, Fer01, Joe13, KD20, PS22]. Early-Phase [BQP^{+}15]. Earth [GS87a, JMOG99, SMMS24, Wor94]. Earthquake [CGT^{+}23, CCC^{+}07]. Easter [Bli88a]. Easy [MD99f]. EasyToy [LMZ11]. Ebb [Dil04]. Eckert [Ano19a]. Eco [Bar15]. Eco-Feedback [Bar15]. Economies [Sin19e]. Economist [Wil84]. Ecosystem [BNA^{+}22, BHPS12, Bas14]. Eddy [HYL^{+}22]. Edge [Ano22-30, CCM15, Gas16, Wei85, Mil93, Ano19f, Ano20c]. Edge-Aware [Gas16]. Edge-Based [CCM15, Wei85]. Edges [RTF^{+}05]. edging [HBW95]. Editable [LMZ11, CR91]. Editing [DBV89, GHQ04a, GHQ04b, LD08, MD99f, MTC07, NM08, NF91, Ols86, Pat12, SSTP18, VBHS11, GS12]. Editor [Ano93t, II10, Iga11, KZC^{+}11, Lev00, Tau10a, Tau13a, WMG00, Ano94j, Bre96a, Bre97, CR02a, Coo02a, Ear01, Eng97, GP91a, Her96, KNR93, Ma01, Ma06, MBB95, Mun02, Myk88, Nie93a, Par98, Rho93, Rho97b, RSZ96, Sta91b, Sto99, Tau13b, Tho01, Tho02a, TR94, Won99, vL06]. Editor-in-Chief [Ano94j, Her96, Tho02a]. Editorial [ECS^{+}15, Mol18c, Sto07, Sto23a, Sto23b, Sto23c, Sto23d, Sto23e, Sto24a, Sto24b, WB14, Tau10b, Tau11d]. Editors [Tau12c, Dev99a, TF01, AD05, AS10, Ano23-63, AR99, BO09, BC05, BES09, BS16, BDM02, BEG92, BNP99, CH10, CCR10, CES07, DM02a, DS07b, Ear97, EMTTT98, EFD05, ED93, EH03, FT00a, FSK07, FM03, FDC14, FL00, GE97a, GE98, GW98, GHR01, GH97, GP93b, HJ03, Hay96f, IZ07, JB02, KM00, KMS07, KF05, LK03, LS04b, LOB08, LR14, MP99, ML05, MT94, MEW98, Nie92, PP98, PB10, PGS09, RR96, RMD06, RR03, RA93, RS99, RBF95, RBT98, SB06, SC06, Sto15, SP04b, SBE95, TC15, WB00a, YCKF15, vHV09, vWNS10]. Edits [MD99f]. Educating [ZTJ^{+}07]. Education [Ano16c, Ano17a, Ano20v, BC05, CNC^{+}05, CPAV^{+}17, Dev99a, Dev99b, Dom00, DJP^{+}08, GAe^{+}24, HSHS05, Joe13, JMC^{+}02, JLMV06, MM21, PD21, RME14, Ric85, RR86a, RDDY07, SG11, Tak05, Whi97, XDJ23, Rho96, Tau11d]. Educational [AGR21b, CLP^{+}14, CLP^{+}15, DiS14, GWF98, LPMAA^{+}22, WMB85, vD05]. EduClust [FIB^{+}20]. EE [Bal97]. EE/CS [Bal97]. Effect [BQP^{+}19, Sin09d]. Effective [DJP^{+}08, Gus83, HE02, KFH^{+}18, LFMGPG23, Mur84a, RB00, SES05, WKY23, McC83, Rob88]. Effectively [Mac99a]. Effectiveness [GWA84, KCMK20]. Effects [ABS16, BHPB^{+}12, HLCC08, HG13, LX04, MD00f, MTT86, Rok96, She06, Smi84b, WTLM15, WP87, NINT89]. Efficient [ABS^{+}05, AA09, BTM17, Cam91, Cam17, CGIB86, DKS^{+}03, Gal89, Gas16, KHSE98, RR85, Van84, WEWL00, Woo93, XJW18, YZM^{+}11, FJ87]. Efficiently [IKMF12]. Efforts [CBS^{+}22]. Egg [Hec86a, Bli88a]. Egocentric [CLDW16]. Egress [CTJ^{+}17]. Egyptian [YYOiT92]. EIC [Sto23a, Sto23b, Sto23c, Sto23d, Sto23f, Sto24a, Sto24b, Tho02b]. EIC’s [Sto23e]. Elastic [HPG08]. Eleanor [LB20]. Electric [KKTI04, SWD^{+}98]. Electrical [DP82, EKM^{+}22]. Electrifying [Sin15b]. ElectroEncephaloGraphics [MM14]. Electromagnetic [BCMF16, IL83]. Electronic [BG96, CW09, Rho96, Whi95a, MLJ19]. Electronics [SB81]. Electrostatic [BDJ86]. Element [EDC96, IS95, SSB06, TCH^{+}96, TH14, UTMI11, YS83]. Elementary [CRA^{+}18, JMC^{+}02]. Elements [Fer01, RZSP04]. elevation [WF93]. Eliminating [Woo88]. Elimination [VCR00, WG95a, WG95b]. Ellen [CS17a]. Ellipse [Van84]. Ellipse-Drawing [Van84]. ellipses [WR89]. Elliptical [GH86]. Emaps [Erv93]. Embedding [DSK^{+}14, FAECOG19, dOYG04]. Embracing [Ano94g, LRCT22]. Emergency [Han00a, MG14]. Emergency-Medical-Service [MG14]. Emerging [BRCP17, EFD05, LLK^{+}03, Per16, Ree84, SP04b, Str04]. Emily [ARL^{+}10]. Emotion [LM10, MQM22, PB10, Pic00, YL11, ZGC14]. Emotion-Based [YL11]. Emotional [KMCMT09, RRV^{+}15, Smi84b, WSK^{+}19]. Emotions [MSU^{+}21, Sin10c, SH03, TN10]. Emphatic [Hua11]. Empirical [DBV89, SC89]. Employing [RSP15]. Employment [LSB18]. empowers [Sim94d]. Enabled [BBB08, NCO03, LS04b]. Enables [MD00f]. Enabling [CRD^{+}18, DRHL^{+}03, NR03, PTB^{+}15, SSZS21, Tac16, TBO^{+}20]. Encarnacão [WW95]. Encoded [DT81, LJB^{+}13, WBS^{+}05]. Encoding [KSY85, XPH05, YKR^{+}84]. Encodings [Kos22]. Encouraging [Enc11]. Encyclopedia [Bal97]. End [MZL^{+}21]. End-to-End [MZL^{+}21]. Endogenous [JW16]. Enemy [Hay96a]. Energy [BB82, GN83, MMMM21, FJ87]. energy-efficient [FJ87]. Engagement [KWK21, WSP^{+}20]. Engaging [HE02]. Engine [BEW^{+}98, BSA^{+}04, DMJ05, MD99e, MJC08]. Engineer [Jer18a, SB81]. Engineered [BQP^{+}15]. Engineering [AAB^{+}22, AJD13, Ano24j, BH83, CCVA02, CNC^{+}05, EJ98, FH84a, GV81, HS87, HDKS95, MSC89, MB05, MRBW12, Myk88, NK85, Paq96, RSS^{+}09, Ric85, SRE98, XDJ23, WJ93, Ano21a, Ano21b, Ano21c, Ano22c, Ano22d, Ano22e, Ano23b, Ano23c, Ano23d, Ano24a]. Engines [EKM^{+}22, Won08]. Enhance [DEN13, ITW01]. Enhanced [BSP97, GF86, PEMF92, SHY^{+}23, WC91]. Enhancements [MD99f]. Enhances [MD99d, SKMY09, YMK^{+}12]. Enhancing [GDS^{+}18, KPL17, MM21, Sra23]. Enough [WSL^{+}20]. Ensemble [LHFL20, RMP^{+}16, RFA^{+}22, OJ14]. Ensembles [BHS^{+}24, Cro18, GGS^{+}18]. Enter [Wor96b]. Enterprise [KBHP14]. Entertainment [Agr21a, CMZ01, HSHS05, KK02, Lat98, MEW98, Mor98, Ros00b, YSS^{+}16, ZHM^{+}03]. Entertainment-Defense [CMZ01]. Enthuse [SP06]. Entire [Cai20]. Entities [YK95]. Entity [KFBI22]. Entrepreneur [Ano16e]. Entropy [ZHZ^{+}15]. Entropy-Based [ZHZ^{+}15]. Entry [MD99e]. Entry-Level [MD99e]. ENTVis [ZHZ^{+}15]. Environ [RSS^{+}09]. Environment [AS96, BC19, BKRE88, BRG^{+}00, BMR90, CEMD22, CRB^{+}02, CHRU85, CdHM97, Dai89, DRHL^{+}03, DGA^{+}97, EM09, Gre86b, HUS^{+}21, Hee99, HAH^{+}23, JJW^{+}99, MN97, MD00f, MAB05, MSRMH09, MMR^{+}06, PHAS20, PMR^{+}09, PBM^{+}96, RMDT96, San01, SRE98, SGG^{+}22, SZP^{+}00, SKMY09, SDPW20, TWQ^{+}23, UFK^{+}89, VS07, YMK^{+}12, KF94]. Environmental [JW01, KP12, RA93, SCB^{+}04b, SSM^{+}07, Smi84a, Den93, RBR^{+}93]. Environments [AM04, AFYC03, AXP07, AA09, BWA96, BH16, BHG23, BMH98, BvDE^{+}99, BH06, CF16, CWZ99, CMBZ00, CWBS99, Ch’11, CMO^{+}99, ERS13, Ell94a, EBG^{+}10, FSP00, Fis02, FPW^{+}00, GHS99, GMNF97, GAB^{+}18, GSG90, Hai87, HB08, HLCC08, HH95, HGK^{+}04, HBG04, JK09, JXW^{+}08, Ken20, KP18, LJ96, LA20a, LA20b, LSMC04, LSV^{+}02, MRB19, NPB^{+}18, dSNRdLdSN16, PBVH15, RL07, RFK^{+}13, RDH^{+}02, Sty96, Sze96, TSK97, UCK^{+}04, VK17, Wei85, dOYG04]. Envisioning [Gel07a, WWF00]. EPA [RBR^{+}93].

Epidemiology [LRS12, UZM^{+}23].

Epilepsy [WKHL96]. Epinome [LRS12]. Episogram [CLDW16]. Equation [Bli05a, Bli06c, Bli06d, Bli06a, Bli06b, Bli07a, Bli07b]. Equations [Mac90, Man94]. Equivalent [BD89]. Era [BT00]. Ergonomic [CLG^{+}16]. Ergonomics [Mye84b]. Erratum [Ano19v, Ano19w]. Error [KT96, Lev00, Sin04e, Bli94c]. Errors [EKM^{+}22, JDH^{+}89, JS03]. Escher [Sin02f]. Essential [Ma23]. Essentials [Jer18a, Lar11]. esthetics [Hag91b]. Estimated [HNC^{+}04]. Estimating [PWH04, TT00, ZGP^{+}13]. Estimation [DNBC16, GDY^{+}11, SB02, UF11]. ETOMs [WW95]. Euclidean [FD03]. Euler [MS82, Wil85]. Europe [Enc00, Enc23a, Enc23b, Hay95b]. European [EGR94, Enc06, RMM^{+}91]. EVAC [PBM^{+}96]. Evacuation [CTJ^{+}17, MJC08, PB06, She82, WZM^{+}11]. evacuations [Sim95e]. Evaluate [MSDS23, Bro89]. Evaluating [AJK15, BBC^{+}20, CTJ^{+}17, DB16, HWP^{+}19, HKDL22, KTA^{+}02, Liv05, MHR^{+}09, MSC89, OKKP13, PGS^{+}08, RMP^{+}16, ST09, SS20, TM05]. Evaluation [AKFPW20, Ano84, BKS20, Bar84, BDH^{+}89, BEK^{+}03, BH84, FJL22, GHS99, GWA84, GAe^{+}24, KCB^{+}23, LKPM19, LTD16, LCG83, MSB^{+}06, NB93, OKS^{+}17, PGS09, SFD16, SM84, SM04, SRK90, USW17, WXK22, Mil93, NB94]. Evans [MD99d]. Event [CPG^{+}16, SOSG08, Ano14p]. Event-Driven [SOSG08]. Events [GMR^{+}23, KN07, MZC^{+}21, RXB09, SSZP15, Woo15]. Eventual [Sin14b]. Every [Jer18a]. Everyday [Gla03d, Jae17]. Everyone [Pac10, PP98, PMQ21]. Everywhere [BDG^{+}99, BB13]. evidence [Sim93d]. Evolution [Bak88, BEK^{+}03, DD84, Hag91b, HYL^{+}22, Hue83, MTT85c, PBV16, SMG^{+}21, Sin07d, Sin08c]. Evolutionary [CMT^{+}10, CC20, McG93, Wor96a]. Evolving [Ano20w, Ano21-32, Ano21-33, VMJT96]. ex [MM89, Mus90a]. Exact [AH95]. Example [FJP02, KS14, PHE21, Smi87, WC11, OD88]. Example-Based [FJP02, PHE21]. Exascale [Ahr15, Sin19e]. Excavated [ZZ99]. Excavation [KL18]. Excellence [Hay95b, Lea96a, Lea97d]. Exchange [RXB09, SH97]. Execution [PB85]. Executive [VG81]. Exercise [WD15b]. Exercises [WD15c]. Exergaming [SHMN09]. Exhibit [AWB^{+}97, Hay95b]. Exhibition [DOH^{+}20, IS95]. Exhibitions [HdlRL^{+}13]. Exhibits [Gel06]. Existing [DB16]. ExoMars [MGC20]. Exotic [SPW84]. Expanding [JLS15, Kas13, KGB14, MFL08, TLCC17]. Experience [CLP^{+}15, FIB^{+}20, GDS^{+}18, HMC^{+}20, KPL17, LRZ23, RCB^{+}05, Sin16f, Sun21, YSS^{+}16]. Experiences [CZAA23, DRHL^{+}03, Enc16a, HT04, KJA^{+}18, MBSD20, MSU^{+}21, MM21, SADH23, Smi87]. Experiencing [Ano98g, MFG^{+}17]. Experiential [PPVH23, SDM22]. Experiment [DPG18, Fei85, Use98]. Experimental [Lev10, SSB06]. Experiments [BW94, Enc15c, HBW^{+}24, JBCS23, TSF^{+}98, War88]. Expert [Hay97, MSW10, TM05]. Expertise [Kir21, TBO^{+}20]. Experts [Ano15-44, LCB^{+}23]. Explain [vdEAA^{+}23]. Explainability [EKS22, ŽLK22a]. Explainable [CL18]. Explaining [PKN^{+}22, TH01]. Explanations [EAM22, Fei85, NRH^{+}22, YCB^{+}22]. explodes [Ano94d]. Exploiting [GAW86, GP89, MZP^{+}95]. Exploranation [HPS20, YLT18]. Exploration [AAG^{+}22, AGBC09, Bac16, BBT^{+}13, CRD^{+}18, GN83, HBW^{+}14, HB16, II15, JKKM^{+}03, JSDL21, KK94, KP18, LOC^{+}16, MCGS21, MNC^{+}22, RRR^{+}16, RLCI24, SBD^{+}17, SvW09, Sin13d, TH17, UCK^{+}04, YLL10]. Explorations [BMC04, SCS22, SCS23]. Exploratory [FDK12, KFBI22, MZ22, PS09, SKW^{+}12, SGEF^{+}12, WVL^{+}18]. Explore [vdEAA^{+}23]. Explorer [KCW^{+}17, SHS^{+}21]. Exploring [BAB^{+}23a, BHAA^{+}13, DSW07, FF20, Hag90b, JKM23, JAPF07, Kro02, Lea96b, Ma00, MMMM21, MG04, Mun98, Pru90, PG93, Rhe02, RMD06, Rhy11, RSDP01, Sin15c, SOSG08, Swe01, Zha21]. Exposition [Lev00]. Exposure [CTM18]. Express [LM10]. Expression [LSQ^{+}12, LZR^{+}21, Sin08e, WKM^{+}04]. Expressions [Bli00d, EG98]. Expressive [CT04b, Dev13, LM10, LS95]. Extended [AVV20, OKM84, Chi87]. Extending [CS15c, CCF97, Til83b]. Extends [Sim96c]. Extensible [CMBZ00, LS08, MTTF85, MRD06]. Extension [DJP^{+}08, Ks22, PM86]. Extensions [Pic92]. Exterior [LHZ^{+}21]. ExteriorTag [LHZ^{+}21]. Externalize [vdEAA^{+}23]. Extra [Ano98j, Ano98h, Ano98i]. Extracting [Lar11, LYO02, LFPS01, PSX13, Rie93, ZTT99]. Extraction [LF87, LLR09, MGW^{+}18]. Extrapolation [LZR^{+}21]. Extraterrestrial [Pic95]. Extreme [CPA^{+}10, Ma09, WSP12, WSJ^{+}12]. Extreme-Scale [WSP12, WSJ^{+}12]. ExVis [Use98]. Eye [DLN05, Kos22, KW15, MKL97, Whi95a, HKW99]. Eyeglasses [ZGL^{+}17]. Eyes [MGC20, Sin03d, Sim94a].

F [Sin12f]. Fabric [EDC96, MD00f]. Fabricate [SR13]. Fabricating [BRL^{+}17, TMB^{+}13, Wan10]. Fabrication [BKLZ17, BA13, CMPA18, IO07, Mue18, KGM95]. FabSquare [BRL^{+}17]. Facades [Fin08, FZH^{+}13]. Face [CD12, FHH^{+}21, Gla04a, HML83, Mar82, Sin22a, YLR17]. Faces [Bee15, MQM22, PB10, TMB^{+}13, Sim94a]. Facial [EG98, LD08, LYO02, LSQ^{+}12, PHE21, Par82, RP95, RPM17, SBS10, TBP01, VPB^{+}91, ASW93]. Facial-Expression [LSQ^{+}12]. Facilitate [CLG^{+}16, JSDL21, Rei23]. Facilitating [RWB^{+}14]. Facilities [FKKB05]. Facing [SPS99]. Factor [EJ98, LLL^{+}13]. Factor-Based [EJ98]. Factored [WS04]. Factors [Atw84, Bol94, FWC84, Gre84, Liv05, LRZ23, Smi84a, SD83, XLJ^{+}21]. Factory [SBC15]. Fail [LMWC17]. Fail-Safe [LMWC17]. Failure [BQP^{+}19]. Fairness [FS89]. Fake [SK13, Bli88d]. Faking [Bar97b]. Fame [Wea98b]. Famous [MD99f]. Fan [FM89b]. Fang [Use93]. Fantastic [Abo00b]. Fantasy [Cha97a, Pac04]. Farewell [Bar05, Bli94b, Tau12a]. Fast [Ash03, EWS96, Gla84, GPR94, HHKK17, HJCW06, IKSZ03, Kan09, KSSL07, LGW^{+}12, LFT19, LJB^{+}13, MD99e, PQC^{+}12, RKJH99, Rob87, SW83, SK97, Sin10f, UO91, Web08, WW95, GFB87]. Faster [Del97, Del99a, FM89b, FB81, KS97, Gre91a]. Fatal [HMRS06]. Fault [FML23]. Favor [Lea96a]. Fear [HWK^{+}96]. Feasibility [PK87, SC21]. Feature [CM92, HR98, LF84, LTS^{+}13, LWZY19, LLR09, MD00f, Mar83, SW85, She22, SSC^{+}11, SMP^{+}22, WBS^{+}05, Won08]. Feature-Based [CM92]. Featured [Ano20-70]. Features [GZS88, JS10, KTD10, KLR12, KHP12, LF87, Sch17, WMG97, SG90]. Feedback [AVV20, BAB^{+}23a, Bar15, COM15, HH95, SK05]. Feel [MKH20]. Feels [GAe^{+}24, Wor94]. Femoral [GRWN87]. Fernbach [Ano17-31]. Fiber [Lea96c]. Fiction [Ano98g, NMRI06]. Fidelity [ADBR^{+}12, Bee15, CZAA23, MH04, NK14, SBMK20]. FIDS [Kan88]. Field [ASS^{+}21, Ano10d, Bas14, CR15, CMB94, FPH^{+}06, FTAT00, GK05, GAe^{+}24, HTZ^{+}19, HLCC08, HBW^{+}24, Rok96, Sch00, SWD^{+}98, WLH^{+}12, Wri83, YM95, HH91, HPvW94, PT92]. Fields [Bal97, DBV89, FUF06, IL83, JXW^{+}08, KVM^{+}05, KHP12, STW^{+}08, Tay02, UIL^{+}06, DH93, EW96, MJM93]. FieldVis [FOI^{+}07]. Fieldwork [PVV^{+}03]. Fighter [Gha12]. Fighting [CML^{+}23, MP23]. Figurative [CM16]. Figure [AGL87, BMW87, BWP^{+}16, HE82, WH97, Zel82]. Figures [CBD^{+}93, Fet82, Mac90]. Figurines [TMB^{+}13]. Filaments [MWY^{+}15]. Files [HDKS95, WH99, Ano98j]. Fill [Gla01b, Her86]. filling [Sch93a]. Film [Bux05, FH84b, KB14, MTT87b, MM21, Mor98]. Filter [GH86]. Filtering [Gas16, JBBL02, LKPM19, LR90, Sin14c]. Filters [BSP97, CRC^{+}02, Sin05b]. Final [Ano95e, Ano95f, Ano95g, Ano96j]. Financial [MD00f]. Find [Del99a, HMC^{+}20, MG92]. Finding [Cha99, GCS^{+}12, How00, NB04, TBFGC23, War04]. Fine [MWY^{+}15]. Finger [PBVH15]. Fingerprint [Han00b, PSX13, Sim94b]. Fingerprint-based [Han00b]. Fingertip [OSK02]. fingertips [Sim95c]. Finite [EDC96, OATO85, SSB06, TCH^{+}96, UTMI11, YS83]. Finite-Element [EDC96, SSB06]. Fire [Abo98a, FMMS03, KYK^{+}24, RHC95]. Firefighters [TSK97]. fireworks [LTSE92]. First [BEHt82, Die85, DMSJ23, Han00a, Han14, Hee99, HLCC08, Koz12, PRN15, BHK91]. First-Generation [Hee99]. First-Person [HLCC08]. Firsts [Sin12f]. Fit [Ano16-37, Ano16-38, VCBS97, Ano17z]. Fitness [BNA^{+}22]. Fitting [ADBW96, BMC84, BW94, VK17]. Five [Hib04, Jer18a]. Flag [GA97]. Flames [Dio97]. Flat [Bal97, Lea96d, MD99f, Zuc84, Kan88]. Flat-Panel [MD99f, Kan88]. Flatbed [Sch00]. Flatland [MD99f]. Flattening [SM86]. FlatWorld [PNPS03]. Flavor [CZAA23]. Flaws [HMRS06, TMWJK04]. Flexibility [Sin15d]. Flexible [BFB^{+}98, CPCS08, DD04, EWS96, EDC96, KB14, LSQ^{+}12, PBR^{+}14]. Flight [AHDG00, Gla00g, LK95b, Sch81, Sin17d, Yan85, ZMR^{+}88, CG94]. Flipping [LPOLID18]. Floating [Bli97a]. Floating-Point [Bli97a]. Flodar [Ano98j, Tre98]. Flood [Hua21]. Flooding [ZCS^{+}06]. Floor [Del99b]. Florentine [Abo99b, BRM^{+}02]. Flow [Dil04, FOI^{+}07, FMMS03, GCS^{+}12, GHCW21, Gün20, HTZ^{+}19, Hed84, IG98, JS10, KHP12, STW^{+}08, Sin11b, SFL^{+}04, Tre98, War08, YIK85, van93, vHdP94, vdEAA^{+}23, JM95]. Flows [KSRS13, RGG19, HH91]. Fluid [Bro15, CdHM97, Gai04, KC05, NL02, vHdP94, HH91]. Fluid-Based [NL02]. Fluids [DYS^{+}07]. Flurries [SFL^{+}04]. Fly [MHC98]. Flying [HWK^{+}96, MKH20]. Focal [WVL^{+}18]. Focus [Ano95e, Ano95f, Ano95g, Ano96j, Ano14c, Ano15d, Ano15e, Ano16g, CS17a, Ebe05, KMH02, MGN^{+}11, SP15, JSDL21]. Focused [GLC11]. Focusing [GLRK06]. Foe [Bal97]. Fold [Gla98c]. Folk [VVP^{+}11]. Follow [ZWWS19]. Follower [SKN^{+}05]. Following [Dil04]. Fond [Bar05]. Font [MTS^{+}20]. Fonts [HBBG95, HH01]. Food [MMMM21]. Foodsheds [TWQ^{+}23]. Force [ACL^{+}12, HF18, HWP^{+}19, HH95, SK05, Bro89]. Force-Based [ACL^{+}12]. forecasting [TC96]. Forensic [UBS^{+}12]. Forensic-Case [UBS^{+}12]. Forensics [Han00b]. Forest [ZK10]. Forget [Del05]. Form [Cas87, CS07, HT98, Kim84, KKAM84, LW94, LFOI15, SLC24, SB86b, Ano96h, Ano97g, Chi87]. Form-Based [KKAM84]. Formal [Dom00, RDDY07]. Format [Agu86, BE11, FL00]. Formation [SMG^{+}21]. Formations [GD13]. Formgraphics [KKAM84]. Forms [FH85, Sei93]. Formulas [Wil85]. Formulation [Bli03a]. Fortran [Bli94b]. fossil [IP92]. fossils [KDH95b]. Foster [Gai17, SFD16]. Fostering [BABB^{+}21, Dom12, RGRLFM21]. Foundations [HS04, HU83, RET^{+}94]. Fountain [YL09]. Four [Bli04c, Mac94, Wix83]. Four-Dimensional [Wix83]. Fourier [Gla99d, MWM87, SP04a]. Fourth [HH92b]. FoxTrax [Cav97]. Fracas [Sin05c]. Fractal [CR94, Har96, OF14, Sin05c, Sin06b, Sin09c, Sin15c, Sin16a]. Fractals [MD95b, Sin05f, Sin14f, Van10, GBCK93]. fractional [Bli88b]. Fragile [CT09]. Fragmentation [TS13]. Fragments [Sin08b]. Frame [FM85, GF86, PEMF92]. Frame-Buffer [GF86, PEMF92]. Frameless [TLL^{+}96]. Frames [AH95, MED^{+}93]. Framework [BWD^{+}19, CXW^{+}20, DKS^{+}05, DM02b, GTHL99, HW18, JAC^{+}08, KSRF11, KSK21, KGR^{+}18, LGK^{+}06, MD02e, RBD14, SPA^{+}16a, Tak85, VP99, WYHS21, vdEAA^{+}23, MD99e, MD99d]. Franca [SR16]. Fraunhofer [Ear98]. Frederick [Whi23]. Free [AHDG00, Cas87, CS07, COM15, Dai89, DDdSC11, FI83, HT98, Kim84, LW94, Lea96d, PLS^{+}16, SLC24, SB86b, Chi87, EW96]. Free-Form [Cas87, CS07, HT98, Kim84, LW94, SLC24, SB86b, Chi87]. Free-Throw [COM15]. Freeform [Men94]. Freehand [CEMD22, RLOW13]. Freestyle [GD13]. freeze [MED^{+}93]. freeze-frames [MED^{+}93]. Fresh [Ano14d]. Friend [Bal97, Bli04a, Hay96a]. Friendly [Pat12]. Frieze [Gla96a]. Front [Ano14e, Ano14f, Ano15g, Ano15h, Ano15j, Ano16h, Ano16i, Ano16j, Ano16k, Ano16l, Ano16m, Ano17e, Ano17f, Ano17g, Ano17h, Ano18h, Ano18i, Ano18j, Ano18k, Ano18l, Ano18m, Ano19x, Ano19y, Ano19z, Ano19-27, Ano19-28, Ano19-29, Ano20x, Ano20y, Ano20z, Ano20-27, Ano20-28, Ano20-29, Ano21-34, Ano21-35, Ano21-36, Ano21-37, Ano22-31, Ano22-32, Ano22-33, Ano22-34, Ano22-35, Ano22-36, Ano23-35, Ano23-36, Ano23-37, Ano23-38, Ano23-39, Ano23-40, Ano24k, Ano24l, FGR85, GC91, Ano14g, Ano14h, Ano14i, Ano15f, Ano15i, Ano15k, Ano17d, Ano17i, Ano21-38]. Front-to-back [GC91]. Frontier [LCG^{+}02, Swe01, Tur03]. Frontiers [Ma23]. Frugal [RV10]. Fugue [Bli97b]. Fujimoto [Nea89b]. Fukuda [Sch21]. Fukuda-Style [Sch21]. Full [BC87, CXPG19, JF99, KK02, hKKSK17, LaV07, PMR^{+}09, PABS86]. Full-Body [KK02, hKKSK17, PMR^{+}09]. full-function [PABS86]. Full-Spectral [JF99]. Full-Wave [BC87]. Fully [WSKY09]. Fun [Bli96b, Bli98f, Sin09c]. Function [CKPB16, GHQ04a, HP91, LV23, Wil13, PPK05, PT90, PLB^{+}01, SDT^{+}93, Van10, GZM97, Har96, PABS86]. Function-Based [PPK05]. Functional [Eas90, SMWS88, OG91]. Functionality [MD99f]. Functions [CPK06, DMR83, Eng86, FLNR90, MBV17, PP12, Gol83c, MTS91, Rou87]. Fund [Ano97i, Ano23-47, Ano23-48, Ano23-49, Ano24p, Ano22-51, Ano22-52, Ano22-53, Ano22-54, Ano24q]. Fundamentals [Bre96b]. Funding [MML^{+}15]. Funny [Del05]. Further [Bli98e, Ric15]. Fusion [dOGDM^{+}19, Sin04f, WKHL96]. Future [AAB^{+}22, AM90, Arn14b, Bee82, Bli00a, BSHB01, CWZ99, CMZ01, Enc14b, ECS^{+}15, FE97, FIB^{+}20, GOPK18, Har00, Kas00, MKH20, May83, O’S11, OJ14, Per16, PBSB08, RV07, Ros00b, SF20, SGE^{+}18, SB03a, STJ^{+}20, TYK01, WFR^{+}00, Whi95d, Ano14p, Wil91]. Fuzzy [CMT^{+}10, FS86, Lea96e, Zha96b].

Gain [KGR^{+}18]. Gait [van96]. Gaits [AP11]. Galant [Sta17]. Galleries [Gel06]. Gallery [Sin11c, Swe01]. Game [Abo99a, BRD^{+}20, BEW^{+}98, Bli99a, CMT^{+}10, DMJ05, DB16, Del00c, DEN13, Die11, ENDAD13, EH16, GAe^{+}24, KIQ^{+}06, LC16, LPOLID18, MKH20, MTN18, MJC08, NOK^{+}08, RdCFP^{+}20, RRV^{+}15, SSTP18, Sin22c, SOM^{+}10, ZTJ^{+}07]. Gameplay [QCP^{+}10]. Gameplayers [DY09]. Games [AGR21b, BES09, BLH^{+}08, CZP^{+}10, Got09, Hay96e, IKSZ03, SKUT^{+}10, LaV08, LA20b, LFMGPG23, Saw08, SS19, Sch08b, Wea98a, ZED^{+}09]. Gamification [HUS^{+}21]. Gaming [KPL17]. Gamut [NM08, SPA^{+}16a]. Gamut-Mapping [SPA^{+}16a]. gamuts [Rob88]. GANs [Sin21d]. Gap [PBV16, PvW07]. Gaps [vW06]. García [FBML15]. Garden [WWF00]. Gardens [MRL13]. Garment [SCOSL13]. Gas [Ano94g]. Gauge [BWP^{+}16]. Gauge-Figure [BWP^{+}16]. Gaussian [LJR96]. Gaze [JSH16, LM10]. Gears [Sin01b]. Gen [BLH^{+}08]. Genealogy [Sin16a]. General [Ber86, Bli06c, GKB89, KHP^{+}11, KKGH90, RR85]. General-Purpose [GKB89]. Generalization [HB06]. Generalized [Gol84, HF07, RO96, Roc87]. Generated [Gla96b, HM85, JCM07, MIA^{+}08, PF83, Yan85, WFG92]. Generates [MD99e]. Generating [AH95, BB84, GD13, IP92, IMS03, KB82, LBW^{+}17, MHR^{+}09, MSH^{+}08, Pic95, Rok96, SR85, SRL06, SM95, Shi83, YWR09]. Generation [AK84, CCT^{+}20, DD19, Eng97, Gai17, HB84, Hee99, HBBG95, ILC10, JLW08, JSDL21, KSY85, Ler82, Ma07, MOP89, MTS^{+}20, RY06, RO86, RBE08, Rot03, Sch81, WEWL00, YS83, ZLMY19, ZTJ^{+}07, vD05, GFB87, Jam87, Ols86, WS92, Whi87, WR89]. Generative [BRB17, BM24, CCT^{+}20, CM11, Hei94, Shi17, VT24]. Generator [Abo98b]. Generic [Ker91, KSRS13, LV23]. Genomics [LZF^{+}14]. Genre [BRCP17]. Geographic [AAW17, EJ98, MC85, SMP^{+}19]. Geographical [Kub84]. Geolocalized [ISMGG22]. Geolocated [vLBW17]. Geologic [KTD10]. GEOMAP [Kim84]. GEOMAP-III [Kim84]. Geometric [AEFQ96, BD89, BD90, BH84, CS06, DM02b, DR83, Elb05, Gla92, GGG08, GMTF89, Gol00, Hei94, KHP12, KPTS89, Lee08, LS21, MD02e, MGO^{+}16, MK84, Nie92, Ozt15, RSDP01, Ros89b, SG86a, Tak85, Tau12b, UTMI11, WW86, Wil85, Bar81b, Car87, FH85, GJ93, GU86, GBCK93, Kal87, Til81]. geometric-modeling [Kal87]. geometric-modeling/drafting [Kal87]. Geometrical [DM02b, MD02e]. Geometrically [BD90]. Geometry [Bee15, Ben99, Bli05c, Cam17, FD03, FWSN86, GTHL99, Mil99a, Mil99b, RR86b, SW83, SR16, SM86, SG86a, WKS86, Cam91, GG92]. Geometry-Based [Ben99]. Geophysical [GDF^{+}14]. Georges [LFE^{+}18]. Geoscience [JLMV06, SSZS21]. Geoscientists [FPH^{+}06]. Geospatial [DMAS17, GDS^{+}18, HJF06, Jon07, KPSN04, MMR^{+}06, UZM^{+}23]. GeoTime [PCH^{+}07]. Geovisualization [Kra06, MGP^{+}04, RMD06]. GeoWall [JLMV06]. Germany [Enc23a]. Gestalt [RC15]. Gestural [KDMH17, RLOW13, Mak87]. Gesture [BPB14, OSK02, SZP^{+}00, KmKK^{+}18]. gesture-based [KmKK^{+}18]. Gestures [CPAV^{+}17]. Get [Abo99a, Ano15l, Ano20-32, Ano20-30, Ano20-31, Ano21-39, Ano21-40, Ano21-41, Ano21-42, Ano22-37, Ano22-38, Ano22-39, Ano22-40, Ano22-41, Ano22-42, Ano24m, BG96]. Getting [Fol00, Gla02c, Gro06, Par98]. Gharavi [CS19]. Ghost [Bli98a, ILC10]. GI [MBSD20]. Gift [Ano18-37, Ano18-38, Ano18-43]. Gigapixel [BE11, PPKM15]. GII [DB96, GB96b]. Girls [Wea98a]. GIS [Wor94]. give [Sim94a]. Given [FM82]. GKS [BEHt82, BDDH95, GK86, MN84, PM86, Sch86, Spi86]. GKS-3D [PM86]. GKS-94 [BDDH95]. Glance [BFH02]. Glass [BFR21]. Glasses [Del05, Hay95b]. Glassner [BG96, Gla02a, Gla96e, Gla96b, Gla96a, Gla96d, Gla96c, Gla97a, Gla97e, Gla97d, Gla97c, Gla97f, Gla97b, Gla98e, Gla98b, Gla98a, Gla98f, Gla98c, Gla98d, Gla99d, Gla99f, Gla99a, Gla99e, Gla99c, Gla99b, Gla00a, Gla00e, Gla00f, Gla00b, Gla00c, Gla00d, Gla01b, Gla01f, Gla01a, Gla01c, Gla01d, Gla01e, Gla02e, Gla02d, Gla02f, Gla02b, Gla02c, Gla03a, Gla03b, Gla03e, Gla03f, Gla03c, Gla03d, Gla04a, Gla04b, Gla04e, Gla04f, Gla04c, Gla04d, Gla05a, Gla05b]. Global [GB96a, GOPK18, JLJ^{+}10, MYK^{+}08, MED^{+}93, Sim96c, Whi20, ZCBM14]. Globe [ISMGG22]. Glove [MGO^{+}16, SZ94]. Glove-based [SZ94]. Glyph [LMWC17]. Glyphs [CE98]. GMSOLID [BG82, Sar82]. GNN [WYHS21]. Go [Ano18-50, Hay96a, Mac97a, Ano14-39, Ano15-43, MD99e]. Goal [KB82, Lea96a, Mas98, MOD00, NLvLM09]. Goal-Based [MOD00]. Goal-Directed [KB82, NLvLM09]. Going [Gla97a, Sin19f]. Goings [Tau11b]. Golf [LY92]. Goodbye [ST10]. Goodsell [CF16]. Google [Jon07, Zhu12]. Gorilla [AWB^{+}97]. Government [Hay95b, Ros21]. GPS [Sch07, SAHK13]. GPU [Bai09, Bai11, Bai13, IO07, JCM07, RNNTD14, RBE08, SWS07, SGG^{+}22, Ste18]. GPU-Accelerated [RNNTD14, SGG^{+}22]. GPU-Based [JCM07, RBE08, SWS07]. GRAAL [JAC^{+}08]. Grabbing [KP14]. Graceful [Alp93]. Gradient [ZGP^{+}13]. Gradients [CPM^{+}18, KGM95]. Grading [AJK15]. Graduate [EE12]. Graduates [NK94]. Grain [Sin11a]. Grand [CZP^{+}10, FKK^{+}10, WDC^{+}23, vL16]. Grandpa [Bli92c]. Graph [AAB^{+}23, BVD19, DFM^{+}02, HWP^{+}19, HKDL22, HAH^{+}23, IK15, OKS^{+}17, Pat12, RBD14, SR16, SZMJ12, SSZP15, TPR^{+}22, WYHS21, WWW^{+}23, WCG^{+}11, WZQ^{+}19, WKY23, SK91]. Graph-Based [RBD14, SR16, SSZP15]. Graphic [Ano92h, Ano93m, Ano93n, Ano93o, Ano93p, Ano93q, Ano94k, Ano94l, Ano94m, Cai20, GF94, MOP89, Mar83, Mar84, RC85, Sim93b, Sto00, War00, McC83, MF89]. Graphical [Agu86, BU23, DiS14, FPJ04, MTT83, SWD^{+}98, Wil84, Alp93, OD88, PM86]. Graphically [Agr09, Cha09, CW09, GDS^{+}18, Mac09, Saw08, Sto21]. Graphics [Agr18a, Ake89, Ama87, AVV20, AS11, Ang17, Ano03a, Ano04a, Ano16f, Ano19-54, Ano21k, Ano23x, Ano23y, Ano23z, Ano23-27, AM90, Arn14a, Arn14b, ASP^{+}86, BMC84, BC87, BC05, Bal98a, Bar97a, BDJ86, Ber83, BBK82, BT00, BB82, BM86, Bli96d, Bli98c, BH81, BDM02, BM03, BA85, BEHt82, Bor89, BBS09, BEG92, CDLS19, CPAM08, CH81, CWZ99, CB86, CHP21, Car83, CM86, CR02a, Cha85, CG85, Dal19, De 11, Del02, Den93, Dev00, DDG88, DD84, Ear98, Ell94b, Enc00, Enc06, ECS^{+}15, Enc23a, Enc23b, Eng86, Eng00, Eng20, Fer01, Fet82, Fig01b, FWC84, FAB^{+}98, FHH^{+}21, FPT^{+}82, FPI84, FL81, Gar15, GN83, GGB23, GS87a, Gel07a, GB96a, Gla03d, Gol00, GWH05, Got82, GKB89, Gro83, Gus83, GVA86, Hai87, Hal87, HJ03, Han00a, Han01]. Graphics [HH82, HJO86, HD01, Hob81, Hoe12, HW82b, Hue83, Ike84, IS95, JAC^{+}08, KB09, KTA^{+}02, Kas10a, Kas10b, Kas11a, Kas11b, Kas11c, KD20, KKAM84, Kri01, Kri06, Kro02, KR86, LLK^{+}03, LK03, Lav83, LLML16, Lea96d, LR90, LMP89, LA20b, LPMEPCM11, LSFW83, LCG83, Mab82, MC85, Mac99a, ML16, Mac81, Mac83, Mac85, Mac94, MEW98, MD99e, MD99d, Mac00a, MNR^{+}91, Mar83, Mil99b, Moh85, MP23, MM14, Mye84c, Mye85, Nea86, NG96, Pal81, PG04, PZD^{+}18, Pot96, PFA86, PARV05, PK08, Red01, Ree84, RR84, RME14, Rei23, RGA^{+}83, Rho90, Rho93, Rho97a, Rib98, Rie81, RA93, RS99, RMM^{+}91, RR86a, Ros00b, RFN89, SB86a, SvNG89, SDS^{+}11, Saw83, SMWS88, SB97, SG11, SF20, SRL06, She82, SBM86, Ska86, SG86b, SK86, Ste18, SP06, SRS83]. Graphics [SDS95a, Sto21, SBE95, SD83, Tak05, Tau10c, TM88, Tho00, Tur02a, UMO82, USW17, VG84, VG81, Voo89, WMS92, War81b, War81a, War85, WK85, WMB85, Whi97, Whi87, Whi84, WSKY09, Wor96b, XDJ23, Zha22b, de 86, Bli91b, Bli91c, Bli91d, Bli92c, Bli93b, Bre96a, Bri90, Bro89, DS87, FG88b, GB96b, HH92a, Hod92, Jam87, Mac91, MG88b, Pin91, PG93, RKRW87, Rho91, Rho96, SW88, Shi86, Sta91a, Sta91b, SDS95b, WWC91, Woo91a, Gre86a, Hay95b, Hay96b, MTT85a, Pea88]. Graphics-Related [SDS^{+}11]. GraphMatchMaker [TPR^{+}22]. Graphoto [PKM18]. Graphs [HPX^{+}21, HGK10, Mun98, PvW07, RY06, SSTP18, Shi83, Sow00, vLBW17]. Grasping [AKFPW20]. Grass [BPB09]. Grassland [JWHW12]. Gravity [hKKSK17]. Gray [MKFN87]. Gray-Scale [MKFN87]. Grayscale [HBBG95]. Green [LY92]. Greetings [Sto23f]. Grid [AGP^{+}22, BS03, WS90a, WHC^{+}14, FP93, Rou87, Use93, HJ03, JKKM^{+}03, NR03, SB03a, Sin03e]. Grid-Distributed [BS03]. Gridded [DP00]. Grids [ADBR^{+}12, Chr83, DMR83, FS01, ILC10, KS97, MWY^{+}15, RSK^{+}06, Kla93]. Grinstein [LFE^{+}18]. Ground [FZ03, Sim95b]. Groundwater [LHM23]. Group [FY97, GD13, Hil82, Lea96a, MCP^{+}06, TSF^{+}98]. Groups [MZP^{+}95, PE09, Gla96a]. Groupware [HMRS06]. Growing [MFL08]. Growth [KSSL07]. Guard [Tho02b]. Guest [WB14, AD05, AS10, AR99, BO09, BC05, BES09, BS16, BDM02, Bre96a, BEG92, Bre97, BNP99, CR02a, CH10, CCR10, Coo02a, CES07, DM02a, Dev99a, DS07b, Ear97, EMTTT98, Ear01, EFD05, ED93, EH03, Eng97, FT00a, FSK07, FM03, FDC14, FL00, GE97a, GE98, GW98, GHR01, GH97, GP91a, GP93b, HJ03, IZ07, JB02, KM00, KMS07, KNR93, KF05, LK03, LS04b, LOB08, LR14, MP99, Ma01, Ma06, ML05, MT94, MBB95, MEW98, Mun02, Myk88, Nie92, Nie93a, PP98, Par98, PB10, PGS09, Rho93, RR96, Rho97b, RMD06, RR03, RA93, RS99, RBF95, RSZ96, RBT98, SB06, SC06, Sta91b, Sto99, Sto15, SP04b, SBE95, TC15, TF01, TR94, WB00a, Won99, YCKF15, vHV09, vL06, vWNS10]. GUI [Rao97]. Guidance [BBB08, COM15, SKS^{+}96b]. Guide [IFH^{+}03, Mac81, Mac83, VIK^{+}02, WXK22]. Guided [CCT^{+}20, FHH^{+}21, NCQ^{+}11, XZWH14]. Guidelines [HS04, PS09, SD83]. Guides [Hay97, LTB22, MSW10]. Guiding [HBB20]. GUIDON [RC85]. GUIDON-WATCH [RC85]. GUIs [LS08]. guitar [Ano97i]. Gunnery [LLZ13]. Guys [BvRS^{+}11]. GWB [MS82]. GyroWand [HROIK16].

Habitats [MR98]. Hair [XY01, WS92]. HairStudio [XY01]. Halos [Gla96e, Gla96b]. Hamming [LMWC17]. Hand [HFZ^{+}13, MIA^{+}08, KmKK^{+}18, LK95a]. Hand-Drawn [MIA^{+}08]. Hand-Driven [HFZ^{+}13]. Handed [EZM^{+}97]. Handheld [BVRF05, SPT^{+}15]. Handler [Goo86]. Handles [KP14, MD99d]. Handling [BHP07, ESS82]. Handprinted [WB85]. Hands [BPB14, Kas12a, NF11, Sin21b, Sim94a]. Hang [CZP^{+}10]. happily [Ger96]. Happiness [Got09]. Haptic [BSA^{+}04, CT04a, DYS^{+}07, HS04, KSD04, KP18, KW04, KL09, LPLN04, LGM13, LS04b, LS04a, LO05, LLZ13, LX04, MH04, NZS21, SS97, SCB04a, Sin04d]. Haptic-Based [LLZ13]. Haptic-Rendering [KSD04]. Haptics [BDK^{+}04, GHQ04b, SM04, Sin04b]. HaptiStylus [ABS16]. Hard [JB02, Kas12b]. Hardware [ABS^{+}05, Ano94n, Bar91, BFP86, Eng00, GWH05, Hob81, Kas13, KY97, Laz87, MD96c, Mac96, Mac97b, Mac98, MD99f, MD99e, MD99d, Mac99b, MD00f, Mac01, Mac03, Mac04, SR13, WBS^{+}05, Pap96, Sta91a]. Hardware-Assisted [Bar91, KY97, WBS^{+}05]. hardware-independent [Pap96]. Hardware/Software [MD96c]. Harlan [Ano18n]. Harlem [JLC^{+}02]. Harmful [BT07]. Harmonic [MG88a]. Harnessing [Int00, RTF^{+}05, SQG^{+}05]. Hats [Sin18f, Whi94a]. Having [Enc17b]. Haystacks [War04]. Hazard [Hua21]. Hazards [ZR24]. HD [HKW99]. HD-Eye [HKW99]. HDR [SPA^{+}16a]. Head [HCL^{+}20, MD00f, RS14, TG21, SK13]. Head-Motion [TG21]. Head-Mounted [HCL^{+}20]. Head-Pose-Based [RS14]. Heading [PPA14]. Heads [Hay96a]. Healing [Tak05]. Health [CLP18, Han01, LTD16, MGA^{+}21, RPOC21, Saw08]. Healthcare [GB16, PPVH23]. Hearing [GDS^{+}18]. Heart [WOF96]. Heart-Vessel [WOF96]. Hearts [Ano97i]. Heat [SMP^{+}19]. Heidelberg [MMS^{+}91]. Heights [Ano15-44, GE98]. Helical [PR08]. Helicopter [MG14]. Hell [Dio97]. Hello [ST10]. Help [Ano18o, LSB18, NAB^{+}23, Rei17]. Helping [Del02, DASF13]. Helps [Del99a]. Her [Sin11c, Sin12c]. Here [Bar00, Del98b, Bli89c]. Heritage [Arn14a, Arn14b, BT00, EHBPG04, GBT^{+}02, RSZ20, Sin14a, SPM10]. Herzog [Ano08b]. Heterogeneous [AOH^{+}14, PGM19, RFK^{+}13]. Heuristic [SFD16]. Hey [BG96]. Hidden [DOH^{+}20, HE82, MR87, OKM84, Wei23]. Hidden-Surface [MR87, OKM84]. Hierarchical [CCV85, FUF06, GPW^{+}19, Gla00b, GD93, HLE04, ITSK06, LWZ^{+}18, MKFN87, SW88]. Hierarchies [GS87b, SL00]. hierarchy [FH85]. High [ADBR^{+}12, AJD13, Ano15m, BL14, Bee15, BB82, CWWW21, CB86, CZAA23, CCC^{+}07, CNL00, DBS96, Del98a, DASF13, DD84, FT00b, Gas16, HBW95, HKW99, Ike84, JLW08, KKV^{+}02, LS84, MD99e, MTT83, MH04, MQR99, MMP^{+}07, NK14, Rot03, SFF^{+}00, SBMK20, SGC21, TO83, War85, XPH05, DWL^{+}92, Sim95b]. High-Dimensional [Gas16, HKW99]. High-Dynamic [JLW08]. High-Dynamic-Range [XPH05]. High-Fidelity [ADBR^{+}12, CZAA23, MH04, NK14, SBMK20]. High-Level [MTT83, TO83]. High-Performance [CB86, DD84, FT00b, War85]. High-Precision [MMP^{+}07]. High-Quality [Rot03, HBW95]. High-Resolution [CCC^{+}07, KKV^{+}02, MQR99, SFF^{+}00, Sim95b]. High-Speed [DBS96, Ike84, LS84]. High-Tech [MD99e]. Higher [Dwy16, SCS23]. Higher-Order [Dwy16]. Highlights [iAH03, GH97, vWNS10]. Highly [XZWH14]. Highs [Sto21]. Historic [ATS^{+}03]. Historical [MSED23, RV82]. Histories [DOH^{+}20]. History [Ano21i, Ano21j, Ano22k, Ano22l, Ano22m, Ano23j, Ano23k, Ano23l, Ano23m, Ano24d, Ano24e, CR02a, Hag90d, Lor20, Woo91a]. HLS [FS86]. HMDs [CR15]. Hockey [Cav97]. hole [ABB^{+}93]. Holes [HT98, Sch93a]. Holistic [HAH^{+}23]. Holly [II10]. Hollywood [PNPS03]. Holocaust [Wei23]. Holograms [Bim06, PF83]. holography [HH92a, Sch91]. Holon [Sin24b]. Home [Agr21a, Ano15m, Bar15, Lat98, Lea96a]. homes [FJ87]. Homogeneous [ERS13, Rie81, Bli93b]. Homographies [PXC02]. Honor [Hay95b]. Honoring [Ano20-35, Ano20-33, Ano20-34]. Hopefully [SF20]. Hopper [Ano88a]. horizon [Bli95e].

Hosek [Wil13]. HOST

[Ano20-36, Ano20-37]. House

[Ano14s, Ano14-29, Ano15c, Ano15t, Ano15u, Ano15y, Ano15-35, Ano15-38, Ano16d, Ano16g, Ano16r, Ano16s, Ano16t, Ano16u, Ano16v, Ano16y, Ano16x, Ano16z, Ano16-27, Ano16-28, Ano16-29, Ano16-36, Ano16-38, Ano16-42, Ano16-43, Ano16-39, Ano17a, Ano17c, Ano17p, Ano17-27]. HPC [Ano18p, Ano18q]. Hubble [LK95b]. Hulls [RSR83b]. Human [AAAW22, AGL87, Atw84, Bad82, Bad23, Bee15, BRD17, Bol94, BBC^{+}87, BvDE^{+}99, CCP82, CBD^{+}93, CNT^{+}97, CMSEM08, DBV^{+}08, DLEF15, EAM22, EJ98, Enc16a, Fet82, FWC84, Gre84, HS04, HWY16, HE82, IK15, Jer18a, KJA^{+}18, hKKSK17, KB96, Koc94, KW10, LBS^{+}03, Liv05, LK08, MGA^{+}21, MSU^{+}21, MQM22, MKL97, Myn08, NZB00, Paq96, PB10, SWDR16, SRJ99, Sin16f, Sin21b, Smi84a, Sun21, SD83, TK03, Tal97, VCR98, WC11, Wil82, YNH04, YCKF15, aCJB^{+}12, vOW15, KDH95b, MB91, Sim96b, TSH96, Whi95d]. Human-Centered [AAAW22, BvDE^{+}99, Jer18a, Tal97, YCKF15]. Human-Character [WC11]. Human-Computer [Koc94]. Human-Robot [DLEF15]. Human-Virtual [BRD17]. Humanities [BSS20, BEAC^{+}18, CDLS19, CN03]. Humans [BES17, CGMB20, CEMTT98, EMTTT98, EBT98, KMTM^{+}98, Lok06, LCR^{+}14, Tac16, YBL^{+}12]. Humpback [WAPW06]. Humphrey [Ano22-81, Ano23-84, Ano23-85, Ano23-86]. Hungry [Sin19a]. Hurrah [SvNG89]. Hurricane [ZCS^{+}06]. HuSIS [SWDR16]. HutWindows [KM88]. Hybrid [KSD04, KRG^{+}22, Lev90, ME22, RFK^{+}13, RLG^{+}02, SDM22]. Hybrid-Reality [RFK^{+}13]. Hybridized [DOH^{+}20]. Hydrothermal [LSB21]. Hyperbolic [Bli04a, Mun98, OC14, Bli95e]. Hyperfun [CAP^{+}05]. Hypermedia [And98, CLC^{+}96, Cra91]. Hyperscore [FPJ04]. Hyperspectral [ZPG^{+}21]. hyperstream [DH93].

I-WAY [DB96]. I/O [KHP^{+}11]. i3Drive [AP10]. i860 [GKB89]. Ice [ASP^{+}21]. Iconic [Lod83, Mar84]. Icons [YL11, Hag94]. ICU [RRF^{+}23]. Identify [Han00b]. Identifying [CCB22, MGO^{+}16, NAB^{+}23, SSZP15, TCH^{+}96]. Identity [Mar84, REF^{+}06, SCOSL13]. Identity-Differentiating [REF^{+}06]. IDGraphs [RGL^{+}06]. IEEE [Ano20r, Ano22a, Ano23a, DPL^{+}11, GGS^{+}18, LZF^{+}14, OKET12, SKS^{+}13, Ano98f, Ano99, Ano00, Ano03a, Ano04a, Ano13b, Ano14n, Ano14o, Ano14j, Ano14m, Ano14k, Ano14l, Ano14s, Ano14q, Ano14p, Ano14r, Ano14t, Ano15n, Ano15o, Ano15p, Ano15v, Ano15t, Ano15q, Ano15u, Ano15r, Ano15s, Ano15w, Ano15y, Ano15x, Ano16r, Ano16s, Ano16t, Ano16u, Ano16n, Ano16o, Ano16p, Ano16q, Ano16v, Ano16y, Ano16x, Ano16z, Ano16w, Ano16-27, Ano16-28, Ano16-29, Ano17p, Ano17j, Ano17k, Ano17l, Ano17m, Ano17n, Ano17o, Ano17q, Ano18s, Ano18u, Ano18t, Ano18x, Ano18v, Ano18w, Ano18r, Ano18y, Ano19a, Ano19-30, Ano19-31, Ano19-32, Ano19-35, Ano19-36, Ano19-34, Ano19-33, Ano19-37, Ano19-38, Ano19-39, Ano19-40, Ano20r, Ano20-47, Ano20-48, Ano20-43]. IEEE [Ano20-49, Ano20-51, Ano20-44, Ano20-38, Ano20-45, Ano20-41, Ano20-39, Ano20-50, Ano20-42, Ano20-46, Ano20-40, Ano20-52, Ano20-53, Ano21-55, Ano21-43, Ano21-66, Ano21-59, Ano21-44, Ano21-45, Ano21-56, Ano21-46, Ano21-47, Ano21-48, Ano21-49, Ano21-50, Ano21-67, Ano21-51, Ano21-52, Ano21-68, Ano21-63, Ano21-60, Ano21-64, Ano21-57, Ano21-53, Ano21-61, Ano21-58, Ano21-54, Ano21-62, Ano21-65, Ano22-47, Ano22-55, Ano22-51, Ano22-43, Ano22-60, Ano22-44, Ano22-61, Ano22-52, Ano22-48, Ano22-45, Ano22-62, Ano22-53, Ano22-49, Ano22-46, Ano22-63, Ano22-56, Ano22-54, Ano22-58, Ano22-64, Ano22-57, Ano22-50, Ano22-59, Ano22-65, Ano23-28, Ano23-33, Ano23-34, Ano23-29, Ano23-30, Ano23-31, Ano23-32, Ano23-50, Ano23-63, Ano23-64, Ano23-56, Ano23-62, Ano23-51, Ano23-57, Ano23-41, Ano23-65, Ano23-52, Ano23-58, Ano23-42, Ano23-46, Ano23-59, Ano23-47]. IEEE [Ano23-53, Ano23-43, Ano23-48, Ano23-54, Ano23-60, Ano23-44, Ano23-49, Ano23-55, Ano23-61, Ano23-45, Ano23-66, Ano23-67, Ano24h, Ano24i, Ano24p, Ano24t, Ano24r, Ano24n, Ano24q, Ano24s, Ano24u, Ano24o, BHS^{+}24, FGCB23, FS23, Hay96b, Hay96f, LRS09, PBSB08, SMG^{+}21, SLC^{+}23, Sto07, vWNS10, Ano18a, Ano18b, Ano19o, Ano19g, Ano19h, Ano19i, Ano19j, Ano19k, Ano19l, Ano20j, Ano20k, Ano20l, Ano20q, Ano20m, Ano20n, Ano20o, Ano20p, Ano20-32, Ano20-30, Ano20-31, Ano20f, Ano20g, Ano20h, Ano21y, Ano21z, Ano21-27, Ano21-28, Ano21-29, Ano21-39, Ano21-40, Ano21-41, Ano21-42, Ano21i, Ano21j, Ano21k, Ano21l, Ano21m, Ano21n, Ano21o, Ano21p, Ano21q, Ano21t, Ano21r, Ano21s, Ano22b, Ano22z, Ano22-27, Ano22-28, Ano22-37, Ano22-38, Ano22-39, Ano22-40, Ano22-41, Ano22-42]. IEEE [Ano22k, Ano22l, Ano22m, Ano22n, Ano22o, Ano22p, Ano22q, Ano22r, Ano22s, Ano22t, Ano22u, Ano22v, Ano23u, Ano23x, Ano23y, Ano23z, Ano23-27, Ano23j, Ano23k, Ano23l, Ano23m, Ano23n, Ano23o, Ano23p, Ano24m, Ano24d, Ano24e]. If [Cai20, Fre84, MSDS23, Sin02f]. iFeel IM [TN10]. IGES [GG87, Ken81, LK82, Smi83]. II [DPG18, HHH^{+}21, Mus90a, RLIB99, SW88]. III [Kim84]. Ill [Mac90]. Ill-Conditioned [Mac90]. Illuminance [GUFG15]. Illuminating [DOH^{+}20, HH92b, Sin17b]. Illumination [ANS15, MYK^{+}08, RDRS10, SMP^{+}22, WJP22, Whi20, ZCBM14, ZGP^{+}13]. Illusion [SHY^{+}23]. Illusions [SB13]. Illusory [Abo00a]. Illustrating [LWY^{+}20]. Illustration [BSP97, GAB^{+}18, SES05]. Illustrative [HBP^{+}07]. Illustrators [Kon12]. Image [AK84, AL11, AJK15, BWR24, BFK^{+}84, BNP99, CWWW21, CCT^{+}20, CGIB86, CRD^{+}18, CXPG19, Deb02, DG99, DNBC16, ESS82, ERH^{+}11, Eng97, FK87, FS90, dOGDM^{+}19, GLC11, GPW^{+}19, Gla03e, HG83, HNC^{+}04, IKMF12, IWF97, JLW08, VCR00, JF99, KB92, KKGH90, LKS96, Ler82, LW13, LPZ12, LWZY19, LLL^{+}13, LZS04, MTT87a, MKL97, Nea89a, OSRW97, PP95, Rei23, Rie93, RM95, Sch81, Sch21, SMWS88, SKP09, TSLX12, TLC02, TA85, TT00, UCK^{+}04, VT24, VP99, WK85, WKHL96, XJW18, XJ94, XPH05, Yan01, YY16, YBL^{+}12, YMS01, Zha96b, Bro88, GFB87, Har96, Kor87, LH92, MNR^{+}91, WS92, DWL^{+}92]. Image-Based [CXPG19, Deb02, HNC^{+}04, IWF97, LZS04, TLC02, UCK^{+}04, DG99]. Image-Generation [Eng97]. Image-Space [SKP09]. Imaged [DM02a]. Imaged-Based [DM02a]. ImageHive [TSLX12]. Imagery [CTM18, KB88, WFR^{+}00, Yan85]. Images [AFYC03, CFF01, DNBC16, FI83, GJM18, GH86, GZM97, HU83, HM85, Lea96a, Lea96d, LWS98, LS84, MIA^{+}08, MTTF85, MKFN87, OSJ11, PCJ^{+}16, QNT02, RAPL98, RGW05, RAGS01, Rob87, SW83, SLNG07, SPA^{+}16a, SN86, SKN^{+}05, Tit07, TR93, WGZ^{+}18, WN82, YP08, ZJH^{+}18, Zhu19, ZPG^{+}21, FG88a, Ger91, Kor87, NINT89, SR93, WFG92]. Imagination [Coh00, SH03]. Imaging [BKS19, BEG92, CDBL83, Del98b, Dev99b, Dev00, FZY84, FYZ85, Gel07b, GBT^{+}02, Hei18, LSB21, Mas18, MMP^{+}07, PBM^{+}96, RTF^{+}05, RDH^{+}02, SK03, SC21, SDT^{+}93, UBS^{+}12, WH99, BHK91, CG94, Ker91, NF91, Dev99a]. Imagining [FBML15].

Immersion [JLC98]. Immersive

[Ano18c, BH16, Bol94, BMC04, BRL06, CMT^{+}10, Cam16, FFN19, FPH^{+}06, HCL^{+}20, HW18, KCB^{+}23, KND^{+}18, KKF^{+}21, KH14, LBDM19, MR98, MLJ19, MTN18, MG04, NdSM^{+}16, NPB^{+}18, PPKM15, PBSB08, SSM^{+}07, SKMY09, VM07, WHKM^{+}13, YMK^{+}12, ZWWS19, ZED^{+}09, vDFL^{+}00]. Impact [CLP18, GSFT16, Hua21, LHFL20, LRZ23, LSC^{+}17, SF20, Tho00, Wei84]. Impacting [MSU^{+}21]. Impacts [LSVB21]. Impaired [GJM18, PSE^{+}09]. Imperfect [Ger98]. Implementation [ABS^{+}05, PP12]. Implementation-Oriented [PP12]. Implemented [Mam89]. Implementing [BH81]. Implementor [AEB86]. Implicit [Hof93, KS01, NB93, SB03b]. Implicitly [Cha88, HW90]. Improve [BWG24, Kee10]. Improved [HBB20, HB97, KM88, Koz10, PB85, RR84, TH14, XZZL19]. Improvement [JDZ04]. Improving [AMJ^{+}08, AESF21, CNC^{+}05, KDMH17, Pal81, SKW^{+}12, Sin18c, SOM^{+}10]. Improvising [Sin03b, Sin16b]. IMU [HROIK16]. IMU-Based [HROIK16]. In-Class [WD15b, WD15c]. Inbuilt [DL15]. Inc. [Ano19b]. inch [MD99e]. Including [MD96c, SR16]. Inclusion [Gai17]. Incoming [Möl22]. Incompatibility [Lea96d]. Incompatible [Sch86]. Incorporating [RME14, SG11]. Increase [DSA^{+}19]. Increased [LLG^{+}07]. Increasing [Ahr15, AMDM22, Kla93]. Incremental [Bre87, FDK12, Gil94, GI94, Mil93]. Incrementally [SBC15]. Independent [ANS15, WHR19, War81b, War81a, Pap96]. Index [Ano95a, Ano95b, Ano95c, Ano95d, Ano96a, Ano96b, Ano96c, Ano96d, Ano96e, Ano96f, Ano96g, Ano97a, Ano97b, Ano97c, Ano97d, Ano97e, Ano97f, Ano98f, Ano98a, Ano98b, Ano98c, Ano98d, Ano98e, Ano99, Ano00, Ano01, Ano02, Ano03a, Ano04a, Ano05a, Ano06a, Ano08a, Ano10a, Ano11a, LHZ^{+}21, Bar81b]. Indexed [AP11, MTT85b, MTT87a]. Indicators [Ano14b, Ano14a]. Indie [GAe^{+}24]. Individual [ZOC^{+}12]. Individuals [Ano14b, PCG^{+}21, Ano14a]. Indoor [AFYC03, HB08, PCG^{+}21]. Industrial [CP89, GUFG15, KS07, Mye82, Nav04, PTB^{+}15, VSLG22, Goe96]. Industrie [PTB^{+}15, Sto15]. Industries [RBW05]. Industry [Agr21a, CSWB01, CSA22, KD20, Kri04, Lea97d, Mor98, Ped01, SBCvdS04, PZD^{+}18]. Inertia [Gün20]. Inertial [Fox05]. Infants [PP11]. Inference [HSK02, MZ22, Mil93]. Inferences [BWG24]. Inferring [Bli99c]. Infinite [Pie87b, Sin06c, Sin11d, Sin15d]. Infinity [Sin07b]. Inflection [Bli99b]. Influence [HBW^{+}14, MCGS21]. Info [Ano18g, Ano18t]. Information [Ano96i, Ano96k, Ano97h, Ano18u, Ano18v, Ano18w, Ano19-35, Ano19-36, Ano20r, Ano20-47, Ano20-48, Ano20-49, Ano22-58, Ano22-59, Ano23-56, Ano23-57, Ano23-58, Ano23-59, Ano23-60, Ano23-61, Ano24t, Ano24u, BFH02, BBDM98, BWHR99, Che05, Che08, CEH^{+}09, hCRBK98, CE98, CAG^{+}06, CM11, DY09, DR83, Eic05, Enc17a, FZH^{+}13, FJHW00, GB96a, GE97a, GE97b, HT04, IBH^{+}09, JS10, JBBL02, KFH^{+}18, Ker13, KRL97, Kro04, Kub84, MC85, Mac81, Mac83, Mac85, Mac00b, Mun02, PKM18, PS22, Rei17, Rhy03, RS97, SMP^{+}19, SLNG07, SvW09, SSOC01, ST09, SSMH03, TH01, WB00a, WB00b, YY16, Zha07, ZNMT15, BB95, BHMN95, Car96]. Information-Technology-Based [SSMH03]. Information-Theoretic [Che08]. Informational [RFS08]. Informing [SRF^{+}06]. InfoVis [Hay96d, KSK21, KS14, Lea97c]. Infrared [LSFW83, Sch99]. Infrastructure [GB96a, MSH^{+}08, TCB03]. Infused [MS24, MCS22]. Injuries [PP11]. Ink [Del00c, Lee99]. Inking [Ann17]. Inner [BDH^{+}89]. Innovating [KIQ^{+}06]. Innovation [Myn08, PRN15]. Innovations [SH06]. Innovative [Zuc84]. Inpainting [SLC24]. Input [FHKH06, Goo86, MD00f, Mor05, PBVH15, SZ94, TM88, WB85, WTPHK09]. Inscribe [MD99d]. Inselberg [Shn20]. Insertion [ADBW96]. Insight [AY16, CS15c, Cam16, CZGR09, KBHP14, Nor06, RLCI24]. Insights [Agr19, Ahr15, BFR21, CLms15, CCC^{+}07, DMSJ23, Her24, MML^{+}15, NLvLM09, SML^{+}09, WSL^{+}20, vL18]. Insoles [HLC11]. Inspection [BHG23, DCCE19, aKTRL^{+}21]. Inspector [CPCS08]. Inspiration [Arn14a, Arn14b]. Inspired [CS19, JW16]. Instances [MYT09]. Institute [Ear98]. institutional [RBMS17]. Instruction [APB^{+}21, CLC^{+}96, SRS83]. Instructor [CPAV^{+}17, Kon12]. Instrument [VM07, WHKM^{+}13]. Instrumentation [PBM^{+}96]. Instrumented [SRDT01]. Instruments [ASP^{+}86]. Integral [lLK84]. Integrate [KS14]. Integrated [Bee82, EHBPG04, GDS^{+}18, KYK^{+}24, RRR^{+}16, YKR^{+}84, Kal87]. Integrating [Ber99, BHMN95, CHC^{+}97, FJHW00, KK02, Kee10, PMI06, PS09, PFH95, RSS^{+}09, SDS^{+}11]. Integration [Ano20r, Saw08, Smi83, LJR96]. Intel [GKB89, PWW97, Shi86, Wor96b]. Intellectual [LIS^{+}19, Whi95d]. Intelligence [Ano23-33, BHPS12, Bas14, CHH24, Die85, FDC14, LWM14, LSVB21, MS24, RMDT96, XSN^{+}20, Ano23a]. Intelligent [Ano22b, Ano22p, HGK^{+}04, Shu23, Sin22d, Spi86]. Inter [RBMS17]. Inter-institutional [RBMS17]. Interacting [Bai01, EBT98, Enc15b, FBML15, HdlRL^{+}13, SB16, SC06, Str04, WLVL^{+}96, WB14]. Interaction [BVRF05, BRD17, BPB14, BJC18, Cat96, CMO^{+}99, DSP^{+}14, DYS^{+}07, EAM22, End14, ECNZ15, FWC84, FLMZ97, FPW^{+}00, HHLL14, IZ07, II15, Jer18b, JLS15, KMS07, Kee10, KBHT06, KGB14, Kul09, LaV08, LGZ04, LSR^{+}00, LFOI15, MMR^{+}06, Mue18, Mye87, PMI06, SB06, SE19, SS97, SHS^{+}21, WHKM^{+}13, Alp93, WC91]. Interactions [BCMF16, BKt82, CLDW16, DJS^{+}09, KW04, LKPM19, SWDR16, SGE^{+}18, BKtH82]. Interactive [ADBR^{+}12, ABVA08, AP10, AOH^{+}14, AZSW22, BMC84, BBF^{+}98, Bar95, Bar97a, BQP^{+}19, BABB^{+}21, BKM16, BSM^{+}15, BH16, BF93, BAAQ21, BBC^{+}20, BG82, BTTU84, BBT^{+}13, BHP07, CW96, Cro87a, DCCE19, DBV89, DPL^{+}11, DG99, DAG95, ERH^{+}11, Ell94b, ECNZ15, Eng86, FEV^{+}01, FJHW00, FAB^{+}98, Fre86, GN83, GAB^{+}18, Gel06, GSW24, Gir87, Gla02e, Gla02d, GB16, GS88, GHQ04b, HF18, Hai06, Hay96e, HCV^{+}13, HB16, kHHBEH13, HBP^{+}07, Hua11, IM14, Iga19, IIH^{+}13, JWL^{+}18, JBH^{+}10, JXW^{+}08, KIN^{+}05, KK02, Kan88, KFH^{+}18, KWK21, Kas11c, KD20, Ken20, KMM^{+}01, KKM96, KY97, LD99, LV23, LSV^{+}02, MVWB05, MSA^{+}94, ME22, MSK04, MNC^{+}22, Mor05, MMD^{+}19, NOK^{+}08, Nea89a, NMRI06, ONDA20, PLM^{+}19, PM82, Pie87a, PS22, RR95, RLOW13, RGA^{+}83, RNNTD14, RPHW11, REF^{+}06, SKR18, SPMK14]. Interactive [SWvdWvW12, SS01, SBD^{+}17, SS19, SWS07, SKP09, SRL06, SBM86, Ski81, SMMS24, SH03, SK86, SJSK19, SBS10, SGEF^{+}12, Tak05, TSLX12, TH17, TWQ^{+}23, UTMI11, UBS^{+}12, UZM^{+}23, UCK^{+}04, WB85, WC11, WS90b, WL88, WHF01, WW24, XY01, YIK85, YSWW08, YLL10, ZPG^{+}21, dRvWW10, vL06, Ano98g, BB13, Bri90, Bro88, KmKK^{+}18, Nie86, DMSJ23]. Interactively [BHPB^{+}12, CEO^{+}93]. Interactivity [SO95]. Intercomparison [LSC^{+}17]. Interdisciplinarity [LHM23]. Interdisciplinary [RBMS17, RSZ96]. Interest [RGRLFM21]. Interface [Ano18f, Ano18z, Bli05b, BSHB01, EJ98, FKC^{+}05, GPA^{+}17, Gre86a, HZHD03, KM88, KGB14, KRG^{+}22, KPL17, KDMH17, KL09, MD99d, MTT85a, Mar82, Mar84, MCP^{+}06, Mye84a, OBE^{+}84, Ovi03, PFA86, Pra84a, RD90, RC85, RSE99, RZSP04, SZP^{+}00, SBM86, SRS83, SKN^{+}05, TK03, Tak85, TM88, TWB^{+}07, WP87, Wil84, Wil86, KmKK^{+}18, Ols86, Pea88, Rob88, Ano19s, Ano19t]. Interfaces [BH16, BPB14, BCF^{+}08, BSHB01, CG85, CW09, EH03, EI17, FB09, FL12, GSSO05, HdlRL^{+}13, Kas13, LJ15, LGM13, LS04b, Ma00, MSWT14, PARV05, RSP15, SKJ^{+}13, Shn03, SWD^{+}98, SS20, SS15, WFOP85, WWV^{+}10, EFB^{+}94, Mye88, WC91]. Interfacing [Lod83]. Interference [Dia91, SFCD99, Dia94]. Interiors [MG14]. Internal [CBS^{+}22]. International [Ano20r, CPC99, Jar83, BHK91]. Internet [Ano19h, Bai95, Hoe12, MD99e, MD99d, PTB^{+}15]. Interobserver [BWR24]. interpersonal [Phi91]. Interpolation [CW88, DMR83, FM89a, Gar15, GI94, LMD92, McL83, Pie87a, PT00, RCB98, RM95, WH97, WMG00, CW87, FM87, HZB92, Rou87, WF93]. Interpolatory [FM90]. Interpretation [DMSJ23, MD00f, SD07]. Interpreting [MUN^{+}10, VVP^{+}11]. Interrogation [HHS^{+}92]. Intersection [HAG83, LF84, LHM23, PO84, ABB90, DS87, Sug92, Wan93, YSSP91]. Intersections [SB03b, Pat93]. Intersystem [LK82]. Interval [MK84, SF92]. Intervals [WDK17]. Intervention [BNA^{+}22, CZY^{+}13]. Interventions [RNNTD14]. Interview [CF16, CS17a, Nea89b, OBG^{+}23, SCN90]. Interviews3D [BHP07]. Interweaving [KJA^{+}18]. Interwoven [WW07]. Intio [WNAJ92]. Intracardiac [CZY^{+}13]. Intracranial [FZY84]. Intraoperative [SKS^{+}96b]. Intricate [DHC^{+}22, Tit07]. Introduced [MD99f]. Introduces [MD99e].

Introducing [RT07, Sto07, Tau13a, Tau13b, Zhu12]. Introduction [AD05, AS10, AS11, AR99, BO09, BC05, BES09, BSS20, BDM02, Bre96a, BEG92, Bre97, BNP99, CR02a, CH10, CCR10, Coo02a, CES07, DM02a, Dev99a, Dil98a, DS07b, Ear97, EMTTT98, Ear01, EFD05, ED93, EH03, Eng97, FT00a, FSK07, FM03, FL00, GE97a, GE98, GW98, GHR01, GH97, GP93b, HJ03, HJS00, IZ07, JB02, KM00, KMS07, KNR93, KF05, LK03, LS04b, LOB08, MP99, Ma01, Ma06, ML05, MT94, MBB95, MEW98, Mun02, Myk88, Nie92, NVC92, Nie93a, PP98, Par98, PB10, PGS09, PP12, Rho93, RR96, Rho97b, RMD06, RR03, RA93, RS99, RBF95, RSZ96, RBT98, Rus11, SB06, SC06, Sto99, SP04b, SBE95, Tau12b, TF01, TR94, WB00a, Won99, vHV09, vL06, vWNS10, BS16, FDC14, GP91a, LR14, Sei93, Sta91b]. introduction [Sto15, TC15, YCKF15]. Introductory [DiS14, Hai06, SCB04a]. Intruders [TMWJK04]. Intrusion [EWF02, ITSK06, RGL^{+}06, SZZ^{+}18]. INTUIT [DMSJ23]. Intuitive [WC11]. Invariance [ANS15]. Invariant [CT09, SSC^{+}11, SMP^{+}22]. Invariants [Bli02a]. Invasive [BDK^{+}04, BSHW07]. Invent [Ano17c, Ano17b]. Inverse [KB96]. Investigating [GDF^{+}14, MQM22]. Investigation [BBC^{+}87, SMP^{+}19, THB^{+}90]. Invisibility [Nie07, Bli88b]. Invisible [DOH^{+}20, Sin19c]. Inviting [VCS19]. Involving [DSP^{+}14, FDK12]. Ion [RLV^{+}16]. Iowa [CNL00]. IPC [Li22a]. IRI [Ano20r]. Iris [LBS^{+}03, Nic84]. Iron [BvRS^{+}11]. IrPen [HHLL14]. Irradiance [DDdSC11, GSHG98]. Irregular [Gie92, MW82, SGC21]. irregularly [IP92]. ISDN [JOnH93]. Islands [SMP^{+}19]. ISMAR [FGCB23, SR10]. Isn’t [KHHP06]. Isogeometric [AESF21]. ISOS [BDH^{+}89]. isosurfaces [Wal91]. Issue [Agr09, Ano19-54, Ano20s, Ano22-29, BHH^{+}21, BCPS22, Dil98a, Enc17b, GGS22, LQ18, Rus11, Tur02b, Whi97]. Issues [Fis02, Hal87, Mil86, Mil89, Mur84b, Pra84b, Ros94, Smi84a, Sto01, Whi95d, EFB^{+}94, HPvW94, KHK^{+}94, KS94, NBG^{+}94, RET^{+}94, RN94]. Italic [Hua11]. Italy [MFG^{+}17]. Iterated [Van10, Har96]. Iterative [AAG^{+}22, LCW12, SL00]. Itself [Per16].

J2ME [Kri03]. Jabiru [SQG^{+}05]. Jag [FI83]. Jag-Free [FI83]. Jaggy [Bli89e]. Jam [Pac04]. Jantzen [CS17a]. Japan [Kri01, Nea86]. Japanese [Kah94]. Java [MD99f, MD00f, SD99, SQG^{+}05]. Java3D [HJF06]. Jazz [Sin24a]. Jerry [Sin12d]. Jet [SK05]. JHAVÉ [Nap05]. Jim [Bli92e, Bli92f, Ano10b, Bli87, Bli88d, Bli88e, Bli88c, Bli88b, Bli89c, Bli89e, Bli89d, Bli89b, Bli89a, Bli90b, Bli90a, Bli90c, Bli91b, Bli91a, Bli91c, Bli91d, Bli92c, Bli92a, Bli92b, Bli92d, Bli93c, Bli93b, Bli93d, Bli93a, Bli94b, Bli94c, Bli94d, Bli94a, Bli95a, Bli95b, Bli95d, Bli95e, Bli95c, Bli96a, Bli96c, Bli96b, BG96, Bli96d, Bli97b, Bli97a, Bli98a, Bli98f, Bli98e, Bli98c, Bli98d, Bli98b, Bli99c, Bli99b, Bli99a, Bli00a, Bli00c, Bli00b, Bli00d, Bli00e, Bli01a, Bli01b, Bli02a, Bli02b, Bli03e, Bli03a, Bli03b, Bli03c, Bli03d, Bli04a, Bli04b, Bli04c, Bli05b, Bli05d, Bli05a, Bli05c, Bli06c, Bli06d, Bli06a, Bli06b, Bli07a, Bli07b]. Jing [CZP^{+}10]. JNZNBRK [JBCS23]. Job [AAG^{+}22, Ano14c, Ano15d, Ano15e, Ano16g, Ano18-29, Ano19-45]. Jobs [Ano13c, Ano18-27, Ano20-50, Ano21-63, Ano21-64, Ano21-65, Ano22-60, Ano22-61, Ano22-62, Ano22-63, Ano22-64, Ano22-65, Ano23-62]. John [Kas17]. Joins [Hay95b]. Joint [ADZK^{+}21, ASS^{+}21, BBT^{+}13]. Joints [VCR98]. Jon [MS24]. Journal [Ano20-32, Ano20-30, Ano20-31, Ano21-39, Ano21-40, Ano21-41, Ano21-42, Ano22-37, Ano22-39, Ano22-40, Ano22-41, Ano22-42, Ano24m, Ano20f, Ano22-38]. Journalism [KND^{+}18]. Journalists [Rei17]. Journals [Ano14b, Ano14a]. Journeymen [FY97]. Joy [Sin13d]. JPEG [XPH05]. JPL [Hay95b]. Jr [Whi23]. Judge [Hay95b]. Judgment [DPL23]. Juggling [Fig01a]. Julian [LTS^{+}13]. July [HBR^{+}92b]. June [Ano20u]. Jupiter [Ano94d]. Just [vL16]. JustEdit [MD99f].

Kaleidoscope [OWZ17, Bli88c]. KDE [FH21]. Keep [Ano19-43, Ano19-41, Ano19-42, Ano20-54, Nea89c]. Keeping [Whi95a]. Ken [Ano17-31]. Kendall [LZR^{+}21]. Kennedy [Ano17-31]. Kernel [PM86]. Key [IK15, PTB^{+}15, Smi83, TR83, YNH04]. Key-Node-Separated [IK15]. Keyframing [Wea98b]. Keypoint [SLSP22]. Keypoint-Based [SLSP22]. Keyworded [NVC92]. keywords [Bar81b]. Kid [Del98b]. kids [Sim94d]. Killer [Nav04]. Killing [Sin20a]. Kinaesthetic [HMC^{+}20]. Kind [Kos15]. Kinematic [CCP82, Dai89, Til83b, MB91]. Kinetic [WZL^{+}21]. Kineticist [WZL^{+}21]. Kingdom [Del98a]. Kit [Smi87]. Knots [Gla99b, Gla00a]. Knotwork [Gla99c]. Know [Gla98c, Jer18a]. Knowledge [Bec97, CEH^{+}09, CH10, CPG^{+}16, HZHD03, KTD10, KW10, MGP^{+}04, MSW10, dSNRdLdSN16, RC85, SI94, SADH23, TWQ^{+}23, Ano18-40]. Knowledge-Assisted [CH10, CPG^{+}16, KTD10, KW10]. Knowledge-Based [RC85, SI94]. Knowledge-Driven [SADH23]. Koch [PS88]. Kunii [Woz21].

L [Woz21]. L. [Use93]. Lab [CNL00, GK05, LPMAA^{+}22]. Labeled [SSTP18]. Labeling [LFT19, ZSK^{+}21]. Laboratory [SSMH03, Wei23, MLJ19]. Labs [TCE^{+}02, Hay96c]. Lag [TCH^{+}96]. Lamp [WAB^{+}13]. Lance [CS19]. Land [AZSW22]. Landfill [GDF^{+}14]. Landing [MPDR14]. Landscape [Erv93, FBML15, MMR^{+}06, QNT02, SC21, War00]. Landscapes [BCMF16, LFFB20, RGCB07, Sin16c, YZM^{+}11]. Language [CW96, FSB82b, Per16, SG86b, SS20, YIK85, FSB82a, GS88, Pal92, Sto99]. Language-Based [CW96]. Languages [BBK82, SR16]. Laparoscopic [LGK^{+}06]. Laplacian [KSSL07]. Large [ABM^{+}01, Ahr22, AFYC03, AJK15, AGP^{+}22, Bal98a, BWA96, BE11, BB05, BSA^{+}04, BFBK00, CRF^{+}21, DSP^{+}14, DSW07, EHBPG04, EWF02, FS01, FKKB05, FL00, Fyo97, HYL^{+}22, HYN03, HMRS06, HLB^{+}07, JBH^{+}10, JXW^{+}08, Kan09, KPSN04, KHP12, KMM^{+}01, KNK99, KÇC^{+}01, KF05, Ma01, MP01, MSWT14, Mor05, MRBW12, Mun98, PSJ87, RL07, RFK^{+}13, RS14, RCB^{+}05, RW99, SBMK20, SB16, Sen99, SMMS24, SBSW96, VBHS11, VK17, WAB^{+}05, WYM10, War04, WSK^{+}00, WK85, WTR11, YLL10, YZM^{+}11, YWG^{+}10, de 86, KmKK^{+}18, KKM96]. Large-area [de 86]. Large-Crowd [JXW^{+}08]. Large-Display [HMRS06, RCB^{+}05]. Large-Format [BE11, FL00]. Large-Image [WK85]. Large-Scale [ABM^{+}01, Ahr22, AJK15, AGP^{+}22, Bal98a, DSW07, EHBPG04, EWF02, HYL^{+}22, HYN03, HLB^{+}07, JBH^{+}10, Ma01, MP01, MSWT14, MRBW12, RL07, RW99, SBMK20, SBSW96, VBHS11, WAB^{+}05, WYM10, War04, YWG^{+}10]. largest [Bli88a]. Larry [Tau12a]. Laser [Lea96a, MD99f, SPM10]. Lasers [RB93]. Last [AC05, SvNG89]. Latency [Fri20]. Lateral [BHS^{+}24]. Latin [Hay95b]. Launched [MD99e]. Launches [MD99e]. Laurenzo [LCS21]. Law [WW95]. Laws [Gol01]. lay [Sch93b]. Layer [AAB^{+}23, KHP^{+}11]. Layered [Boh95, CMP95, LLY23, OKS^{+}17, YSWW08]. LayerNet [LLY23]. Layers [Fig01a, Sin15d]. Layout [HKDL22]. Layouts [ABVA08, Cam17, HWP^{+}19, IK15]. Lazy [BH84]. Lead [MHTD23, She06]. Leader [PB06]. Leading [Ano15-44]. Learn [Per21]. Learned [BAB^{+}23a, FT16, WCG^{+}11]. Learning [Agr18a, AAAW22, Ano19-54, BRD^{+}20, BKM16, BCPS22, BBC^{+}20, BD21, CPM^{+}18, CL18, DL22, Dom12, DBJEY^{+}02, EH16, GPS^{+}21, GHCW21, HWP^{+}19, HTZ^{+}19, HPX^{+}21, Hay96e, HKDL22, JMOG99, JRM17, KJR^{+}18, KSK21, LFT19, LA20a, Ma07, Ma23, MSW10, MB05, MM21, MMD^{+}19, dSNRdLdSN16, PPVH23, Pea91, QCP^{+}10, RdCFP^{+}20, Rou09, SBRCD17, SHS^{+}21, SW22, TWQ^{+}23, WYHS21, Wan22, WFZB21, Zhu19]. Learning-Based [PPVH23]. Leave [OCS20]. Left [Fol00, Sin12b, Wri83]. Legacy [SK13]. Legal [Lav83]. Legged [Gir87]. Legibility [GUFG15]. Legible [CBZ^{+}08]. Length [GP90]. Lens [BSP97]. Lenses [BH06, YCB05]. Lesion [GPS^{+}21]. Less [Pie87c]. Lessons [BAB^{+}23a, Bas14, FT16, MLJ19, May83, WCG^{+}11]. Letter [Lev00, WMG00]. Letters [Ano93t, Ano93r, Ano93s, Ano18-28, Ano19-44, Ano18y, Ano19-38, Ano19g, Ano22n]. Level [MZC^{+}21, MD99e, MTT83, SSTP18, SLSP22, TO83, Whi04, YCB^{+}22]. LevelMerge [SSTP18]. Leveraging [EE12, Sun21]. Lexcube [SMMS24]. Liberator [Ano94g]. Library [Ano18-50, FPNVHH17, Ano14-39, Ano15-43]. Licensing [Whi95d]. Lidar [Rot03, MBV17]. Life [CLP^{+}14, How00, Jac15, KM00, MD99e, NPH^{+}09, PTR^{+}98, Sin03a, Sin03b, Sin12e, Sin14e]. Light [BDC^{+}12, CWWW21, Cha98a, DKS^{+}03, Del00a, HG86, JBCS23, PWH04, RDRS10, Sin02c, SFCD99, VB24, VG84, WCS18, WLH^{+}12, Wor93]. Light-Source [VG84]. Lighting [BPB09, DM02a, Deb02, DAG95, IKSZ03, Pic92, SMDBR22, SSH^{+}98, Str90, XNSW19]. Lightning [Gla00e, Gla00f, PS22].

LightPainter [CTM18]. Lightweight [RLCI24, TPA24]. Lightyear [MD99e]. Like [CPN09, GAe^{+}24, MKH20]. Lillian [Sin12f]. Limb [CKPB16, VCBS97]. Limit [DB16, Gla05b]. Limited [VK17]. Limits [BU23]. Line [Bli91b, Bli03a, Bli04c, BB00, Bre87, CWB02, COK97, DW87, Fre86, Gil94, HS01, Lea96a, MD00f, Max90, Pan90, Rob89, RR85, SL00, Til81, Wri90, PB87]. Line-Clipping [RR85]. Line-Drawing [Pan90]. Line/polygon [Til81]. Linear [GWW86, Gol03, GI94, Pic92, RY06, SSB06, UF11, ZZ05]. Lines [Bli03e, Bli03a, Bli03b, Bli03c, Bli03d, Bli04a, Bli04b, Bli04c, HE82, PSX13, SR85, Sin18b, SP04a, DH93, Sim93c]. Lineups [DBV^{+}08]. Lingua [SR16]. Linking [LLR09]. Linux [MD00f]. Lip [WD15a]. Lip-Sync [WD15a]. Liquid [FH84b, SKN^{+}15]. Lisa [BGM^{+}07]. Literacy [BHH^{+}21, CCB22, CRA^{+}18, FJL22]. Literalism [Smi87]. Literally [KMH02]. Literary [FBML15]. Literature [PD21]. Little [BvRS^{+}11]. Live [ABWS05, BHM^{+}22, BJC18, CBD10, SG11, Sin23e, WD15a, ZSK^{+}21]. Live-Captured [ZSK^{+}21]. Liver [KLR12, RBBS06]. LiverAnatomyExplorer [BMS^{+}13]. Living [BRM10, ZED^{+}09, aCJB^{+}12]. Load [Whi94b]. Loading [LJB^{+}13]. Local [CXW^{+}20, KLR12, YCB^{+}22, Sim93a]. Locales [BWA96]. Localization [GJM18, SMP^{+}22]. Locally [GG92]. Locating [PO84]. Locomotion [Iwa99, hKKSK17, KB96]. Lofted [FB89]. Logarithmic [WO13]. Logic [GF86, PEMF92, PJJF96]. Logic-Enhanced [GF86, PEMF92]. Long [CTM18, PS09, SAB^{+}23, Sin23e]. Long-Exposure [CTM18]. Long-Term [PS09]. Longs [RV07]. Look [Ano14d, BDG^{+}99, HAG83, Kos22]. Look-Up [HAG83]. Looking [Ano18-29, Ano19-45, BFR21, Del98b, Hag90c, Kas12b, SF20, Bli88e]. Looks [Ano97i]. Lookup [FK87]. Loops [SB06]. Loose [Lai01]. Loosely [Sin24b]. Lorensen [Ano20-62]. Lost [Sim96c]. Love [Ano18-39]. Low [CWWW21, Fri20, GJM18, JAC^{+}08, MD99f, MD99e, PCJ^{+}16, SGC21, WSKY09, YST84, ZYJ^{+}21]. Low-Altitude [PCJ^{+}16]. Low-Cost [PCJ^{+}16, SGC21, YST84]. Low-Dose-Rate [ZYJ^{+}21]. Low-Latency [Fri20]. Low-Light [CWWW21]. Low-Power [JAC^{+}08, WSKY09]. Low-Priced [MD99e]. Lower [VCBS97]. Lower-Limb [VCBS97]. lowly [Bli96a]. LumiConSense [KB14]. Luminance [TH14]. Lumisight [KIN^{+}05]. Lungs [KW10]. Lutz [LB20]. Lyapunov [Pic90].

m [MSU^{+}16]. M1C2 [Enc15c]. Machina [MM89, Mus90a]. Machine [Agr18a, AAAW22, BCPS22, BBC^{+}20, EH16, LF87, Ma07, Ma23, MSW10, SBRCD17, SHS^{+}21, Sin06d, Sin20b, Sin20f]. Machine-Learning-Driven [SBRCD17]. Machines [Pan90]. Machining [CM92, JDH^{+}89]. Machover [Mac97a, vD12]. Macondo [FBML15]. Macros [OD88]. macroscope [Bow93]. Macrostructured [DG99]. MacSpin [DDG88]. Mad [Sin08a]. Madden [LTS^{+}13]. Made [CSMT03, GS12]. Made-to-Measure [CSMT03]. Magazine [Ano22b, Ano22p]. MAGES [ZPL^{+}23]. Magic [BSP97, Bli00e, BH06, Hec86a, Kul09, MPP81, Smi87, YCB05, Del98a, DJM13]. MagicBook [BKP01]. Magnetic [Chr83]. Magnetohydrodynamic [FOI^{+}07]. Magnifiers [BEK^{+}03]. Mahalanobis [FH21]. mail [MD99d]. Maintain [Fyo97]. Maintainability [BSA^{+}04]. Maintaining [Hur82]. Maintenance [DFP^{+}11]. Major [Ano19m, Ano23q, Ano23r]. Make [Shn03, XDJ23, Bli95c]. Maker [SR13]. Makes [MD99f, MD99d]. Making [BB13, BJP23, BDG^{+}99, CCC18, DOH^{+}20, Fry83, Hec86a, Jae17, JKM23, KMCMT09, LB20, MR99, MHTD23, MM14, OBG^{+}23, Ozt15, PGV20, PKN^{+}22, RBMS17, RA93, Sin19c, SEAKC21, TR83, WS09a, WS09b, WLC^{+}12, WCS18, Whi96c, Yea97a, Den93]. Man [Ano20-62, Lil81]. Man-Model [Lil81]. Manage [Ano19-53, NAB^{+}23]. Management [AHDG00, BB05, CE98, Cro96, DB03, DR83, EZM^{+}97, GAe^{+}24, KM88, LAC^{+}14, LHM23, MD99e, MGW^{+}18, NMRI06, OBE^{+}84, Pal81, QCP^{+}10, Tak85, UMO82, VG81, WKY23]. Manager [Goo86, Mye88]. Manages [MD99e]. Managing [DR83, Ma00, SSW^{+}07, SKJ^{+}13]. ManEx [EKM^{+}22]. Manikins [MG14]. Manipulate [Zhu19]. Manipulating [GU86, RP95, GG92, SR93]. Manipulation [CW96, EBN13, FK87, HKSK17, KP14, LS84, LiA10, VP16, FB93, PT92, UO91]. Manipulator [Agb95]. Manned [Gla00g]. Mantle [ME22]. Manual [DLEF15]. Manufacture [Pra84a, Sim94c]. Manufacturing [Bai95, BHM^{+}22, Boh95, CMP95, CSA22, DHC^{+}22, GGS22, HLC11, LF87, Sto15, YKR^{+}84, RKRW87]. Manuscripts [Gla98d]. Many [Bli99b, Bli99a, Bli00b, Fig01a, OCS20, WSL^{+}20, Bli89d, Whi94a]. Map [BT07, BH11, BBDZ08, HB06, Lin99, Ros89b, Zac84]. Mapping [ABS^{+}05, Bow93, CLG^{+}16, FS86, FPBB08, Gel07b, Gre86b, Hec86b, Hua21, KJR^{+}18, Lea97a, ODW15, PT90, RO86, SB86a, SPA^{+}16a, WD15b, aE01, Bli90a]. Mappings [BS86, OC14]. Maps [BCA^{+}20, BBDZ08, CMT^{+}10, CG84, HGK10, LC16, Mam89, Nur08, SWvdWvW12, SAHK13, YM95, YZM^{+}11, Zhu12, War88]. MapShaper.org [HB06]. Marathon [YIÇ09]. Marbling [JCM07, LJJ^{+}12, LJJ^{+}17, XMJ08]. March [Hay96b]. Marching [Ano20-62, Lor20]. Marine [KP12]. Marker [ANS15]. Market [NB04, Pra84b, Wor96b, Koh93]. Marking [WMB85]. Markup [MD99d]. Márquez [FBML15]. Mars [SDPW20]. MarsCAPE [SDPW20]. Martin [Ebe17]. Maryland [PS22]. Mask [WEWL00]. Mass [HJCW06]. Mass-Spring [HJCW06]. Masses [ASP^{+}21, AB14b, LaV08, Sch07]. Massive [AM04, BHP07, DGY07, GCS^{+}12, RLIB99]. Massive-Model [DGY07]. Massively [MP01, MSU^{+}16]. Master [MGN^{+}11, Sin09a, Wor93]. Masthead [Ano13d, Ano14u, Ano14v, Ano14w, Ano14x, Ano14y, Ano14z, Ano15z, Ano15-27, Ano15-28, Ano15-30, Ano15-31, Ano15-32, Ano15-29, Ano16-30, Ano16-31, Ano16-32, Ano16-33, Ano16-34, Ano16-35, Ano17r, Ano17s, Ano17t, Ano17u, Ano17v, Ano17w, Ano18-34, Ano18-30, Ano18-31, Ano18-32, Ano18-33, Ano19-46, Ano19-47, Ano19-48, Ano19-49, Ano19-50, Ano19-51, Ano20-55, Ano20-56, Ano20-57, Ano20-58, Ano20-59, Ano20-60, Ano21-69, Ano21-70, Ano21-71, Ano21-72, Ano21-73, Ano22-66, Ano22-67, Ano22-68, Ano22-69, Ano22-70, Ano22-71, Ano23-68, Ano23-69, Ano23-70, Ano23-71, Ano23-72, Ano23-73, Ano24v, Ano24w]. Match [SJSK19]. Matches [Lea96a, SJB^{+}16]. Matching [CRC^{+}02, MD99f, TPR^{+}22]. Material [PR12]. Materials [CXW^{+}20, HG13, TNTM07, Wan10]. Math [OF14, OC14, ODW15, OZH15, OWZ17, WO13]. Mathematical [Coh83, LaV07, LJJ^{+}12, LJJ^{+}17]. Mathematics [Cha99, Rou09, Hag94]. Matrices [Bli02b, Bli03b, Fil84]. Matrix [Bli00e, FH84b, WMFM13, Bli96a]. Matt [Hag90b]. Matter [Kri00, KH14, Rhy03, SBD^{+}17]. Matters [Ano14-28, Ano14-29, Sto12, VB24, Ano14-27]. Matting [JAPF07]. Mature [BT00]. Maturing [GH97]. Mauchly [Ano19a]. Maximizing [SBC15]. Maya [Abo00c]. MCAE [GZS88]. McCormack [MS24]. Me [Bli88d]. Mean [DBDH05, PGA13, Bli92c]. Meaningful [NLvLM09, VCS19]. Means [HAG83]. Measure [CSMT03, SJSK19]. Measurement [aKTRL^{+}21]. Measurements [EKM^{+}22]. Measures [RFS08]. Measuring [KVM^{+}05, MIA^{+}08, Nor06, dSNRdLdSN16, WLL^{+}16]. Mechanical [BA85, Myk88, SPJ99]. Mechanicals [Sin01b]. Mechanism [BKRE88, TM88, Til83b]. Media [Ear01, FZH^{+}13, FE97, Fry83, dOGDM^{+}19, Gol12, Kri04, LWM14, MRD06, Ped01, PZD^{+}18, SH06, Swe01, Tur02b, Yan01, YY16]. Medial [KPN05]. Medial-Axis-Based [KPN05]. MediaTable [dRvWW10]. Mediated [VT24]. Mediating [MCP^{+}06]. Medical [BWR24, CFF01, CJS^{+}19, Dev99a, Dev99b, DCLK03, GSG^{+}21, GPR94, HU83, VCR00, Lea96d, LPZ12, MVWB05, MG14, NTS^{+}07, Sim96c, SES05, THB^{+}90, ZPL^{+}23, Ker91, MNR^{+}91, NF91]. Medical-Image [LPZ12]. Medicine [Del98b, Dev00, PF83, Rho90, Rho93, Rho97a, Rho91]. Medium [Sin10f, Sin18e, DH91]. Medusa [Sin22e]. Meeting [FKKB05, FE97, LRS09]. Meets [Ken20, Kos22]. MegaMol [KGR^{+}18]. Melding [BC19]. Members [Tau11d]. Membership [Ano13e, Ano14-27, Ano14-28, Ano14-29, Ano16-37, Ano16-38, Ano18-36, Ano18-35, Ano18-37, Ano18-38, LV23, Ano17z, Ano18-40]. Memoriam [Ebe17, Enc15a, Kas17]. Memories [PEMF92]. Memory [DDdSC11, DSKH95, GF86, KMCMT09, KB88, SKS96a, WK85, Whi84]. Menagerie [PT89]. Mental [Smi84b]. Menus [KKG^{+}13]. Merges [Mar07]. Merging [FZ03, SSTP18]. Mesh [GS12, GXMW15, KT96, LPZ12, XNSW19, YS83]. Mesh-Based [XNSW19]. Meshes [Gie92, KSK00]. Meshing [ABS^{+}05, LTG92]. Mesoscale [GAB^{+}18]. Message [Ano94j, Hay96f, Her96, Tho01, Tho02a, Tho02b]. Metacomputing [LKS96]. Metafile [HJO86]. Metallic [BWWM12]. Metamorphosis [KSK00, OOHR14]. Metaphor [BK06, FMM06, FFN19, KP14]. Metaverse [ZPL^{+}23, WSSS23]. meteorological [Sch93b, Tre95]. Meteorology [Gel07a, PSJ87]. Meter [YBL^{+}12]. Method [BDH^{+}89, BKt82, CXW^{+}20, CK85b, FM82, Fre84, Hor84, Ks22, lLK84, OATO85, PLS^{+}16, Pie87b, RSDP01, Shi83, SSZP15, UTMI11, WL88, YSW^{+}23, BKtH82, WS92]. methodology [Rob91]. Methods [ATS^{+}03, BFH86, CP13, DEN13, ELK97, GCS94, Gün20, Hal99, HKDL22, JDH^{+}89, Kro04, MQM22, MK84, ST09, TLCC17, WW08, FM87, Pin91]. Metro [BBDZ08]. Michael [SCS22]. Michel [RvW21]. Michelangelo [BRM^{+}02]. Michigan [FE97]. Micro [Tho84]. Micro-Cap [Tho84]. Microchannels [KC05]. Microcomputer [Cha85, Moh85]. Microcomputers [Lic84]. Microprocessor [GK86]. Microscope [Sin21f]. Microscopy [BHAA^{+}13]. Microsoft [Hay95b, PH98, PH99, Whi95d]. Mid [SPM10]. Mid-Range [SPM10]. Midair [RSP15]. Middle [CNL00]. Miebach [MCS22]. Mike [Whi94a]. Millennium [DeF00, Pri00, Rhy00, Sow00, Sta91b]. Miller [Whi94a]. Millions [WC11]. Mills [Ano18n]. Mimics [Ano98g]. Mind [Sin13e, Sin20f]. Mine [RW99]. Minecraft [RWB^{+}14]. MineSet [Bec97]. Ming [WWF00]. Minimal [HD01]. Minimally [BDK^{+}04, BSHW07]. Mining [Agr19, FTAT00, HKW99, KPSN04, Won99]. minute [Sim95d]. Mira [MTT83]. Mirages [BTL90, Mus90b]. MIRANIM [MTTF85]. mirroring [Kor87]. Mirrors [KB09]. Mission [SSRH95]. Missions [BHY18, GA97]. Misuse [EWF02]. Mita [MD99f]. Mitsubishi [MD99f]. Mixed [BK06, FGCB23, HSHS05, Ks22, LCR^{+}14, ML05, TYK01, WW24, ZYJ^{+}21]. Mixed-Reality [BK06, LCR^{+}14]. Mixing [KC05, Sin01a, SR10]. ML [vdEAA^{+}23]. MMX [Bli97b, PWW97]. Mobile [BLH^{+}08, BBB08, BCI08, CPAM08, Dev13, DL15, DD04, HPG08, HBP^{+}07, JBBL02, Kri03, LS08, LPMEPCM11, MD00f, MSED23, Nur08, OK09, Pie06, PARV05, PK08, Sch07, SKJ^{+}13, SB16, SCH05, WS09a, WS09b, WSKY09, ZTJ^{+}07, Ano14p]. Mobility [Hua21]. Mobrex [BCI08]. Mock [CWZ99]. MockUp [MD99f]. Mode [BQP^{+}19]. Model [AGR21b, ADBW96, BRB17, BRM^{+}02, CM86, CT09, DSK^{+}14, DGY07, DPA^{+}11, EWS96, EH99, Gol83b, HWY16, LM10, LO94, LK95a, Lil81, LSFW83, Wil13, MD99e, MMS^{+}91, MMR^{+}06, MBV17, MMD^{+}19, NL02, PBR^{+}14, SBRCD17, SM84, SHY^{+}23, Str90, WAH^{+}21, WD15a, WLVL^{+}96, XZZL19, YYiTI92, YCB^{+}22, GG87, JM95, MB91, RKRW87, RN94, Sim95b]. Model-based [LK95a, RN94]. Modeled [Ano97i]. Modeler [Coh83, MS82, SVB97, YT85]. Modelers [Hay95b, Hil82, Mil89]. Modeling [ATS^{+}03, ARS17, AJH^{+}12, BL14, BES17, BWWM12, BA13, BHG23, BG82, Bre97, CAP^{+}05, CS85, Cas87, CC20, CMP95, CIF99, CS06, Coo02b, CMPA18, Dal19, DM02a, DG99, Doo82, DMP93, EDC96, Elb05, FD03, FMMS03, FFA^{+}97, Gol00, GAW86, GS94, GM12, GJ13, HB84, Hay96b, Hed84, HYN03, Iga19, IKMF12, Jos11, KNRB12, LD99, LS21, MD99e, MD00f, MPDR14, Mas98, MMZ94, MP89, MMR^{+}06, MKKJ96, Mye82, NTH02, NG96, Nie93b, OSJ11, PJC88, PPK05, Pat12, PB06, PE09, Pie06, PP11, PPA14, Pra84a, RWB^{+}14, RV82, RV83, RR92, Rho97b, SC89, Sar90, SMWS88, Sin03c, Spa16b, SBS10, Sto99, Tak85, Til83b, TSUC86, Tur03, Tur88, UTMI11, WW86, WK94, WW08, WMV^{+}08, Wei85, Whi04, Wil85, WFOP85, WMG97, Bre96a, Chi87, Gor83, GS88, GT93]. modeling [KS94, MCW93, NFHL91, NBG^{+}94, Sim95e, TR94]. modeling/drafting [Kal87]. Models [Arb90, AGL87, Bad82, BRD^{+}20, BDC^{+}12, Bar84, Ben99, Boh95, CPCS08, CBZ^{+}08, CEO^{+}93, DNS^{+}15, FM85, FZ03, FM22, GMNF97, GSN09, HR98, HS87, IFH^{+}03, JWL^{+}18, KTA^{+}02, KMS07, KPTS89, LBD^{+}92, MSH^{+}08, Mil86, Mil88b, PSX13, Par82, PT90, Pic92, RWF02, RSDP01, Rot03, RGCB07, SADC02, SSW^{+}07, SD07, SM95, SRDT01, Sim97, SK86, aKTRL^{+}21, TW88, WZM^{+}11, Woo88, YWR09, YLR17, ASW93, Gol86, KKM96, MTT91, PT92, SG90, WF93]. Modern [RME14]. Modes [LPLN04, LFFB20]. Modification [SR97]. Modified [YS83]. Modulation [JBCS23, TH14]. Module [MD99e]. Moebius [Sin01b]. Moiré [Gla97b]. Mold [BRL^{+}17, IO07]. Molecular [GSSO05, KDH^{+}95a, Max83, MG88a, SK86, VBW94, Pal92]. Molecules [Sim95c]. Molli [SGEF^{+}12]. Mona [BGM^{+}07]. Monitor [MD99e]. Monitoring [BHM^{+}22]. Monitors [Ano14d, MD99f]. Monochromats [RGW05]. Monocular [MMP^{+}07]. Monolithic [KB92]. Monoscopic [BWP^{+}16]. Monte [XP05]. Montréal [MTT87b]. MOOC [QC15]. Mood [YP08]. Mood-Transferring [YP08]. Moon [Sin20d]. Morgan [BC19]. Morphing [LWS98, SHY^{+}23]. Morphology [VCR00]. Morphology-Based [VCR00]. Morse [SKK91]. Mosaicing [ACO00]. Mosaics [HSK02, Sze96]. Motifs [SES05]. Motion [ABWS05, BRD^{+}20, Bre97, CEMTT98, Dai89, Del98c, DLN05, Gir87, GD11, HFZ^{+}13, HHKK17, HEW17, KK02, KB82, LD08, LLZ13, Mac90, MTT85c, OZH00, Pot98, RCB98, RPM17, Sin10c, TG21, WF02, WH97, Wil87a, Wil82, WTR11, ZWWS19, vBE12, GS88]. Motion-Shape [HEW17]. Motions [PR08, VCR98]. Motivated [Dom12, Hib04].

Motivation [DSA^{+}19, ST09]. Motor [Zel82]. Mountains [Sin04c]. Mounted [Fox05, HCL^{+}20, US17]. Mouse [FPW^{+}00, LZF^{+}14, WJ88]. Move [DASF13, HMC^{+}20, Wor94]. Movement [ABS16, CCP82, Ebe05, ZGC14, EW96]. Movements [NLvLM09]. Moves [JSH16, Rib98, Whi95d, ZD98]. Moves-on-Stills [JSH16]. Movie [MTT86, MCGS21]. Movies [KZC^{+}11, WKB11]. Moving [BKP01, Ger96, Kos23, LaV07, Lea96a, ML05, Mil12, SWvdWvW12, WW86]. MPEG [Lea96a, NCO03, Whi95d]. MPEG-4 [Lea96a, NCO03]. MR [GZM97]. Mullican [Hag90b]. Multi [Ano15-44, GK86, RBE08]. Multi-Core [Ano15-44]. Multi-Microprocessor [GK86]. Multi-Volume [RBE08]. Multiactor [KSRF11]. Multiaxial [PT94]. Multicast [Fis02, MZP^{+}95]. Multicomputers [BMR90, Ell94b]. Multidimensional [DKS^{+}05, KK94, LLR09, RCB98, SCB^{+}04b]. Multidisciplinarity [EE12]. Multifaceted [RFA^{+}22]. Multifield [BHS^{+}24, LLR09, WBS^{+}05]. Multiflash [RTF^{+}05]. Multifrequency [CWBS99]. Multigrid [LGW^{+}12]. Multilabel [NRH^{+}22]. Multilevel [SBHM97, WZL^{+}21]. Multimedia [CCR10, CTW^{+}10, GP91a, GGV92, JOnH93, MSED23, PWW97, RR95, RMDT96, Sim94d, WSKY09, Woo91b, dRvWW10, Cla91, DSP91, GP91b, Pea91, Phi91, Rei91]. Multimodal [CMO^{+}99, EH03, GPS^{+}21, KBHT06, Ovi03, RC15, SJSK19, WKHL96, WVL^{+}18]. Multimodel [DCCE19]. Multiperspective [SK03]. Multiple [BMW87, CBD^{+}93, DL22, GL82, GH86, HG13, LWS98, Mil07, MKKJ96, Sin18f, Tay02, UIL^{+}06, YIK85, HO91]. Multiple-Scattering [HG13]. Multiple-Write [GL82]. multiplexed [Hod92]. Multiplexing [OOHR14]. Multiprocessor [GP89, GGV92, HDF^{+}92]. Multiprojector [SFF^{+}00]. Multipurpose [KKG^{+}13]. Multiresolution [OSRW97]. Multiscale [CD12, SD07, UZM^{+}23, YLR17]. Multiscaled [DY07]. Multiscope [NRH^{+}22]. Multisensor [Ste91]. Multisized [DY07]. Multisketch [ADZK^{+}21]. Multispectral [NK14]. Multitouchless [KGB14]. Multiuser [BWA96, BMC04, DGA^{+}97, EBG^{+}10, EYY03, MRD06, REF^{+}06, WHKM^{+}13]. Multivalue [LPK05]. Multivalued [MUN^{+}10]. Multivariate [BF93, Int00, War04, WG95b, MTS91, NFHL91]. Multiview [LYO02]. Multiway [JLC98]. mummy [YYOiT92]. Murmurations [CS15a]. Muse [Mar07]. Museum [BBB08, HCC05, Nea89a, Zha22b, ZLK^{+}22b, PCZ^{+}09]. Museums [Gel06, Bal98c]. Music [Sin03d, Tit07, YL11]. Musical [BH16, VM07, YL09]. Musical-Fountain [YL09]. MVP [GGV92]. mWorld [DGA^{+}97]. My [Bli05c, Hay96a, Hib04, Bli88d]. myComputer [Ano13f]. myCS [Ano16-36, Ano17x, Ano17y, Ano18-39]. Mystery [Abo99b, Del98c]. Myth [Ano20-62, Bre84].

’n [Abo00c, Bar00, SC12, PS86]. N-Dimensional [PS86]. Name [Jon89, WHSM17]. Naming [BBK82]. Nano [HPS20]. Nanometer [YBL^{+}12]. Nanoscale [RA06]. Nanyang [Sou04]. NAPLPS [Cha85]. Narrations [ND05]. Narrative [ALD^{+}06, BJC18, BAE^{+}18, HYL^{+}22, KWK21, WKM^{+}04, vL06]. NASA [BHY18, Del97, Mil03]. Nathalie [MCS22]. Nathan [SCS23]. National [CB86, SAB^{+}23]. Natura [MM89, Mus90a]. Natural [AL11, GPA^{+}17, Int00, JLS15, LJ15, Mil88a, NP17, OSJ11, SCS22, SVT10, SS20, WJP22, WTR11, Zhu19, MG92].

Natural-Image [AL11]. NaturaSketch [OSJ11]. Nature [Blo86, RSZ96, Sin17b, Sin18d]. Naval [ZD98]. Navigation [GSN09, HLCC08, JXW^{+}08, MYPT09, PKJ^{+}07, RLOW13, Rod14, WTPHK09, ZYJ^{+}21, dHPP10]. NBS [Ska86]. NCGA [Ano88b]. Ndebele [LMPW01]. NDT [IL83]. Near [AGL87, GMTF89, Kri05, LKS96, Sch00, XNSW19]. Near-Field [Sch00]. Near-Real-Time [AGL87, LKS96]. Necessarily [Sch86]. Neck [PP11]. Need [Ano20w, Ano21-32, Ano21-33, Dom00, Joe13, KS08, Rei17, RDDY07, Bli89c]. Needles [War04]. Needs [Jer18a]. Negawatt [BRM10]. Negotiation [Hay96d]. Neighborhoods [DY07]. NeRF [SLC24]. NeRF-In [SLC24]. nerve [ASK94, SAO^{+}95]. Net [Gla97c, MD99f, JMT^{+}01]. NetCDF [RD90]. Network [CXW^{+}20, CPN09, CXPG19, DGTK07, Dwy16, FAL^{+}06, GP93a, GLRK06, HLW13, ITSK06, KFBI22, LLY23, MD99e, MZL^{+}21, RFN89, RGCB07, She82, SSH^{+}98, SLSP22, Tre98, Wei84, WKY23, YSW^{+}23, ZHZ^{+}15, BEW91, Eic96a]. Network-Based [RGCB07]. Network-Transparent [RFN89]. Networked [FSP00, HILW98, Tac98]. Networking [Ano18x, Ano18-41, CNL00]. Networks [Bac16, CCT^{+}20, DBS96, DeF00, DD19, LFV^{+}97, LWZY19, MW82, PHE21, TPA24, WMFM13, WZQ^{+}19, ZPG^{+}21, RKRW87]. Neural [CXW^{+}20, DD19, HHKK17, PHE21, She22, ZPG^{+}21]. Neuroanatomy [SM86]. Neurological [HS04]. NeuroNet [SKS^{+}96b]. Neuroscience [JBH^{+}10]. Neurosurgery [RGA^{+}83]. Neurosurgical [RNNTD14]. NeuroTrace [JBH^{+}10]. Newdle [YLL10]. News [Ano92h, Ano93m, Ano93n, Ano93o, Ano93p, Ano93q, Ano94k, Ano94l, Ano94m, Ano97i, Ano98g, Bal97, Bal98c, Bal98b, Hay96a, Hay95b, Hay96d, IS95, ISMGG22, KND^{+}18, Lea96d, Lea97c, Lea97d, Mac97a, Whi95d, WW95, Wor96b, YLL10]. Newspaper [OBG^{+}23]. Newton [Dia94]. Next [Agr09, BLH^{+}08, DeF00, Gai17, JBD^{+}21, JS03, LCG^{+}02, LR14, Ma07, Mye85, Rhy00, RRB^{+}14, ZTJ^{+}07, vD05, Sta91b]. Next-Gen [BLH^{+}08]. Next-Generation [Gai17, vD05]. Nexus [MMMM21]. Nice [Bli93c]. Night [LLY23, Sch99]. NIH [MJM^{+}06]. No [Hay95b, WF02]. Noble [Bar04]. Node [IK15]. Noise [SKBF12, Tur02a, XP05]. Nominate [Ano20-61]. Nominations [Ano16c, Ano17a, Ano19m, Ano20v, Ano20-35, Ano20-34, Ano23q, Ano23r]. Nominees [Ano16c, Ano17a]. Nonaliasing [Fan86]. Nonconvex [lLK84]. Nondissipative [XMJ08]. Nonhuman [KP11]. Nonintersecting [SS88]. Noninvasive [WKHL96]. Nonlinear [LR90, YCB05]. Nonmanifold [YK95, CR91]. Nonparametric [JAPF07]. Nonphotorealistic [DCLK03, FM03, Her24, LS95]. Nonpinhole [RPHW11]. Nonrational [PT03]. Nonrigid [CD12]. Nonrigid-Deformation [CD12]. Nonuniform [PBR^{+}14, RSK^{+}06, YSWW08]. Nonuniform-Layered [YSWW08]. Nonwovens [VSLG22]. Normal [FZY84, Gla97e, JDZ04, SADC02]. Normalization [GMTF89]. Normals [NK14]. Note [And98, Fil84, Ger98, HB97, LR90, Mus90b, She82, WGI15, Wri98]. Note-Taking [WGI15]. Notebook [BG96, Gla02a, Gla96e, Gla96b, Gla96a, Gla96d, Gla96c, Gla97a, Gla97e, Gla97d, Gla97c, Gla97f, Gla97b, Gla98e, Gla98b, Gla98a, Gla98f, Gla98c, Gla98d, Gla99d, Gla99f, Gla99a, Gla99e, Gla99c, Gla99b, Gla00a, Gla00e, Gla00f, Gla00b, Gla00c, Gla00d, Gla01b, Gla01f, Gla01a, Gla01c, Gla01d, Gla01e, Gla02e, Gla02d, Gla02f, Gla02b, Gla02c, Gla03a, Gla03b, Gla03e, Gla03f, Gla03c, Gla03d, Gla04a, Gla04b, Gla04e, Gla04f, Gla04c, Gla04d, Gla05a, Gla05b]. Notebooks [Kos23]. Notes [Ano10c, Mac97a, Wri83]. Nothing [Sin13f]. Notion [WS90b]. Novel [ABS16, DSP^{+}14, SZZ^{+}18, XP05]. Novelty [AMDM22]. Novice [FPJ04]. Novices [II22]. NPSNET [CMBZ00]. NPSNET-V [CMBZ00]. NRC [Wor96b]. NSF [MJM^{+}06, Ano97i]. nSpace [PCH^{+}07]. NTSC [Bli93c, DS87]. nu [Nie86]. nu-spline [Nie86]. Nuclear [AMJ^{+}08]. Nucleus [CHY99]. Numbers [DH91, Sto05]. Numeric [MD99e]. Numerical [Mab82, MHTD23]. Numerically [JDH^{+}89]. Numerics [Bli07b]. NURBS [BS96, Far92, LW94, Pie91, SR97]. NURBS-Based [LW94]. NVE [FEV^{+}01]. NYSE [Del99b].

O [Gla99e, KHP^{+}11]. Oak [Sim95a].

Object [AKFPW20, BKSS07, GS87b, Gol84, HKSK17, KP14, NL02, SC89, SPMS04, SW85, SHY^{+}23, Sil95, SLSP22, TPA24, WJ88, EW96, KS94]. Object-Oriented [SC89, Sil95]. Objective [USW17, WGZ^{+}18]. Objects [BRL^{+}17, BC87, Bli05c, BH84, CCV85, CW96, CS07, DT81, FGR85, GWW86, HS87, JK09, LMA^{+}17, LFPS01, MD99e, MOP89, MSF07, MK84, OATO85, PRdJ07, PS86, Rie93, RZSP04, SS97, SWvdWvW12, SS88, SKP09, TUH^{+}90, TT84, WLL^{+}16, WWM87, YYiTI92, YY99, UO91]. Oblique [KKGH90]. observation [Hay95a]. Observations [CRA^{+}18]. Observing [LZF^{+}14]. Occluding [CS07, Her24]. Occlusion [AM04, AXP07, GSF08, SKUT^{+}10]. Occlusion-Resistant [AXP07]. Occupancy [TPA24]. Ocean [DPA^{+}11, GMNF97, HYL^{+}22, MWM87, JM95].

Oceanographic [RFA^{+}22, Ros89a]. OCME [GS12]. October [Ano15c]. Octree [DT81, FA85, KSY85, YKR^{+}84, YKFT84]. Octree-Encoded [DT81]. Octree-Related [YKFT84]. Octrees [GWW86]. Ocularist [LBS^{+}03]. Off [Hay95b, JAPF07, PLB^{+}01, Sin01c, SO95, SN86]. Off-Center [JAPF07]. Offenhuber [OCS20]. Offered [MD99e]. Offers [MD99f, MD99e, MD99d, Sim95d]. Office [BW00, GOPK18, Lea96a, LWM14, MD99d, MD00f, WFR^{+}00]. officer [ZPA95]. Offline [TS22]. Offset [ELK97]. Offsets [TH84]. Offshore [CSWB01]. Oh [Bli05c]. Ohio [CHP21]. Oil [Del99a, LLG^{+}07, MMN^{+}21, RRR^{+}16]. Okino [MD99d]. Old [Abo99b, Bre90, DS07a]. Older [LTD16]. Olfactory [NOK^{+}08]. Olympic [PCZ^{+}09]. Olympics [She06, TC96]. Omitted [HE82]. Omnimax [GH86]. On-Campus [BYP21]. On-Line [Lea96a]. On-the-Fly [MHC98]. One [Ano18-40, AL11, Arn14a, CWWW21, Dom12, Hag90d, HBB20, LLY23, Rob87, Sin03c]. One-Dimensional [Rob87]. One-Semester [Dom12]. One-Step [LLY23]. One-Way [Arn14a]. Online [BYP21, BvDE^{+}99, CSMT03, Gel07b, PFB06, TN10, YLL10]. Only [Lev00]. onto [YZM^{+}11]. Ontologies [CCE^{+}14]. Open [Ano19-43, Ano19-41, Ano19-42, Ano20-32, Ano20-30, Ano20-31, Ano20f, Ano20-54, Ano21-39, Ano21-40, Ano21-41, Ano21-42, Ano22-37, Ano22-38, Ano22-39, Ano22-40, Ano22-41, Ano22-42, Ano24m, DMJ05, GSG^{+}21, Gla99a, SSM^{+}07, Sin23f, UZM^{+}23]. Open-Source [SSM^{+}07]. OpenGL [PHAS20, RME14, RV07]. OpenGL-Based [PHAS20]. Opening [Gor96]. Opens [Mac97a]. OpenSpace [BHY18, BAE^{+}18]. Opera [Wor93]. Operation [GPA^{+}17, Chi87]. Operations [Arb90, DFP^{+}11, HHK89, LSMC04, PS86, Rob87, TSUC86, YT84, Car87]. Operator [AMJ^{+}08, XP05, Gor83]. Operators [AGP^{+}22, CGT^{+}23, MS82, PZD^{+}18]. Opponent [BBS09]. Opportunities [Ano20w, Ano21-32, Ano21-33, Bar15, BM24, CS15c, GB16, Li22b, Ma09, Mac00b, NAB^{+}23, Pac10]. Optical [BBF^{+}98, GBT^{+}02, MMP^{+}07, PTR^{+}98, TLL^{+}96, WW95]. Optics [Ano98h]. Optimal [Bli89b, LLL^{+}13, SM86]. Optimal-Scaling-Factor [LLL^{+}13]. Optimization [KRG^{+}22, NMRI06, Tau12b, HO91]. Optimization-Based [NMRI06]. Optimized [Red01, Sun21]. Optimizing [Bli00d, CTJ^{+}17, PLM^{+}19, PA18, SB97, VSLG22]. Options [Ano16-37, Ano16-38, Ano19-43, Ano19-41, Ano19-42, Ano20-54, Ano17z]. Oracle [CXW^{+}20, YSW^{+}23]. Oral [Han01]. Oranges [RSZ96]. Order [Dwy16, DH93]. Organic [IMS03, VCS19]. Organization [CRD^{+}18]. Organizing [Jon07, LC16]. oRGB [BBS09]. Oriental [Lee99]. Orientation [Ale21, HB08, YNA99]. Orientations [Lee08]. Oriented [DL16, Kos16, MTTF85, PP12, SC89, Sil95]. Origami [Gla96d, Gla96c]. Origins [Enc23a, Enc23b, Möl20, Whi20, Cro87b, McG93]. Orthogonal [LD08]. Orthogonal-Blendshape-Based [LD08]. Orthopedic [QCP^{+}10, SSS00]. Orthophotography [Lea97a]. oscillating [IP92]. Oscillation [LTS^{+}13]. Osmose [Sim96a]. Ostium [MGW^{+}18]. O’Sullivan [Tau12e]. Other [Gre86b]. Our [AAB^{+}22, BC19, Bli04a, LCB^{+}23, OCS20, SF20, Tau10c, Wat06, CNL00, Sim94a]. Out-Of-Core [FS01, GS12]. Out-of-Home [Lat98]. Outcomes [MM21]. Outdoor [Pie06, RLG^{+}02, WJP22, YNA99]. Outer [BDH^{+}89, NTB^{+}20]. Outgoing [Möl22]. Outstanding [Sin16d]. Overcoming [Gel08]. overlay [Wan93]. Overload [CAG^{+}06]. Overview [BDDH95, CS85]. Own [KGR^{+}18, Sin11c].

P [Use93, Whi23]. Pace [Kri01]. Pacific [vWNS10, FKC^{+}05]. PacificViz [Ano13b]. Pack [MD99f]. Package [RR86a]. Packet [CAG^{+}06]. Pact [Del00c]. PADL [Bro82]. PADL-2 [Bro82]. page [SWS91]. Pages [XHC^{+}17]. Pain [KGR^{+}18, SKJ^{+}13]. Paint [GV85]. Painter [Sin12d]. Painting [CT04b, Del00a, HE02, HN85, LMPW01, Lee99, Shi17, Sin04d, Zha07]. Pair [HM85]. Pairs [DPL23]. Pairwise [FAECOG19]. Paleontology [Fig01b]. Palette [CW09, Ks22, MSK04]. Palmer [Sin12d]. Pandemics [CML^{+}23, MP23]. Panel [MD99f, MD00f, Tau11c, Zuc84, Kan88]. Panels [EBN13]. Panoramas [WHF01]. Panoramic [SF24, WW24, XZZL19]. Paper [KSM85, KGK94, Lar11, Lev00, Sch08a, Zha22a, Bli87]. Papers [Ano11b, Ano14n, Ano14o, Ano15v, Ano18c, Ano19o, Ano20q, Ano20m, Ano20r, Ano20n, Ano20o, Ano20p, Ano20-41, Ano20-42, Ano21z, Ano21-27, Ano21-28, Ano21-29, Ano21-55, Ano21-56, Ano21-57, Ano21-58, Ano22-28, Ano22-47, Ano22o, Ano22-48, Ano22-49, Ano22-50, Ano23-28, Ano23-33, Ano23-34, Ano23-29, Ano23-30, Ano23x, Ano23-31, Ano23y, Ano23z, Ano23-32, Ano23-27, Ano24h, Ano24i, LIS^{+}19, WCE^{+}17, Pea88]. Parabolic [Pav85]. Paraboloid [Bli04a]. Paradigm [Koc94, LaV07, YLT18]. Paradigms [JPSM13]. Parallax [TG21]. Parallel [ABM^{+}01, BHPB^{+}12, Elb01, GP93a, HB97, ILC10, KHP^{+}11, MPHK94, MP01, MCEF94, MR87, Neu94, Pan90, PLM^{+}19, Whi94b, WHC94, XZWH14, YSSP91, YIÇ09, BBP94, HM95, Wan93]. Parallel-Visualization [KHP^{+}11]. Parallelepipeds [FA85]. Parallelization [Wri90]. Parameter [HBW^{+}14, SW22, VSLG22]. Parameterizable [HH01]. Parameterization [AESF21, Gla97f, GXMW15]. Parameterized [Par82, van96].

Parameters [BBC^{+}87, LYO02, RB00, TT00, WK90, ZGP^{+}13]. Parametric [BD89, BD90, Bli99b, Bli99a, Bli00b, GP90, HAG83, LF84, Lin99, LMD92, PB87, WWL^{+}21, ZJH^{+}18]. parametrically [Roc87]. Parametrization [BS96]. Parcel [Sin15e]. Park [SS01]. Part [Bai11, Bai13, BS86, Bli92b, Bli99b, Bli99a, Bli00b, Bli00e, Bli01a, Bli04a, Bli04b, Bli04c, Bli06c, Bli06d, Bli06a, Bli06b, Bli07a, Bli07b, DM02b, ENDAD13, Gla02a, Gla99c, Gla99b, Gla00a, Gla00e, Gla00f, Gla00c, Gla00d, Gla01c, Gla01d, Gla01e, Gla02d, Gla03a, Gla03b, Gla04f, Gla04c, Gla04d, Gla05a, MD02e, OF14, OC14, ODW15, OZH15, OWZ17, PH98, SDS95a, WG95a, WG95b, Bli94a, KmKK^{+}18, SDS95b, Arn14a, Arn14b, Bli03e, Bli03a, Bli03b, Bli03c, Bli03d, Enc23a, Enc23b, Gla02e, HHH^{+}21, PH99, WS09a, WS09b, YMK^{+}12]. Participating [dOGDM^{+}19]. Participatory [SML^{+}09, Woo15]. Particle [DPG18, EWS96, Gos90, RSP15, SFL^{+}04, WYG^{+}12]. Particle-System [EWS96]. Particles [van93, Bli91d]. Partitioning [JWL^{+}18, SS88]. partly [Her86]. Partnership [Rho97a]. Parts [Boh95, EDC96, JWL^{+}18, SS88]. Pass [Sin15e]. Passive [Bee15, OZH00]. Past [May83, O’S11, ODE^{+}13, PKT01, SKN^{+}05, Wil91, Rho91]. Patch [BW94, FM89a, LLL^{+}13]. Patch-Boundary [BW94]. Patch-wise [LLL^{+}13]. Patches [AL11, Cas87, EH99, PK87, SA85, Ste84]. Patent [WW95]. Path [GPW^{+}19, Tau10c]. Pathologic [FZY84, SSZP15]. Paths [PKN^{+}22]. Pathway [LWZ^{+}18]. Pathways [CGJ^{+}17]. Patient [RPOC21]. Patients [CKPB16, RRF^{+}23]. Pattern [Vos85]. Patterns [ADZK^{+}21, CWGC90, Gla97b, Jer18b, LZF^{+}14, MHTD23, NB04, OF14, OC14, ODW15, SS19, SD07, WO13]. Pausch [SK13]. PC [BEW^{+}98, Cav87, DS87, Hal87, JOnH93, Jam87, Laz87, MD99f, MD00f, Rib98, WPLM01]. PC-CAD [Cav87, Hal87]. PCs [GRW99, Hay96d, PWW97]. Peace [Hay96d]. Peacekeeping [LSMC04]. Peak [RV07]. Peano [WN82]. Pebbles [Sin04c]. Pedagogy [BFR21, Die11]. Pedestrian [Fox05]. Pedestrians [PE09]. Pediatric [RRF^{+}23]. Peephole [BK06]. Peer [BAB^{+}23a, BFR21]. Pen [HHLL14, KL09]. Penrose [Gla98f]. People [GJM18, IS95, MFM^{+}16, Shu23]. Per-Element [TH14]. Perceived [AMDM22]. Perceiving [WFG92]. Perception [CS17b, DPA^{+}11, GFW^{+}16, HE02, LLML16, MKH20, MZ22, SHY^{+}23, Sto12, Sun21, TBP01, TF01, XZWH14, EFB^{+}94]. Perceptions [Lec17]. Perceptive [BTD^{+}03, LSR^{+}00]. Perceptual [DL22, EH03, HW82a, ITW01, LWZY19, SKBF12, SB13, WGZ^{+}18, War08, BHMN95, Rob88]. Perceptually [HBBG95, Red01]. Perchance [Sin20b]. Perfect [Ano16d, WWF00, Wil91]. Perfection [Abo98b]. Performance [Ala92, BRG^{+}00, BKV^{+}10, BBC^{+}87, CS19, CB86, DD84, FT00b, GRB^{+}85, HML83, Hoe12, HKSK17, MMD^{+}19, NB94, PLM^{+}19, Pra84b, SO95, SH07, WP87, War85, Bro89, DWL^{+}92, HM95]. Performance-Based [HKSK17, SH07]. Performances [BH16, LFFB20]. Performative [Her23]. Perils [Gla97f]. Peripheral [Cav87]. Peripherals [MD99f]. Permanent [Sin19f]. Perq [HW82b]. Persia [Abo99a]. Persistent [LJ96, SKJ^{+}13]. Person [HLCC08]. Personal [CLms15, GM87, Hec86a, KWK21, KW15, Mue18, Per21, RLCI24, Tho84, TLCC17, TC15, WTLM15, Woo15, YST84, de 86]. Personality [DPA^{+}11]. Personalization [SCOSL13]. Personalized [FH19, TMB^{+}13]. Perspective [Ang17, Bee82, COM15, DFL^{+}18, Enc06, FH84a, GH86, HPS20, Kir21, Kri00, MLJ19, Mar84, MTN18, Mye82, Rob87, YCB05, Bli93b, RMM^{+}91]. Perspectives [BCF^{+}08, KN07]. Pervasive [Ano20j, Ano20k, Ano20l, Ano21y, Ano22z, Ano22-27, Ano22q, Ano23u, DB03, TCB03]. PET [Ger91]. Petascale [BHAA^{+}13]. Petra [VFLJ02]. PEX [RFN89, SRK90]. Phantom [SS97]. Phantom-Based [SS97]. Pharmaceutical [SBCvdS04]. Phase [AP11, BQP^{+}15]. Phase-Indexed [AP11]. Phases [NLvLM09]. PhD [Car16]. Phenomena [BBF^{+}98, CS15a, Mil88a]. Phenotypical [LZF^{+}14]. PHIGS [AEB86, Sch86, SBM86, TM88]. PHIGS-Based [TM88]. Philosophy [OBG^{+}23]. Phobias [JAM^{+}05, WAB^{+}13]. Phone [BBB08]. Phone-Enabled [BBB08]. Phones [HB08, Sch07, WS09a, WS09b]. Photo [MD99f, aKTRL^{+}21]. Photo-Realistic [aKTRL^{+}21]. PhotoCloud [BBT^{+}13]. Photograph [MTC07]. Photographic [DBV^{+}08]. Photographs [SRDT01, Ano98g]. Photography [DS07b, Lev10, PIBV11]. Photometric [NK14, XNSW19]. Photopolymer [BRL^{+}17]. Photorealistic [ARL^{+}10, CG94, HCC05, TK03]. Photos [LBW^{+}17]. Photosketcher [ERH^{+}11]. Phylogenetic [HBW^{+}14]. Physical [BNA^{+}22, CS15a, DMAS17, HAB^{+}01, HW18, JBCS23, KCMK20, LFOI15, PCG^{+}21, RZSP04, Smi84b, SDPW20, VK17, WLVL^{+}96]. Physical/Biological [WLVL^{+}96]. Physicality [Off20]. Physicalization [DMSJ23, HHH^{+}20, HHH^{+}21, KWK21, KJA^{+}18, PGV20, Per21]. Physicalizations [MSED23]. Physically [CPN09, CPK06, LX04, RL07, TW88]. Physician [GRB^{+}85]. Physics [ASP^{+}21, BB82, DPG18, HZ11, Kan09, hKKSK17, MQR99, SL11, VBHS11]. Physics-Based [HZ11, hKKSK17, VBHS11]. Physiological [HS04, Mur84a]. Physiology [TA85]. Picks [Hay95b]. Pictorial [Fei85].

Picture [Gla02c, LSB18, Par98, PM82, SM84, Woe83, Wor83]. Picture-Plane [SM84]. Picture-Processing [PM82]. Pictures [Han14, Lea97a, Sim97, Whi95c]. Pie [CWZ99]. Pieces [Gla02f]. Piecewise [PR08]. Piegl [Use93]. Pietà [Abo99b, BRM^{+}02]. Pinhole [PBR^{+}14]. Pins [JPSM13]. Pioneering [Swe01]. PIONS [BM86]. Pipeline [Bli96d, FPI84, Mye84c, Bli91b, Bli91c, Bli91d, Bli92c, Bli93b]. Pitfalls [EAM22]. Pixel [AMS91, Bli91c, Bli05d, Bre96b, DY07, DSKH95, Hag90d, Mam89, GFB87]. Pixel-Processing [Bre96b]. Pixel-selected [AMS91]. PixelClipper [WSP^{+}20]. Pixels [Bli98b, Lea96c]. Placement [Ch’11, Jon89, SK97, WJ88]. Plagues [Lea96d]. Plan [Sin17f]. Planar [Cho95, McL83, SB02, Wan10]. Plane [LF84, SM84, MG92]. Planes [Rie81]. Planetary [FPH^{+}06, GS87a]. Planning [BTTU84, Del00b, DPL^{+}11, FAZH17, FFA^{+}97, KCW^{+}17, MP89, MJC08, RBBS06, RNNTD14, WZM^{+}11, WKHL96, YMDO93, Zha96a, SSRH95]. Plans [CTJ^{+}17]. Plant [AMJ^{+}08, DP82, MD99f]. Plants [Ch’11, LD99, Pru90]. plasma [PCS95]. Platform [AGP^{+}22, BHP07, CJS^{+}19, OBO^{+}23, SADH23, VPVG21]. Platforms [BRD^{+}20, Lev10]. Platonic [Gla96d]. Plausible [PE09]. Play [Gha12]. Player [AY16, DEN13]. Players [PPA14]. Playground [Rou09]. Playing [DB16, Lec17]. Pleasing [PKM18]. Pleasure [Bli98f]. Plenoptic [GLC11]. Plot [Ks22, Ks22]. Plots [FM89b, HML83, TS97]. Plotting [BDJ86, HDF^{+}92, HP91, Ks22]. Plumes [LSB21]. Plush [LMZ11]. PMUVis [AGP^{+}22]. Pneumatic [HF18]. Podcast [Ano22-72, Ano22-73, Ano22-74, Ano23-74, Ano23-75, Ano23-76, Ano23-77]. Podiy [II22]. Poetry [BRG^{+}00, Sin14d]. Point [Bli97a, BA85, CW88, DD04, Fre84, Gal89, GGS^{+}18, GHQ04a, GHQ04b, JDZ04, KPSN04, LCW12, MZL^{+}21, PSX13, PMI06, PG04, SPMS04, SBMK20, SGG^{+}22, SBC15, XNSW19, Gro06, CW87, Coq87, Sim93c]. Point-Based [DD04, PG04, SPMS04, SBMK20, GGS^{+}18]. Point-Cloud-Based [PSX13]. Pointing [AA09]. Points [Bli99b, Pie87b, PT00, Sin04a]. Poisson [FHH^{+}21]. polar [Sei93]. Polarization [SKS^{+}13, WK90]. Polarized [ZGP^{+}13]. Polarizing [BWWM12]. Policy [KNRB12]. Political [CKJC12, WW95]. Pollutants [PCG^{+}21]. PolyCover [Ale21]. Polygon [Bar91, GPR94, GG83, Lev90, OATO85, Van10, Whi94b, HBW95, Til81, Wan93]. Polygon-Based [Van10]. Polygon-To-Rectangle [GG83]. Polygonal [IFH^{+}03, KT96, Lue01, ODW15, SP04a]. Polygonalization [HW90]. Polygons [Gla99d]. Polyhedra [lLK84]. Polyhedral [HHK89, SS88, Tur88]. Polymorph [LWS98]. Polynomial [Bli00e, Bli01a, FM82, Man94]. Polytopes [OWZ17]. Pong [BRL06]. pool [RHC95]. Pooling [ABWS05]. Pop [Gla02e, Gla02d]. Pop-up [Gla02e, Gla02d]. Populace [BO09, PD09]. Porcelain [CCT^{+}20]. Portability [Bri90, SQG^{+}05]. Portable [BFB^{+}98, GOPK18, Lea96d, MD00f]. Portfolios [MML^{+}15]. Portrait [CBC^{+}20, ZJH^{+}18]. Pose [GDY^{+}11, RS14, SB02, SLSP22, TCB03]. Pose-Aware [TCB03]. Poses [WC11]. Posing [KDMH17, WC11, WTR11]. Position [BTM17]. Position-Based [BTM17]. Positioning [BMW87, HB08, RDRS10]. Possibilities [Hua11, Lea96b, Sin06c]. Post [CGT^{+}23, Mac09]. Post-Earthquake [CGT^{+}23]. Post-Traumatic [Mac09]. Posters [JPSM13]. Postgraduate [ZD98]. Poststroke [DSA^{+}19]. posture [LK95a]. Potential [LaV07, Myn08, PD21, SF20].

Potter [CBB^{+}14]. Pottery [CCC18, DHC^{+}22]. PotteryGo [CCC18]. Pouches [II22]. Power [AMJ^{+}08, Ahr15, AGP^{+}22, JAC^{+}08, LCC^{+}96, PS09, Sin11e, Sin20c, WHC^{+}14, WSKY09]. Powerful [MMZ94]. Powering [WFZB21]. Practical [BBS09, Die11, GS12, SKBF12, SL11, WS09a, WS09b, WLC^{+}12, WD15a]. Practice [FL81, MRB19, Rei23, Sin15a, WMV^{+}08, WGI15, Bli94a]. Practices [Dom12, KBHP14, TBFGC23]. Practicing [LRS09]. Practitioner [Kir21]. Pragmatic [ST09]. Prajna [Swi10]. Preach [LRS09]. Precision [MMP^{+}07, PCJ^{+}16]. Precomputed [DKS^{+}03, DYS^{+}07]. Predicting [JSH16, WJP22]. Prediction [EH99, GPS^{+}21, WLL^{+}09]. Predictions [VP16]. Premier [Tur02b]. Premo [DiL15]. premultiplied [Bli96b]. Prepare [Ano17-28, Ano17-29, Ano17-30]. Preparing [Ren15]. Preprint [Ano20-63]. Preprocessing [SSG95]. Prescription [ZGL^{+}17]. Presenccia [SFT^{+}07]. Presence [BTD^{+}03, SFT^{+}07, Sra23, BRD17]. Present [Jar83]. Presentation [Hue83, Kos16]. Presentation-Oriented [Kos16]. Presentations [CLms15, MD99e, Wri83]. Presenter [NZB00]. Presenting [RXB09, SHY^{+}23]. Preservation [HJCW06]. Preserving [Bar81a, CWL^{+}10, DMR83, FM89a, GXMW15, LO05, McL83, MZL^{+}21, RGW05, SLNG07, FM87, Rou87]. Presto [FY97]. Pretense [Whi95b]. Pretrained [SLC24]. Pretty [Han14, Sim97]. Priced [MD99f, MD99e]. Primer [SDS95a, SDS95b]. Primitives [BRD^{+}20, CS06, Woo88, DSP91]. Prince [Abo99a]. Principle [Jon07, Nie07, OWZ17]. Principles [hCRBK98, HH82, Mur84a, NFMD90, ND05, RC15, War81b, War81a, War88]. Printable [JWL^{+}18]. Printed [EH99, WLL^{+}16]. Printer [MD99f]. Printing [Ano15-33, BRL^{+}17, BA13, BU23, MFM^{+}16, RBK^{+}13].

Prints [Lea96a, US17]. Prior [dOGDM^{+}19]. Priority [FSP00]. Priors [SLC24]. Prism [Sin12e, Kor87, WS92]. prism-based [WS92]. Privacy [Ano19-52, Ano22-72, Ano22-73, Ano22-74, Ano23-74, Ano23-75, Ano23-76, Ano23-77, Ano14t, Ano19i, Ano19j, Ano21l, Ano21m, Ano22r, Ano23n, Ano23o]. Prize [FKK^{+}10]. Probabilistic [GPW^{+}19]. Probe [IL83, MG04]. Problem [GASF94, GV85, GRvL^{+}23, Hor84, MTT85c, RPOC21, RdCFP^{+}20]. Problem-Solving [RdCFP^{+}20]. Problematic [Gla97f]. Problems [Bli98c, Che05, Fol00, Hib04, Joh04, MC85, MJS20, vD05]. Procedural [BHG23, Hed84, Pat12, WW08, WMV^{+}08, GS88]. Procedurally [FB89, WBS^{+}05]. Procedure [CR94, Fre86, SR85]. Procedures [KL18, NTS^{+}07]. Process [Ber83, Bra86, CK85b, FM86, Swi10, Mak87]. Processes [DJS^{+}09, EAM22, SOSG08, GG87]. Processing [AM84, AJK15, BFK^{+}84, Bor89, Bre96b, CCG^{+}04, DSKH95, KB88, KB92, LKS96, LPZ12, LBSS85, LS21, MK84, PM82, SMWS88, Spa16b, Tau12b, Wix83, Zha22d]. Processor [FPI84, Gol84, GKB89, GVA86, IM98, MD99e, WMS92, YT85, ASP^{+}86]. Processors [MD99e, HO91, Saw08]. Procrastination [ST09]. producing [GU86]. Product [Ano95a, Ano95b, Ano95c, Ano95d, Ano96b, Ano96c, Ano96d, Ano96e, Ano96f, Ano96g, Ano97b, Ano97c, Ano97d, Ano97e, Ano97f, Ano98a, Ano98b, Ano98c, Ano98d, Ano98e, Bli03e, Cai20, CCVA02, DB03, HML83, SR16, SH97, WJ93]. Production [CPA^{+}10, Coo02b, FFA^{+}97, Hee99, II22, Lea96d, RRR^{+}16, TR83, Woe83, Wor83, Zha96a]. Productivity [BA85, Bra86, Hal87, MD99f, Pal81, Wei84]. Products [Ano88c, Ano93u, Ano94n, Ano04b, Ano04c, Ano04d, Ano04e, Ano04f, Ano05b, Ano05c, Ano05d, Ano05h, Ano05e, Ano05f, Ano05g, Ano06c, Ano06d, Ano06e, Ano06f, Ano06g, Ano06h, Ano07c, Ano07b, Ano07d, Ano09b, Ano09c, Ano09d, Ano09e, Ano09f, Ano09g, Ano11d, Ano11e, Ano12b, Dil96, Dil97, Dil98b, Dil99, Dil01, Enc04, Kas08, Kas12c, Lea96a, MD95a, MD96c, MD96a, MD96b, Mac96, MD97a, MD97b, Mac97b, MD98a, MD98b, MD98c, MD98d, MD98e, MD98f, Mac98, MD99f, MD99e, MD99d, MD99a, MD99b, MD99c, Mac99b, MD00f, MD00b, MD00c, MD00d, MD00e, MD00a, MD01a, MD01b, MD01c, MD01d, MD01e, MD01f, Mac01, MD02a, MD02b, MD02c, MD02d, ME02, Mac02, ME03a, ME03b, ME03c, ME03d, ME03e, ME03f, Mac03, ME04, Mac04]. Products [MK07a, MK07b, MK07c, Wan10, Zin08a, Zin08b, Zin08c, Zin08d]. Professional [Ano24g, CK85a, Ano21u, Ano22w, Ano23v, Ano23w, Ano24f]. Professionals [MD99e]. Profile [PGV20]. Profiler [WHSM17]. Profiles [TH84]. Program [LRS09, MD00f, Mei85, Saw83]. Programmable [CJS^{+}19, GWH05, MD00f, WSKY09]. Programming [BMR90, PARV05, RWB^{+}14, TS13, TS22]. Programs [DiS14, Doo82, HM95]. Progress [RMG16, Woo15, vDFL^{+}00]. Progression [Fet82]. Progressive [LJB^{+}13, NR03, OKKP13, SKBF12]. Progressive-Rendering [OKKP13]. Project [CK85a, CDLS19, RBMS17, ARL^{+}10, Ano98h, JMOG99, Pit93, SFT^{+}07, TSF^{+}98]. Projected [WS90a, WFR^{+}00]. Projecting [Bal98a]. Projection [EBBS00, Gre86b, HJS00]. Projection-based [HJS00]. Projections [YCB05]. Projective [HBG04, Rie81]. Projector [BVRF05, MD99f, MD99d, MSWT14]. Projector-Based [MSWT14]. Projectors [MD00f]. Projects [Ber99, BKP01, BSHB01, Bry97, CWZ99, CMBZ00, CSA22, CMO^{+}99, CKP96, DB96, Del00c, DSP^{+}14, Dio97, EBT98, EBBS00, FLMZ97, FWSN86, FFA^{+}97, FPW^{+}00, GMNF97, Goe96, GRW99, HILW98, JW01, JLC98, KKV^{+}02, KKM96, LMPW01, LD97, LRS09, MR98, MD99d, MRBW12, RSS^{+}09, RBW05, RWF02, RDO^{+}96, Ros97, RDDT97, Rou09, SS97, SS01, SK03, She04, SPS99, SSS00, SD99, Tac98, WS09a, WS09b, WWG97, ZD98]. Prolog [FWSN86, GWA84, Jon89]. Prompt [WW95]. Propagations [GPW^{+}19]. Properties [lLK84, WJP22]. Property [Whi95d]. Proportional [TLT11]. Proportional-Derivative [TLT11]. Proposal [Hai87, Lea96a]. Proposals [Ano14b, Ano20s, Ano22-29, Ano14a]. propositional [PJJF96]. Propping [NZS21]. Prosodic [BRB17]. Prostate [BSM^{+}15]. Prostheses [GRWN87]. Prosthesis [VCBS97, RKRW87]. Protecting [VP99]. Protection [Lav83]. Protein [AEFQ96, SGEF^{+}12]. proteins [OG91]. Protocol [Agu86]. Protocols [BS03]. Prototype [CMSEM08, GAe^{+}24, VM07]. Prototypes [MSF07]. Prototyping [Bai95, KRB11, MSRMH09, TBP01, UVa95, ZP95]. Provenance [BGM19, BWD^{+}19, FJKTX19, MRB19, XAJK^{+}15]. Provides [Hay96d, Lea97b, MD99d]. Providing [HH95, KKAM84]. Proview [MD99f, MD99e]. Proxies [NZS21]. Pruning [GMTF89]. Pseudoperspective [Gol01]. Psychological [HAB^{+}01]. Psychophysical [HS04, FG88a]. Public [BHY18, BAE^{+}18, CT09, FZH^{+}13, HdlRL^{+}13, HHY20, HCV^{+}13, HPS20, KKG^{+}13, LB20, RLOW13, RS14, SPS99, WSP^{+}20, BB13]. Publication [JPSM13]. Publications [Ano23-63]. Published [Ano20-32, Ano20-30, Ano20-31, Ano21-39, Ano21-40, Ano21-41, Ano21-42, Ano22-37, Ano22-38, Ano22-39, Ano22-40, Ano22-41, Ano22-42, Ano24m]. Publisher [MD99e]. Publishing [MD99e, DFL87, MD99d]. Puck [Cav97]. Pulsatile [SFL^{+}04]. Pulsing [CM16]. PUMA [PLM^{+}19]. PUMA-V [PLM^{+}19]. Pump [OKET12]. Puppet [Gha12]. Puppetry [Stu98]. Pure [Sin16e]. Purpose [GKB89]. Pursuing [AM90, CS15b, Kir21]. Pursuit [Got09]. Pushing [BU23, LCS21, Ma23, Ric15]. Put [CW96]. Putting [Agr18b, Gla02f, Han14, Sim96b]. Puzzles [RSDP01]. PVG [FHH^{+}21]. Pyramid [LWZY19]. Pyramidal [De 89]. Python [AJH^{+}12].

Quad [Cam17]. Quadratic [Bli05a, Bli06d, CW88, GF86, CW87]. Quadric [Gol83b, Mil88b]. Quadric-Surface [Mil88b]. Quadrics [Gol83a, MG92]. Quadrilateral [GXMW15, Pan15]. Quadtree [YS83]. Qualities [DiS14]. Quality [AESF21, Ano23-84, Ano23-85, Ano23-86, CCVA02, Got82, LLML16, Rot03, SGC21, VSLG22, WGZ^{+}18, ZYL^{+}17, FG88a, HBW95]. Quantification [KHP12]. Quantified [CLms15]. Quantitative [AHP^{+}10]. Quantitatively [BAB^{+}23a]. Quantities [LV23]. Quantization [Bli94c, XJ94]. Quantum [Ano20-52, Ano20-53, Ano23-34, Ano23-66, Ano23-67, BAB^{+}23b, Gla01c, Gla01d, Gla01e, vOW15]. Quartets [Sin23c]. Quartic [Bli02a]. Quasi [LMWC17]. Quasi-Hamming [LMWC17]. Quaternion [ZZ05]. Queries [FDK12]. Query [ADZK^{+}21, SJSK19]. Questionnaires [SPT^{+}15]. Queue [Enc17b]. Quick [FM82]. Quo [KH14]. QuteVis [ADZK^{+}21].

R&D [ZHM^{+}03]. Races [Kop03]. Racing [HS01]. RADAR [DMSJ23]. Radial [MW82, MBV17, SZZ^{+}18]. Radiance [Wil13]. Radiation [CWGC90, Mur84b].

Radical [Sin21e, SH06]. Radioactive [GAe^{+}24]. Radiosity [CGIB86, GSG90, GCS94, LTG92, Gre91b]. Railway [LAC^{+}14]. Rainbow [Ano22-72, Ano22-73, Ano22-74, Ano23-74, Ano23-75, Ano23-76, Ano23-77, BT07, WSS23]. Rainbows [Sin20a]. Raising [Bad23]. Rama [Gha12]. Ramakrishna [Ano19c, Ano19d, Ano20b]. Random [TH14]. Randy [SK13]. Range [HNC^{+}04, JLW08, SPM10, XPH05]. Rank [XHC^{+}17]. Ranking [CPG^{+}16]. Ranks [PBV16]. Rapid [Bai95, BDC^{+}12, CG85, MSRMH09, Rob89, UVa95, ZP95]. Rapid-Prototyping [MSRMH09]. Rare [SSZP15]. Raster [Chr83, Eas90, FPT^{+}82, FI83, GH86, HDF^{+}92, HW82b, MR87, Rie93, RR86a, Whi84]. Rastering [Bre87]. Rasterizing [Tau94]. Rate [ZYJ^{+}21]. Rates [SOM^{+}10, Gre91a]. Ratio [ANS15, Bli98d]. Rational [Bar93, Bli99b, Bli00b, EK99, Pie87a, PT89, Til83a, CP92]. Rau [Ano19c, Ano19d, Ano20b]. Ray [AH95, BTL90, Dia91, Dia94, DSW07, FTI86, Gla84, Gla88a, GS87b, GP89, KS97, Lev90, LAM01, MMS^{+}91, MMZ94, MSRMH09, Mus90b, PB85, PK87, RBE08, SKBF12, SF24, SB86b, WK90, WW24, WKS86, YCK92, AMS91, BBP94, CH95, Gre91b, Woo93]. Ray-Casting-Based [MSRMH09]. Ray-Traced [AH95]. Ray-Tracing [FTI86, PB85]. Raycasting [HROIK16]. Rayground [VPVG21]. Rays [Mac97a]. Re [Sim96c]. Re-creating [Sim96c]. Reach [Kri05]. Reaching [HS01, LCI^{+}20]. reaction [CR95]. reaction-diffusion [CR95]. Read [Lar11, Mol18a]. Readability [HWP^{+}19]. Reading [BKS20, Har00]. Readings [BSS20]. Real [ACL^{+}12, AGL87, AB14a, BWWM12, BCPK02, Bli05b, BPB09, Bro99, BFW99, CZAA23, CSWB01, CdHM97, CT04b, Coo02a, DYS^{+}07, ERS13, EBG^{+}10, GE97b, GMTF89, GRW99, GRB^{+}85, Gos90, GFW^{+}16, GDY^{+}11, HE82, Jai00, JWSG10, KMTM^{+}98, KVM^{+}05, KMS07, Kri05, SKUT^{+}10, Lea97b, LKS96, LMA^{+}17, ML05, MD99f, MD99d, MRB19, MCSH21, MYPT09, MHC98, Mil12, MMR^{+}06, MYK^{+}08, NTH02, NL01, OSK02, PWH04, PFS03, PD21, Rao97, RTF^{+}05, SKR18, SSB06, SMP^{+}22, Tac98, TYK01, TG21, UCK^{+}04, Wee02, Zha21, Fan86, MF89, BW00]. Real-Time [ACL^{+}12, BFW99, CSWB01, CdHM97, CT04b, DYS^{+}07, ERS13, EBG^{+}10, GMTF89, GRW99, GRB^{+}85, GFW^{+}16, GDY^{+}11, HE82, JWSG10, KMTM^{+}98, KVM^{+}05, KMS07, Kri05, SKUT^{+}10, LMA^{+}17, MD99f, MYPT09, MHC98, Mil12, MMR^{+}06, MYK^{+}08, NTH02, OSK02, PFS03, SSB06, Wee02, SKR18, Tac98, Fan86, MF89]. Real-to-Synthetic [SMP^{+}22]. Real-World [BWWM12, MRB19, RTF^{+}05, TG21, UCK^{+}04, Zha21]. Realism [Abo00a, Ama87, Kul09, SKMY09, YMK^{+}12]. Realistic [Ch’11, CT04a, CGIB86, HG83, IKSZ03, JWSG10, JF99, KMTM^{+}98, KYK^{+}24, LYO02, MTTF85, OBG^{+}23, SKMY09, SBS10, Str90, aKTRL^{+}21, TLCC17, TR93, Web08, Wil87b, Yan21, YMK^{+}12]. Realities [Enc15b, FGCB23, LTB22, Sin12a, SR10, Sim94c]. Reality [AKFPW20, AWB^{+}97, Ano98g, Ano16b, AR99, AB14b, ABB^{+}01, BN95, BKV^{+}10, BKP01, BPB14, Bre84, Bro99, BVD19, BK06, CLP^{+}14, CML^{+}23, CEMD22, CGT^{+}23, CZP^{+}10, CNL00, DFP^{+}11, DBS96, DJM13, DSA^{+}19, DPG18, DOH^{+}20, EGR94, EBBS00, FLMZ97, FCS22, FLSG98, Got09, GDY^{+}11, HROIK16, HILW98, HSHS05, IHL18, Jai00, JAM^{+}05, JBBL02, KN07, KYK^{+}24, KKV^{+}02, KBHT06, KLL^{+}15, LTB22, LO94, Lea97b, LD97, LLG^{+}07, LOB08, Liv05, LRZ23, LCR^{+}14, LFMGPG23, ML05, MRL13, MZP^{+}95, Mac09, MD99e, MD99d, MSU^{+}21, MSWT14, MSDS23, MKH20, MKKJ96, MSED23, Nav04, NTS^{+}07, NdSM^{+}16, NCQ^{+}11, NZS21, NTB^{+}20, OBG^{+}23, PHE21, PCZ^{+}09, PPKM15, Per16, Pie06, PD21, PBSB08, PXC02, RGRLFM21, RRF^{+}23, RFK^{+}13, RBW05, RBBS06, RLG^{+}02, RMR^{+}19, Rok96, RBF95, Ros00a, SPT^{+}15, SSW^{+}07, SF20, SGC21]. Reality [SGE^{+}18, SC21, SVT10, Shn03, SPS99, Smi87, SD99, SDM22, Sra23, Sto99, Sun21, SK05, TYK01, TSK97, TG21, UVa95, VIK^{+}02, WS09a, WS09b, WW95, WW24, YNA99, ZYJ^{+}21, Bal98c, Hag91a, Kah94, Kru95, MT94, Pap96, RBOV94]. Realization [Spi86]. Realizing [Iwa99, SFT^{+}07]. Really [Rhy03, WPO96]. Realm [Sin02a]. Realtime [SS19]. Reasoning [DJS^{+}09, EAM22, Gla98b, Swi10]. Reborn [RBT98]. Reception [Lea96e]. Receptors [AYS98]. Recipe [Kir21, Pic89]. Recognition [CD12, HR98, MD00f, NF11, OSK02, RS14, TS22, Vos85, WB85, ZGC14, Ano15l, Sim94a]. Recognizer [TO83]. Recognizing [SG90]. Recollections [Ano08b]. Reconfiguration [ABVA08]. Reconstructing [GHCW21, HTZ^{+}19, SRDT01, Woo91a]. Reconstruction [BGSC85, Bee15, EHBPG04, HWY16, KS20, KDH95b, KW10, KKGH90, LMP89, MBV17, PT94, RR03, TPA24, WOF96, YMS01, PFH95]. Recording [PVV^{+}03, DS87]. Recovering [DJS^{+}09]. Recovery [Bar97a, CD12, LLG^{+}07, MMN^{+}21, ZZ99]. Rectangle [GG83]. rectangular [Nie86]. Recurrent [DD19, XZZL19, Har96]. Red [GA97]. Redesign [Sin21a]. Redirected [NPB^{+}18, SB13]. Redistribution [GSG90]. Redo [TSUC86]. Reduce [NdSM^{+}16]. Reduced [hKKSK17, Voo89, PABS86]. Reduced-Complexity [Voo89]. reduced-function [PABS86]. Reducing [CR15, PT00, TCH^{+}96, Tur02a]. Reduction [HTZ^{+}19, SKBF12, XP05, CP92]. Redundant [Woo88]. Reeb [SK91].

Reference [CM86, Sch11]. refinement [Rou87]. Reflectance [BWWM12, CPK06, OOHR14, WJP22]. Reflecting [Sin12e]. Reflection [Gla98e, MD18, TT00, WKM^{+}04]. Reflections [CRA^{+}18, CSA22, Kd21]. Reflex [SBHM97]. Refraction [BHPB^{+}12]. Region [FMM06, WWL^{+}21, YNH04]. Region-Display [FMM06]. Regions [Pav85]. Registration [BN95, CD12, TUH^{+}90, YNA99, MNR^{+}91]. Regression [DCCE19]. Regressive [PHE21]. Regular [OWZ17]. Regulation [RdCFP^{+}20, RRV^{+}15]. Rehabilitation [DSA^{+}19, Tak05]. Rejects [Hay95b]. Rejoining [YSW^{+}23]. Related [SDS^{+}11, YKFT84]. Relationship [IK15]. Relationships [Sin17a, WFG92]. Relaxation [GCS94]. Relaxing [Sim96a]. Release [Sin17a]. Released [MD99f, MD99e]. Releases [MD99e, MD99d, Wor96b]. Reliable [Li22a]. Relics [ZZ99]. relies [Bro89]. Relighting [WHF01]. Remaking [Hag90d]. Remedy [Hay96e]. Remember [KMCMT09]. Remembering [Ano20-62, LFE^{+}18, Whi23]. Remembrances [Ano08b]. Remeshing [PFB06, Pan15, XZWH14]. Remote [APB^{+}21, BBT^{+}13, FT00b, GPA^{+}17, GA97, KP18, LS08, LD97, PBM^{+}96, Tac98, VPVG21, ZSK^{+}21]. Removal [MR87]. Removing [Boh95]. Renaissance [Fig00]. Rendered [YZM^{+}11]. Rendering [ADBR^{+}12, AFYC03, iAH03, Bai01, BL14, BHPB^{+}12, BPB09, Cha97a, CJS^{+}19, CT04a, DM02a, DGY07, DCLK03, DD04, DW87, Elb01, FS01, FM03, Fri20, GP93a, GF86, GMTF89, GJ13, Hai06, Her03, Her24, HNC^{+}04, HBP^{+}07, IM98, JDZ04, KS98, KSD04, KMM^{+}01, KY97, SKUT^{+}10, LS95, LKV90, LPLN04, Lev90, Lev00, LS04b, LS04a, LO05, LTSE92, LX04, MP01, MD99f, MAB05, MYPT09, ME22, MCEF94, MD95b, NTH02, Neu94, NFMD90, Nie07, OKKP13, Pan94, PQC^{+}12, RB00, RPHW11, SPMS04, SCB04a, SBMK20, SKP09, SB03b, SFCD99, TLC02, THB^{+}90, UO93, VB24, VPVG21, Wee02, Whi94b, WHC94, WMG00, WW24, WPLM01, Yan21, YSWW08, Bli93a, DWL^{+}92, MPHK94, PB87, RHC95, SSG95, WPO96]. Renderings [Dev13]. RenderMan [AM90, HC21]. Renders [MD99e]. Rendez [MTT87b]. Rendez-Vous [MTT87b]. Repellers [PS88]. Replace [Gla03e]. Replicating [Bal98c]. reply [Use93]. Report [APB^{+}21, Atw84, FGCB23, FIB^{+}20, KJA^{+}18, PBSB08, Ric85, Wor96b, vDFL^{+}00, MJM^{+}06]. Reports [PCG^{+}21]. Repositories [PKJ^{+}07]. Repository [KÇC^{+}01]. Representation [BD85, Cam17, CNT^{+}97, HKDL22, HMC^{+}20, KSD04, Max83, RC15, SR95, She22, Til83a, WWL^{+}21, XZZL19, Bar83, CR91, GD93, Mak87]. Representations [DPL23, HB16, HW18, KSY85, MKFN87, MHTD23, Men94, MMZ94, SCB^{+}04b, Wil87c, Mas18, Rob91]. Representative [GHCW21]. Representing [Bar93, CCV85, CK85b, FB89, GZS88, Lee08, Pie87b, Sto05, WW07, ZTT99]. Reproducibility [FF20, Rei23]. Reproduction [MD99e, RGW05, SPA^{+}16a, TR93]. Reprojection [AH95]. Reptile [ZLK^{+}22b]. Repulsive [HF18]. Repurpose [Sin14e]. Requirements [BHK91, GVA86]. Requires [SR13]. Rescue [CGT^{+}23, Sin23c]. Research [Ano14b, Ano19-53, Ano20-63, BMC04, CK85a, CPAM08, CMZ01, CSA22, CPAV^{+}17, Del00c, DEN13, Ear98, ENDAD13, EFB^{+}94, Enc15c, GPW^{+}19, GAe^{+}24, HF07, HPvW94, IIH^{+}13, Joe13, Joh04, JLMV06, KHK^{+}94, Kee10, Kub84, KS94, Lar11, Lea96a, LIS^{+}19, LCB^{+}23, LHM23, MFL08, Ma23, MN97, MD18, Mol18b, MML^{+}15, MJM^{+}06, MM14, NBG^{+}94, NPB^{+}18, Rao97, RV83, Ric15, RET^{+}94, Ros94, RN94, Sim96c, Sim97, Sto21, TLCC17, TBP01, UZM^{+}23, WLC^{+}12, WSK^{+}00, Wil86, XAJK^{+}15, ZNMT15, vD05, Ano14a]. Researcher [Ros21]. Researchers [Hay96c]. Resections [DPL^{+}11]. Reshaping [ZJH^{+}18]. Resident [Sin19d]. residues [MV93]. Resistant [AXP07]. Resists [MD99e]. Resolution [ASS^{+}21, CCC^{+}07, CXPG19, FJP02, GJM18, JAPF07, KKV^{+}02, MQR99, SFF^{+}00, SW85, SO95, Sim95b]. Respectable [WF02]. Respiration [KLR12]. Response [Han00a, MMN^{+}21, SKMY09, WSK^{+}19, YMK^{+}12]. Responsive [UTMI11, KF94, Ros97, RDDT97, WWG97]. Restoration [CCG^{+}04, DNBC16, dOGDM^{+}19]. Resultants [WG95a, WG95b]. Results [AM84, Gla97c, WAH^{+}21]. Retargeting [LLL^{+}13, SLNG07]. Rethinking [Mue18]. Retinal [MKL97]. Retrieval [BKSS07, HEW17, RSE99, SS17, XZZL19, YL11]. retrieved [Car96]. Retrofitting [IBH^{+}09]. Return [Bli89e, SW85]. Reuse [AZSW22, Ano20r]. Revelations [WNAJ92]. Review [BFR21, ENDAD13, LS95, PD21, MD99e]. Reviewer [Ano12a]. Reviewers [Ano03b, Ano04g, Ano06b, Ano07a, Ano08c, Ano09a, Ano14-33, Ano15a]. Reviews [TM05]. Revisited [Abo00c, Fig00]. Revisiting [RDDY07]. Revisualizing [Sin09e]. Revolution [Ano15-45, Gol83a, KNR93, Pie87b, PT03, Ric85, Sin05a]. Revolutionizing [Han01]. RF [Cro96]. RGB [Gla89, SLC24]. RGB-D [SLC24]. RGBD [WW24]. Rheumatoid [ZCBM14]. Rhinestone [Iga11]. Rib [KW10]. Ribarsky [Ebe17]. Rich [And98, CE98, GLC11, Hoe12, JLJ^{+}10, RLCI24]. richer [Mas18]. Rides [SS01]. Ridge [PSX13, Sim95a]. Right [Sin12b, Sin18e, Bli95c]. Rigid [Bar95, TW88]. Ring [CPN09, VS14]. Ring-Like [CPN09]. Ringing [RV10]. rings [Dia94]. Ripped [SN86]. Risk [Ano16-43]. Risk-Based [Ano16-43]. River [MD99e]. RNNbow [CPM^{+}18]. RNNs [CPM^{+}18]. Road [FR00, RV07, Sch99, SAB^{+}23, Whi95a]. RoboJockey [YSS^{+}16]. Robot [Abo98a, DLEF15, LJ15, LMA^{+}17, MP89, RM89]. Robot-Assisted [LMA^{+}17]. Robotic [Hay95b, II15]. Robotics [Tac98]. Robots [GPA^{+}17, YSS^{+}16]. Robust [HHK89, LSQ^{+}12, Ovi03]. Robustness [Mil99a]. Rock [Abo00c, Ano14p, Ano14-30, Ano14-31, Ano14-32, Ano15-35, Ano15-34, Ano15-33, Ano15-36, Ano16-42, Ano16-40, Ano16-41, Ano16-43, Bar00, Sin05b]. Role [BABB^{+}21, DHC^{+}22, Gai04, MRB19, Ska86, WCE^{+}17, GB96b]. Roll [Abo00c]. Romanesque [Kri03]. Rooke [Wor96a]. Room [BK06]. Ropes [Bar97b]. Rosenblum [Tau12a]. rotating [Kor87]. Rotation [Fil84, SSC^{+}11]. Rotation-Invariant [SSC^{+}11]. Rotational [ABS16]. Rotations [Lee08, Tau11a]. Round [JMOG99]. rounding [Chi87]. Routing [FR00]. ROV [Agb95]. ROV-Manipulator [Agb95]. Rover [MGC20]. RP [SBE95]. Rubric [BAB^{+}23a]. Rugby [LC16]. Rule [BF93]. Rule-Based [BF93]. Run [SL00, Sim95e]. Runs [SL00].

S [Ano22-81, Ano23-84, Ano23-85, Ano23-86, Bli04c, ZPL^{+}23]. Saemangeum [KP12]. Safe [LMWC17]. Safety [LAC^{+}14]. SAGE [CG85]. Sales [Hay95b]. Salience [WLL^{+}16]. Saliency [DV21, SLNG07]. Sally [WCS18]. Salon [Wea98a]. Salt [MMGG12]. Same [Bli99a, BDG^{+}99, Tay02]. sample [Ano98j]. Sampled [WK94]. Sampling [AFYC03, KVHD17, LW13, PBR^{+}14]. Samurai [Kri03]. Sankey [MMMM21]. SAPPHIRE [Mye84a]. Satellite [GJM18, LKS96]. Satellites [Lea96d]. Saturation [Sin12f]. Say [Gla99e, SS20].

Scalable [CRF^{+}21, IM98, KB14, KVHD17, LCC^{+}00, MWY^{+}15, SZMJ12, WPLM01]. Scalar [GHQ04a]. Scalar-Function-Driven [GHQ04a]. Scale [ABM^{+}01, Ahr22, AJK15, AGP^{+}22, Bal98a, DSW07, EHBPG04, EWF02, HYL^{+}22, HYN03, HLB^{+}07, JBH^{+}10, Ma01, MP01, Ma09, MKFN87, MSWT14, MRBW12, RL07, RW99, SBMK20, SBSW96, VBHS11, WAB^{+}05, WYM10, War04, WSP12, WSJ^{+}12, YWG^{+}10, vOW15, Sim93b]. Scales [LSC^{+}17, YBL^{+}12, LH92]. Scaling [BNA^{+}22, CPA^{+}10, GE98, LLL^{+}13, AES93]. Scan [BE11, DW87, GPR94, Max90, Pav85, RB93, Rob89, SBMK20, PB87]. Scan-Line [DW87, Max90, Rob89, PB87]. Scandinavian [Oia82]. Scanline [Cro87a]. Scanned [CPCS08, TH17]. Scanner [BFB^{+}98, Sch00, VPB^{+}91]. Scanners [SPM10]. Scanning [HGMT17, PLS^{+}16, Paq96, Roc87]. Scape [HBG04]. Scary [SR10]. Scattered [BMC84, HLE04, Nie93b, RM95, NFHL91, Tre95]. Scattering [BHPB^{+}12, ERS13, GJ13, HG13, SKP09]. Scatterplot [Koz10]. Scatterplots [FH19, FH21, XLJ^{+}21]. Scenarios [LSC^{+}17]. Scene [Gla01a, MD00f, Sow00]. Scenes [Enc14c, IKMF12, JLW08, MWM87, NTH02, NCO03, SGC21, TS22, WJP22, ZSK^{+}21]. Scheduling [AAG^{+}22, FSP00, Ste18]. Schemata [Woo85]. Schematic [TO83]. Scheme [ABS^{+}05, LMWC17]. schemes [GD93]. School [AJD13, Ano16f, BL14, BNA^{+}22, CRA^{+}18, JMC^{+}02, ZD98]. School-Based [BNA^{+}22]. Schoolers [DASF13]. Schools [CNL00]. Schwartz [Sin12f]. Schwarz [ODW15]. Science [AAW17, Ano20r, Ano21a, Ano21b, Ano21c, Ano22c, Ano22d, Ano22e, Ano23b, Ano23c, Ano23d, Ano24j, Ano24a, CS15b, CS19, CBGH19, CKJC12, DiS14, Gai17, Han14, Ker13, MB05, MGC20, PL19, RBS21, SM91, Sin08a, Sin13b, Sin15f, Sin17e, VCS19, XDJ23, Yea97b, YLT18, vOW15]. Sciences [GS87a]. Scientific [AHP^{+}10, Ahr15, Ahr22, Ano87, BHS^{+}24, Cam16, CJS^{+}19, Enc16b, GV81, GGS^{+}18, Gro83, HB16, Joh04, Kee10, KJA^{+}18, LPLN04, LZF^{+}14, LCC^{+}96, Ma00, MLF^{+}12, MAB05, NVC92, Rei23, RD90, Rhy00, Rhy03, Ros94, SKR18, SMG^{+}21, TBO^{+}20, UFK^{+}89, UZM^{+}23, WTPHK09, WFL^{+}00, vDFL^{+}00, NBG^{+}94]. Scientist [PA18]. SciVis [KS14]. Score [SML^{+}09, Wor93]. Scores [PBV16]. scouting [Woo91b]. Scratches [MD99e]. Screen [BEK^{+}03, Cro87a, ERS13, Lea96a, WGZ^{+}18, Bli96c]. Screen-Area [Cro87a]. Screen-Space [ERS13]. Screening [SOM^{+}10]. Screens [EBBS00, RSP15]. ScrewBender [PR08]. Scroll [Bar00]. Sculpted [MCS22]. Sculpting [MFM^{+}16]. Sculpture [Abo98b, WZL^{+}21, McG93]. Sculptured [JDH^{+}89, Mil86]. Sculptures [Kd21, Nea88]. Sea [AFYC03, VS14]. Seafloor [LSB21]. SeaMaster [Agb95]. Seamless [FLMZ97]. Seamlessly [BKP01]. Search [Ano14c, Ano15d, Ano15e, Ano16g, Gla03e, HE02, SJSK19, Won08]. Seashell [Pic89]. seashells [IP92]. Second [CLP^{+}14, FM82, NPH^{+}09, DH93, Whi87]. Second-Degree [FM82]. second-order [DH93]. Secondary [OZH00]. Secrets [How00, HMRS06, Sin05e]. sections [MG92, SK91]. Secure [Ano16-39, Ano17-27]. Security [Ano14t, Ano16-43, Ano19i, Ano19j, Ano19-52, Ano21l, Ano21m, Ano22r, Ano22-72, Ano22-73, Ano22-74, Ano23n, Ano23o, Ano23-74, Ano23-75, Ano23-76, Ano23-77, BNP99, CAG^{+}06, Kro02, RXB09]. Sedbon [SCS22]. See [DBDH05, Gla99e, LSB18, Sim95e, WPO96]. Seeds [Sin16f]. Seeing [HH08, Kro04, MGC20, SH06, SP06, ZK10]. Seek [Ano23-63]. Seen [TYK01]. Sees [Sin13e]. Segment [Sin03c]. Segmentation [BWR24, KTD10, KS20, LFT19, LLY23, TS22, YNH04]. Segmenting [Sen99]. Seismic [WL88, Sim95b]. selected [AMS91, Pea88]. Selecting [BH81]. Selection [AA09, DCCE19, SVT10, Wor96a]. Selections [FH21]. Selective [Ano92i, Ano92j]. Selectively [AYS98]. Self [CR94, CS07, Dom12, LC16, RA06, USW17]. Self-Assembly [RA06]. Self-Evaluation [USW17]. Self-Motivated [Dom12]. Self-Occluding [CS07]. Self-Organizing [LC16]. Self-Similar [CR94]. Selfer [CLms15]. Selfie [CBC^{+}20]. Selikoff [SCS23]. Semantic [AVV20, End14, ECNZ15, ITW01, LLY23, LHZ^{+}21, LFPS01, MZC^{+}21, MSH^{+}08, PvW07]. Semantics [CRB^{+}02, GPW^{+}19, Spa16b]. Semester [Dom12]. Semiconductor [BEG92]. Sensation [LO05]. Sensation-Preserving [LO05]. Sense [Ozt15, SHY^{+}23, Sra23]. Sensed [MGA^{+}21]. Sensemaking [FJKTX19, XAJK^{+}15, XSN^{+}20]. Senses [Lec17]. Sensing [CML^{+}23, PBVH15, SWD^{+}98]. Sensing-Based [CML^{+}23]. Sensor [KVM^{+}05, KB14]. Sensors [DL15, Fox05, Mil12]. Sentiment [BFR21]. Seoul [Cab89]. Separated [IK15]. Sequence [DD19, PP95]. Sequence-to-Sequence [DD19]. Sequences [MKKJ96, OSRW97, QNT02, RO86, Rod14, War88]. Sequencing [Pot96]. Séquin [Abo98b]. Séquin-Collins [Abo98b]. Seri [Gha12]. Serial [Sin10d]. Series [LWY^{+}20, XHC^{+}17]. Serious [BRD^{+}20, BES09, LA20b, LFMGPG23, MTN18, RdCFP^{+}20, RRV^{+}15, ZED^{+}09]. Service [Ano20-51, Ano21-66, Ano21-67, Ano21-68, Ano23-64, Ano23-65, HB06, MG14]. Services [Ano19-39, Ano19-40, Fry83, Hay96a]. Set [Arb90, CB86, FA85, GHQ04a, GHQ04b, HHK89, LBD^{+}92, MSC89, PNPS03, PRdJ07, SM84, Woo88, Car87, Lea97c]. Set-Design [PNPS03]. Set-Theoretic [LBD^{+}92, SM84, Woo88]. Sets [BDH^{+}89, BHP07, CR94, CW88, DP00, FS86, JBH^{+}10, KPSN04, KMM^{+}01, KNK99, Kri01, KÇC^{+}01, MP01, OCS20, Sen99, TBB^{+}23, Whi04, WFL^{+}00, CW87, RN94]. Setting [Tit07]. Settlement [Hay95b]. Settling [SKN^{+}15]. Seven [HF07]. Seymour [Ano17-31]. SFF [YSW^{+}23]. SFF-Siam [YSW^{+}23]. Shaded [CE83, CG84, SW83]. Shader [AS11, RBE08]. Shader-Based [AS11]. Shaders [Bai09, Bai11, Bai13, CJS^{+}19]. Shading [CHRU85, MB94, MHC98, YYiTI92, NINT89]. Shadow [Ber86, Del98c, Gha12, HG86, WD15b, WD15c, WPF90, Bli88d, Woo93]. Shadowing [Rob89]. Shadows [AKFPW20, BB84, MTC07]. Shanshui [Shi17]. Shape [Ale21, Bli06a, Bre97, CAP^{+}05, CP89, CRC^{+}02, DPL23, DMR83, FM89a, Gla97g, HZB92, HH01, HEW17, Jos11, KS07, LFOI15, LZR^{+}21, Mac97a, MSF07, McL83, MZL^{+}21, PPK05, RMP^{+}16, RB00, RSE99, RP95, SS17, SR97, Sch17, SR95, SHY^{+}23, Spa16b, Wil87c, WZQ^{+}19, YT84, GU86, SG90, Spa16b]. Shape-Based [RSE99, HZB92]. Shape-Preserving [DMR83, FM89a, McL83, MZL^{+}21]. Shapes [Cho95, CMPA18, DL22, FB81, Jac15, MGO^{+}16, Sar90, SPW84, GJ93, Her86, KS94, Pie87c, SRS91]. Shaping [AAB^{+}22, Cha98a, YSWW08]. Share [Ano19-53, Ano20-63, PB94]. Shared [DDdSC11, LRIC15, LA20a, LFFB20, MBSD20]. Shared-Memory [DDdSC11]. Sharing [HLB^{+}07, JMT^{+}01]. Sharpening [SP15]. Sharper [Bal97]. Shedding [Sin02c]. Sheep [CGMB20]. Shell [Men94, UO93]. Shelves [ASP^{+}21]. sheriff [Sim93a]. Shifting [SHY^{+}23]. Shiny [Bli00a]. Ship [Gos90, RSR83b, VS07]. Shipboard [RDO^{+}96]. Shiphulls [RSR83a]. Shipped [MD99f]. Shoe [Fox05]. Shoe-Mounted [Fox05]. Shoes [KRB11]. Shop [CPK06]. Short [And98, Ger98, HB97, Pic89, SBS10, Wri98]. Short-Term [SBS10]. Shortcut [KKG^{+}13]. Shot [CWWW21]. Show [vHdP94]. Showcase [BFSE01]. Showing [Sin01c]. Shows [Hay95b, YL09]. Shrinking [Kas13]. Shut [Gla99a]. Shuttle [Mil03]. Siam [YSW^{+}23]. Siamese [YSW^{+}23]. Sick [Pot98]. Sided [HT98, Sch93a]. Sides [Hag90c]. Sidney [Ano17-31]. SIGGRAPH [Ano10c, Bli87, Kas12a, KWJ23, Sin23e, Bli95d, BG96, Hay96a, SP04b]. SIGGRAPH2015 [Ano15-37, Ano15-38]. Sights [MFG^{+}17]. sign [Ker91]. Signals [KB09]. Signature [HEW17]. Signatures [WFL^{+}00]. Significance [Gla97d]. Signs [Gla97d]. Silhouette [IFH^{+}03]. Silicon [FH84b, WNAJ92, Ake89, Hay95b]. Silver [WF02]. Similar [CR94]. Similarity [FMM06, SJSK19]. Simple [Nea89c, SPMS04, SR97, Bow93]. Simplification [CBZ^{+}08, KT96, Lue01, YZM^{+}11]. Simulate [JWHW12]. Simulated [BMH98, Fet82, KP11]. Simulates [MD99d]. Simulating [AP11, HW95, IL83, KLR12, Lee99, MMGG12, SKN^{+}15, VMJT96]. Simulation [AMJ^{+}08, ACO00, ARS17, Bar95, Bar04, BDK^{+}04, Bro15, BRL06, CCP82, CEMD22, CF93, CZAA23, CdHM97, CEO^{+}93, CKP96, GA97, GRW99, GGS^{+}18, GD11, KSSL07, KL18, LTS^{+}13, Li22a, MD99f, MD99e, MD99d, MSA^{+}94, MWY^{+}15, MSB^{+}06, NPH^{+}09, PFB06, PD09, PVML96, RM89, RFA^{+}22, Rho97b, RBS21, Sch81, SBD^{+}17, SS19, SSB06, SOM^{+}10, Til83b, VS14, Web08, Yan85, CG94, Gre91b]. Simulation-Based [GA97, SOM^{+}10]. Simulations [AHP^{+}10, Ahr15, CPN09, CCC^{+}07, CKJC12, GCS^{+}12, GD13, HH08, KSRF11, Kri05, KHSE98, LHFL20, MA98, MTN18, MJC08, OZH00, RLV^{+}16, SRE98, SWS07, SSH^{+}98, SH03, SBHM97, SW22, TCH^{+}96, TCE^{+}02, WYG^{+}12, Wil82, YIÇ09, YWG^{+}10, Gre91a]. Simulator [Agb95, AP10, CZY^{+}13, EHA^{+}14, LGW^{+}12, NCQ^{+}11, WMB85, SAB^{+}23]. Simulators [BSHW07, LGK^{+}06, LGM13, SB06, ZMR^{+}88]. Single [CWWW21, CC20, CXPG19, DNBC16, dOGDM^{+}19, GGV92, IKMF12, TT00]. Single-Chip [GGV92]. Single-Shot [CWWW21]. Site [KJA^{+}18]. Site-Specific [KJA^{+}18]. Sites [ATS^{+}03, EHBPG04, SPM10, VIK^{+}02]. Sitting [ZR24]. Situ [CBGH19, EI17, Ma09, Mor16, PA18, RLV^{+}16, YWG^{+}10, HYL^{+}22]. Situated [BHM^{+}22, JPSM13, SDM22]. Situation [Gla97e, SLC^{+}23]. Situational [FGW00, RDDT97]. Situations [ADZK^{+}21]. Six [WJ88]. Six-Dimensional [WJ88]. Sixth [Ano20a]. Size [Sto12]. Skateboard [GFW^{+}16]. Skeleton [SR95, WWL^{+}21]. Skeleton-Based [WWL^{+}21]. Skeletons [WZL^{+}21]. Sketch [ERH^{+}11, FH19, FH21, IZ07, KS07, LFT19, NF11, SS17, SD07, She22, TS22, TWB^{+}07, ZLMY19, SC12]. Sketch-Based [ERH^{+}11, FH19, FH21, IZ07, KS07, SS17, TWB^{+}07]. Sketch-n-Stretch [SC12]. Sketched [KS20]. Sketches [LBW^{+}17, OSJ11]. Sketching [CS07, CW09, LaV07, LMZ11, SC12]. Sketchpad [FPJ04]. Skill [Ano18-35, RSL^{+}19, SKW^{+}12]. SkillChoice [Ano18-36]. Skills [Ano20w, Ano21-32, Ano21-33, CNC^{+}05, EM09, Lok06, PPVH23, SFD16, SG11]. Skin [AYS98, JWSG10, KKTI04]. Skinny [BvRS^{+}11]. Skull [DNS^{+}15]. Sky [SBRCD17]. Skylight [Wil13]. Skyline [MD99f]. Slave [Kd21]. Slide [Fry83, Hop83, TR83, Wri83]. Slide-Making [TR83]. Slim [MD00f]. Slim-Line [MD00f]. Slime [Sin22c]. Slow [CS15c]. SM [MD99f]. Small [MTS^{+}20, TSF^{+}98, WAB^{+}13]. Small-Animal [WAB^{+}13]. smallest [Ano97i]. Smart [AD05, AGP^{+}22, CML^{+}23, LTJ18, MD00f, SKS96a, SBC15, TWQ^{+}23, HGK^{+}04]. Smart-Its [HGK^{+}04]. Smarter [Abo99a]. Smartphone [CBC^{+}20, MGA^{+}21]. Smartphone-Sensed [MGA^{+}21]. SmartTouch [KKTI04]. Smells [MFG^{+}17]. Smile [Han01]. Smiles [She05]. Smoke [FMMS03, MWY^{+}15, ZWWS19]. Smooth [Her24, Ste84, VBW94]. Smoothing [PR08]. snags [Bli93a]. Snapshot [PIBV11]. Snapshots [Ano10d]. Snow [SF24, VB24]. Snowstorm [Bli98a]. SNR [CWWW21]. Soap [Gla00c, Gla00d]. Soapbox [Whi87]. SoC [WSKY09]. Soccer [PPA14, SJB^{+}16, SJSK19]. Social [Ano18x, Ano18-41, CLDW16, CML^{+}23, EM09, FCS22, KB09, LGZ04, LWM14, PIBV11, Sch08b, WLL^{+}09, WMFM13]. Society [Ano20-51, Ano20-50, Ano21-67, Ano21-68, Ano21-63, Ano21-64, Ano21-65, Ano22-51, Ano22-60, Ano22-61, Ano22-52, Ano22-62, Ano22-53, Ano22-63, Ano22-54, Ano22-65, Ano23-62, Ano23-65, Ano24n, Ano24q, Ano24o, Ano96i, Ano97h, Ano14s, Ano14q, Ano14p, Ano14r, Ano15w, Ano15y, Ano15x, Ano16y, Ano16x, Ano16z, Ano16w, Ano16-27, Ano17q, Ano18s, Ano18u, Ano18t, Ano18x, Ano18v, Ano18w, Ano18r, Ano18y, Ano18-28, Ano19-30, Ano19-31, Ano19-32, Ano19-35, Ano19-36, Ano19-34, Ano19-33, Ano19-38, Ano19-44, Ano20-47, Ano20-48, Ano20-43, Ano20-49, Ano20-44, Ano20-38, Ano20-45, Ano20-41, Ano20-39, Ano20-42, Ano20-46, Ano20-40, Ano20f, Ano21-55, Ano21-43, Ano21-66, Ano21-59, Ano21-44, Ano21-45, Ano21-56, Ano21-46, Ano21-47, Ano21-48, Ano21-49, Ano21-50, Ano21-51, Ano21-52, Ano21-60, Ano21-57, Ano21-53]. Society [Ano21-61, Ano21-58, Ano21-54, Ano21-62, Ano22-38, Ano22-47, Ano22-55, Ano22-43, Ano22-44, Ano22-48, Ano22-45, Ano22-49, Ano22-46, Ano22-56, Ano22-58, Ano22-64, Ano22-57, Ano22-50, Ano22-59, Ano23-28, Ano23-29, Ano23-30, Ano23-31, Ano23-32, Ano23-50, Ano23-63, Ano23-64, Ano23-56, Ano23-51, Ano23-57, Ano23-41, Ano23-52, Ano23-58, Ano23-42, Ano23-46, Ano23-59, Ano23-47, Ano23-53, Ano23-43, Ano23-48, Ano23-54, Ano23-60, Ano23-44, Ano23-49, Ano23-55, Ano23-61, Ano23-45, Ano24h, Ano24i, Ano24p, Ano24t, Ano24r, Ano24s, Ano24u, Ano20-32, Ano20-30, Ano20-31, Ano21-39, Ano21-40, Ano21-41, Ano21-42, Ano22-37, Ano22-39, Ano22-40, Ano22-41, Ano22-42, Ano22o, Ano24m]. Sociotechnical [MJS20]. Soft [BB84, MTC07, NL02, PFB06, SG11, Sin09f, WWM87]. Soft-Object [NL02]. softkeys [BM88]. Software [Ano88a, Ano94n, Ano23-84, Ano23-85, Ano23-86, CPA^{+}10, CE98, Dil96, Dil97, Dil98b, Dil99, Dil01, Enc04, GAW86, HH82, Hay95b, Hay97, JXW^{+}08, KLL^{+}15, Lar10, Lav83, Laz87, LOB08, LCG83, MP01, MD96c, MD99f, MD99e, MD99d, MD00f, Mac03, MOP89, MRBW12, SR13, War81b, War81a, War85, Wea98a, Whi95d, vD05, Bri90]. Soil [AM84, MV93]. Solar [Del97, FML23, Gla96e, Gla96b, Wil13, SVB97]. Solar-Radiance [Wil13]. Solid [Bar95, CS85, Cas87, CK85b, DMP93, Hay96b, Hil82, KPTS89, LBD^{+}92, LJJ^{+}17, MS82, MMZ94, Mil86, Mil88b, Mil89, MP89, Mye82, PCOS10, Pra84a, RV82, RV83, RR92, RR86b, SC89, Til83b, TSUC86, Tur88, TR94, Wei85, WFOP85, Woo88, WKS86, WWM87, YT85, Cam91, CR95, GT93, KKM96, SG90]. Solid-Texture [PCOS10]. Solids [AESF21, Arb90, BG82, Gla96d, Gla96c, HHK89, Kim84, LW13, Men94, WW86, Chi87, CH95]. Solution [MD99f]. Solutions [FM82, MD99f, Ree84, Sad16, MD99f]. Solve [Abo99b, Bli06c, Bli06d, Bli06a, Bli06b, Bli07a, Bli07b, Bli05a]. Solver [LGW^{+}12]. Solving [GASF94, Hor84, Man94, RdCFP^{+}20, XNSW19]. Some [Coh83, HS87]. Someone [Sin23a]. Something [Bre90, Han01, Sin13f]. Sonar [CWBS99]. Sonic [SGE^{+}18]. Soon [BW00]. Sophisticated [SRS83]. Sorting [MCEF94]. Sound [Coo02b, DBJEY^{+}02, Ken20, MCGS21, RL07]. Sounds [Coo02a, MFG^{+}17]. Source [DMJ05, Mac85, RDRS10, SSM^{+}07, VG84]. Sources [Mac81, Mac83]. Southern [MFG^{+}17]. Space [ASS^{+}21, BB05, BPB14, Bli03e, Bli03a, Bli03b, Bli03c, Bli03d, Bli04a, Bli04b, Bli04c, BBS09, CC20, Cha98a, ERS13, EYY03, Gla84, GRvL^{+}23, JKM23, KCW^{+}17, KY97, Lea96b, LW13, LLR09, LZR^{+}21, LK95b, MMMM21, Mil03, Mun98, NM08, NTB^{+}20, SWDR16, SKP09, SSC^{+}11, Sim96c, Sin06e, SW22, VK17, VSLG22, WKS86, ABB^{+}93, CH95, PB87]. space-based [PB87]. space-times [ABB^{+}93]. Spacecraft [MPDR14]. SpacePlane [SBJ^{+}01]. Spaces [FZH^{+}13, Gol00, HHY20, MG04, PSE^{+}09, SPJ99, Wee02, Rob88]. Spacetime [Gla88a]. Sparse [DP00, Gie92, HWY16, KLR12, KFH^{+}18, LSQ^{+}12]. Spatial [BH16, BBDM98, CCV85, EI17, FL12, Jer18b, KPL17, LaV08, LKPM19, LFOI15, LPK05, LSC^{+}17, MSWT14, Rob89, SKJ^{+}13, SLC^{+}23, SO95, SS15, TM88, WHKM^{+}13, WTPHK09, ZYL^{+}17, dHPP10, Fan86, PFH95, WFG92]. Spatial-Mechanism [TM88]. Spatializations [FMM06]. Spatialized [PHAS20]. Spatially [RBK^{+}13, WWV^{+}10]. Spatio [DFL^{+}18]. Spatio-Temporal [DFL^{+}18]. Spatiotemporal [Bee15, LOC^{+}16]. Speaking [Agr09, Cha09, CW09, Mac09, Saw08, SPS99, Sto21]. Speaks [MSDS23]. Special [Ano03b, Ano04g, Ano06b, Ano07a, Ano08c, Ano09a, Ano14-33, Ano16f, Ano19-54, Ano20s, Ano22-29, BHH^{+}21, BD85, BCPS22, Dil98a, GGS22, LF84, LQ18, MTT86, Mar83, Rus11, Tak05]. Specific [BH11, KJA^{+}18, Tre99, Eng86]. Specification [Car83]. Specifying [DMR83]. Spectating [LRZ23]. Spectra [SFCD99]. Spectral [BCMF16, BP06, Hal99, JF99]. Spectrum [Gla89, HLW13, VS14]. Spectrum-Based [HLW13]. Speech [BHS11, SZP^{+}00, WD15a]. Speech-Driven [WD15a]. Speech/Gesture [SZP^{+}00]. Speed [DBS96, Gla05b, Hay96d, Ike84, LS84, Mil99a, Mil03, PB85]. Speeding [AM91, HS01]. Speeds [MD99f, SK02]. Sphere [Bal98b, FLNR90, Bli95a, Bli95e]. sphere.2 [Bli95b]. Spherical [FUF06, MG88a, ZGP^{+}13]. Spherical-Gradient [ZGP^{+}13]. Sphericons [Sin01b]. Spill [MMN^{+}21]. Spinal [SC21, SBHM97]. spine [MB91]. Spines [KRG^{+}22]. Spiral [Sin09d]. Spirals [WO13]. SpiroSurface [HF18]. Splatting [HLE04]. Spline [DMR83, Fre84, HN85, PT89, RSR83a, RSR83b, SR85, SB86b, WWL^{+}21, BD85, GU86, Nie86, YSSP91]. Splines [BD90, FM90, GS94, HP91, Pav85, PT03, Rie22, Til83a, Bar93, CP92, Gol86]. Splitting [AAB^{+}23, MMP^{+}07]. Spontaneous [LSR^{+}00]. Sports [BS16, BKV^{+}10, CPG^{+}16, LCB^{+}23, LRZ23, VP16, Woo15]. Sportswear [CLG^{+}16]. Spray [Pan94]. Spread [BHS^{+}24]. Spreading [Enc23b]. Spreadsheets [hCRBK98, MD99d]. Spring [HJCW06, TNTM07]. Spring-Bead [TNTM07]. Springs [Bar97b, Kan09]. Sprints [BYP21]. Ssecrett [JBH^{+}10]. Stable [JK09, TLT11]. Stage [SG11]. Staking [Mye84c]. Standard [BEHt82, Hai87, SPT^{+}15, SBM86, War85, BHK91, Lea97c, CPC99]. Standards [HILW98, LLK^{+}03, PBSB08, Pra84b, Ska86, SG86b, Wil91]. Standing [ZR24]. Star [Abo00c, SR95]. Star-Skeleton [SR95].

Stars [Ano14-30, Ano14-31, Ano14-32, Ano15-35, Ano15-34, Ano15-33, Ano15-36, Ano16-42, Ano16-40, Ano16-41, Ano16-43, Ano14p]. Start [Joe13]. Starting [HS01, Hur82]. StatCast [LOC^{+}16]. State [Ano10d, Bux05, CPAM08, CHP21, Dev13, Gel07b, KW04, OOHR14, SGE^{+}18, vL18, Mac97a]. State-Surface [KW04]. Statement [Gor96]. Static [vLBW17]. Station [SB81]. Stations [Lil81]. Statistical [CH81, UF11]. Statistics [CWB02, PS09]. Status [RV83, Ric85]. Staudhammer [Kas17]. Stay [Ano15-39, Ano18-42, Ano19-55, Bar00]. Steep [RV07]. Steering [DCCE19]. Steiner [SA85]. Stem [GRWN87, HBW^{+}24]. Stencils [II10]. Step [BB00, CE83, Gil94, GI94, JS03, LLY23, SL00, WR89]. Steps [Whi95d]. Stereo [BWP^{+}16, HM85, MVWB05, NK14, PSJ87, XNSW19]. Stereoscopic [Gro83, HBG04, JLMV06, Sch00, SWS07, Hod92]. Stereotactic [RGA^{+}83, TLL^{+}96]. Steven [Wor96a]. Stiffness [SPMK14]. Still [Bal97, BT07, KH14, Lea96e, QNT02, SKN^{+}15, XPH05]. Still-Image [XPH05]. Stills [JSH16]. Stimulating [AYS98]. Stimuli [CPAV^{+}17]. Stipple [MIA^{+}08]. Stippling [PFS03]. Stochastic [FM85, HB84, SKBF12]. Stock [NB04, Koh93]. Stone [Sin05e]. stones [DFL87]. Stop [HFZ^{+}13]. Stop-Motion [HFZ^{+}13]. Storage [Ahr15, SCH05, WW95]. Store [CSMT03]. Stories [Bli05b, LRIC15, Sin17c, ZE23]. Storm [ZCS^{+}06, BB95]. Storm-Surge [ZCS^{+}06]. Story [LRIC15, RY06, SH03]. StorySpace [WKM^{+}04]. Storytelling [BAAQ21, CM11, Kos15, MLF^{+}12, SS19, ZE23]. STRAD [FPNVHH17]. Straight [BB00, Gil94]. Straight-Line [BB00, Gil94]. Straßer [Enc15a]. Strategies [Ano15-44, DB16, SS20, UIL^{+}06]. StratomeX [TLS^{+}14]. Stream [RGL^{+}06].

Streaming [ABM^{+}01, ACL^{+}12, GTHL99, ILC10, Kri04, NCO03]. Streamlines [GHCW21, HTZ^{+}19]. Streams [ABWS05, BHAA^{+}13]. Street [Gha12]. Stress [KRG^{+}22, Mac09, Smi84b]. Stretch [SC12]. String [Gla99f]. Stringing [Sin05f]. Stripes [Elb01, MR87]. Stripped [Hay96d]. Stripped-Down [Hay96d]. Stroke [CKPB16, Her03, KS20, TS13]. Stroke-Based [Her03, KS20]. Structural [GSSO05, GGS^{+}18, HPX^{+}21, MSF07]. Structure [CCV85, De 89, SMG^{+}21, Woo85, Zac84]. Structured [And98, ILC10, RSK^{+}06]. Structures [AEFQ96, FZY84, GCS^{+}12, KB82, OKS^{+}17, SM95, STW^{+}08, SB02, Wei85, YKFT84, Ala92, NB94, SW88]. STSRNet [ASS^{+}21]. Student [Ano18-37, Ano18-38, Ano18-43, DSP^{+}14]. Students [AJK15, BL14, Ker13, LSB18, Mei85, Per21, SP06, YY16]. Studierstube [Ano98i]. Studies [DFQ12, Her23, KHI^{+}03, MP99, PS09, SDS^{+}11, She82]. Studio [CBC^{+}20]. Studios [FY97]. Study [AY16, AOH^{+}14, ENDAD13, KM00, KND^{+}18, LCG83, LWM14, MKH20, PSJ87, PCH^{+}07, SF20, SC21, SD83, WXK22, WBA^{+}96, ŽLK22a, Cra91, Til81]. Studying [ADZK^{+}21, DEN13, RRV^{+}15, SWDR16, ZY18]. Stuff [Kas12b]. Style [GUFG15, HHKK17, MQM22, Sch21, ZLMY19]. Styles [LC16, WC91]. Styling [KS07]. Stylized [iAH03]. Stylus [ABS16]. Subatomic [DPG18, Sin02a]. Subdivision [ABS^{+}05, Gla84, KS98, RV10, WS04, Car87, CH95]. Sublinear [UF11]. Submarine [CBD10]. Subpixel [LKV90]. Subpixelic [Bli91d]. SUBPLEX [YCB^{+}22]. Subpopulation [YCB^{+}22]. Subset [MTS^{+}20]. Substance [Lea97d]. substitution [Gla92]. Substrate [JBCS23]. Subsurface [FPBB08, SKP09]. Subtractions [Tau10b]. Subtypes [TLS^{+}14]. Success [HMRS06]. Successes [MC85]. Successful [TR83]. Suite [MD99d].

Sulci [ZTT99]. Summarization [CLDW16, TSLX12]. Summarize [Rod14]. Summary [Bro82, MJM^{+}06, RV82]. Summer [BL14, Mol18a]. Sun [Gla96e, Gla96b, MD99f]. Sunburst [Rod14]. SunScreen [FML23]. Super [ASS^{+}21, Bli93c, FJP02, JAPF07]. Super-Resolution [ASS^{+}21, FJP02, JAPF07]. Supercomputer [Bor89, IS95]. Supercomputing [BvRS^{+}11, GS87a, JLC98]. Superfaces [KT96]. Superman [Koz12]. Supernovae [Tur03]. Superquadric [FB81]. Superquadrics [Bar81a]. supersampling [Kla93]. Superworkstation [Ake89]. Superworkstations [SvNG89]. Supplied [BC19]. Support [AZSW22, BQP^{+}15, Cam16, CG85, Cro96, DFP^{+}11, FAZH17, JOG^{+}17, LGZ04, MGP^{+}04, MD99d, MD00f, MN84, SEAKC21, VS07, YMDO93, Sim93a]. Supported [PPVH23]. Supporting [AAG^{+}22, BWA96, Fis21, HBG04, KWK21, LD97, LJ96, Nap05, SvW09, WSP^{+}20, WKM^{+}04]. Supports [YIÇ09]. Supposed [Sim96a]. Supreme [Whi95d, Hay96a]. Surface [Ale21, ADBW96, BD85, BFH86, BFP86, BHG23, Bro15, Cas87, CM92, CHRU85, CT04a, De 89, DY09, DMP93, FM89a, GF86, HHS^{+}92, KVM^{+}05, KSD04, KW04, KRG^{+}22, LLR09, LZS04, LMD92, MD00f, Mil88b, MHC98, MR87, MIEL00, NK14, NTM^{+}01, NB93, OKM84, Pat93, PT90, PT94, PT00, Rob89, Sar82, SR85, SA85, SM95, SKK91, SH07, Ste84, Tay02, Til83a, US17, VS14, WJP22, Wei85, Woe83, Wor83, XZWH14, aCJB^{+}12, van93, ABB90, Gor83, Nie86, PT92, VPB^{+}91]. Surface-Mounted [US17]. Surface-to-surface [Pat93]. surface/surface [ABB90]. Surfaces [BMC84, Bak88, Bar93, BWWM12, FS89, FB89, Gol83b, GHQ04a, GHQ04b, HW90, HAG83, HN85, IIH^{+}13, JDH^{+}89, KS01, Kim84, LF84, Lev88, Max83, MG88a, Men94, Mil86, PO84, Pie87b, PT03, RSR83a, RSR83b, RV10, REF^{+}06, Sar90, SM86, Sed90a, Sed90b, SGC21, SM04, SB03b, SB86b, Tau94, TF97, TUH^{+}90, VBW94, WS90a, WS04, Whi04, WW07, YKR^{+}84, Bar83, Chi87, GG92, GU86, Hof93, MV93, Pic90, PB87, Roc87, Sch93a]. SurfCuit [US17]. Surge [ZCS^{+}06]. Surgery [AKME^{+}90, BSHW07, CF93, CEO^{+}93, EHA^{+}14, LMA^{+}17, MSB^{+}06, QCP^{+}10, RBBS06, RR96, She05, SSS00, TLL^{+}96, WKHL96]. Surgical [BDK^{+}04, BTM17, BMS^{+}13, BTTU84, PFB06, PVML96, SKW^{+}12, ZYJ^{+}21]. Surname [MNC^{+}22]. Surrealism [Sin03f]. Surrendering [Sin02d]. Surveillance [BKM16, MHR^{+}09, dHPP10]. Survey [Ama87, Ano84, AM84, BCF^{+}08, CEMTT98, Doo82, Far92, Hec86b, Her03, Lue01, Pie91, PCOS10, SZ94, WB00b, WPF90, YWR09, ZGC14, vBE12, Bar83, Car96]. Surveying [ZE23]. Sushi [OBG^{+}23]. Sustainability [CBS^{+}22]. Sustainable [Ano16-29, Ano18b, Ano19k, Ano19l, Ano20h, Ano21t, Ano21r, Ano21s, Ano22u, Ano22v, Ano23p, BRM10, FAZH17]. Sutherland [MD99d]. SVGA [MD99d]. swap [MPHK94]. Sweep [MW82, Vos85]. Sweep-To-CSG [Vos85]. sweeping [Coq87]. Swept [WW86]. Swimming [FKC^{+}05]. Swordplay [KIQ^{+}06]. Symbiosis [Arn14a]. Symbol [KS20]. Symbolic [lLK84, PBVH15, BHMN95]. Symphonies [Tit07]. Symposium [Mac97a, PGS^{+}08]. Sync [WD15a]. Synced [RB93]. Synchronization [SAHK13]. Synchronized [dOYG04]. Syndrome [RMR^{+}19]. Syndromic [MHR^{+}09]. Syndromic-Surveillance [MHR^{+}09]. Synergistic [LPLN04]. Synesthesia [Sin01a]. Synopsis [Ano87]. Synoptic [SFL^{+}04, WSL^{+}20]. Syntactic [AVV20]. Syntax [ITW01]. Synthesis [BKRE88, CGIB86, CXPG19, DLN05, DY07, ERH^{+}11, HG83, JF99, LBS^{+}03, LSQ^{+}12, MTT87a, MH04, MWM87, OK09, Pic89, PCOS10, RL07, SMDBR22, WH97, van96]. Synthesize [Zhu19]. Synthesized [MF89, ZWWS19]. Synthesizer [Gla97g]. Synthesizing [DBJEY^{+}02, RPM17, YLR17]. Synthetic [BKS19, Gla00e, Gla00f, Lea96b, Ler82, MHR^{+}09, MTT87b, Mus99, Pic00, SMP^{+}22, SBSW96, MTT91]. System [Agb95, AM04, ACL^{+}12, AZSW22, AGP^{+}22, ASP^{+}86, BQP^{+}15, BBK82, BSM^{+}15, BM86, BH81, BWHR99, Cav97, CCC18, CG85, CDBL83, CPAV^{+}17, Del97, DR83, EWS96, EYY03, Eng86, Eng20, FAZH17, FPT^{+}82, FPI84, FTI86, Gos90, HF18, HFZ^{+}13, HJCW06, Hur82, II22, JSDL21, JCM07, JRM17, KCW^{+}17, KND^{+}18, KBHT06, KM88, Kub84, Ler82, LCC^{+}00, LD08, LAC^{+}14, MZC^{+}21, MMN^{+}21, MC85, MD99f, MD00f, MTT83, MTTF85, MP89, NK85, PIBV11, PB94, RRR^{+}16, RGA^{+}83, RC85, RFN89, SI94, SB03a, SHMN09, Ski81, SMMS24, SSO^{+}12, SVB97, Tak85, Tho84, UMO82, VMJT96, WKB11, Wei84, YST84, ZY18, ZCS^{+}06, ZGL^{+}17, ZYJ^{+}21, Ano98j, GG87, GS88, Pap96, Phi91, Kan88, PH98, PH99, Pie06, PM86, UFK^{+}89, YMDO93]. Systems [Ano88a, BN95, Bee82, BMC04, BBC^{+}20, Car83, DP82, EWF02, FML23, Fyo97, GZS88, Gre84, HJS00, II15, JW16, KFH^{+}18, KS14, KPTS89, Liv05, MD99f, Man94, OZH15, PLS^{+}16, PM82, PCJ^{+}16, Ree84, SCS22, SBC15, Smi83, SSMH03, Tal97, TR83, Til83b, Van10, WFOP85, Bli95b, CR95, Har96, Kal87, Ker91, MF89, Atw84, MD99e, Ano22b, Ano22p].

T. [Use93]. Table [Ano11c, Ano13g, Ano14-34, Ano14-35, Ano14-36, Ano14-37, Ano14-38, Ano15-40, Ano15-41, Ano15-42, Ano16-44, Ano16-45, Ano16-46, Ano16-47, Ano16-48, Ano17-32, Ano17-33, Ano17-34, Ano17-35, Ano17-36, Ano17-37, Ano18-44, Ano18-45, Ano18-46, Ano18-47, Ano18-48, Ano18-49, Ano19-56, Ano19-57, Ano19-58, Ano19-59, Ano19-60, Ano20-64, Ano20-65, Ano20-66, Ano20-67, Ano20-68, Ano20-69, Ano21-74, Ano21-75, Ano21-76, Ano21-77, Ano21-78, Ano21-79, Ano22-75, Ano22-76, Ano22-77, Ano22-78, Ano22-79, Ano22-80, Ano23-78, Ano23-79, Ano23-80, Ano23-81, Ano23-82, Ano23-83, Ano24x, Ano24y, BRL06, FK87, KIN^{+}05]. Tables [HAG83, MRD06, Bli90b]. Tablet [HHLL14]. Tabletop [Gel06, kHHBEH13, KIN^{+}05, MCP^{+}06, PMI06, SCH05]. Tabletops [HF18, SC06, SRF^{+}06]. Tablets [AL02]. Tabular [JSDL21]. Tackling [SJSK19]. Tactile [AYS98, DMSJ23, IWF97, NTM^{+}01]. Take [Ano14-39, Ano15-43, Ano15-44, Ano18-50, Mac97a]. Taken [KMH02]. Takes [Sch99]. Taking [Cha97b, GWF98, Sin17d, WGI15]. Tale [BvRS^{+}11, Bli03d]. Talk [Off20]. Talking [LTB22]. Tangencies [Fre86]. tangent [MG92]. Tangible [CW09, GSSO05, HCC05, Mas98, MMR^{+}06]. Tank [LLZ13]. Tapestry [Kos15]. Tapping [SH03]. Taps [PBVH15]. Target [Sch08b, Bal98c]. Targets [Abo98a, MD99d]. Task [BWD^{+}19, KCMK20, SO95, Tre99]. Task-Specific [Tre99]. Tasks [AZSW22, BWP^{+}16, DCCE19, KCMK20, OKKP13]. Taxonomy [AGR21b, GMR^{+}23, Jer18b, Mye88]. Taxonomy-Driven [AGR21b]. TCP [Lea96d]. TDZ [MD99d]. Tea [Sin13b]. Teaching [AS11, Ang17, BHH^{+}21, BMS^{+}13, BD21, CDLS19, EM09, FCS22, FIB^{+}20, Gre84, Kon12, Lok06, MAB05, Mei85, MRBW12, PB94, RSL^{+}19, Tak05, VPVG21, Wan22, WH99, ZNMT15]. Team [Hay95b, LK95b, LCR^{+}14, VG81]. Teams [Han00a]. Teamwork [SG11]. teapot [Cro87b]. Tech [Ano15m, Ano18-29, Ano19-45, Del98a, Mac97a, MD99e, Blo86].

Technical [Bro82, Kon12, LRZ23, MPP81, Myn08]. Technique [BP06, GL82, KSD04, Lea96d, Mam89, SR97, Coq87, Fan86, Roc87]. Techniques [BSS20, Cro81, DGY07, DT81, EHBPG04, Eng97, FWC84, HH82, Hay96d, HBB20, HB16, HM85, II15, Ike84, Jer18b, KDH^{+}95a, Kas11c, KB82, Kos16, Kul09, LS95, LTD16, MHR^{+}09, Mue18, Mye87, NFMD90, NG96, PNPS03, RR86b, Sed90a, Sed90b, Vos85, Zel82, vBE12, Car87]. Technological [BAB^{+}23b, Sou04, WLL^{+}09, WW95]. Technologies [Ano15b, Ano17c, CSMT03, Eas90, EFD05, KYK^{+}24, KPL17, LLK^{+}03, Mac97a, Per16, PZD^{+}18, SP04b, Str04, WSSS23, Zuc84, Ano17b]. Technology [Agr21a, Ahr22, Ano20-70, Bal97, BHS11, Bli93c, BEG92, CS19, CKP96, Del99a, DiS14, Ebe05, Enc11, JXW^{+}08, Kas10a, Kas10b, Kas11a, Kas11b, KLL^{+}15, Lat98, Lea96a, LOB08, Ma07, MJS20, Mar07, MTN18, Ped01, PTB^{+}15, RBMS17, SBE95, SSMH03, WP87, WKM^{+}04, DMSJ23]. Technology-Based [Lat98]. Tectonics [HW18]. Tele [Bai95, JLC98, KND^{+}18]. Tele-Immersion [JLC98]. Tele-Immersive [KND^{+}18]. Tele-Manufacturing [Bai95]. Telecommunication [KNK99]. Teleconferencing [BCPK02]. Telemedicine [Lea96e]. Teleoperated [Abo98a]. Teleoperation [DLEF15, SKN^{+}05, SAHK13]. Telepresence [SSO^{+}12]. Telerobotic [Hay95b]. Telescope [LK95b]. Television [Hay96a]. Telexistence [Tac98, Tac16]. Tell [Sch17, ZE23]. Telling [LRIC15]. Template [AWM10]. Template-Based [AWM10]. templates [Ols86]. temples [Mak87]. Temporal [ABWS05, DFL^{+}18, FPNVHH17, NRH^{+}22, PBV16, SD07, SLC^{+}23, WTPHK09, XSN^{+}20]. Temporally [PIBV11]. Ten [Bli98c, Dil98b, Fol00, GSG^{+}21, KBF05, Mac98, MD00a, Nie93a].

Tennis [AY16, BRL06]. TennisViewer [JB97]. tense [Wil91]. Tension [FM89a, FM90, Smi87, FM87]. Tension-Compatible [FM89a]. Tensions [Mor16]. Tensor [Bli01a, Bli01b, Bli02a, KRG^{+}22, DH93, HPvW94]. Term [PS09, SBS10]. Terminal [DD84, Ike84]. Terminals [Mur84b, Mye84b]. Terrain [EBG^{+}10, Kla89, MBV17, Pic95, RLIB99, RWF02, Sin04c, Sin15c, VBHS11, CG94]. TerraVision [RLIB99]. TerraVista [MD99f]. Terrestrial [MBV17]. Terrorism [Del02]. Tessellations [AM84]. Test [Fu24, Zha22c]. Testbed [DB96, HG83]. Testing [Ano97i, BDG^{+}99, HG86, MPDR14, PJC24, GG87]. Texas [ASP^{+}86]. Texels [JWHW12]. Texram [SKS96a]. Text [ABWS05, ACL^{+}12, BFR21, BFK^{+}84, DiS14, DL16, GUFG15, RSE99, VT24, ŽLK22a]. Text-Based [DiS14]. Text-to-Image [VT24]. textiles [Bre96a]. Texture [Ash03, BS86, DLN05, DY07, Elb05, FPBB08, HB84, Hec86b, KMM^{+}01, KL09, MHC98, OK09, PCOS10, SK97, YM95, Bli90a]. Texture-Based [KMM^{+}01]. Textured [CBZ^{+}08]. Textures [CT04a, DG99, DBJEY^{+}02, Gla00b, IWF97, Int00, IMS03, JCM07, Lai01, MB94, OSRW97, TBP01, YSWW08]. Texturing [SKS96a, WWM87]. Thanks [Ano03b, Ano04g, Ano06b, Ano07a, Ano08c, Ano09a, Ano12a, Ano14-33, Tau12e]. Theater [AJD13, SG11, BG96]. theaters [Bal98c]. Theatre [BJC18]. Their [AMDM22, GSFT16, MKFN87, TBFGC23]. Them [KMCMT09, Wan10]. Theme [SS01]. Themes [O’S11]. Theoretic [Che08, LBD^{+}92, SM84, Woo88]. Theoretical [CGJ^{+}17]. Theory [BC87, Bra86, CR94, JS10, LZR^{+}21, Mei85, Rhy11, War88, War08, SKK91, Bli94d]. Therapeutic [WAB^{+}13]. Therapy [Lea97b, RR96]. There [Bli99b, Bli99a, Bli00b, Cha97b, DeF00, Fol00, Ken20, Rhe02, Bli89d]. Thermal [FML23]. Thermo [YNH04]. Thermo-Key [YNH04]. Thermometer [WDK17]. Thesis [Car16]. Thin [FH84b, KB14]. Thin-Film [FH84b, KB14]. Thing [LR14, RRB^{+}14]. Things [HS01, LFFB20, Sin06f]. Thinking [LFFB20, RWB^{+}14]. Third [Kas11c, Jam87]. Thomas [Ano10b, MD99e, MD99d]. Thou [MR99]. Thoughts [Bux05, Wea99b]. Thousand [Whi95c]. Threaded [MSU^{+}16]. Three [BD89, BvRS^{+}11, Bli95c, CDBL83, Gal89, Gol03, Gro83, JKM23, MTT85c, Nak84, PM86, Sin16b, Sto05, XZZL19, vHdP94, FP95]. Three-Dimensional [CDBL83, Gro83, MTT85c, Nak84, PM86, XZZL19]. Three-Point [Gal89]. thrill [Sim95d]. Throw [COM15]. Throwing [Mac97a]. Thumb [PBVH15]. Thumb-to-Finger [PBVH15]. Thunder [Gla00e, Gla00f, PS22]. ThunderPunch [KmKK^{+}18]. TI [Wor96b]. Tile [BE11]. Tile-Scan [BE11]. Tiled [CSI86, HJS00, Sto01]. Tilers [NB93]. Tiling [Cam17, Gla98a, Gla98f, Sch21]. Tilings [OF14, OZH15]. Timber [WMB85]. Timber-Marking [WMB85]. TIMCOGS [WMB85]. Time [ACL^{+}12, ASS^{+}21, AGL87, BPB09, BFW99, CS15c, CSWB01, CdHM97, CT04b, DB16, DYS^{+}07, DAG95, DL16, ERS13, EBG^{+}10, Fu24, GCS^{+}12, GMTF89, GRW99, GRB^{+}85, Gos90, GFW^{+}16, GDY^{+}11, Hag90d, HE82, Hod92, JWHW12, JWSG10, KMTM^{+}98, KVM^{+}05, KMS07, Kri05, KHHP06, SKUT^{+}10, LKS96, LMA^{+}17, LWY^{+}20, MD99f, MD99d, MYPT09, MHC98, Mil12, MMR^{+}06, MYK^{+}08, NTH02, OOHR14, OSK02, PWH04, PFS03, PA18, SSB06, STW^{+}08, Sin03c, Sin19b, SKN^{+}05, SAHK13, UF11, WYM10, Wee02, Won08, XHC^{+}17, Zha22c, Fan86, MF89, SKR18, Tac98]. Time-Dependent [STW^{+}08]. Time-multiplexed [Hod92]. Time-Oriented [DL16]. Time-Series [LWY^{+}20]. Time-Varying [GCS^{+}12, JWHW12, WYM10]. Times [Rib98, ABB^{+}93]. TimeSets [XSN^{+}20]. Tinkertoys [Bli04b]. Tinmith [Pie06]. Tissue [PFB06, SSB06]. Titanate [SKS^{+}13]. Title [GWF98]. Titles [MD99e]. Today [Ano20-61, GHR01]. Together [CS15a, Gla02f, RBMS17]. Tokyo [KR86]. tolerance [GT93]. Tolerances [GZS88]. Tomás [LCS21]. Tomography [NFMD90]. Tone [TR93, WN82]. Tool [AWM10, Ano07b, BMS^{+}13, Dwy16, FOI^{+}07, HN85, JOG^{+}17, LTB16, MD99f, MD99e, MD00f, Nea88, NRH^{+}22, PLS^{+}16, PPVH23, RRF^{+}23, SBSW96, WHR19, WLL^{+}09, XY01, ZHZ^{+}15, GG87, Ols86, SRS91]. Tool-Free [PLS^{+}16]. Toolbox [MD99f]. Toolbox/SM [MD99f]. Toolkit [Dye90, MSU^{+}16, WHSM17, Wan22]. Toolkits [DFM^{+}02]. Tools [Abo99a, Ano05e, Ano05f, Ano05g, Ano06c, Ano06d, Ano06e, Ano06f, Ano06g, Ano06h, Ano07c, Ano07d, Ano09b, Ano09c, Ano09d, Ano09e, Ano09f, Ano09g, Ano11d, Ano11e, Ano12b, Arn14b, AHDG00, BH83, CH81, Coh83, CKJC12, GSN09, Hay96d, JKKM^{+}03, JBH^{+}10, Kas08, Kas12c, KRL97, MD99f, MK07a, MK07b, MK07c, MSK04, Mil03, PM82, PJC24, SSM^{+}07, Shu23, Sim97, WAB^{+}05, Wix83, Zin08a, Zin08b, Zin08c, Zin08d, ZP95, NF91, MD99f, Zhu12]. Top [Ano94n, Ano05h, Ano20-70, Che05, Dil96, Dil97, Dil98b, Dil99, Dil01, Enc04, Fol00, Hay95b, Hib04, Joh04, MD96c, Mac96, Mac97b, Mac98, Mac99b, MD00a, Mac01, Mac02, Mac03, Mac04, Nie93a, WSJ^{+}12, XHC^{+}17, KmKK^{+}18]. Top-Viewed [XHC^{+}17]. Topic [DL16]. Topic- [DL16]. Topics [Ma23, VPVG21]. Topological [Dal19, KSY85, SM95]. Topology [FTAT00, Gün20, HSK02, HLW13, YK95, HH91]. Tornado [UMO82]. Torque [ABS16]. torso [MB91]. Torus [Iwa99]. Tosiyasu [Woz21]. ToT [Fu24]. Touch [Enc16a, GHQ04b, HdlRL^{+}13, KKTI04, Lec17, PMI06, SRF^{+}06, Sin22e, LS04b]. Touch-Based [GHQ04b, HdlRL^{+}13]. Touch-Enabled [LS04b]. Touchscreen [MD99e]. Tour [Hai06, HCL^{+}20, Nea86]. Tournament [VP16]. Toy [LMZ11]. Traced [AH95]. Tracer [Lev90]. Traces [OCS20]. Tracing [BTL90, Dia91, DSW07, FTI86, Gla84, Gla88a, GS87b, GP89, KS97, LAM01, MMS^{+}91, Mus90b, OBO^{+}23, PB85, PK87, SKBF12, SF24, SB86b, WK90, WW24, WKS86, YCK92, AMS91, BBP94, CH95, Dia94, Gre91b, Woo93]. Tracker [MD00f]. Tracking [AP11, ANS15, Bro15, Cav97, Cha88, Fox05, JB02, KW15, LTS^{+}13, LMA^{+}17, OSK02, PXC02, RLG^{+}02, SLSP22, TH14, WF02, YNA99]. Traction [KVM^{+}05]. Trade [Gla01f, Jar83, SO95]. Trading [Del99b, NB04]. Traditional [VVP^{+}11]. Traditions [Wor95]. Traffic [ADZK^{+}21, AHDG00, CCM15, CM16, FR00, GLRK06, GSFT16, Tre98, ZHZ^{+}15]. Trail [Del98c]. Train [TSK97]. Trainable [TS13]. Trained [PB06]. Trainer [KCMK20]. Training [AMJ^{+}08, BDK^{+}04, BSHW07, BKM16, CGT^{+}23, CCC18, COM15, DFP^{+}11, HSHS05, KYK^{+}24, KCMK20, LGW^{+}12, LLZ13, LK95b, LSMC04, LCR^{+}14, Mor98, MSB^{+}06, NCQ^{+}11, PFB06, PHAS20, RMR^{+}19, Sad16, Saw83, SKW^{+}12, SSS00, ZPA95, ZED^{+}09, ZPL^{+}23, SSRH95]. Trajectory [JRM17]. Transactions [Ano16-28, Ano16-29, Ano18a, Ano18b, Ano19o, Ano19k, Ano19l, Ano20q, Ano20m, Ano20n, Ano20o, Ano20p, Ano20g, Ano20h, Ano21z, Ano21-27, Ano21-28, Ano21-29, Ano21n, Ano21o, Ano21p, Ano21q, Ano21t, Ano21r, Ano21s, Ano22-28, Ano22s, Ano22t, Ano22u, Ano22v, Ano23p, Hay96b]. Transatlantic [MN97]. Transcontinental [LJ96]. Transfer [Ash03, FHH^{+}21, HHKK17, KPTS89, LK82, LZS04, PLB^{+}01, RAGS01, SCOSL13, WFOP85, ZLMY19]. Transferring [YP08]. Transform [SMP^{+}22, Bli93b, Fan86]. Transformation [Agr21a, CT09, Ros89b, Sin18a, Sin20c]. Transformations [Bar81a, GWW86, Gol03, Rob89, ZZ05, AES93]. Transforming [FZH^{+}13, KB09, LRIC15, SC21, Sin14f, TBP01, VT24]. Transforms [Bli99c, Kri04, Mur93]. Transient [Hei18, RLCI24]. Transistors [FH84b]. Transition [PvW07, ZPL^{+}23]. Translated [MR99]. Translating [Hag94, vOW15]. Translation [Ken81, LWZY19, MOP89]. Translator [Kon12]. Translators [MD99d]. Translucency [JWSG10]. Translucent [HG13, SKP09]. Transmission [VB24]. Transparency [Mam89, Nie07]. Transparent [KB14, RFN89]. Transplantation [vBE12]. Transport [BDC^{+}12, DKS^{+}03, STW^{+}08]. Transportable [MD99e]. Transportation [ADZK^{+}21, Pac10]. Traumatic [Mac09]. Travel [Sin10e]. Traveling [Abo00b, Gla88b]. Traversed [GBCK93]. Treadmill [Iwa99]. Treat [JAM^{+}05]. Treating [HAB^{+}01, Mac09, WAB^{+}13]. Treatment [RRF^{+}23]. Treatments [Ano97i]. Tree [AK84, DBJEY^{+}02, GMTF89, Sch11, SRDT01]. Treemaps [Hei03]. Trees [HBW^{+}14, JB97, MMP^{+}07, Web08, ZK10, GC91]. Treevis.net [Sch11]. Trenches [APB^{+}21]. Trend [MNC^{+}22]. Trends [Ahr22, Ano20-70, Gar15, Oia82, Ped01]. Triage [Bli90b]. Trial [Abo98a, CLP^{+}15, Sin04e, Sim93d]. Trial-and-Error [Sin04e]. Triangle [De 89, KS98]. Triangle-Based [De 89]. Triangles [PQC^{+}12]. Triangular [Gla98d, GS94, KSK00]. Triangulated [MW82]. Triangulation [KS01, PT94, Woe83, Wor83, YMS01, FP93, FP95, Sug92, Use93]. Triangulations [Sch93c]. Tribute [Shn20]. Trick [GFW^{+}16]. Tricks [Bli97a, Gla01f, Sin11d]. Trigger [EAM22]. trigonal [WS92].

Trillion [AL11]. Trimmed [Cas87]. Trip [Bli96d, Dio97, GAe^{+}24, Bli91b, Bli91c, Bli91d, Bli92c, Bli93b]. Triple [GI94]. Triple-Step [GI94]. Triplet [Gla89]. Tripling [SML^{+}09]. Triumph [Sin14b]. True [Bar97a, FPT^{+}82, PJC88, WWC91]. true-color [WWC91]. Truga001 [IM98]. Trust [BABB^{+}21, vdEAA^{+}23]. Trusted [Rei23]. truth [Bli90a]. Try [CLP^{+}15, ZGL^{+}17]. Try-On [CLP^{+}15, ZGL^{+}17]. Tsunami [Kri05]. tubes [Bli89b]. Tumor [DPL^{+}11]. Tuned [HBBG95]. Tunnel [BL92, SOSG08, Use98]. Tunnels [DOH^{+}20]. turbulence [PCS95]. Turbulent [GCS^{+}12]. Turbulent-Flow [GCS^{+}12]. Turn [Han00a]. Turning [CBC^{+}20]. Turnkey [BH81]. Turntable [PLS^{+}16]. Turntable-Based [PLS^{+}16]. Tutorial [Dev99b, DGY07, Far92, HJS00, Hod92, Jos11, LPZ12, Mac99a, Nad99, SAH00, Ste07, UF11, YWR09, Gla92]. Tutorials [Sin11e]. TV [Hay95b, Hay96c, Lea97c, San01]. TVCG [Hay96b]. TViews [MRD06]. Twitter [ŽLK22a]. Two [BS86, Bli03b, Bli03d, CKJC12, EZM^{+}97, ESS82, Gol83b, HBB20, MKFN87, PO84, SB86a, TH84]. Two-Bit [SB86a]. Two-Dimensional [ESS82, TH84]. Two-Handed [EZM^{+}97]. Two-Way [MKFN87]. Type [Hua11]. Types [MTT83]. Typographic [MTS^{+}20]. Typology [vLBW17].

Ubi [KL09]. Ubi-Pen [KL09]. Ubiq [FCS22]. Ubiquitous [AJH^{+}12, CJS^{+}19, Tac16]. UIMS [OD88]. Ultrafractal [Sin07e]. Ultrascale [AS10, JLJ^{+}10]. Ultrasonic [ACO00, SSG95]. Ultrasonograms [Nak84]. Ultrasound [NE93, NCQ^{+}11, SRJ99]. Ultrasound-Guided [NCQ^{+}11]. Unbound [RBF95]. Uncalibrated [HWY16]. Uncanny [Gel08]. Uncertainty [BJP23, FDK12, GMR^{+}23, HNC^{+}04, JS03, MHTD23, PGA13, PGM19, SCB^{+}04b, TBB^{+}23, vLBW17]. Uncertainty-Aware [PGM19]. Unconstrained [RDRS10]. Undergraduates [Ren15]. Underground [MSH^{+}08, Gel07a]. Undersea [Sim95d]. Understand [LSB18, WYG^{+}12, YCB^{+}22]. Understanding [BQP^{+}19, CS15c, FMMS03, GDS^{+}18, Kro04, LF87, LFFB20, MUN^{+}10, PKN^{+}22, SADC02, SFT^{+}07, TLCC17, WGI15, ZOC^{+}12]. Underwater [CWBS99, DNBC16, GPA^{+}17, KL18, SCB^{+}04b, WAPW06, aE01, vL16, Ste91]. Undo [TSUC86]. Unexpected [CES07]. Unfolding [Bac16]. Unification [LF87, Pal85]. Unified [SS19, STJ^{+}20, YT84]. Uniform [Chr83, FP93, Use93]. Unifying [WJ93]. Union [FE97]. Unique [MYT09, RR95]. Uniting [Sch08a]. Unity [Sin06a]. Univariate [RO86, War88]. Universe [Mus99, GF94]. University [GAe^{+}24, Sou04, CHP21, Del00c, FE97, PS22]. Unknown [WJP22]. Unlimited [Ano18-40]. Unlocking [Myn08, Sin03e]. Unmanned [WPL22]. Unorganized [LCW12]. Unscripted [ALD^{+}06]. Unseen [BC19, Yea97a]. Unsolved [Bli98c, Che05]. Unsteady [GHCW21, OKET12]. Unstructured [FS01, RO96]. Untitled [Sin07d]. Untouchable [KKTI04]. Unwrapping [AL02]. Update [Ano92i, Ano92j, Bli89a]. Updated [Mac83, MD99e]. Updates [MD99d]. Updating [SBC15]. Upgrade [MD00f]. Upon [Bli98e, Gla98e]. Upper [CKPB16, KmKK^{+}18]. upper-body-part [KmKK^{+}18]. Upper-Limb [CKPB16]. Uppers [Bli92a, Bli92b]. uPy [AJH^{+}12]. Urban [ABVA08, CC20, CBZ^{+}08, Del00b, DFL^{+}18, HYN03, Hua21, KCW^{+}17, KJR^{+}18, Kub84, PJC24, RWF02, SMP^{+}19, WW08, WMV^{+}08, ZY18, KCW^{+}17]. Urn [Gol86]. urnful [Gol83c]. Usability [SPT^{+}15]. Usable [ECNZ15, DS87]. Usage [KKG^{+}13]. Use [Ano14b, AM84, CLP^{+}14, Enc11, JLS15, LR90, MTT83, Mur84a, Smi84a, SRS83, Ano14a, Rob88]. Used [MTN18]. useful [GP91a]. User [Ano84, Bli05b, BA85, BCF^{+}08, BCI08, CCT^{+}20, DEN13, DMSJ23, DJS^{+}09, ENDAD13, EJ98, EI17, FH84a, FB09, GHS99, Hay97, Her23, KM88, KHI^{+}03, KGB14, KPL17, LJ15, LRZ23, Ma00, Mab82, MSWT14, Mye84a, OBE^{+}84, Pat12, PvW07, RSP15, RCB^{+}05, Rod14, RZSP04, SDS^{+}11, Tak85, WP87, Wil86, WWV^{+}10, EFB^{+}94, Mye88, Ols86, Rob88, WC91]. User-Centered [GHS99]. User-Defined [PvW07]. User-Friendly [Pat12]. User-Guided [CCT^{+}20]. User/Artist [Hay97]. Users [Cav87, JBD^{+}21, MSDS23, MCGS21, TBFGC23, ZOC^{+}12, ZSK^{+}21]. Uses [HS01, Wor93]. Using [ADZK^{+}21, ANS15, ABM^{+}01, ABWS05, ABS^{+}05, AL11, BMC84, Bai09, Bai11, Bai13, BRD^{+}20, BKRE88, BQP^{+}19, BDH^{+}89, BVRF05, Bec97, BS03, BKV^{+}10, BSP97, BA85, BWG24, CD12, CF16, CR15, CWBS99, CJS^{+}19, CLC^{+}96, CCT^{+}20, CKPB16, DKS^{+}03, DLN05, DL15, DiS14, DNS^{+}15, DSA^{+}19, DD19, DY07, EH16, FH21, FS86, FKKB05, FH84b, FG88b, FTAT00, Fyo97, GLC11, GPW^{+}19, GMTF89, GH86, HF18, HTZ^{+}19, HB84, HT04, HB08, HJF06, HP91, HEW17, HKSK17, IO07, ITW01, JS10, JSH16, Jar83, JLJ^{+}10, JK09, JWHW12, JXW^{+}08, JAM^{+}05, KS98, KLR12, KK94, KP14, hKKSK17, KS97, KL18, KRG^{+}22, KPL17, LC16, LM10, Lar10, LCC^{+}00, LMZ11, LLZ13, LK08, LSFW83, LSMC04, LSVB21, MLF^{+}12, Mac99a, MQM22, MG14, MG92, MMR^{+}06, Moh85, MJC08, MML^{+}15, MR87, MBV17, NdSM^{+}16, ND05]. Using [OWZ17, PHE21, PH98, PH99, PBV16, Pie87b, PCJ^{+}16, RBBS06, RGL^{+}06, RKJH99, Rob87, RDDT97, RPM17, RW99, SFD16, SADC02, Sch00, SD07, SR95, SSC^{+}11, SHMN09, SC12, SP04a, Spa16b, SH97, SPM10, SB13, SP06, SAHK13, TSK97, TS22, TLS^{+}14, Tur88, VPVG21, Vos85, VP16, WYHS21, WZL^{+}21, WJ88, WSW08, Wil87b, WN82, XDJ23, YKR^{+}84, YSWW08, YLR17, YMS01, ZZ05, ZLMY19, ZE23, ZCBM14, Car87, CP92, EW96, FP93, MPHK94, Mil93, RBOV94, Use93]. Utility [FZH^{+}13]. Utilization [BAB^{+}23a]. UTM [WPL22]. UV [BRL^{+}17].

V [BRL06, CMBZ00, PLM^{+}19, XY01]. V-HairStudio [XY01]. V-Pong [BRL06]. V.2.0 [Lea96a]. V1.10 [MD99f]. VADE [JJW^{+}99]. Vadis [KH14]. Vail [Atw84]. Validating [TCE^{+}02, GG87]. Valley [Gel08, PSX13, Sin20e]. Valley-Ridge [PSX13]. Value [CS15b, HMC^{+}20, KKF^{+}21, WSK^{+}19, Wri98]. Values [DY09, Enc14d]. Vamping [Sin07e]. Variability [BWR24, SBS10]. Variables [Pet12]. Variational [GT93]. Variations [Abo98b]. Varied [MYPT09]. Variety [SP15]. Various [Bar84, Ano98j, NINT89]. Varying [GCS^{+}12, JWHW12, RBK^{+}13, WYM10]. Vascular [GGG08]. VAST [PGS^{+}08, SLC^{+}23, LWM14, PCH^{+}07, TPR^{+}22]. Vauquois [DOH^{+}20]. VBBTs [CKPB16]. VDT [Smi84a, Smi84b]. VE [MR98]. VECnt [MD99e]. Vector [ASS^{+}21, Bli00d, CMB94, Dal19, Eas90, FHH^{+}21, HTZ^{+}19, JXW^{+}08, Mil99a, Mil99b, MR87, UIL^{+}06, YM95, YZM^{+}11, EW96, HH91, HPvW94, Wan93]. Vectorization [PB85]. Vectorized [DW87]. Vectors [Bli05c, PWH04]. Vegetable [Wea99a]. Vehicle [LD97, Sin10e, SAHK13]. Vehicles [SGG^{+}22, WPL22]. Venn [Gla03f]. Verbal [KCMK20]. Verbs [RCB98]. Verifiable [KS08]. Verification [YY99]. Verifying [AHP^{+}10, BWWM12]. Versa [AB14a]. Version [Ber86, MD99f, MD99e]. Versus [KCMK20, DBV^{+}08, Eas90, PH98, PH99].

Vertices [AAB^{+}23]. Very [PSJ87]. VEs [CHC^{+}97, Goe96, Hay96d, Iwa99, SBSW96]. Vesna [VCS19]. Vessel [WOF96]. Vessels [BGSC85, CM16]. Vestibular [CHC^{+}97]. VFX [Abo00a]. vg500 [MD99f]. VHF [SRJ99]. Via [BC19, CPM^{+}18, HTZ^{+}19, LK82, LHZ^{+}21, OCS20, AHP^{+}10, FK87, GWW86, GHCW21, HPX^{+}21, HSK02, MD99d, RMP^{+}16, SCOSL13, XZZL19]. Vibratory [IWF97]. Vibrotactile [KL09, MGO^{+}16]. Vice [AB14a]. Victoria [VCS19]. Victories [She06]. Video [BN95, BFW99, Bux05, DB16, EH16, HFZ^{+}13, IKSZ03, Kri04, LaV08, LYO02, LFPS01, LFMGPG23, LWY^{+}20, MD99d, MKKJ96, Mur84b, Mye84b, NTH02, NCO03, ONDA20, QNT02, Saw08, SLNG07, She06, Sze96, WW24, YNH04, dHPP10, Bli90c, DS87, MED^{+}93]. Video-Based [BN95, HFZ^{+}13, NTH02]. Video-Enabled [NCO03]. Videocassette [Moh85]. Videodisc [Whi95d]. Videos [CTM18, DV21, FHH^{+}21, PD21]. Videotex [Cha85]. View [AEB86, Ano98h, Ano98i, Ano98j, AR99, BB05, BA85, Che08, CXPG19, FM22, KIN^{+}05, KKV^{+}02, MD99d, NR03, XZZL19, YCB05, KmKK^{+}18, OG91]. View-Dependent [FM22, KIN^{+}05, NR03]. Viewed [XHC^{+}17]. Viewer [KZC^{+}11, JSH16]. Viewer-Centric [KZC^{+}11]. Viewing [AEFQ96, CCF97, DB03, GWW86, GSN09, kHHBEH13, Kas00, RC85, AES93]. Viewpoint [GSF08]. Viewpoint-Based [GSF08]. Viewpoints [Bai01, BFH02, Bet00, CZGR09, CSWB01, CEH^{+}09, Dom00, HD01, Int00, Lai01, LLR09, Ma00, MQR99, PLB^{+}01, Shn03, ST09, Tay02, Whi99, Won08]. viewport [Bli92c]. Views [DL22, FZ03, GH86, HWY16, LHFL20, Rob87, SH97, WSL^{+}20]. Village [Sim96c]. Virtual [AMJ^{+}08, AKFPW20, AWB^{+}97, AS96, Ano94g, Ano20u, AA09, AR99, AB14b, BO09, Bad23, BNA^{+}22, BRB17, BWA96, BES17, BRG^{+}00, BRD17, BKV^{+}10, BFSE01, BHG23, BMH98, Bro99, BvDE^{+}99, BH06, BL92, Bry97, CPCS08, CLP^{+}14, CLP^{+}15, CEMD22, CNT^{+}97, CWZ99, CMBZ00, CGMB20, CGT^{+}23, CEMTT98, CZAA23, CBD10, CdHM97, CZP^{+}10, CCC18, Ch’11, CS17b, CT04b, Coo02a, CNL00, DBV^{+}08, DRHL^{+}03, DBS96, DGA^{+}97, DSA^{+}19, Dio97, DASF13, DJP^{+}08, DV21, DPG18, DOH^{+}20, EMTTT98, EM09, EG98, Ell94a, EBT98, EGR94, EBBS00, EYY03, FSP00, Fis02, FLMZ97, FCS22, FPW^{+}00, GHS99, GMNF97, Got09, GAe^{+}24, GOPK18, HCL^{+}20, HZHD03, HAH^{+}23, HLCC08, HH95, HBW^{+}24, HILW98, HCC05, HKSK17, IHL18, JJW^{+}99, JK09, JXW^{+}08, JLC^{+}02, KMTM^{+}98, KYK^{+}24, KMCMT09, KGK94, KP14, KLL^{+}15, KCMK20, KRB11, LTB22, LO94]. Virtual [Lea96b, Lea97b, LD97, LJ96, LA20a, LA20b, LLG^{+}07, LOB08, LRS09, LSMC04, Lok06, LSV^{+}02, LFMGPG23, LPMAA^{+}22, Mac09, MT94, MD99e, MTT86, Mam89, MSU^{+}21, MGO^{+}16, MKH20, MSA^{+}94, MFM^{+}16, MG14, MM21, MKL97, Nad99, NdSM^{+}16, NP17, NCQ^{+}11, NPB^{+}18, NZS21, NZB00, dSNRdLdSN16, OBG^{+}23, PHE21, PCZ^{+}09, PKT01, PSE^{+}09, PHAS20, PE09, PTR^{+}98, PD21, PBSB08, PBM^{+}96, PFA86, RL07, RGRLFM21, RRF^{+}23, RBBS06, RMDT96, RSDP01, RMR^{+}19, Rok96, RBF95, Ros00a, SS97, SK13, SRE98, SG11, SGC21, SGE^{+}18, SADH23, SC21, She04, She05, SVT10, SHY^{+}23, Sim93d, SHMN09, Sin11c, SPS99, SKMY09, SSS00, Sou04, SD99, Sra23, Sty96, Sun21, SK05, Sze96, TK03, TYK01, TSK97, TG21, TPA24, TWB^{+}07, UVa95, VS07, VK17, VMJT96, Wea98a, WLVL^{+}96, WSSS23, YIÇ09, YMK^{+}12, ZGL^{+}17]. Virtual [Zha22b, ZZ99, ZLK^{+}22b, dOYG04, Bal98c, Kah94, KF94, Kru95, Pap96, PFH95, RBOV94, Bry97, CC20, Sto99, SBJ^{+}01]. Virtual-Reality [DPG18]. Virtual-Wear [CLP^{+}15]. VirtualDesk [FFN19]. Virtuality [BKP01]. Virtualized [KN07]. Virtually [Cha97b, HWK^{+}96, Tac16]. Vis [CS15c, Bal98b]. VisAdapt [JOG^{+}17]. Visage [KRL97]. Viscoelastic [SSB06, TNTM07]. VisDB [KK94]. Visenet [LFV^{+}97]. VisHikers [WXK22]. Visibility [CWWW21, DKS^{+}03, Gün20, Hor84, WS90a]. Visible [BFP86, DOH^{+}20, RWF02, SW85, Sin19c, XDJ23, Yea97a, Sim96b, TSH96, Whi95d]. Vision [Ano23-46, CWZ99, De 11, Fer01, FAB^{+}98, KVM^{+}05, Lec17, Sch99, SKN^{+}05, MG88b, Agr21a, MD99e]. Vision-Based [KVM^{+}05]. Visions [Enc06, Sin02a]. VisLitE [FJL22]. Visual [Agr19, AD05, ACO00, AAG^{+}22, AAAW22, AGBC09, AOH^{+}14, Bac16, BRB17, BHPS12, Bas14, BQP^{+}15, BQP^{+}19, BHM^{+}22, BSM^{+}15, BBC^{+}87, BBC^{+}20, CLDW16, CLP18, CR15, Cha09, CZGR09, Che08, CTW^{+}10, CP13, CL18, CCC^{+}07, CPG^{+}16, COM15, CRD^{+}18, CR02b, Cro18, DGTK07, DSK^{+}14, Die85, DFL^{+}18, DL16, DLEF15, Dwy16, EKM^{+}22, EE12, EBN13, End14, EI17, FML23, FAL^{+}06, GCS^{+}12, GMR^{+}23, GRW99, GGS^{+}18, GDS^{+}18, HPX^{+}21, HYL^{+}22, HBW^{+}14, HKW99, Hua21, JS10, Jae17, JKKM^{+}03, JSDL21, KCW^{+}17, KFH^{+}18, KPSN04, KFBI22, Kro02, KW15, KVHD17, LKPM19, LPLN04, LGM13, LMA^{+}17, Li22b, LS04a, LAC^{+}14, LRS12, LTB16, MZC^{+}21, MMN^{+}21, MSW10, MGO14, MD99e, MD00f, MHR^{+}09, MGA^{+}21, MHTD23, MB05, MZ22, MNC^{+}22, MGN^{+}11, MMD^{+}19, ND05, NRH^{+}22, OBO^{+}23, PGS^{+}08, PGS09, PTB^{+}15, QC15, RMG16, RFA^{+}22]. Visual [RR84, Ren15, RSE99, RBD14, Sad16, SBCvdS04, SMG^{+}21, SCB^{+}04b, SMWS88, SMP^{+}19, SGC21, SZP^{+}00, SZMJ12, SLC^{+}23, She06, Sin04f, Sin18c, SKMY09, Sto15, SEAKC21, Sun21, SW22, Swi10, TH14, TMWJK04, TC06, Tit07, TC15, TPR^{+}22, VSLG22, WLL^{+}16, WYG^{+}12, WNAJ92, Won99, WT04, WLL^{+}09, WSP12, WSJ^{+}12, WMFM13, WHC^{+}14, WDC^{+}23, XLJ^{+}21, XZWH14, YLL10, YMK^{+}12, YCB^{+}22, ZY18, Zha22d, ZHZ^{+}15, ZYL^{+}17, vL18, vL16, BF93]. Visual-Analysis [Swi10]. Visual-Analytics [LAC^{+}14, LRS12, MHR^{+}09, PGS09, RBD14]. Visual-Computing [SZP^{+}00]. Visual-Interactive [BSM^{+}15]. Visual-Perception-Guided [XZWH14]. Visual-Quality [SGC21]. Visual/Haptic [LPLN04]. Visualization [AAB^{+}22, Abo99b, APB^{+}21, Agr09, ABM^{+}01, AS10, AHP^{+}10, Ahr22, AWM10, ACL^{+}12, AGR21b, ASS^{+}21, ABB^{+}93, Ano87, Ano92k, Ano92l, Ano92m, Ano93v, Ano19-61, AHDG00, BHH^{+}21, Bai01, Bai09, Bai11, Bai13, BB95, BBF^{+}98, BAB^{+}23a, BKS20, BRM10, BS16, BM24, DFQ12, BFR21, BABB^{+}21, BFH02, BSS20, Bet00, BvRS^{+}11, BAB^{+}23b, BHAA^{+}13, BYP21, BCPS22, BMR90, BAE^{+}18, BWG18, BHS^{+}24, BJP23, BEAC^{+}18, BD21, CRF^{+}21, CCG^{+}04, CCB22, CS15a, Cam16, CHW21, CCE^{+}14, CR95, CZGR09, CSWB01, CJS^{+}19, Che05, CEH^{+}09, CH10, CGJ^{+}17, CRA^{+}18, hCRBK98, CPA^{+}10, CBGH19, CLms15, CSA22, CG94, CNC^{+}05, CAG^{+}06, CMB94, Cro96, CKJC12, CM11, CCM15, CM16, CWL^{+}10, DHC^{+}22, Del00b, DSK^{+}14, DKS^{+}05, Dev99b, DPL^{+}11, DMAS17, DFM^{+}02, Dom00, Dom12, DPG18, Dwy16, Dye90, Eic96b, Eic05, Enc11, ECS^{+}15, Enc16b]. Visualization [Enc17a, EKS22, FGW00, FF20, FJL22, FDK12, Fis21, FM86, FMMS03, FT00b, FT16, FLSG98, Fyo97, Gai04, Gai17, Ger91, GB96a, GE97a, GE97b, Ger98, Gie92, GSG^{+}21, GSW23, GSW24, GGS22, GB16, GGS^{+}18, GFW^{+}16, Gün20, HM95, HHY20, Hay96b, Hee99, Hei03, HKDL22, HAH^{+}23, HMHM18, HO91, HD01, HT04, HJF06, HSAKD98, HLW13, HLB^{+}07, ISMGG22, Int00, IBH^{+}09, IIH^{+}13, ITSK06, IK15, JBD^{+}21, JBH^{+}10, JLJ^{+}10, JOG^{+}17, Joh04, JLMV06, KTD10, KDH^{+}95a, KC05, KM00, KNR93, KCB^{+}23, Kee10, KK94, KQM13, KHP^{+}11, Ker13, KS08, KM13, Kir21, Koh93, KNRB12, KDHL08, Kos16, Kos22, Kos23, KNK99, KKF^{+}21, KRG^{+}22, KGM95, KGR^{+}18, KF94, KH14, KÇC^{+}01, KHSE98, Lai01, LFV^{+}97, Lar10, Lar11, LTJ18, LPLN04, LTD16, LCG^{+}02, LTS^{+}13, LIS^{+}19, LCI^{+}20, LMWC17, LZF^{+}14, LWZ^{+}18, Lin99]. Visualization [LCB^{+}23, LQ18, LLR09, LCC^{+}96, LRZ23, LPMAA^{+}22, LY92, LRCT22, LHM23, MP99, Ma00, Ma01, Ma06, Ma07, MFL08, Ma09, MLF^{+}12, Ma23, MJM93, MD99f, MD99e, MD99d, Mac00b, MRB19, MAB05, MLJ19, Mat96, MVWB05, MCW93, MA98, MQR99, ME22, MD18, MG04, Mil12, MKKJ96, MZ22, MML^{+}15, Mor16, MSU^{+}16, MKL97, MED^{+}93, MRBW12, Mun02, MJM^{+}06, NPH^{+}09, Nap05, NE93, NK94, Nor06, NR03, NAB^{+}23, OKM84, OKET12, ODE^{+}13, Pac10, PW97, PP98, PLM^{+}19, PKM18, PCS95, PA18, PRN15, PS09, PL19, PLB^{+}01, PS22, PGA13, PGM19, PJJF96, PM97, RMP^{+}16, RRR^{+}16, Rei17, Rhe02, Rhy00, Rhy03, Rhy11, RBOV94, RKJH99, RR03, Ric15, RSK^{+}06, RGG19, RNNTD14, RBS21, RRB^{+}14, Rod14, RS97, RDH^{+}02, RB92, Ros94, RG95, RGCB07, RLV^{+}16, RDDY07, RW99, RSL^{+}19, SE19, SKR18, SFD16]. Visualization [SKS^{+}13, SMG^{+}21, Sch93b, SKW^{+}12, Sch91, Sch11, SWS07, SI94, SB03a, SSM^{+}07, SZZ^{+}18, Shn03, SvW09, Sil95, Sin09e, SFL^{+}04, SAO^{+}95, SMMS24, ST09, SS20, Ste07, Ste91, SGEF^{+}12, SBHM97, SSMH03, SSZS21, Tay02, TLCC17, TSH96, TC15, Tre95, TC96, TS97, Tre98, Tre99, TBO^{+}20, UFK^{+}89, UIL^{+}06, UBS^{+}12, UZM^{+}23, VCR98, VCBS97, WAH^{+}21, WHR19, WSK^{+}19, Wan22, War08, WCE^{+}17, WSK^{+}00, WSW08, WBS^{+}05, WWW^{+}23, WWG97, WPL22, Whi99, WGI15, WL88, WFL^{+}00, Won08, Wri97, Wri98, WVL^{+}18, WFZB21, WKY23, YNF^{+}92, YWG^{+}10, YCKF15, Zha96a, ZCS^{+}06, Zha07, Zha21, ZE23, ZCBM14, Zhu12, ZPG^{+}21, ZOC^{+}12, vDFL^{+}00, vD05, vHV09, van93, vdEAA^{+}23, vLBW17, BEW91, Eic96a, Erv93, GB96b, HPvW94, JM95, KHK^{+}94, NBG^{+}94, OJ14, Pal92, Phi91, RET^{+}94, RN94, SSRH95, UO91, YYOiT92, Hay96d]. Visualization [Lea97c, NVC92, vWNS10]. Visualizations [BF93, BS03, BWG24, BVD19, DD19, GRvL^{+}23, Han14, HE02, KKVL05, LTB22, Ma00, MSRMH09, Rei23, ST09, TM05, WSP^{+}20, WTLM15, WTPHK09, ŽLK22a]. Visualize [Bli02b, CF16, Koz12, LSVB21, PBV16, SEAKC21]. Visualizer [CPCS08]. Visualizes [Del97]. Visualizing [ASP^{+}21, ABWS05, AL02, And98, ABB^{+}93, ASK94, Bak88, BCMF16, BGM^{+}07, BGM19, BCI08, Car96, CPM^{+}18, CBS^{+}22, CWGC90, CWBS99, CHY99, DH93, DP00, FPNVHH17, FOI^{+}07, FLNR90, GMNF97, Ger91, GPS^{+}21, GDF^{+}14, GSFT16, HM95, HH91, HGK10, IG98, JS03, KJR^{+}18, KP12, Kla89, Koh93, KNK99, KGM95, LC16, LSB21, LPK05, Ma00, MP01, MJM93, MQR99, MTS91, Mil07, MV93, NFHL91, NL01, OKS^{+}17, OKET12, OWZ17, PRdJ07, Pic90, PW14, PvW07, RXB09, RFK^{+}13, RLIB99, RBR^{+}93, Rob88, Ros89a, SPJ99, SKS^{+}13, Sch93b, SAH00, SRJ99, Sen99, STW^{+}08, Sin17e, Tay02, TBB^{+}23, WYM10, WAPW06, WGP97, Wil82, WDK17, Woo15, XHC^{+}17, YBL^{+}12, Zha96a, aCJB^{+}12, Dia94, TSH96]. Visually [ADZK^{+}21, BWR24, GJM18, LRIC15]. Visuohaptic [MSB^{+}06]. VisVisual [Wan22]. Vitality [ZY18]. VitalVizor [ZY18]. Vivid [Sin03f]. Viz [Sin20e]. VizSim [Del99a]. VLNet [CNT^{+}97]. VLSI [GVA86, MN84, PJC88, PK87, Shi86]. vocabulary [McC83]. Voice [TBFGC23]. Vol [Ano98f, Ano99, Ano02, Ano03a, Ano04a, Ano05a, Ano06a]. Volcanic [GSFT16, RG95]. Volume [AWM10, Ano00, Bai01, BHAA^{+}13, Boh95, CJS^{+}19, CBB^{+}14, CW93, DCLK03, FTAT00, Gie92, GP93a, GSHG98, HBR^{+}92a, HJCW06, KMM^{+}01, Lev88, Lev90, Lev00, MD99f, MD99d, MB94, MSRMH09, Mur93, Neu94, NR03, RDRS10, RBE08, RHC95, SM84, SP15, SES05, TR83, WW86, WK94, WD15c, WMG00, YNF^{+}92, ZCBM14, Gor83, HBR^{+}92b, KHK^{+}94, MPHK94, SSG95]. Volume-based [HBR^{+}92a, HBR^{+}92b]. Volume-Rendering [Neu94]. Volume-Sampled [WK94]. Volumes [ACO00, Ber86, LGLD86, Pic90]. Volumetric [EZM^{+}97, SKUT^{+}10, NFMD90, SB97, Sen99, WBS^{+}05, WL88, UO91]. Volunteer [Ano20-51, Ano21-66, Ano21-67, Ano21-68, Ano23-64, Ano23-65]. Volunteered [SMP^{+}19]. Voreen [MSRMH09]. Voronoi [AM84, Cho95, LBD^{+}92]. Vortex [KH98, MWY^{+}15]. Vortical [OKET12]. Vorticity [BHS^{+}24]. Vorticity-Driven [BHS^{+}24]. Vous [MTT87b]. Voxar [Pit93]. Voxel [CMP95, FGR85, GWW86, KB88, LHZ^{+}21]. Voxel-Based [CMP95, FGR85, GWW86, KB88]. Voxelization [COK97]. Voyage [CNL00, Sin06b]. Voyages [Kd21, GF94]. VPark [JMT^{+}01]. VPS [SW22]. VR [Agr18b, Ano94g, Bal97, BSHW07, Ber99, BKP01, BHMN95, BSHB01, Bry97, CWZ99, CMBZ00, Cha97b, CZY^{+}13, CMO^{+}99, COM15, CKP96, CNL00, CN03, DB96, Del00c, Dio97, EBT98, EBBS00, FKC^{+}05, FLMZ97, FFA^{+}97, FPW^{+}00, GMNF97, Goe96, GHR01, GRW99, HAB^{+}01, HILW98, IS95, Jer18a, JW01, JXW^{+}08, JLC98, JMOG99, JMC^{+}02, KIQ^{+}06, KKVL05, KKV^{+}02, KKM96, LaV08, LMPW01, LGK^{+}06, LFV^{+}97, Lea97d, LGM13, Lec17, LD97, LRS09, LK08, MR98, MD00f, MCSH21, MBSD20, MFG^{+}17, MTN18, MJC08, MYK^{+}08, PNPS03, PMQ21, PD21, PBSB08, PBVH15, RSS^{+}09, RWF02, RSDP01, RRV^{+}15, RSZ96, RDO^{+}96, Ros97, RDDT97, RBT98, Rou09, SS97, SS01, SKJ^{+}13, SF20, SK03, SHY^{+}23, Sim95d, Sim95e, Sim96a, SPS99, SSS00, SD99, SDM22, SSRH95, Tac98, aKTRL^{+}21, UVa95, VK17, VFLJ02, WS09a]. VR [WS09b, WSW08, WWG97, ZPA95, Zha22b, ZR24, ZPL^{+}23, ZSK^{+}21, ZD98, vDFL^{+}00]. VR-Based [BSHW07, RRV^{+}15, ZSK^{+}21]. VR/AR [Agr18b, PBVH15, WSW08, Zha22b]. VRAIS [Bal98c, GH97, Lea97d]. VRML [Ano98j, BDG^{+}99, CPC99, GTHL99, JW01, Law99, Lea96a, MR99, Nad99, RLIB99, RW99, Whi96a]. VRMosaic [AS96]. VTK [CJS^{+}19, MSU^{+}16, SAH00]. VTK-m [MSU^{+}16]. Vulcan [AGBC09].

W [Bli98f]. Wafer [Lin99]. Wait

[DDdSC11]. Wait-Free [DDdSC11]. Wake [Sin22f]. Wakes [Gos90]. Walk [EBBS00, IDJW19]. Walk-up [EBBS00]. Walking [NPB^{+}18, SB13]. Walkthroughs [RAPL98, KKM96]. Wall [LCC^{+}00, Whi99]. Walled [MRL13]. WALLoE [Kan09]. Walls [SFF^{+}00, WAB^{+}05]. Wanna [Abo00c]. Wanted [Ano18o]. War [DOH^{+}20, Whi95d]. Warning [RM95]. Warping [BBDZ08, FS90, KBGK17, XJW18, Zha96b]. WarpIV [RLV^{+}16]. Was [Lev00]. Waste [GAe^{+}24]. Watch [Ano15-44, Koz12, Nie93a, RC85]. Water [ASP^{+}21, FAZH17, LHFL20, MMMM21, NTB^{+}20]. Watermarking [Ben99, BFW99, CT09, SP04a, YY99]. Watts [Ano22-81, Ano23-84, Ano23-85, Ano23-86]. Wave [BC87, Enc13, Kas11c, NTM^{+}01, VS14, SADH23]. Wavefront [Hay95b]. Wavelet [DBJEY^{+}02, Mur93]. Wavelets [SDS95a, SDS95b]. Waves [MM14]. Way [Arn14a, MKFN87, Wor93, DB96]. Ways [Hay96a, vHdP94]. Wear [CLP^{+}15]. Wearable [Ano15-45, BBDM98, GFW^{+}16, KBHT06, PGV20]. Wearables [Ano15-36]. Wearing [Sin18f]. Weather [KJA^{+}18, TC96, Whi96d, Gel07a]. weatherman [Whi96d]. Weaving [Gla02a, Gla03a, Gla03b, IM14]. Weavy [IM14]. Web [BDC^{+}12, LFMGPG23, AS96, Ano94d, Ano98h, Ano98i, Ano98j, CFF01, CRF^{+}21, CAP^{+}05, FPNVHH17, Ger96, HB06, Hay96f, Hay96d, HJF06, JKKM^{+}03, KSRS13, Kri01, LLK^{+}03, LK03, Law99, Lea96a, LJB^{+}13, LPMEPCM11, Mac13, MD99f, MD99e, MD99d, RS97, SADH23, Sim95a, Sto00, Whi96a, Whi96d, Whi97, Won08]. Web-Based [CFF01, CRF^{+}21, CAP^{+}05, FPNVHH17, HJF06, JKKM^{+}03, RS97, SADH23]. Weber [WCS18]. WebGL [Ang17, BMS^{+}13]. WebGL-Based [BMS^{+}13]. Week [Ano23-34, Ano23-66, Ano23-67, Ano20-52, Ano20-53]. Weight [SHY^{+}23, ZJH^{+}18]. Weight-change [ZJH^{+}18]. Weighted [GH86, WWW^{+}23]. Weighting [XJW18]. Welcome [Sto08, Tho02a]. Well [CLP18]. Westenberg [RvW21]. Western [Hay95b]. Whales [WAPW06]. Wheel [CBB^{+}14, FPNVHH17]. Where [Bli88e, Sch07, Sim93c]. Wherever [Ano18-50, Ano14-39, Ano15-43]. Whether [FM82]. Which [EAM22, SF20, Sin10e]. WHIM [Goo86]. Whirled [Bli02b]. Whisper [Sin09f]. White [RB93]. Wicked [RPOC21]. Wide [Ger96, GDY^{+}11, Sim95a, Whi96d]. Wide-Area [GDY^{+}11]. Widely [SKJ^{+}13]. Wider [Bal97]. Widgets [REF^{+}06]. Wiimote [WWV^{+}10]. Wikipedia [XHC^{+}17]. Wild [GSW23, LRZ23]. Wildfire [BHS^{+}24, MA98, PGM19, SF24]. Wildlife [ZLK^{+}22b]. Wilkie [Wil13]. Will [Per16, SF20, Wor96b, Bal97]. William [Ebe17]. WIMP [vD00]. Win [Sin23b].

Wind [BL92, Gel07a, Use98]. Windchill [MD99e]. Winding [SAB^{+}23]. Window [Goo86, PH98, PH99, Pin91, Mye88]. Windows [Ber83, CSI86, PH98, PH99]. Winds [Mac97a]. Windtunnel [Bry97]. Winner [BHS^{+}24, DPL^{+}11, GGS^{+}18, LZF^{+}14, OKET12, SKS^{+}13, SMG^{+}21, SLC^{+}23]. Winners [FKK^{+}10]. Wins [WW95]. Wired [Swe01]. Wireless [Lea96a, Kri01]. Wiring [Ski81]. wise [LLL^{+}13]. Withering [JWHW12]. within [AS96, Kri05, MR98, VK17]. Without [Sin17f, SDM22, Whi95b, PGA13, Whi96d]. Witkin [De 11]. Wolfgang [Enc15a]. Woman [Del98c]. Women [Ano20-35, Ano20-33, Ano20-34]. Wonder [Sin15f]. Wonderful [Bli90c]. Wonderland [WGP97]. Wondrous [SR10]. Word [CWL^{+}10, LBSS85]. Wordle [JLS15]. WordlePlus [JLS15]. Words [Whi95c]. Work [Agr18b, Cha97b, FPH^{+}06, Hib04, KBHP14, LJ96, Lil81, MD99e, Mil03, NdSM^{+}16, Sin11f, Sin12c, Smi84b, TM05, KF94, Sim96b]. Workbench [Coh83, KKAM84, LSR^{+}00, LRS12, Bry97, KF94, Ros97, RDDT97, WWG97]. Workflow [BWR24, NTS^{+}07, WLC^{+}12]. Workflow-Driven [NTS^{+}07]. Workflows [Kee10]. Working [BW00, CLP^{+}14, Sin10f]. Workload [Hei03]. Works [KM13]. Worksheets [BD21]. Workshop [AJD13, Atw84, ENDAD13, PBSB08]. Workspace [FJHW00, STJ^{+}20, WBA^{+}96]. Workstation [BKRE88, CK85a, FH84a, GK86, GRB^{+}85, Kel84, Nic84, Pra84b, Smi84a, Spi86, WK85, Wil84]. Workstations [GP93a, HW82b, MD99e, MSC89, Ree84, PABS86, RMM^{+}91]. World [AAB^{+}22, Ano15-44, Ano17c, Ano19-39, Ano19-40, Ano20-63, BWWM12, BCPK02, Bli05b, CMT^{+}10, DASF13, Gel07b, GE97b, Ger98, Gre86b, Hag90b, IS95, ML05, MRB19, NL01, Pen10, PTR^{+}98, Rao97, RTF^{+}05, Sin02c, Sin14c, Sin21c, TG21, TH01, UCK^{+}04, Zha21, ZED^{+}09, ZPL^{+}23, dOYG04, vOW15, Ano97i, Ano17b, Bli88a, Bli90c, Mas18, DOH^{+}20, Ger96, Sim95a, Whi96d]. Worlds [Abo00b, BDG^{+}99, Coo02a, JW01, Lea96a, Mac09, Nad99, SK13, Sch08a, Swe01, TYK01, WSSS23]. Worldview [Kal87]. Worldwide [Sin02e]. Worm [TS97, GR94]. Worth [Whi95c]. Worthiness [KHSE98]. Worthy [Ano20-61]. Woven [ADBW96]. Wrangling [BGM19]. Write [GL82, Bli87]. Writing [MIEL00]. wrongs [Bli95c]. WWW [Sim96c]. WYSBOAVRTWYG [Bli91a].

X [Mac97a, PH98, PH99]. X-Rays [Mac97a]. XFEM [JK09]. Xflow [KSRS13]. XML3D [KSRS13]. Xroads [Ano15-38].

Ya [Abo00c]. YaQ [MYPT09]. Year [Abo99b, Ang17, Ano94n, Bal98b, Her96, MLJ19, Tau12d]. Years [KGR^{+}18, Mye85, NPB^{+}18, ODE^{+}13, PW14, SAB^{+}23]. York [Blo86]. Young [MFM^{+}16]. Yuan [WWF00, WWF00].

Z [BFP86, DW87, Rao97]. Z-Buffer [BFP86, DW87]. Z-GUI [Rao97]. Zen [Sin20f]. Zero [Boh95]. Zero-Volume [Boh95]. Zoom [Hay95b].

References

Argelaguet:2009:EPS

[AA09] Ferran Argelaguet and Carlos Andujar. Efficient 3D pointing selection in cluttered virtual environments. IEEE Computer Graphics and Applications, 29(6): 34–43, November/December 2009. CODEN ICGADZ.

ISSN 0272-1716 (print), 1558-1756 (electronic).

Andrienko:2022:VAH

[AAAW22] Natalia Andrienko, Gennady Andrienko, Linara Adilova, and Stefan Wrobel. Visual analytics for humancentered machine learning. IEEE Computer Graphics and Applications, 42(1): 123–133, January/February 2022. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Abdelaal:2022:VAE

[AAB^{+}22] Moataz Abdelaal, Felix Amtsberg, Michael Becher, Rebeca Duque Estrada, Fabian Kannenberg, Aimée Sousa Calepso, Hans Jakob Wagner, Guido Reina, Michael Sedlmair, Achim Menges, and Daniel Weiskopf. Visualization for architecture, engineering, and construction: Shaping the future of our built world. IEEE Computer Graphics and Applications, 42(2):10–20, March/ April 2022. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Ahmed:2023:SVL

[AAB^{+}23] Reyan Ahmed, Patrizio Angelini, Michael A. Bekos, Giuseppe Di Battista, Michael Kaufmann, Philipp Kindermann, Stephen Kobourov, Martin Nöllenburg, Antonios Symvonis, Anaïs Villedieu, and Markus Wallinger. Splitting vertices in 2-layer graph drawings. IEEE Computer Graphics and Applications, 43(3):24– 35, May/June 2023. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Andrienko:2022:SVE

[AAG^{+}22] Gennady Andrienko, Natalia Andrienko, Jose Manuel Cordero Garcia, Dirk Hecker, and George A. Vouros. Supporting visual exploration of iterative job scheduling. IEEE Computer Graphics and Applications, 42 (3):74–86, May/June 2022. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Andrienko:2017:GDS

[AAW17] Gennady Andrienko, Natalia Andrienko, and Robert Weibel. Geographic data science. IEEE Computer Graphics and Applications, 37(5):15–17, September/ October 2017. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2017/05/mcg2017050015.
html.

Avila:2014:RDV

[AB14a] Lisa Avila and Mike Bailey. From real to digital, and vice versa. IEEE Computer Graphics and Applications, 34(4):80–c3, July/ August 2014. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Avila:2014:VRM

[AB14b] Lisa Avila and Mike Bailey. Virtual reality for the masses. IEEE Computer Graphics and Applications, 34(5):103–104, September/ October 2014. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL http://csdl.
                                        computer.org/csdl/mags/
                                        cg/2014/05/mcg2014050103.
                                        html.

Aziz:1990:BSS

[ABB90] Nadim M. Aziz, Reda Bata, and Sudarshan Bhat. Bézier surface/surface intersection. IEEE Computer Graphics and Applications, 10(1):50– 58, January/February 1990. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anninos:1993:VBV

[ABB^{+}93] Peter Anninos, Mark Bajuk, David Bernstein, Edward Seidel, Larry Smarr, and David Hobill. Visualization blackboard: Visualizing black hole space-times. IEEE Computer Graphics and Applications, 13(1):12– 14, January/February 1993.

CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Azuma:2001:RAA

[ABB^{+}01] Ronald Azuma, Yohan Baillot, Reinhold Behringer, Steven Feiner, Simon Julier, and Blair MacIntyre. Recent advances in augmented reality. IEEE Computer Graphics and Applications, 21(6):34–47, November/ December 2001. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://dlib.
computer.org/cg/books/
cg2001/g6034abs.htm;
http://dlib.computer.org/
cg/books/cg2001/pdf/g6034.
pdf.

Ahrens:2001:LSD

[ABM^{+}01] James Ahrens, Kristi Brislawn, Ken Martin, Berk Geveci, C. Charles Law, and Michael Papka. Largescale data visualization using parallel data streaming. IEEE Computer Graphics and Applications, 21(4):34– 41, July/August 2001. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

computer.org/cga/cg2001/
g4034abs.htm;   http:/
/dlib.computer.org/cg/
books/cg2001/pdf/g4034.
pdf.

Abouaf:1998:ATF

[Abo98a] Jeffrey Abouaf. Applications: Trial by fire: Teleoperated robot targets Chernobyl. IEEE Computer Graphics and Applications, 18(4):10–14, July/August 1998. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Abouaff:1998:AVP

[Abo98b]  Jeffrey Abouaff.  Ap-

plications: Variations on perfection: The Séquin-Collins sculpture generator. IEEE Computer Graphics and Applications, 18(6):15–20, November/ December 1998. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL http://dlib.
                                        computer.org/cg/books/
                                        cg1998/pdf/g6015.pdf.

Abouaf:1999:ACA

[Abo99a] Jeffrey Abouaf. About the cover: Adventure game tools get smarter — “Prince of Persia 3D”. IEEE Computer Graphics and Applications, 19(4):4–5, July/ August 1999. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL http://dlib.
                                        computer.org/cg/books/
                                        cg1999/pdf/g4004.pdf.

Abouaf:1999:AFP

[Abo99b]  Jeffrey Abouaf.  Appli-

cations: The Florentine Pietà: Can visualization solve the 450-year-old mystery? IEEE Computer Graphics and Applications, 19(1):6–10, January/ February 1999. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://dlib.
computer.org/cg/books/
cg1999/pdf/g1006.pdf.

Abouaf:2000:ACC

[Abo00a] Jeffrey Abouaf. About the cover: Creating illusory realism through VFX. IEEE Computer Graphics and Applications, 20(4):4–5, July/ August 2000. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://dlib.
computer.org/cg/books/
cg2000/pdf/g4004.pdf.

Abouaf:2000:ACT

[Abo00b] Jeffrey Abouaf. About the cover: Traveling in fantastic worlds. IEEE Computer Graphics and Applications, 20(3):4–5, May/ June 2000. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://dlib.
computer.org/cg/books/
cg2000/pdf/g3004.pdf.

Abouaf:2000:MYW

[Abo00c] Jeffrey Abouaf. Maya: “So ya wanna be a rock ’n roll star” revisited. IEEE Computer Graphics and Applications, 20(2):7–11, March/ April 2000. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://dlib.
computer.org/cg/books/
cg2000/extras/g2007x1.
mov;   http://dlib.computer.
org/cg/books/cg2000/extras/
g2007x2.mov;   http://
dlib.computer.org/cg/books/
cg2000/pdf/g2007.pdf.

Amor:2005:MSE

[ABS^{+}05] Margarita Amor, Montserrat Boo, Wolfgang Strasser, Johannes Hirche, and Michael Doggett. A meshing scheme for efficient hardware implementation of butterfly subdivision using displacement mapping. IEEE Computer Graphics and Applications, 25(2):46–59, March/ April 2005. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Arasan:2016:HNS

[ABS16] Atakan Arasan, Cagatay Basdogan, and Tevfik Metin Sezgin. HaptiStylus: A novel stylus for conveying movement and rotational torque effects. IEEE Computer Graphics and Applications, 36(1):30–41, January/ February 2016. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://www.
computer.org/csdl/mags/

cg/2016/01/mcg2016010030-
abs.html.

Aliaga:2008:IRU

[ABVA08] Daniel G. Aliaga, Bedřich Beneš, Carlos A. Vanegas, and Nathan Andrysco. Interactive reconfiguration of urban layouts. IEEE Computer Graphics and Applications, 28(3):38–47, May/ June 2008. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Albrecht-Buehler:2005:VLT

[ABWS05] Conrad Albrecht-Buehler, Benjamin Watson, and David A. Shamma. Visualizing live text streams using motion and temporal pooling. IEEE Computer Graphics and Applications, 25(3):52–59, May/ June 2005. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Angesleva:2005:LC

[AC05] Jussi Angesleva and Ross Cooper. Last clock. IEEE Computer Graphics and Applications, 25(1):20–23, January/February 2005. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

apCenydd:2012:VSL

[aCJB^{+}12] Llyr ap Cenydd, Nigel W.

John, Marina Bloj, Annette Walter, and Nicholas I. Phillips. Visualizing the surface of a living human brain. IEEE Computer Graphics and Applications, 32(2): 55–65, March/April 2012. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Alsakran:2012:RTV

[ACL^{+}12] Jamal Alsakran, Yang Chen, Dongning Luo, Ye Zhao, Jing Yang, Wenwen Dou, and Shixia Liu. Real-time visualization of streaming text with a force-based dynamic system. IEEE Computer Graphics and Applications, 32(1):34–45, January/ February 2012. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Aiger:2000:MUV

[ACO00] Dror Aiger and Daniel Cohen-Or. Mosaicing ultrasonic volumes for visual simulation. IEEE Computer Graphics and Applications, 20(2):53–61, March/ April 2000. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

computer.org/cga/cg2000/
g2053abs.htm;   http:/
/dlib.computer.org/cg/
books/cg2000/pdf/g2053.
pdf.

Agrawala:2005:GEI

[AD05] Maneesh Agrawala and Fredo Durand. Guest Editors’ introduction: Smart depiction for visual communication. IEEE Computer Graphics and Applications, 25(3):20–21, May/ June 2005. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cg/2005/03/g3020.pdf.

Aggarwal:2012:HFI

[ADBR^{+}12] Vibhor Aggarwal, Kurt Debattista, Thomas Bashford-Rogers, Piotr Dubla, and Alan Chalmers. Highfidelity interactive rendering on desktop grids. IEEE Computer Graphics and Applications, 32(3):24–36, May/June 2012. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Aono:1996:FWC

[ADBW96] Masaki Aono, Paolo Denti, David E. Breen, and Michael J. Wozny. Fitting a woven cloth model to a curved surface: Dart insertion. IEEE Computer Graphics and Applications, 16(5):60–70, September 1996. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Addison:1995:DBD

[Add95] R. Addison. “Detour: Brain deconstruction ahead”. IEEE Computer Graphics and Applications, 15(2):14–17, March/April 1995. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

AL-Dohuki:2021:QVS

[ADZK^{+}21] S. AL-Dohuki, Y. Zhao, F. Kamw, J. Yang, X. Ye, and W. Chen. QuteVis: Visually studying transportation patterns using multisketch query of joint traffic situations. IEEE Computer Graphics and Applications, 41(2):35–48, March/ April 2021. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

amEmde:2001:AMU

[aE01] Barbara Anne am Emde. Applications: 3D mapping of underwater caves. IEEE Computer Graphics and Applications, 21(2):14–20, March/April 2001. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL http://dlib.
                                        computer.org/cg/books/
                                        cg2001/pdf/g2014.pdf.

Abi-Ezzi:1986:IVP

[AEB86] Salim S. Abi-Ezzi and Albert J. Bunshaft. An implementor’s view of PHIGS.

IEEE Computer Graphics and Applications, 6(2):12– 23, February 1986. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Akkiraju:1996:VGP

[AEFQ96] Nataraj Akkiraju, Herbert Edelsbrunner, Ping Fu, and Jiang Qian. Viewing geometric protein structures from inside a CAVE. IEEE Computer Graphics and Applications, 16(4):58–61, July/August 1996. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Abi-Ezzi:1993:SBV

[AES93] Salim S. Abi-Ezzi and Leon A. Shirman. The scaling behavior of viewing transformations. IEEE Computer Graphics and Applications, 13(3):48–54, May/June 1993. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Altenhofen:2021:AIP

[AESF21] Christian Altenhofen, Tobias Ewald, André Stork, and Dieter W. Fellner. Analyzing and improving the parameterization quality of Catmull–Clark solids for isogeometric analysis. IEEE Computer Graphics and Applications, 41(3):34–47, May/June 2021. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Aliaga:2003:SID

[AFYC03] Daniel G. Aliaga, Thomas Funkhouser, Dimah Yanovsky, and Ingrid Carlbom. Sea of images: a dense sampling approach for rendering large indoor environments. IEEE Computer Graphics and Applications, 23(6):22–30, November/December 2003. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http:

                                        //csdl.computer.org/dl/
                                        mags/cg/2003/06/g6022.
                                        pdf;   http://csdl.computer.
                                        org/dl/mags/cg/2003/06/
                                        g6022abs.htm.

Agba:1995:SRS

[Agb95] Emmanuel I. Agba. SeaMaster: An ROV-Manipulator system simulator. IEEE Computer Graphics and Applications, 15(1):24–31, January/February 1995. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Andrysco:2009:AVE

[AGBC09] Nathan Andrysco, Kevin Robert Gurney, Bedrich Beneš, and Kathy Corbin. Applications: Visual exploration of the Vulcan CO_{2} data. IEEE Computer Graphics and Applications, 29(1):6– 11, January/February 2009. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Armstrong:1987:NCH

[AGL87] William W. Armstrong, Mark Green, and Robert Lake. Near-real-time control of human figure models. IEEE Computer Graphics and Applications, 7(6): 52–61, June 1987. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Arunkumar:2022:PLS

[AGP^{+}22] Anjana Arunkumar, Nitin Gupta, Andrea Pinceti, Lalitha Sankar, and Chris Bryan. PMUVis: a largescale platform to assist power system operators in a smart grid. IEEE Computer Graphics and Applications, 42(6):84–95, November/December 2022. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Agrawal:2009:GSC

[Agr09]  Amit Agrawal.  Graph-

ically speaking: Coming next issue: Collaborative visualization. IEEE Computer Graphics and Applications, 29(4):81–85, July/ August 2009. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Agrawal:2018:AML

[Agr18a] Amit Agrawal. Application of machine learning to computer graphics. IEEE Computer Graphics and Applications, 38(4):93–96, July/ August 2018. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        cg/2018/04/mcg2018040093.
                                        html.

Agrawal:2018:PVA

[Agr18b] Amit Agrawal. Putting VR/ AR to work. IEEE Computer Graphics and Applications, 38(1):115–118, January/February 2018. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL https:/

                                        /www.computer.org/csdl/
                                        mags/cg/2018/01/mcg2018010115.
                                        html.

Agrawal:2019:MIV

[Agr19]  Amit Agrawal.  Mining

insights from visual assets. IEEE Computer Graphics and Applications, 39(1):107–109, January/ February 2019. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL https:

                                        //ieeexplore.ieee.org/
                                        document/8663647/.

Agrawal:2021:DVA

[Agr21a] A. Agrawal. Dolby Vision: Advancing the technology of cinema and home entertainment transformation of an industry. IEEE Computer Graphics and Applications, 41(2):96–98, March/ April 2021. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Amabili:2021:TDM

[AGR21b] Lorenzo Amabili, Kuhu Gupta, and Renata Georgia Raidou. A taxonomydriven model for designing educational games in visualization. IEEE Computer Graphics and Applications, 41(6):71–79, November/December 2021. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Aguilar:1986:FGC

[Agu86] Lorenzo Aguilar. A format for a graphical communications protocol. IEEE Computer Graphics and Applications, 6(3):52–62, March/ April 1986. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Adelson:1995:GER

[AH95] Stephen J. Adelson and Larry F. Hodges. Generating exact ray-traced animation frames by reprojection. IEEE Computer Graphics and Applications, 15(3):43– 52, May/June 1995. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Azuma:2000:VTF

[AHDG00] Ronald Azuma, Howard Neely III, Michael Daily, and Ryan Geiss. Visualization tools for free flight air-traffic management. IEEE Computer Graphics and Applications, 20(5):32–36, September/ October 2000. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

                                        computer.org/cga/cg2000/
                                        g5032abs.htm;   http:/
                                        /dlib.computer.org/cg/
                                        books/cg2000/pdf/g5032.
                                        pdf.

Ahrens:2010:VSS

[AHP^{+}10] James Ahrens, Katrin Heitmann, Mark Petersen, Jonathan Woodring, Sean Williams, Patricia Fasel, Christine Ahrens, Chung-Hsing Hsu, and Berk Geveci. Verifying scientific simulations via comparative and quantitative visualization. IEEE Computer Graphics and Applications, 30(6): 16–28, November/December 2010. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Ahrens:2015:ISD

[Ahr15] James Ahrens. Increasing scientific data insights about exascale class simulations under power and storage constraints. IEEE Computer Graphics and Applications, 35(2):8–11, March/ April 2015. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
cg/2015/02/mcg2015020008-
abs.html.

Ahrens:2022:TTC

[Ahr22] James Ahrens. Technology trends and challenges for large-scale scientific visualization. IEEE Computer Graphics and Applications, 42(4):114–119, July/ August 2022. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Alford:2013:AWT

[AJD13] Jennifer Ginger Alford, Lucas Jacob, and Paul Dietz. Animatronics workshop: A theater + engineering collaboration at a high school. IEEE Computer Graphics and Applications, 33(6):9–13, November/December 2013. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Autin:2012:UUC

[AJH^{+}12] Ludovic Autin, Graham Johnson, Johan Hake, Arthur Olson, and Michel Sanner. uPy: a ubiquitous CG Python API with biological-modeling applications. IEEE Computer Graphics and Applications, 32(5):50–61, September/ October 2012. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Artner:2015:EGS

[AJK15] Nicole M. Artner, Ines Janusch, and Walter G. Kropatsch. Evaluating and grading students in large-scale image processing courses. IEEE Computer Graphics and Applications, 35(5):101–c3, September/October 2015. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL http://www.
                                        computer.org/csdl/mags/
                                        cg/2015/05/mcg2015050101-
                                        abs.html.

Aono:1984:BTI

[AK84] Masaki Aono and Tosiyasu L. Kunii. Botanical tree image generation. IEEE Computer Graphics and Applications, 4(5):10–29, 32–34, May/June 1984. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Akeley:1989:SGS

[Ake89] Kurt Akeley. The Silicon Graphics 4D/240GTX superworkstation. IEEE Computer Graphics and Applications, 9(4):71–83, July/ August 1989. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Al-Kalbani:2020:EDS

[AKFPW20] M. Al-Kalbani, M. Frutos-Pascual, and I. Williams. Evaluation of drop shadows for virtual object grasping in augmented reality. IEEE Computer Graphics and Applications, 40(4): 10–25, July/August 2020. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Adams:1990:CS

[AKME^{+}90] Ludwig Adams, Werner Krybus, Dietrich Meyer-Ebrecht, Rainer Rueger, Joachim M. Gilsbach, Ralph Moesges, and Georg Schloendorff. Computer-assisted surgery. IEEE Computer Graphics and Applications, 10(3):43–51, May/ June 1990. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Tadeja:2021:MIP

[aKTRL^{+}21] S awomir Konrad Tadeja, Wojciech Rydlewicz, Yupu Lu, Tomasz Bubas, Maciej Rydlewicz, and Per Ola Kristensson. Measurement and inspection of photorealistic 3-D VR models.

IEEE Computer Graphics and Applications, 41(6):143– 151, November/December 2021. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anderson:2002:UVC

[AL02] Sean Anderson and Marc Levoy. Unwrapping and visualizing cuneiform tablets. IEEE Computer Graphics and Applications, 22(6): 82–88, November/December 2002. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL

                                        http://dlib.computer.org/
                                        cg/books/cg2002/pdf/g6082.
                                        pdf;   http://www.computer.
                                        org/cga/cg2002/g6082abs.
                                        htm.

Arietta:2011:BUD

[AL11] Sean Arietta and Jason Lawrence. Building and using a database of one trillion natural-image patches. IEEE Computer Graphics and Applications, 31(1):9– 19, January/February 2011. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Ala:1992:PAB

[Ala92] Seshagiri Rao Ala. Performance anomalies in boundary data structures. IEEE Computer Graphics and Applications, 12(2):49–58, March/April 1992. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Aylett:2006:UNA

[ALD^{+}06] Ruth Aylett, Sandy Louchart, Joao Dias, Ana Paiva, Marco Vala, Sarah Woods, and Lynne Hall. Unscripted narrative for affectively driven characters. IEEE Computer Graphics and Applications, 26(3):42– 52, May/June 2006. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Alexa:2021:PSA

[Ale21] Marc Alexa. PolyCover:

Shape approximating with discrete surface orientation. IEEE Computer Graphics and Applications, 41(3):85– 95, May/June 2021. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Alpert:1993:GIG

[Alp93] Sherman R. Alpert. Graceful interaction with graphical constraints. IEEE Computer Graphics and Applications, 13(2):82–91, March/ April 1993. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Arnold:1984:UVT

[AM84] D. B. Arnold and W. J.

Milne. The use of Voronoi tessellations in processing soil survey results. IEEE Computer Graphics and Applications, 4(3):22–28, March/April 1984. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Apodaca:1990:RPF

[AM90] Anthony A. Apodaca and M. W. Mantle. Render-Man: Pursuing the future of graphics. IEEE Computer Graphics and Applications, 10(4):44–49, July/ August 1990. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Angel:1991:SBA

[AM91] Edward Angel and Don Morrison. Speeding up Bresenham’s algorithm. IEEE Computer Graphics and Applications, 11(6):16–17, November/December 1991. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Aila:2004:DOC

[AM04] Timo Aila and Ville Miettinen. dPVS: An occlusion culling system for massive dynamic environments. IEEE Computer Graphics and Applications, 24(2):86– 97, March/April 2004. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

csdl.computer.org/comp/

mags/cg/2004/02/g2086abs.
htm;   http://csdl.computer.
org/dl/mags/cg/2004/02/
g2086.htm;   http://csdl.
computer.org/dl/mags/cg/
2004/02/g2086.pdf.

Amanatides:1987:RCG

[Ama87]  John Amanatides.  Re-

alism in computer graphics: a survey. IEEE Computer Graphics and Applications, 7(1):44–56, January/ February 1987. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Araujo:2022:PCC

[AMDM22] Victor Araujo, Julia Melgare, Bruna Martini Dalmoro, and Soraia Raupp Musse. Is the perceived comfort with CG characters increasing with their novelty? IEEE Computer Graphics and Applications, 42(1):32– 46, January/February 2022. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Aghina:2008:VCD

[AMJ^{+}08] Maurício Alves C. Aghina, Antônio Carlos A. Mól, Carlos Alexandre F. Jorge, Cláudio M. N. A. Pereira, Thiago F. B. Varela, Gerson G. Cunha, and Luiz Landau. Virtual control desks for nuclear power plant simulation: Improving operator training. IEEE Computer Graphics and Applications, 28(4):6–9, July/ August 2008. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Akimoto:1991:PRT

[AMS91] Takaaki Akimoto, Kenji Mase, and Yasuhito Suenaga. Pixel-selected ray tracing. IEEE Computer Graphics and Applications, 11(4):14–22, July/August 1991. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Andrews:1998:SNV

[And98] Keith Andrews. Short note: Visualizing rich, structured hypermedia. IEEE Computer Graphics and Applications, 18(4):40–42, July/ August 1998. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

                                        computer.org/cga/cg1998/
                                        g4040abs.htm.

Angel:2017:CTC

[Ang17] Ed Angel. The case for teaching computer graphics with WebGL: A 25year perspective. IEEE Computer Graphics and Applications, 37(2):106–112, March/April 2017. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/

cg/2017/02/mcg2017020106-
abs.html.

Annett:2017:DIC

[Ann17] Michelle Annett. (Digitally) inking in the 21st century. IEEE Computer Graphics and Applications, 37(1):92–99, January/ February 2017. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2017/01/mcg2017010092-
abs.html.

Anonymous:1984:CCU

[Ano84] Anonymous. CAD/CAM user survey evaluation. IEEE Computer Graphics and Applications, 4(2):21– 24, February 1984. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:1987:VSC

[Ano87] Anonymous. Visualization in scientific computing — a synopsis. IEEE Computer Graphics and Applications, 7(7):61–70, July/ August 1987. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:1988:HAS

[Ano88a] Anonymous. Hopper and ACM software systems awards. IEEE Computer Graphics and Applications, 8(2):90–91, March/ April 1988. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL http://www.
                                        computer.org/csdl/mags/
                                        cg/02/mcg1988020090-abs.
                                        html.

Anonymous:1988:N

[Ano88b] Anonymous. NCGA 88. IEEE Computer Graphics and Applications, 8(2):8–14, March/April 1988. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL http://www.
                                        computer.org/csdl/mags/
                                        cg/02/mcg1988020008-abs.
                                        html.

Anonymous:1988:NP

[Ano88c] Anonymous. New products. IEEE Computer Graphics and Applications, 8(2):92– 93, March/April 1988. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http:/

                                        /www.computer.org/csdl/
                                        mags/cg/02/mcg1988020092-
                                        abs.html.

Anonymous:1992:ACa

[Ano92a] Anonymous. About the cover. IEEE Computer Graphics and Applications, 12(5):4–??, September 1992. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:1992:ACb

[Ano92b] Anonymous. About the cover. IEEE Computer Graphics and Applications, 12(6):4–??, November/December 1992. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:1992:A

[Ano92c] Anonymous. Applications.

IEEE Computer Graphics and Applications, 12(6): 13–??, November/December 1992. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:1992:C

[Ano92d] Anonymous. Conferences.

IEEE Computer Graphics and Applications, 12(6): 90–??, November/December 1992. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:1992:DDa

[Ano92e] Anonymous. Displays on display. IEEE Computer Graphics and Applications, 12(4):8–??, July/August 1992. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:1992:DDb

[Ano92f] Anonymous. Displays on display. IEEE Computer Graphics and Applications, 12(5):8–??, September 1992. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:1992:DDc

[Ano92g]  Anonymous.  Displays

on display. IEEE Computer Graphics and Applications, 12(6):7–??, November/December 1992. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:1992:GN

[Ano92h] Anonymous. Graphic news. IEEE Computer Graphics and Applications, 12(6): 89–??, November/December 1992. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:1992:SUa

[Ano92i] Anonymous. Selective update. IEEE Computer Graphics and Applications, 12(4):14–??, July/August 1992. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:1992:SUb

[Ano92j]  Anonymous.  Selective

update. IEEE Computer Graphics and Applications, 12(5):14–??, September 1992. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:1992:VBa

[Ano92k] Anonymous. Visualization blackboard. IEEE Computer Graphics and Applications, 12(4):11–??, July/ August 1992. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:1992:VBb

[Ano92l] Anonymous. Visualization blackboard. IEEE Computer Graphics and Applications, 12(5):12–??, September 1992. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:1992:VBc

[Ano92m] Anonymous. Visualization blackboard. IEEE Computer Graphics and Applications, 12(6):10–??, November/December 1992. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:1993:ACa

[Ano93a] Anonymous. About the cover. IEEE Computer Graphics and Applications, 13(1):6–??, January/February 1993. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:1993:ACb

[Ano93b] Anonymous. About the cover. IEEE Computer Graphics and Applications, 13(2):5–??, March/April 1993. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:1993:ACc

[Ano93c] Anonymous. About the cover. IEEE Computer Graphics and Applications, 13(4):4–??, July/August 1993. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:1993:ACd

[Ano93d] Anonymous. About the cover. IEEE Computer Graphics and Applications, 13(5):4–??, September 1993. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:1993:ACe

[Ano93e] Anonymous. About the cover. IEEE Computer Graphics and Applications, 13(6):5–??, November/December 1993. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:1993:Aa

[Ano93f] Anonymous. Applications. IEEE Computer Graphics and Applications, 13(1):101– ??, January/February 1993. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:1993:Ae

[Ano93g] Anonymous. Applications. IEEE Computer Graphics and Applications, 13(6): 93–??, November/December 1993. CODEN ICGADZ.

ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:1993:Ca

[Ano93h] Anonymous. Conferences.

IEEE Computer Graphics and Applications, 13(1):102– ??, January/February 1993. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:1993:Cb

[Ano93i] Anonymous. Conferences.

IEEE Computer Graphics and Applications, 13(5):89– ??, September 1993. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:1993:Cc

[Ano93j] Anonymous. Conferences.

IEEE Computer Graphics and Applications, 13(6): 99–??, November/December 1993. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:1993:D

[Ano93k]  Anonymous.  Displays.

IEEE Computer Graphics and Applications, 13(6): 9–??, November/December 1993. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:1993:DD

[Ano93l] Anonymous. Displays on display. IEEE Computer Graphics and Applications, 13(2):8–??, March/April 1993. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:1993:GNa

[Ano93m] Anonymous. Graphic news. IEEE Computer Graphics and Applications, 13(1):100– ??, January/February 1993. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:1993:GNb

[Ano93n] Anonymous. Graphic news. IEEE Computer Graphics and Applications, 13(2): 92–??, March/April 1993. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:1993:GNc

[Ano93o] Anonymous. Graphic news. IEEE Computer Graphics and Applications, 13(4): 76–??, July/August 1993. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:1993:GNd

[Ano93p] Anonymous. Graphic news. IEEE Computer Graphics and Applications, 13(5):86– ??, September 1993. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:1993:GNe

[Ano93q] Anonymous. Graphic news.

IEEE Computer Graphics and Applications, 13(6): 98–??, November/December 1993. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:1993:La

[Ano93r] Anonymous. Letters. IEEE Computer Graphics and Applications, 13(1):4–??, January/February 1993. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:1993:Lb

[Ano93s] Anonymous. Letters. IEEE Computer Graphics and Applications, 13(2):4–??, March/April 1993. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:1993:LE

[Ano93t]  Anonymous.  Letters to

the editor. IEEE Computer Graphics and Applications, 13(6):4–??, November/December 1993. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:1993:NP

[Ano93u] Anonymous. New products.

IEEE Computer Graphics and Applications, 13(1):106– ??, January/February 1993.

CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:1993:VB

[Ano93v] Anonymous. Visualization blackboard. IEEE Computer Graphics and Applications, 13(6):13–??, November/December 1993. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:1994:ACb

[Ano94a] Anonymous. About the cover. IEEE Computer Graphics and Applications, 14(2):4–??, March/April 1994. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:1994:ACc

[Ano94b] Anonymous. About the cover. IEEE Computer Graphics and Applications, 14(3):4–??, May/June 1994. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:1994:A

[Ano94c] Anonymous. Applications. IEEE Computer Graphics and Applications, 14(1):10– ??, January/February 1994. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:1994:CEJ

[Ano94d] Anonymous. Comet explodes on Jupiter — and the Web. IEEE Computer Graphics and Applications, 14(6):12–13, November/December 1994. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:1994:Ca

[Ano94e] Anonymous. Conferences.

IEEE Computer Graphics and Applications, 14(1):80– ??, January/February 1994. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:1994:Cb

[Ano94f] Anonymous. Conferences.

IEEE Computer Graphics and Applications, 14(2): 93–??, March/April 1994. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:1994:CCC

[Ano94g] Anonymous. Conferences:

Clearing the clouds: Virtual gas dynamics; embracing VR as a liberator. IEEE Computer Graphics and Applications, 14 (3):89–??, May/June 1994. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:1994:DDb

[Ano94h] Anonymous. Displays on display. IEEE Computer Graphics and Applications, 14(2):7–??, March/April 1994. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:1994:DDc

[Ano94i] Anonymous. Displays on display. IEEE Computer Graphics and Applications, 14(3):8–??, May/June 1994. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:1994:EM

[Ano94j]  Anonymous.  Editor-in-

Chief’s message. IEEE Computer Graphics and Applications, 14(3):86–??, May/June 1994. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:1994:GNa

[Ano94k] Anonymous. Graphic news. IEEE Computer Graphics and Applications, 14(1):79– ??, January/February 1994. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:1994:GNb

[Ano94l] Anonymous. Graphic news. IEEE Computer Graphics and Applications, 14(2): 90–??, March/April 1994.

CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:1994:GNc

[Ano94m] Anonymous. Graphic news.

IEEE Computer Graphics and Applications, 14(3):83– ??, May/June 1994. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:1994:THS

[Ano94n] Anonymous. Top 10 hardware and software products of the year. IEEE Computer Graphics and Applications, 14(1):86–??, January/ February 1994. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:1995:APIa

[Ano95a] Anonymous. Advertiser/ product index. IEEE Computer Graphics and Applications, 15(3):96, May/ June 1995. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:1995:APIb

[Ano95b] Anonymous. Advertiser/ product index. IEEE Computer Graphics and Applications, 15(5):96, July/ August 1995. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:1995:APIc

[Ano95c] Anonymous. Advertiser/ product index. IEEE Computer Graphics and Applications, 15(5):104–??, September 1995. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:1995:APId

[Ano95d] Anonymous. Advertiser/ product index. IEEE Computer Graphics and Applications, 15(6):112–??, November/December 1995. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:1995:FFPa

[Ano95e] Anonymous. Final focus. IEEE Computer Graphics and Applications, 15(3):C3– ??, May/June 1995. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:1995:FFPb

[Ano95f] Anonymous. Final focus. IEEE Computer Graphics and Applications, 15(5):C3, July/August 1995. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:1995:FFPc

[Ano95g] Anonymous. Final focus. IEEE Computer Graphics and Applications, 15(6):C3, November/December 1995. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:1995:Ca

[Ano95h] Anonymous. On the calendar. IEEE Computer Graphics and Applications, 15 (3):89–??, May/June 1995. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:1995:Cb

[Ano95i] Anonymous. On the calendar. IEEE Computer Graphics and Applications, 15(5):92, July/August 1995. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:1995:Cc

[Ano95j] Anonymous. On the calendar. IEEE Computer Graphics and Applications, 15 (5):97–??, September 1995. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:1995:Cd

[Ano95k] Anonymous. On the calendar. IEEE Computer Graphics and Applications, 15(6): 96–??, November/December 1995. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:1996:AI

[Ano96a]  Anonymous.  1996 an-

nual index. IEEE Computer Graphics and Applications, 16(6):85–95, November/December 1996. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:1996:APIa

[Ano96b] Anonymous. Advertiser/ product index. IEEE Computer Graphics and Applications, 16(1):96, January/ February 1996. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:1996:APIb

[Ano96c] Anonymous. Advertiser/ product index. IEEE Computer Graphics and Applications, 16(2):96, March/ April 1996. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:1996:APIc

[Ano96d] Anonymous. Advertiser/ product index. IEEE Computer Graphics and Applications, 16(3):96, May/ June 1996. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:1996:APId

[Ano96e] Anonymous. Advertiser/ product index. IEEE Computer Graphics and Applications, 16(4):96, July/ August 1996. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:1996:APIe

[Ano96f] Anonymous. Advertiser/ product index. IEEE Computer Graphics and Applications, 16(5):96, September 1996. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:1996:APIf

[Ano96g] Anonymous. Advertiser/ product index. IEEE Computer Graphics and Applications, 16(6):96, November/ December 1996. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:1996:CF

[Ano96h]  Anonymous.  Change-of-

address form. IEEE Computer Graphics and Applications, 16(2):1–??, March/ April 1996. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:1996:CSI

[Ano96i] Anonymous. Computer Society information. IEEE Computer Graphics and Applications, 16(2):76, March/ April 1996. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:1996:FF

[Ano96j] Anonymous. Final focus. IEEE Computer Graphics and Applications, 16(1):C3– ??, January/February 1996. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:1996:ICA

[Ano96k] Anonymous. Information for CG&A authors. IEEE Computer Graphics and Applications, 16(2):C3–??, March/ April 1996. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:1997:AI

[Ano97a] Anonymous. 1997 annual index. IEEE Computer Graphics and Applications, 17(6):110–119, November/ December 1997. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://dlib.
computer.org/cg/books/
cg1997/pdf/g6110.pdf.

Anonymous:1997:APIa

[Ano97b] Anonymous. Advertiser/ product index. IEEE Computer Graphics and Applications, 17(2):96, March/ April 1997. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:1997:APIb

[Ano97c] Anonymous. Advertiser/ product index. IEEE Computer Graphics and Applications, 17(3):88, May/ June 1997. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:1997:APIc

[Ano97d] Anonymous. Advertiser/ product index. IEEE Computer Graphics and Applications, 17(4):96, July/ August 1997. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:1997:APId

[Ano97e] Anonymous. Advertiser/ product index. IEEE Computer Graphics and Applications, 17(5):88, September/October 1997. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:1997:APIe

[Ano97f] Anonymous. Advertiser/ product index. IEEE Computer Graphics and Applications, 17(6):120, November/ December 1997. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:1997:CF

[Ano97g]  Anonymous.  Change-of-

address form. IEEE Computer Graphics and Applications, 17(4):1–??, July/ August 1997. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:1997:CSI

[Ano97h] Anonymous. Computer Society information. IEEE Computer Graphics and Applications, 17(4):C3, July/ August 1997. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:1997:NTT

[Ano97i] Anonymous. In the news: Testing treatments on modeled hearts; Cornell carves the world’s smallest guitar, NSF looks for new centers to fund. IEEE Computer Graphics and Applications, 17(5):84, September/October 1997. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://dlib.
computer.org/cg/books/
cg1997/pdf/g5084.pdf.

Anonymous:1998:APIa

[Ano98a] Anonymous. Advertiser/ product index. IEEE Computer Graphics and Applications, 18(2):112, March/ April 1998. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:1998:APIb

[Ano98b] Anonymous. Advertiser/ product index. IEEE Computer Graphics and Applications, 18(3):96, May/ June 1998. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://dlib.
computer.org/cg/books/
cg1998/pdf/g3096.pdf.

Anonymous:1998:APIc

[Ano98c] Anonymous. Advertiser/ product index. IEEE Computer Graphics and Applications, 18(4):96, July/ August 1998. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:1998:APId

[Ano98d] Anonymous. Advertiser/ product index. IEEE Computer Graphics and Applications, 18(5):96, September/October 1998. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://dlib.
computer.org/cg/books/
cg1998/pdf/g5096.pdf.

Anonymous:1998:APIe

[Ano98e] Anonymous. Advertiser/ product index. IEEE Computer Graphics and Applications, 18(6):86, November/ December 1998. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://dlib.
computer.org/cg/books/
cg1998/pdf/g6081.pdf.

Anonymous:1998:AII

[Ano98f] Anonymous. Annual index, IEEE CG&A vol.

18. IEEE Computer Graphics and Applications, 18(6):87–96, November/ December 1998. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://dlib.
computer.org/cg/books/
cg1998/pdf/g6087.pdf.

Anonymous:1998:NIP

[Ano98g] Anonymous. In the news: Experiencing interactive photographs; animation of reality mimics fiction. IEEE Computer Graphics and Applications, 18(4):15, July/ August 1998. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http:/

/www.sandia.gov/media/
comethit.htm.

Anonymous:1998:WEVa

[Ano98h] Anonymous. Web extra: View animations from: the Optics Project. IEEE Computer Graphics and Applications, 18(4):??, July/ August 1998. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://www.
apple.com/quicktime.

Anonymous:1998:WEVb

[Ano98i]  Anonymous.  Web ex-

tra: View animations of Studierstube. IEEE Computer Graphics and Applications, 18(4):??, July/ August 1998. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://www.
apple.com/quicktime.

Anonymous:1998:WEVc

[Ano98j] Anonymous. Web extra:

View sample VRML files of various displays from the Flodar system. IEEE Computer Graphics and Applications, 18(5):??, September/October 1998. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http:

//www.computer.org/cga/
cg1998/extras/g5extra.
htm.

Anonymous:1999:IIC

[Ano99]  Anonymous.  1999 in-

dex: IEEE CG&A vol.

19. IEEE Computer Graphics and Applications, 19 (6):101–112, November/ December 1999. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://dlib.
computer.org/cg/books/
cg1999/pdf/g6101.pdf.

Anonymous:2000:IIC

[Ano00]  Anonymous.  2000 in-

dex, IEEE CG&A volume 20. IEEE Computer Graphics and Applications, 20(6):114–128, November/ December 2000. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://dlib.
computer.org/cg/books/
cg2000/pdf/g6114.pdf.

Anonymous:2001:AI

[Ano01] Anonymous. Annual index. IEEE Computer Graphics and Applications, 21(6):87–96, November/ December 2001. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://dlib.
computer.org/cg/books/
cg2001/g6087abs.htm;
http://dlib.computer.org/
cg/books/cg2001/pdf/g6087.
pdf;   http://dlib.computer.
org/cg/books/cg2002/pdf/
g1.zip.

Anonymous:2002:AIV

[Ano02] Anonymous. 2002 annual index, vol. 22. IEEE Computer Graphics and Applications, 22(6):119–131, November/December 2002. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL

http://www.computer.org/
cga/cg2002/g6119.pdf.

Anonymous:2003:ICG

[Ano03a] Anonymous. IEEE Computer Graphics and Applications 2003 annual index vol. 23. IEEE Computer Graphics and Applications, 23(6):102–111, November/ December 2003. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http:

//csdl.computer.org/dl/
mags/cg/2003/06/g6102.
pdf.

Anonymous:2003:STC

[Ano03b] Anonymous. Special thanks to CG&A’s reviewers. IEEE Computer Graphics and Applications, 23(1):24–25, January/February 2003. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http:

//dlib.computer.org/cg/
books/cg2003/pdf/g1024.
pdf.

Anonymous:2004:AII

[Ano04a]  Anonymous.  2004 an-

nual index IEEE computer graphics and applications vol. 24. IEEE Computer Graphics and Applications, 24(6):86–96, November/ December 2004. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cg/2004/06/g6086.pdf;
http://csdl.computer.org/
dl/mags/cg/2004/06/g6086.
htm.

Anonymous:2004:NPa

[Ano04b] Anonymous. New products.

IEEE Computer Graphics and Applications, 24(2):109– 111, March/April 2004. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http:

                                        //csdl.computer.org/dl/
                                        mags/cg/2004/02/g2109.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/cg/2004/02/
                                        g2109.pdf.

Anonymous:2004:NPb

[Ano04c] Anonymous. New products. IEEE Computer Graphics and Applications, 24(3):101– 104, May/June 2004. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http:

                                        //csdl.computer.org/dl/
                                        mags/cg/2004/03/g3101.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/cg/2004/03/
                                        g3101.pdf.

Anonymous:2004:NPc

[Ano04d] Anonymous. New products. IEEE Computer Graphics and Applications, 24(4):103– 104, July/August 2004. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

                                        csdl.computer.org/comp/
                                        mags/cg/2004/04/g4103.
                                        pdf;   http://csdl.computer.
                                        org/dl/mags/cg/2004/04/
                                        g4103.htm.

Anonymous:2004:NPd

[Ano04e] Anonymous. New products. IEEE Computer Graphics and Applications, 24(5):107–108, September/ October 2004. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cg/2004/05/g5107.pdf;
http://csdl.computer.org/
dl/mags/cg/2004/05/g5107.
htm.

Anonymous:2004:NPe

[Ano04f] Anonymous. New products. IEEE Computer Graphics and Applications, 24(6):75–77, November/ December 2004. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cg/2004/06/g6075.pdf;
http://csdl.computer.org/
dl/mags/cg/2004/06/g6075.
htm.

Anonymous:2004:STC

[Ano04g] Anonymous. Special thanks to CG&A’s reviewers. IEEE Computer Graphics and Applications, 24(1):22–23, January/February 2004. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

csdl.computer.org/comp/
mags/cg/2004/01/g1022.
pdf.

Anonymous:2005:AIV

[Ano05a] Anonymous. 2005 annual index, vol. 25. IEEE Computer Graphics and Applications, 25(6):87–95, November/December 2005. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

                                        csdl.computer.org/comp/
                                        mags/cg/2005/06/g6087.
                                        pdf.

Anonymous:2005:NPa

[Ano05b] Anonymous. New products. IEEE Computer Graphics and Applications, 25(1):94– 95, January/February 2005. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL

                                        http://csdl.computer.org/
                                        comp/mags/cg/2005/01/g1094.
                                        pdf.

Anonymous:2005:NPb

[Ano05c] Anonymous. New products. IEEE Computer Graphics and Applications, 25(2):107– 108, March/April 2005. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

                                        csdl.computer.org/comp/
                                        mags/cg/2005/02/g2107.
                                        pdf.

Anonymous:2005:NPc

[Ano05d] Anonymous. New products. IEEE Computer Graphics and Applications, 25(3):94– 95, May/June 2005. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

                                        csdl.computer.org/comp/
                                        mags/cg/2005/03/g3094.
                                        pdf.

Anonymous:2005:TPb

[Ano05e]  Anonymous.  Tools and

products. IEEE Computer Graphics and Applications, 25(4):86–88, July/ August 2005. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cg/2005/04/g4086.pdf.

Anonymous:2005:TPc

[Ano05f]  Anonymous.  Tools and

products. IEEE Computer Graphics and Applications, 25(5):93–95, September/October 2005. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cg/2005/05/g5093.pdf.

Anonymous:2005:TPd

[Ano05g]  Anonymous.  Tools and

products. IEEE Computer Graphics and Applications, 25(6):84–86, November/December 2005. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

csdl.computer.org/comp/
mags/cg/2005/06/g6084.
pdf.

Anonymous:2005:TPa

[Ano05h] Anonymous. Top products of 2004. IEEE Computer Graphics and Applications, 25(1):96, January/ February 2005. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cg/2005/01/g1096.pdf.

Anonymous:2006:AIV

[Ano06a] Anonymous. 2006 annual index, vol. 26. IEEE Computer Graphics and Applications, 26(6):105–112, November/December 2006. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL

http://bell.computer.org/
dlcomments/;   http://
csdl.computer.org/comp/
mags/cg/2006/06/g6105.
pdf.

Anonymous:2006:STC

[Ano06b] Anonymous. Special thanks to CG&A’s reviewers. IEEE Computer Graphics and Applications, 26(1):6–8, January/February 2006. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

csdl.computer.org/comp/
mags/cg/2006/01/g1006.
pdf.

Anonymous:2006:TPa

[Ano06c]  Anonymous.  Tools and

products. IEEE Computer Graphics and Applications, 26(1):94–95, January/ February 2006. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cg/2006/01/g1094.pdf.

Anonymous:2006:TPb

[Ano06d]  Anonymous.  Tools and

products. IEEE Computer Graphics and Applications, 26(2):92–95, March/ April 2006. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://bell.
computer.org/dlcomments/
;   http://csdl.computer.
org/comp/mags/cg/2006/
02/g2092.pdf.

Anonymous:2006:TPc

[Ano06e]  Anonymous.  Tools and

products. IEEE Computer Graphics and Applications, 26(3):94–96, May/ June 2006. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://bell.
computer.org/dlcomments/
;   http://csdl.computer.
org/comp/mags/cg/2006/
03/g3094.pdf.

Anonymous:2006:TPd

[Ano06f]  Anonymous.  Tools and

products. IEEE Computer Graphics and Applications, 26(4):101–104, July/ August 2006. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://bell.
computer.org/dlcomments/
;   http://csdl.computer.

                                        org/comp/mags/cg/2006/
                                        04/g4101.pdf.

Anonymous:2006:TPe

[Ano06g]  Anonymous.  Tools and

products. IEEE Computer Graphics and Applications, 26(5):94–95, September/October 2006. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL http://csdl.
                                        computer.org/comp/mags/
                                        cg/2006/05/g5094.pdf.

Anonymous:2006:TPf

[Ano06h]  Anonymous.  Tools and

products. IEEE Computer Graphics and Applications, 26(6):103–104, November/ December 2006. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL http://csdl.
                                        computer.org/comp/mags/
                                        cg/2006/06/g6103.pdf.

Anonymous:2007:STC

[Ano07a] Anonymous. Special thanks to CG&A’s reviewers. IEEE Computer Graphics and Applications, 27(1):6–8, January/February 2007. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

                                        csdl.computer.org/comp/
                                        mags/cg/2007/01/g1006.
                                        pdf.

Anonymous:2007:TPb

[Ano07b]  Anonymous.  Tool and

products. IEEE Computer Graphics and Applications, 27(2):85–87, March/ April 2007. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cg/2007/02/g2085.pdf.

Anonymous:2007:TPa

[Ano07c]  Anonymous.  Tools and

products. IEEE Computer Graphics and Applications, 27(1):97–99, January/ February 2007. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cg/2007/01/g1097.pdf.

Anonymous:2007:TPc

[Ano07d]  Anonymous.  Tools and

products. IEEE Computer Graphics and Applications, 27(3):96, 94–95, May/ June 2007. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://bell.
computer.org/dlcomments/
;   http://csdl.computer.
org/comp/mags/cg/2007/
03/g3096.pdf.

Anonymous:2008:AI

[Ano08a] Anonymous. Annual index.

IEEE Computer Graphics and Applications, 28(6):0, November/December 2008. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2008:BHR

[Ano08b]  Anonymous.  Bert Her-

zog, 1929–2008: Remembrances and recollections. IEEE Computer Graphics and Applications, 28(5):6– 9, September/October 2008. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL

                                        http://csdl.computer.org/
                                        comp/mags/cg/2008/05/mcg2008050006.
                                        pdf.

Anonymous:2008:STC

[Ano08c] Anonymous. Special thanks to CG&A’s reviewers. IEEE Computer Graphics and Applications, 28(1):6–9, January/February 2008. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

                                        csdl.computer.org/comp/
                                        mags/cg/2008/01/mcg2008010006.
                                        pdf.

Anonymous:2009:STC

[Ano09a] Anonymous. Special thanks to CG&A’s reviewers. IEEE Computer Graphics and Applications, 29(2):11–13, March/April 2009. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2009:TPa

[Ano09b]  Anonymous.  Tools and

products. IEEE Computer Graphics and Applications, 29(1):79–81, January/ February 2009. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2009:TPb

[Ano09c]  Anonymous.  Tools and

products. IEEE Computer Graphics and Applications, 29(2):88, c3, March/April 2009. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2009:TPc

[Ano09d]  Anonymous.  Tools and

products. IEEE Computer Graphics and Applications, 29(3):96, c3, May/ June 2009. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2009:TPd

[Ano09e]  Anonymous.  Tools and

products. IEEE Computer Graphics and Applications, 29(4):92–93, July/ August 2009. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2009:TPe

[Ano09f]  Anonymous.  Tools and

products. IEEE Computer Graphics and Applications, 29(5):101–103, September/ October 2009. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2009:TPf

[Ano09g]  Anonymous.  Tools and

products. IEEE Computer Graphics and Applications, 29(6):96, c3, November/ December 2009. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2010:AI

[Ano10a] Anonymous. Annual index. IEEE Computer Graphics and Applications, 30(1): ??, January/February 2010. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2010:JT

[Ano10b] Anonymous. Jim Thomas, 1946–2010. IEEE Computer Graphics and Applications, 30(6):10–13, November/December 2010. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2010:NS

[Ano10c] Anonymous. Notes from SIGGRAPH 2010. IEEE Computer Graphics and Applications, 30(5):88, September/October 2010. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2010:SSF

[Ano10d]  Anonymous.  Snapshots

of the state of the field.

IEEE Computer Graphics and Applications, 30(2):6–7, March/April 2010. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2011:AI

[Ano11a] Anonymous. Annual index.

IEEE Computer Graphics and Applications, 31(6):ii, November/December 2011. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2011:CP

[Ano11b] Anonymous. Call for papers.

IEEE Computer Graphics and Applications, 31(6):67, December 2011. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2011:TC

[Ano11c] Anonymous. Table of contents. IEEE Computer Graphics and Applications, 31(6):1–2, December 2011. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2011:TPa

[Ano11d]  Anonymous.  Tools and

products. IEEE Computer Graphics and Applications, 31(2):86–87, March/ April 2011. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2011:TPb

[Ano11e]  Anonymous.  Tools and

products. IEEE Computer Graphics and Applications, 31(5):95–96, September/October 2011. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2012:RT

[Ano12a]  Anonymous.  Reviewer

thanks. IEEE Computer Graphics and Applications, 32(1), January/February 2012. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2012:TP

[Ano12b]  Anonymous.  Tools and

products. IEEE Computer Graphics and Applications, 32(2):96, March/April 2012. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2013:C

[Ano13a] Anonymous. Certification. IEEE Computer Graphics and Applications, 33(6):47, November/December 2013. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2013:IP

[Ano13b] Anonymous. IEEE PacificViz 2014. IEEE Computer Graphics and Applications, 33(6):69, November/ December 2013. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2013:JB

[Ano13c] Anonymous. Jobs board.

IEEE Computer Graphics and Applications, 33(6):35, November/December 2013. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2013:Mb

[Ano13d] Anonymous. [Masthead].

IEEE Computer Graphics and Applications, 33(6):1, November/December 2013. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2013:Mc

[Ano13e] Anonymous. Membership.

IEEE Computer Graphics and Applications, 33(6):c3, November/December 2013. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2013:Ma

[Ano13f] Anonymous. myComputer.

IEEE Computer Graphics and Applications, 33(6):c2, November/December 2013. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2013:TC

[Ano13g]  Anonymous.  Table of

contents. IEEE Computer Graphics and Applications, 33(6):2–3, November/ December 2013. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2014:AUBa

[Ano14a] Anonymous. Appropriate use of bibliometric indicators for the assessment of journals, research proposals, and individuals. IEEE Computer Graphics and Applications, 34(2):87–88, March/ April 2014. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2014:AUBb

[Ano14b] Anonymous. Appropriate use of bibliometric indicators for the assessment of journals, research proposals, and individuals. IEEE Computer Graphics and Applications, 34(3):87–88, May/ June 2014. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2014:FYJ

[Ano14c]  Anonymous.  Focus on

your job search [advertisement]. IEEE Computer Graphics and Applications, 34(5):35, September/October 2014. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://csdl.
computer.org/csdl/mags/
cg/2014/05/mcg2014050035.
pdf.

Anonymous:2014:FLM

[Ano14d] Anonymous. A fresh look at monitors. IEEE Computer Graphics and Applications, 34(6):88–c3, November/ December 2014. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://csdl.
computer.org/csdl/mags/
cg/2014/06/mcg2014060088.
html.

Anonymous:2014:FCa

[Ano14e] Anonymous. Front cover.

IEEE Computer Graphics and Applications, 34(1): c1, January/February 2014. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2014:FCb

[Ano14f] Anonymous. Front cover.

IEEE Computer Graphics and Applications, 34(2):c1, March/April 2014. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2014:FCc

[Ano14g] Anonymous. [Front cover].

IEEE Computer Graphics and Applications, 34(3):c1, May/June 2014. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2014:FCd

[Ano14h] Anonymous. [Front cover]. IEEE Computer Graphics and Applications, 34(5):c1, September/October 2014. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL

                                        http://csdl.computer.org/
                                        csdl/mags/cg/2014/05/mcg20140500c1.
                                        pdf.

Anonymous:2014:FCe

[Ano14i] Anonymous. [Front cover]. IEEE Computer Graphics and Applications, 34(6):c1, November/December 2014. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL

                                        http://csdl.computer.org/
                                        csdl/mags/cg/2014/06/mcg20140600c1.
                                        pdf.

Anonymous:2014:ICCc

[Ano14j] Anonymous. IEEE CG&A call for articles. IEEE Computer Graphics and Applications, 34(5):c2, September/October 2014. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL http://csdl.
                                        computer.org/csdl/mags/
                                        cg/2014/05/mcg20140500c2.
                                        pdf.

Anonymous:2014:ICCe

[Ano14k] Anonymous. IEEE CG&A call for articles. IEEE Computer Graphics and Applications, 34(6):c2, November/ December 2014. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://csdl.
computer.org/csdl/mags/
cg/2014/06/mcg20140600c2.
pdf.

Anonymous:2014:ICCf

[Ano14l] Anonymous. IEEE CG&A call for articles. IEEE Computer Graphics and Applications, 34(6):c4, November/ December 2014. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://csdl.
computer.org/csdl/mags/
cg/2014/06/mcg20140600c4.
pdf.

Anonymous:2014:ICCd

[Ano14m] Anonymous. IEEE CG&A calls for articles. IEEE Computer Graphics and Applications, 34(5):25, September/October 2014. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://csdl.
computer.org/csdl/mags/
cg/2014/05/mcg2014050025.
pdf.

Anonymous:2014:ICCa

[Ano14n] Anonymous. IEEE CG&A calls for papers. IEEE Computer Graphics and Applications, 34(2):57, March/ April 2014. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2014:ICCb

[Ano14o] Anonymous. IEEE CG&A calls for papers. IEEE Computer Graphics and Applications, 34(3):19, May/ June 2014. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2014:ICSc

[Ano14p] Anonymous. IEEE Computer Society — rock stars of mobile cloud [future event]. IEEE Computer Graphics and Applications, 34(2):c4, March/April 2014. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2014:ICSb

[Ano14q] Anonymous. IEEE Computer Society [advertisement]. IEEE Computer Graphics and Applications, 34(2):c3, March/April 2014. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2014:ICSd

[Ano14r] Anonymous. IEEE Computer Society [advertisement]. IEEE Computer Graphics and Applications, 34(4):c2, July/August 2014. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2014:ICSa

[Ano14s] Anonymous. IEEE Computer Society house advertisement. IEEE Computer Graphics and Applications, 34(1):c2, January/ February 2014. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2014:ISP

[Ano14t]  Anonymous.  IEEE Se-

curity & Privacy [advertisement]. IEEE Computer Graphics and Applications, 34(5):41, September/October 2014. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://csdl.
computer.org/csdl/mags/
cg/2014/05/mcg2014050041.
pdf.

Anonymous:2014:Ma

[Ano14u]  Anonymous.  Masthead.

IEEE Computer Graphics and Applications, 34(1): 1, January/February 2014. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2014:Mb

[Ano14v]  Anonymous.  Masthead.

IEEE Computer Graphics and Applications, 34(2):1, March/April 2014. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2014:Mc

[Ano14w]  Anonymous.  Masthead.

IEEE Computer Graphics and Applications, 34(3):1, May/June 2014. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2014:Md

[Ano14x] Anonymous. [Masthead]. IEEE Computer Graphics and Applications, 34(4):1, July/August 2014. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2014:Me

[Ano14y] Anonymous. [Masthead]. IEEE Computer Graphics and Applications, 34(5):1, September/October 2014. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL

                                        http://csdl.computer.org/
                                        csdl/mags/cg/2014/05/mcg2014050001.
                                        pdf.

Anonymous:2014:Mf

[Ano14z]  Anonymous.  Masthead.

IEEE Computer Graphics and Applications, 34(6):1, November/December 2014. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL

http://csdl.computer.org/
csdl/mags/cg/2014/06/mcg2014060001.
pdf.

Anonymous:2014:MMAa

[Ano14-27]  Anonymous.  Member-

ship matters [advertisement]. IEEE Computer Graphics and Applications, 34(2):c2, March/April 2014. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2014:MMAb

[Ano14-28] Anonymous. Membership matters [advertisement]. IEEE Computer Graphics and Applications, 34(5):c3, September/October 2014. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL

http://csdl.computer.org/
csdl/mags/cg/2014/05/mcg20140500c3.
pdf.

Anonymous:2014:MMH

[Ano14-29] Anonymous. Membership matters house advertisement. IEEE Computer Graphics and Applications, 34(1):c4, January/February 2014. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2014:RSB

[Ano14-30] Anonymous. Rock stars of big data analytics [advertisement]. IEEE Computer Graphics and Applications, 34(5):c4, September/October 2014. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://csdl.
computer.org/csdl/mags/
cg/2014/05/mcg20140500c4.
pdf.

Anonymous:2014:RSCa

[Ano14-31] Anonymous. Rock stars of cybersecurity [advertisement]. IEEE Computer Graphics and Applications, 34(4):c4, July/August 2014. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2014:RSCb

[Ano14-32] Anonymous. Rock stars of cybersecurity [advertisement]. IEEE Computer Graphics and Applications, 34(5):7, September/ October 2014. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://csdl.
computer.org/csdl/mags/
cg/2014/05/mcg2014050007.
pdf.

Anonymous:2014:STC

[Ano14-33] Anonymous. Special thanks to CG&A’s reviewers. IEEE Computer Graphics and Applications, 34(1):9, January/ February 2014. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2014:TCa

[Ano14-34]  Anonymous.  Table of

contents. IEEE Computer Graphics and Applications, 34(1):2–3, January/ February 2014. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2014:TCb

[Ano14-35]  Anonymous.  Table of

contents. IEEE Computer Graphics and Applications, 34(2):2–3, March/ April 2014. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2014:TCc

[Ano14-36] Anonymous. Table of contents. IEEE Computer Graphics and Applications, 34(3):2–3, May/June 2014. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2014:TCd

[Ano14-37]  Anonymous.  Table of

contents. IEEE Computer Graphics and Applications, 34(4):2–3, July/ August 2014. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2014:TCe

[Ano14-38]  Anonymous.  Table of

contents. IEEE Computer Graphics and Applications, 34(5):2–3, September/October 2014. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://csdl.
computer.org/csdl/mags/
cg/2014/05/mcg2014050002.
pdf.

Anonymous:2014:TCL

[Ano14-39] Anonymous. Take the CS Library wherever you go! [advertisement]. IEEE Computer Graphics and Applications, 34(5):51, September/October 2014. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://csdl.
computer.org/csdl/mags/
cg/2014/05/mcg2014050051.
pdf.

Anonymous:2015:R

[Ano15a]  Anonymous.  2014 re-

viewers. IEEE Computer Graphics and Applications, 35(1):19, January/ February 2015. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://csdl.
computer.org/csdl/mags/
cg/2015/01/mcg2015010019.
html.

Anonymous:2015:ADT

[Ano15b] Anonymous. Advanced display technologies. IEEE Computer Graphics and Applications, 35(1):96, January/February 2015. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

csdl.computer.org/csdl/
mags/cg/2015/01/mcg2015010096-
abs.html.

Anonymous:2015:DOH

[Ano15c]  Anonymous.  Deadline:

31 October house advertisement. IEEE Computer Graphics and Applications, 35(5):c4, September/October 2015. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
cg/2015/05/mcg20150500c4.
pdf.

Anonymous:2015:FYJa

[Ano15d]  Anonymous.  Focus on

your job search [advertisement]. IEEE Computer Graphics and Applications, 35(1):c3, January/ February 2015. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://csdl.
computer.org/csdl/mags/
cg/2015/01/mcg20150100c3.
pdf.

Anonymous:2015:FYJb

[Ano15e] Anonymous. Focus on your job search [Advertisement]. IEEE Computer Graphics and Applications, 35(4):13, July/August 2015. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL http://www.
                                        computer.org/csdl/mags/
                                        cg/2015/04/mcg2015040013.
                                        pdf.

Anonymous:2015:FCa

[Ano15f] Anonymous. [Front cover]. IEEE Computer Graphics and Applications, 35(1): c1, January/February 2015. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL

                                        http://csdl.computer.org/
                                        csdl/mags/cg/2015/01/mcg20150100c1.
                                        pdf.

Anonymous:2015:FCb

[Ano15g] Anonymous. Front cover. IEEE Computer Graphics and Applications, 35(2):c1, March/April 2015. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL http://www.
                                        computer.org/csdl/mags/
                                        cg/2015/02/mcg20150200c1.
                                        pdf.

Anonymous:2015:FCc

[Ano15h] Anonymous. Front cover. IEEE Computer Graphics and Applications, 35(3):c1, May/June 2015. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL http://www.
                                        computer.org/csdl/mags/
                                        cg/2015/03/mcg20150300c1.
                                        pdf.

Anonymous:2015:FCd

[Ano15i] Anonymous. [Front cover].

IEEE Computer Graphics and Applications, 35(4):c1, July/August 2015. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
cg/2015/04/mcg20150400c1.
pdf.

Anonymous:2015:FCe

[Ano15j] Anonymous. Front cover.

IEEE Computer Graphics and Applications, 35(5):c1, September/October 2015. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL

http://www.computer.org/
csdl/mags/cg/2015/05/mcg20150500c1.
pdf.

Anonymous:2015:FCf

[Ano15k] Anonymous. [Front cover].

IEEE Computer Graphics and Applications, 35(6):c1, November/December 2015. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL

http://www.computer.org/
csdl/mags/cg/2015/06/mcg20150600c1.
pdf.

Anonymous:2015:GRY

[Ano15l] Anonymous. Get the recognition you deserve [advertisement]. IEEE Computer Graphics and Applications, 35(6):c3, November/ December 2015. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
cg/2015/06/mcg20150600c3.
pdf.

Anonymous:2015:HTH

[Ano15m]  Anonymous.  High tech

@ home. IEEE Computer Graphics and Applications, 35(3):8–9, May/ June 2015. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
cg/2015/03/mcg2015030008.
html.

Anonymous:2015:ICCa

[Ano15n] Anonymous. IEEE CG&A call for articles. IEEE Computer Graphics and Applications, 35(1):c2, January/ February 2015. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://csdl.
computer.org/csdl/mags/
cg/2015/01/mcg20150100c2.
pdf.

Anonymous:2015:ICCb

[Ano15o] Anonymous. IEEE CG&A call for articles. IEEE Computer Graphics and Applications, 35(2):15, March/ April 2015. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://www.

computer.org/csdl/mags/
cg/2015/02/mcg2015020015.
pdf.

Anonymous:2015:ICCc

[Ano15p] Anonymous. IEEE CG&A call for articles. IEEE Computer Graphics and Applications, 35(2):c2, March/ April 2015. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
cg/2015/02/mcg20150200c2.
pdf.

Anonymous:2015:ICCf

[Ano15q] Anonymous. IEEE CG&A call for articles. IEEE Computer Graphics and Applications, 35(4):c2, July/ August 2015. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
cg/2015/04/mcg20150400c2.
pdf.

Anonymous:2015:ICCh

[Ano15r] Anonymous. IEEE CG&A call for articles. IEEE Computer Graphics and Applications, 35(6):29, November/ December 2015. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
cg/2015/06/mcg2015060029.
pdf.

Anonymous:2015:ICCi

[Ano15s] Anonymous. IEEE CG&A call for articles. IEEE Computer Graphics and Applications, 35(6):41, November/ December 2015. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL http://www.
                                        computer.org/csdl/mags/
                                        cg/2015/06/mcg2015060041.
                                        pdf.

Anonymous:2015:ICCe

[Ano15t] Anonymous. IEEE CG&A call for articles house advertisement. IEEE Computer Graphics and Applications, 35(3):c2, May/ June 2015. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL http://www.
                                        computer.org/csdl/mags/
                                        cg/2015/03/mcg20150300c2.
                                        pdf.

Anonymous:2015:ICCg

[Ano15u] Anonymous. IEEE CG&A call for articles: House advertisement. IEEE Computer Graphics and Applications, 35(5):91, September/October 2015. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL http://www.
                                        computer.org/csdl/mags/
                                        cg/2015/05/mcg2015050091.
                                        pdf.

Anonymous:2015:ICCd

[Ano15v] Anonymous. IEEE Cloud Computing call for papers. IEEE Computer Graphics and Applications, 35(2):c3, March/April 2015. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
cg/2015/02/mcg20150200c3.
pdf.

Anonymous:2015:ICSa

[Ano15w] Anonymous. IEEE Computer Society [advertisement]. IEEE Computer Graphics and Applications, 35(1):82, January/February 2015. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL

http://csdl.computer.org/
csdl/mags/cg/2015/01/mcg2015010082.
pdf.

Anonymous:2015:ICSc

[Ano15x] Anonymous. IEEE Computer Society [Advertisement]. IEEE Computer Graphics and Applications, 35(4):107, July/August 2015. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL

http://www.computer.org/
csdl/mags/cg/2015/04/mcg2015040107.
pdf.

Anonymous:2015:ICSb

[Ano15y] Anonymous. IEEE Computer Society house advertisement. IEEE Computer Graphics and Applications, 35(2):79, March/ April 2015. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
cg/2015/02/mcg2015020079.
pdf.

Anonymous:2015:Ma

[Ano15z] Anonymous. [Masthead]. IEEE Computer Graphics and Applications, 35(1): 1, January/February 2015. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL

http://csdl.computer.org/
csdl/mags/cg/2015/01/mcg2015010001.
pdf.

Anonymous:2015:Mb

[Ano15-27]  Anonymous.  Masthead.

IEEE Computer Graphics and Applications, 35(2):1, March/April 2015. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
cg/2015/02/mcg2015020001.
pdf.

Anonymous:2015:Mc

[Ano15-28]  Anonymous.  Masthead.

IEEE Computer Graphics and Applications, 35(3):1, May/June 2015. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://www.

computer.org/csdl/mags/
cg/2015/03/mcg2015030001.
pdf.

Anonymous:2015:Mh

[Ano15-29]  Anonymous.  Masthead.

IEEE Computer Graphics and Applications, 35(3):1, May/June 2015. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
cg/2015/03/mcg2015030001.
pdf.

Anonymous:2015:Md

[Ano15-30] Anonymous. [Masthead].

IEEE Computer Graphics and Applications, 35(4):1, July/August 2015. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
cg/2015/04/mcg2015040001.
pdf.

Anonymous:2015:Me

[Ano15-31]  Anonymous.  Masthead.

IEEE Computer Graphics and Applications, 35(5):4, September/October 2015. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL

http://www.computer.org/
csdl/mags/cg/2015/05/mcg2015050004.
pdf.

Anonymous:2015:Mf

[Ano15-32] Anonymous. [Masthead].

IEEE Computer Graphics and Applications, 35(6):5, November/December 2015. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL

http://www.computer.org/
csdl/mags/cg/2015/06/mcg2015060005.
pdf.

Anonymous:2015:RSP

[Ano15-33] Anonymous. Rock stars of 3D printing [advertisement]. IEEE Computer Graphics and Applications, 35(1):c4, January/ February 2015. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://csdl.
computer.org/csdl/mags/
cg/2015/01/mcg20150100c4.
pdf.

Anonymous:2015:RSCb

[Ano15-34] Anonymous. Rock stars of cybersecurity [Advertisement]. IEEE Computer Graphics and Applications, 35(4):c4, July/ August 2015. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
cg/2015/04/mcg20150400c4.
pdf.

Anonymous:2015:RSCa

[Ano15-35] Anonymous. Rock stars of cybersecurity house advertisement. IEEE Computer Graphics and Applications, 35(3):c4, May/ June 2015. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
cg/2015/03/mcg20150300c4.
pdf.

Anonymous:2015:RSW

[Ano15-36] Anonymous. Rock stars of wearables [Advertisement]. IEEE Computer Graphics and Applications, 35(4):63, July/August 2015. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
cg/2015/04/mcg2015040063.
pdf.

Anonymous:2015:SA

[Ano15-37] Anonymous. SIGGRAPH2015 advertisement. IEEE Computer Graphics and Applications, 35(3):7, May/ June 2015. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
cg/2015/03/mcg2015030007.
pdf.

Anonymous:2015:SXD

[Ano15-38] Anonymous. SIGGRAPH2015 xroads of discovery house advertisement. IEEE Computer Graphics and Applications, 35(2):c4, March/ April 2015. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL http://www.
                                        computer.org/csdl/mags/
                                        cg/2015/02/mcg20150200c4.
                                        pdf.

Anonymous:2015:SCA

[Ano15-39] Anonymous. Stay connected [advertisement]. IEEE Computer Graphics and Applications, 35(1):83, January/ February 2015. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL http://csdl.
                                        computer.org/csdl/mags/
                                        cg/2015/01/mcg2015010083.
                                        pdf.

Anonymous:2015:TCa

[Ano15-40]  Anonymous.  Table of

contents. IEEE Computer Graphics and Applications, 35(1):2–3, January/ February 2015. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL http://csdl.
                                        computer.org/csdl/mags/
                                        cg/2015/01/mcg2015010002.
                                        pdf.

Anonymous:2015:TCb

[Ano15-41]  Anonymous.  Table of

contents. IEEE Computer Graphics and Applications, 35(4):2–3, July/ August 2015. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL http://www.
                                        computer.org/csdl/mags/
                                        cg/2015/04/mcg2015040002.
                                        pdf.

Anonymous:2015:TCc

[Ano15-42]  Anonymous.  Table of

contents. IEEE Computer Graphics and Applications, 35(6):c2–1, November/December 2015. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http:/

/www.computer.org/csdl/
mags/cg/2015/06/mcg20150600c2.
pdf.

Anonymous:2015:TCL

[Ano15-43]  Anonymous.  Take the

CS Library wherever you go! [Advertisement]. IEEE Computer Graphics and Applications, 35(4):c3, July/ August 2015. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
cg/2015/04/mcg20150400c3.
pdf.

Anonymous:2015:WWL

[Ano15-44] Anonymous. Watch the world’s leading experts take multi-core strategies to new heights [Advertisement]. IEEE Computer Graphics and Applications, 35(6):c4, November/ December 2015. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
cg/2015/06/mcg20150600c4.
pdf.

Anonymous:2015:WR

[Ano15-45] Anonymous. The wearable revolution. IEEE Computer Graphics and Applications, 35(2):104, March/ April 2015. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL http://www.
                                        computer.org/csdl/mags/
                                        cg/2015/02/mcg2015020104-
                                        abs.html.

Anonymous:2016:ADA

[Ano16a] Anonymous. Art in the digital age. IEEE Computer Graphics and Applications, 36(4):6–7, July/ August 2016. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        cg/2016/04/mcg2016040006-
                                        abs.html.

Anonymous:2016:AYR

[Ano16b]  Anonymous.  Augment

your reality. IEEE Computer Graphics and Applications, 36(1):6–7, January/ February 2016. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL http://www.
                                        computer.org/csdl/mags/
                                        cg/2016/01/mcg2016010006.
                                        html.

Anonymous:2016:CNE

[Ano16c] Anonymous. Call for nominees education awards nominations. IEEE Computer Graphics and Applications, 36(5):61, September/October 2016. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2016/05/mcg2016050061.
pdf.

Anonymous:2016:CPC

[Ano16d] Anonymous. CiSE is the perfect combination house advertisement. IEEE Computer Graphics and Applications, 36(1):29, January/ February 2016. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
cg/2016/01/mcg2016010029.
pdf.

Anonymous:2016:CEA

[Ano16e] Anonymous. Computer entrepreneur award. IEEE Computer Graphics and Applications, 36(5):94, September/October 2016. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2016/05/mcg2016050094.
pdf.

Anonymous:2016:CGB

[Ano16f] Anonymous. A computer graphics back-to-school special. IEEE Computer Graphics and Applications, 36(5):95–96, September/ October 2016. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        cg/2016/05/mcg2016050095.
                                        html.

Anonymous:2016:FYJ

[Ano16g]  Anonymous.  Focus on

your job search house advertisement. IEEE Computer Graphics and Applications, 36(6):c4, November/ December 2016. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        cg/2016/06/mcg20160600c4.
                                        pdf.

Anonymous:2016:FCa

[Ano16h] Anonymous. Front cover. IEEE Computer Graphics and Applications, 36(1): c1, January/February 2016. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL

                                        http://www.computer.org/
                                        csdl/mags/cg/2016/01/mcg20160100c1.
                                        pdf.

Anonymous:2016:FCb

[Ano16i] Anonymous. Front cover. IEEE Computer Graphics and Applications, 36(2):c1, March/April 2016. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/

cg/2016/02/mcg20160200c1.
pdf.

Anonymous:2016:FCc

[Ano16j] Anonymous. Front cover.

IEEE Computer Graphics and Applications, 36(3):c1, May/June 2016. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2016/03/mcg20160300c1.
pdf.

Anonymous:2016:FCd

[Ano16k] Anonymous. Front cover.

IEEE Computer Graphics and Applications, 36(4):c1, July/August 2016. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2016/04/mcg20160400c1.
pdf.

Anonymous:2016:FCe

[Ano16l] Anonymous. Front cover.

IEEE Computer Graphics and Applications, 36(5):c1, September/October 2016. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL

https://www.computer.org/
csdl/mags/cg/2016/05/mcg20160500c1.
pdf.

Anonymous:2016:FCf

[Ano16m] Anonymous. Front cover.

IEEE Computer Graphics and Applications, 36(6):c1, November/December 2016. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL

https://www.computer.org/
csdl/mags/cg/2016/06/mcg20160600c1.
pdf.

Anonymous:2016:ICCe

[Ano16n] Anonymous. IEEE CG&A call for articles. IEEE Computer Graphics and Applications, 36(4):91, July/ August 2016. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2016/04/mcg2016040091.
pdf.

Anonymous:2016:ICCf

[Ano16o] Anonymous. IEEE CG&A call for articles. IEEE Computer Graphics and Applications, 36(4):c2, July/ August 2016. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2016/04/mcg20160400c2.
pdf.

Anonymous:2016:ICCg

[Ano16p] Anonymous. IEEE CG&A call for articles. IEEE Computer Graphics and Applications, 36(5):27, September/October 2016. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.

computer.org/csdl/mags/
cg/2016/05/mcg2016050027.
pdf.

Anonymous:2016:ICCh

[Ano16q] Anonymous. IEEE CG&A call for articles. IEEE Computer Graphics and Applications, 36(5):c2, September/October 2016. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2016/05/mcg20160500c2.
pdf.

Anonymous:2016:ICCa

[Ano16r] Anonymous. IEEE CG&A call for articles house advertisement. IEEE Computer Graphics and Applications, 36(1):15, January/ February 2016. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
cg/2016/01/mcg2016010015.
pdf.

Anonymous:2016:ICCb

[Ano16s] Anonymous. IEEE CG&A call for articles house advertisement. IEEE Computer Graphics and Applications, 36(2):c2, March/ April 2016. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/

                                        cg/2016/02/mcg20160200c2.
                                        pdf.

Anonymous:2016:ICCc

[Ano16t] Anonymous. IEEE CG&A call for articles house advertisement. IEEE Computer Graphics and Applications, 36(3):59, May/ June 2016. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        cg/2016/03/mcg2016030059.
                                        pdf.

Anonymous:2016:ICCd

[Ano16u] Anonymous. IEEE CG&A call for articles house advertisement. IEEE Computer Graphics and Applications, 36(3):c2, May/ June 2016. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        cg/2016/03/mcg20160300c2.
                                        pdf.

Anonymous:2016:ICCi

[Ano16v] Anonymous. IEEE CG&A call for articles house advertisement. IEEE Computer Graphics and Applications, 36(6):c2, November/ December 2016. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/

cg/2016/06/mcg20160600c2.
pdf.

Anonymous:2016:ICSd

[Ano16w] Anonymous. IEEE Computer Society. IEEE Computer Graphics and Applications, 36(4):c3, July/ August 2016. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2016/04/mcg20160400c3.
pdf.

Anonymous:2016:ICSb

[Ano16x] Anonymous. IEEE Computer Society: Be at the center of it all house advertisement. IEEE Computer Graphics and Applications, 36(1):c4, January/ February 2016. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
cg/2016/01/mcg20160100c4.
pdf.

Anonymous:2016:ICSa

[Ano16y] Anonymous. IEEE Computer Society house advertisement. IEEE Computer Graphics and Applications, 36(1):79, January/ February 2016. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://www.
computer.org/csdl/mags/

                                        cg/2016/01/mcg2016010079.
                                        pdf.

Anonymous:2016:ICSc

[Ano16z] Anonymous. IEEE Computer Society house advertisement. IEEE Computer Graphics and Applications, 36(3):83, May/ June 2016. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        cg/2016/03/mcg2016030083.
                                        pdf.

Anonymous:2016:ICSe

[Ano16-27] Anonymous. IEEE Computer Society house advertisement. IEEE Computer Graphics and Applications, 36(6):96, November/ December 2016. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        cg/2016/06/mcg2016060096.
                                        pdf.

Anonymous:2016:ITB

[Ano16-28] Anonymous. IEEE Transactions on Big Data house advertisement. IEEE Computer Graphics and Applications, 36(1):91, January/ February 2016. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL http://www.
                                        computer.org/csdl/mags/

cg/2016/01/mcg2016010091.
pdf.

Anonymous:2016:ITS

[Ano16-29] Anonymous. IEEE Transactions on Sustainable Computing house advertisement. IEEE Computer Graphics and Applications, 36(1): c3, January/February 2016. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL

http://www.computer.org/
csdl/mags/cg/2016/01/mcg20160100c3.
pdf.

Anonymous:2016:Ma

[Ano16-30]  Anonymous.  Masthead.

IEEE Computer Graphics and Applications, 36(1): 5, January/February 2016. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL

http://www.computer.org/
csdl/mags/cg/2016/01/mcg2016010005.
pdf.

Anonymous:2016:Mb

[Ano16-31]  Anonymous.  Masthead.

IEEE Computer Graphics and Applications, 36(2):c3, March/April 2016. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2016/02/mcg20160200c3.
pdf.

Anonymous:2016:Mc

[Ano16-32]  Anonymous.  Masthead.

IEEE Computer Graphics and Applications, 36(3):c3, May/June 2016. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        cg/2016/03/mcg20160300c3.
                                        pdf.

Anonymous:2016:Md

[Ano16-33]  Anonymous.  Masthead.

IEEE Computer Graphics and Applications, 36(4):1, July/August 2016. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        cg/2016/04/mcg2016040001.
                                        pdf.

Anonymous:2016:Me

[Ano16-34]  Anonymous.  Masthead.

IEEE Computer Graphics and Applications, 36(5):1, September/October 2016. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL

                                        https://www.computer.org/
                                        csdl/mags/cg/2016/05/mcg2016050001.
                                        pdf.

Anonymous:2016:Mf

[Ano16-35]  Anonymous.  Masthead.

IEEE Computer Graphics and Applications, 36(6):1, November/December 2016. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL

https://www.computer.org/
csdl/mags/cg/2016/06/mcg2016060001.
pdf.

Anonymous:2016:MHA

[Ano16-36] Anonymous. myCS house advertisement. IEEE Computer Graphics and Applications, 36(6):69, November/ December 2016. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2016/06/mcg2016060069.
pdf.

Anonymous:2016:NMOa

[Ano16-37] Anonymous. New membership options for a better fit. IEEE Computer Graphics and Applications, 36(5):c3, September/October 2016. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2016/05/mcg20160500c3.
pdf.

Anonymous:2016:NMOb

[Ano16-38] Anonymous. New membership options for a better fit house advertisement. IEEE Computer Graphics and Applications, 36(6):c3, November/December 2016. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL https:/

                                        /www.computer.org/csdl/
                                        mags/cg/2016/06/mcg20160600c3.
                                        pdf.

Anonymous:2016:SHA

[Ano16-39] Anonymous. Not so secure house advertisement. IEEE Computer Graphics and Applications, 36(6):57, November/December 2016. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL https:/

                                        /www.computer.org/csdl/
                                        mags/cg/2016/06/mcg2016060057.
                                        pdf.

Anonymous:2016:RSBb

[Ano16-40] Anonymous. Rock stars of big data. IEEE Computer Graphics and Applications, 36(4):c4, July/ August 2016. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        cg/2016/04/mcg20160400c4.
                                        pdf.

Anonymous:2016:RSBc

[Ano16-41] Anonymous. Rock stars of big data. IEEE Computer Graphics and Applications, 36(5):c4, September/October 2016. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        cg/2016/05/mcg20160500c4.
                                        pdf.

Anonymous:2016:RSBa

[Ano16-42] Anonymous. Rock stars of big data house advertisement. IEEE Computer Graphics and Applications, 36(3):c4, May/ June 2016. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2016/03/mcg20160300c4.
pdf.

Anonymous:2016:RSR

[Ano16-43] Anonymous. Rock stars of risk-based security house advertisement. IEEE Computer Graphics and Applications, 36(2):c4, March/ April 2016. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2016/02/mcg20160200c4.
pdf.

Anonymous:2016:TCb

[Ano16-44]  Anonymous.  Table of

contents. IEEE Computer Graphics and Applications, 36(2):1–2, March/ April 2016. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2016/02/mcg2016020001.
pdf.

Anonymous:2016:TCc

[Ano16-45]  Anonymous.  Table of

contents. IEEE Computer Graphics and Applications, 36(3):1–2, May/ June 2016. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        cg/2016/03/mcg2016030001.
                                        pdf.

Anonymous:2016:TCd

[Ano16-46]  Anonymous.  Table of

contents. IEEE Computer Graphics and Applications, 36(4):2–3, July/ August 2016. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        cg/2016/04/mcg2016040002.
                                        pdf.

Anonymous:2016:TCe

[Ano16-47]  Anonymous.  Table of

contents. IEEE Computer Graphics and Applications, 36(5):2–3, September/October 2016. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        cg/2016/05/mcg2016050002.
                                        pdf.

Anonymous:2016:TCf

[Ano16-48]  Anonymous.  Table of

contents. IEEE Computer Graphics and Applications, 36(6):2–3, November/ December 2016. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2016/06/mcg2016060002.
pdf.

Anonymous:2017:CNE

[Ano17a]  Anonymous.  Call for

nominees education awards nominations house advertisement. IEEE Computer Graphics and Applications, 37(2):c3, March/ April 2017. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2017/02/mcg20170200c3.
pdf.

Anonymous:2017:CYIa

[Ano17b] Anonymous. Can you invent a better world through technologies? IEEE Computer Graphics and Applications, 37(1):c3, January/ February 2017. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2017/01/mcg20170100c3.
pdf.

Anonymous:2017:CYIb

[Ano17c] Anonymous. Can you invent a better world through technologies? house advertisement. IEEE Computer Graphics and Applications, 37(2):11, March/ April 2017. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        cg/2017/02/mcg2017020011.
                                        pdf.

Anonymous:2017:FCa

[Ano17d] Anonymous. [Front cover]. IEEE Computer Graphics and Applications, 37(1): c1, January/February 2017. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL

                                        https://www.computer.org/
                                        csdl/mags/cg/2017/01/mcg20170100c1.
                                        pdf.

Anonymous:2017:FCb

[Ano17e] Anonymous. Front cover. IEEE Computer Graphics and Applications, 37(2):c1, March/April 2017. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        cg/2017/02/mcg20170200c1.
                                        pdf.

Anonymous:2017:FCc

[Ano17f] Anonymous. Front cover. IEEE Computer Graphics and Applications, 37(3):c1, May/June 2017. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL https://www.

computer.org/csdl/mags/
cg/2017/03/mcg20170300c1.
pdf.

Anonymous:2017:FCd

[Ano17g] Anonymous. Front cover.

IEEE Computer Graphics and Applications, 37(4):c1, July/August 2017. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2017/04/mcg20170400c1.
pdf.

Anonymous:2017:FCe

[Ano17h] Anonymous. Front cover.

IEEE Computer Graphics and Applications, 37(5):c1, September/October 2017. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL

https://www.computer.org/
csdl/mags/cg/2017/05/mcg20170500c1.
pdf.

Anonymous:2017:FCf

[Ano17i] Anonymous. [Front cover].

IEEE Computer Graphics and Applications, 37(6):c1, November/December 2017. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL

https://www.computer.org/
csdl/mags/cg/2017/06/mcg20170600c1.
pdf.

Anonymous:2017:ICCb

[Ano17j] Anonymous. IEEE CG&A call for articles. IEEE Computer Graphics and Applications, 37(3):c2, May/ June 2017. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2017/03/mcg20170300c2.
pdf.

Anonymous:2017:ICCc

[Ano17k] Anonymous. IEEE CG&A call for articles. IEEE Computer Graphics and Applications, 37(3):c3, May/ June 2017. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2017/03/mcg20170300c3.
pdf.

Anonymous:2017:ICCd

[Ano17l] Anonymous. IEEE CGA call for articles. IEEE Computer Graphics and Applications, 37(4):c2, July/ August 2017. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2017/04/mcg20170400c2.
pdf.

Anonymous:2017:ICCe

[Ano17m] Anonymous. IEEE CGA call for articles. IEEE Computer Graphics and Applications, 37(4):c3, July/ August 2017. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2017/04/mcg20170400c3.
pdf.

Anonymous:2017:ICCf

[Ano17n] Anonymous. IEEE CGA call for articles. IEEE Computer Graphics and Applications, 37(5):c3, September/October 2017. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2017/05/mcg20170500c3.
pdf.

Anonymous:2017:ICCg

[Ano17o] Anonymous. IEEE CG&A call for articles. IEEE Computer Graphics and Applications, 37(6):c3, November/ December 2017. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2017/06/mcg20170600c3.
pdf.

Anonymous:2017:ICCa

[Ano17p] Anonymous. IEEE CG&A call for articles house advertisement. IEEE Computer Graphics and Applications, 37(2):c2, March/ April 2017. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/

                                        cg/2017/02/mcg20170200c2.
                                        pdf.

Anonymous:2017:ICS

[Ano17q] Anonymous. IEEE Computer Society. IEEE Computer Graphics and Applications, 37(4):94, July/ August 2017. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        cg/2017/04/mcg2017040094.
                                        pdf.

Anonymous:2017:Ma

[Ano17r]  Anonymous.  Masthead.

IEEE Computer Graphics and Applications, 37(1): 1, January/February 2017. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL

                                        https://www.computer.org/
                                        csdl/mags/cg/2017/01/mcg2017010001.
                                        pdf.

Anonymous:2017:Mc

[Ano17s]  Anonymous.  Masthead.

IEEE Computer Graphics and Applications, 37(2):1, March/April 2017. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        cg/2017/02/mcg2017020001.
                                        pdf.

Anonymous:2017:Md

[Ano17t]  Anonymous.  Masthead.

IEEE Computer Graphics and Applications, 37(3):1, May/June 2017. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2017/03/mcg2017030001.
pdf.

Anonymous:2017:Me

[Ano17u]  Anonymous.  Masthead.

IEEE Computer Graphics and Applications, 37(4):1, July/August 2017. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2017/04/mcg2017040001.
pdf.

Anonymous:2017:Mf

[Ano17v]  Anonymous.  Masthead.

IEEE Computer Graphics and Applications, 37(5):c2, September/October 2017. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL

https://www.computer.org/
csdl/mags/cg/2017/05/mcg20170500c2.
pdf.

Anonymous:2017:Mh

[Ano17w] Anonymous. [Masthead].

IEEE Computer Graphics and Applications, 37(6):c2, November/December 2017. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL

https://www.computer.org/

                                        csdl/mags/cg/2017/06/mcg20170600c2.
                                        pdf.

Anonymous:2017:Mb

[Ano17x] Anonymous. myCS. IEEE Computer Graphics and Applications, 37(1):79, January/February 2017. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL https:/

                                        /www.computer.org/csdl/
                                        mags/cg/2017/01/mcg2017010079.
                                        pdf.

Anonymous:2017:Mg

[Ano17y] Anonymous. myCS. IEEE Computer Graphics and Applications, 37(6):102, November/December 2017. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL https:/

                                        /www.computer.org/csdl/
                                        mags/cg/2017/06/mcg2017060102.
                                        pdf.

Anonymous:2017:NMO

[Ano17z] Anonymous. New membership options for a better fit. IEEE Computer Graphics and Applications, 37(1):69, January/ February 2017. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        cg/2017/01/mcg2017010069.
                                        pdf.

Anonymous:2017:SHA

[Ano17-27] Anonymous. Not so secure house advertisement. IEEE Computer Graphics and Applications, 37(2):19, March/ April 2017. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2017/02/mcg2017020019.
pdf.

Anonymous:2017:PCa

[Ano17-28] Anonymous. Prepare to connect. IEEE Computer Graphics and Applications, 37(3):c4, May/ June 2017. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2017/03/mcg20170300c4.
pdf.

Anonymous:2017:PCb

[Ano17-29] Anonymous. Prepare to connect. IEEE Computer Graphics and Applications, 37(5):c4, September/October 2017. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2017/05/mcg20170500c4.
pdf.

Anonymous:2017:PCc

[Ano17-30] Anonymous. Prepare to connect. IEEE Computer Graphics and Applications, 37(6):c4, November/ December 2017. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2017/06/mcg20170600c4.
pdf.

Anonymous:2017:SCS

[Ano17-31] Anonymous. Seymour Cray, Sidney Fernbach, & Ken Kennedy Awards. IEEE Computer Graphics and Applications, 37(1):c4, January/February 2017. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL https:/

/www.computer.org/csdl/
mags/cg/2017/01/mcg20170100c4.
pdf.

Anonymous:2017:TCa

[Ano17-32]  Anonymous.  Table of

contents. IEEE Computer Graphics and Applications, 37(1):2–3, January/ February 2017. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2017/01/mcg2017010002.
pdf.

Anonymous:2017:TCb

[Ano17-33]  Anonymous.  Table of

contents. IEEE Computer Graphics and Applications, 37(2):2–3, March/ April 2017. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2017/02/mcg2017020002.
pdf.

Anonymous:2017:TCc

[Ano17-34]  Anonymous.  Table of

contents. IEEE Computer Graphics and Applications, 37(3):2–3, May/ June 2017. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2017/03/mcg2017030002.
pdf.

Anonymous:2017:TCd

[Ano17-35]  Anonymous.  Table of

contents. IEEE Computer Graphics and Applications, 37(4):2–3, July/ August 2017. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2017/04/mcg2017040002.
pdf.

Anonymous:2017:TCe

[Ano17-36]  Anonymous.  Table of

contents. IEEE Computer Graphics and Applications, 37(5):1–2, September/October 2017. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/

                                        cg/2017/05/mcg2017050001.
                                        pdf.

Anonymous:2017:TCf

[Ano17-37]  Anonymous.  Table of

contents. IEEE Computer Graphics and Applications, 37(6):1–2, November/ December 2017. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        cg/2017/06/mcg2017060001.
                                        pdf.

Anonymous:2018:ITB

[Ano18a] Anonymous. IEEE Transactions on Big Data. IEEE Computer Graphics and Applications, 38(6):134, ???? 2018. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2018:ITS

[Ano18b] Anonymous. IEEE Transactions on Sustainable Computing. IEEE Computer Graphics and Applications, 38(6):135, ???? 2018. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2018:CPI

[Ano18c] Anonymous. Call for papers: Immersive analytics. IEEE Computer Graphics and Applications, 38(2):1, March/April 2018. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2018/02/mcg2018020001.
pdf.

Anonymous:2018:CBA

[Ano18d] Anonymous. Charles Babbage Award ad. IEEE Computer Graphics and Applications, 38(3):c2, May/ June 2018. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2018/03/mcg20180300c2.
pdf.

Anonymous:2018:C

[Ano18e]  Anonymous.  COMP-

SAC 2018. IEEE Computer Graphics and Applications, 38(1):c3, January/ February 2018. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2018/01/mcg20180100c3.
pdf.

Anonymous:2018:CIA

[Ano18f] Anonymous. Connect on interface ad. IEEE Computer Graphics and Applications, 38(4):c4, July/ August 2018. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/

                                        cg/2018/04/mcg20180400c4.
                                        pdf.

Anonymous:2018:CI

[Ano18g] Anonymous. CS info. IEEE Computer Graphics and Applications, 38(1):128, January/February 2018. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL https:/

                                        /www.computer.org/csdl/
                                        mags/cg/2018/01/mcg2018010128.
                                        pdf.

Anonymous:2018:FCa

[Ano18h] Anonymous. Front cover. IEEE Computer Graphics and Applications, 38(1): c1, January/February 2018. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL

                                        https://www.computer.org/
                                        csdl/mags/cg/2018/01/mcg20180100c1.
                                        pdf.

Anonymous:2018:FCb

[Ano18i] Anonymous. Front cover. IEEE Computer Graphics and Applications, 38(2):c1, March/April 2018. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        cg/2018/02/mcg20180200c1.
                                        pdf.

Anonymous:2018:FCc

[Ano18j] Anonymous. Front cover. IEEE Computer Graphics and Applications, 38(3):c1, May/June 2018. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2018/03/mcg20180300c1.
pdf.

Anonymous:2018:FCd

[Ano18k] Anonymous. Front cover.

IEEE Computer Graphics and Applications, 38(4):c1, July/August 2018. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2018/04/mcg20180400c1.
pdf.

Anonymous:2018:FCe

[Ano18l] Anonymous. Front cover.

IEEE Computer Graphics and Applications, 38(5):c1, September/October 2018. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL

https://www.computer.org/
csdl/mags/cg/2018/05/mcg20180500c1.
pdf.

Anonymous:2018:FCf

[Ano18m] Anonymous. Front cover.

IEEE Computer Graphics and Applications, 38(6):C1, ???? 2018. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2018:HDM

[Ano18n]  Anonymous.  Harlan D.

Mills Award ad. IEEE Computer Graphics and Applications, 38(4):1, July/ August 2018. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        cg/2018/04/mcg2018040001.
                                        pdf.

Anonymous:2018:HWA

[Ano18o] Anonymous. Help wanted ads. IEEE Computer Graphics and Applications, 38 (5):c2, September/October 2018. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL

                                        https://www.computer.org/
                                        csdl/mags/cg/2018/05/mcg20180500c2.
                                        pdf.

Anonymous:2018:HA

[Ano18p] Anonymous. HPC award. IEEE Computer Graphics and Applications, 38(2):c2, March/April 2018. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        cg/2018/02/mcg20180200c2.
                                        pdf.

Anonymous:2018:HAA

[Ano18q] Anonymous. HPC award ad. IEEE Computer Graphics and Applications, 38(3): 1, May/June 2018. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2018/03/mcg2018030001.
pdf.

Anonymous:2018:ICSg

[Ano18r] Anonymous. IEEE Computer Society. IEEE Computer Graphics and Applications, 38(6):136, ???? 2018. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2018:ICSa

[Ano18s] Anonymous. IEEE Computer Society: Be at the center of it all. IEEE Computer Graphics and Applications, 38(1):c2, January/ February 2018. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2018/01/mcg20180100c2.
pdf.

Anonymous:2018:ICSc

[Ano18t] Anonymous. IEEE Computer Society info. IEEE Computer Graphics and Applications, 38(3):118, May/ June 2018. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2018/03/mcg2018030118.
pdf.

Anonymous:2018:ICSb

[Ano18u] Anonymous. IEEE Computer Society information. IEEE Computer Graphics and Applications, 38(2):132, March/April 2018. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        cg/2018/02/mcg2018020132.
                                        pdf.

Anonymous:2018:ICSe

[Ano18v] Anonymous. IEEE Computer Society information. IEEE Computer Graphics and Applications, 38(4):17, July/August 2018. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        cg/2018/04/mcg2018040017.
                                        pdf.

Anonymous:2018:ICSf

[Ano18w] Anonymous. IEEE Computer Society information. IEEE Computer Graphics and Applications, 38(5):c3, September/October 2018. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL

                                        https://www.computer.org/
                                        csdl/mags/cg/2018/05/mcg20180500c3.
                                        pdf.

Anonymous:2018:ICSd

[Ano18x] Anonymous. IEEE Computer Society social networking ad. IEEE Computer Graphics and Applications, 38(4):9, July/ August 2018. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2018/04/mcg2018040009.
pdf.

Anonymous:2018:ILC

[Ano18y] Anonymous. IEEE letters of the Computer Society. IEEE Computer Graphics and Applications, 38(6):C2, ???? 2018. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2018:IA

[Ano18z] Anonymous. Interface ad.

IEEE Computer Graphics and Applications, 38(2):c4, March/April 2018. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2018/02/mcg20180200c4.
pdf.

Anonymous:2018:JBA

[Ano18-27] Anonymous. Jobs board ad.

IEEE Computer Graphics and Applications, 38(3):149, May/June 2018. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/

                                        cg/2018/03/mcg2018030149.
                                        pdf.

Anonymous:2018:LCS

[Ano18-28] Anonymous. Letters of the Computer Society ad. IEEE Computer Graphics and Applications, 38(4):c2, July/ August 2018. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        cg/2018/04/mcg20180400c2.
                                        pdf.

Anonymous:2018:LBT

[Ano18-29] Anonymous. Looking for the BEST tech job for you? IEEE Computer Graphics and Applications, 38(6):1, ???? 2018. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2018:Ma

[Ano18-30]  Anonymous.  Masthead.

IEEE Computer Graphics and Applications, 38(1): 1, January/February 2018. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL

                                        https://www.computer.org/
                                        csdl/mags/cg/2018/01/mcg2018010001.
                                        pdf.

Anonymous:2018:Mb

[Ano18-31]  Anonymous.  Masthead.

IEEE Computer Graphics and Applications, 38(2):7, March/April 2018. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2018/02/mcg2018020007.
pdf.

Anonymous:2018:Mc

[Ano18-32]  Anonymous.  Masthead.

IEEE Computer Graphics and Applications, 38(4):5, July/August 2018. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2018/04/mcg2018040005.
pdf.

Anonymous:2018:Md

[Ano18-33]  Anonymous.  Masthead.

IEEE Computer Graphics and Applications, 38(5):1, September/October 2018. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL

https://www.computer.org/
csdl/mags/cg/2018/05/mcg2018050001.
pdf.

Anonymous:2018:M

[Ano18-34]  Anonymous.  Masthead.

IEEE Computer Graphics and Applications, 38(6):4, ???? 2018. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2018:MSC

[Ano18-35] Anonymous. Membership skill choice ad. IEEE Computer Graphics and Applications, 38(2):102, March/ April 2018. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2018/02/mcg2018020102.
pdf.

Anonymous:2018:MSA

[Ano18-36] Anonymous. Membership SkillChoice ad. IEEE Computer Graphics and Applications, 38(4):c3, July/ August 2018. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2018/04/mcg20180400c3.
pdf.

Anonymous:2018:MSGa

[Ano18-37] Anonymous. Membership student gift ad. IEEE Computer Graphics and Applications, 38(2):c3, March/ April 2018. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2018/02/mcg20180200c3.
pdf.

Anonymous:2018:MSGb

[Ano18-38] Anonymous. Membership student gift ad. IEEE Computer Graphics and Applications, 38(4):37, July/ August 2018. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2018/04/mcg2018040037.
pdf.

Anonymous:2018:MML

[Ano18-39] Anonymous. myCS: More to love. IEEE Computer Graphics and Applications, 38(1):c4, January/ February 2018. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2018/01/mcg20180100c4.
pdf.

Anonymous:2018:OMU

[Ano18-40] Anonymous. One membership. Unlimited knowledge. IEEE Computer Graphics and Applications, 38(6):C3, ???? 2018. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2018:SNA

[Ano18-41] Anonymous. Social networking ad. IEEE Computer Graphics and Applications, 38(5):c4, September/October 2018. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2018/05/mcg20180500c4.
pdf.

Anonymous:2018:SC

[Ano18-42]  Anonymous.  Stay con-

nected. IEEE Computer Graphics and Applications, 38(6):C4, ???? 2018. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2018:SGA

[Ano18-43]  Anonymous.  Student

gift ad. IEEE Computer Graphics and Applications, 38(3):c4, May/ June 2018. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        cg/2018/03/mcg20180300c4.
                                        pdf.

Anonymous:2018:TCa

[Ano18-44]  Anonymous.  Table of

contents. IEEE Computer Graphics and Applications, 38(1):2–3, January/ February 2018. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        cg/2018/01/mcg2018010002.
                                        pdf.

Anonymous:2018:TCb

[Ano18-45]  Anonymous.  Table of

contents. IEEE Computer Graphics and Applications, 38(2):2–3, March/ April 2018. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2018/02/mcg2018020002.
pdf.

Anonymous:2018:TCc

[Ano18-46]  Anonymous.  Table of

contents. IEEE Computer Graphics and Applications, 38(3):2–3, May/ June 2018. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2018/03/mcg2018030002.
pdf.

Anonymous:2018:TCd

[Ano18-47]  Anonymous.  Table of

contents. IEEE Computer Graphics and Applications, 38(4):2–3, July/ August 2018. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2018/04/mcg2018040002.
pdf.

Anonymous:2018:TCe

[Ano18-48]  Anonymous.  Table of

contents. IEEE Computer Graphics and Applications, 38(5):2–3, September/October 2018. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/

                                        cg/2018/05/mcg2018050002.
                                        pdf.

Anonymous:2018:TCf

[Ano18-49] Anonymous. Table of contents. IEEE Computer Graphics and Applications, 38(6):2–3, ???? 2018. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2018:TCL

[Ano18-50] Anonymous. Take the CS library wherever you go ad. IEEE Computer Graphics and Applications, 38(2):121, March/April 2018. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        cg/2018/02/mcg2018020121.
                                        pdf.

Anonymous:2019:AIC

[Ano19a] Anonymous. ACM/IEEE CS Eckert–Mauchly Award. IEEE Computer Graphics and Applications, 39(1): C4, January/February 2019. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2019:AI

[Ano19b] Anonymous. Apple Inc. IEEE Computer Graphics and Applications, 39(6):45, November/December 2019. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2019:BRRa

[Ano19c] Anonymous. B. Ramakrishna Rau Award. IEEE Computer Graphics and Applications, 39(1):110, January/February 2019. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2019:BRRb

[Ano19d] Anonymous. B. Ramakrishna Rau Award. IEEE Computer Graphics and Applications, 39(2):C4, March/ April 2019. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2019:BD

[Ano19e] Anonymous. Big data. IEEE Computer Graphics and Applications, 39(2):24, March/ April 2019. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2019:CE

[Ano19f]  Anonymous.  Comput-

ing Edge. IEEE Computer Graphics and Applications, 39(5):C4, September/October 2019. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2019:ICA

[Ano19g] Anonymous. IEEE Computer Architecture Letters.

IEEE Computer Graphics and Applications, 39(4):96, July/August 2019. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2019:IIC

[Ano19h] Anonymous. IEEE Internet Computing. IEEE Computer Graphics and Applications, 39(4):C3, July/ August 2019. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2019:ISPa

[Ano19i] Anonymous. IEEE Security & Privacy. IEEE Computer Graphics and Applications, 39(2):C2, March/ April 2019. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2019:ISPb

[Ano19j] Anonymous. IEEE Security & Privacy. IEEE Computer Graphics and Applications, 39(3):C2, May/June 2019. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2019:ITSa

[Ano19k] Anonymous. IEEE Transactions on Sustainable Computing. IEEE Computer Graphics and Applications, 39(2):1, March/April 2019. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2019:ITSb

[Ano19l] Anonymous. IEEE Transactions on Sustainable Computing. IEEE Computer Graphics and Applications, 39(6):29, November/ December 2019. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2019:CMA

[Ano19m] Anonymous. Call for 2019 major awards nominations. IEEE Computer Graphics and Applications, 39(5):C3, September/October 2019. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2019:CAa

[Ano19n]  Anonymous.  Call for

articles. IEEE Computer Graphics and Applications, 39(6):85, November/ December 2019. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2019:CPI

[Ano19o] Anonymous. Call for papers: IEEE Transactions on Computers. IEEE Computer Graphics and Applications, 39(6):C2, November/ December 2019. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2019:Cb

[Ano19p] Anonymous. COMPSAC 2019. IEEE Computer Graphics and Applications, 39(3):C3, May/June 2019. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2019:Cc

[Ano19q] Anonymous. ComputingEdge. IEEE Computer Graphics and Applications, 39(4):1, July/August 2019. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2019:Cd

[Ano19r] Anonymous. ComputingEdge. IEEE Computer Graphics and Applications, 39(6):C3, November/December 2019. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2019:CIa

[Ano19s] Anonymous. Connect on Interface. IEEE Computer Graphics and Applications, 39(1):C2, January/February 2019. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2019:CIb

[Ano19t] Anonymous. Connect on Interface. IEEE Computer Graphics and Applications, 39(2):C3, March/ April 2019. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2019:Ca

[Ano19u]  Anonymous.  Contents.

IEEE Computer Graphics and Applications, 39(1):2– 3, January/February 2019. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2019:Ea

[Ano19v]  Anonymous.  Erratum.

IEEE Computer Graphics and Applications, 39 (1):106, January/February 2019. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL

https://ieeexplore.ieee.
org/document/8667503/.

Anonymous:2019:Eb

[Ano19w]  Anonymous.  Erratum.

IEEE Computer Graphics and Applications, 39(6):120– 122, November/December 2019. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2019:FCa

[Ano19x] Anonymous. Front cover.

IEEE Computer Graphics and Applications, 39(1): C1, January/February 2019. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2019:FCb

[Ano19y] Anonymous. Front cover. IEEE Computer Graphics and Applications, 39(2):C1, March/April 2019. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2019:FCc

[Ano19z] Anonymous. Front cover. IEEE Computer Graphics and Applications, 39(3):C1, May/June 2019. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2019:FCd

[Ano19-27] Anonymous. Front cover. IEEE Computer Graphics and Applications, 39(4):C1, July/August 2019. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2019:FCe

[Ano19-28] Anonymous. Front cover. IEEE Computer Graphics and Applications, 39(5):C1, September/October 2019. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2019:FCf

[Ano19-29] Anonymous. Front cover. IEEE Computer Graphics and Applications, 39(6):C1, November/December 2019. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2019:ICSa

[Ano19-30] Anonymous. IEEE Computer Society. IEEE Computer Graphics and Applications, 39(1):112, January/ February 2019. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2019:ICSb

[Ano19-31] Anonymous. IEEE Computer Society. IEEE Computer Graphics and Applications, 39(2):112, March/ April 2019. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2019:ICSc

[Ano19-32] Anonymous. IEEE Computer Society. IEEE Computer Graphics and Applications, 39(3):96, May/ June 2019. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2019:ICSg

[Ano19-33] Anonymous. IEEE Computer Society. IEEE Computer Graphics and Applications, 39(6):123, November/ December 2019. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2019:ICSf

[Ano19-34] Anonymous. IEEE Computer Society has you covered! IEEE Computer Graphics and Applications, 39(6):1, November/ December 2019. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2019:ICSd

[Ano19-35] Anonymous. IEEE Computer Society information. IEEE Computer Graphics and Applications, 39(4):95, July/August 2019. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2019:ICSe

[Ano19-36] Anonymous. IEEE Computer Society information. IEEE Computer Graphics and Applications, 39(5):104, September/October 2019. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2019:ID

[Ano19-37] Anonymous. IEEE DataPort. IEEE Computer Graphics and Applications, 39(5):1, September/October 2019. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2019:ILC

[Ano19-38] Anonymous. IEEE letters of the Computer Society. IEEE Computer Graphics and Applications, 39(1):111, January/February 2019. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2019:IWCa

[Ano19-39] Anonymous. IEEE World Congress on Services 2019. IEEE Computer Graphics and Applications, 39(3):C4, May/June 2019. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2019:IWCb

[Ano19-40] Anonymous. IEEE World Congress on Services 2019. IEEE Computer Graphics and Applications, 39(4):C4, July/August 2019. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2019:KYCa

[Ano19-41] Anonymous. Keep your career options open. IEEE Computer Graphics and Applications, 39(4):C2, July/ August 2019. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2019:KYCb

[Ano19-42] Anonymous. Keep your career options open. IEEE Computer Graphics and Applications, 39(5):C2, September/October 2019. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2019:KYC

[Ano19-43] Anonymous. Keep your career options open. IEEE Computer Graphics and Applications, 39(6):C4, November/December 2019. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2019:LCS

[Ano19-44] Anonymous. Letters of The Computer Society. IEEE Computer Graphics and Applications, 39(3):1, May/ June 2019. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2019:LBT

[Ano19-45] Anonymous. Looking for the best tech job for YOU. IEEE Computer Graphics and Applications, 39(1): C3, January/February 2019. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2019:Ma

[Ano19-46]  Anonymous.  Masthead.

IEEE Computer Graphics and Applications, 39(1): 4, January/February 2019. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL

https://ieeexplore.ieee.
org/document/8663514/.

Anonymous:2019:Mb

[Ano19-47]  Anonymous.  Masthead.

IEEE Computer Graphics and Applications, 39(2):4, March/April 2019. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL https:

//ieeexplore.ieee.org/
document/8672996/.

Anonymous:2019:Mc

[Ano19-48]  Anonymous.  Masthead.

IEEE Computer Graphics and Applications, 39(3):4, May/June 2019. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2019:Md

[Ano19-49]  Anonymous.  Masthead.

IEEE Computer Graphics and Applications, 39(4):4, July/August 2019. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2019:Me

[Ano19-50]  Anonymous.  Masthead.

IEEE Computer Graphics and Applications, 39(5):4, September/October 2019. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2019:Mf

[Ano19-51]  Anonymous.  Masthead.

IEEE Computer Graphics and Applications, 39(6):4, November/December 2019. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2019:SP

[Ano19-52] Anonymous. Security & privacy. IEEE Computer Graphics and Applications, 39(6):16, November/ December 2019. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2019:SMY

[Ano19-53] Anonymous. Share and manage your research data. IEEE Computer Graphics and Applications, 39(6):124, November/December 2019. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2019:SID

[Ano19-54] Anonymous. Special issue: Deep learning in computer graphics. IEEE Computer Graphics and Applications, 39(2):25, March/ April 2019. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL https:

//ieeexplore.ieee.org/
document/8673960/.

Anonymous:2019:SC

[Ano19-55]  Anonymous.  Stay con-

nected. IEEE Computer Graphics and Applications, 39(1):1, January/February 2019. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2019:TCa

[Ano19-56]  Anonymous.  Table of

contents. IEEE Computer Graphics and Applications, 39(2):2–3, March/ April 2019. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2019:TCb

[Ano19-57] Anonymous. Table of contents. IEEE Computer Graphics and Applications, 39(3):2–3, May/June 2019. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2019:TCc

[Ano19-58]  Anonymous.  Table of

contents. IEEE Computer Graphics and Applications, 39(4):2–3, July/ August 2019. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2019:TCd

[Ano19-59]  Anonymous.  Table of

contents. IEEE Computer Graphics and Applications, 39(5):2–3, September/October 2019. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2019:TCe

[Ano19-60]  Anonymous.  Table of

contents. IEEE Computer Graphics and Applications, 39(6):2–3, November/ December 2019. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2019:V

[Ano19-61] Anonymous. Visualization 4.0. IEEE Computer Graphics and Applications, 39(6): 8–16, November/December 2019. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2020:AS

[Ano20a] Anonymous. Apple-Sixth. IEEE Computer Graphics and Applications, 40(1): 55, January/February 2020. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2020:BRR

[Ano20b] Anonymous. B. Ramakrishna Rau Award. IEEE Computer Graphics and Applications, 40(2):C3, March/ April 2020. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2020:CE

[Ano20c] Anonymous. Computing Edge. IEEE Computer Graphics and Applications, 40(1):C4, January/February 2020. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2020:Cb

[Ano20d] Anonymous. ComputingEdge.

IEEE Computer Graphics and Applications, 40(5):C4, September/October 2020. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2020:Cc

[Ano20e] Anonymous. ComputingEdge.

IEEE Computer Graphics and Applications, 40(6):C4, November/December 2020. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2020:IOJ

[Ano20f] Anonymous. IEEE Open Journal of the Computer Society. IEEE Computer Graphics and Applications, 40(1):1, January/February 2020. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2020:ITB

[Ano20g] Anonymous. IEEE Transactions on Big Data. IEEE Computer Graphics and Applications, 40(2):120, March/ April 2020. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2020:ITS

[Ano20h] Anonymous. IEEE Transactions on Sustainable Computing. IEEE Computer Graphics and Applications, 40(6):75, November/ December 2020. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2020:CAa

[Ano20i] Anonymous. Call for articles. IEEE Computer Graphics and Applications, 40(4):95, July/August 2020. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2020:CAIa

[Ano20j] Anonymous. Call for articles: IEEE Pervasive Computing. IEEE Computer Graphics and Applications, 40(2):102, March/ April 2020. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2020:CAIb

[Ano20k] Anonymous. Call for articles: IEEE Pervasive Computing. IEEE Computer Graphics and Applications, 40(3):104, May/June 2020. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2020:CAIc

[Ano20l] Anonymous. Call for articles: IEEE Pervasive Computing. IEEE Computer Graphics and Applications, 40(4):70, July/August 2020. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2020:CPIa

[Ano20m] Anonymous. Call for papers: IEEE Transactions on Computers. IEEE Computer Graphics and Applications, 40(1):111, January/ February 2020. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2020:CPIc

[Ano20n] Anonymous. Call for papers:

IEEE Transactions on Computers. IEEE Computer Graphics and Applications, 40(2):90, March/April 2020. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2020:CPId

[Ano20o] Anonymous. Call for papers:

IEEE Transactions on Computers. IEEE Computer Graphics and Applications, 40(3):16, May/June 2020. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2020:CPIe

[Ano20p] Anonymous. Call for papers: IEEE Transactions on Computers. IEEE Computer Graphics and Applications, 40(4):50, July/August 2020. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2020:CPI

[Ano20q] Anonymous. Call for papers: IEEE Transactions on Computers. IEEE Computer Graphics and Applications, 40(6):87, November/ December 2020. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2020:CPIb

[Ano20r] Anonymous. Call for papers: IEEE 21st International Conference on Information Reuse and Integration for Data Science (IEEE IRI 2020). IEEE Computer Graphics and Applications, 40(2):C2, March/ April 2020. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2020:CSI

[Ano20s] Anonymous. Call for special issue proposals. IEEE Computer Graphics and Applications, 40(5):128, September/October 2020. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2020:Ca

[Ano20t] Anonymous. COMPSAC 2020. IEEE Computer Graphics and Applications, 40(1):C2, January/February 2020. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2020:CVJ

[Ano20u] Anonymous. CVPR Virtual June 14–19. IEEE Computer Graphics and Applications, 40(6):C2, November/ December 2020. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2020:EAN

[Ano20v]  Anonymous.  Education

awards nominations. IEEE Computer Graphics and Applications, 40(2):1, March/ April 2020. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2020:ECO

[Ano20w] Anonymous. Evolving career opportunities need your skills. IEEE Computer Graphics and Applications, 40(5):1, September/October 2020. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2020:FCa

[Ano20x] Anonymous. Front cover. IEEE Computer Graphics and Applications, 40(1): C1, January/February 2020. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2020:FCb

[Ano20y] Anonymous. Front cover. IEEE Computer Graphics and Applications, 40(2):C1, March/April 2020. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2020:FCc

[Ano20z] Anonymous. Front cover. IEEE Computer Graphics and Applications, 40(3):C1, May/June 2020. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2020:FCd

[Ano20-27] Anonymous. Front cover. IEEE Computer Graphics and Applications, 40(4):C1, July/August 2020. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2020:FCe

[Ano20-28] Anonymous. Front cover. IEEE Computer Graphics and Applications, 40(5):C1, September/October 2020. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2020:FCf

[Ano20-29] Anonymous. Front cover.

IEEE Computer Graphics and Applications, 40(6):C1, November/December 2020. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2020:GPNa

[Ano20-30] Anonymous. Get published in the new IEEE Open Journal of the Computer Society. IEEE Computer Graphics and Applications, 40(3):C4, May/June 2020. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2020:GPNb

[Ano20-31] Anonymous. Get published in the new IEEE Open Journal of the Computer Society. IEEE Computer Graphics and Applications, 40(5):C3, September/October 2020. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2020:GPN

[Ano20-32] Anonymous. Get published in the new IEEE Open Journal of the Computer Society. IEEE Computer Graphics and Applications, 40(6):C3, November/December 2020. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2020:HWCa

[Ano20-33]  Anonymous.  Honoring

women in computing. IEEE Computer Graphics and Applications, 40(3):72, May/ June 2020. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2020:HWCb

[Ano20-34]  Anonymous.  Honoring

women in computing: Call for nominations. IEEE Computer Graphics and Applications, 40(4):1, July/ August 2020. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2020:HWC

[Ano20-35]  Anonymous.  Honoring

women in computing: Call for nominations. IEEE Computer Graphics and Applications, 40(6):98, November/December 2020. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2020:Ha

[Ano20-36] Anonymous. HOST 2020. IEEE Computer Graphics and Applications, 40(3):C3, May/June 2020. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2020:Hb

[Ano20-37] Anonymous. Host 2020.

IEEE Computer Graphics and Applications, 40(4):C3, July/August 2020. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2020:ICSg

[Ano20-38] Anonymous. IEEE Computer Society. IEEE Computer Graphics and Applications, 40(4):104, July/ August 2020. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2020:ICSj

[Ano20-39] Anonymous. IEEE Computer Society. IEEE Computer Graphics and Applications, 40(5):127, September/October 2020. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2020:ICSn

[Ano20-40] Anonymous. IEEE Computer Society. IEEE Computer Graphics and Applications, 40(6):100, November/ December 2020. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2020:ICSi

[Ano20-41] Anonymous. IEEE Computer Society: Call for papers. IEEE Computer Graphics and Applications, 40(5):25, September/October 2020. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2020:ICSl

[Ano20-42] Anonymous. IEEE Computer Society: Call for papers. IEEE Computer Graphics and Applications, 40(6):37, November/ December 2020. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2020:ICSc

[Ano20-43] Anonymous. IEEE Computer Society has you covered! IEEE Computer Graphics and Applications, 40(3):C2, May/June 2020. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2020:ICSf

[Ano20-44] Anonymous. IEEE Computer Society has you covered! IEEE Computer Graphics and Applications, 40(4):C2, July/ August 2020. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2020:ICSh

[Ano20-45] Anonymous. IEEE Computer Society has you covered! IEEE Computer Graphics and Applications, 40(5):C2, September/October 2020. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2020:ICSm

[Ano20-46] Anonymous. IEEE Computer Society has you covered! IEEE Computer Graphics and Applications, 40(6):97, November/ December 2020. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2020:ICSa

[Ano20-47] Anonymous. IEEE Computer Society information. IEEE Computer Graphics and Applications, 40 (1):148, January/February 2020. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2020:ICSb

[Ano20-48] Anonymous. IEEE Computer Society information. IEEE Computer Graphics and Applications, 40(2):119, March/April 2020. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2020:ICSd

[Ano20-49] Anonymous. IEEE Computer Society information. IEEE Computer Graphics and Applications, 40(3):123, May/June 2020. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2020:ICSk

[Ano20-50] Anonymous. IEEE Computer Society Jobs Board. IEEE Computer Graphics and Applications, 40(6):1, November/December 2020. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2020:ICSe

[Ano20-51] Anonymous. IEEE Computer Society Volunteer Service Awards. IEEE Computer Graphics and Applications, 40(3):124, May/ June 2020. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2020:IQWa

[Ano20-52] Anonymous. IEEE Quantum Week. IEEE Computer Graphics and Applications, 40(3):1, May/June 2020. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2020:IQWb

[Ano20-53] Anonymous. IEEE Quantum Week. IEEE Computer Graphics and Applications, 40(4):C4, July/ August 2020. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2020:KYC

[Ano20-54] Anonymous. Keep your career options open. IEEE Computer Graphics and Applications, 40(1):C3, January/February 2020. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2020:Ma

[Ano20-55]  Anonymous.  Masthead.

IEEE Computer Graphics and Applications, 40(1): 4, January/February 2020. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2020:Mb

[Ano20-56]  Anonymous.  Masthead.

IEEE Computer Graphics and Applications, 40(2):4, March/April 2020. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2020:Mc

[Ano20-57]  Anonymous.  Masthead.

IEEE Computer Graphics and Applications, 40(3):4, May/June 2020. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2020:Md

[Ano20-58]  Anonymous.  Masthead.

IEEE Computer Graphics and Applications, 40(4):4, July/August 2020. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2020:Me

[Ano20-59]  Anonymous.  Masthead.

IEEE Computer Graphics and Applications, 40(5):4, September/October 2020. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2020:Mf

[Ano20-60]  Anonymous.  Masthead.

IEEE Computer Graphics and Applications, 40(6):4, November/December 2020. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2020:NWC

[Ano20-61]  Anonymous.  Nominate

a worthy colleague today! IEEE Computer Graphics and Applications, 40(2):C4, March/April 2020. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2020:RBL

[Ano20-62] Anonymous. Remembering Bill Lorensen: The man, the myth, and marching cubes. IEEE Computer Graphics and Applications, 40(2): 112–118, March/April 2020. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2020:SYP

[Ano20-63] Anonymous. Share your preprint research with the world! IEEE Computer Graphics and Applications, 40(6):99, November/ December 2020. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2020:TCa

[Ano20-64]  Anonymous.  Table of

contents. IEEE Computer Graphics and Applications, 40(1):2–3, January/ February 2020. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2020:TCb

[Ano20-65]  Anonymous.  Table of

contents. IEEE Computer Graphics and Applications, 40(2):2–3, March/ April 2020. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2020:TCc

[Ano20-66] Anonymous. Table of contents. IEEE Computer Graphics and Applications, 40(3):2–3, May/June 2020. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2020:TCd

[Ano20-67]  Anonymous.  Table of

contents. IEEE Computer Graphics and Applications, 40(4):2–3, July/ August 2020. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2020:TCe

[Ano20-68]  Anonymous.  Table of

contents. IEEE Computer Graphics and Applications, 40(5):2–3, September/October 2020. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2020:TCf

[Ano20-69]  Anonymous.  Table of

contents. IEEE Computer Graphics and Applications, 40(6):2–3, November/ December 2020. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2020:TTT

[Ano20-70] Anonymous. Top technology trends for 2020 featured in Computer. IEEE Computer Graphics and Applications, 40(1):39, January/ February 2020. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2021:CSEa

[Ano21a] Anonymous. Computing in Science & Engineering. IEEE Computer Graphics and Applications, 41(3):95, May/June 2021. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2021:CSEb

[Ano21b] Anonymous. Computing in Science & Engineering. IEEE Computer Graphics and Applications, 41(5):137, September/October 2021. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2021:CSEc

[Ano21c] Anonymous. Computing in Science & Engineering. IEEE Computer Graphics and Applications, 41(6):142, November/December 2021. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2021:Ca

[Ano21d] Anonymous. ComputingEdge.

IEEE Computer Graphics and Applications, 41(1): C2, January/February 2021. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2021:Cb

[Ano21e] Anonymous. ComputingEdge.

IEEE Computer Graphics and Applications, 41(3):C2, May/June 2021. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2021:Cc

[Ano21f] Anonymous. ComputingEdge. IEEE Computer Graphics and Applications, 41(4):C2, July/August 2021. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2021:Ce

[Ano21g] Anonymous. ComputingEdge. IEEE Computer Graphics and Applications, 41(5):C2, September/October 2021. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2021:Cg

[Ano21h] Anonymous. ComputingEdge. IEEE Computer Graphics and Applications, 41(6):C2, November/December 2021. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2021:IAHa

[Ano21i] Anonymous. IEEE Annals of the History of Computing. IEEE Computer Graphics and Applications, 41(5):123, September/October 2021. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2021:IAHb

[Ano21j] Anonymous. IEEE Annals of the History of Computing. IEEE Computer Graphics and Applications, 41(6):186, November/December 2021. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2021:ICG

[Ano21k] Anonymous. IEEE Computer Graphics and Applications. IEEE Computer Graphics and Applications, 41(2):4, March/April 2021. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2021:ISPa

[Ano21l] Anonymous. IEEE Security & Privacy. IEEE Computer Graphics and Applications, 41(1):62, January/ February 2021. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2021:ISPb

[Ano21m] Anonymous. IEEE Security & Privacy. IEEE Computer Graphics and Applications, 41(3):58, May/June 2021. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2021:ITBa

[Ano21n] Anonymous. IEEE Transactions on Big Data. IEEE Computer Graphics and Applications, 41(1):48, January/February 2021. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2021:ITBb

[Ano21o] Anonymous. IEEE Transactions on Big Data. IEEE Computer Graphics and Applications, 41(2):88, March/ April 2021. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2021:ITBc

[Ano21p] Anonymous. IEEE Transactions on Big Data. IEEE Computer Graphics and Applications, 41(3):19, May/ June 2021. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2021:ITBd

[Ano21q] Anonymous. IEEE Transactions on Big Data. IEEE Computer Graphics and Applications, 41(6):24, November/December 2021. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2021:ITSa

[Ano21r] Anonymous. IEEE Transactions on Sustainable Computing. IEEE Computer Graphics and Applications, 41(1):108, January/ February 2021. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2021:ITSb

[Ano21s] Anonymous. IEEE Transactions on Sustainable Computing. IEEE Computer Graphics and Applications, 41(2):24, March/April 2021. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2021:ITS

[Ano21t] Anonymous. IEEE Transactions on Sustainable Computing. IEEE Computer Graphics and Applications, 41(4):135, July/ August 2021. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2021:IPC

[Ano21u] Anonymous. IT Professional: Call for articles. IEEE Computer Graphics and Applications, 41(6):58, November/December 2021. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2021:CAa

[Ano21v] Anonymous. Call for articles. IEEE Computer Graphics and Applications, 41(1):98, January/February 2021. CODEN ICGADZ.

ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2021:CAb

[Ano21w] Anonymous. Call for articles. IEEE Computer Graphics and Applications, 41(3):70, May/June 2021. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2021:CAc

[Ano21x] Anonymous. Call for articles. IEEE Computer Graphics and Applications, 41(3):84, May/June 2021. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2021:CAI

[Ano21y] Anonymous. Call for articles: IEEE Pervasive Computing. IEEE Computer Graphics and Applications, 41(1):86, January/February 2021. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2021:CPIa

[Ano21z] Anonymous. Call for papers: IEEE Transactions on Computers. IEEE Computer Graphics and Applications, 41(1):34, January/ February 2021. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2021:CPIb

[Ano21-27] Anonymous. Call for papers:

IEEE Transactions on Computers. IEEE Computer Graphics and Applications, 41(2):62, March/April 2021. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2021:CPIc

[Ano21-28] Anonymous. Call for papers:

IEEE Transactions on Computers. IEEE Computer Graphics and Applications, 41(3):7, May/June 2021. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2021:CPId

[Ano21-29] Anonymous. Call for papers: IEEE Transactions on Computers. IEEE Computer Graphics and Applications, 41(4):136, July/ August 2021. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2021:Ch

[Ano21-30]  Anonymous.  COMP-

SAC 2022. IEEE Computer Graphics and Applications, 41(6):1, November/ December 2021. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2021:Cf

[Ano21-31] Anonymous. Cover. IEEE Computer Graphics and Applications, 41(5):143, September/October 2021. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2021:ECOa

[Ano21-32] Anonymous. Evolving career opportunities need your skills. IEEE Computer Graphics and Applications, 41(1):1, January/February 2021. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2021:ECOb

[Ano21-33] Anonymous. Evolving career opportunities need your skills. IEEE Computer Graphics and Applications, 41(2):1, March/April 2021. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2021:FCa

[Ano21-34] Anonymous. Front cover. IEEE Computer Graphics and Applications, 41(1): C1, January/February 2021. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2021:FCb

[Ano21-35] Anonymous. Front cover. IEEE Computer Graphics and Applications, 41(3):C1, May/June 2021. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2021:FCc

[Ano21-36] Anonymous. Front cover.

IEEE Computer Graphics and Applications, 41(4):C1, July/August 2021. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2021:FCd

[Ano21-37] Anonymous. Front cover.

IEEE Computer Graphics and Applications, 41(5):C1, September/October 2021. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2021:FCe

[Ano21-38] Anonymous. [Front cover].

IEEE Computer Graphics and Applications, 41(6):C1, November/December 2021. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2021:GPIa

[Ano21-39] Anonymous. Get published in the IEEE Open Journal of the Computer Society. IEEE Computer Graphics and Applications, 41(4):1, July/August 2021. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2021:GPIb

[Ano21-40] Anonymous. Get published in the IEEE Open Journal of the Computer Society. IEEE Computer Graphics and Applications, 41(5):C3, September/October 2021. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2021:GPIc

[Ano21-41] Anonymous. Get published in the IEEE Open Journal of the Computer Society. IEEE Computer Graphics and Applications, 41(6):C3, November/December 2021. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2021:GPN

[Ano21-42] Anonymous. Get published in the new IEEE Open Journal of the Computer Society. IEEE Computer Graphics and Applications, 41(1): C3, January/February 2021. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2021:ICSb

[Ano21-43] Anonymous. IEEE Computer Society. IEEE Computer Graphics and Applications, 41(1):106, January/ February 2021. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2021:ICSe

[Ano21-44] Anonymous. IEEE Computer Society. IEEE Computer Graphics and Applications, 41(2):C1, March/ April 2021. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2021:ICSf

[Ano21-45] Anonymous. IEEE Computer Society. IEEE Computer Graphics and Applications, 41(2):C2, March/ April 2021. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2021:ICSh

[Ano21-46] Anonymous. IEEE Computer Society. IEEE Computer Graphics and Applications, 41(2):48, March/ April 2021. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2021:ICSi

[Ano21-47] Anonymous. IEEE Computer Society. IEEE Computer Graphics and Applications, 41(2):119, March/ April 2021. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2021:ICSj

[Ano21-48] Anonymous. IEEE Computer Society. IEEE Computer Graphics and Applications, 41(2):120, March/ April 2021. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2021:ICSk

[Ano21-49] Anonymous. IEEE Computer Society. IEEE Computer Graphics and Applications, 41(2):C3, March/ April 2021. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2021:ICSl

[Ano21-50] Anonymous. IEEE Computer Society. IEEE Computer Graphics and Applications, 41(2):C4, March/ April 2021. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2021:ICSn

[Ano21-51] Anonymous. IEEE Computer Society. IEEE Computer Graphics and Applications, 41(3):139, May/ June 2021. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2021:ICSo

[Ano21-52] Anonymous. IEEE Computer Society. IEEE Computer Graphics and Applications, 41(4):133, July/ August 2021. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2021:ICSu

[Ano21-53] Anonymous. IEEE Computer Society. IEEE Computer Graphics and Applications, 41(5):144, September/October 2021. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2021:ICSx

[Ano21-54] Anonymous. IEEE Computer Society. IEEE Computer Graphics and Applications, 41(6):199, November/ December 2021. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2021:ICSa

[Ano21-55] Anonymous. IEEE Computer Society: Call for papers. IEEE Computer Graphics and Applications, 41(1):7, January/February 2021. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2021:ICSg

[Ano21-56] Anonymous. IEEE Computer Society call for papers. IEEE Computer Graphics and Applications, 41(2):7, March/April 2021. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2021:ICSt

[Ano21-57] Anonymous. IEEE Computer Society: Call for papers. IEEE Computer Graphics and Applications, 41(5):78, September/October 2021. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2021:ICSw

[Ano21-58] Anonymous. IEEE Computer Society: Call for papers. IEEE Computer Graphics and Applications, 41(6):197, November/ December 2021. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2021:ICSd

[Ano21-59] Anonymous. IEEE Computer Society has you covered! IEEE Computer Graphics and Applications, 41(1):C4, January/February 2021. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2021:ICSr

[Ano21-60] Anonymous. IEEE Computer Society has you covered! IEEE Computer Graphics and Applications, 41(4):C4, July/ August 2021. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2021:ICSv

[Ano21-61] Anonymous. IEEE Computer Society has you covered! IEEE Computer Graphics and Applications, 41(5):C4, September/October 2021. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2021:ICSy

[Ano21-62] Anonymous. IEEE Computer Society has you covered! IEEE Computer Graphics and Applications, 41(6):200, November/ December 2021. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2021:ICSq

[Ano21-63] Anonymous. IEEE Computer Society Jobs Board. IEEE Computer Graphics and Applications, 41(4):C3, July/August 2021. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2021:ICSs

[Ano21-64] Anonymous. IEEE Computer Society Jobs Board. IEEE Computer Graphics and Applications, 41(5):1, September/October 2021. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2021:ICSz

[Ano21-65] Anonymous. IEEE Computer Society Jobs Board. IEEE Computer Graphics and Applications, 41(6):C4, November/December 2021. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2021:ICSc

[Ano21-66] Anonymous. IEEE Computer Society: Volunteer service awards. IEEE Computer Graphics and Applications, 41(1):107, January/ February 2021. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2021:ICSm

[Ano21-67] Anonymous. IEEE Computer Society Volunteer Service Awards. IEEE Computer Graphics and Applications, 41(3):1, May/June 2021. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2021:ICSp

[Ano21-68] Anonymous. IEEE Computer Society: Volunteer Service Awards. IEEE Computer Graphics and Applications, 41(4):134, July/ August 2021. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2021:Ma

[Ano21-69]  Anonymous.  Masthead.

IEEE Computer Graphics and Applications, 41(1): 4, January/February 2021. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2021:Mb

[Ano21-70]  Anonymous.  Masthead.

IEEE Computer Graphics and Applications, 41(3):4, May/June 2021. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2021:Mc

[Ano21-71]  Anonymous.  Masthead.

IEEE Computer Graphics and Applications, 41(4):4, July/August 2021. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2021:Md

[Ano21-72]  Anonymous.  Masthead.

IEEE Computer Graphics and Applications, 41(5):4, September/October 2021. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2021:Me

[Ano21-73] Anonymous. [Masthead].

IEEE Computer Graphics and Applications, 41(6):4, November/December 2021. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2021:TCa

[Ano21-74]  Anonymous.  Table of

contents. IEEE Computer Graphics and Applications, 41(1):2–3, January/ February 2021. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2021:TCb

[Ano21-75]  Anonymous.  Table of

contents. IEEE Computer Graphics and Applications, 41(2):2–3, March/ April 2021. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2021:TCc

[Ano21-76] Anonymous. Table of contents. IEEE Computer Graphics and Applications, 41(3):2–3, May/June 2021. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2021:TCd

[Ano21-77]  Anonymous.  Table of

contents. IEEE Computer Graphics and Applications, 41(4):2–3, July/ August 2021. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2021:TCe

[Ano21-78]  Anonymous.  Table of

contents. IEEE Computer Graphics and Applications, 41(5):2–3, September/October 2021. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2021:TCf

[Ano21-79]  Anonymous.  Table of

contents. IEEE Computer Graphics and Applications, 41(6):2–3, November/ December 2021. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2022:ICC

[Ano22a] Anonymous. 2023 IEEE CS:

Charles Babbage Award. IEEE Computer Graphics and Applications, 42(2):115, March/April 2022. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2022:AII

[Ano22b] Anonymous. Awards: IEEE Intelligent Systems magazine’s. IEEE Computer Graphics and Applications, 42(2):1, March/April 2022. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2022:CSEa

[Ano22c] Anonymous. Computing in Science & Engineering.

IEEE Computer Graphics and Applications, 42(1): 72, January/February 2022. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2022:CSEb

[Ano22d] Anonymous. Computing in Science & Engineering. IEEE Computer Graphics and Applications, 42(2):44, March/April 2022. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2022:CSEc

[Ano22e] Anonymous. Computing in Science & Engineering. IEEE Computer Graphics and Applications, 42(6):23, November/December 2022. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2022:Cb

[Ano22f] Anonymous. ComputingEdge. IEEE Computer Graphics and Applications, 42(1): C2, January/February 2022. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2022:Cc

[Ano22g] Anonymous. ComputingEdge. IEEE Computer Graphics and Applications, 42(2):C2, March/April 2022. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2022:Cd

[Ano22h] Anonymous. ComputingEdge.

IEEE Computer Graphics and Applications, 42(3):C2, May/June 2022. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2022:Ce

[Ano22i] Anonymous. ComputingEdge.

IEEE Computer Graphics and Applications, 42(4):C2, July/August 2022. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2022:Cf

[Ano22j] Anonymous. ComputingEdge.

IEEE Computer Graphics and Applications, 42(5):C2, September/October 2022. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2022:IAHa

[Ano22k] Anonymous. IEEE Annals of the History of Computing. IEEE Computer Graphics and Applications, 42(1): 55, January/February 2022. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2022:IAHb

[Ano22l] Anonymous. IEEE Annals of the History of Computing.

IEEE Computer Graphics and Applications, 42(2):32, March/April 2022. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2022:IAHc

[Ano22m] Anonymous. IEEE Annals of the History of Computing. IEEE Computer Graphics and Applications, 42(6):10, November/December 2022. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2022:ICA

[Ano22n] Anonymous. IEEE Computer Architecture Letters. IEEE Computer Graphics and Applications, 42(5):99, September/October 2022. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2022:ICSf

[Ano22o] Anonymous. IEEE Computer Society: Call for papers. IEEE Computer Graphics and Applications, 42(2):7, March/April 2022. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2022:IIS

[Ano22p] Anonymous. IEEE Intelligent Systems magazine. IEEE Computer Graphics and Applications, 42(3):115, May/June 2022. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2022:IPCb

[Ano22q] Anonymous. IEEE Pervasive Computing call for articles. IEEE Computer Graphics and Applications, 42(6):122, November/ December 2022. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2022:ISP

[Ano22r] Anonymous. IEEE Security & Privacy. IEEE Computer Graphics and Applications, 42(5):18, September/October 2022. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2022:ITBa

[Ano22s] Anonymous. IEEE Transactions on Big Data. IEEE Computer Graphics and Applications, 42(4):51, July/ August 2022. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2022:ITBb

[Ano22t] Anonymous. IEEE Transactions on Big Data. IEEE Computer Graphics and Applications, 42(6):125, November/December 2022. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2022:ITSa

[Ano22u] Anonymous. IEEE Transactions on Sustainable Computing. IEEE Computer Graphics and Applications, 42(4):113, July/ August 2022. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2022:ITSb

[Ano22v] Anonymous. IEEE Transactions on Sustainable Computing. IEEE Computer Graphics and Applications, 42(6):106, November/ December 2022. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2022:IPCa

[Ano22w] Anonymous. IT Professional call for articles. IEEE Computer Graphics and Applications, 42(6):36, November/ December 2022. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2022:CAa

[Ano22x]  Anonymous.  Call for

articles. IEEE Computer Graphics and Applications, 42(1):115, January/ February 2022. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2022:CAb

[Ano22y] Anonymous. Call for articles. IEEE Computer Graphics and Applications, 42(2):88, March/April 2022. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2022:CAIa

[Ano22z] Anonymous. Call for articles: IEEE Pervasive Computing. IEEE Computer Graphics and Applications, 42(1):133, January/ February 2022. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2022:CAIb

[Ano22-27] Anonymous. Call for articles: IEEE Pervasive Computing. IEEE Computer Graphics and Applications, 42(5):50, September/October 2022. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2022:CPI

[Ano22-28] Anonymous. Call for papers:

IEEE Transactions on Computers. IEEE Computer Graphics and Applications, 42(4):19, July/August 2022. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2022:CSI

[Ano22-29] Anonymous. Call for special issue proposals. IEEE Computer Graphics and Applications, 42(1):1, January/ February 2022. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2022:CEa

[Ano22-30]  Anonymous.  Comput-

ing edge. IEEE Computer Graphics and Applications, 42(6):C2, November/ December 2022. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2022:FCa

[Ano22-31] Anonymous. Front cover. IEEE Computer Graphics and Applications, 42(1): C1, January/February 2022. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2022:FCb

[Ano22-32] Anonymous. Front cover. IEEE Computer Graphics and Applications, 42(2):C1, March/April 2022. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2022:FCc

[Ano22-33] Anonymous. Front cover. IEEE Computer Graphics and Applications, 42(3):C1, May/June 2022. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2022:FCd

[Ano22-34] Anonymous. Front cover.

IEEE Computer Graphics and Applications, 42(4):C1, July/August 2022. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2022:FCe

[Ano22-35] Anonymous. Front cover.

IEEE Computer Graphics and Applications, 42(5):C1, September/October 2022. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2022:FCf

[Ano22-36] Anonymous. Front cover.

IEEE Computer Graphics and Applications, 42(6):C1, November/December 2022. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2022:GPNa

[Ano22-37] Anonymous. Get published in the new IEEE Open Journal of the Computer Society. IEEE Computer Graphics and Applications, 42(1): C3, January/February 2022. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2022:GPNb

[Ano22-38] Anonymous. Get published in the new IEEE Open Journal of the Computer Society. IEEE Computer Graphics and Applications, 42(2):C3, March/April 2022. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2022:GPNc

[Ano22-39] Anonymous. Get published in the new IEEE Open Journal of the Computer Society. IEEE Computer Graphics and Applications, 42(3):C3, May/June 2022. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2022:GPNd

[Ano22-40] Anonymous. Get published in the new IEEE Open Journal of the Computer Society. IEEE Computer Graphics and Applications, 42(4):C3, July/August 2022. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2022:GPNe

[Ano22-41] Anonymous. Get published in the new IEEE Open Journal of the Computer Society. IEEE Computer Graphics and Applications, 42(5):C3, September/October 2022. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2022:GPNf

[Ano22-42] Anonymous. Get published in the new IEEE Open Journal of the Computer Society. IEEE Computer Graphics and Applications, 42(6):C3, November/December 2022. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2022:ICSd

[Ano22-43] Anonymous. IEEE Computer Society. IEEE Computer Graphics and Applications, 42(1):136, January/ February 2022. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2022:ICSg

[Ano22-44] Anonymous. IEEE Computer Society. IEEE Computer Graphics and Applications, 42(2):116, March/ April 2022. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2022:ICSk

[Ano22-45] Anonymous. IEEE Computer Society. IEEE Computer Graphics and Applications, 42(3):116, May/ June 2022. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2022:ICSo

[Ano22-46] Anonymous. IEEE Computer Society. IEEE Computer Graphics and Applications, 42(4):120, July/ August 2022. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2022:ICSa

[Ano22-47] Anonymous. IEEE Computer Society: Call for papers. IEEE Computer Graphics and Applications, 42(1):16, January/February 2022. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2022:ICSj

[Ano22-48] Anonymous. IEEE Computer Society: Call for papers. IEEE Computer Graphics and Applications, 42(3):28, May/June 2022. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2022:ICSn

[Ano22-49] Anonymous. IEEE Computer Society: Call for papers. IEEE Computer Graphics and Applications, 42(4):39, July/August 2022. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2022:ICSv

[Ano22-50] Anonymous. IEEE Computer Society call for papers.

IEEE Computer Graphics and Applications, 42(6):8, November/December 2022. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2022:ICSc

[Ano22-51] Anonymous. IEEE Computer Society D&I Fund. IEEE Computer Graphics and Applications, 42 (1):135, January/February 2022. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2022:ICSi

[Ano22-52] Anonymous. IEEE Computer Society D&I Fund. IEEE Computer Graphics and Applications, 42(3):1, May/June 2022. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2022:ICSm

[Ano22-53] Anonymous. IEEE Computer Society D&I Fund. IEEE Computer Graphics and Applications, 42(4):1, July/August 2022. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2022:ICSr

[Ano22-54] Anonymous. IEEE Computer Society D&I Fund. IEEE Computer Graphics and Applications, 42(5):98, September/October 2022.

CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2022:ICSb

[Ano22-55] Anonymous. IEEE Computer Society has you covered! IEEE Computer Graphics and Applications, 42(1):134, January/ February 2022. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2022:ICSq

[Ano22-56] Anonymous. IEEE Computer Society has you covered! IEEE Computer Graphics and Applications, 42(5):1, September/October 2022. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2022:ICSu

[Ano22-57] Anonymous. IEEE Computer Society has you covered! IEEE Computer Graphics and Applications, 42(6):1, November/ December 2022. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2022:ICSs

[Ano22-58] Anonymous. IEEE Computer Society information. IEEE Computer Graphics and Applications, 42(5):100, September/October 2022. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2022:ICSw

[Ano22-59] Anonymous. IEEE Computer Society information. IEEE Computer Graphics and Applications, 42(6):126, November/December 2022. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2022:ICSe

[Ano22-60] Anonymous. IEEE Computer Society Jobs Board. IEEE Computer Graphics and Applications, 42(1): C4, January/February 2022. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2022:ICSh

[Ano22-61] Anonymous. IEEE Computer Society Jobs Board. IEEE Computer Graphics and Applications, 42(2):C4, March/April 2022. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2022:ICSl

[Ano22-62] Anonymous. IEEE Computer Society Jobs Board. IEEE Computer Graphics and Applications, 42(3):C4, May/June 2022. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2022:ICSp

[Ano22-63] Anonymous. IEEE Computer Society Jobs Board. IEEE Computer Graphics and Applications, 42(4):C4, July/August 2022. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2022:ICSt

[Ano22-64] Anonymous. IEEE Computer Society jobs board. IEEE Computer Graphics and Applications, 42(5):C4, September/October 2022. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2022:ICSx

[Ano22-65] Anonymous. IEEE Computer Society Jobs Board. IEEE Computer Graphics and Applications, 42(6):C4, November/December 2022. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2022:Ma

[Ano22-66]  Anonymous.  Masthead.

IEEE Computer Graphics and Applications, 42(1): 4, January/February 2022. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2022:Mb

[Ano22-67]  Anonymous.  Masthead.

IEEE Computer Graphics and Applications, 42(2):4, March/April 2022. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2022:Mc

[Ano22-68]  Anonymous.  Masthead.

IEEE Computer Graphics and Applications, 42(3):4, May/June 2022. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2022:Md

[Ano22-69]  Anonymous.  Masthead.

IEEE Computer Graphics and Applications, 42(4):4, July/August 2022. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2022:Me

[Ano22-70]  Anonymous.  Masthead.

IEEE Computer Graphics and Applications, 42(5):4, September/October 2022. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2022:Mf

[Ano22-71]  Anonymous.  Masthead.

IEEE Computer Graphics and Applications, 42(6):4, November/December 2022. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2022:RCSa

[Ano22-72]  Anonymous.  Over the

rainbow: 21st century security & privacy podcast. IEEE Computer Graphics and Applications, 42(2):100, March/April 2022. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2022:RCSb

[Ano22-73] Anonymous. Over the rainbow: 21st century security & privacy podcast. IEEE Computer Graphics and Applications, 42(5):75, September/October 2022. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2022:RCSc

[Ano22-74]  Anonymous.  Over the

rainbow: 21st century security & privacy podcast. IEEE Computer Graphics and Applications, 42(6):46, November/December 2022. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2022:TCa

[Ano22-75]  Anonymous.  Table of

contents. IEEE Computer Graphics and Applications, 42(1):2–3, January/ February 2022. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2022:TCb

[Ano22-76]  Anonymous.  Table of

contents. IEEE Computer Graphics and Applications, 42(2):2–3, March/ April 2022. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2022:TCc

[Ano22-77] Anonymous. Table of contents. IEEE Computer Graphics and Applications, 42(3):2–3, May/June 2022. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2022:TCd

[Ano22-78]  Anonymous.  Table of

contents. IEEE Computer Graphics and Applications, 42(4):2–3, July/ August 2022. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2022:TCe

[Ano22-79]  Anonymous.  Table of

contents. IEEE Computer Graphics and Applications, 42(5):2–3, September/October 2022. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2022:TCf

[Ano22-80]  Anonymous.  Table of

contents. IEEE Computer Graphics and Applications, 42(6):2–3, November/ December 2022. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2022:WHA

[Ano22-81]  Anonymous.  Watts S.

Humphrey Award. IEEE Computer Graphics and Applications, 42(4):79, July/ August 2022. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2023:ICA

[Ano23a] Anonymous. 2023 IEEE Conference on Article Intelligence. IEEE Computer Graphics and Applications, 43(3):C2, May/June 2023. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2023:CSEa

[Ano23b] Anonymous. Computing in Science & Engineering. IEEE Computer Graphics and Applications, 43(2):68, March/April 2023. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2023:CSEb

[Ano23c] Anonymous. Computing in Science & Engineering. IEEE Computer Graphics and Applications, 43(3):47, May/June 2023. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2023:CSEc

[Ano23d] Anonymous. Computing in Science & Engineering. IEEE Computer Graphics and Applications, 43(6):86, November/December 2023. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2023:Ca

[Ano23e] Anonymous. ComputingEdge.

IEEE Computer Graphics and Applications, 43(2):C4, March/April 2023. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2023:Cb

[Ano23f] Anonymous. ComputingEdge.

IEEE Computer Graphics and Applications, 43(3):C4, May/June 2023. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2023:Cd

[Ano23g] Anonymous. ComputingEdge.

IEEE Computer Graphics and Applications, 43(4):C4, July/August 2023. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2023:Ce

[Ano23h] Anonymous. ComputingEdge. IEEE Computer Graphics and Applications, 43(5):C4, September/October 2023. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2023:Cf

[Ano23i] Anonymous. ComputingEdge. IEEE Computer Graphics and Applications, 43(6):C4, November/December 2023. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2023:IAHa

[Ano23j] Anonymous. IEEE Annals of the History of Computing. IEEE Computer Graphics and Applications, 43(1): 38, January/February 2023. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2023:IAHb

[Ano23k] Anonymous. IEEE Annals of the History of Computing. IEEE Computer Graphics and Applications, 43(3):110, May/June 2023. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2023:IAHc

[Ano23l] Anonymous. IEEE Annals of the History of Computing. IEEE Computer Graphics and Applications, 43(5):121, September/October 2023. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2023:IAHd

[Ano23m] Anonymous. IEEE Annals of the History of Computing. IEEE Computer Graphics and Applications, 43(6):63, November/December 2023. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2023:ISPa

[Ano23n] Anonymous. IEEE Security & Privacy. IEEE Computer Graphics and Applications, 43(1):111, January/ February 2023. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2023:ISPb

[Ano23o] Anonymous. IEEE Security & Privacy. IEEE Computer Graphics and Applications, 43(3):87, May/June 2023. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2023:ITS

[Ano23p] Anonymous. IEEE Transactions on Sustainable Computing. IEEE Computer Graphics and Applications, 43(3):83, May/June 2023. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2023:CMAa

[Ano23q] Anonymous. Call for 2023 major awards nominations. IEEE Computer Graphics and Applications, 43(2):114, March/April 2023. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2023:CMAb

[Ano23r] Anonymous. Call for 2023 major awards nominations. IEEE Computer Graphics and Applications, 43(4):131, July/August 2023. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2023:CAb

[Ano23s] Anonymous. Call for articles. IEEE Computer Graphics and Applications, 43(3):60, May/June 2023. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2023:CAc

[Ano23t]  Anonymous.  Call for

articles. IEEE Computer Graphics and Applications, 43(4):110, July/ August 2023. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2023:CAIb

[Ano23u] Anonymous. Call for articles: IEEE Pervasive Computing. IEEE Computer Graphics and Applications, 43(3):11, May/June 2023. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2023:CAIa

[Ano23v] Anonymous. Call for articles: IT Professional. IEEE Computer Graphics and Applications, 43(2):100, March/April 2023. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2023:CAIc

[Ano23w] Anonymous. Call for articles: IT Professional. IEEE Computer Graphics and Applications, 43(6):140, November/December 2023. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2023:CPIf

[Ano23x] Anonymous. Call for papers:

IEEE Computer Graphics and Applications. IEEE Computer Graphics and Applications, 43(5):1, September/October 2023. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2023:CPIh

[Ano23y]  Anonymous.  Call for

papers: IEEE Computer Graphics and Applications. IEEE Computer Graphics and Applications, 43(5):131, September/October 2023. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2023:CPIi

[Ano23z] Anonymous. Call for papers: IEEE Computer Graphics and Applications. IEEE Computer Graphics and Applications, 43(6):1, November/December 2023. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2023:CPIk

[Ano23-27]  Anonymous.  Call for

papers: IEEE Computer Graphics and Applications. IEEE Computer Graphics and Applications, 43(6):144, November/December 2023. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2023:CPIa

[Ano23-28] Anonymous. Call for papers: IEEE Computer Society. IEEE Computer Graphics and Applications, 43(1): 9, January/February 2023. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2023:CPId

[Ano23-29] Anonymous. Call for papers:

IEEE Computer Society. IEEE Computer Graphics and Applications, 43(2):7, March/April 2023. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2023:CPIe

[Ano23-30] Anonymous. Call for papers: IEEE Computer Society. IEEE Computer Graphics and Applications, 43(3): 7, May/June 2023. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2023:CPIg

[Ano23-31] Anonymous. Call for papers:

IEEE Computer Society. IEEE Computer Graphics and Applications, 43(5):90, September/October 2023. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2023:CPIj

[Ano23-32] Anonymous. Call for papers:

IEEE Computer Society. IEEE Computer Graphics and Applications, 43(6):32, November/December 2023. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2023:CPIb

[Ano23-33] Anonymous. Call for papers:

IEEE conference on artificial intelligence. IEEE Computer Graphics and Applications, 43(1):C4, January/ February 2023. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2023:CPIc

[Ano23-34] Anonymous. Call for papers: IEEE quantum week. IEEE Computer Graphics and Applications, 43(2):1, March/April 2023. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2023:FCa

[Ano23-35] Anonymous. Front cover. IEEE Computer Graphics and Applications, 43(1): C1, January/February 2023. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2023:FCb

[Ano23-36] Anonymous. Front cover. IEEE Computer Graphics and Applications, 43(2):C1, March/April 2023. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2023:FCc

[Ano23-37] Anonymous. Front cover. IEEE Computer Graphics and Applications, 43(3):C1, May/June 2023. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2023:FCd

[Ano23-38] Anonymous. Front cover.

IEEE Computer Graphics and Applications, 43(4):C1, July/August 2023. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2023:FCe

[Ano23-39] Anonymous. Front cover.

IEEE Computer Graphics and Applications, 43(5):C1, September/October 2023. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2023:FCf

[Ano23-40] Anonymous. Front cover.

IEEE Computer Graphics and Applications, 43(6):C1, November/December 2023. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2023:ICSh

[Ano23-41] Anonymous. IEEE Computer Society career center. IEEE Computer Graphics and Applications, 43(2):C3, March/April 2023. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2023:ICSl

[Ano23-42] Anonymous. IEEE Computer Society career center.

IEEE Computer Graphics and Applications, 43(3):C3, May/June 2023. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2023:ICSq

[Ano23-43] Anonymous. IEEE Computer Society career center. IEEE Computer Graphics and Applications, 43(4):C3, July/August 2023. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2023:ICSu

[Ano23-44] Anonymous. IEEE Computer Society career center. IEEE Computer Graphics and Applications, 43(5):C3, September/October 2023. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2023:ICSy

[Ano23-45] Anonymous. IEEE Computer Society career center. IEEE Computer Graphics and Applications, 43(6):C3, November/December 2023. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2023:ICSm

[Ano23-46] Anonymous. IEEE Computer Society conferences on computer vision. IEEE Computer Graphics and Applications, 43(4):1, July/ August 2023. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2023:ICSo

[Ano23-47] Anonymous. IEEE Computer Society D&I fund. IEEE Computer Graphics and Applications, 43(4):130, July/August 2023. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2023:ICSr

[Ano23-48] Anonymous. IEEE Computer Society D&I fund. IEEE Computer Graphics and Applications, 43(5):C2, September/October 2023. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2023:ICSv

[Ano23-49] Anonymous. IEEE Computer Society D&I fund. IEEE Computer Graphics and Applications, 43(6):C2, November/December 2023. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2023:ICSa

[Ano23-50] Anonymous. IEEE Computer Society has you covered! IEEE Computer Graphics and Applications, 43(1):C2, January/February 2023. CODEN ICGADZ.

ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2023:ICSf

[Ano23-51] Anonymous. IEEE Computer Society has you covered! IEEE Computer Graphics and Applications, 43(2):C2, March/ April 2023. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2023:ICSj

[Ano23-52] Anonymous. IEEE Computer Society has you covered! IEEE Computer Graphics and Applications, 43(3):123, May/June 2023. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2023:ICSp

[Ano23-53] Anonymous. IEEE Computer Society has you covered! IEEE Computer Graphics and Applications, 43(4):132, July/ August 2023. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2023:ICSs

[Ano23-54] Anonymous. IEEE Computer Society has you covered! IEEE Computer Graphics and Applications, 43(5):82, September/October 2023. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2023:ICSw

[Ano23-55] Anonymous. IEEE Computer Society has you covered! IEEE Computer Graphics and Applications, 43(6):21, November/ December 2023. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2023:ICSd

[Ano23-56] Anonymous. IEEE Computer Society information. IEEE Computer Graphics and Applications, 43 (1):112, January/February 2023. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2023:ICSg

[Ano23-57] Anonymous. IEEE Computer Society information. IEEE Computer Graphics and Applications, 43(2):113, March/April 2023. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2023:ICSk

[Ano23-58] Anonymous. IEEE Computer Society information. IEEE Computer Graphics and Applications, 43(3):124, May/June 2023. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2023:ICSn

[Ano23-59] Anonymous. IEEE Computer Society information. IEEE Computer Graphics and Applications, 43(4):129, July/August 2023. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2023:ICSt

[Ano23-60] Anonymous. IEEE Computer Society information. IEEE Computer Graphics and Applications, 43(5):132, September/October 2023. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2023:ICSx

[Ano23-61] Anonymous. IEEE Computer Society information. IEEE Computer Graphics and Applications, 43(6):143, November/December 2023. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2023:ICSe

[Ano23-62] Anonymous. IEEE Computer Society Jobs Board. IEEE Computer Graphics and Applications, 43(1): C3, January/February 2023. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2023:ICSb

[Ano23-63] Anonymous. IEEE Computer Society: Publications seek 2024 Editors in Chief. IEEE Computer Graphics and Applications, 43(1): 1, January/February 2023. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2023:ICSc

[Ano23-64] Anonymous. IEEE Computer Society volunteer service awards. IEEE Computer Graphics and Applications, 43(1):110, January/ February 2023. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2023:ICSi

[Ano23-65] Anonymous. IEEE Computer Society Volunteer Service Awards. IEEE Computer Graphics and Applications, 43(3):122, May/ June 2023. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2023:IQWa

[Ano23-66] Anonymous. IEEE quantum week. IEEE Computer Graphics and Applications, 43(3):1, May/June 2023. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2023:IQWb

[Ano23-67] Anonymous. IEEE quantum week. IEEE Computer Graphics and Applications, 43(4):C2, July/ August 2023. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2023:Ma

[Ano23-68]  Anonymous.  Masthead.

IEEE Computer Graphics and Applications, 43(1): 4, January/February 2023. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2023:Mb

[Ano23-69]  Anonymous.  Masthead.

IEEE Computer Graphics and Applications, 43(2):4, March/April 2023. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2023:Mc

[Ano23-70]  Anonymous.  Masthead.

IEEE Computer Graphics and Applications, 43(3):4, May/June 2023. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2023:Md

[Ano23-71]  Anonymous.  Masthead.

IEEE Computer Graphics and Applications, 43(4):4, July/August 2023. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2023:Me

[Ano23-72]  Anonymous.  Masthead.

IEEE Computer Graphics and Applications, 43(5):4, September/October 2023. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2023:Mf

[Ano23-73]  Anonymous.  Masthead.

IEEE Computer Graphics and Applications, 43(6):4, November/December 2023. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2023:RCSa

[Ano23-74] Anonymous. Over the rainbow: 21st century security & privacy podcast. IEEE Computer Graphics and Applications, 43(3):23, May/ June 2023. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2023:RCSb

[Ano23-75] Anonymous. Over the rainbow: 21st century security & privacy podcast. IEEE Computer Graphics and Applications, 43(4):96, July/ August 2023. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2023:RCSc

[Ano23-76] Anonymous. Over the rainbow: 21st century security & privacy podcast. IEEE Computer Graphics and Applications, 43(5):98, September/October 2023. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2023:RCSd

[Ano23-77]  Anonymous.  Over the

rainbow: 21st century security & privacy podcast. IEEE Computer Graphics and Applications, 43(6):45, November/December 2023. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2023:TCa

[Ano23-78]  Anonymous.  Table of

contents. IEEE Computer Graphics and Applications, 43(1):2–3, January/ February 2023. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2023:TCb

[Ano23-79]  Anonymous.  Table of

contents. IEEE Computer Graphics and Applications, 43(2):2–3, March/ April 2023. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2023:TCc

[Ano23-80] Anonymous. Table of contents. IEEE Computer Graphics and Applications, 43(3):2–3, May/June 2023. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2023:TCd

[Ano23-81]  Anonymous.  Table of

contents. IEEE Computer Graphics and Applications, 43(4):2–3, July/ August 2023. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2023:TCe

[Ano23-82]  Anonymous.  Table of

contents. IEEE Computer Graphics and Applications, 43(5):2–3, September/October 2023. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2023:TCf

[Ano23-83]  Anonymous.  Table of

contents. IEEE Computer Graphics and Applications, 43(6):2–3, November/ December 2023. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2023:WHSa

[Ano23-84]  Anonymous.  Watts S.

Humphrey Software Quality Award. IEEE Computer Graphics and Applications, 43(2):22, March/April 2023. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2023:WHSb

[Ano23-85]  Anonymous.  Watts S.

Humphrey Software Quality Award. IEEE Computer Graphics and Applications, 43(3):9, May/June 2023. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2023:WHSc

[Ano23-86]  Anonymous.  Watts S.

Humphrey Software Quality Award. IEEE Computer Graphics and Applications, 43(4):80, July/August 2023. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2024:CSEb

[Ano24a] Anonymous. Computing in Science & Engineering. IEEE Computer Graphics and Applications, 44(2):72, March/April 2024. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2024:Ca

[Ano24b] Anonymous. ComputingEdge. IEEE Computer Graphics and Applications, 44(1): C4, January/February 2024. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2024:Cb

[Ano24c] Anonymous. ComputingEdge.

IEEE Computer Graphics and Applications, 44(2):C4, March/April 2024. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2024:IAHa

[Ano24d] Anonymous. IEEE Annals of the History of Computing. IEEE Computer Graphics and Applications, 44(1): 39, January/February 2024. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2024:IAHb

[Ano24e] Anonymous. IEEE Annals of the History of Computing. IEEE Computer Graphics and Applications, 44(2):9, March/April 2024. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2024:CAIb

[Ano24f] Anonymous. Call for articles: IT Professional. IEEE Computer Graphics and Applications, 44(2):118, March/April 2024. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2024:CAIa

[Ano24g] Anonymous. Call for articles: IT professional. IEEE Computer Graphics and Applications, 44(1):104, January/February 2024. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2024:CPIa

[Ano24h] Anonymous. Call for papers: IEEE Computer Society. IEEE Computer Graphics and Applications, 44(1): 37, January/February 2024. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2024:CPIb

[Ano24i] Anonymous. Call for papers: IEEE Computer Society. IEEE Computer Graphics and Applications, 44(2):7, March/April 2024. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2024:CSEa

[Ano24j] Anonymous. Computing in science & engineering. IEEE Computer Graphics and Applications, 44(1): 85, January/February 2024. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2024:FCa

[Ano24k] Anonymous. Front cover. IEEE Computer Graphics and Applications, 44(1): C1, January/February 2024. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2024:FCb

[Ano24l] Anonymous. Front cover.

IEEE Computer Graphics and Applications, 44(2):C1, March/April 2024. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2024:GPN

[Ano24m] Anonymous. Get published in the new IEEE Open Journal of the Computer Society. IEEE Computer Graphics and Applications, 44(2):120, March/April 2024. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2024:ICSd

[Ano24n] Anonymous. IEEE Computer Society Career Center. IEEE Computer Graphics and Applications, 44(1): C3, January/February 2024. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2024:ICSh

[Ano24o] Anonymous. IEEE Computer Society Career Center. IEEE Computer Graphics and Applications, 44(2):C3, March/April 2024. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2024:ICSa

[Ano24p] Anonymous. IEEE Computer Society D&I fund. IEEE Computer Graphics and Applications, 44(1): C2, January/February 2024. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2024:ICSe

[Ano24q] Anonymous. IEEE Computer Society D&I Fund. IEEE Computer Graphics and Applications, 44(2):C2, March/April 2024. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2024:ICSc

[Ano24r] Anonymous. IEEE Computer Society has you covered! IEEE Computer Graphics and Applications, 44(1):7, January/February 2024. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2024:ICSf

[Ano24s] Anonymous. IEEE Computer Society has you covered! IEEE Computer Graphics and Applications, 44(2):1, March/April 2024. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2024:ICSb

[Ano24t] Anonymous. IEEE Computer Society information.

IEEE Computer Graphics and Applications, 44(1): 1, January/February 2024. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2024:ICSg

[Ano24u] Anonymous. IEEE Computer Society information. IEEE Computer Graphics and Applications, 44(2):119, March/April 2024. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2024:Ma

[Ano24v]  Anonymous.  Masthead.

IEEE Computer Graphics and Applications, 44(1): 4, January/February 2024. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2024:Mb

[Ano24w]  Anonymous.  Masthead.

IEEE Computer Graphics and Applications, 44(2):4, March/April 2024. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2024:TCa

[Ano24x]  Anonymous.  Table of

contents. IEEE Computer Graphics and Applications, 44(1):2–3, January/ February 2024. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2024:TCb

[Ano24y]  Anonymous.  Table of

contents. IEEE Computer Graphics and Applications, 44(2):2–3, March/ April 2024. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Agnus:2015:IIA

[ANS15] Vincent Agnus, Stephane Nicolau, and Luc Soler. Illumination independent and accurate marker tracking using cross-ratio invariance. IEEE Computer Graphics and Applications, 35(5):22–33, September/October 2015. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
cg/2015/05/mcg2015050022-
abs.html.

Angelelli:2014:IVA

[AOH^{+}14] Paolo Angelelli, Steffen Oeltze, Judit Haasz, Cagatay Turkay, Erlend Hodneland, Arvid Lundervold, Astri J. Lundervold, Bernhard Preim, and Helwig Hauser. Interactive visual analysis of heterogeneous cohort-study data. IEEE Computer Graphics and Applications, 34(5):70–82, September/ October 2014. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://csdl.
computer.org/csdl/mags/
cg/2014/05/mcg2014050070-
abs.html.

Ambro:2010:IID

[AP10] Miha Ambroz and Ivan Prebil. i3Drive, a 3D interactive driving simulator. IEEE Computer Graphics and Applications, 30(2): 86–92, March/April 2010. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Abe:2011:SGP

[AP11] Yeuhi Abe and Jovan Popović. Simulating 2D gaits with a phase-indexed tracking controller. IEEE Computer Graphics and Applications, 31(4):22–33, July/August 2011. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Aerts:2021:RID

[APB^{+}21] Jan Aerts, Jannes Peeters, Jelmer Bot, Danai Kafetzaki, and Houda Lamqaddam. Remote instruction for data visualization design: a report from the trenches. IEEE Computer Graphics and Applications, 41(6): 15–24, November/December 2021. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Astheimer:1999:GEI

[AR99] Peter Astheimer and Lawrence Rosenblum. Guest Editors’ introduction: a business view of virtual reality. IEEE Computer Graphics and Applications, 19(6):28–29, November/ December 1999. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL http://dlib.
                                        computer.org/cg/books/
                                        cg1999/pdf/g6028.pdf.

Arbab:1990:SMB

[Arb90] Farhad Arbab. Set models and Boolean operations for solids and assemblies. IEEE Computer Graphics and Applications, 10(6):76–86, November/December 1990. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Alexander:2010:DEP

[ARL^{+}10] Oleg Alexander, Mike Rogers, William Lambeth, Jen-Yuan Chiang, Wan-Chun Ma, Chuan-Chang Wang, and Paul Debevec. The Digital Emily Project: Achieving a photorealistic digital actor. IEEE Computer Graphics and Applications, 30(4): 20–31, July/August 2010. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Arnold:2014:CGCa

[Arn14a] David Arnold. Computer graphics and cultural heritage: From one-way inspiration to symbiosis, Part 1. IEEE Computer Graphics and Applications, 34 (3):76–86, May/June 2014. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Arnold:2014:CGCb

[Arn14b] David Arnold. Computer graphics and cultural heritage, Part 2: Continuing inspiration for future tools. IEEE Computer Graphics and Applications, 34(4): 70–79, July/August 2014. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Augsdorfer:2017:CMS

[ARS17] Ursula H. Augsdörfer and Andreas Riffnaller-Schiefer. On the convergence of modeling and simulation. IEEE Computer Graphics and Applications, 37(4):8–13, July/ August 2017. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2017/04/mcg2017040008-
abs.html.

Angus:1996:VWA

[AS96] I. G. Angus and H. A. Sowizral. VRMosaic: Web access from within a virtual environment. IEEE Computer Graphics and Applications, 16(3):6–10, May/June 1996. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Ahrens:2010:GEI

[AS10] James Ahrens and Han-Wei Shen. Guest Editors’ introduction: Ultrascale visualization. IEEE Computer Graphics and Applications, 30(3):20–21, May/ June 2010. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Angel:2011:TSB

[AS11] Ed Angel and Dave Shreiner. Teaching a shader-based introduction to computer graphics. IEEE Computer Graphics and Applications, 31(2):9–13, March/ April 2011. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Ashikhmin:2003:FTT

[Ash03] Michael Ashikhmin. Fast texture transfer. IEEE Computer Graphics and Applications, 23(4):38–43, July/August 2003. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cg/2003/04/g4038abs.htm;

http://csdl.computer.
org/dl/mags/cg/2003/04/
g4038.htm;   http://csdl.
computer.org/dl/mags/cg/
2003/04/g4038.pdf.

Avila:1994:VNC

[ASK94] R. S. Avila, L. M. Sobierajski, and A. E. Kaufman. Visualizing nerve cells. IEEE Computer Graphics and Applications, 14(5):11– 13, September 1994. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Asal:1986:TIG

[ASP^{+}86] Michael D. Asal, Graham Short, Tom Preston, Richard Simpson, Derek Roskell, and Karl M. Guttag. The Texas Instruments 34010 Graphics System Processor. IEEE Computer Graphics and Applications, 6(10):24–39, October 1986. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Abram:2021:AWM

[ASP^{+}21] G. Abram, F. Samsel, M. R.

Petersen, X. Asay-Davis, D. Comeau, and S. F. Price. Antarctic water masses and ice shelves: Visualizing the physics. IEEE Computer Graphics and Applications, 41(1):35–41, January/ February 2021. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

An:2021:SDJ

[ASS^{+}21] Yifei An, Han-Wei Shen, Guihua Shan, Guan Li, and Jun Liu. STSRNet: Deep joint space time superresolution for vector field visualization. IEEE Computer Graphics and Applications, 41(6):122–132, November/ December 2021. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Akimoto:1993:ACF

[ASW93] Takaaki Akimoto, Yasuhito Suenaga, and Richard S. Wallace. Automatic creation of 3D facial models. IEEE Computer Graphics and Applications, 13 (5):16–22, September 1993. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Allen:2003:NMD

[ATS^{+}03] Peter K. Allen, Alejandro Troccoli, Benjamin Smith, Stephen Murray, Ioannis Stamos, and Marius Leordeanu. New methods for digital modeling of historic sites. IEEE Computer Graphics and Applications, 23(6):32–41, November/December 2003. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL

g6032abs.htm;   http://
csdl.computer.org/comp/
mags/cg/2003/06/extras/
g6032x1.avi;   http://
csdl.computer.org/dl/mags/
cg/2003/06/g6032.pdf.

Atwood:1984:RVW

[Atw84] Michael E. Atwood. Report on the Vail Workshop on Human Factors in Computer Systems. IEEE Computer Graphics and Applications, 4(12):48–66, December 1984. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Andujar:2020:SSA

[AVV20] C. Andujar, C. R. Vijulie, and À. Vinacua. Syntactic and semantic analysis for extended feedback on computer graphics assignments. IEEE Computer Graphics and Applications, 40(3):105–111, May/ June 2020. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Allison:1997:VRG

[AWB^{+}97] Don Allison, Brian Wills, Doug Bowman, Jean Wineman, and Larry F. Hodges. The virtual reality gorilla exhibit. IEEE Computer Graphics and Applications, 17(6):30–38, November/December 1997. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

                                        computer.org/cga/cg1997/
                                        g6030abs.htm.

Akiba:2010:ATB

[AWM10] Hiroshi Akiba, Chaoli Wang, and Kwan-Liu Ma. AniViz: a template-based animation tool for volume visualization. IEEE Computer Graphics and Applications, 30(5):61–71, September/October 2010. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Aliaga:2007:ORC

[AXP07] Daniel G. Aliaga, Yi Xu, and Voicu Popescu. Occlusionresistant camera design for acquiring active environments. IEEE Computer Graphics and Applications, 27(5):68–78, September/ October 2007. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Albrecht:2016:DIC

[AY16] Kim Albrecht and Burcu Yucesoy. Designing for insight: A case study from tennis player analysis. IEEE Computer Graphics and Applications, 36(4):102–108, July/August 2016. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/

cg/2016/04/mcg2016040102-
abs.html.

Asamura:1998:SSS

[AYS98] Naoya Asamura, Nozomu Yokoyama, and Hiroyuki Shinoda. Selectively stimulating skin receptors for tactile display. IEEE Computer Graphics and Applications, 18(6):32–37, November/December 1998. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

computer.org/cga/cg1998/
g6032abs.htm;   http:/
/dlib.computer.org/cg/
books/cg1998/pdf/g6032.
pdf.

Annanias:2022:IDS

[AZSW22] Yves Annanias, Dirk Zeckzer, Gerik Scheuermann, and Daniel Wiegreffe. An interactive decision support system for land reuse tasks. IEEE Computer Graphics and Applications, 42(6): 72–83, November/December 2022. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Boltz:1985:MDP

[BA85] Richard J. Boltz and J. Thomas Avery, Jr. Mechanical design productivity using CAD graphics — a user’s point of view. IEEE Computer Graphics and Applications, 5(2):40–44, February 1985. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Bickel:2013:CAF

[BA13] Bernd Bickel and Marc Alexa. Computational aspects of fabrication: Modeling, design, and 3D printing. IEEE Computer Graphics and Applications, 33(6): 24–25, November/December 2013. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Bogucka:2021:DIS

[BAAQ21] Edyta Paulina Bogucka, Bon Adriel Aseniero, Luca Maria Aiello, and Daniele Quercia. The Dreamcatcher: Interactive storytelling of dreams. IEEE Computer Graphics and Applications, 41(3): 105–112, May/June 2021. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Barajas:2023:LLQ

[BAB^{+}23a] Daniel J. Barajas, Xornam S. Apedoe, David G. Brizan, Alark P. Joshi, and Sophie J. Engle. Lessons learned from quantitatively exploring visualization rubric utilization for peer feedback. IEEE Computer Graphics and Applications, 43(1):10–21, January/ February 2023. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Bethel:2023:QCV

[BAB^{+}23b] E. Wes Bethel, Mercy G.

Amankwah, Jan Balewski, Roel Van Beeumen, Daan Camps, Daniel Huang, and Talita Perciano. Quantum computing and visualization: a disruptive technological change ahead. IEEE Computer Graphics and Applications, 43(6):101–111, November/December 2023. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Beauxis-Aussalet:2021:RIV

[BABB^{+}21] Emma Beauxis-Aussalet, Michael Behrisch, Rita Borgo, Duen Horng Chau, Christopher Collins, David Ebert, Mennatallah El-Assady, Alex Endert, Daniel A. Keim, Jörn Kohlhammer, Daniela Oelke, Jaakko Peltonen, Maria Riveiro, Tobias Schreck, Hendrik Strobelt, and Jarke J. van Wijk. The role of interactive visualization in fostering trust in AI. IEEE Computer Graphics and Applications, 41(6): 7–12, November/December 2021. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Bach:2016:UDN

[Bac16] Benjamin Bach. Unfolding dynamic networks for visual exploration. IEEE Computer Graphics and Applications, 36(2):74–82, March/ April 2016. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        cg/2016/02/mcg2016020074-
                                        abs.html.

Badler:1982:HBM

[Bad82]  N. I. Badler.  Human

body models and animation. IEEE Computer Graphics and Applications, 2(11):6–7, November/December 1982. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Badler:2000:A

[Bad00] Norman I. Badler. Animation 2000++. IEEE Computer Graphics and Applications, 20(1):28–29, January/ February 2000. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL http://dlib.
                                        computer.org/cg/books/
                                        cg2000/pdf/g1028.pdf.

Badler:2023:RVH

[Bad23] Norman I. Badler. On raising a virtual human. IEEE Computer Graphics and Applications, 43(6):133–140, November/December 2023. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Bock:2018:OCN

[BAE^{+}18] Alexander Bock, Emil Axelsson, Carter Emmart, Masha Kuznetsova, Charles Hansen, and Anders Ynnerman. OpenSpace: Changing the narrative of public dissemination in astronomical visualization from what to how. IEEE Computer Graphics and Applications, 38(3):44–57, May/ June 2018. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2018/03/mcg2018030044-
abs.html.

Bailey:1995:TRP

[Bai95] Michael J. Bailey. Telemanufacturing: Rapid prototyping on the Internet. IEEE Computer Graphics and Applications, 15(6): 20–26, November/December 1995. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Bailey:2001:VVI

[Bai01]  Mike Bailey.  Visualiza-

tion viewpoints: Interacting with direct volume rendering. IEEE Computer Graphics and Applications, 21(1):10–13, January/ February 2001. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://dlib.
computer.org/cg/books/
cg2001/pdf/g1010.pdf.

Bailey:2009:UGS

[Bai09] Mike Bailey. Using GPU shaders for visualization. IEEE Computer Graphics and Applications, 29(5): 96–100, September/October 2009. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Bailey:2011:UGS

[Bai11] Mike Bailey. Using GPU shaders for visualization, part 2. IEEE Computer Graphics and Applications, 31(2):67–73, March/April 2011. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Bailey:2013:UGS

[Bai13] Mike Bailey. Using GPU shaders for visualization, part 3. IEEE Computer Graphics and Applications, 33(3):5–11, May/June 2013. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Baker:1988:BVC

[Bak88] H. Harlyn Baker. Building, visualizing, and computing on surfaces of evolution. IEEE Computer Graphics and Applications, 8(4):31– 41, July/August 1988. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Baldwin:1997:NVF

[Bal97] Robin Baldwin. In the news: VR: Friend or foe? flat displays sharper, wider — but still not cheap; EE/CS encyclopedia will cover the fields; technology briefs. IEEE Computer Graphics and Applications, 17(6):102–103, November/ December 1997. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://dlib.
computer.org/cg/books/
cg1997/pdf/g6102.pdf.

Baldwin:1998:ACP

[Bal98a]  Robin Baldwin.  About

the cover: Projecting largescale graphics. IEEE Computer Graphics and Applications, 18(2):4–5, March/ April 1998. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://dlib.
computer.org/cg/books/
cg1998/pdf/g2004.pdf.

Baldwin:1998:NVY

[Bal98b] Robin Baldwin. In the news: Vis 97 — year of the sphere. IEEE Computer Graphics and Applications, 18(1):7, January/ February 1998. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

computer.org/cga/cg1998/
g1007abs.htm;   http:/
/dlib.computer.org/cg/
books/cg1998/pdf/g1007.
pdf.

Baldwin:1998:NVA

[Bal98c] Robin Baldwin. In the news: VRAIS 98 advances virtual reality; virtual theaters target arcades, museums; replicating artifacts in digital 3-D. IEEE Computer Graphics and Applications, 18(3):6–7, May/ June 1998. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL http://dlib.
                                        computer.org/cg/books/
                                        cg1998/pdf/g3006.pdf.

Barr:1981:SAT

[Bar81a] Alan H. Barr. Superquadrics and angle-preserving transformations. IEEE Computer Graphics and Applications, 1(1):11–23, January/ February 1981. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Barsky:1981:CGD

[Bar81b] B. A. Barsky. Computeraided geometric design: a bibliography with keywords and classified index. IEEE Computer Graphics and Applications, 1(3):67–106, July/August 1981. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Barnhill:1983:SRD

[Bar83] Robert E. Barnhill. A survey of the representation and design of surfaces. IEEE Computer Graphics and Applications, 3(7):9–16, October 1983. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Barsky:1984:DEV

[Bar84] Brian A. Barsky. A description and evaluation of various 3-D models. IEEE Computer Graphics and Applications, 4(1):38–52, January/ February 1984. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Barkans:1991:HPA

[Bar91] Anthony C. Barkans. Hardwareassisted polygon antialiasing. IEEE Computer Graphics and Applications, 11(1): 80–88, January/February 1991. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Barsky:1993:RBR

[Bar93] Brian A. Barsky. Rational beta-splines for representing curves and surfaces. IEEE Computer Graphics and Applications, 13(6):24–32, November/December 1993. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Baraff:1995:ISS

[Bar95] David Baraff. Interactive simulation of solid rigid bodies. IEEE Computer Graphics and Applications, 15 (3):63–75, May/June 1995. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Barkans:1997:CRT

[Bar97a] Anthony C. Barkans. Color recovery: True color 8-Bit interactive graphics. IEEE Computer Graphics and Applications, 17(1):67–77, January/February 1997. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Barzel:1997:FDR

[Bar97b] Ronen Barzel. Faking dynamics of ropes and springs. IEEE Computer Graphics and Applications, 17(3):31– 39, May/June 1997. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Bartlett:2000:RSH

[Bar00] Joel F. Bartlett. Rock ’n’ scroll is here to stay. IEEE Computer Graphics and Applications, 20(3):40– 45, May/June 2000. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

computer.org/cga/cg2000/
g3040abs.htm;   http:/
/dlib.computer.org/cg/
books/cg2000/pdf/g3040.
pdf.

Barbalet:2004:NAS

[Bar04] Tom Barbalet. Noble ape simulation. IEEE Computer Graphics and Applications, 24(2):6–12, March/ April 2004. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http:/

/csdl.computer.org/dl/
mags/cg/2004/02/g2006.
htm;   http://csdl.computer.
org/dl/mags/cg/2004/02/
g2006.pdf.

Barsky:2005:FF

[Bar05] Brian A. Barsky. A fond farewell. IEEE Computer Graphics and Applications, 25(4):6–7, July/ August 2005. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cg/2005/04/g4006.pdf.

Bartram:2015:DCO

[Bar15]  Lyn Bartram.  Design

challenges and opportunities for eco-feedback in the home. IEEE Computer Graphics and Applications, 35(4):52–62, July/ August 2015. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
cg/2015/04/mcg2015040052-
abs.html.

Basole:2014:VBE

[Bas14]  Rahul C. Basole.  Vi-

sual business ecosystem intelligence: Lessons from the field. IEEE Computer Graphics and Applications, 34(5):26–34, September/October 2014. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL http://csdl.
                                        computer.org/csdl/mags/
                                        cg/2014/05/mcg2014050026-
                                        abs.html.

Bettels:1982:CGH

[BB82] Juergen Bettels and Rene Brun. Computer graphics in high energy physics. IEEE Computer Graphics and Applications, 2(3):61– 64, 66–68, 70–72, July/ August 1982. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Brotman:1984:GSS

[BB84] Lynne Shapiro Brotman and Norman I. Badler. Generating soft shadows with a depth buffer algorithm. IEEE Computer Graphics and Applications, 4(10):5– 12, October 1984. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Baker:1995:VBA

[BB95] M. Pauline Baker and Colleen Bushell. Visualization blackboard: After the storm: considerations for information visualization. IEEE Computer Graphics and Applications, 15(3):12–15, May/ June 1995. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Boyer:2000:AAS

[BB00] Vincent Boyer and Jean-Jacques Bourdin. Autoadaptive step straight-line algorithm. IEEE Computer Graphics and Applications, 20(5):67–69, September/October 2000. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

computer.org/cga/cg2000/
g5067abs.htm;   http:/
/dlib.computer.org/cg/
books/cg2000/pdf/g5067.
pdf.

Bezerianos:2005:VSM

[BB05] Anastasia Bezerianos and Ravin Balakrishnan. View and space management on large displays. IEEE Computer Graphics and Applications, 25(4):34–43, July/ August 2005. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Boring:2013:MPD

[BB13] Sebastian Boring and Dominikus Baur. Making public displays interactive everywhere. IEEE Computer Graphics and Applications, 33(2):28–36, March/ April 2013. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Bruns:2008:MPE

[BBB08] Erich Bruns, Benjamin Brombach, and Oliver Bimber. Mobile phone-enabled museum guidance with adaptive classification. IEEE Computer Graphics and Applications, 28(4):98–102, July/August 2008. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Booth:1987:PHV

[BBC^{+}87] Kellogg S. Booth, M. Phillip Bryden, William B. Cowan, Michael F. Morgan, and Brian L. Plante. On the parameters of human visual performance: an investigation of the benefits of antialiasing. IEEE Computer Graphics and Applications, 7(9):34–41, September 1987. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Boukhelifa:2020:CEI

[BBC^{+}20] N. Boukhelifa, A. Bezerianos, R. Chang, C. Collins, S. Drucker, A. Endert, J. Hullman, C. North, and M. Sedlmair. Challenges in evaluating interactive visual machine learning systems. IEEE Computer Graphics and Applications, 40(6): 88–96, November/December 2020. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Billinghurst:1998:SID

[BBDM98] Mark Billinghurst, Jerry Bowskill, Nick Dyer, and Jason Morphett. Spatial information displays on a wearable computer. IEEE Computer Graphics and Applications, 18(6):24–31, November/December 1998. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

computer.org/cga/cg1998/
g6024abs.htm;   http:/
/dlib.computer.org/cg/
books/cg1998/pdf/g6024.
pdf.

Bottger:2008:MWA

[BBDZ08] Joachim Böttger, Ulrik Brandes, Oliver Deussen, and Hendrik Ziezold. Map warping for the annotation of metro maps. IEEE Computer Graphics and Applications, 28(5):56–65, September/October 2008. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Banks:1998:IVO

[BBF^{+}98] David C. Banks, Jay M.

Brown, John T. Foley, Kiril N. Vidimce, and Ming-Hoe Kiu. Interactive 3D visualization of optical phenomena. IEEE Computer Graphics and Applications, 18(4):66–69, July/ August 1998. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

                                        computer.org/cga/cg1998/
                                        g4066abs.htm;   http:/
                                        /dlib.computer.org/cg/
                                        books/cg1998/pdf/g4066.
                                        pdf.

Berk:1982:NCS

[BBK82] Toby Berk, Lee Brownston, and Arie Kaufman. A new color-naming system for graphics languages. IEEE Computer Graphics and Applications, 2(3):37–42, 44, May/June 1982. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Badouel:1994:DDC

[BBP94] Didier Badouel, Kadi Bouatouch, and Thierry Priol. Distributing data and control for ray tracing in parallel. IEEE Computer Graphics and Applications, 14(4): 69–77, July/August 1994. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Bratkova:2009:OPO

[BBS09] Margarita Bratkova, Solomon Boulos, and Peter Shirley. oRGB: a practical opponent color space for computer graphics. IEEE Computer Graphics and Applications, 29(1):42–55, January/ February 2009. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Brivio:2013:PIR

[BBT^{+}13] Paolo Brivio, Luca Benedetti, Marco Tarini, Federico Ponchio, Paolo Cignoni, and Roberto Scopigno. Photo-Cloud: Interactive remote exploration of joint 2D and 3D datasets. IEEE Computer Graphics and Applications, 33(2):86–96, c3, March/April 2013. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Bahar:1987:FTA

[BC87] Ezekiel Bahar and Swapan Chakrabarti. Fullwave theory applied to computer-aided graphics for 3D objects. IEEE Computer Graphics and Applications, 7(7):46–60, July/ August 1987. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Bailey:2005:GEI

[BC05] Mike Bailey and Steve Cunningham. Guest Editors’ introduction: Computer graphics in education. IEEE Computer Graphics and Applications, 25(5):23, September/October 2005. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL http://csdl.
                                        computer.org/comp/mags/
                                        cg/2005/05/g5023.pdf.

Barnard:2019:MBM

[BC19] M. Barnard and B. D. Campbell. Morgan Barnard: Melding our environment and the unseen supplied via data. IEEE Computer Graphics and Applications, 39(6):102–107, November/ December 2019. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Bogucka:2020:CDC

[BCA^{+}20] E. P. Bogucka, M. Constantinides, L. M. Aiello, D. Quercia, W. So, and M. Bancilhon. Cartographic design of cultural maps. IEEE Computer Graphics and Applications, 40(6): 12–20, November/December 2020. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Bowman:2008:SUI

[BCF^{+}08] Doug A. Bowman, Sabine Coquillart, Bernd Froehlich, Michitaka Hirose, Yoshifumi Kitamura, Kiyoshi Kiyokawa, and Wolfgang Stuerzlinger. Survey: 3D user interfaces: New directions and perspectives.

IEEE Computer Graphics and Applications, 28(6): 20–36, November/December 2008. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Burigat:2008:MVU

[BCI08] Stefano Burigat, Luca Chittaro, and Lucio Ieronutti. Mobrex: Visualizing users’ mobile browsing behaviors. IEEE Computer Graphics and Applications, 28(1):24– 32, January/February 2008. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Balogh:2016:SLV

[BCMF16] Brett Balogh, Anil Camci, Paul Murray, and Angus G. Forbes. Spectral landscapes: Visualizing electromagnetic interactions. IEEE Computer Graphics and Applications, 36(5):7–11, September/October 2016. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2016/05/mcg2016050007-
abs.html.

Billinghurst:2002:RWT

[BCPK02] Mark Billinghurst, Adrian Cheok, Simon Prince, and Hirokazu Kato. Real world teleconferencing. IEEE Computer Graphics and Applications, 22(6):11–13, November/December 2002.

CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL

                                        http://dlib.computer.org/
                                        cg/books/cg2002/pdf/g6011.
                                        pdf.

Bikakis:2022:SIM

[BCPS22] Nikos Bikakis, Panos K. Chrysanthis, George Papastefanatos, and Tobias Schreck. Special issue on machine learning approaches in big data visualization. IEEE Computer Graphics and Applications, 42(3):39–40, May/ June 2022. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Barsky:1985:BSS

[BD85] Brian A. Barsky and Tony D. DeRose. The beta2-spline: a special case of the betaspline curve and surface representation. IEEE Computer Graphics and Applications, 5(9):46–58, September 1985. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Barsky:1989:GCP

[BD89] Brian A. Barsky and Tony D. DeRose. Geometric continuity of parametric curves: Three equivalent characterizations. IEEE Computer Graphics and Applications, 9(6):60–69 (or 60–68??), November/December 1989. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Barsky:1990:GCP

[BD90] Brian A. Barsky and Tony D.

DeRose. Geometric continuity of parametric curves: Constructions of geometrically continuous splines. IEEE Computer Graphics and Applications, 10(1):60– 68, January/February 1990. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Byrd:2021:AWT

[BD21] Vetria L. Byrd and Nicole Dwenger. Activity worksheets for teaching and learning data visualization. IEEE Computer Graphics and Applications, 41(6): 25–36, November/December 2021. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Baranoski:2012:RDL

[BDC^{+}12] Gladimir V. G. Baranoski, Thomas Dimson, Tenn F. Chen, Bradley Kimmel, Daniel Yim, and Erik Miranda. Rapid dissemination of light transport models on the web. IEEE Computer Graphics and Applications, 32(3):10–15, May/ June 2012. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Brodlie:1995:GO

[BDDH95] Ken W. Brodlie, Ljiljana B. Damnjanovic, David A. Duce, and F. Robert A. Hopgood. GKS-94: An overview. IEEE Computer Graphics and Applications, 15(6):64–71, November/December 1995. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Brady:1999:VTM

[BDG^{+}99] Mary Brady, Alden Dima, Len Gebase, Michael Kass, Carmelo Montanez-Rivera, and Lynne Rosenthal. VRML testing: Making VRML worlds look the same everywhere. IEEE Computer Graphics and Applications, 19(2):59–67, March/ April 1999. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

computer.org/cga/cg1999/
g2059abs.htm;   http:/
/dlib.computer.org/cg/
books/cg1999/pdf/g2059.
pdf.

Beacon:1989:BEU

[BDH^{+}89] Geof R. Beacon, J. R. Dodsworth, S. E. Howe, R. Gordon Oliver, and Anthony Saia. Boundary evaluation using inner and outer sets: The ISOS method.

IEEE Computer Graphics and Applications, 9(2):39– 51, March/April 1989. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Ben-Dor:1986:NGC

[BDJ86] Avner Ben-Dor and Brian Robert Jones. New graphics controller for electrostatic plotting. IEEE Computer Graphics and Applications, 6(1):16–25, January/ February 1986. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Basdogan:2004:HMI

[BDK^{+}04] Cagatay Basdogan, Suvranu De, Jung Kim, Manivannan Muniyandi, Hyun Kim, and Mandayam A. Srinivasan. Haptics in minimally invasive surgical simulation and training. IEEE Computer Graphics and Applications, 24(2):56–64, March/ April 2004. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cg/2004/02/g2056abs.htm;
http://csdl.computer.
org/dl/mags/cg/2004/02/
g2056.htm;   http://csdl.
computer.org/dl/mags/cg/
2004/02/g2056.pdf.

Bliss:2002:GEI

[BDM02] Frank Bliss, John Dill, and Carl Machover. Guest Editors’ introduction: Graphics in advanced computeraided design. IEEE Computer Graphics and Applications, 22(3):22–23, May/ June 2002. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

computer.org/cga/cg2002/
g3022abs.htm;   http:/
/dlib.computer.org/cg/
books/cg2002/pdf/g3022.
pdf.

Ben-Ezra:2011:DGL

[BE11] Moshe Ben-Ezra. A digital gigapixel large-format tilescan camera. IEEE Computer Graphics and Applications, 31(1):49–61, January/ February 2011. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Bradley:2018:VDH

[BEAC^{+}18] Adam James Bradley, Mennatallah El-Assady, Katharine Coles, Eric Alexander, Min Chen, Christopher Collins, Stefan Jänicke, and David Joseph Wrisley. Visualization and the digital humanities:. IEEE Computer Graphics and Applications, 38(6):26–38, ???? 2018. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL https:

//ieeexplore.ieee.org/
document/8617736/.

Becker:1997:UMK

[Bec97] Barry G. Becker. Using MineSet for knowledge discovery. IEEE Computer Graphics and Applications, 17(4): 75–78, July/August 1997. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Beeby:1982:FIC

[Bee82]  W. Beeby.  The future

of integrated CAD/CAM systems: the Boeing perspective. IEEE Computer Graphics and Applications, 2(1):51–56, January/ February 1982. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Beeler:2015:PSG

[Bee15] Thabo Beeler. Passive spatiotemporal geometry reconstruction of human faces at high fidelity. IEEE Computer Graphics and Applications, 35(3):82–90, May/ June 2015. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
cg/2015/03/mcg2015030082-
abs.html.

Bresenham:1992:GEI

[BEG92] Jack Bresenham, Nick England, and Karl Guttag. Guest Editors’ introduction: Semiconductor technology for graphics and imaging.

IEEE Computer Graphics and Applications, 12(6): 40–??, November/December 1992. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Bono:1982:GFG

[BEHt82] Peter R. Bono, Jose L. Encarnacao, F. Robert A. Hopgood, and Paul J. W. ten Hagen. GKS — the first graphics standard. IEEE Computer Graphics and Applications, 2(3):9–15, 17–19, 21–23, July/August 1982. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Blenkhorn:2003:SME

[BEK^{+}03] Paul Blenkhorn, Gareth Evans, Alasdair King, Sri Hastuti Kurniawan, and Alistair Sutcliffe. Screen magnifiers: Evolution and evaluation. IEEE Computer Graphics and Applications, 23(5):54–61, September/October 2003. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http:

//csdl.computer.org/dl/
mags/cg/2003/05/g5054.
pdf;   http://csdl.computer.
org/dl/mags/cg/2003/05/
g5054abs.htm.

Benedens:1999:GBW

[Ben99] Oliver Benedens. Geometrybased watermarking of 3D models. IEEE Computer Graphics and Applications, 19(1):46–55, January/ February 1999. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

computer.org/cga/cg1999/
g1046abs.htm;   http:/
/dlib.computer.org/cg/
books/cg1999/pdf/g1046.
pdf.

Berg:1983:CGD

[Ber83] Charles R. Berg. Computer graphics displays: Windows for process control. IEEE Computer Graphics and Applications, 3(3):43–48, 50, 52–55, May/June 1983. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Bergeron:1986:GVC

[Ber86] P. Bergeron. A general version of Crow’s shadow volumes. IEEE Computer Graphics and Applications, 6(9):17–28, September 1986. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Berta:1999:PVI

[Ber99] Julien Berta. Projects in VR: Integrating VR and CAD. IEEE Computer Graphics and Applications, 19(5):14–19, September/ October 1999. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://dlib.

                                        computer.org/cg/books/
                                        cg1999/pdf/g5014.pdf.

Barnes:2009:GEI

[BES09] Tiffany Barnes, L. Miguel Encarnação, and Christopher D. Shaw. Guest Editors’ introduction: Serious games. IEEE Computer Graphics and Applications, 29(2):18–19, March/ April 2009. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Bender:2017:MVH

[BES17] Jan Bender, Kenny Erleben, and Barbara Solenthaler. Modeling virtual humans. IEEE Computer Graphics and Applications, 37(6):26–27, November/ December 2017. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        cg/2017/06/mcg2017060026.
                                        html.

Bethel:2000:VVV

[Bet00]  Wes Bethel.  Visualiza-

tion viewpoints: Visualization dot com. IEEE Computer Graphics and Applications, 20(3):17–20, May/ June 2000. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL http://dlib.
                                        computer.org/cg/books/
                                        cg2000/pdf/g3017.pdf.

Becker:1991:BNV

[BEW91] Richard A. Becker, Stephen G.

Eick, and Allan R. Wilks. Basics of network visualization. IEEE Computer Graphics and Applications, 11(3):12–14, May/ June 1991. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Bishop:1998:DPG

[BEW^{+}98] Lars Bishop, Dave Eberly, Turner Whitted, Mark Finch, and Michael Shantz. Designing a PC game engine. IEEE Computer Graphics and Applications, 18(1):46–53, January/ February 1998. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

computer.org/cga/cg1998/
g1046abs.htm;   http:/
/dlib.computer.org/cg/
books/cg1998/pdf/g1046.
pdf.

Beshers:1993:AVR

[BF93] Clifford Beshers and Steven Feiner. Auto Visual: Rulebased design of interactive multivariate visualizations. IEEE Computer Graphics and Applications, 13(4): 41–49, July/August 1993. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Borghese:1998:AFP

[BFB^{+}98] Nunzio Alberto Borghese, Giancarlo Ferrigno, Guido Baroni, Antonio Pedotti, Stefano Ferrari, and Riccardo Savarè. Autoscan: a flexible and portable 3D scanner. IEEE Computer Graphics and Applications, 18(3):38–41, May/ June 1998. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

computer.org/cga/cg1998/
g3038abs.htm;   http:/
/dlib.computer.org/cg/
books/cg1998/pdf/g3038.
pdf.

Buxton:2000:LDA

[BFBK00] William Buxton, George Fitzmaurice, Ravin Balakrishnan, and Gordon Kurtenbach. Large displays in automotive design. IEEE Computer Graphics and Applications, 20(4):68–75, July/ August 2000. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

computer.org/cga/cg2000/
g4068abs.htm;   http:/
/dlib.computer.org/cg/
books/cg2000/pdf/g4068.
pdf.

Beck:1986:SAM

[BFH86] James M. Beck, Rida T. Farouki, and John K. Hinds. Surface analysis methods. IEEE Computer Graphics and Applications, 6(12):18– 36, December 1986. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Bell:2002:VVI

[BFH02] Blaine Bell, Steven Feiner, and Tobias Höllerer. Visualization viewpoints: Information at a glance. IEEE Computer Graphics and Applications, 22(4):6–9, July/ August 2002. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

computer.org/cga/cg2002/
g4006abs.htm;   http:/
/dlib.computer.org/cg/
books/cg2002/pdf/g4006.
pdf.

Blomberg:1984:NAT

[BFK^{+}84] Lars Blomberg, Kerstin Freckner, Bjorn Kruse, Gunilla Lonnemark, Staffan Romberger, and Yngve Sunblad. A new approach to text and image processing. IEEE Computer Graphics and Applications, 4(7): 12–22, July/August 1984. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Booth:1986:HAZ

[BFP86] Kellogg S. Booth, David R.

Forsey, and Alan W. Paeth. Hardware assistance for Z-Buffer visible surface algorithms. IEEE Computer Graphics and Applications, 6(11):31–39, November/December 1986. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Beasley:2021:TLG

[BFR21] Zachariah J. Beasley, Alon Friedman, and Paul Rosen. Through the looking glass: Insights into visualization pedagogy through sentiment analysis of peer review text. IEEE Computer Graphics and Applications, 41(6): 59–70, November/December 2021. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Bimber:2001:VS

[BFSE01] Oliver Bimber, Bernd Fröhlich, Dieter Schmalstieg, and L. Miguel Encarnação. The virtual showcase. IEEE Computer Graphics and Applications, 21(6):48–55, November/December 2001. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL

                                        http://dlib.computer.org/
                                        cg/books/cg2001/g6048abs.
                                        htm;   http://dlib.computer.
                                        org/cg/books/cg2001/pdf/
                                        g6048.pdf.

Busch:1999:DWC

[BFW99] Christoph Busch, Wolfgang Funk, and Stephen Wolthusen. Digital watermarking: From concepts to real-time video applications. IEEE Computer Graphics and Applications, 19(1):25–35, January/ February 1999. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

computer.org/cga/cg1999/
g1025abs.htm;   http:/
/dlib.computer.org/cg/
books/cg1999/pdf/g1025.
pdf.

Boyse:1982:GIM

[BG82] John W. Boyse and Jack E.

Gilchrist. GMSOLID: Interactive modeling for design and analysis of solids. IEEE Computer Graphics and Applications, 2(2):27–30, 32, 34–40, March/April 1982. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Blinn:1996:JBCd

[BG96] James F. Blinn and Andrew Glassner. Jim Blinn corners Andrew Glassner’s notebook: Hey, buddy, how do I get into the Siggraph Electronic Theater. IEEE Computer Graphics and Applications, 16(6):72–75, November/December 1996. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Borgeat:2007:VAM

[BGM^{+}07] Louis Borgeat, Guy Godin, Philippe Massicotte, Guillaume Poirier, François Blais, and J.-Angelo Beraldin. Visualizing and analyzing the Mona Lisa. IEEE Computer Graphics and Applications, 27(6):60–68, November/December 2007. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL

                                        http://csdl.computer.org/
                                        comp/mags/cg/2007/extras/
                                        mcg2007060060x1.mov.

Bors:2019:CVP

[BGM19] C. Bors, T. Gschwandtner, and S. Miksch. Capturing and visualizing provenance from data wrangling. IEEE Computer Graphics and Applications, 39(6):61–75, November/December 2019. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Barillot:1985:RCB

[BGSC85] Christian Barillot, Bernard Gibaud, Jean-Marie M. Scarabin, and Jean-Louis L. Coatrieux. 3D reconstruction of cerebral blood vessels. IEEE Computer Graphics and Applications, 5(12):13–19, December 1985. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Bliss:1981:SIT

[BH81] Frank W. Bliss and George M. Hyman. Selecting and implementing a turnkey graphics system. IEEE Computer Graphics and Applications, 1(2):55–62, 64, 70, April 1981. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Bruggere:1983:TCE

[BH83] Thomas H. Bruggere and Elizabeth Hollomon. Tools for computer-aided engineering. IEEE Computer Graphics and Applications, 3(9):48–53, December 1983. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Burton:1984:LEG

[BH84] F. Warren Burton and Matthew M. Huntbach. Lazy evaluation of geometric objects. IEEE Computer Graphics and Applications, 4(1):28–33, January/ February 1984. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Brown:2006:MLA

[BH06] Leonard D. Brown and Hong Hua. Magic lenses for augmented virtual environments. IEEE Computer Graphics and Applications, 26(4):64–73, July/ August 2006. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Borland:2011:CSC

[BH11] David Borland and Alan Huber. Collaborationspecific color-map design. IEEE Computer Graphics and Applications, 31(4):7– 11, July/August 2011. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Berthaut:2016:SII

[BH16] Florent Berthaut and Martin Hachet. Spatial interfaces and interactive 3D environments for immersive musical performances. IEEE Computer Graphics and Applications, 36(5):82–87, September/ October 2016. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        cg/2016/05/mcg2016050082-
                                        abs.html.

Beyer:2013:ECP

[BHAA^{+}13] Johanna Beyer, Markus Hadwiger, Ali Al-Awami, Won-Ki Jeong, Narayanan Kasthuri, Jeff W. Lichtman, and Hanspeter Pfister. Exploring the Connectome: Petascale volume visualization of microscopy data streams. IEEE Computer Graphics and Applications, 33(4):50–61, July/ August 2013. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Bosnar:2023:PDM

[BHG23] Lovro Bosnar, Hans Hagen, and Petra Gospodnetic. Procedural defect modeling for virtual surface inspection environments. IEEE Computer Graphics and Applications, 43(2):13–22, March/ April 2023. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Bach:2021:SIV

[BHH^{+}21] Benjamin Bach, Samuel Huron, Uta Hinrichs, Jonathan C. Roberts, and Sheelagh Carpendale. Special issue on visualization teaching and literacy. IEEE Computer Graphics and Applications, 41(6):13–14, November/December 2021. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Blum:1991:RFI

[BHK91] Christof Blum, Georg Rainer Hofmann, and Detlef Kromker. Requirements for the first international imaging standard. IEEE Computer Graphics and Applications, 11(2):61–70, March/April 1991. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Becher:2022:SVA

[BHM^{+}22] Michael Becher, Dominik Herr, Christoph Müller, Kuno Kurzhals, Guido Reina, Lena Wagner, Thomas Ertl, and Daniel Weiskopf. Situated visual analysis and live monitoring for manufacturing. IEEE Computer Graphics and Applications, 42(2):33–44, March/ April 2022. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Bolter:1995:VBI

[BHMN95] J. Bolter, L. F. Hodges, T. Meyer, and A. Nichols. VR blackboard: Integrating perceptual and symbolic information in VR. IEEE Computer Graphics and Applications, 15(4):8–11, July/ August 1995. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Bruderlin:2007:IPI

[BHP07] Beat Brüderlin, Mathias Heyer, and Sebastian Pfützner. Interviews3D: a platform for interactive handling of massive data sets. IEEE Computer Graphics and Applications, 27(6): 48–59, November/December 2007. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Bernabei:2012:PAI

[BHPB^{+}12] Daniele Bernabei, Ajit Hakke-Patil, Francesco Banterle, Marco Di Benedetto, Fabio Ganovelli, Sumanta Pattanaik, and Roberto Scopigno. A parallel architecture for interactively rendering scattering and refraction effects. IEEE Computer Graphics and Applications, 32(2):34–43, March/ April 2012. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Basole:2012:VAC

[BHPS12] Rahul C. Basole, Mengdie Hu, Pritesh Patel, and John T. Stasko. Visual analytics for convergingbusiness-ecosystem intelligence. IEEE Computer Graphics and Applications, 32(1):92–96, January/ February 2012. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Berger:2011:CCS

[BHS11] Michael A. Berger, Gregor Hofer, and Hiroshi Shimodaira. Carnival — combining speech technology and computer animation. IEEE Computer Graphics and Applications, 31(5): 80–89, September/October 2011. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Borrelli:2024:ISV

[BHS^{+}24] Gabriel Borrelli, Lars Hagemann, Jannik Steinkühler, Adrian Derstroff, Marina Evers, Karim Huesmann, Simon Leistikow, Hennes Rave, Reyhaneh Sabbagh Gol, and Lars Linsen. 2022 IEEE scientific visualization contest winner: Multifield analysis of vorticitydriven lateral spread in wildfire ensembles. IEEE Computer Graphics and Applications, 44(1):40–49, January/ February 2024. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Bock:2018:OBN

[BHY18] Alexander Bock, Charles Hansen, and Anders Ynnerman. OpenSpace: Bringing NASA missions to the public. IEEE Computer Graphics and Applications, 38(5):112–118, September/ October 2018. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        cg/2018/05/mcg2018050112-
                                        abs.html.

Bimber:2006:AH

[Bim06] Oliver Bimber. Augmenting holograms. IEEE Computer Graphics and Applications, 26(5):12–17, September/October 2006. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cg/2006/extras/g5012x1.
avi;   http://www.holographics.
de/.

Bluff:2018:INA

[BJC18] Andrew Bluff, Andrew Johnston, and David Clarkson. Interaction, narrative and animation in live theatre. IEEE Computer Graphics and Applications, 38(2):8–14, March/ April 2018. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2018/02/mcg2018020008.
html.

Boukhelifa:2023:VDM

[BJP23] Nadia Boukhelifa, Chris R.

Johnson, and Kristin Potter. Visualization and decision making design under uncertainty. IEEE Computer Graphics and Applications, 43(5):23–25, September/October 2023. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Butz:2006:APM

[BK06] Andreas Butz and Antonio Krüger. Applying the peephole metaphor in a mixedreality room. IEEE Computer Graphics and Applications, 26(1):56–63, January/ February 2006. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Benes:2017:CDF

[BKLZ17] Bedrich Benes, David J. Kasik, Wilmot Li, and Hao Zhang. Computational design and fabrication. IEEE Computer Graphics and Applications, 37(3):32–33, May/June 2017. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        cg/2017/03/mcg2017030032.
                                        html.

Bera:2016:ICB

[BKM16] Aniket Bera, Sujeong Kim, and Dinesh Manocha. Interactive crowd-behavior learning for surveillance and training. IEEE Computer Graphics and Applications, 36(6):37–45, November/ December 2016. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        cg/2016/06/mcg2016060037-
                                        abs.html.

Billinghurst:2001:PVM

[BKP01] Mark Billinghurst, Hirokazu Kato, and Ivan Poupyrev. Projects in VR: The Magic-Book — moving seamlessly between reality and virtuality. IEEE Computer Graphics and Applications, 21(3):

6–8, May/June 2001. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http:

//dlib.computer.org/cg/
books/cg2001/pdf/g3006.
pdf.

Barris:1988:MSU

[BKRE88] Wesley C. Barris, Sridhar Kota, Donald R. Riley, and Arthur G. Erdman. Mechanism synthesis using the workstation environment. IEEE Computer Graphics and Applications, 8(2):39–50, March/ April 1988. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
cg/02/00000502-abs.html;
http://www.computer.org/
csdl/mags/cg/02/mcg1988020039-
abs.html.

Bimber:2019:SAI

[BKS19] O. Bimber, I. Kurmi, and D. C. Schedl. Synthetic aperture imaging with drones. IEEE Computer Graphics and Applications, 39(3):8–15, May/June 2019. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Bares:2020:CRV

[BKS20] A. Bares, D. F. Keefe, and F. Samsel. Close reading for visualization evaluation. IEEE Computer Graphics and Applications, 40(4): 84–95, July/August 2020. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Bustos:2007:CBO

[BKSS07] Benjamin Bustos, Daniel Keim, Dietmar Saupe, and Tobias Schreck. Contentbased 3D object retrieval. IEEE Computer Graphics and Applications, 27(4): 22–27, July/August 2007. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Borufka:1982:DCM

[BKt82] H. G. Borufka, Herbert W. Kuhlmann, and Paul J. W. ten Hagen. Dialogue cells: a method for defining interactions. IEEE Computer Graphics and Applications, 2(5):25–27, 29–30, 32–33, July/August 1982. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Borufka:1982:DCA

[BKtH82] H. G. Borufka, H. W. Kuhlmann, and P. J. W. ten Hagen. Dialogue cells: a method for defining interactions. IEEE Computer Graphics and Applications, 2(7):25–33, July/ August 1982. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Bideau:2010:UVR

[BKV^{+}10] Benoit Bideau, Richard Kulpa, Nicolas Vignais, Sébastien Brault, Franck Multon, and Cathy Craig. Using virtual reality to analyze sports performance. IEEE Computer Graphics and Applications, 30(2): 14–21, March/April 2010. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Bryson:1992:VWT

[BL92] Steve Bryson and Creon Levit. The virtual wind tunnel. IEEE Computer Graphics and Applications, 12(4): 25–34, July/August 1992. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Bailey:2014:SBC

[BL14] Mike Bailey and Cathy Law. A summer Blender camp: Modeling, rendering, and animation for high school students. IEEE Computer Graphics and Applications, 34(1):65–67, January/ February 2014. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Broll:2008:TNG

[BLH^{+}08] Wolfgang Broll, Irma Lindt, Iris Herbst, Jan Ohlenburg, Anne-Kathrin Braun, and Richard Wetzel. Toward next-gen mobile AR games. IEEE Computer Graphics and Applications, 28(4): 40–48, July/August 2008. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Blinn:1987:JBC

[Bli87] James F. Blinn. Jim Blinn’s corner: How to write a paper for SIGGRAPH. IEEE Computer Graphics and Applications, 7(12):62–64, December 1987. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Blinn:1988:WLE

[Bli88a] J. F. Blinn. The world’s largest Easter egg and what came out of it. IEEE Computer Graphics and Applications, 8(2):16–23, March/ April 1988. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
cg/02/00000499-abs.html;
 http://www.computer.org/
csdl/mags/cg/02/mcg1988020016n-
abs.html.

Blinn:1988:JBCd

[Bli88b] James Blinn. Jim Blinn’s corner — fractional invisibility. IEEE Computer Graphics and Applications, 8(6): 77–84, November/December 1988. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Blinn:1988:JBCc

[Bli88c]  James F. Blinn.  Jim

Blinn’s corner — the 3D kaleidoscope. IEEE Computer Graphics and Applications, 8(5):9–14, September/October 1988. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Blinn:1988:JBCa

[Bli88d]  James F. Blinn.  Jim

Blinn’s corner: Me and my (fake) shadow. IEEE Computer Graphics and Applications, 8(1):82–86, January/ February 1988. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Blinn:1988:JBCb

[Bli88e] Jim Blinn. Jim Blinn’s corner: Where am I? what am I looking at? IEEE Computer Graphics and Applications, 8(4):76–81, July/ August 1988. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Blinn:1989:JBCe

[Bli89a]  James F. Blinn.  Jim

Blinn’s corner — cubic curve update. IEEE Computer Graphics and Applications, 9(6):70–72, November/December 1989. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Blinn:1989:JBCd

[Bli89b] James F. Blinn. Jim Blinn’s corner — optimal tubes. IEEE Computer Graphics and Applications, 9(5):8–13, September 1989. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Blinn:1989:JBCa

[Bli89c] James F. Blinn. Jim Blinn’s corner — what we need around here is more aliasing. IEEE Computer Graphics and Applications, 9(1):75– 77, January/February 1989. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Blinn:1989:JBCc

[Bli89d] James F. Blinn. Jim Blinn’s corner: How many different cubic curves are there? IEEE Computer Graphics and Applications, 9(3):78– 83, May/June 1989. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Blinn:1989:JBCb

[Bli89e] James F. Blinn. Jim Blinn’s corner: Return of the jaggy. IEEE Computer Graphics and Applications, 9(2):82– 89, March/April 1989. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Blinn:1990:JBCb

[Bli90a] James F. Blinn. Jim Blinn’s corner — the truth about texture mapping. IEEE Computer Graphics and Applications, 10(2):78–83, March/April 1990. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Blinn:1990:JBCa

[Bli90b] James F. Blinn. Jim Blinn’s corner — triage tables. IEEE Computer Graphics and Applications, 10(1):70– 75, January/February 1990. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Blinn:1990:JBCc

[Bli90c] Jim Blinn. Jim Blinn’s corner: Wonderful world of video. IEEE Computer Graphics and Applications, 10(3):83–87, May/ June 1990. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Blinn:1991:JBCb

[Bli91a]  James F. Blinn.  Jim

Blinn’s corner — WYS-BOAVRTWYG. IEEE Computer Graphics and Applications, 11(3):85–92, May/June 1991. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Blinn:1991:JBCa

[Bli91b] James F. Blinn. Jim Blinn’s corner: a trip down the graphics pipeline: Line clipping. IEEE Computer Graphics and Applications, 11(1):98–105, January/February 1991. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Blinn:1991:JBCc

[Bli91c] James F. Blinn. Jim Blinn’s corner: a trip down the graphics pipeline: Pixel coordinates. IEEE Computer Graphics and Applications, 11(4):81–85, July/ August 1991. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Blinn:1991:JBCd

[Bli91d] James F. Blinn. Jim Blinn’s corner: a trip down the graphics pipeline: Subpixelic particles. IEEE Computer Graphics and Applications, 11(5):86–90, September 1991. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Blinn:1992:JBCb

[Bli92a] James F. Blinn. Jim Blinn’s corner — uppers and downers. IEEE Computer Graphics and Applications, 12(2): 85–91, March/April 1992.

CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Blinn:1992:JBCc

[Bli92b] James F. Blinn. Jim Blinn’s corner — uppers and downers: Part 2. IEEE Computer Graphics and Applications, 12(3):80–85, May/ June 1992. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Blinn:1992:JBCa

[Bli92c]  James F. Blinn.  Jim

Blinn’s corner: a trip down the graphics pipeline: Grandpa, what does “viewport” mean? IEEE Computer Graphics and Applications, 12(1):83–87, January/ February 1992. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Blinn:1992:JBCd

[Bli92d] Jim Blinn. Jim Blinn’s corner. IEEE Computer Graphics and Applications, 12(4): 89–??, July/August 1992. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:1992:JBCb

[Bli92e] Jim Blinn. Jim Blinn’s corner. IEEE Computer Graphics and Applications, 12(5): 106–??, September 1992. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:1992:JBCc

[Bli92f] Jim Blinn. Jim Blinn’s corner. IEEE Computer Graphics and Applications, 12(6): 87–??, November/December 1992. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Blinn:1993:JBCd

[Bli93a] J. F. Blinn. Jim Blinn’s corner: Backface culling snags (rendering algorithm). IEEE Computer Graphics and Applications, 13(6):94–97, November/December 1993. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Blinn:1993:JBCb

[Bli93b]  James F. Blinn.  Jim

Blinn’s corner: a trip down the graphics pipeline: The homogeneous perspective transform. IEEE Computer Graphics and Applications, 13(3):75–80, May/ June 1993. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Blinn:1993:JBCa

[Bli93c] James F. Blinn. Jim Blinn’s corner: NTSC: Nice technology, super color. IEEE Computer Graphics and Applications, 13(2):17–23, March/ April 1993. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Blinn:1993:JBCc

[Bli93d] James F. Blinn. Jim Blinn’s corner: What’s the deal with the DCT? IEEE Computer Graphics and Applications, 13(4):78–83, July/ August 1993. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Blinn:1994:JBCd

[Bli94a] J. F. Blinn. Jim Blinn’s corner: Composting, part 2: practice. IEEE Computer Graphics and Applications, 14(6):78–82, November/December 1994. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Blinn:1994:JBCa

[Bli94b] J. F. Blinn. Jim Blinn’s corner: Farewell to Fortran. IEEE Computer Graphics and Applications, 14(2): 86–89, March/April 1994. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Blinn:1994:JBCb

[Bli94c] J. F. Blinn. Jim Blinn’s corner: Quantization error and dithering. IEEE Computer Graphics and Applications, 14(4):78–82, July/ August 1994. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Blinn:1994:JBCc

[Bli94d]  James F. Blinn.  Jim

Blinn’s corner: Compositing. 1. Theory. IEEE Computer Graphics and Applications, 14(5):83–87, September 1994. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Blinn:1995:JBCa

[Bli95a] J. F. Blinn. Jim Blinn’s corner: How to draw a sphere.

1. IEEE Computer Graphics and Applications, 15(1):78– 83, January/February 1995. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Blinn:1995:JBCb

[Bli95b] J. F. Blinn. Jim Blinn’s corner: How to draw a sphere.2. Coordinate systems. IEEE Computer Graphics and Applications, 15(2):70–76, March/April 1995. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Blinn:1995:JBCe

[Bli95c] J. F. Blinn. Jim Blinn’s corner: Three wrongs make a right. IEEE Computer Graphics and Applications, 15(6):90–93, November/December 1995. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Blinn:1995:JBCc

[Bli95d] Jim Blinn. Jim Blinn’s corner: How to attend a Siggraph Conference. IEEE Computer Graphics and Applications, 15(4):86–88, July/August 1995. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Blinn:1995:JBCd

[Bli95e] Jim F. Blinn. Jim Blinn’s corner: How to draw a sphere. 3. The hyperbolic horizon. IEEE Computer Graphics and Applications, 15(5):87–93, September 1995. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Blinn:1996:JBCa

[Bli96a] J. Blinn. Jim Blinn’s corner: Consider the lowly 2 × 2 matrix. IEEE Computer Graphics and Applications, 16(2):82–88, March/ April 1996. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Blinn:1996:JBCc

[Bli96b] J. Blinn. Jim Blinn’s corner: Fun with premultiplied alpha. IEEE Computer Graphics and Applications, 16(5):86–89, September 1996. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Blinn:1996:JBCb

[Bli96c] J. F. Blinn. Jim Blinn’s corner: Calculating screen coverage. IEEE Computer Graphics and Applications, 16(3):84–88, May/ June 1996. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Blinn:1996:JBCe

[Bli96d] Jim Blinn. Jim Blinn’s Corner: a Trip Down the Graphics Pipeline. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 1996. ISBN 1-55860-387-5. vii + 214 pp. LCCN T385.B585 1996. This book is a collection of columns from the journal IEEE Computer Graphics and Applications.

Blinn:1997:JBCb

[Bli97a]  James F. Blinn.  Jim

Blinn’s corner: Floatingpoint tricks. IEEE Computer Graphics and Applications, 17(4):80–82, July/ August 1997. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). Discusses use of IEEE 754 single-precision floating-point bit patterns as integers for implementations of fast, but low-accuracy, functions useful in computer graphics.

Blinn:1997:JBCa

[Bli97b] James F. Blinn. Jim Blinn’s corner: Fugue for MMX. IEEE Computer Graphics and Applications, 17(2):88– 93, March/April 1997. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). Makes several cogent comments about deficiencies in the Intel MMX pixel-processing instruction set [PWW97] for use in image compositing.

Blinn:1998:JBCa

[Bli98a] Jim Blinn. Jim Blinn’s corner: a ghost in a snowstorm. IEEE Computer Graphics and Applications, 18(1):79–84, January/ February 1998. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

computer.org/cga/cg1998/
g1079abs.htm;   http:/
/dlib.computer.org/cg/
books/cg1998/pdf/g1079.
pdf.

Blinn:1998:JBCf

[Bli98b] Jim Blinn. Jim Blinn’s Corner: Dirty Pixels. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 1998. ISBN 1-55860-455-3.

256 pp. LCCN T385.B586 1998. This book is a collection of columns from the journal IEEE Computer Graphics and Applications.

Blinn:1998:JBCd

[Bli98c] Jim Blinn. Jim Blinn’s corner: Ten more unsolved problems in computer graphics. IEEE Computer Graphics and Applications, 18(5):86–89, September/October 1998. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL http://dlib.
                                        computer.org/cg/books/
                                        cg1998/pdf/g5086.pdf.

Blinn:1998:JBCe

[Bli98d] Jim Blinn. Jim Blinn’s corner: The cross ratio. IEEE Computer Graphics and Applications, 18(6):78–80, November/ December 1998. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL http://dlib.
                                        computer.org/cg/books/
                                        cg1998/pdf/g6078.pdf.

Blinn:1998:JBCc

[Bli98e] Jim Blinn. Jim Blinn’s corner: Upon further consideration. IEEE Computer Graphics and Applications, 18(4):87–91, July/ August 1998. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Blinn:1998:JBCb

[Bli98f] Jim Blinn. Jim Blinn’s corner: W pleasure, W fun. IEEE Computer Graphics and Applications, 18(3):78– 82, May/June 1998. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http:

//dlib.computer.org/cg/
books/cg1998/pdf/g3078.
pdf.

Blinn:1999:JBCc

[Bli99a] James F. Blinn. Jim Blinn’s corner: How many different parametric cubic curves are there? part 2: The “Same” game. IEEE Computer Graphics and Applications, 19(6):88–92, November/December 1999. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http:

//dlib.computer.org/cg/
books/cg1999/pdf/g6088.

pdf. See printing error correction [Bli00a].

Blinn:1999:JBCb

[Bli99b] Jim Blinn. Jim Blinn’s corner: How many rational parametric cubic curves are there?: Part 1: Inflection points. IEEE Computer Graphics and Applications, 19(4):84–87, July/ August 1999. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://dlib.

                                        computer.org/cg/books/
                                        cg1999/pdf/g4084.pdf.

Blinn:1999:JBCa

[Bli99c] Jim Blinn. Jim Blinn’s corner: Inferring transforms. IEEE Computer Graphics and Applications, 19(3):93– 98, May/June 1999. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http:

                                        //dlib.computer.org/cg/
                                        books/cg1999/pdf/g3093.
                                        pdf.

Blinn:2000:JBCa

[Bli00a]  James F. Blinn.  Jim

Blinn’s corner: a bright, shiny future. IEEE Computer Graphics and Applications, 20(1):18–19, January/ February 2000. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL http://dlib.
                                        computer.org/cg/books/
                                        cg2000/pdf/g1018.pdf.

Blinn:2000:JBCc

[Bli00b] James F. Blinn. Jim Blinn’s corner: How many different rational parametric cubic curves are there? part 3: The catalog. IEEE Computer Graphics and Applications, 20(2):85–88, March/ April 2000. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL http://dlib.
                                        computer.org/cg/books/
                                        cg2000/pdf/g2085.pdf.

Blinn:2000:JBCb

[Bli00c] James F. Blinn. Jim Blinn’s corner … correction. IEEE Computer Graphics and Applications, 20(1):69, January/February 2000. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http:

//dlib.computer.org/cg/
books/cg2000/pdf/g1069.

pdf. See [Bli99a].

Blinn:2000:JBCd

[Bli00d] James F. Blinn. Jim Blinn’s corner: Optimizing C++ vector expressions. IEEE Computer Graphics and Applications, 20(4):97–103, July/August 2000. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://dlib.
computer.org/cg/books/
cg2000/pdf/g4097.pdf.

Blinn:2000:JBCe

[Bli00e] Jim Blinn. Jim Blinn’s corner: Polynomial discriminants part 1: Matrix magic. IEEE Computer Graphics and Applications, 20(6):94–98, November/December 2000. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http:

//dlib.computer.org/cg/
books/cg2000/pdf/g6094.
pdf.

Blinn:2001:JBCa

[Bli01a] Jim Blinn. Jim Blinn’s corner: Polynomial discriminants: Part 2: Tensor diagrams. IEEE Computer Graphics and Applications, 21(1):86–92, January/ February 2001. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://dlib.
computer.org/cg/books/
cg2001/pdf/g1086.pdf.

Blinn:2001:JBCb

[Bli01b] Jim Blinn. Jim Blinn’s corner: Tensor contraction in C++. IEEE Computer Graphics and Applications, 21(2):88–92, March/ April 2001. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://dlib.
computer.org/cg/books/
cg2001/pdf/g2088.pdf.

Blinn:2002:JBCa

[Bli02a] James F. Blinn. Jim Blinn’s corner: Quartic discriminants and tensor invariants. IEEE Computer Graphics and Applications, 22(2):86– 91, March/April 2002. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

computer.org/cga/cg2002/
g2086abs.htm;   http:/
/dlib.computer.org/cg/
books/cg2002/pdf/g2086.
pdf.

Blinn:2002:JBCb

[Bli02b] James F. Blinn. Jim Blinn’s corner: Visualize whirled 2 × 2 matrices. IEEE Computer Graphics and Applications, 22(4):98–102, July/ August 2002. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

computer.org/cga/cg2002/
g4098abs.htm;   http:/
/dlib.computer.org/cg/
books/cg2002/pdf/g4098.
pdf.

Blinn:2003:JBCb

[Bli03a] James F. Blinn. Jim Blinn’s corner: Lines in space: Part 2: The line formulation. IEEE Computer Graphics and Applications, 23(3):72– 79, May/June 2003. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http:

//dlib.computer.org/cg/
books/cg2003/pdf/g3072.
pdf.

Blinn:2003:JBCc

[Bli03b] James F. Blinn. Jim Blinn’s corner: Lines in space: Part 3: The two matrices. IEEE Computer Graphics and Applications, 23(4):96– 101, July/August 2003. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http:

//csdl.computer.org/dl/
mags/cg/2003/04/g4096.
htm;   http://csdl.computer.

                                        org/dl/mags/cg/2003/04/
                                        g4096.pdf.

Blinn:2003:JBCd

[Bli03c]  James F. Blinn.  Jim

Blinn’s corner: Lines in space: Part 4: Back to the diagrams. IEEE Computer Graphics and Applications, 23(5):84–93, September/October 2003. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http:

                                        //csdl.computer.org/dl/
                                        mags/cg/2003/05/g5084.
                                        pdf.

Blinn:2003:JBCe

[Bli03d]  James F. Blinn.  Jim

Blinn’s corner: Lines in space: Part 5: a tale of two lines. IEEE Computer Graphics and Applications, 23(6):84–97, November/December 2003. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http:

                                        //csdl.computer.org/dl/
                                        mags/cg/2003/06/g6084.
                                        pdf.

Blinn:2003:JBCa

[Bli03e] Jim Blinn. Jim Blinn’s corner: Lines in space: Part 1: The 4D cross product. IEEE Computer Graphics and Applications, 23(2):84– 91, March/April 2003. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http:

//dlib.computer.org/cg/
books/cg2003/pdf/g2084.
pdf.

Blinn:2004:JBCa

[Bli04a]  James F. Blinn.  Jim

Blinn’s corner: Lines in space—part 6: Our friend the hyperbolic paraboloid. IEEE Computer Graphics and Applications, 24(3):92– 100, May/June 2004. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http:

//csdl.computer.org/dl/
mags/cg/2004/03/g3092.
htm;   http://csdl.computer.
org/dl/mags/cg/2004/03/
g3092.pdf.

Blinn:2004:JBCb

[Bli04b] James F. Blinn. Jim Blinn’s corner: Lines in space, part 7: The algebra of Tinkertoys. IEEE Computer Graphics and Applications, 24(4):96–102, July/ August 2004. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http:/

/csdl.computer.org/dl/
mags/cg/2004/04/g4096.
htm;   http://csdl.computer.
org/dl/mags/cg/2004/04/
g4096.pdf.

Blinn:2004:JBCc

[Bli04c] James F. Blinn. Jim Blinn’s corner: Lines in space, part 8: Line(s) through four lines. IEEE Computer Graphics and Applications, 24(5):100–106, September/ October 2004. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

                                        csdl.computer.org/comp/
                                        mags/cg/2004/05/extras/
                                        g5100x1.htm;   http://
                                        csdl.computer.org/dl/mags/
                                        cg/2004/05/g5100.htm;
                                        http://csdl.computer.org/
                                        dl/mags/cg/2004/05/g5100.
                                        pdf.

Blinn:2005:JBCc

[Bli05a] James F. Blinn. Jim Blinn’s corner: How to solve a quadratic equation. IEEE Computer Graphics and Applications, 25(6):76–79, November/December 2005. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Blinn:2005:JBCa

[Bli05b] James F. Blinn. Jim Blinn’s corner: User interface stories from the real world. IEEE Computer Graphics and Applications, 25(1):92– 93, January/February 2005. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Blinn:2005:VGO

[Bli05c] James F. Blinn. Jim Blinn’s corner: Vectors and geometry and objects, oh my! IEEE Computer Graphics and Applications, 25(3):84– 93, May/June 2005. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Blinn:2005:JBCb

[Bli05d]  James F. Blinn.  Jim

Blinn’s corner: What is a pixel? IEEE Computer Graphics and Applications, 25(5):82–87, September/October 2005. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Blinn:2006:JBCb

[Bli06a] James F. Blinn. Jim Blinn’s corner: How to solve a cubic equation, part 1: The shape of the discriminant. IEEE Computer Graphics and Applications, 26(3):84– 93, May/June 2006. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Blinn:2006:JBCc

[Bli06b] James F. Blinn. Jim Blinn’s corner: How to solve a cubic equation, part 2: The 1¯1 case. IEEE Computer Graphics and Applications, 26(4):90–100, July/August 2006. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Blinn:2006:HSC

[Bli06c] James F. Blinn. Jim Blinn’s corner: How to solve a cubic equation, part 3: General depression and a new covariant. IEEE Computer Graphics and Applications, 26(6):92–102, November/ December 2006. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Blinn:2006:JBCa

[Bli06d] James F. Blinn. Jim Blinn’s corner: How to solve a quadratic equation, part 2. IEEE Computer Graphics and Applications, 26(2): 82–87, March/April 2006. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Blinn:2007:JBCa

[Bli07a] James F. Blinn. Jim Blinn’s corner: How to solve a cubic equation, part 4: The 111 case. IEEE Computer Graphics and Applications, 27(1):100–103, January/ February 2007. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://bell.
computer.org/dlcomments/

.

Blinn:2007:JBCb

[Bli07b] James F. Blinn. Jim Blinn’s corner: How to solve a cubic equation, part 5: Back to numerics. IEEE Computer Graphics and Applications, 27(3):78–89, May/ June 2007. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Bloomenthal:1986:NAN

[Blo86] Jules Bloomenthal. Nature at New York Tech. IEEE Computer Graphics and Applications, 6(5):4–5, May/ June 1986. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Bettels:1986:PGS

[BM86] J. Bettels and D. R. Myers. The PIONS graphics system. IEEE Computer Graphics and Applications, 6(7):30–8, July/ August 1986. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Bly:1988:CS

[BM88] Sara A. Bly and William R.

Mallgren. Cycling softkeys. IEEE Computer Graphics and Applications, 8(5): 96–98, September/October 1988. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Boier-Martin:2003:AG

[BM03] Ioana M. Boier-Martin.

Adaptive graphics. IEEE Computer Graphics and Applications, 23(1):6–10, January/February 2003. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http:

//dlib.computer.org/cg/
books/cg2003/pdf/g1006.
pdf.

Basole:2024:GAV

[BM24] Rahul C. Basole and Timothy Major. Generative AI for visualization: Opportunities and challenges. IEEE Computer Graphics and Applications, 44(2): 55–64, March/April 2024. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Bacchelli-Montefusco:1984:UIG

[BMC84] Laura Bacchelli-Montefusco and Giulio Casciola. Using interactive graphics for fitting surfaces to scattered data. IEEE Computer Graphics and Applications, 4(7):43–45, July/ August 1984. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Bolas:2004:NRE

[BMC04] Mark Bolas, Ian McDowall, and Dan Corr. New research and explorations into multiuser immersive display systems. IEEE Computer Graphics and Applications, 24(1):18–21, January/February 2004. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http:

//csdl.computer.org/dl/

mags/cg/2004/01/g1018.
pdf.

Brogan:1998:DSC

[BMH98] David C. Brogan, Ronald A.

Metoyer, and Jessica K. Hodgins. Dynamically simulated characters in virtual environments. IEEE Computer Graphics and Applications, 18(5):58–69, September/October 1998. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

computer.org/cga/cg1998/
g5058abs.htm;   http:/
/dlib.computer.org/cg/
books/cg1998/pdf/g5058.
pdf.

Bishop:1990:VPE

[BMR90] Gary Bishop, Mark Monger, and Paul Ramsey. A visualization programming environment for multicomputers. IEEE Computer Graphics and Applications, 10(4): 50–58, July/August 1990. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Birr:2013:LWB

[BMS^{+}13] Steven Birr, Jeanette Monch, Dirk Sommerfeld, Uta Preim, and Bernhard Preim. The LiverAnatomyExplorer: A WebGL-based surgical teaching tool. IEEE Computer Graphics and Applications, 33(5):48–58, September/October 2013. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Badler:1987:AFP

[BMW87] Norman I. Badler, Kamran H. Manoochehri, and Graham Walters. Articulated figure positioning by multiple constraints. IEEE Computer Graphics and Applications, 7(6):28–38, June 1987. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Bajura:1995:DRC

[BN95] Michael Bajura and Ulrich Neumann. Dynamic registration correction in video-based augmented reality systems. IEEE Computer Graphics and Applications, 15(5):52–60, September 1995. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Ball:2022:SVF

[BNA^{+}22] Catherine Ball, Eric Novotny, Sun Joo Ahn, Lindsay Hahn, Michael D. Schmidt, Stephen L. Rathbun, and Kyle Johnsen. Scaling the virtual fitness buddy ecosystem as a school-based physical activity intervention for children. IEEE Computer Graphics and Applications, 42(1):105–115, January/February 2022. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Busch:1999:GEI

[BNP99] Christoph Busch, Klara Nahrstedt, and Ioannis Pitas. Guest Editors’ introduction: Image security. IEEE Computer Graphics and Applications, 19(1):16– 17, January/February 1999. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL

http://dlib.computer.org/
cg/books/cg1999/pdf/g1016.
pdf.

Badler:2009:GEI

[BO09] Norman I. Badler and Carol O’Sullivan. Guest Editors’ introduction: Virtual populace. IEEE Computer Graphics and Applications, 29(4):17–18, July/ August 2009. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Bohn:1995:RZP

[Boh95] Jan Helge Bohn. Removing zero-volume parts from CAD models for layered manufacturing. IEEE Computer Graphics and Applications, 15(6):27–34, November/December 1995. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Bolas:1994:HFD

[Bol94] Mark T. Bolas. Human factors in the design of an immersive display. IEEE Computer Graphics and Applications, 14(1):55–59, January/ February 1994. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Borden:1989:GPG

[Bor89] Bruce S. Borden. Graphics processing on a graphics supercomputer. IEEE Computer Graphics and Applications, 9(4):56–62, July/ August 1989. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Bowen:1993:DDM

[Bow93] William Bowen. Displays on display: Mapping US Census data with a simple “macroscope”. IEEE Computer Graphics and Applications, 13(5):9–11, September 1993. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Bochko:2006:SCA

[BP06] Vladimir Bochko and Jussi Parkkinen. A spectral color analysis and colorization technique. IEEE Computer Graphics and Applications, 26(5):74–82, September/October 2006. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Boulanger:2009:RGR

[BPB09] Kévin Boulanger, Sumanta N.

Pattanaik, and Kadi Bouatouch. Rendering grass in real time with dynamic lighting. IEEE Computer Graphics and Applications, 29(1):32–41, January/ February 2009. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Billinghurst:2014:HSG

[BPB14] Mark Billinghurst, Tham Piumsomboon, and Huidong Bai. Hands in space: Gesture interaction with augmented-reality interfaces. IEEE Computer Graphics and Applications, 34(1):77–80, January/February 2014. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Basole:2015:VAE

[BQP^{+}15] Rahul C. Basole, Ahsan Qamar, Hyunwoo Park, Christiaan J. J. Paredis, and Leon F. McGinnis. Visual analytics for early-phase complex engineered system design support. IEEE Computer Graphics and Applications, 35(2):41–51, March/ April 2015. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
cg/2015/02/mcg2015020041-
abs.html.

Basole:2019:UFM

[BQP^{+}19] R. C. Basole, A. Qamar, B. Pal, M. Corral, M. Meinhart, and A. Narechania. Understanding failure mode effect analysis data using interactive visual analytics. IEEE Computer Graphics and Applications, 39(6): 17–26, November/December 2019. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Brady:1986:TPC

[Bra86] James T. Brady. A theory of productivity in the creative process. IEEE Computer Graphics and Applications, 6(5):25–34, May/June 1986. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Barbulescu:2017:GAV

[BRB17] Adela Barbulescu, Remi Ronfard, and Gerard Bailly. A generative audio-visual prosodic model for virtual actors. IEEE Computer Graphics and Applications, 37(6):40–51, November/ December 2017. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2017/06/mcg2017060040-
abs.html.

Bach:2017:EGD

[BRCP17] Benjamin Bach, Nathalie Henry Riche, Sheelagh Carpendale, and Hanspeter Pfister. The emerging genre of data comics. IEEE Computer Graphics and Applications, 37(3):6–13, May/ June 2017. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2017/03/mcg2017030006-
abs.html.

Bevacqua:2017:BCP

[BRD17] Elisabetta Bevacqua, Romain Richard, and Pierre De Loor. Believability and copresence in human-virtual character interaction. IEEE Computer Graphics and Applications, 37(4):17–29, July/August 2017. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2017/04/mcg2017040017-
abs.html.

Bakalos:2020:MPC

[BRD^{+}20] N. Bakalos, I. Rallis, N. Doulamis, A. Doulamis, A. Voulodimos, and V. Vescoukis. Motion primitives classification using deep learning models for serious game platforms. IEEE Computer Graphics and Applications, 40(4):26–38, July/ August 2020. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Brentano:1984:DCM

[Bre84] Lewis Brentano. Distributed CAD/CAM: Myth and reality. IEEE Computer Graphics and Applications, 4(8):18–22, August 1984. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Bresenham:1987:AIL

[Bre87] Jack E. Bresenham. Ambiguities in incremental line rastering. IEEE Computer Graphics and Applications, 7(5):31–43, May/June 1987. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Bresenham:1990:SON

[Bre90] Jack Bresenham. Something old, new, borrowed, and blue. IEEE Computer Graphics and Applications, 10(5):42–44, September 1990. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Breen:1996:CGT

[Bre96a] David E. Breen. Guest Editor’s introduction: Computer graphics in textiles and apparel modeling. IEEE Computer Graphics and Applications, 16(5):26–27, September 1996. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Bresenham:1996:PF

[Bre96b] Jack Bresenham. Pixelprocessing fundamentals. IEEE Computer Graphics and Applications, 16(1):74– 82, January/February 1996. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Bresenham:1997:GEI

[Bre97] Jack Bresenham. Guest Editor’s introduction: Shape and motion modeling. IEEE Computer Graphics and Applications, 17(3):21, May/ June 1997. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Benford:2000:PPC

[BRG^{+}00] Steve Benford, Gail Reynard, Chris Greenhalgh, Dave Snowdon, and Adrian Bullock. A poetry performance in a collaborative virtual environment. IEEE Computer Graphics and Applications, 20(3):66– 75, May/June 2000. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

computer.org/cga/cg2000/
g3066abs.htm;   http:/
/dlib.computer.org/cg/
books/cg2000/pdf/g3066.
pdf.

Brittain:1990:PIG

[Bri90] Donald L. Brittain. Portability of interactive graphics software. IEEE Computer Graphics and Applications, 10(4):70–75, July/ August 1990. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Brunnett:2006:VPI

[BRL06] Guido Brunnett, Stephan Rusdorf, and Mario Lorenz. V-Pong: An immersive table tennis simulation. IEEE Computer Graphics and Applications, 26(4):10–13, July/August 2006. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Babaei:2017:FFP

[BRL^{+}17] Vahid Babaei, Javier Ramos, Yongquan Lu, Guillermo Webster, and Wojciech Matusik. FabSquare: Fabricating photopolymer objects by mold 3D printing and UV curing. IEEE Computer Graphics and Applications, 37(3):34–42, May/ June 2017. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        cg/2017/03/mcg2017030034-
                                        abs.html.

Bernardini:2002:BDM

[BRM^{+}02] Fausto Bernardini, Holly Rushmeier, Ioana M. Martin, Joshua Mittleman, and Gabriel Taubin. Building a digital model of Michelangelo’s Florentine Pietà. IEEE Computer Graphics and Applications, 22(1):59– 67, January/February 2002. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL

http://dlib.computer.org/
cg/books/cg2002/g1059abs.
htm;   http://dlib.computer.
org/cg/books/cg2002/pdf/
g1059.pdf.

Bartram:2010:CNV

[BRM10] Lyn Bartram, Johnny Rodgers, and Kevin Muise. Chasing the negawatt: Visualization for sustainable living. IEEE Computer Graphics and Applications, 30(3):8–14, May/ June 2010. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Brown:1982:PTS

[Bro82] Christopher M. Brown.

PADL-2: a technical summary. IEEE Computer Graphics and Applications, 2(2):68–84, March/April 1982. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Brown:1988:II

[Bro88] Maxine D. Brown. The interactive image. IEEE Computer Graphics and Applications, 8(1):10–16, January/ February 1988. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Brown:1989:UAF

[Bro89] Mark Brown. US Air Force relies on computer graphics to evaluate combat aircraft performance. IEEE Computer Graphics and Applications, 9(1):10–11, January/ February 1989. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Brooks:1999:WRA

[Bro99] Frederick P. Brooks, Jr. What’s real about virtual reality? IEEE Computer Graphics and Applications, 19(6):16–27, November/December 1999. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

                                        computer.org/cga/cg1999/
                                        g6016abs.htm;   http:/
                                        /dlib.computer.org/cg/
                                        books/cg1999/pdf/g6016.
                                        pdf.

Brochu:2015:TBS

[Bro15] Tyson Brochu. Toward better surface tracking for fluid simulation. IEEE Computer Graphics and Applications, 35(1):74–81, January/ February 2015. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://csdl.
computer.org/csdl/mags/
cg/2015/01/mcg2015010074-
abs.html.

Bryson:1997:PVV

[Bry97] Steve Bryson. Projects in VR: The virtual windtunnel on the Virtual Workbench. IEEE Computer Graphics and Applications, 17(4):15, July/August 1997. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Bier:1986:TPT

[BS86] E. A. Bier and K. R. Sloan, Jr. Two part texture mappings. IEEE Computer Graphics and Applications, 6(9):40–53, September 1986. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Blanc:1996:APC

[BS96] Carole Blanc and Christophe Schlick. Accurate parametrization of conics by NURBS. IEEE Computer Graphics and Applications, 16(6): 64–71, November/December 1996. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Bethel:2003:GDV

[BS03] E. Wes Bethel and John Shalf. Grid-distributed visualizations using connectionless protocols. IEEE Computer Graphics and Applications, 23(2):51–59, March/ April 2003. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://dlib.
computer.org/cg/books/
cg2003/pdf/g2051.pdf;
http://www.computer.org/
cga/cg2003/g2051abs.htm.

Basole:2016:SDV

[BS16] Rahul C. Basole and Dietmar Saupe. Sports data visualization [Guest Editors’ introduction]. IEEE Computer Graphics and Applications, 36(5):24–26, September/October 2016. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2016/05/mcg2016050024.
html.

Borro:2004:LHD

[BSA^{+}04] Diego Borro, Joan Savall, Aiert Amundarain, Jorge Juan Gil, Alejandro García-Alonso, and Luis Matey. A large haptic device for aircraft engine maintainability. IEEE Computer Graphics and Applications, 24(6):70–74, November/ December 2004. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http:

//csdl.computer.org/dl/
mags/cg/2004/06/g6070.
htm;   http://csdl.computer.
org/dl/mags/cg/2004/06/
g6070.pdf.

Broll:2001:PVI

[BSHB01] Wolfgang Broll, Leonie Schäfer, Tobias Höllerer, and Doug Bowman. Projects in VR: Interface with angels: The future of VR and AR interfaces. IEEE Computer Graphics and Applications, 21(6):14–17, November/ December 2001. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://dlib.
computer.org/cg/books/
cg2001/g6014abs.htm;
http://dlib.computer.org/
cg/books/cg2001/pdf/g6014.
pdf.

Basdogan:2007:VBS

[BSHW07] Cagatay Basdogan, Mert Sedef, Matthias Harders, and Stefan Wesarg. VRbased simulators for training in minimally invasive surgery. IEEE Computer Graphics and Applications, 27(2):54–66, March/April 2007. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Bernard:2015:VIS

[BSM^{+}15] Jurgen Bernard, David Sessler, Thorsten May, Thorsten Schlomm, Dirk Pehrke, and Jorn Kohlhammer. A visual-interactive system for prostate cancer cohort analysis. IEEE Computer Graphics and Applications, 35(3):44–55, May/ June 2015. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
cg/2015/03/mcg2015030044-
abs.html.

Bier:1997:EIU

[BSP97] Eric Bier, Maureen Stone, and Ken Pier. Enhanced illustration using magic lens filters. IEEE Computer Graphics and Applications, 17(6):62–70, November/December 1997. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

computer.org/cga/cg1997/
g6062abs.htm.

Benito-Santos:2020:DDI

[BSS20] A. Benito-Santos and R. T. Sánchez. A data-driven introduction to authors, readings, and techniques in visualization for the digital humanities. IEEE Computer Graphics and Applications, 40(3):45–57, May/ June 2020. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Berndt:2000:CHM

[BT00] Erhard Berndt and José Carlos Teixeira. Cultural heritage in the mature era of computer graphics. IEEE Computer Graphics and Applications, 20(1):36–37, January/February 2000. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http:

//dlib.computer.org/cg/
books/cg2000/pdf/g1036.
pdf.

Borland:2007:RCM

[BT07] David Borland and Russell M. Taylor II. Rainbow color map (Still) considered harmful. IEEE Computer Graphics and Applications, 27(2):14–17, March/ April 2007. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Bentley:2003:PP

[BTD^{+}03] Frank Bentley, Konrad Tollmar, David Demirdjian, Kimberle Koile, and Trevor Darrell. Perceptive presence. IEEE Computer Graphics and Applications, 23(5):26–36, September/October 2003. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http:

//csdl.computer.org/dl/
mags/cg/2003/05/g5026.
pdf;   http://csdl.computer.

                                        org/dl/mags/cg/2003/05/
                                        g5026abs.htm.

Berger:1990:RTM

[BTL90] Marc Berger, Terry Trout, and Nancy Levit. Ray tracing mirages. IEEE Computer Graphics and Applications, 10(3):36–41, May/ June 1990. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Berndt:2017:ESC

[BTM17] Iago Berndt, Rafael Torchelsen, and Anderson Maciel. Efficient surgical cutting with position-based dynamics. IEEE Computer Graphics and Applications, 37(3):24– 31, May/June 2017. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL https:/

                                        /www.computer.org/csdl/
                                        mags/cg/2017/03/mcg2017030024-
                                        abs.html.

Brewster:1984:ISP

[BTTU84] Linda J. Brewster, Sushma S. Trivedi, Heang K. Tuy, and Jayaram K. Udupa. Interactive surgical planning. IEEE Computer Graphics and Applications, 4(3):31– 40, March/April 1984. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Brunton:2023:CPL

[BU23] Alan Brunton and Philipp Urban. Cuttlefish: Pushing the limits of graphical 3-D printing. IEEE Computer Graphics and Applications, 43(5):114–121, September/ October 2023. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Buxton:2005:TSC

[Bux05] Bill Buxton. Thoughts on the state of 3D CG in film and video. IEEE Computer Graphics and Applications, 25(3):80–83, May/ June 2005. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Bschel:2019:ARG

[BVD19] W. Büschel, S. Vogt, and R. Dachselt. Augmented reality graph visualizations. IEEE Computer Graphics and Applications, 39(3):29– 40, May/June 2019. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Brown:1999:HCC

[BvDE^{+}99] Judith R. Brown, Andy van Dam, Rae Earnshaw, José Encarnação, Richard Guedj, Jennifer Preece, Ben Shneiderman, and John Vince. Human-centered computing, online communities, and virtual environments. IEEE Computer Graphics and Applications, 19(6):70–74, November/December 1999. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

                                        computer.org/cga/cg1999/
                                        g6070abs.htm;   http:/
                                        /dlib.computer.org/cg/
                                        books/cg1999/pdf/g6070.
                                        pdf.

Beardsley:2005:IUH

[BVRF05] Paul Beardsley, Jeroen Van Baar, Ramesh Raskar, and Clifton Forlines. Interaction using a handheld projector. IEEE Computer Graphics and Applications, 25(1):39– 43, January/February 2005. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Bethel:2011:VSC

[BvRS^{+}11] E. Wes Bethel, John van Rosendale, Dale Southard, Kelly Gaither, Hank Childs, Eric Brugger, and Sean Ahern. Visualization at supercomputing centers: The tale of little big iron and the three skinny guys. IEEE Computer Graphics and Applications, 31(1):90–95, January/February 2011. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Bartels:1994:ECP

[BW94] Richard H. Bartels and David R. Warn. Experiments with curvaturecontinuous patch-boundary fitting. IEEE Computer Graphics and Applications, 14(5):64–73, September 1994. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Bishop:2000:WOR

[BW00] Gary Bishop and Greg Welch. Working in the office of “Real soon now”. IEEE Computer Graphics and Applications, 20(4):76– 78, July/August 2000. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

computer.org/cga/cg2000/
g4076abs.htm;   http:/
/dlib.computer.org/cg/
books/cg2000/pdf/g4076.
pdf.

Barrus:1996:LSL

[BWA96] John W. Barrus, Richard C.

Waters, and David B. Anderson. Locales: Supporting large multiuser virtual environments. IEEE Computer Graphics and Applications, 16(6):50–57, November/December 1996. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Bors:2019:PTA

[BWD^{+}19] C. Bors, J. Wenskovitch, M. Dowling, S. Attfield, L. Battle, A. Endert, O. Kulyk, and R. S. Laramee. A provenance task abstraction framework. IEEE Computer Graphics and Applications, 39(6):46–60, November/December 2019. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Borland:2018:CV

[BWG18] David Borland, Wenyuan Wang, and David Gotz. Contextual visualization. IEEE Computer Graphics and Applications, 38(6):17– 23, ???? 2018. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL https:

                                        //ieeexplore.ieee.org/
                                        document/8617767/.

Borland:2024:UCI

[BWG24] David Borland, Arran Zeyu Wang, and David Gotz. Using counterfactuals to improve causal inferences from visualizations. IEEE Computer Graphics and Applications, 44(1):95–104, January/February 2024. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Brinkley:1999:DAI

[BWHR99] James F. Brinkley, Benjamin A. Wong, Kevin P. Hinshaw, and Cornelius Rosse. Design of an anatomy information system. IEEE Computer Graphics and Applications, 19(3):38–48, May/June 1999. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

computer.org/cga/cg1999/
g3038abs.htm;   http:/
/dlib.computer.org/cg/
books/cg1999/pdf/g3038.
pdf.

Bernhard:2016:AGF

[BWP^{+}16] Matthias Bernhard, Manuela Waldner, Paskal Plank, Veronika Solteszova, and Ivan Viola. The accuracy of gauge-figure tasks in monoscopic and stereo displays. IEEE Computer Graphics and Applications, 36(4):56–66, July/ August 2016. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2016/04/mcg2016040056-
abs.html.

Bayat:2024:WVA

[BWR24] Hannah Clara Bayat, Manuela Waldner, and Renata G. Raidou. A workflow to visually assess interobserver variability in medical image segmentation. IEEE Computer Graphics and Applications, 44(1):86–94, January/ February 2024. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Berger:2012:MVP

[BWWM12] Kai Berger, Andrea Weidlich, Alexander Wilkie, and Marcus Magnor. Modeling and verifying the polarizing reflectance of real-world metallic surfaces. IEEE Computer Graphics and Applications, 32(2):24–33, March/April 2012. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Beyer:2021:VDS

[BYP21] Johanna Beyer, Yalong Yang, and Hanspeter Pfister. Visualization design sprints for online and oncampus courses. IEEE Computer Graphics and Applications, 41(6):37–47, November/December 2021. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Cabezas:1989:SB

[Cab89] Roger Cabezas. From Seoul to Barcelona. IEEE Computer Graphics and Applications, 9(4):6–9, July/ August 1989. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Conti:2006:CSI

[CAG^{+}06] Gregory Conti, Kulsoom Abdullah, Julian Grizzard, John Stasko, John A. Copeland, Mustaque Ahamad, Henry L. Owen, and Chris Lee. Countering security information overload through alert and packet visualization. IEEE Computer Graphics and Applications, 26(2):60–70, March/ April 2006. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Cairo:2020:IAG

[Cai20] A. Cairo. If anything on this graphic causes confusion, discard the entire product. IEEE Computer Graphics and Applications, 40(2): 91–97, March/April 2020. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Cameron:1991:EBC

[Cam91] Stephen Cameron. Efficient bounds in constructive solid geometry. IEEE Computer Graphics and Applications, 11(3):68–74, May/ June 1991. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Campbell:2016:IVS

[Cam16] Bruce D. Campbell. Immersive visualization to support scientific insight. IEEE Computer Graphics and Applications, 36(3):17–21, May/June 2016. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2016/03/mcg2016030017-
abs.html.

Campen:2017:TBQ

[Cam17] Marcel Campen. Tiling the bunny: Quad layouts for efficient 3D geometry representation. IEEE Computer Graphics and Applications, 37(3):88–95, May/ June 2017. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        cg/2017/03/mcg2017030088-
                                        abs.html.

Cartwright:2005:WBS

[CAP^{+}05] Richard Cartwright, Valery Adzhiev, Alexander A. Pasko, Yuichiro Goto, and Tosiyasu L. Kunii. Webbased shape modeling with Hyperfun. IEEE Computer Graphics and Applications, 25(2):60–69, March/ April 2005. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Carson:1983:SCG

[Car83] George S. Carson. The specification of computer graphics systems. IEEE Computer Graphics and Applications, 3(6):27–30, 32–36, 38–41, September 1983. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Carlbom:1987:AGS

[Car87] Ingrid B. Carlbom. An algorithm for geometric set operations using cellular subdivision techniques. IEEE Computer Graphics and Applications, 7(5):44–55, May/ June 1987. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Card:1996:VRI

[Car96] Stuart K. Card. Visualizing retrieved information: a survey. IEEE Computer Graphics and Applications, 16(2): 63–67, March/April 1996. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Card:2016:PTD

[Car16] Stuart K. Card. The PhD thesis deconstructed. IEEE Computer Graphics and Applications, 36(4):92–101, July/August 2016. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2016/04/mcg2016040092.
html.

Casale:1987:FSM

[Cas87] Malcolm S. Casale. Freeform solid modeling with trimmed surface patches. IEEE Computer Graphics and Applications, 7(1):33– 43, January/February 1987. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Catarci:1996:ID

[Cat96] Tiziana Catarci. Interaction with databases. IEEE Computer Graphics and Applications, 16(2):67–69, March/ April 1996. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Cavin:1987:PCP

[Cav87] Doyle K. Cavin. Peripheral considerations for PC-CAD users. IEEE Computer Graphics and Applications, 7(10):32–38, October 1987. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Cavallero:1997:FHP

[Cav97] Rick Cavallero. The Fox-Trax hockey puck tracking system. IEEE Computer Graphics and Applications, 17(2):6–12, March/ April 1997. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Carinalli:1986:NAG

[CB86] Charles Carinalli and John Blair. National’s advanced graphics chip set for highperformance graphics. IEEE Computer Graphics and Applications, 6(10):40–48, October 1986. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Cho:2014:VDP

[CBB^{+}14] Sungmin Cho, Dongyoub Baek, Seung-Yeob Baek, Kunwoo Lee, and Hyunwoo Bang. 3D volume drawing on a potter’s wheel. IEEE Computer Graphics and Applications, 34(3):50– 58, May/June 2014. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Capece:2020:TSS

[CBC^{+}20] N. Capece, F. Banterle, P. Cignoni, F. Ganovelli, and U. Erra. Turning a smartphone selfie into a studio portrait. IEEE Computer Graphics and Applications, 40(1):140–147, January/February 2020. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Calvert:1993:DAM

[CBD^{+}93] Tom Calvert, Armin Bruderlin, John Dill, Thecla Schiphorst, and Chris Welman. Desktop animation of multiple human figures. IEEE Computer Graphics and Applications, 13(3):18– 26, May/June 1993. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Chapman:2010:WAL

[CBD10] Paul Chapman, Kim Bale, and Pierre Drap. We all live in a virtual submarine. IEEE Computer Graphics and Applications, 30(1):85– 89, January/February 2010. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Childs:2019:SVC

[CBGH19] H. Childs, J. Bennett, C. Garth, and B. Hentschel. In situ visualization for computational science. IEEE Computer Graphics and Applications, 39(6):76–85, November/December 2019. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Ceccarini:2022:VIS

[CBS^{+}22] Chiara Ceccarini, Edyta Paulina Bogucka, Indira Sen, Marios Constantinides, Catia Prandi, and Daniele Quercia. Visualizing internal sustainability efforts in big companies. IEEE Computer Graphics and Applications, 42(3):87–98, May/ June 2022. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Chang:2008:LST

[CBZ^{+}08] Remco Chang, Thomas Butkiewicz, Caroline Ziemkiewicz, Zachary Wartell, Nancy Pollard, and William Ribarsky. Legible simplification of textured urban models. IEEE Computer Graphics and Applications, 28 (3):27–36, May/June 2008. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Chandler:2020:MVA

[CC20] T. Chandler and A. Clulow.

Modeling Virtual Angkor: An evolutionary approach to a single urban space. IEEE Computer Graphics and Applications, 40(3):9–16, May/ June 2020. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Camba:2022:IDC

[CCB22] Jorge D. Camba, Pedro Company, and Vetria L. Byrd. Identifying deception as a critical component of visualization literacy. IEEE Computer Graphics and Applications, 42(1): 116–122, January/February 2022. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Chourasia:2007:VIH

[CCC^{+}07] Amit Chourasia, Steve Cutchin, Yifeng Cui, Reagan W. Moore, Kim Olsen, Steven M. Day, J. Bernard Minster, Philip Maechling, and Thomas H. Jordan. Visual insights into highresolution earthquake simulations. IEEE Computer Graphics and Applications, 27(5):28–34, September/ October 2007. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL http://csdl.
                                        computer.org/comp/mags/
                                        cg/2007/extras/g5028x1.
                                        mov.

Chiang:2018:PVP

[CCC18] Pei-Ying Chiang, Han-Yu Chang, and Yung-Ju Chang. PotteryGo: A virtual pottery making training system. IEEE Computer Graphics and Applications, 38(2):74–88, March/ April 2018. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        cg/2018/02/mcg2018020074-
                                        abs.html.

Carpendale:2014:OBD

[CCE^{+}14] Sheelagh Carpendale, Min Chen, Daniel Evanko, Nils Gehlenborg, Carsten Gorg, Larry Hunter, Francis Rowland, Margaret-Anne Storey, and Hendrik Strobelt. Ontologies in biological data visualization. IEEE Computer Graphics and Applications, 34(2):8–15, March/ April 2014. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Carpendale:1997:EDV

[CCF97] M. Sheelagh T. Carpendale, David J. Cowperthwaite, and F. David Fracchia. Extending distortion viewing from 2D to 3D. IEEE Computer Graphics and Applications, 17(4):42–51, July/ August 1997. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Callieri:2004:VDP

[CCG^{+}04] Marco Callieri, Paolo Cignoni, Fabio Ganovelli, Gaetano Impoco, Claudio Montani, Paolo Pingi, Federico Ponchio, and Roberto Scopigno. Visualization and 3D data processing in the David restoration. IEEE Computer Graphics and Applications, 24(2):16–21, March/ April 2004. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http:/

/csdl.computer.org/dl/
mags/cg/2004/02/g2016.
htm;   http://csdl.computer.
org/dl/mags/cg/2004/02/
g2016.pdf.

Cruz:2015:CAE

[CCM15] Pedro Cruz, Antonio Cruz, and Penousal Machado. Contiguous animated edgebased cartograms for traffic visualization. IEEE Computer Graphics and Applications, 35(5):76–83, September/October 2015. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://www.

                                        computer.org/csdl/mags/
                                        cg/2015/05/mcg2015050076-
                                        abs.html.

Calvert:1982:AKS

[CCP82] T. W. Calvert, J. Chapman, and A. Patla. Aspects of the kinematic simulation of human movement. IEEE Computer Graphics and Applications, 2(9):41–50, November/December 1982. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Chinchor:2010:GEI

[CCR10] Nancy A. Chinchor, Michael G. Christel, and William Ribarsky. Guest Editors’ introduction: Multimedia analytics. IEEE Computer Graphics and Applications, 30(5): 18–19, September/October 2010. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Chen:2020:CPI

[CCT^{+}20] S. S. Chen, H. Cui, P. Tan, X. Sun, Y. Ji, and H. Duh. Cantonese porcelain image generation using user-guided generative adversarial networks. IEEE Computer Graphics and Applications, 40(5):100–107, September/ October 2020. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Carlbom:1985:HDS

[CCV85] Ingrid Carlbom, Indranil Cakravarty, and David Vanderschel. A hierarchical data structure for representing the spatial decomposition of 3-D objects. IEEE Computer Graphics and Applications, 5(4):24–31, April 1985. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Contero:2002:PDQ

[CCVA02] Manuel Contero, Pedro Company, Carlos Vila, and Nuria Aleixos. Product data quality and collaborative engineering. IEEE Computer Graphics and Applications, 22(3):32–42, May/ June 2002. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

computer.org/cga/cg2002/
g3032abs.htm;   http:/
/dlib.computer.org/cg/
books/cg2002/pdf/g3032.
pdf.

Cai:2012:NDR

[CD12] Liang Cai and Feipeng Da.

Nonrigid-deformation recovery for 3D face recognition using multiscale registration. IEEE Computer Graphics and Applications, 32(3):37–45, May/ June 2012. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Cook:1983:TDS

[CDBL83] Larry T. Cook, Samuel J. Dwyer III, Solomon Batnitzky, and Kyo Rak Lee. A three-dimensional display system for diagnostic imaging applications. IEEE Computer Graphics and Applications, 3(5):13–19, August 1983. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Chen:1997:RFS

[CdHM97] Jim X. Chen, Niels da Vitoria Lobo, Charles E. Hughes, and J. Michael Moshell. Real-time fluid simulation in a dynamic virtual environment. IEEE Computer Graphics and Applications, 17(3):52–61, May/ June 1997. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Cantoni:2019:TCG

[CDLS19] Virginio Cantoni, Piercarlo Dondi, Luca Lombardi, and Alessandra Setti. Teaching computer graphics through a digital humanities project. IEEE Computer Graphics and Applications, 39(2):89– 94, March/April 2019. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL https:

                                        //ieeexplore.ieee.org/
                                        document/8672984/.

Chuang:1983:SCA

[CE83] Richard Chuang and Glenn Entis. 3-D shaded computer animation — step by step. IEEE Computer Graphics and Applications, 3(9):18– 25, December 1983. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Chuah:1998:IRG

[CE98] Mei C. Chuah and Stephen G.

Eick. Information rich glyphs for software management data. IEEE Computer Graphics and Applications, 18(4):24–29, July/ August 1998. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

computer.org/cga/cg1998/
g4024abs.htm.

Chen:2009:VVD

[CEH^{+}09] Min Chen, David Ebert, Hans Hagen, Robert S. Laramee, Robert van Liere, Kwan-Liu Ma, William Ribarsky, Gerik Scheuermann, and Deborah Silver. Visualization viewpoints: Data, information, and knowledge in visualization. IEEE Computer Graphics and Applications, 29(1):12–19, January/ February 2009. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Capece:2022:BAS

[CEMD22] Nicola Capece, Ugo Erra, Gilda Manfredi, and Rocco Di Bello. BoidVR: an agent simulation environment based on freehand and virtual reality. IEEE Computer Graphics and Applications, 42(6):107–115, November/December 2022. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Cavazza:1998:SMC

[CEMTT98] Marc Cavazza, Rae Earnshaw, Nadia Magnenat-Thalmann, and Daniel Thalmann. Survey: Motion control of virtual humans. IEEE Computer Graphics and Applications, 18(5):24–31, September/ October 1998. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

computer.org/cga/cg1998/
g5024abs.htm;   http:/
/dlib.computer.org/cg/
books/cg1998/pdf/g5024.
pdf.

Cover:1993:IDM

[CEO^{+}93] Steven Cover, Norberto Ezquerra, James F. O’Brien, Richard Rowe, Thomas Gadacz, and Ellen Palm. Interactively deformable models for surgery simulation. IEEE Computer Graphics and Applications, 13(6): 68–75, November/December 1993. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Cook:2007:GEI

[CES07] Kris Cook, Rae Earnshaw, and John Stasko. Guest Editors’ introduction: Discovering the unexpected. IEEE Computer Graphics and Applications, 27(5):15–19, September/October 2007. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cg/2007/05/g5015.pdf.

Caponetti:1993:CSB

[CF93] Laura Caponetti and Anna Maria Fanelli. Computer-aided simulation for bone surgery. IEEE Computer Graphics and Applications, 13(6): 86–92, November/December 1993. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Campbell:2016:UAV

[CF16] Bruce D. Campbell and Samsel Francesca. Using art to visualize cellular environments: An interview with David Goodsell. IEEE Computer Graphics and Applications, 36(6):6–10, November/December 2016. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL https:/

/www.computer.org/csdl/

                                        mags/cg/2016/06/mcg2016060006-
                                        abs.html.

Cai:2001:WBD

[CFF01] Weidong Cai, Dagan Feng, and Roger Fulton. Webbased digital medical images. IEEE Computer Graphics and Applications, 21(1):44–47, January/February 2001. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL

                                        http://computer.org/cga/
                                        cg2001/g1044abs.htm;
                                        http://dlib.computer.org/
                                        cg/books/cg2001/pdf/g1044.
                                        pdf.

Coquillart:1984:SDD

[CG84] Sabine Coquillart and Michel Gangnet. Shaded display of digital maps. IEEE Computer Graphics and Applications, 4(7):35–42, July/ August 1984. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Clemons:1985:SSA

[CG85] Eric K. Clemons and Arnold J. Greenfield. The SAGE system architecture: a system for the rapid development of graphics interfaces for decision support. IEEE Computer Graphics and Applications, 5(11): 38–50, November/December 1985. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Cohen:1994:VBP

[CG94] D. Cohen and C. Gotsman. Visualization blackboard: Photorealistic terrain imaging and flight simulation. IEEE Computer Graphics and Applications, 14(2):10–12, March/April 1994. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Cohen:1986:ERA

[CGIB86] Michael F. Cohen, Donald P. Greenberg, David S. Immel, and Philip J. Brock. An efficient radiosity approach for realistic image synthesis. IEEE Computer Graphics and Applications, 6(3):26–35, March/ April 1986. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Chen:2017:PTA

[CGJ^{+}17] Min Chen, Georges Grinstein, Chris R. Johnson, Jessie Kennedy, and Melanie Tory. Pathways for theoretical advances in visualization. IEEE Computer Graphics and Applications, 37(4):103–112, July/ August 2017. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2017/04/mcg2017040103-
abs.html.

Carrozzino:2020:DVH

[CGMB20] M. A. Carrozzino, R. Galdieri, O. M. Machidon, and M. Bergamasco. Do virtual humans dream of digital sheep? IEEE Computer Graphics and Applications, 40(4):71–83, July/ August 2020. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Carrozzino:2023:VRT

[CGT^{+}23] Marcello A. Carrozzino, Giovanni Giuliodori, Camilla Tanca, Chiara Evangelista, and Massimo Bergamasco. Virtual reality training for post-earthquake rescue operators. IEEE Computer Graphics and Applications, 43(3):61–70, May/ June 2023. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Caporal:1981:TAS

[CH81] P. M. Caporal and G. J. Hahn. Tools for automated statistical graphics. IEEE Computer Graphics and Applications, 1(10):72–82, October 1981. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Chuang:1995:NSS

[CH95] Jung-Hong Chuang and Weun-Jier Hwang. A new space subdivision for ray tracing CSG solids. IEEE Computer Graphics and Applications, 15(6):56–62, November/December 1995. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Chen:2010:GEI

[CH10] Min Chen and Hans Hagen.

Guest Editors’ introduction: Knowledge-assisted visualization. IEEE Computer Graphics and Applications, 30(1):15–16, January/ February 2010. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Chng:2011:RPP

[Ch’11] Eugene Ch’ng. Realistic placement of plants for virtual environments. IEEE Computer Graphics and Applications, 31(4):66–77, July/August 2011. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Chang:1985:MGA

[Cha85] Keith Y. Chang. Microcomputer graphics and applications with NAPLPS Videotex. IEEE Computer Graphics and Applications, 5(6): 21–33, June 1985. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Chandler:1988:TAI

[Cha88] Richard E. Chandler. A tracking algorithm for implicitly defined curves. IEEE Computer Graphics and Applications, 8(2):83–89, March/April 1988. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL http://www.
                                        computer.org/csdl/mags/
                                        cg/02/00000506-abs.html;
                                         http://www.computer.org/
                                        csdl/mags/cg/02/mcg1988020083-
                                        abs.html.

Charles:1997:ACR

[Cha97a]  John Charles.  About

the cover: Rendering on a fantasy. IEEE Computer Graphics and Applications, 17(5):4–5, September/October 1997. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL http://dlib.
                                        computer.org/cg/books/
                                        cg1997/pdf/g5004.pdf.

Charles:1997:ACV

[Cha97b] John Charles. About the cover: Virtually there: Taking VR to work. IEEE Computer Graphics and Applications, 17(6):4–5, November/ December 1997. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL http://dlib.
                                        computer.org/cg/books/
                                        cg1997/pdf/g6004.pdf.

Charles:1998:ACL

[Cha98a] John Charles. About the cover: Light shaping digital space. IEEE Computer Graphics and Applications, 18(3):4–5, May/ June 1998. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://dlib.
computer.org/cg/books/
cg1998/pdf/g3004.pdf.

Charles:1998:ACA

[Cha98b]  John Charles.  About

the cover: The art of chance. IEEE Computer Graphics and Applications, 18(5):4–5, September/October 1998. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://dlib.
computer.org/cg/books/
cg1998/pdf/g5004.pdf.

Charles:1999:ACF

[Cha99] John Charles. About the cover: Finding beauty in mathematics. IEEE Computer Graphics and Applications, 19(1):4–5, January/ February 1999. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://dlib.
computer.org/cg/books/
cg1999/pdf/g1004.pdf.

Chabot:2009:GSD

[Cha09] Christian Chabot. Graphically speaking: Demystifying visual analytics. IEEE Computer Graphics and Applications, 29(2):84–87, March/April 2009. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Cress:1997:IVD

[CHC^{+}97] Jeffrey D. Cress, Lawrence J. Hettinger, James A. Cunningham, Gary E. Riccio, Michael W. Haas, and Grant R. McMillan. Integrating vestibular displays for VEs and airborne applications. IEEE Computer Graphics and Applications, 17(6):46–52, November/December 1997. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

computer.org/cga/cg1997/
g6046abs.htm.

Chen:2005:TUI

[Che05] Chaomei Chen. Top 10 unsolved information visualization problems. IEEE Computer Graphics and Applications, 25(4):12–16, July/ August 2005. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Chen:2008:ITV

[Che08] Chaomei Chen. An informationtheoretic view of visual analytics. IEEE Computer Graphics and Applications, 28(1):18–23, January/ February 2008. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Campbell:2024:AAI

[CHH24] Bruce Donald Campbell, Nicholas Hedley, and Aaron Hertzmann. Art and artificial intelligence. IEEE Computer Graphics and Applications, 44(2):10–11, March/ April 2024. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Chiyokura:1987:ERO

[Chi87] Hiroaki Chiyokura. An extended rounding operation for modeling solids with free-form surfaces. IEEE Computer Graphics and Applications, 7(12):27–36, December 1987. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Chou:1995:VDP

[Cho95] Jin J. Chou. Voronoi diagrams for planar shapes. IEEE Computer Graphics and Applications, 15(2): 52–59, March/April 1995. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Carlson:2021:CGA

[CHP21] Wayne Carlson, Ronald Hackathorn, and Richard Parent. Computer graphics and animation at The Ohio State University. IEEE Computer Graphics and Applications, 41(3):8–17, May/ June 2021. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Christensen:1983:UGR

[Chr83] Albert H. J. Christensen. Uniform grids from raster data (magnetic anomalies). IEEE Computer Graphics and Applications, 3(6):44– 53, September 1983. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Chen:1985:SSC

[CHRU85] Lih-Shyang Chen, Gabor T. Herman, R. Anthony Reynolds, and Jayaram K. Udupa. Surface shading in the Cuberille environment. IEEE Computer Graphics and Applications, 5(12):33– 43, December 1985. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Campbell:2021:VCC

[CHW21] B. D. Campbell, N. Hedley, and C. Weaver. Visualization of climate change. IEEE Computer Graphics and Applications, 41(1):15– 16, January/February 2021. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Cook:1999:VAN

[CHY99] Norman D. Cook, Takefumi Hayashi, and Nobuaki Yoshida. Visualizing the atomic nucleus. IEEE Computer Graphics and Applications, 19(5):54–60, September/October 1999. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

computer.org/cga/cg1999/
g5054abs.htm;   http:/
/dlib.computer.org/cg/
books/cg1999/pdf/g5054.
pdf.

Chenney:1999:DMC

[CIF99] Stephen Chenney, Jeffrey Ichnowski, and David Forsyth. Dynamics modeling and culling. IEEE Computer Graphics and Applications, 19(2):79–87, March/ April 1999. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

computer.org/cga/cg1999/
g2079abs.htm;   http:/
/dlib.computer.org/cg/
books/cg1999/pdf/g2079.
pdf.

Chaudhary:2019:CPU

[CJS^{+}19] Aashish Chaudhary, Sankhesh J.

Jhaveri, Alvaro Sanchez, Lisa S. Avila, Kenneth M. Martin, Allison Vacanti, Marcus D. Hanwell, and Will Schroeder. Crossplatform ubiquitous volume rendering using programmable shaders in VTK for scientific and medical visualization. IEEE Computer Graphics and Applications, 39(1):26–43, January/February 2019. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL https:

                                        //ieeexplore.ieee.org/
                                        document/8663650/.

Campbell:1985:PWR

[CK85a] Roy H. Campbell and William J. Kubitz. Professional workstation research project. IEEE Computer Graphics and Applications, 6(5):17–24, May/June 1985. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Chiyokura:1985:MRS

[CK85b] Hiroaki Chiyokura and Fumihiko Kimura. A method of representing the solid design process. IEEE Computer Graphics and Applications, 5(4):32–41, April 1985. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Crouser:2012:TVT

[CKJC12] R. Jordan Crouser, Daniel E. Kee, Dong Hyun Jeong, and Remco Chang. Two visualization tools for analyzing agent-based simulations in political science. IEEE Computer Graphics and Applications, 32(1):67– 77, January/February 2012. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Cremer:1996:PVD

[CKP96] J. Cremer, J. Kearney, and Y. Papelis. Projects in VR: Driving simulation: Challenges for VR technology. IEEE Computer Graphics and Applications, 16(5):16– 20, September 1996. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Cho:2016:ULF

[CKPB16] Sungmin Cho, Won-Seok Kim, Nam-Jong Paik, and Hyunwoo Bang. Upperlimb function assessment using VBBTs for stroke patients. IEEE Computer Graphics and Applications, 36(1):70–78, January/ February 2016. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
cg/2016/01/mcg2016010070-
abs.html.

Choo:2018:VAE

[CL18] Jaegul Choo and Shixia Liu.

Visual analytics for explainable deep learning. IEEE Computer Graphics and Applications, 38(4):84–92, July/August 2018. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2018/04/mcg2018040084-
abs.html.

Clark:1991:DM

[Cla91] David R. Clark. The demise of multimedia. IEEE Computer Graphics and Applications, 11(4):75–80, July/ August 1991. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Chen:1996:UHC

[CLC^{+}96] Lih-Shyang Chen, Pei-Wen Liu, Ku-Yaw Chang, Jong-Ping Chen, Su-Chou Chen, Hong-Chow Hong, and Jain Liu. Using hypermedia in computer-aided instruction. IEEE Computer Graphics and Applications, 16(3):52– 57, May/June 1996. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Cao:2016:EVS

[CLDW16] Nan Cao, Yu-Ru Lin, Fan Du, and Dashun Wang. Episogram: Visual summarization of egocentric social interactions. IEEE Computer Graphics and Applications, 36(5):72–81, September/October 2016. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2016/05/mcg2016050072-
abs.html.

Cao:2016:CBM

[CLG^{+}16] Mingliang Cao, Yi Li, Yueping Guo, Lei Yao, and Zhigeng Pan. Customized body mapping to facilitate the ergonomic design of sportswear. IEEE Computer Graphics and Applications, 36(6):70–77, November/ December 2016. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2016/06/mcg2016060070-
abs.html.

Choe:2015:CVI

[CLms15] Eun Kyoung Choe, Bongshin Lee, and m.c. schraefel. Characterizing visualization insights from quantified selfers’ personal data presentations. IEEE Computer Graphics and Applications, 35(4):28–37, July/ August 2015. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
cg/2015/04/mcg2015040028-
abs.html.

Cao:2014:CEU

[CLP^{+}14] Mingliang Cao, Yi Li, Zhigeng Pan, Josephine Csete, Shu Sun, Jie Li, and Yu Liu. Creative educational use of virtual reality: Working with second life. IEEE Computer Graphics and Applications, 34(5):83–87, September/October 2014. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL http://csdl.
                                        computer.org/csdl/mags/
                                        cg/2014/05/mcg2014050083-
                                        abs.html.

Cao:2015:EVW

[CLP^{+}15] Mingliang Cao, Yi Li, Zhigeng Pan, Josephine Csete, Shu Sun, Jie Li, and Yu Liu. Educational virtual-wear trial: More than a virtual try-on experience. IEEE Computer Graphics and Applications, 35(6):83–89, November/ December 2015. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL http://www.
                                        computer.org/csdl/mags/
                                        cg/2015/06/mcg2015060083-
                                        abs.html.

Cao:2018:TVA

[CLP18] Mingliang Cao, Yi Li, and Zhigeng Pan. Toward visual avatars that dress you well and impact your health. IEEE Computer Graphics and Applications, 38(2):22–27, March/ April 2018. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        cg/2018/02/mcg2018020022-
                                        abs.html.

Carson:1986:RMC

[CM86] George S. Carson and Eileen McGinnis. The reference model for computer graphics. IEEE Computer Graphics and Applications, 6(8): 17–23, August 1986. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Cavendish:1992:FSD

[CM92] James C. Cavendish and Samuel P. Marin. Featurebased surface design and machining. IEEE Computer Graphics and Applications, 12(5):61–??, September 1992. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Cruz:2011:GSI

[CM11] Pedro Cruz and Penousal Machado. Generative storytelling for information visualization. IEEE Computer Graphics and Applications, 31(2):80–85, March/ April 2011. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Cruz:2016:PBV

[CM16] Pedro Cruz and Penousal Machado. Pulsing blood vessels: A figurative approach to traffic visualization. IEEE Computer Graphics and Applications, 36(2):16–21, March/ April 2016. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        cg/2016/02/mcg2016020016-
                                        abs.html.

Crawfis:1994:VFV

[CMB94] Roger Crawfis, Nelson Max, and Barry Becker. Vector field visualization. IEEE Computer Graphics and Applications, 14(5):50–56, September 1994. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Capps:2000:PVN

[CMBZ00] Michael Capps, Don Mc-Gregor, Don Brutzman, and Michale Zyda. Projects in VR: NPSNET-V: a new beginning for dynamically extensible virtual environments. IEEE Computer Graphics and Applications, 20(5):12–15, September/ October 2000. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL extras/

                                        g5012x1.pdf;   http://
                                        dlib.computer.org/cg/books/
                                        cg2000/pdf/g5012.pdf.

Cao:2023:FPA

[CML^{+}23] Jacky Cao, Hassan Mehmood, Xiaoli Liu, Sasu Tarkoma, Ekaterina Gilman, and Xiang Su. Fighting pandemics with augmented reality and smart sensing-based social distancing. IEEE Computer Graphics and Applications, 43(1):65–75, January/ February 2023. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Cohen:1999:PVM

[CMO^{+}99] Philip Cohen, David McGee, Sharon Oviatt, Lizhong Wu, Joshua Clow, Robert King, Simon Julier, and Lawrence Rosenblum. Projects in VR: Multimodal interaction for 2D and 3D environments. IEEE Computer Graphics and Applications, 19(4):10–13, July/ August 1999. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://dlib.
computer.org/cg/books/
cg1999/pdf/g4010.pdf.

Chandru:1995:VML

[CMP95] Vijay Chandru, Swami Manohar, and C. Edmond Prakash. Voxel-based modeling for layered manufacturing. IEEE Computer Graphics and Applications, 15(6):42–47, November/December 1995. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Corker-Marin:2018:CMA

[CMPA18] Quentin Corker-Marin, Alexander Pasko, and Valery Adzhiev. 4D cubism: Modeling, animation, and fabrication of artistic shapes. IEEE Computer Graphics and Applications, 38(3):131– 139, May/June 2018. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL https:/

                                        /www.computer.org/csdl/
                                        mags/cg/2018/03/mcg2018030131-
                                        abs.html.

Carvalho:2008:PCH

[CMSEM08] Elizabeth Carvalho, Adérito Marcos, Maribel Yasmina Santos, and João Espregueira-Mendes. A prototype for cartographic human body analysis. IEEE Computer Graphics and Applications, 28(5):16–21, September/ October 2008. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Cai:2010:CIG

[CMT^{+}10] Yundong Cai, Chunyan Miao, Ah-Hwee Tan, Zhiqi Shen, and Boyang Li. Creating an immersive game world with evolutionary fuzzy cognitive maps. IEEE Computer Graphics and Applications, 30(2):58–70, March/April 2010. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Capps:2001:FED

[CMZ01] Michael Capps, Perry Mc-Dowell, and Michael Zyda. A future for entertainmentdefense research collaboration. IEEE Computer Graphics and Applications, 21(1):37–43, January/ February 2001. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

computer.org/cga/cg2001/
g1037abs.htm;   http:/
/dlib.computer.org/cg/
books/cg2001/pdf/g1037.
pdf.

Cruz-Neira:2003:CHN

[CN03] Carolina Cruz-Neira. Computational humanities: The new challenge for VR. IEEE Computer Graphics and Applications, 23(3):10–13, May/June 2003. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://dlib.
computer.org/cg/books/
cg2003/pdf/g3010.pdf.

Contero:2005:IVS

[CNC^{+}05] Manuel Contero, Ferran Naya, Pedro Company, Jose Luis Saorin, and Julian Conesa. Improving visualization skills in engineering education. IEEE Computer Graphics and Applications, 25(5):24–31, September/October 2005. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Cruz-Neira:2000:VVR

[CNL00] Carolina Cruz-Neira and Gary Lindahl. A voyage into virtual reality: Networking our VR lab to Iowa middle schools and high schools. IEEE Computer Graphics and Applications, 20(2):16–19, March/ April 2000. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL http://dlib.
                                        computer.org/cg/books/
                                        cg2000/pdf/g2016.pdf.

Capin:1997:VHR

[CNT^{+}97] Tolga K. Capin, Hansrudi Noser, Daniel Thalmann, Igor Sunday Pandzic, and Nadia Magnenat Thalmann. Virtual human representation and communication in VLNet. IEEE Computer Graphics and Applications, 17(2):42–53, March/ April 1997. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Cohen:1983:SMT

[Coh83] Elaine Cohen. Some mathematical tools for a modeler’s workbench. IEEE Computer Graphics and Applications, 3(7):63–66, October 1983. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Cohen:2000:IA

[Coh00] Michael Cohen. Imagination amplification. IEEE Computer Graphics and Applications, 20(1):54–55, January/ February 2000. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://dlib.
computer.org/cg/books/
cg2000/pdf/g1054.pdf.

Cohen-Or:1997:LVC

[COK97] Daniel Cohen-Or and Arie Kaufman. 3D line voxelization and connectivity control. IEEE Computer Graphics and Applications, 17(6):80–87, November/December 1997. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

computer.org/cga/cg1997/
g6080abs.htm.

Covaci:2015:VPF

[COM15] Alexandra Covaci, Anne-Helene Olivier, and Franck Multon. Visual perspective and feedback guidance for VR free-throw training. IEEE Computer Graphics and Applications, 35(5):55–65, September/ October 2015. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
cg/2015/05/mcg2015050055-
abs.html.

Cook:2002:GEI

[Coo02a] Perry R. Cook. Guest Editor’s introduction: Virtual worlds, real sounds. IEEE Computer Graphics and Applications, 22(4):22, July/ August 2002. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

                                        computer.org/cga/cg2002/
                                        g4022abs.htm;   http:/
                                        /dlib.computer.org/cg/
                                        books/cg2002/pdf/g4022.
                                        pdf.

Cook:2002:SPM

[Coo02b] Perry R. Cook. Sound production and modeling. IEEE Computer Graphics and Applications, 22(4):23– 27, July/August 2002. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

                                        computer.org/cga/cg2002/
                                        g4023abs.htm;   http:/
                                        /dlib.computer.org/cg/
                                        books/cg2002/pdf/g4023.
                                        pdf.

Coquillart:1987:CPB

[Coq87]  Sabine Coquillart.  A

control-point-based sweeping technique. IEEE Computer Graphics and Applications, 7(11):36–44, November/December 1987. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Chen:1989:SAA

[CP89] Shenchang Eric Chen and Richard E. Parent. Shape averaging and its applications to industrial design.

IEEE Computer Graphics and Applications, 9(1):47– 54, January/February 1989. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Chou:1992:DRU

[CP92] Jin J. Chou and Les A.

Piegl. Data reduction using cubic rational B-splines. IEEE Computer Graphics and Applications, 12(3):60– 68, May/June 1992. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Choo:2013:CCM

[CP13] Jaegul Choo and Haesun Park. Customizing computational methods for visual analytics with big data. IEEE Computer Graphics and Applications, 33(4): 22–28, July/August 2013. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Childs:2010:ESP

[CPA^{+}10] Hank Childs, David Pugmire, Sean Ahern, Brad Whitlock, Mark Howison, Prabhat, Gunther H. Weber, and E. Wes Bethel. Extreme scaling of production visualization software on diverse architectures. IEEE Computer Graphics and Applications, 30(3):22–31, May/ June 2010. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Capin:2008:SAM

[CPAM08] Tolga Capin, Kari Pulli, and Tomas Akenine-Möller. The state of the art in mobile graphics research. IEEE Computer Graphics and Applications, 28(4): 74–84, July/August 2008. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Cui:2017:ASS

[CPAV^{+}17] Jian Cui, Voicu Popescu, Nicoletta Adamo-Villani, Susan Wagner Cook, Katherine A. Duggan, and Howard S. Friedman. Animation stimuli system for research on instructor gestures in education. IEEE Computer Graphics and Applications, 37(4):72–83, July/ August 2017. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        cg/2017/04/mcg2017040072-
                                        abs.html.

Carson:1999:DVI

[CPC99] George S. Carson, Richard F. Puk, and Rikk Carey. Developing the VRML 97 International Standard. IEEE Computer Graphics and Applications, 19(2):52–58, March/April 1999. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

computer.org/cga/cg1999/
g2052abs.htm;   http:/
/dlib.computer.org/cg/
books/cg1999/pdf/g2052.
pdf.

Callieri:2008:VIF

[CPCS08] Marco Callieri, Federico Ponchio, Paolo Cignoni, and Roberto Scopigno. Virtual inspector: a flexible visualizer for dense 3D scanned models. IEEE Computer Graphics and Applications, 28(1):44–54, January/February 2008. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Chung:2016:KAR

[CPG^{+}16] David H. S. Chung, Matthew L.

Parry, Iwan W. Griffiths, Robert S. Laramee, Rhodri Bown, Philip A. Legg, and Min Chen. Knowledgeassisted ranking: A visual analytic application for sports event data. IEEE Computer Graphics and Applications, 36(3):72–82, May/June 2016. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2016/03/mcg2016030072-
abs.html.

Colbert:2006:BSC

[CPK06] Mark Colbert, Sumanta Pattanaik, and Jaroslav Krivánek. BRDF-Shop: Creating physically correct bidirectional reflectance distribution functions. IEEE Computer Graphics and Applications, 26(1):30–36, January/February 2006. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

csdl.computer.org/comp/
mags/cg/2006/extras/g1030x1.
mov;   http://csdl.computer.
org/comp/mags/cg/2006/
extras/g1030x2.mov.

Cashman:2018:RVL

[CPM^{+}18] Dylan Cashman, Geneviève Patterson, Abigail Mosca, Nathan Watts, Shannon Robinson, and Remco Chang. RNNbow: Visualizing learning via backpropagation gradients in RNNs. IEEE Computer Graphics and Applications, 38(6):39–50, ???? 2018. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL https:

//ieeexplore.ieee.org/
document/8617747/.

Cheong:2009:PBC

[CPN09] Joono Cheong, Jongwoo Park, and Silviu-Iulian Niculescu. Physically based collaborative simulations under ring-like network configurations. IEEE Computer Graphics and Applications, 29(4):64–80, July/ August 2009. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Crocker:1991:ENB

[CR91] Gary A. Crocker and William F. Reinke. An editable nonmanifold boundary representation. IEEE Computer Graphics and Applications, 11(2):39–51, March/April 1991. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Casey:1994:SFS

[CR94] Stephen D. Casey and Nicholas F. Reingold. Selfsimilar fractal sets: Theory and procedure. IEEE Computer Graphics and Applications, 14(3):73–82, May/ June 1994. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Chambers:1995:VBV

[CR95] P. Chambers and A. Rockwood. Visualization blackboard: Visualization of solid reaction-diffusion systems. IEEE Computer Graphics and Applications, 15(5):7– 11, September 1995. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Chalmers:2002:GEI

[CR02a] Alan Chalmers and Holly Rushmeier. Guest Editor’s introduction: Computer graphics in art history and archaeology. IEEE Computer Graphics and Applications, 22(5):22–23, September/October 2002. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://dlib.
computer.org/cg/books/
cg2002/pdf/g5022.pdf.

Crossno:2002:VD

[CR02b] Patricia Crossno and David Rogers. Visual debugging. IEEE Computer Graphics and Applications, 22(6): 6–10, November/December 2002. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL

http://dlib.computer.org/
cg/books/cg2002/pdf/g6006.
pdf.

Carnegie:2015:RVD

[CR15] Kieran Carnegie and Taehyun Rhee. Reducing visual discomfort with HMDs using dynamic depth of field. IEEE Computer Graphics and Applications, 35(5):34–41, September/ October 2015. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
cg/2015/05/mcg2015050034-
abs.html.

Crane:1991:HSA

[Cra91] Gregory Crane. Hypermedia and the study of ancient culture. IEEE Computer Graphics and Applications, 11(4):45–51, July/ August 1991. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Chevalier:2018:ORV

[CRA^{+}18] Fanny Chevalier, Nathalie Henry Riche, Basak Alper, Catherine Plaisant, Jeremy Boy, and Niklas Elmqvist. Observations and reflections on visualization literacy in elementary school. IEEE Computer Graphics and Applications, 38(3):21–29, May/ June 2018. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2018/03/mcg2018030021-
abs.html.

Cera:2002:CEA

[CRB^{+}02] Christopher D. Cera, William C.

Regli, Ilya Braude, Yuri Shapirstein, and Cheryl V. Foster. A collaborative 3D environment for authoring design semantics. IEEE Computer Graphics and Applications, 22(3):43– 55, May/June 2002. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

computer.org/cga/cg2002/
g3043abs.htm;   http:/
/dlib.computer.org/cg/
books/cg2002/pdf/g3043.
pdf.

Corney:2002:CFS

[CRC^{+}02] Jonathan Corney, Heather Rea, Doug Clark, John Pritchard, Michael Breaks, and Roddy MacLeod. Coarse filters for shape matching. IEEE Computer Graphics and Applications, 22(3):65– 74, May/June 2002. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

                                        computer.org/cga/cg2002/
                                        g3065abs.htm;   http:/
                                        /dlib.computer.org/cg/
                                        books/cg2002/pdf/g3065.
                                        pdf.

Crissaff:2018:AEV

[CRD^{+}18] Lhaylla Crissaff, Louisa Wood Ruby, Samantha Deutch, R. Luke DuBois, Jean-Daniel Fekete, Juliana Freire, and Claudio Silva. ARIES: Enabling visual exploration and organization of art image collections. IEEE Computer Graphics and Applications, 38(1):91–108, January/February 2018. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL https:/

                                        /www.computer.org/csdl/
                                        mags/cg/2018/01/mcg2018010091-
                                        abs.html.

Caillou:2021:CWB

[CRF^{+}21] P. Caillou, J. Renault, J. D. Fekete, A. C. Letournel, and M. Sebag. Cartolabe: A Web-based scalable visualization of large document collections. IEEE Computer Graphics and Applications, 41(2):76–88, March/ April 2021. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Crow:1981:CAT

[Cro81] Franklin C. Crow. A comparison of antialiasing techniques. IEEE Computer Graphics and Applications, 1(1):40–43, 46–48, January/ February 1981. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Crocker:1987:SCI

[Cro87a] Gary A. Crocker. Screenarea coherence for interactive scanline display algorithms. IEEE Computer Graphics and Applications, 7(9):10–17, September 1987. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Crow:1987:OT

[Cro87b] Franklin C. Crow. The origins of the teapot. IEEE Computer Graphics and Applications, 7(1):8–19, January/February 1987. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Cronin:1996:RVS

[Cro96] T. Cronin. RF visualization to support airborne collection management. IEEE Computer Graphics and Applications, 16(3):11–13, May/June 1996. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Crossno:2018:CVA

[Cro18] Patricia Crossno. Challenges in visual analysis of ensembles. IEEE Computer Graphics and Applications, 38(2):122–131, March/April 2018. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2018/02/mcg2018020122-
abs.html.

Casale:1985:OAS

[CS85] Malcolm S. Casale and Edward L. Stanton. An overview of analytic solid modeling. IEEE Computer Graphics and Applications, 5(2):45–56, February 1985. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Coleman:2006:CGC

[CS06] Patrick Coleman and Karan Singh. Cords: Geometric curve primitives for modeling contact. IEEE Computer Graphics and Applications, 26(3):72–79, May/ June 2006. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Cordier:2007:FFS

[CS07] Frederic Cordier and Hyewon Seo. Free-form sketching of self-occluding objects. IEEE Computer Graphics and Applications, 27(1):50– 59, January/February 2007. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Campbell:2015:MDT

[CS15a] Bruce D. Campbell and Francesca Samsel. Murmurations: Drawing together art, visualization, and physical phenomena. IEEE Computer Graphics and Applications, 35(4):8–12, July/ August 2015. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
cg/2015/04/mcg2015040008-
abs.html.

Campbell:2015:PVA

[CS15b] Bruce D. Campbell and Francesca Samsel. Pursuing value in art-science collaborations. IEEE Computer Graphics and Applications, 35(1):6–11, January/ February 2015. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://csdl.

computer.org/csdl/mags/
cg/2015/01/mcg2015010006-
abs.html.

Campbell:2015:SVE

[CS15c] Bruce D. Campbell and Francesca Samsel. Slow vis: Extending opportunities for insight and understanding over time. IEEE Computer Graphics and Applications, 35(6):6–10, November/December 2015. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http:/

/www.computer.org/csdl/
mags/cg/2015/06/mcg2015060006-
abs.html.

Campbell:2017:CFI

[CS17a] Bruce D. Campbell and Francesca Samsel. Coming into focus: An interview with Ellen Jantzen. IEEE Computer Graphics and Applications, 37(5):5– 8, September/October 2017. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL

https://www.computer.org/
csdl/mags/cg/2017/05/mcg2017050005-
abs.html.

Chollet:2017:PVA

[CS17b] Mathieu Chollet and Stefan Scherer. Perception of virtual audiences. IEEE Computer Graphics and Applications, 37(4):50–59, July/ August 2017. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2017/04/mcg2017040050-
abs.html.

Campbell:2019:LGP

[CS19] Bruce Campbell and Francesca Samsel. Lance Gharavi: Performance inspired science + technology. IEEE Computer Graphics and Applications, 39(2):8–13, March/April 2019. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL https:

//ieeexplore.ieee.org/
document/8673007/.

Cibulski:2022:RVR

[CSA22] Lena Cibulski, Johanna Schmidt, and Wolfgang Aigner. Reflections on visualization research projects in the manufacturing industry. IEEE Computer Graphics and Applications, 42(2): 21–32, March/April 2022. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Cohen:1986:CBT

[CSI86] Ellis S. Cohen, Edward T.

Smith, and Lee A. Iverson. Constraint-based tiled windows. IEEE Computer Graphics and Applications, 6(5):35–45, May/June 1986. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Cordier:2003:MMT

[CSMT03] Frederic Cordier, Hyewon Seo, and Nadia Magnenat-Thalmann. Made-to-measure technologies for an online clothing store. IEEE Computer Graphics and Applications, 23(1):38–48, January/February 2003. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http:

//dlib.computer.org/cg/
books/cg2003/pdf/g1038.
pdf;   http://www.computer.
org/cga/cg2003/g1038abs.
htm.

Csuri:1991:AA

[Csu91] Charles A. Csuri. Art and animation. IEEE Computer Graphics and Applications, 11(1):30–35, January/ February 1991. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Chapman:2001:VVR

[CSWB01] Paul Chapman, Peter Stevens, Derek Wills, and Graham Brookes. Visualization viewpoints: Real-time visualization in the offshore industry. IEEE Computer Graphics and Applications, 21(4):6– 10, July/August 2001. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http:

//dlib.computer.org/cg/
books/cg2001/pdf/g4006.
pdf.

Choi:2004:TRH

[CT04a] Seungmoon Choi and Hong Z.

Tan. Toward realistic haptic rendering of surface textures. IEEE Computer Graphics and Applications, 24(2):40–47, March/ April 2004. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cg/2004/02/g2040abs.htm;
http://csdl.computer.
org/dl/mags/cg/2004/02/
g2040.htm;   http://csdl.
computer.org/dl/mags/cg/
2004/02/g2040.pdf.

Chu:2004:RTP

[CT04b] Nelson S. H. Chu and Chiew-Lan Tai. Realtime painting with an expressive virtual Chinese brush. IEEE Computer Graphics and Applications, 24(5):76–85, September/ October 2004. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http:

//csdl.computer.org/dl/
mags/cg/2004/05/g5076.
htm;   http://csdl.computer.
org/dl/mags/cg/2004/05/
g5076.pdf.

Chou:2009:ATI

[CT09] Chang-Min Chou and Din-Chang Tseng. Affinetransformation-invariant public fragile watermarking for 3D model authentication.

IEEE Computer Graphics and Applications, 29(2): 72–79, March/April 2009. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Cassol:2017:EOE

[CTJ^{+}17] Vincius J. Cassol, Estevao Smania Testa, Claudio Rosito Jung, Muhammad Usman, Petros Faloutsos, Glen Berseth, Mubbasir Kapadia, Norman I. Badler, and Soraia Raupp Musse. Evaluating and optimizing evacuation plans for crowd egress. IEEE Computer Graphics and Applications, 37(4):60–71, July/ August 2017. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2017/04/mcg2017040060-
abs.html.

Chen:2018:LCL

[CTM18] Yi-Ling Chen, Zhenyu Tang, and Kwan-Liu Ma. Light-Painter: Creating longexposure imagery from videos. IEEE Computer Graphics and Applications, 38(4):27–36, July/ August 2018. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2018/04/mcg2018040027-
abs.html.

Chinchor:2010:MAV

[CTW^{+}10] Nancy A. Chinchor, James J.

Thomas, Pak Chung Wong, Michael G. Christel, and William Ribarsky. Multimedia analysis + visual analytics = multimedia analytics. IEEE Computer Graphics and Applications, 30(5): 52–60, September/October 2010. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Cendes:1987:QIA

[CW87] Z. Cendes and S. Wong. C^{1} quadratic interpolation over arbitrary point sets. IEEE Computer Graphics and Applications, 7(10):8–16, October 1987. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Cendes:1988:CQI

[CW88] Zoltan J. Cendes and Steven H. Wong. C^{1} quadratic interpolation over arbitrary point sets. IEEE Computer Graphics and Applications, 3(1):8–16, January/February 1988. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Clifton:1993:DVD

[CW93] T. E. Clifton III and Fred L.

Wefer. Direct volume display devices. IEEE Computer Graphics and Applications, 13(4):57–65, July/ August 1993. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Clay:1996:PLI

[CW96] Sharon Rose Clay and Jane Wilhelms. Put: Languagebased interactive manipulation of objects. IEEE Computer Graphics and Applications, 16(2):31–39, March/ April 1996. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Cottam:2009:GSS

[CW09] Matthew Cottam and Katie Wray. Graphically speaking: Sketching tangible interfaces: Creating an electronic palette for the design community. IEEE Computer Graphics and Applications, 29(3):90–95, May/ June 2009. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Chen:2002:ASL

[CWB02] Jim Chen, Xusheng Wang, and Jack Bresenham. The analysis and statistics of line distribution. IEEE Computer Graphics and Applications, 22(6):100–107, November/December 2002. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL

http://dlib.computer.org/

cg/books/cg2002/pdf/g6100.
pdf;   http://www.computer.
org/cga/cg2002/g6100abs.
htm.

Chapman:1999:VUE

[CWBS99] Paul Chapman, Derek Wills, Graham Brookes, and Peter Stevens. Visualizing underwater environments using multifrequency sonar. IEEE Computer Graphics and Applications, 19(5):61–65, September/ October 1999. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

computer.org/cga/cg1999/
g5061abs.htm;   http:/
/dlib.computer.org/cg/
books/cg1999/pdf/g5061.
pdf.

Chakrabarti:1990:VRP

[CWGC90] Swapan Chakrabarti, J. C.

Wong, Sivaprasad Gogineni, and Sungho Cho. Visualizing radiation patterns of antennas. IEEE Computer Graphics and Applications, 10(1):41–49, January/ February 1990. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Cui:2010:CPD

[CWL^{+}10] Weiwei Cui, Yingcai Wu, Shixia Liu, Furu Wei, Michelle Zhou, and Huamin Qu. Context-preserving, dynamic word cloud visualization. IEEE Computer Graphics and Applications, 30(6):42–53, November/December 2010. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Calvert:2005:ACD

[CWRF05] Tom Calvert, Lars Wilke, Rhonda Ryman, and Ilene Fox. Applications of computers to dance. IEEE Computer Graphics and Applications, 25(2):6–12, March/ April 2005. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Cao:2021:HVS

[CWWW21] Wei Cao, Shiqian Wu, Dianwei Wang, and Jiaxin Wu. A high visibility and SNR image from one single-shot lowlight image. IEEE Computer Graphics and Applications, 41(5):124–137, September/ October 2021. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Capps:1999:PVC

[CWZ99] Michael Capps, Kent Watsen, and Michale Zyda. Projects in VR: Cyberspace and mock apple pie: a vision of the future of graphics and virtual environments. IEEE Computer Graphics and Applications, 19(6):8–11, November/December 1999. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http:

//dlib.computer.org/cg/
books/cg1999/pdf/g6008.
pdf.

Cun:2019:DAF

[CXPG19] Xiaodong Cun, Feng Xu, Chi-Man Pun, and Hao Gao. Depth-assisted full resolution network for single image-based view synthesis. IEEE Computer Graphics and Applications, 39(2):52– 64, March/April 2019. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL https:

//ieeexplore.ieee.org/
document/8565998/.

Chen:2020:CMO

[CXW^{+}20] S. Chen, H. Xu, G. Weizhe, L. Xuxin, and M. Bofeng. A classification method of oracle materials based on local convolutional neural network framework. IEEE Computer Graphics and Applications, 40(3):32–44, May/June 2020. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Chalmers:2023:VFH

[CZAA23] Alan Chalmers, Danel Zholzhanova, Tarunya Arun, and Ali Asadipour. Virtual flavor: High-fidelity simulation of real flavor experiences. IEEE Computer Graphics and Applications, 43(2):23–31, March/ April 2023. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Chang:2009:VVD

[CZGR09] Remco Chang, Caroline Ziemkiewicz, Tera Marie Green, and William Ribarsky. Visualization viewpoints: Defining insight for visual analytics. IEEE Computer Graphics and Applications, 29(2):14–17, March/ April 2009. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Chen:2010:AGV

[CZP^{+}10] Wenzhi Chen, Mingmin Zhang, Zhigeng Pan, Gengdai Liu, Huaqing Shen, Shengnan Chen, and Yong Liu. Animations, games, and virtual reality for the Jing–Hang Grand Canal. IEEE Computer Graphics and Applications, 30(3):84– 88, May/June 2010. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Chiang:2013:VSI

[CZY^{+}13] Patricia Chiang, Jianmin Zheng, You Yu, Koon Hou Mak, Chee Kong Chui, and Yiyu Cai. A VR simulator for intracardiac intervention. IEEE Computer Graphics and Applications, 33(1):44–57, January/ February 2013. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Dorsey:1995:IDC

[DAG95] Julie Dorsey, James Arvo, and Donald Greenberg. Interactive design of complex time dependent lighting. IEEE Computer Graphics and Applications, 15(2): 26–36, March/April 1995. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Dai:1989:CMA

[Dai89] Fan Dai. Collision-free motion of an articulated kinematic chain in a dynamic environment. IEEE Computer Graphics and Applications, 9(1):70–74, January/ February 1989. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Dalstein:2019:TMV

[Dal19] B. Dalstein. Topological modeling for vector graphics. IEEE Computer Graphics and Applications, 39 (3):86–95, May/June 2019. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Domik:2013:HHS

[DASF13] Gitta Domik, Stephan Arens, Peter Stilow, and Hauke Friedrich. Helping high schoolers move the (virtual) world. IEEE Computer Graphics and Applications, 33(1):70–74, January/ February 2013. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

DeFanti:1996:PVG

[DB96] Tom DeFanti and Maxine Brown. Projects in VR: GII testbed applications on the I-WAY. IEEE Computer Graphics and Applications, 16(4):14–17, July/ August 1996. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

DAmora:2003:PVP

[DB03] Bruce D’Amora and Fausto Bernardini. Pervasive 3D viewing for product data management. IEEE Computer Graphics and Applications, 23(2):14–19, March/ April 2003. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL http://dlib.
                                        computer.org/cg/books/
                                        cg2003/pdf/g2014.pdf.

Davies:2016:EES

[DB16] Bryan Davies and Edwin Blake. Evaluating existing strategies to limit video game playing time. IEEE Computer Graphics and Applications, 36(2):47–57, March/April 2016. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2016/02/mcg2016020047-
abs.html.

Duke:2005:DYS

[DBDH05] David J. Duke, Ken W.

Brodlie, David A. Duce, and Ivan Herman. Do you see what I mean? IEEE Computer Graphics and Applications, 25(3):6–9, May/ June 2005. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Dubnov:2002:SST

[DBJEY^{+}02] Shlomo Dubnov, Ziv Bar-Joseph, Ran El-Yaniv, Dani Lischinski, and Michael Werman. Synthesizing sound textures through wavelet tree learning. IEEE Computer Graphics and Applications, 22(4):38–48, July/August 2002. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

computer.org/cga/cg2002/
g4038abs.htm;   http:/
/dlib.computer.org/cg/
books/cg2002/pdf/g4038.
pdf.

DeFanti:1996:VRH

[DBS96] Thomas A. DeFanti, Maxine D. Brown, and Rick Stevens. Virtual reality over high-speed networks. IEEE Computer Graphics and Applications, 16(4): 42–43, July/August 1996. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Dickinson:1989:IEC

[DBV89] Robert R. Dickinson, Richard H. Bartels, and Allan H. Vermeulen. The interactive editing and contouring of empirical fields. IEEE Computer Graphics and Applications, 9(3):34–43, May/ June 1989. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Daugherty:2008:VHV

[DBV^{+}08] Brent Daugherty, Sabarish Babu, Lori Van Wallendael, Brian Cutler, and Larry F. Hodges. Virtual human versus human administration of photographic lineups. IEEE Computer Graphics and Applications, 28(6):65–75, November/December 2008. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Das:2019:BIM

[DCCE19] S. Das, D. Cashman, R. Chang, and A. Endert. BEAMES: Interactive multimodel steering, selection, and inspection for regression tasks. IEEE Computer Graphics and Applications, 39(5):20–32, September/October 2019. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Dong:2003:NRM

[DCLK03] Feng Dong, Gordon J. Clapworthy, Hai Lin, and Meleagros A. Krokos. Nonphotorealistic rendering of medical volume data. IEEE Computer Graphics and Applications, 23(4):44–52, July/August 2003. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cg/2003/04/g4044abs.htm;
http://csdl.computer.
org/dl/mags/cg/2003/04/
g4044.htm;   http://csdl.
computer.org/dl/mags/cg/
2003/04/g4044.pdf.

Doornink:1984:AEH

[DD84] Douglas J. Doornink and John C. Dalrymple. Architectural evolution of a highperformance graphics terminal. IEEE Computer Graphics and Applications, 4(4): 47–54, April 1984. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Duguet:2004:FPB

[DD04] Florent Duguet and George Drettakis. Flexible pointbased rendering on mobile devices. IEEE Computer Graphics and Applications, 24(4):57–63, July/ August 2004. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http:/

/csdl.computer.org/dl/
mags/cg/2004/04/g4057.
htm;   http://csdl.computer.
org/dl/mags/cg/2004/04/
g4057.pdf.

Dibia:2019:DAG

[DD19] V. Dibia and Ç. Demiralp. Data2Vis: Automatic generation of data visualizations using sequenceto-sequence recurrent neural networks. IEEE Computer Graphics and Applications, 39(5):33–46, September/October 2019. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Debattista:2011:WFS

[DDdSC11] Kurt Debattista, Piotr Dubla, Luis Paulo Peixoto dos Santos, and Alan Chalmers. Wait-free sharedmemory irradiance caching. IEEE Computer Graphics and Applications, 31(5): 66–78, September/October 2011. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Donoho:1988:MDG

[DDG88] Andrew W. Donoho, David L.

Donoho, and Miriam Gasko. MacSpin: Dynamic graphics on a desktop computer. IEEE Computer Graphics and Applications, 8(4):51– 58, July/August 1988. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

deBruyne:1986:CLG

[de 86] P. de Bruyne. A compact large-area graphics digitizer for personal computers. IEEE Computer Graphics and Applications, 6(12):49– 53, December 1986. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

DeFloriani:1989:PDS

[De 89] Leila De Floriani. A pyramidal data structure for triangle-based surface description. IEEE Computer Graphics and Applications, 9(2):67–78, March/ April 1989. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

DeRose:2011:AWC

[De 11]  Tony De Rose.  Andy

Witkin: From computer vision to computer graphics. IEEE Computer Graphics and Applications, 31 (1):102–104, c3, January/ February 2011. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Debevec:2002:IBL

[Deb02]  Paul Debevec.  Image-

based lighting. IEEE Computer Graphics and Applications, 22(2):26–34, March/ April 2002. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

computer.org/cga/cg2002/
g2026abs.htm;   http:/
/dlib.computer.org/cg/
books/cg2002/pdf/g2026.
pdf.

DeFanti:2000:BTB

[DeF00] Tom DeFanti. Better than being there: Next millennium networks. IEEE Computer Graphics and Applications, 20(1):60–61, January/ February 2000. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://dlib.
computer.org/cg/books/
cg2000/pdf/g1060.pdf.

Delaney:1997:AFB

[Del97]  Ben Delaney.  Applica-

tions: Faster and better and cheaper: NASA visualizes the solar system. IEEE Computer Graphics and Applications, 17(6):10–15, November/ December 1997. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://dlib.

computer.org/cg/books/
cg1997/pdf/g6010.pdf.

Delaney:1998:ACH

[Del98a]  Ben Delaney.  About

the cover: High tech in the Magic Kingdom. IEEE Computer Graphics and Applications, 18(6):4–5, November/December 1998. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL

http://dlib.computer.org/
cg/books/cg1998/pdf/g6004.
pdf.

Delaney:1998:AIM

[Del98b]  Ben Delaney.  Applica-

tions: Imaging in medicine — here’s looking in you, kid. IEEE Computer Graphics and Applications, 18(3): 12–19, May/June 1998. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http:

//dlib.computer.org/cg/
books/cg1998/pdf/g3012.
pdf.

Delaney:1998:ATS

[Del98c] Ben Delaney. Applications:

On the trail of the shadow woman: The mystery of motion capture. IEEE Computer Graphics and Applications, 18(5):14–19, September/October 1998. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://dlib.

                                        computer.org/cg/books/
                                        cg1998/pdf/g5014.pdf.

Delaney:1999:AVT

[Del99a] Ben Delaney. Applications: VizSim technology helps find oil faster. IEEE Computer Graphics and Applications, 19(2):10–16, March/ April 1999. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL http://dlib.
                                        computer.org/cg/books/
                                        cg1999/pdf/g2010.pdf.

Delaney:1999:NTF

[Del99b] Ben Delaney. The NYSE’s 3D trading floor. IEEE Computer Graphics and Applications, 19(6):12–15, November/December 1999. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL

                                        http://dlib.computer.org/
                                        cg/books/cg1999/pdf/g6012.
                                        pdf.

Delaney:2000:ACP

[Del00a] Ben Delaney. About the cover: Painting with light. IEEE Computer Graphics and Applications, 20(6):4–5, November/December 2000. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL

                                        http://dlib.computer.org/
                                        cg/books/cg2000/pdf/g6004.
                                        pdf.

Delaney:2000:AVU

[Del00b] Ben Delaney. Applications:

Visualization in urban planning: They didn’t build LA in a day. IEEE Computer Graphics and Applications, 20(3):10–16, May/ June 2000. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://dlib.
computer.org/cg/books/
cg2000/pdf/g3010.pdf.

Delaney:2000:PVG

[Del00c] Ben Delaney. Projects in VR: This is NOT a game: US Army, University of California ink pact for research center. IEEE Computer Graphics and Applications, 20(1):6–9, January/ February 2000. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://dlib.
computer.org/cg/books/
cg2000/pdf/g1006.pdf.

Delaney:2002:CGH

[Del02] Ben Delaney. Computer graphics: Helping to cope with terrorism. IEEE Computer Graphics and Applications, 22(2):16–23, March/ April 2002. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

computer.org/cga/cg2002/
g2016abs.htm;   http://
computer.org/cga/cg2002/
pdf/g2016.pdf;   http:

                                        //dlib.computer.org/cg/
                                        books/cg2002/pdf/g2016.
                                        pdf.

Delaney:2005:FFG

[Del05] Ben Delaney. Forget the funny glasses. IEEE Computer Graphics and Applications, 25(3):14–19, May/ June 2005. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Denzer:1993:GED

[Den93] Ralf Denzer. Graphics for environmental decision making. IEEE Computer Graphics and Applications, 13(2): 58–64, March/April 1993. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Desurvire:2013:MGU

[DEN13] Heather Desurvire and Magy Seif El-Nasr. Methods for game user research: Studying player behavior to enhance game design. IEEE Computer Graphics and Applications, 33(4): 82–87, July/August 2013. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Dev:1999:GEI

[Dev99a] Parvati Dev. Guest Editor’s introduction: Imaging in medical education. IEEE Computer Graphics and Applications, 19(3):20, May/June 1999. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://dlib.
computer.org/cg/books/
cg1999/pdf/g3020.pdf.

Dev:1999:TIV

[Dev99b]  Parvati Dev.  Tutorial:

Imaging and visualization in medical education. IEEE Computer Graphics and Applications, 19(3):21–31, May/June 1999. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

computer.org/cga/cg1999/
g3021abs.htm;   http:/
/dlib.computer.org/cg/
books/cg1999/pdf/g3021.
pdf.

Dev:2000:GIM

[Dev00] Parvati Dev. Graphics and imaging in medicine. IEEE Computer Graphics and Applications, 20(1):24–25, January/February 2000. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http:

//dlib.computer.org/cg/
books/cg2000/pdf/g1024.
pdf.

Dev:2013:MER

[Dev13] Kapil Dev. Mobile expressive renderings: The state of the art. IEEE Computer Graphics and Applications, 33(3):22–31, May/ June 2013. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

DeFanti:1987:CPB

[DFL87] Thomas A. DeFanti, Richard Frankel, and Lawrence A. Leske. A call for the publishing of Blt-stones. IEEE Computer Graphics and Applications, 7(10):39–49, October 1987. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Doraiswamy:2018:STU

[DFL^{+}18] Harish Doraiswamy, Juliana Freire, Marcos Lage, Fabio Miranda, and Claudio Silva. Spatio-temporal urban data analysis: A visual analytics perspective. IEEE Computer Graphics and Applications, 38(5):26–35, September/October 2018. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        cg/2018/05/mcg2018050026-
                                        abs.html.

Dogrusoz:2002:GVT

[DFM^{+}02] Ugur Dogrusoz, Qingwen Feng, Brendan Madden, Michael Doorley, and Arne Frick. Graph visualization toolkits. IEEE Computer Graphics and Applications, 22(1):30–37, January/ February 2002. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://dlib.
computer.org/cg/books/
cg2002/g1030abs.htm;
http://dlib.computer.org/
cg/books/cg2002/pdf/g1030.
pdf.

DeCrescenzio:2011:ARA

[DFP^{+}11] Francesca De Crescenzio, Massimiliano Fantini, Franco Persiani, Luigi Di Stefano, Pietro Azzari, and Samuele Salti. Augmented reality for aircraft maintenance training and operations support. IEEE Computer Graphics and Applications, 31(1):96–101, January/February 2011. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Battista:2012:VAD

[DFQ12] Giuseppe Di Battista, Jean-Daniel Fekete, and Huamin Qu. Visualization applications and design studies. IEEE Computer Graphics and Applications, 32(1):20– 21, January/February 2012. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Dischler:1999:IIB

[DG99] Jean-Michel Dischler and Djamchid Ghazanfarpour. Interactive image-based modeling of macrostructured textures. IEEE Computer Graphics and Applications, 19(1):66–74, January/ February 1999. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

                                        computer.org/cga/cg1999/
                                        g1066abs.htm;   http:/
                                        /dlib.computer.org/cg/
                                        books/cg1999/pdf/g1066.
                                        pdf.

Dias:1997:MMV

[DGA^{+}97] Jose Miguel Salles Dias, Ricardo Galli, Antonio Carlos Almeida, Carlos A. C. Belo, and Jose Manuel Rebordao. mWorld: a multiuser 3D virtual environment. IEEE Computer Graphics and Applications, 17(2):55–65, March/April 1997. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

DAmico:2007:VDC

[DGTK07] Anita D. D’Amico, John R. Goodall, Daniel R. Tesone, and Jason K. Kopylec. Visual discovery in computer network defense. IEEE Computer Graphics and Applications, 27(5):20–27, September/October 2007. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Dietrich:2007:MMR

[DGY07] Andreas Dietrich, Enrico Gobbetti, and Sung-Eui Yoon. Massive-model rendering techniques: a tutorial. IEEE Computer Graphics and Applications, 27(6): 20–34, November/December 2007. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Davenport:1991:NMC

[DH91] Glorianna Davenport and Jonathan D. Harber. Numbers: a medium that counts. IEEE Computer Graphics and Applications, 11(4): 39–44, July/August 1991. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Delmarcelle:1993:VST

[DH93] Thierry Delmarcelle and Lambertus Hesselink. Visualizing second-order tensor fields with hyperstream lines. IEEE Computer Graphics and Applications, 13(4):25–33, July/August 1993. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Dashti:2022:RIP

[DHC^{+}22] Sarah Dashti, Fiaz Hussain, Fiona Carroll, Edmond Prakash, and Andres Navarro-Newball. Role of intricate pottery visualization in ceramic manufacturing. IEEE Computer Graphics and Applications, 42(5):90–97, September/October 2022. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

deHaan:2010:SNC

[dHPP10] Gerwin de Haan, Huib Piguillet, and Frits Post. Spatial navigation for context-aware video surveillance. IEEE Computer Graphics and Applications, 30(5):20–31, September/October 2010. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Dias:1991:RTI

[Dia91] Maria Lurdes Dias. Ray tracing interference color. IEEE Computer Graphics and Applications, 11(2): 54–60, March/April 1991. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Dias:1994:RTI

[Dia94] Maria Lurdes Dias. Ray tracing interference color: visualizing Newton’s rings. IEEE Computer Graphics and Applications, 14(3):17– 20, May/June 1994. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Dietrich:1985:VIF

[Die85] Frank Dietrich. Visual intelligence: the first decade of computer art (1965–1975). IEEE Computer Graphics and Applications, 5(7):33– 45, July/August 1985. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Dietrich:2000:DAD

[Die00] Frank Dietrich. 3D dreams — the art of David Em. IEEE Computer Graphics and Applications, 20(2):4–5, March/April 2000. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://dlib.
computer.org/cg/books/
cg2000/pdf/g2004.pdf.

Diefenbach:2011:PGD

[Die11] Paul J. Diefenbach. Practical game design and development pedagogy. IEEE Computer Graphics and Applications, 31(3):84–88, May/June 2011. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Dill:1996:TSP

[Dil96] John Dill. Top 10 software products for 1995. IEEE Computer Graphics and Applications, 16(1):92, January/February 1996. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Dill:1997:TSP

[Dil97] John Dill. The top 10 software products of 1996. IEEE Computer Graphics and Applications, 17(1): 89, January/February 1997.

CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Dill:1998:SII

[Dil98a] John Dill. Special introduction to this issue. IEEE Computer Graphics and Applications, 18(4):16, July/ August 1998. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Dill:1998:TTS

[Dil98b] John Dill. Top ten software products of 1997. IEEE Computer Graphics and Applications, 18(1):94, January/February 1998. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

computer.org/cga/cg1998/
g1094abs.htm;   http:/
/dlib.computer.org/cg/
books/cg1998/pdf/g1094.
pdf.

Dill:1999:TSN

[Dil99] John Dill. Top 10 software: New products of 1998. IEEE Computer Graphics and Applications, 19(1):93, January/February 1999. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http:

//dlib.computer.org/cg/
books/cg1999/pdf/g1093.
pdf.

Dill:2001:TSP

[Dil01] John Dill. Top 10 software products of 2000. IEEE Computer Graphics and Applications, 21(1):93–94, January/February 2001. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http:

//dlib.computer.org/cg/
books/cg2001/pdf/g1094.
pdf.

Dill:2004:FEF

[Dil04] John C. Dill. Following the ebb and flow at CG&A. IEEE Computer Graphics and Applications, 24(2):c2, March/April 2004. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cg/2004/02/g20c2.pdf;
http://csdl.computer.org/
dl/mags/cg/2004/02/g20c2.
htm.

DiLorenzo:2015:PDA

[DiL15] Paul Carmen DiLorenzo.

Premo: DreamWorks Animation’s new approach to animation. IEEE Computer Graphics and Applications, 35(4):14–21, July/ August 2015. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
cg/2015/04/mcg2015040014-
abs.html.

Dionisio:1997:PVV

[Dio97] José Dionisio. Projects in VR: Virtual Hell: a trip through the flames. IEEE Computer Graphics and Applications, 17(3):11–14, May/June 1997. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

DiSalvo:2014:GQE

[DiS14]  Betsy DiSalvo.  Graph-

ical qualities of educational technology: Using drag-and-drop and textbased programs for introductory computer science. IEEE Computer Graphics and Applications, 34(6):12–15, November/ December 2014. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL http://csdl.
                                        computer.org/csdl/mags/
                                        cg/2014/06/mcg2014060012-
                                        abs.html.

Demuynck:2013:MCN

[DJM13] Olivier Demuynck, José, and Manuel Menéndez. Magic Cards: A new augmentedreality approach. IEEE Computer Graphics and Applications, 33(1):12–19, January/February 2013. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Dondera:2008:VCE

[DJP^{+}08] Radu Dondera, Chun Jia, Voicu Popescu, Cristina Nita-Rotaru, Melissa Dark, and Cynthia S. York. Virtual classroom extension for effective distance education. IEEE Computer Graphics and Applications, 28(1):64– 74, January/February 2008. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Dou:2009:RRP

[DJS^{+}09] Wenwen Dou, Dong Hyun Jeong, Felesia Stukes, William Ribarsky, Heather Richter Lipford, and Remco Chang. Recovering reasoning processes from user interactions. IEEE Computer Graphics and Applications, 29(3):52–61, May/ June 2009. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Daubert:2003:ELT

[DKS^{+}03] Katja Daubert, Jan Kautz, Hans-Peter Seidel, Wolfgang Heidrich, and Jean-Michel Dischler. Efficient light transport using precomputed visibility. IEEE Computer Graphics and Applications, 23(3):28–37, May/ June 2003. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://dlib.
computer.org/cg/books/
cg2003/pdf/g3028.pdf;

                                        http://www.computer.org/
                                        cga/cg2003/g3028abs.htm.

Dennis:2005:DVF

[DKS^{+}05] Brent Dennis, Sarat Kocherlakota, Amit Sawant, Laura Tateosian, and Christopher G. Healey. Designing a visualization framework for multidimensional data. IEEE Computer Graphics and Applications, 25(6): 10–15, November/December 2005. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Dev:2015:DDC

[DL15] Kapil Dev and Manfred Lau. Democratizing digital content creation using mobile devices with inbuilt sensors. IEEE Computer Graphics and Applications, 35(1):84–94, January/ February 2015. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL http://csdl.
                                        computer.org/csdl/mags/
                                        cg/2015/01/mcg2015010084-
                                        abs.html.

Dou:2016:TTO

[DL16] Wenwen Dou and Shixia Liu. Topic- and timeoriented visual text analysis. IEEE Computer Graphics and Applications, 36(4):8– 13, July/August 2016. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL https:/

/www.computer.org/csdl/
mags/cg/2016/04/mcg2016040008-
abs.html.

Dev:2022:LPA

[DL22] Kapil Dev and Manfred Lau.

Learning perceptual aesthetics of 3-D shapes from multiple views. IEEE Computer Graphics and Applications, 42(1):20–31, January/ February 2022. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Dunser:2015:VMC

[DLEF15] Andreas Dunser, Martin Lochner, Ulrich Engelke, and David Rozado Fernandez. Visual and manual control for human-robot teleoperation. IEEE Computer Graphics and Applications, 35(3):22–32, May/ June 2015. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
cg/2015/03/mcg2015030022-
abs.html.

Deng:2005:AEM

[DLN05] Zhigang Deng, J. P. Lewis, and Ulrich Neumann. Automated eye motion using texture synthesis. IEEE Computer Graphics and Applications, 25(2):24–30, March/ April 2005. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Debevec:2002:GEI

[DM02a] Paul Debevec and Leonard McMillan. Guest Editors’ introduction: Imagedbased modeling, rendering, and lighting. IEEE Computer Graphics and Applications, 22(2):24–25, March/ April 2002. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

                                        computer.org/cga/cg2002/
                                        g2024abs.htm;   http:/
                                        /dlib.computer.org/cg/
                                        books/cg2002/pdf/g2024.
                                        pdf.

Dorst:2002:GAC

[DM02b] Leo Dorst and Stephen Mann. Geometric algebra: a computational framework for geometrical applications (Part 1). IEEE Computer Graphics and Applications, 22(3):24–31, May/ June 2002. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

                                        computer.org/cga/cg2002/
                                        g3024abs.htm;   http:/
                                        /dlib.computer.org/cg/
                                        books/cg2002/pdf/g3024.
                                        pdf.

Djavaherpour:2017:PVG

[DMAS17] Hessam Djavaherpour, Ali Mahdavi-Amiri, and Faramarz F. Samavati. Physical visualization of geospatial datasets. IEEE Computer Graphics and Applications, 37(3):61–69, May/ June 2017. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2017/03/mcg2017030061-
abs.html.

Darken:2005:DOS

[DMJ05] Rudy Darken, Perry Mc-Dowell, and Erik Johnson. The Delta3D open source game engine. IEEE Computer Graphics and Applications, 25(3):10–12, May/ June 2005. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Dutta:1993:CSS

[DMP93] Debasish Dutta, Ralph R.

Martin, and Michael J. Pratt. Cyclides in surface and solid modeling. IEEE Computer Graphics and Applications, 13(1):53– 59, January/February 1993. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Dodd:1983:SSI

[DMR83] S. L. Dodd, D. F. McAllister, and J. A. Roulier. Shape-preserving spline interpolation for specifying bivariate functions on grids. IEEE Computer Graphics and Applications, 3(6):70– 79, September 1983. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Djavaherpour:2023:FII

[DMSJ23] Hessam Djavaherpour, Lynn Moorman, Faramarz Samavati, and Yvonne Jansen. First insights into INTUIT: an INteractive Tactile Physicalization for User Interpretation of RADAR Technology. IEEE Computer Graphics and Applications, 43(5): 91–98, September/October 2023. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Drews:2016:UDE

[DNBC16] Paulo L. J. Drews, Erickson R. Nascimento, Silvia S. C. Botelho, and Mario Fernando Montenegro Campos. Underwater depth estimation and image restoration based on single images. IEEE Computer Graphics and Applications, 36(2):24–35, March/ April 2016. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        cg/2016/02/mcg2016020024-
                                        abs.html.

Dias:2015:CCA

[DNS^{+}15] Paulo Dias, Luis Neves, Daniel Santos, Catarina Coelho, Maria Teresa Ferreira, Helder Santos, Samuel Silva, and Beatriz Sousa Santos. CraMs: Craniometric analysis application using 3D skull models. IEEE Computer Graphics and Applications, 35(6):11–17, November/ December 2015. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
cg/2015/06/mcg2015060011-
abs.html.

Gaya:2019:SIR

[dOGDM^{+}19] Joel Felipe de Oliveira Gaya, Amanda Duarte, Felipe Codevilla Moraes, Paulo Drews-Jr, and Silvia Silva da Costa Botelho. Single image restoration for participating media based on prior fusion. IEEE Computer Graphics and Applications, 39(1):71–83, January/February 2019. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL https:

//ieeexplore.ieee.org/
document/8663517/.

Duer:2020:MIV

[DOH^{+}20] Z. Duer, T. Ogle, D. Hicks, S. Fralin, T. Tucker, and R. Yu. Making the invisible visible: Illuminating the hidden histories of the World War I tunnels at Vauquois through a hybridized virtual reality exhibition. IEEE Computer Graphics and Applications, 40(4):39–50, July/ August 2020. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Domik:2000:VVD

[Dom00] Gitta Domik. Visualization viewpoints: Do we need formal education in visualization? IEEE Computer Graphics and Applications, 20(4):16–19, July/ August 2000. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL http://dlib.
                                        computer.org/cg/books/
                                        cg2000/pdf/g4016.pdf.

Domik:2012:FCS

[Dom12]  Gitta Domik.  Foster-

ing collaboration and selfmotivated learning: Best practices in a one-semester visualization course. IEEE Computer Graphics and Applications, 32(1):87–91, January/February 2012. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Dooley:1982:AMP

[Doo82]  Marianne Dooley.  An-

thropometric modeling programs — a survey. IEEE Computer Graphics and Applications, 2(9):17–25, November/December 1982. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

deOliveira:2004:SWE

[dOYG04] Jauvane C. de Oliveira, Seok Jong Yu, and Nicolas D. Georganas. Synchronized world embedding in virtual environments. IEEE Computer Graphics and Applications, 24(4):73– 83, July/August 2004. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http:

//csdl.computer.org/dl/
mags/cg/2004/04/g4073.
htm;   http://csdl.computer.
org/dl/mags/cg/2004/04/
g4073.pdf.

Dawson:1982:CDP

[DP82] Fred T. Dawson and Robert H.

Philhower. Computer-aided design of plant electrical systems. IEEE Computer Graphics and Applications, 2(1):31–34, 36–37, January/ February 1982. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Djurcilov:2000:VSG

[DP00] Suzana Djurcilov and Alex Pang. Visualizing sparse gridded data sets. IEEE Computer Graphics and Applications, 20(5):52–57, September/October 2000. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL

http://computer.org/cga/
cg2000/g5052abs.htm;

                                        http://dlib.computer.org/
                                        cg/books/cg2000/pdf/g5052.
                                        pdf.

Durupinar:2011:HOP

[DPA^{+}11] Funda Durupinar, Nuria Pelechano, Jan M. Allbeck, Ugur Gudukbay, and Norman I. Badler. How the ocean personality model affects the perception of crowds. IEEE Computer Graphics and Applications, 31(3):22–31, May/ June 2011. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Duer:2018:BVR

[DPG18] Zach Duer, Leo Piilonen, and George Glasson. Belle2VR: A virtual-reality visualization of subatomic particle physics in the Belle II experiment. IEEE Computer Graphics and Applications, 38(3):33–43, May/ June 2018. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        cg/2018/03/mcg2018030033-
                                        abs.html.

Diepenbrock:2011:IVC

[DPL^{+}11] Stefan Diepenbrock, Jorg-Stefan Prassni, Florian Lindemann, Hans-Werner Bothe, and Timo Ropinski. 2010 IEEE Visualization Contest winner: Interactive planning for brain tumor resections. IEEE Computer Graphics and Applications, 31(5):6–13, September/October 2011. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Dev:2023:CSR

[DPL23] Kapil Dev, Letizia Palumbo, and Manfred Lau. Comparing shape representations for the aesthetic judgment of 3-D shape pairs. IEEE Computer Graphics and Applications, 43(6):33–45, November/December 2023. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Dube:1983:MGI

[DR83] R. Peter Dube and Marcia Rivers Smith. Managing geometric information with a database management system. IEEE Computer Graphics and Applications, 3(7):57–62, October 1983. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Davis:2003:ECV

[DRHL^{+}03] Larry Davis, Jannick Rolland, Felix Hamza-Lup, Yonggang Ha, Jack Norfleet, Beth Pettitt, and Celina Imielinska. Enabling a continuum of virtual environment experiences. IEEE Computer Graphics and Applications, 23(2):10–12, March/April 2003. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL http://dlib.
                                        computer.org/cg/books/
                                        cg2003/pdf/g2010.pdf.

deRooij:2010:MIC

[dRvWW10] Ork de Rooij, Jarke J. van Wijk, and Marcel Worring. MediaTable: Interactive categorization of multimedia collections. IEEE Computer Graphics and Applications, 30(5):42–51, September/October 2010. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

DeFanti:1987:UIP

[DS87] Thomas A. DeFanti and Daniel J. Sandin. The usable intersection of PC graphics and NTSC video recording. IEEE Computer Graphics and Applications, 7(10):50– 58, October 1987. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Dill:2007:NO

[DS07a] John C. Dill and Maureen C. Stone. In with the new, out with the old. IEEE Computer Graphics and Applications, 27(2):6, March/ April 2007. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cg/2007/02/g2006.pdf.

Durand:2007:GEI

[DS07b] Frédo Durand and Richard Szeliski. Guest Editors’ introduction: Computational photography. IEEE Computer Graphics and Applications, 27(2):21–22, March/ April 2007. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cg/2007/02/g2021.pdf.

Dias:2019:UVR

[DSA^{+}19] Paulo Dias, Ricardo Silva, Paula Amorim, Jorge Laíns, Eulália Roque, Inês Serôdio, Fátima Pereira, Beatriz Sousa Santos, and Mike Potel. Using virtual reality to increase motivation in poststroke rehabilitation. IEEE Computer Graphics and Applications, 39(1):64–70, January/February 2019. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL https:

//ieeexplore.ieee.org/
document/8663653/.

Demiralp:2014:VEM

[DSK^{+}14] Cagatay Demiralp, Carlos E. Scheidegger, Gordon L. Kindlmann, David H. Laidlaw, and Jeffrey Heer. Visual embedding: A model for visualization. IEEE Computer Graphics and Applications, 34(1):10–15, January/February 2014. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Donovan:1995:PPM

[DSKH95] Walt Donovan, Paolo Sabella, Ihtisham Kabir, and Michael M. Hsieh. Pixel processing in a memory controller. IEEE Computer Graphics and Applications, 15(1):51– 61, January/February 1995. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Nunes:2016:MKA

[dSNRdLdSN16] Eunice P. dos Santos Nunes, Licinio G. Roque, and Fatima de Lourdes dos Santos Nunes. Measuring knowledge acquisition in 3D virtual learning environments. IEEE Computer Graphics and Applications, 36(2):58–67, March/ April 2016. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2016/02/mcg2016020058-
abs.html.

Davenport:1991:CPM

[DSP91] Glorianna Davenport, Thomas Aguierre Smith, and Natalio Pincever. Cinematic primitives for multimedia. IEEE Computer Graphics and Applications, 11(4):67–74, July/ August 1991. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Dias:2014:SPI

[DSP^{+}14] Paulo Dias, Tiago Sousa, Joao Parracho, Igor Cardoso, Andre Monteiro, and Beatriz Sousa Santos. Student projects involving novel interaction with large displays. IEEE Computer Graphics and Applications, 34(2):80–86, March/April 2014. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Dietrich:2007:EBR

[DSW07] Andreas Dietrich, Abe Stephens, and Ingo Wald. Exploring a Boeing 777: Ray tracing large-scale CAD data. IEEE Computer Graphics and Applications, 27(6):36–46, November/ December 2007. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Doctor:1981:DTO

[DT81] Louis J. Doctor and John G.

Torborg. Display techniques for octree-encoded objects. IEEE Computer Graphics and Applications, 1(3):29– 38, July/August 1981. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Du:2021:SCV

[DV21] Ruofei Du and Amitabh Varshney. Saliency computation for virtual cinematography in 360^{∘} videos. IEEE Computer Graphics and Applications, 41(4): 99–106, July/August 2021. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Dyer:1987:VSZ

[DW87] Scott Dyer and Scott Whitman. A vectorized scanline Z-buffer rendering algorithm. IEEE Computer Graphics and Applications, 7(7):34–45, July/ August 1987. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Dunnett:1992:ICH

[DWL^{+}92] Graham J. Dunnett, Martin White, Paul F. Lister, Richard L. Grimsdale, and France Glemot. The Image chip for high performance 3D rendering. IEEE Computer Graphics and Applications, 12(6):41–52, November/December 1992. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Dwyer:2016:NVH

[Dwy16] Tim Dwyer. Network visualization as a higher-order visual analysis tool. IEEE Computer Graphics and Applications, 36(6):78–85, November/December 2016. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL

https://www.computer.org/
csdl/mags/cg/2016/06/mcg2016060078-
abs.html.

Dong:2007:MTS

[DY07] Feng Dong and Xujiong Ye.

Multiscaled texture synthesis using multisized pixel neighborhoods. IEEE Computer Graphics and Applications, 27(3):41–47, May/ June 2007. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Dixit:2009:DSI

[DY09] Priyesh N. Dixit and G. Michael Youngblood. Discovering 3D surface information values from gameplayers. IEEE Computer Graphics and Applications, 29(2):30–38, March/April 2009. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Dyer:1990:DTV

[Dye90] D. Scott Dyer. A dataflow toolkit for visualization. IEEE Computer Graphics and Applications, 10(4): 60–69, July/August 1990. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Dobashi:2007:PAR

[DYS^{+}07] Yoshinori Dobashi, Tsuyoshi Yamamoto, Makoto Sato, Shoichi Hasegawa, Mitsuaki Kato, and Tomoyuki Nishita. A precomputed approach for real-time haptic interaction with fluids. IEEE Computer Graphics and Applications, 27 (3):90–92, May/June 2007. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

El-Assady:2022:WBR

[EAM22] Mennatallah El-Assady and Caterina Moruzzi. Which biases and reasoning pitfalls do explanations trigger? Decomposing communication processes in human AI interaction. IEEE Computer Graphics and Applications, 42(6):11–23, November/December 2022. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Earnshaw:1997:GEI

[Ear97] Rae Earnshaw. Guest Editors’ introduction. IEEE Computer Graphics and Applications, 17(2):30–31, March/April 1997. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Earnshaw:1998:AFI

[Ear98] Rae Earnshaw. Applications: Fraunhofer Institute:

Building on a decade of computer graphics research. IEEE Computer Graphics and Applications, 18(2):10– 16, March/April 1998. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http:

//dlib.computer.org/cg/
books/cg1998/pdf/g2010.
pdf.

Earnshaw:2001:GEI

[Ear01] Rae Earnshaw. Guest Editor’s introduction: Digital media. IEEE Computer Graphics and Applications, 21(1):14–17, January/ February 2001. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://dlib.
computer.org/cg/books/
cg2001/pdf/g1014.pdf.

Eastman:1990:VVR

[Eas90] Charles M. Eastman. Vector versus raster: a functional comparison of drawing technologies. IEEE Computer Graphics and Applications, 10(5):68–80, September 1990. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Encarnacao:2000:PVW

[EBBS00] L. Miguel Encarnação, Robert J. Barton III, Oliver Bimber, and Dieter Schmalstieg. Projects in VR: Walk-up VR: Virtual reality beyond projection screens. IEEE Computer Graphics and Applications, 20(6):19–23, November/ December 2000. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL http://dlib.
                                        computer.org/cg/books/
                                        cg2000/pdf/g6019.pdf.

Ebenreuter:2005:DMF

[Ebe05] Natalie Ebenreuter. Dance movement: a focus on the technology. IEEE Computer Graphics and Applications, 25(6):80–83, November/December 2005. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Ebert:2017:MMW

[Ebe17] David S. Ebert. In memoriam: Martin William Ribarsky. IEEE Computer Graphics and Applications, 37(4):6–7, July/ August 2017. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        cg/2017/04/mcg2017040006.
                                        html.

Ellis:2010:DTM

[EBG^{+}10] Christopher Ellis, Pavel Babenko, Brian Goldiez, Jason Daly, and Glenn A. Martin. Dynamic terrain for multiuser real-time environments. IEEE Computer Graphics and Applications, 30(1):80–84, January/ February 2010. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Endert:2013:BCP

[EBN13] Alex Endert, Lauren Bradel, and Chris North. Beyond control panels: Direct manipulation for visual analytics. IEEE Computer Graphics and Applications, 33 (4):6–13, July/August 2013. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Emering:1998:PVI

[EBT98] Luc Emering, Roman Boulic, and Daniel Thalmann. Projects in VR: Interacting with virtual humans through body actions. IEEE Computer Graphics and Applications, 18(1):8–11, January/February 1998. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

computer.org/cga/cg1998/
g1008abs.htm;   http:/
/dlib.computer.org/cg/
books/cg1998/pdf/g1008.
pdf.

Endert:2015:SIC

[ECNZ15] Alex Endert, Remco Chang, Chris North, and Michelle Zhou. Semantic interaction: Coupling cognition and computation through usable interactive analytics. IEEE Computer Graphics and Applications, 35(4):94–99, July/August 2015. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
cg/2015/04/mcg2015040094-
abs.html.

Encarnacao:2015:FDC

[ECS^{+}15] L. Miguel Encarnacao, Yung-Yu Chuang, Andre Stork, Dave Kasik, Theresa-Marie Rhyne, Lisa Avila, Jorn Kohlhammer, L. Miguel Encarnacao, Joseph J. LaViola, Melanie Tory, John Dill, Gitta Domik, G. Scott Owen, and Pak Chung Wong. Future directions in computer graphics and visualization: From CG&A’s Editorial Board. IEEE Computer Graphics and Applications, 35(1):20–32, January/ February 2015. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://csdl.
computer.org/csdl/mags/
cg/2015/01/mcg2015010020.
html.

Encarnacao:1993:GEI

[ED93] Jose Encarnacao and John C. Dill. Guest Editors’ introduction. IEEE Computer Graphics and Applications, 13(5):15–??, September 1993. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Eischen:1996:FEM

[EDC96] Jeffrey W. Eischen, Shigan Deng, and Timothy G. Clapp. Finite-element modeling and control of flexible fabric parts. IEEE Computer Graphics and Applications, 16(5):71–80, September 1996. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Elmqvist:2012:LMV

[EE12] Niklas Elmqvist and David S.

Ebert. Leveraging multidisciplinarity in a visual analytics graduate course. IEEE Computer Graphics and Applications, 32(3):84– 87, May/June 2012. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Encarnacao:1994:RIP

[EFB^{+}94] J. Encarnacao, J. Foley, S. Bryson, S. K. Feiner, and N. Gershon. Research issues in perception and user interfaces. IEEE Computer Graphics and Applications, 14(2):67–69, March/ April 1994. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Elliott-Famularo:2005:GEI

[EFD05] Heather Elliott-Famularo and Paul Dietz. Guest Editors’ introduction: Emerging technologies 2004. IEEE Computer Graphics and Applications, 25(1):18–19, January/February 2005. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

csdl.computer.org/comp/
mags/cg/2005/01/g1018.
pdf.

Eisert:1998:AFE

[EG98] Peter Eisert and Bernd Girod. Analyzing facial expressions for virtual conferencing. IEEE Computer Graphics and Applications, 18(5):70–78, September/October 1998. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

computer.org/cga/cg1998/
g5070abs.htm;   http:/
/dlib.computer.org/cg/
books/cg1998/pdf/g5070.
pdf.

Encarnacao:1994:EAV

[EGR94] Jose Encarnacao, Martin Gobel, and Lawrence Roseblum. European activities in virtual reality. IEEE Computer Graphics and Applications, 14(1):66–74, January/ February 1994. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Emmel:1999:TCP

[EH99] Patrick Emmel and Roger David Hersch. Towards a color prediction model for printed patches. IEEE Computer Graphics and Applications, 19(4):54–60, July/ August 1999. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

computer.org/cga/cg1999/
g4054abs.htm;   http:/
/dlib.computer.org/cg/
books/cg1999/pdf/g4054.
pdf.

Encarnacao:2003:GEI

[EH03] L. Miguel Encarnação and Lawrence J. Hettinger. Guest Editors’ introduction: Perceptual multimodal interfaces. IEEE Computer Graphics and Applications, 23(5):24–25, September/October 2003. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http:

//csdl.computer.org/dl/
mags/cg/2003/05/g5024.
pdf.

Erdem:2016:ACA

[EH16] Ali Naci Erdem and Ugur Halici. Applying computational aesthetics to a video game application using machine learning. IEEE Computer Graphics and Applications, 36(4):23–33, July/ August 2016. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        cg/2016/04/mcg2016040023-
                                        abs.html.

Echegaray:2014:BSS

[EHA^{+}14] Goretti Echegaray, Imanol Herrera, Iker Aguinaga, Carlos Buchart, and Diego Borro. A brain surgery simulator. IEEE Computer Graphics and Applications, 34(3):12–18, May/ June 2014. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

El-Hakim:2004:DRL

[EHBPG04] Sabry F. El-Hakim, J.- Angelo Beraldin, Michel Picard, and Guy Godin. Detailed 3D reconstruction of large-scale heritage sites with integrated techniques. IEEE Computer Graphics and Applications, 24(3):21–29, May/ June 2004. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL http://csdl.
                                        computer.org/comp/mags/
                                        cg/2004/03/g3021abs.htm;
                                         http://csdl.computer.
                                        org/dl/mags/cg/2004/03/
                                        g3021.htm;   http://csdl.
                                        computer.org/dl/mags/cg/
                                        2004/03/g3021.pdf.

Ens:2017:SAI

[EI17] Barrett Ens and Pourang Irani. Spatial analytic interfaces: Spatial user interfaces for in situ visual analytics. IEEE Computer Graphics and Applications, 37(2):66–79, March/ April 2017. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2017/02/mcg2017020066-
abs.html.

Eick:1996:ANV

[Eic96a] Stephen G. Eick. Aspects of network visualization. IEEE Computer Graphics and Applications, 16(2): 69–72, March/April 1996. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Eick:1996:AV

[Eic96b] Stephen G. Eick. Aspects of visualization. IEEE Computer Graphics and Applications, 16(2):60–??, March/ April 1996. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Eick:2005:IV

[Eic05] Stephen G. Eick. Information visualization at 10. IEEE Computer Graphics and Applications, 25(1):12– 14, January/February 2005. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Elvins:1998:EHF

[EJ98] T. Todd Elvins and Ramesh Jain. Engineering a human factor-based geographic user interface. IEEE Computer Graphics and Applications, 18(3):66–77, May/ June 1998. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

computer.org/cga/cg1998/
g3066abs.htm;   http:/
/dlib.computer.org/cg/
books/cg1998/pdf/g3066.
pdf.

Elber:1999:CRB

[EK99] Gershon Elber and Myung-Soo Kim. Computing rational bisectors. IEEE Computer Graphics and Applications, 19(6):76–81, November/December 1999. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

computer.org/cga/cg1999/
g6076abs.htm;   http:/
/dlib.computer.org/cg/
books/cg1999/pdf/g6076.
pdf.

Eirich:2022:MVA

[EKM^{+}22] Joscha Eirich, Georgios Koutroulis, Belgin Mutlu, Dominik Jäckle, Roman Kern, Tobias Schreck, and Jürgen Bernard. ManEx: The visual analysis of measurements for the assessment of errors in electrical engines. IEEE Computer Graphics and Applications, 42(2):68–80, March/ April 2022. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Encarnacao:2022:VAE

[EKS22] L. Miguel Encarnação, Jörn Kohlhammer, and Chad A. Steed. Visualization for AI explainability. IEEE Computer Graphics and Applications, 42(6):9–10, November/December 2022. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Elber:2001:RPS

[Elb01] Gershon Elber. Rendering with parallel stripes. IEEE Computer Graphics and Applications, 21(3):44– 52, May/June 2001. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

computer.org/cga/cg2001/
g3044abs.htm;   http:/
/dlib.computer.org/cg/
books/cg2001/pdf/g3044.
pdf.

Elber:2005:GTM

[Elb05] Gershon Elber. Geometric texture modeling. IEEE Computer Graphics and Applications, 25(4):66–76, July/August 2005. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Elber:1997:COC

[ELK97] Gershon Elber, In-Kwon Lee, and Myung-Soo Kim. Comparing offset curve approximation methods. IEEE Computer Graphics and Applications, 17(3):62–71, May/June 1997. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Ellis:1994:WVE

[Ell94a] Stephen R. Ellis. What are virtual environments? IEEE Computer Graphics and Applications, 14(1):17– 22, January/February 1994. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Ellsworth:1994:NAI

[Ell94b] David A. Ellsworth. A new algorithm for interactive graphics on multicomputers. IEEE Computer Graphics and Applications, 14(4): 33–40, July/August 1994. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Ehrlich:2009:AVE

[EM09] Justin A. Ehrlich and James R. Miller. Applications: a virtual environment for teaching social skills: AViSSS. IEEE Computer Graphics and Applications, 29(4):10–16, July/ August 2009. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Earnshaw:1998:GEI

[EMTTT98] Rae Earnshaw, Nadia Magnenat-Thalmann, Demetri Terzopoulos, and Daniel Thalmann. Guest Editors’ introduction: Computer animation for virtual humans. IEEE Computer Graphics and Applications, 18(5):20–23, September/ October 1998. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://dlib.
computer.org/cg/books/
cg1998/pdf/g5020.pdf.

Encarnacao:2000:CGE

[Enc00] José L. Encarnação. Computer graphics in Europe. IEEE Computer Graphics and Applications, 20(1):62– 63, January/February 2000. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL

http://dlib.computer.org/
cg/books/cg2000/pdf/g1062.
pdf.

Encarnacao:2004:TSP

[Enc04] Miguel Encarnaçao. Top software products of 2003. IEEE Computer Graphics and Applications, 24(1): 93, January/February 2004. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL

http://csdl.computer.org/

                                        dl/mags/cg/2004/01/g1092.
                                        pdf.

Encarnacao:2006:CGV

[Enc06] José L. Encarnacao. Computer graphics visions and challenges: a European perspective. IEEE Computer Graphics and Applications, 26(4):83–89, July/ August 2006. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Encarnacao:2011:EUV

[Enc11] L. Miguel Encarnaçao. Encouraging the use of visualization technology. IEEE Computer Graphics and Applications, 31(2):6–8, March/April 2011. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Encarnacao:2013:NW

[Enc13] L. Miguel Encarnacao. A new wave. IEEE Computer Graphics and Applications, 33(6):6–8, November/ December 2013. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Encarnacao:2014:B

[Enc14a] L. Miguel Encarnacao. 3D is back! IEEE Computer Graphics and Applications, 34(3):6–7, May/June 2014. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Encarnacao:2014:BF

[Enc14b] L. Miguel Encarnacao. Back to the future. IEEE Computer Graphics and Applications, 34(1):6–8, January/ February 2014. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Encarnacao:2014:BS

[Enc14c] L. Miguel Encarnacao. Behind the scenes. IEEE Computer Graphics and Applications, 34(5):6, September/ October 2014. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://csdl.
computer.org/csdl/mags/
cg/2014/05/mcg2014050006.
html.

Encarnacao:2014:CV

[Enc14d] L. Miguel Encarnacao.

Core values. IEEE Computer Graphics and Applications, 34(2):6–7, March/ April 2014. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Encarnacao:2015:MWS

[Enc15a] L. Miguel Encarnacao.

In memoriam: Wolfgang Straßer (1941–2015). IEEE Computer Graphics and Applications, 35(2):6–7, March/ April 2015. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
cg/2015/02/mcg2015020006.
html.

Encarnacao:2015:IDR

[Enc15b] L. Miguel Encarnacao. Interacting with diverse realities. IEEE Computer Graphics and Applications, 35(3):6, May/ June 2015. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
cg/2015/03/mcg2015030006.
html.

Encarnacao:2015:MEC

[Enc15c] L. Miguel Encarnacao. M1C2: Experiments in community building through collaborative research dissemination. IEEE Computer Graphics and Applications, 35(6):4, November/ December 2015. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
cg/2015/06/mcg2015060004.
html.

Encarnacao:2016:HTD

[Enc16a] L. Miguel Encarnacao. Human touch in digital experiences. IEEE Computer Graphics and Applications, 36(1):4, January/ February 2016. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
cg/2016/01/mcg2016010004.
html.

Encarnacao:2016:SV

[Enc16b] L. Miguel Encarnacao. Scientific visualization. IEEE Computer Graphics and Applications, 36(3):3, May/ June 2016. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2016/03/mcg2016030003.
html.

Encarnacao:2017:IV

[Enc17a] L. Miguel Encarnacao.

Information visualization. IEEE Computer Graphics and Applications, 37(2):6–7, March/April 2017. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2017/02/mcg2017020006.
html.

Encarnacao:2017:QHI

[Enc17b] L. Miguel Encarnacao. The queue is having an issue. IEEE Computer Graphics and Applications, 37(1):6– 7, January/February 2017. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL

https://www.computer.org/
csdl/mags/cg/2017/01/mcg2017010006.
html.

Encarnacao:2023:OCGa

[Enc23a] José L. Encarnação. The origins of computer graphics in Europe: Part 1: The beginnings in Germany. IEEE Computer Graphics and Applications, 43(2):101–112, March/April 2023. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Encarnacao:2023:OCGb

[Enc23b] José L. Encarnação. The origins of computer graphics in Europe: Part 2. The early spreading of computer graphics in Europe. IEEE Computer Graphics and Applications, 43(3):111–121, May/June 2023. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Endert:2014:SIV

[End14] Alex Endert. Semantic interaction for visual analytics: Toward coupling cognition and computation. IEEE Computer Graphics and Applications, 34(4):8–15, July/ August 2014. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

El-Nasr:2013:GAG

[ENDAD13] Magy Seif El-Nasr, Heather Desurvire, Bardia Aghabeigi, and Anders Drachen. Game analytics for game user research, part 1: A workshop review and case study. IEEE Computer Graphics and Applications, 33(2): 6–11, March/April 2013. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

England:1986:GSA

[Eng86] Nick England. A graphics system architecture for interactive applicationspecific display functions. IEEE Computer Graphics and Applications, 6(1):60– 70, January/February 1986. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

England:1997:GEI

[Eng97] Nick England. Guest Editor’s introduction: New image-generation techniques. IEEE Computer Graphics and Applications, 17(1): 39, January/February 1997. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

England:2000:GH

[Eng00] Nick England. Graphics hardware. IEEE Computer Graphics and Applications, 20(1):46–47, January/ February 2000. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://dlib.
computer.org/cg/books/
cg2000/pdf/g1046.pdf.

England:2020:GS

[Eng20] N. England. The graphics system for the 80’s. IEEE Computer Graphics and Applications, 40(3):112–119, May/June 2020. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Eitz:2011:PIS

[ERH^{+}11] Mathias Eitz, Ronald Richter, Kristian Hildebrand, Tamy Boubekeur, and Marc Alexa. Photosketcher: Interactive sketch-based image synthesis. IEEE Computer Graphics and Applications, 31(6): 56–66, November/December 2011. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Elek:2013:RTS

[ERS13] Oskar Elek, Tobias Ritschel, and Hans-Peter Seidel. Realtime screen-space scattering in homogeneous environments. IEEE Computer Graphics and Applications, 33(3):53–65, May/ June 2013. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Ervin:1993:LVE

[Erv93] Stephen M. Ervin. Landscape visualization with Emaps. IEEE Computer Graphics and Applications, 13(2):28–33, March/April 1993. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Edmonds:1982:IHT

[ESS82] Ernest A. Edmonds, Andre Schappo, and Stephen A. R. Scrivener. Image handling in two-dimensional design. IEEE Computer Graphics and Applications, 2(3):75–78, 80–88, July/ August 1982. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Egbert:1996:COM

[EW96] Parris K. Egbert and Scott H. Winkler. Collisionfree object movement using vector fields. IEEE Computer Graphics and Applications, 16(4):18–24, July/ August 1996. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Erbacher:2002:IMD

[EWF02] Robert F. Erbacher, Kenneth L. Walker, and Deborah A. Frincke. Intrusion and misuse detection in large-scale systems. IEEE Computer Graphics and Applications, 22(1):38–48, January/February 2002. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http:

//dlib.computer.org/cg/
books/cg2002/g1038abs.

                                        htm;   http://dlib.computer.
                                        org/cg/books/cg2002/pdf/
                                        g1038.pdf.

Eberhardt:1996:FFP

[EWS96] Bernhard Eberhardt, Andreas Weber, and Wolfgang Strasser. A fast, flexible, particle-system model for cloth draping. IEEE Computer Graphics and Applications, 16(5):52–59, September 1996. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Endo:2003:DMV

[EYY03] Mamoru Endo, Takami Yasuda, and Shigeki Yokoi. A distributed multiuser virtual space system. IEEE Computer Graphics and Applications, 23(1):50–57, January/ February 2003. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL http://dlib.
                                        computer.org/cg/books/
                                        cg2003/pdf/g1050.pdf;
                                        http://www.computer.org/
                                        cga/cg2003/g1050abs.htm.

Ebert:1997:TVD

[EZM^{+}97] David S. Ebert, Amen Zwa, Ethan L. Miller, Christopher D. Shaw, and D. Aaron Roberts. Two-handed volumetric document corpus management. IEEE Computer Graphics and Applications, 17(4):60–62, July/ August 1997. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Franklin:1985:BOS

[FA85] Wm. Randolph Franklin and Varol Akman. Building an octree from a set of parallelepipeds. IEEE Computer Graphics and Applications, 5(10):58–64, October 1985. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Freeman:1998:CVI

[FAB^{+}98] William T. Freeman, David B.

Anderson, Paul A. Beardsley, Chris N. Dodge, Michal Roth, Craig D. Weissman, William S. Yerazunis, Hiroshi Kage, Kazuo Kyuma, Yasunari Miyake, and Ken ichi Tanaka. Computer vision for interactive computer graphics. IEEE Computer Graphics and Applications, 18(3):42–53, May/ June 1998. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

computer.org/cga/cg1998/
g3042abs.htm;   http:/
/dlib.computer.org/cg/
books/cg1998/pdf/g3042.
pdf.

Fogel:2019:CDD

[FAECOG19] S. Fogel, H. Averbuch-Elor, D. Cohen-Or, and J. Goldberger. Clustering-driven deep embedding with pairwise constraints. IEEE Computer Graphics and Applications, 39(4):16–27, July/August 2019. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Foresti:2006:VCN

[FAL^{+}06] Stefano Foresti, James Agutter, Yarden Livnat, Shaun Moon, and Robert Erbacher. Visual correlation of network alerts. IEEE Computer Graphics and Applications, 26(2): 48–59, March/April 2006. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Fant:1986:NRS

[Fan86] Karl M. Fant. A nonaliasing, real-time spatial transform technique. IEEE Computer Graphics and Applications, 6(1):71–80, January/ February 1986. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Farin:1992:CNT

[Far92] Gerald Farin. From conics to NURBS: a tutorial and survey. IEEE Computer Graphics and Applications, 12 (5):78–??, September 1992. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Freund:2017:DSS

[FAZH17] Alina Freund, Nazli Yonca Aydin, Dirk Zeckzer, and Hans Hagen. A decisionsupport system for sustainable water distribution system planning. IEEE Computer Graphics and Applications, 37(1):44–55, January/ February 2017. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2017/01/mcg2017010044-
abs.html.

Franklin:1981:FCS

[FB81] Wm. Randolph Franklin and Alan H. Barr. Faster calculation of superquadric shapes. IEEE Computer Graphics and Applications, 1(3):41–47, July/August 1981. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Filip:1989:PRL

[FB89] Daniel J. Filip and Thomas W.

Ball. Procedurally representing lofted surfaces. IEEE Computer Graphics and Applications, 9(6):27– 33, November/December 1989. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Fowler:1993:CCM

[FB93] Barry Fowler and Richard Bartels. Constraint-based curve manipulation. IEEE Computer Graphics and Applications, 13(5):43–49, September 1993. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Froehlich:2009:UI

[FB09] Bernd Froehlich and Doug Bowman. 3D user interfaces. IEEE Computer Graphics and Applications, 29(6): 20–21, November/December 2009. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Forbes:2015:IMI

[FBML15] Angus Graeme Forbes, Andres Burbano, Paul Murray, and George Legrady. Imagining Macondo: Interacting with García Márquez’s literary landscape. IEEE Computer Graphics and Applications, 35(5):12–19, September/October 2015. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
cg/2015/05/mcg2015050012-
abs.html.

Friston:2022:TSV

[FCS22] Sebastian Friston, Ben Congdon, and Anthony Steed. Teaching social virtual reality with Ubiq. IEEE Computer Graphics and Applications, 42(6):116–122, November/December 2022. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Fontijne:2003:MEG

[FD03] Daniel Fontijne and Leo Dorst. Modeling 3D Euclidean geometry. IEEE Computer Graphics and Applications, 23(2):68–78, March/April 2003. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://dlib.
computer.org/cg/books/
cg2003/pdf/g2068.pdf;
http://www.computer.org/
cga/cg2003/g2068abs.htm.

Fisher:2014:BIA

[FDC14] Danyel Fisher, Steven Drucker, and Mary Czerwinski. Business intelligence analytics [Guest Editors’ introduction]. IEEE Computer Graphics and Applications, 34(5):22–24, September/ October 2014. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://csdl.
computer.org/csdl/mags/
cg/2014/05/mcg2014050022.
html.

Fisher:2012:EVI

[FDK12] Danyel Fisher, Steven M.

Drucker, and A. Christian Konig. Exploratory visualization involving incremental, approximate database queries and uncertainty. IEEE Computer Graphics and Applications, 32(4):

55–62, July/August 2012. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Flanagan:1997:AMF

[FE97] William P. Flanagan and Rae Earnshaw. Applications: Meeting the future at the University of Michigan Media Union. IEEE Computer Graphics and Applications, 17(3):15–19, May/ June 1997. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Feiner:1985:AEA

[Fei85] Steven Feiner. Apex: an experiment in the automated creation of pictorial explanations. IEEE Computer Graphics and Applications, 5(11):29–37, November/December 1985. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Ferwerda:2001:EEV

[Fer01] James A. Ferwerda. Elements of early vision for computer graphics. IEEE Computer Graphics and Applications, 21(5):22–33, September/October 2001. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL

                                        http://dlib.computer.org/
                                        cg/books/cg2001/g5022abs.
                                        htm;   http://dlib.computer.

org/cg/books/cg2001/pdf/
g5022.pdf.

Fetter:1982:PHF

[Fet82]  William A. Fetter.  A

progression of human figures simulated by computer graphics. IEEE Computer Graphics and Applications, 2(9):9–13, November/ December 1982. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Flerackers:2001:CBI

[FEV^{+}01] Chris Flerackers, Rae Earnshaw, Gert Vansichem, Frank Van Reeth, and Frank Alsema. Creating broadcast interactive drama in an NVE. IEEE Computer Graphics and Applications, 21(1):56–60, January/ February 2001. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

computer.org/cga/cg2001/
g1056abs.htm;   http:/
/dlib.computer.org/cg/
books/cg2001/pdf/g1056.
pdf.

Fekete:2020:ERV

[FF20] J. Fekete and J. Freire. Exploring reproducibility in visualization. IEEE Computer Graphics and Applications, 40(5):108–119, September/ October 2020. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Frohlich:1997:PVC

[FFA^{+}97] Bernd Fröhlich, Mrtin Fischer, Maneesh Agrawala, Andrew Beers, and Pat Hanrahan. Projects in VR: Collaborative production modeling and planning. IEEE Computer Graphics and Applications, 17(4):13–15, July/August 1997. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Filho:2019:CIA

[FFN19] J. A. W. Filho, C. M. D. S. Freitas, and L. Nedel. Comfortable immersive analytics with the VirtualDesk metaphor. IEEE Computer Graphics and Applications, 39(3):41–53, May/ June 2019. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Ferwerda:1988:PAA

[FG88a] James A. Ferwerda and Donald P. Greenberg. A psychophysical approach to assessing the quality of antialiased images. IEEE Computer Graphics and Applications, 8(5):85–95, September/October 1988. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Foley:1988:UCC

[FG88b] James D. Foley and Jack Grimes. Using color in computer graphics. IEEE Computer Graphics and Applications, 8(5):25–27, September/October 1988. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Fiorentino:2023:IIR

[FGCB23] Michele Fiorentino, Frank Guan, Yiyu Cai, and Mark Billinghurst. IEEE ISMAR 2022 report: Toward better mixed realities. IEEE Computer Graphics and Applications, 43(3):84–87, May/ June 2023. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Frieder:1985:BBD

[FGR85] Gideon Frieder, Dan Gordon, and R. Anthony Reynolds. Back-to-front (BTF) display of voxelbased objects. IEEE Computer Graphics and Applications, 5(1):52–60, January/ February 1985. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Feibush:2000:VSA

[FGW00] Eliot Feibush, Nikhil Gagvani, and Daniel Williams. Visualization for situational awareness. IEEE Computer Graphics and Applications, 20(5):38–45, September/October 2000. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

computer.org/cga/cg2000/
g5038abs.htm;   http:/
/dlib.computer.org/cg/
books/cg2000/pdf/g5038.
pdf.

Fisher:1984:EWU

[FH84a] John M. Fisher and Joseph D. Halford. Engineering workstation: a user’s perspective. IEEE Computer Graphics and Applications, 4(8):25– 29, August 1984. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Flasck:1984:ALC

[FH84b] Richard A. Flasck and Scott H. Holmberg. Activematrix liquid crystal displays using amorphous silicon thin-film transistors. IEEE Computer Graphics and Applications, 4(4):19– 22, April 1984. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Farouki:1985:HGF

[FH85] Rida T. Farouki and John K. Hinds. A hierarchy of geometric forms. IEEE Computer Graphics and Applications, 5(5):51–78, May/ June 1985. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Fan:2019:PSB

[FH19] C. Fan and H. Hauser.

Personalized sketch-based brushing in scatterplots. IEEE Computer Graphics and Applications, 39(4): 28–39, July/August 2019. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Fan:2021:SBS

[FH21] Chaoran Fan and Helwig Hauser. On sketchbased selections from scatterplots using KDE, compared to Mahalanobis and CNN Brushing. IEEE Computer Graphics and Applications, 41(5):67–78, September/October 2021. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Fu:2021:PVG

[FHH^{+}21] Qian Fu, Ying He, Fei Hou, Qian Sun, Anxiang Zeng, Zhenchuan Huang, Juyong Zhang, and Yong-Jin Liu. Poisson Vector Graphics (PVG)-guided face color transfer in videos. IEEE Computer Graphics and Applications, 41(6):152–163, November/December 2021. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Frohlich:2006:ID

[FHKH06] Bernd Fröhlich, Jan Hochstrate, Alexander Kulik, and Anke Huckauf. On 3D input devices. IEEE Computer Graphics and Applications, 26(2):15–19, March/April 2006. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Fujimoto:1983:JIR

[FI83] Akira Fujimoto and Kansei Iwata. Jag-free images on raster displays. IEEE Computer Graphics and Applications, 3(9):26–34, December 1983. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Fuchs:2020:TCA

[FIB^{+}20] J. Fuchs, P. Isenberg, A. Bezerianos, M. Miller, and D. A. Keim. Teaching clustering algorithms with EduClust: Experience report and future directions. IEEE Computer Graphics and Applications, 40(2): 98–102, March/April 2020. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Figgins:2000:ACR

[Fig00] Stephen Figgins. About the cover: Renaissance revisited. IEEE Computer Graphics and Applications, 20(5):4–5, September/October 2000. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://dlib.
computer.org/cg/books/
cg2000/pdf/g5004.pdf.

Figgins:2001:ACJ

[Fig01a] Stephen Figgins. About the cover: Juggling the many layers of collage. IEEE Computer Graphics and Applications, 21(1):4–5, January/ February 2001. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://dlib.
computer.org/cg/books/
cg2001/pdf/g1004.pdf.

Figgins:2001:GCP

[Fig01b] Stephen Figgins. Graphics in computational paleontology. IEEE Computer Graphics and Applications, 21(6):6–13, November/December 2001. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http:

//dlib.computer.org/cg/
books/cg2001/g6006abs.
htm;   http://dlib.computer.
org/cg/books/cg2001/pdf/
g6006.pdf.

Figgins:2002:BC

[Fig02] Stephen Figgins. Biomechanics and the cyberhuman. IEEE Computer Graphics and Applications, 22(6):14–20, November/ December 2002. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL http://dlib.
                                        computer.org/cg/books/
                                        cg2002/pdf/g6014.pdf.

Fillmore:1984:NRM

[Fil84] Jay P. Fillmore. A note on rotation matrices. IEEE Computer Graphics and Applications, 4(2):30–33, February 1984. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Finkenzeller:2008:DBF

[Fin08] Dieter Finkenzeller. Detailed building facades. IEEE Computer Graphics and Applications, 28(3):58– 66, May/June 2008. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Fisher:2002:MIC

[Fis02] Hugh Fisher. Multicast issues for collaborative virtual environments. IEEE Computer Graphics and Applications, 22(5):68–75, September/October 2002. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL http://dlib.
                                        computer.org/cg/books/
                                        cg2002/pdf/g5068.pdf;
                                        http://www.computer.org/
                                        cga/cg2002/g5068abs.htm.

Fisher:2021:BSD

[Fis21] D. A. Fisher. BubbleUp: Supporting DevOps with data visualization. IEEE Computer Graphics and Applications, 41(1):99–105, January/February 2021. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Franklin:1987:ABC

[FJ87] Mark J. Franklin and Robert B. Jerard. Application briefs — CAD for architectural design of energyefficient homes. IEEE Computer Graphics and Applications, 7(5):8–13, May/ June 1987. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Fox:2000:IIA

[FJHW00] Armando Fox, Brad Johanson, Pat Hanrahan, and Terry Winograd. Integrating information appliances into an interactive workspace. IEEE Computer Graphics and Applications, 20(3):54– 65, May/June 2000. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

computer.org/cga/cg2000/
g3054abs.htm;   http:/
/dlib.computer.org/cg/
books/cg2000/pdf/g3054.
pdf.

Fekete:2019:PAS

[FJKTX19] J. Fekete, T. J. Jankun-Kelly, M. Tory, and K. Xu. Provenance analysis for sensemaking. IEEE Computer Graphics and Applications, 39(6):27–29, November/December 2019. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Firat:2022:PVV

[FJL22] Elif E. Firat, Alark Joshi, and Robert S. Laramee. Vis-LitE: Visualization literacy and evaluation. IEEE Computer Graphics and Applications, 42(3):99–107, May/ June 2022. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Freeman:2002:EBS

[FJP02] William T. Freeman, Thouis R. Jones, and Egon C. Pasztor. Example-based superresolution. IEEE Computer Graphics and Applications, 22(2):56–65, March/ April 2002. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

                                        computer.org/cga/cg2002/
                                        g2056abs.htm;   http:/
                                        /dlib.computer.org/cg/
                                        books/cg2002/pdf/g2056.
                                        pdf.

Foley:1987:ICL

[FK87] James D. Foley and Won Chul Kim. Image composition via lookup table manipulation. IEEE Computer Graphics and Applications, 7(11):

26–35, November/December 1987. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Fels:2005:SAP

[FKC^{+}05] Sidney Fels, Yuichiro Kinoshita, Tzu-Pei Grace Chen, Yasufumi Takama, Steve Yohanan, Sachiyo Takahashi, Ashley Gadd, and Kenji Funahashi. Swimming across the Pacific: a VR swimming interface. IEEE Computer Graphics and Applications, 25(1):24– 31, January/February 2005. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Figueroa:2010:CGP

[FKK^{+}10] Pablo Figueroa, Yoshifumi Kitamura, Sebastien Kuntz, Lode Vanacken, Steven Maesen, Tom De Weyer, Sofie Notelaers, Johanna Renny Octavia, Anastasiia Beznosyk, Karin Coninx, Felipe Bacim, Regis Kopper, Anamary Leal, Tao Ni, and Doug A. Bowman. 3DUI 2010 Contest grand prize winners. IEEE Computer Graphics and Applications, 30(6):86–96, c3, November/December 2010. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Fitzmaurice:2005:CMF

[FKKB05] George Fitzmaurice, Azam Khan, Gordon Kurtenbach, and Graham Binks. Cinematic meeting facilities using large displays. IEEE Computer Graphics and Applications, 25(4):17–21, July/August 2005. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Fullenwider:1981:CGP

[FL81] Donald R. Fullenwider and James P. Lefever. Computer graphics and the practice of architecture. IEEE Computer Graphics and Applications, 1(4):18–26, October 1981. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Funkhouser:2000:GEI

[FL00] Thomas Funkhouser and Kai Li. Guest Editors’ introduction: Large-format displays. IEEE Computer Graphics and Applications, 20(4):20–21, July/ August 2000. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL http://dlib.
                                        computer.org/cg/books/
                                        cg2000/pdf/g4020.pdf.

Froehlich:2012:NDS

[FL12] Bernd Froehlich and Mark A. Livingston. A new department: Spatial interfaces. IEEE Computer Graphics and Applications, 32(5):3, September/October 2012. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Forsberg:1997:PVS

[FLMZ97] Andrew S. Forsberg, Joseph J.

LaViola, Jr., Lee Markosian, and Robert C. Zeleznik. Projects in VR: Seamless interaction in virtual reality. IEEE Computer Graphics and Applications, 17(6):6–9, November/ December 1997. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://dlib.
computer.org/cg/books/
cg1997/pdf/g6006.pdf.

Foley:1990:VFS

[FLNR90] Thomas A. Foley, David A.

Lane, Gregory M. Nielson, and Ramamani Ramaraj. Visualizing functions over a sphere. IEEE Computer Graphics and Applications, 10(1):32–40, January/ February 1990. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Fuhrmann:1998:CVA

[FLSG98] Anton Fuhrmann, Helwig Lösffelmann, Dieter Schmalstieg, and Michael Gervautz. Collaborative visualization in augmented reality. IEEE Computer Graphics and Applications, 18(4):54– 59, July/August 1998. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

                                        computer.org/cga/cg1998/
                                        g4054abs.htm;   http:/
                                        /dlib.computer.org/cg/
                                        books/cg1998/pdf/g4054.
                                        pdf.

Field:1982:QMD

[FM82] David A. Field and Alexander P. Morgan. A quick method for determining whether a second-degree polynomial has solutions in a given box. IEEE Computer Graphics and Applications, 2(3):65–66, 68, May/ June 1982. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Fournier:1985:FBA

[FM85] Alain Fournier and Thomas Milligan. Frame buffer algorithms for stochastic models. IEEE Computer Graphics and Applications, 5(10):40–46, October 1985. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Foley:1986:DPV

[FM86] James D. Foley and Charles F. McMath. Dynamic process visualization. IEEE Computer Graphics and Applications, 6(3):16–25, March/ April 1986. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Fletcher:1987:ATM

[FM87] G. Yates Fletcher and David F. McAllister. An analysis of tension methods for convexity-preserving interpolation. IEEE Computer Graphics and Applications, 7(8):7–14, August 1987. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Fletcher:1989:TPS

[FM89a] G. Yates Fletcher and David F. McAllister. A tension-compatible patch for shape-preserving surface interpolation. IEEE Computer Graphics and Applications, 9(3):45–55, May/ June 1989. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Fowell:1989:FPF

[FM89b] Richard A. Fowell and David D. McNeil. Faster plots by fan data-compression. IEEE Computer Graphics and Applications, 9(2):58– 66, March/April 1989. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Fletcher:1990:ATA

[FM90] Yates Fletcher and David F.

McAllister. Automatic tension adjustment for interpolatory splines. IEEE Computer Graphics and Applications, 10(1):10–17, January/ February 1990. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Finkelstein:2003:GEI

[FM03] Adam Finkelstein and Lee Markosian. Guest Editors’ introduction: Nonphotorealistic rendering. IEEE Computer Graphics and Applications, 23(4):26–27, July/August 2003. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cg/2003/04/g4026.pdf;
http://csdl.computer.org/
dl/mags/cg/2003/04/g4026.
htm.

Fukusato:2022:VDD

[FM22] Tsukasa Fukusato and Akinobu Maejima. Viewdependent deformation for 2.5-D cartoon models. IEEE Computer Graphics and Applications, 42(5):66–75, September/October 2022. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Feierl:2023:SVF

[FML23] Lukas Feierl, Torsten Möller, and Peter Luidolt. Sun-Screen: Visual fault detection for solar–thermal systems. IEEE Computer Graphics and Applications, 43(6):87–100, November/ December 2023. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Fabrikant:2006:DSM

[FMM06] Sara Irina Fabrikant, Daniel R.

Montello, and David M. Mark. The distancesimilarity metaphor in regiondisplay spatializations. IEEE Computer Graphics and Applications, 26(4):34–44, July/August 2006. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Forney:2003:UFS

[FMMS03] Glenn P. Forney, Daniel Madrzykowski, Kevin B. McGrattan, and Laurel Sheppard. Understanding fire and smoke flow through modeling and visualization. IEEE Computer Graphics and Applications, 23(4):6– 13, July/August 2003. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

csdl.computer.org/comp/
mags/cg/2003/04/extras/
g4006.htm;   http://csdl.
computer.org/comp/mags/
cg/2003/04/g4006abs.htm;
 http://csdl.computer.
org/dl/mags/cg/2003/04/
g4006.htm;   http://csdl.

                                        computer.org/dl/mags/cg/
                                        2003/04/g4006.pdf.

Field:2007:FTV

[FOI^{+}07] Blayne Field, Sean O’Neill, Victoria Interrante, Thomas W. Jones, and Timothy Urness. FieldVis: a tool for visualizing astrophysical magnetohydrodynamic flow. IEEE Computer Graphics and Applications, 27(1):9– 13, January/February 2007. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Foley:2000:GTT

[Fol00] Jim Foley. Getting there: The ten top problems left. IEEE Computer Graphics and Applications, 20(1):66– 68, January/February 2000. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL

                                        http://dlib.computer.org/
                                        cg/books/cg2000/pdf/g1066.
                                        pdf.

Foley:2014:BB

[Fol14] Jim Foley. The best and brightest. IEEE Computer Graphics and Applications, 34(4):6–7, July/ August 2014. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Foxlin:2005:PTS

[Fox05]  Eric Foxlin.  Pedestrian

tracking with shoe-mounted inertial sensors. IEEE Computer Graphics and Applications, 25(6):38–46, November/December 2005. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Fang:1993:DTU

[FP93] Tsung-Pao Fang and Les A.

Piegl. Delaunay triangulation using a uniform grid. IEEE Computer Graphics and Applications, 13(3):36– 47, May/June 1993. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Fang:1995:DTT

[FP95] Tsung-Pao Fang and Les A.

Piegl. Delaunay triangulation in three dimensions. IEEE Computer Graphics and Applications, 15(5):62– 69, September 1995. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Francois:2008:STM

[FPBB08] Guillaume François, Sumanta Pattanaik, Kadi Bouatouch, and Gaspard Breton. Subsurface texture mapping. IEEE Computer Graphics and Applications, 28(1):34– 42, January/February 2008. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Forsberg:2006:AIF

[FPH^{+}06] Andrew Forsberg, Prabhat, Graff Haley, Andrew Bragdon, Joseph Levy, Caleb I. Fassett, David Shean, James W. Head III, Sarah Milkovich, and Mark Duchaineau. Adviser: Immersive field work for planetary geoscientists. IEEE Computer Graphics and Applications, 26(4):46–54, July/August 2006. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Fujimoto:1984:GDS

[FPI84] Akira Fujimoto, Christopher G. Perrott, and Kansei Iwata. A 3-D graphics display system with depth buffer and pipeline processor. IEEE Computer Graphics and Applications, 4(6): 11–23, June 1984. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Farbood:2004:HGS

[FPJ04] Morwaread M. Farbood, Egon Pasztor, and Kevin Jennings. Hyperscore: a graphical sketchpad for novice composers. IEEE Computer Graphics and Applications, 24(1):50–54, January/February 2004. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

csdl.computer.org/comp/

mags/cg/2004/01/g1050abs.
htm;   http://csdl.computer.
org/dl/mags/cg/2004/01/
g1050.pdf.

Fernondez-Prieto:2017:SWW

[FPNVHH17] Diana Fernondez-Prieto, Carol Naranjo-Valero, Jose Tiberio Hernandez, and Hans Hagen. STRAD Wheel: Webbased library for visualizing temporal data. IEEE Computer Graphics and Applications, 37(2):99–105, March/ April 2017. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2017/02/mcg2017020099-
abs.html.

Fuchs:1982:ATD

[FPT^{+}82] Henry Fuchs, Stephen M.

Pizer, Li Ching Tsai, Sandra H. Bloomberg, and E. Ralph Heinz. Adding a true 3-D display to a raster graphics system. IEEE Computer Graphics and Applications, 2(7):73–78, September 1982. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Frohlich:2000:PVC

[FPW^{+}00] Bernd Fröhlich, John Plate, Jürgen Wind, Gerold Wesche, and Margin Göbel. Projects in VR: Cubic-mouse-based interaction in virtual environments. IEEE Computer Graphics and Applications, 20(4):12–15, July/ August 2000. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL http://dlib.
                                        computer.org/cg/books/
                                        cg2000/pdf/g4012.pdf.

Fawcett:2000:ARR

[FR00] John Fawcett and Peter Robinson. Adaptive routing for road traffic. IEEE Computer Graphics and Applications, 20(3):46–53, May/ June 2000. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

                                        computer.org/cga/cg2000/
                                        g3046abs.htm;   http:/
                                        /dlib.computer.org/cg/
                                        books/cg2000/pdf/g3046.
                                        pdf.

Fredricks:1984:AMD

[Fre84] David A. Fredricks. Algebraic method to determine if a point is on a spline. IEEE Computer Graphics and Applications, 4(2):52– 53, February 1984. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Freund:1986:IPC

[Fre86] Dwight D. Freund. An interactive procedure for constructing line and circle tangencies. IEEE Computer Graphics and Applications, 6(4):59–63, April 1986. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Friston:2020:LLR

[Fri20] S. Friston. Low-latency rendering with dataflow architectures. IEEE Computer Graphics and Applications, 40(3):94–104, May/ June 2020. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Frye:1983:CSM

[Fry83] Norma Tanguay Frye. Computer slide making in a media services department. IEEE Computer Graphics and Applications, 3(4):47– 52, July/August 1983. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Farhoosh:1986:CHM

[FS86] Hamid Farhoosh and Gunther Schrack. CNS-HLS mapping using fuzzy sets. IEEE Computer Graphics and Applications, 6(6):28– 35, June 1986. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Farin:1989:CFC

[FS89] Gerald Farin and Nickolas Sapidis. Curvature and the fairness of curves and surfaces. IEEE Computer Graphics and Applications, 9(2):52–57, March/ April 1989. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Frederick:1990:CIW

[FS90] Carl Frederick and Eric L. Schwartz. Conformal image warping. IEEE Computer Graphics and Applications, 10(2):54–61, March/ April 1990. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Farias:2001:CRL

[FS01] Ricardo Farias and Cláudio T. Silva. Out-of-core rendering of large, unstructured grids. IEEE Computer Graphics and Applications, 21(4):42– 50, July/August 2001. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

computer.org/cga/cg2001/
g4042abs.htm;   http:/
/dlib.computer.org/cg/
books/cg2001/pdf/g4042.
pdf.

Fu:2023:ICA

[FS23] Chi-Wing Fu and André Stork. 2022/2023 IEEE CG&A awards. IEEE Computer Graphics and Applications, 43(6):141–142, November/December 2023. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Feiner:1982:DAD

[FSB82a] S. Feiner, D. Salesin, and T. Banchoff. Dial: a diagramatic animation language. IEEE Computer Graphics and Applications, 2(9):43–54, September 1982. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Feiner:1982:DDA

[FSB82b] Steven Feiner, David Salesin, and Thomas Banchoff. DIAL: a diagrammatic animation language. IEEE Computer Graphics and Applications, 2(7):43–46, 48– 54, September 1982. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Fellner:2007:GEI

[FSK07] Dieter W. Fellner, Dietmar Saupe, and Harald Krottmaier. Guest Editors’ introduction: 3D documents. IEEE Computer Graphics and Applications, 27(4):20–21, July/ August 2007. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cg/2007/04/g4020.pdf.

Faisstnauer:2000:PSN

[FSP00] Chris Faisstnauer, Dieter Schmalstieg, and Werner Purgathofer. Priority scheduling for networked virtual environments. IEEE Computer Graphics and Applications, 20(6):66–75, November/December 2000. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL

http://computer.org/cga/
cg2000/g6066abs.htm;
http://dlib.computer.org/
cg/books/cg2000/pdf/g6066.
pdf.

Feiner:2000:GEI

[FT00a] Steven Feiner and Daniel Thalmann. Guest Editors’ introduction. IEEE Computer Graphics and Applications, 20(6):24–25, November/December 2000. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http:

//dlib.computer.org/cg/
books/cg2000/pdf/g6024.
pdf.

Friesen:2000:RHP

[FT00b] Jerrold A. Friesen and Thomas D. Tarman. Remote high-performance visualization and collaboration. IEEE Computer Graphics and Applications, 20(4):45– 49, July/August 2000. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

computer.org/cga/cg2000/
g4045abs.htm;   http:/
/dlib.computer.org/cg/

books/cg2000/pdf/g4045.
pdf.

Froese:2016:LLD

[FT16] Maria-Elena Froese and Melanie Tory. Lessons learned from designing visualization dashboards. IEEE Computer Graphics and Applications, 36(2):83–89, March/April 2016. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2016/02/mcg2016020083-
abs.html.

Fujishiro:2000:VDM

[FTAT00] Issei Fujishiro, Yuriko Takeshima, Taeko Azuma, and Shigeo Takahashi. Volume data mining using 3D field topology analysis. IEEE Computer Graphics and Applications, 20(5):46–51, September/October 2000. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

computer.org/cga/cg2000/
g5046abs.htm;   http:/
/dlib.computer.org/cg/
books/cg2000/pdf/g5046.
pdf.

Fujimoto:1986:AAR

[FTI86] Akira Fujimoto, Takayuki Tanaka, and Kansei Iwata. Arts: Accelerated raytracing system. IEEE Computer Graphics and Applications, 6(4):16–26, April 1986. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Fu:2024:TTT

[Fu24] Chi Wing Fu. The test of time (ToT) awards. IEEE Computer Graphics and Applications, 44(2):117–118, March/April 2024. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Funfzig:2006:HSD

[FUF06] Christoph Fünfzig, Torsten Ullrich, and Dieter W. Fellner. Hierarchical spherical distance fields for collision detection. IEEE Computer Graphics and Applications, 26(1):64–74, January/ February 2006. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Foley:1984:HFC

[FWC84] James D. Foley, Victor L. Wallace, and Peggy Chan. The human factors of computer graphics interaction techniques. IEEE Computer Graphics and Applications, 4(11):13–48, November/December 1984. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Franklin:1986:PGP

[FWSN86] Wm. Randolph Franklin, Peter Y. F. Wu, Sumitro Samaddar, and Margaret Nichols. Prolog and geometry projects. IEEE Computer Graphics and Applications, 6(11):46–55, November/December 1986. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Flanagan:1997:ACP

[FY97] William P. Flanagan and Rogbin Yeary. About the cover: Presto Studios: a group of journeymen. IEEE Computer Graphics and Applications, 17(3):4–5, May/ June 1997. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Fyock:1997:UVM

[Fyo97] Daniel E. Fyock. Using visualization to maintain large computer systems. IEEE Computer Graphics and Applications, 17(4):73–74, July/August 1997. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Farrell:1985:ACI

[FYZ85] Edward J. Farrell, Wen C.

Yang, and Rosario A. Zappulla. Animating 3D CT imaging. IEEE Computer Graphics and Applications, 5(12):26–32, December 1985. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Fruh:2003:CCM

[FZ03] Christian Früh and Avideh Zakhor. Constructing 3D city models by merging aerial and ground views. IEEE Computer Graphics and Applications, 23(6):52–61, November/ December 2003. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http:

//csdl.computer.org/dl/
mags/cg/2003/06/g6052.
pdf;   http://csdl.computer.
org/dl/mags/cg/2003/06/
g6052abs.htm.

Fischer:2013:BIU

[FZH^{+}13] Patrick Tobias Fischer, Christian Zollner, Thilo Hoffmann, Sebastian Piatza, and Eva Hornecker. Beyond information and utility: Transforming public spaces with media facades. IEEE Computer Graphics and Applications, 33(2): 38–46, March/April 2013. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Farrell:1984:CIN

[FZY84] Edward J. Farrell, Rosario Zappulla, and Wen C. Yang. Color 3-D imaging of normal and pathologic intracranial structures. IEEE Computer Graphics and Applications, 4(9):5–17, September 1984. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Gardner:1997:SBR

[GA97] Michael T. Gardner and Philip Amburn. Simulationbased remote debriefing for red flag missions. IEEE Computer Graphics and Applications, 17(5):30–39, September/October 1997. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL

http://dlib.computer.org/
cg/books/cg1997/pdf/g5030.
pdf.

Gardner:2018:CII

[GAB^{+}18] Adam Gardner, Ludovic Autin, Brett Barbaro, Arthur J. Olson, and David S. Goodsell. CellPAINT: Interactive illustration of dynamic mesoscale cellular environments. IEEE Computer Graphics and Applications, 38(6):51–66, ???? 2018. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL https:

//ieeexplore.ieee.org/
document/8617734/.

Graebling:2024:FLI

[GAe^{+}24] Nico Graebling, Melanie Althaus, Özgür Ozan en, Thomas Reimann, Tuanny Cajuhi, Gerik Scheuermann, Olaf Kolditz, and Karsten Rink. Feels like an indie game evaluation of a virtual field trip prototype on radioactive waste management research for university education. IEEE Computer Graphics and Applications, 44(1):13–24, January/ February 2024. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Gaither:2004:VRA

[Gai04]  Kelly Gaither.  Visual-

ization’s role in analyzing computational fluid dynamics data. IEEE Computer Graphics and Applications, 24(3):13–15, May/ June 2004. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http:/

/csdl.computer.org/dl/
mags/cg/2004/03/g3013.
htm;   http://csdl.computer.
org/dl/mags/cg/2004/03/
g3013.pdf.

Gaither:2017:HVC

[Gai17]  Kelly Gaither.  How

visualization can foster diversity and inclusion in next-generation science. IEEE Computer Graphics and Applications, 37 (5):106–112, September/ October 2017. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/

cg/2017/05/mcg2017050106-
abs.html.

Galton:1989:ETA

[Gal89] Ian Galton. An efficient three-point arc algorithm. IEEE Computer Graphics and Applications, 9(6):44– 49, November/December 1989. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Garcia:2015:TCI

[Gar15] Francisco Gonzalez Garcia.

Trends in continuity and interpolation for computer graphics. IEEE Computer Graphics and Applications, 35(6):76–82, November/ December 2015. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
cg/2015/06/mcg2015060076-
abs.html.

Gastal:2016:EHD

[Gas16] Eduardo Simoes Lopes Gastal. Efficient highdimensional, edge-aware filtering. IEEE Computer Graphics and Applications, 36(6):86–95, November/ December 2016. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2016/06/mcg2016060086-
abs.html.

Garcia-Alonso:1994:SCD

[GASF94] Alejandro Garcia-Alonso, Nicolas Serrano, and Juan Flaquer. Solving the collision detection problem. IEEE Computer Graphics and Applications, 14(3):36– 43, May/June 1994. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Grant:1986:ECM

[GAW86] Eric Grant, Phil Amburn, and Turner Whitted. Exploiting classes in modeling and display software. IEEE Computer Graphics and Applications, 6(11):13–20, November/December 1986. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Gershon:1996:CGV

[GB96a] Nahum Gershon and Judith R. Brown. Computer graphics and visualization in the global information infrastructure. IEEE Computer Graphics and Applications, 16(2):60–61, March/ April 1996. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Gershon:1996:RCG

[GB96b] Nahum Gershon and Judith R. Brown. The role of computer graphics and visualization in the GII.

IEEE Computer Graphics and Applications, 16(2): 61–63, March/April 1996. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Gotz:2016:DDH

[GB16] David Gotz and David Borland. Data-driven healthcare: Challenges and opportunities for interactive visualization. IEEE Computer Graphics and Applications, 36(3):90–96, May/ June 2016. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2016/03/mcg2016030090-
abs.html.

Gujar:1993:TGF

[GBCK93] Uday G. Gujar, Virendra C. Bhavsar, Stephen Y. M. Choi, and Prem K. Kalra. Traversed geometric fractals. IEEE Computer Graphics and Applications, 13(5):61–67, September 1993. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Godin:2002:AOI

[GBT^{+}02] Guy Godin, J.-Angelo Beraldin, John Taylor, Luc Cournoyer, Marc Rioux, Sabry El-Hakim, Réjean Baribeau, François Blais, Pierre Boulanger, Jacques Domey, and Michel Picard. Active optical 3D imaging for heritage applications. IEEE Computer Graphics and Applications, 22(5):24–36, September/October 2002. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL http://dlib.
                                        computer.org/cg/books/
                                        cg2002/pdf/g5024.pdf;
                                        http://www.computer.org/
                                        cga/cg2002/g5024abs.htm.

Gordon:1991:FDB

[GC91] Dan Gordon and Shuhong Chen. Front-to-back display of BSP trees. IEEE Computer Graphics and Applications, 11(5):79–85, September 1991. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Gortler:1994:RRM

[GCS94] Steven Gortler, Michael F. Cohen, and Philipp Slusallek. Radiosity and relaxation methods. IEEE Computer Graphics and Applications, 14(6):48–58, November/December 1994. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Gaither:2012:VAF

[GCS^{+}12] Kelly P. Gaither, Hank Childs, Karl W. Schulz, Cyrus Harrison, William Barth, Diego Donzis, and Pui-Kuen Yeung. Visual analytics for finding critical structures in massive time-varying turbulent-flow simulations. IEEE Computer Graphics and Applications, 32(4):34–45, July/ August 2012. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Gunther:1993:HSC

[GD93] Oliver Gunther and Salvador Dominguez. Hierarchical schemes for curve representation. IEEE Computer Graphics and Applications, 13(3):55–63, May/ June 1993. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Gu:2011:CAM

[GD11] Qin Gu and Zhigang Deng.

Context-aware motion diversification for crowd simulation. IEEE Computer Graphics and Applications, 31(5):54–65, September/ October 2011. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Gu:2013:GFG

[GD13] Qin Gu and Zhigang Deng.

Generating freestyle group formations in agent-based crowd simulations. IEEE Computer Graphics and Applications, 33(1):20–31, January/February 2013. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Goncalves:2014:ILC

[GDF^{+}14] Vitor Goncalves, Paulo Dias, Maria Joao Fontoura, Rui Moura, and Beatriz Sousa Santos. Investigating landfill contamination by visualizing geophysical data. IEEE Computer Graphics and Applications, 34(1):16–21, January/ February 2014. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Gune:2018:GHE

[GDS^{+}18] Aditya Gune, Raffaele De Amicis, Bruno Simoes, Christopher A. Sanchez, and H. Onan Demirel. Graphically hearing: Enhancing understanding of geospatial data through an integrated auditory and visual experience. IEEE Computer Graphics and Applications, 38(4):18–26, July/ August 2018. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        cg/2018/04/mcg2018040018-
                                        abs.html.

Guan:2011:RTC

[GDY^{+}11] Tao Guan, Liya Duan, Junqing Yu, Yongjian Chen, and Xu Zhang. Real-time camera pose estimation for wide-area augmented reality applications. IEEE Computer Graphics and Applications, 31(3):56–68, May/ June 2011. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Gershon:1997:GEI

[GE97a] Nahum Gershon and Stephen G.

Eick. Guest Editors’ introduction: Information visualization. IEEE Computer Graphics and Applications, 17(4):29–31, July/ August 1997. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Gershon:1997:IVA

[GE97b] Nahum Gershon and Stephen G.

Eick. Information visualization applications in the real world. IEEE Computer Graphics and Applications, 17(4):66, July/August 1997. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Gershon:1998:GEI

[GE98] Nahum Gershon and Stephen G.

Eick. Guest Editors’ introduction: Scaling to new heights. IEEE Computer Graphics and Applications, 18(4):16–17, July/August 1998. CODEN ICGADZ.

ISSN 0272-1716 (print), 1558-1756 (electronic).

Geller:2006:ITE

[Gel06] Tom Geller. Interactive tabletop exhibits in museums and galleries. IEEE Computer Graphics and Applications, 26(5):6–11, September/October 2006. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Geller:2007:EWM

[Gel07a]  Tom Geller.  Envision-

ing the wind: Meteorology graphics at Weather Underground. IEEE Computer Graphics and Applications, 27(5):92–97, September/October 2007. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Geller:2007:IWS

[Gel07b] Tom Geller. Imaging the world: The state of online mapping. IEEE Computer Graphics and Applications, 27(2):8–13, March/ April 2007. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Geller:2008:OUV

[Gel08] Tom Geller. Overcoming the uncanny valley. IEEE Computer Graphics and Applications, 28(4):11–17, July/ August 2008. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Gershon:1991:VBV

[Ger91] Nahum D. Gershon. Visualization blackboard — visualizing 3D PET images. IEEE Computer Graphics and Applications, 11(5):11– 13, September 1991. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Gershon:1996:MHT

[Ger96] Nahum Gershon. Moving happily through the World Wide Web. IEEE Computer Graphics and Applications, 16(2):72–75, March/ April 1996. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Gershon:1998:SNV

[Ger98] Nahum Gershon. Short note: Visualization of an imperfect world. IEEE Computer Graphics and Applications, 18(4):43–45, July/ August 1998. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

computer.org/cga/cg1998/
g4043abs.htm.

Goldfeather:1986:QSR

[GF86] Jack Goldfeather and Henry Fuchs. Quadratic surface rendering on a logicenhanced frame-buffer memory. IEEE Computer Graphics and Applications, 6(1): 48–59, January/February 1986. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Geller:1994:GVT

[GF94] M. J. Geller and E. E. Falco. Graphic voyages through the universe. IEEE Computer Graphics and Applications, 14(6):7–11, November/December 1994. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Goris:1987:CPC

[GFB87] Andy Goris, Bob Fredrickson, and Harold L. Baeverstad, Jr. A configurable pixel cache for fast image generation. IEEE Computer Graphics and Applications, 7(3):24–32, March/ April 1987. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Groh:2016:WRT

[GFW^{+}16] Benjamin H. Groh, Jasmin Flaschka, Markus Wirth, Thomas Kautz, Martin Fleckenstein, and Bjoern M. Eskofier. Wearable realtime skateboard trick visualization and its community perception. IEEE Computer Graphics and Applications, 36(5):12–18, September/October 2016. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2016/05/mcg2016050012-
abs.html.

Gourley:1983:PRC

[GG83] Kevin D. Gourley and Douglas M. Green. A polygonto-rectangle conversion algorithm. IEEE Computer Graphics and Applications, 3(1):31–32, 34–36, January/ February 1983. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Grabowski:1987:IMC

[GG87] H. Grabowski and R. Glatz.

IGES model comparison system: a tool for testing and validating IGES processes. IEEE Computer Graphics and Applications, 7(11): 47–57, November/December 1987. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Georgiades:1992:LMG

[GG92] Priamos N. Georgiades and Donald P. Greenberg. Locally manipulating the geometry of curved surfaces. IEEE Computer Graphics and Applications, 12(1):54– 64, January/February 1992. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Garrison:2023:CAB

[GGB23] Laura A. Garrison, David S. Goodsell, and Stefan Bruckner. Changing aesthetics in biomolecular graphics. IEEE Computer Graphics and Applications, 43 (3):94–101, May/June 2023. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Goel:2008:AVG

[GGG08] Vikash Ravi Goel, Roy K. Greenberg, and Donald P. Greenberg. Automated vascular geometric analysis of aortic aneurysms. IEEE Computer Graphics and Applications, 28(3):76–86, May/June 2008. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Gralka:2018:ISV

[GGS^{+}18] Patrick Gralka, Sebastian Grottel, Joachim Staib, Karsten Schatz, Grzegorz Karch, Manuel Hirschler, Michael Krone, Guido Reina, Stefan Gumhold, and Thomas Ertl. 2016 IEEE Scientific Visualization Contest winner: Visual and structural analysis of point-based simulation ensembles. IEEE Computer Graphics and Applications, 38(3):106–117, May/ June 2018. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2018/03/mcg2018030106-
abs.html.

Gospodnetic:2022:SIV

[GGS22] Petra Gospodnetic, Christina Gillmann, and Gerik Scheuermann. Special issue on visualization in manufacturing. IEEE Computer Graphics and Applications, 42(2):8–9, March/April 2022. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Guttag:1992:SMM

[GGV92] Karl Guttag, Robert J.

Gove, and Jerry R. Van Aken. A single-chip multiprocessor for multimedia: The MVP. IEEE Computer Graphics and Applications, 12(6):53–64, November/December 1992. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Greene:1986:CRO

[GH86] Ned Greene and Paul S.

Heckbert. Creating raster Omnimax images from multiple perspective views using the elliptical weighted average filter. IEEE Computer Graphics and Applications, 6(6):21–27, June 1986. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Green:1997:GEI

[GH97] Mark Green and Larry Hodges. Guest Editors’ introduction: VRAIS 97 highlights the maturing of a discipline. IEEE Computer Graphics and Applications, 17(6):21–22, November/December 1997. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http:

                                        //dlib.computer.org/cg/
                                        books/cg1997/pdf/g6021.
                                        pdf.

Ghani:2012:SRC

[Gha12] Dahlan Bin Abdul Ghani. Seri Rama: Converting a shadow play puppet to street fighter. IEEE Computer Graphics and Applications, 32(1):8–11, January/ February 2012. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Gu:2021:RUF

[GHCW21] Pengfei Gu, Jun Han, Danny Z. Chen, and Chaoli Wang. Reconstructing unsteady flow data from representative streamlines via diffusion and deep-learning-based denoising. IEEE Computer Graphics and Applications, 41(6):111–121, November/December 2021. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Guo:2004:SFD

[GHQ04a] Xiaohu Guo, Jing Hua, and Hong Qin. Scalar-function-driven editing on point set surfaces. IEEE Computer Graphics and Applications, 24(4):43–52, July/ August 2004. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http:/

/csdl.computer.org/dl/
mags/cg/2004/04/g4043.
htm;   http://csdl.computer.
org/dl/mags/cg/2004/04/
g4043.pdf.

Guo:2004:TBH

[GHQ04b] Xiaohu Guo, Jing Hua, and Hong Qin. Touchbased haptics for interactive editing on point set surfaces. IEEE Computer Graphics and Applications, 24(6):31–39, November/December 2004. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http:

//csdl.computer.org/dl/
mags/cg/2004/06/g6031.
htm;   http://csdl.computer.
org/dl/mags/cg/2004/06/
g6031.pdf.

Goebel:2001:GEI

[GHR01] Martin Goebel, Michitaka Hirose, and Lawrence Rosenblum. Guest Editors’ introduction: Today’s VR. IEEE Computer Graphics and Applications, 21(6):22–24, November/ December 2001. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL http://dlib.
                                        computer.org/cg/books/
                                        cg2001/g6022abs.htm;
                                        http://dlib.computer.org/
                                        cg/books/cg2001/pdf/g6022.
                                        pdf.

Gabbard:1999:UCD

[GHS99] Joseph L. Gabbard, Deborah Hix, and J. Edward Swan II. Usercentered design and evaluation of virtual environments. IEEE Computer Graphics and Applications, 19(6):51–59, November/ December 1999. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

                                        computer.org/cga/cg1999/
                                        g6051abs.htm;   http:/
                                        /dlib.computer.org/cg/
                                        books/cg1999/pdf/g6051.
                                        pdf.

Graham:1994:DTI

[GI94] Phil Graham and S. Sitharama Iyengar. Double- and triplestep incremental linear interpolation. IEEE Computer Graphics and Applications, 14(3):49–53, May/ June 1994. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Giertsen:1992:VVS

[Gie92] Christopher Giertsen. Volume visualization of sparse irregular meshes. IEEE Computer Graphics and Applications, 12(2):40–48, March/April 1992. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Gill:1994:NIS

[Gil94] Graeme W. Gill. N-step incremental straight-line algorithms. IEEE Computer Graphics and Applications, 14(3):66–72, May/ June 1994. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Girard:1987:IDC

[Gir87] Michael Girard. Interactive design of 3D computeranimated legged animal motion. IEEE Computer Graphics and Applications, 7(6):39–51, June 1987. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Ghosh:1993:AGS

[GJ93] Pijush K. Ghosh and Pradeep K. Jain. An algebra of geometric shapes. IEEE Computer Graphics and Applications, 13(5):50– 59, September 1993. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Gutierrez:2013:SAM

[GJ13] Diego Gutierrez and Wojciech Jarosz. Scattering: Acquisition, modeling, and rendering. IEEE Computer Graphics and Applications, 33(3):32–33, May/ June 2013. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Ghilardi:2018:CLL

[GJM18] Marcelo Cabral Ghilardi, Julio Jacques Junior, and Isabel Manssour. Crosswalk localization from low resolution satellite images to assist visually impaired people. IEEE Computer Graphics and Applications, 38(1):30–46, January/ February 2018. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        cg/2018/01/mcg2018010030-
                                        abs.html.

Goebel:1986:MGW

[GK86] M. Goebel and D. Kroemker. A multi-microprocessor GKS workstation. IEEE Computer Graphics and Applications, 6(7):54–60, July/ August 1986. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Gershon:2005:LF

[GK05] Nahum Gershon and Jake Kolojejchick. From the lab to the field. IEEE Computer Graphics and Applications, 25(5):6–7, September/October 2005. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cg/2005/05/g5006.pdf.

Grimes:1989:IIP

[GKB89] Jack D. Grimes, Les Kohn, and Rajeev Bharadhwaj. The Intel i860 64-bit processor: a general-purpose CPU with 3D graphics capabilities. IEEE Computer Graphics and Applications, 9(4):85–94, July/ August 1989. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Gemballa:1982:MBT

[GL82] Renate Gemballa and Rolf Lindner. The multiple-write bus technique. IEEE Computer Graphics and Applications, 2(7):33–34, 36–41, September 1982. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Glassner:1984:SSF

[Gla84] Andrew S. Glassner. Space subdivision for fast ray tracing. IEEE Computer Graphics and Applications, 4(10):15–22, October 1984. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Glassner:1988:SRT

[Gla88a] Andrew S. Glassner. Spacetime ray tracing for animation. IEEE Computer Graphics and Applications, 8(2):60–70, March/ April 1988. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL http://www.
                                        computer.org/csdl/mags/
                                        cg/02/00000504-abs.html;
                                         http://www.computer.org/
                                        csdl/mags/cg/02/mcg1988020060-
                                        abs.html.

Glassner:1988:TD

[Gla88b] Andrew S. Glassner. The traveling dinosaur. IEEE Computer Graphics and Applications, 8(3):6–9, May/ June 1988. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Glassner:1989:HDS

[Gla89] Andrew S. Glassner. How to derive a spectrum from an RGB triplet. IEEE Computer Graphics and Applications, 9(4):95–99, July/ August 1989. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Glassner:1992:GST

[Gla92] Andrew S. Glassner. Geometric substitution: a tutorial. IEEE Computer Graphics and Applications, 12(1): 22–36, January/February 1992. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Glassner:1996:AGNc

[Gla96a]  A. Glassner.  Andrew

Glassner’s notebook: Frieze groups. IEEE Computer Graphics and Applications, 16(3):78–83, May/ June 1996. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Glassner:1996:AGNb

[Gla96b]  Andrew Glassner.  An-

drew Glassner’s notebook: Computer-generated solar halos and sun dogs. IEEE Computer Graphics and Applications, 16(2):77–81, March/April 1996. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Glassner:1996:AGNe

[Gla96c] Andrew Glassner. Andrew Glassner’s notebook: More origami solids. IEEE Computer Graphics and Applications, 16(5):81–85, September 1996. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Glassner:1996:AGNd

[Gla96d]  Andrew Glassner.  An-

drew Glassner’s notebook: Origami platonic solids. IEEE Computer Graphics and Applications, 16(4): 85–91, July/August 1996. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Glassner:1996:AGNa

[Gla96e] Andrew Glassner. Andrew Glassner’s notebook: Solar halos and sun dogs. IEEE Computer Graphics and Applications, 16(1):83– 87, January/February 1996. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Glassner:1997:AGNa

[Gla97a] Andrew Glassner. Andrew Glassner’s notebook — going the distance. IEEE Computer Graphics and Applications, 17(1):78–84, January/February 1997. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Glassner:1997:AGNf

[Gla97b] Andrew Glassner. Andrew Glassner’s notebook: Inside moiré patterns. IEEE Computer Graphics and Applications, 17(6):97–101, November/December 1997. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Glassner:1997:AGNd

[Gla97c]  Andrew Glassner.  An-

drew Glassner’s notebook: Net results. IEEE Computer Graphics and Applications, 17(4):85–89, July/ August 1997. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Glassner:1997:AGNc

[Gla97d] Andrew Glassner. Andrew Glassner’s notebook: Signs of significance. IEEE Computer Graphics and Applications, 17(3):79–82, May/ June 1997. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Glassner:1997:AGNb

[Gla97e] Andrew Glassner. Andrew Glassner’s notebook: Situation normal. IEEE Computer Graphics and Applications, 17(2):83–87, March/ April 1997. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Glassner:1997:AGNe

[Gla97f] Andrew Glassner. Andrew Glassner’s notebook: The perils of problematic parameterization. IEEE Computer Graphics and Applications, 17(5):78–83, September/October 1997. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Glassner:1997:SS

[Gla97g] Andrew S. Glassner. A shape synthesizer. IEEE Computer Graphics and Applications, 17(3):40–51, May/June 1997. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Glassner:1998:AGNc

[Gla98a] Andrew Glassner. Andrew Glassner’s notebook: Aperiodic tiling. IEEE Computer Graphics and Applications, 18(3):83–90, May/ June 1998. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL http://dlib.
                                        computer.org/cg/books/
                                        cg1998/pdf/g3083.pdf.

Glassner:1998:AGNb

[Gla98b] Andrew Glassner. Andrew Glassner’s notebook: Circular reasoning. IEEE Computer Graphics and Applications, 18(2):104–108, March/April 1998. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL http://dlib.
                                        computer.org/cg/books/
                                        cg1998/pdf/g2104.pdf.

Glassner:1998:AGNe

[Gla98c] Andrew Glassner. Andrew Glassner’s notebook: Know when to fold. IEEE Computer Graphics and Applications, 18(5):79–85, September/October 1998. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://dlib.
computer.org/cg/books/
cg1998/pdf/g5079.pdf.

Glassner:1998:AGNf

[Gla98d]  Andrew Glassner.  An-

drew Glassner’s notebook: The triangular manuscripts. IEEE Computer Graphics and Applications, 18(6): 74–77, November/December 1998. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL

http://dlib.computer.org/
cg/books/cg1998/pdf/g6074.
pdf.

Glassner:1998:AGNa

[Gla98e] Andrew Glassner. Andrew Glassner’s notebook: Upon reflection. IEEE Computer Graphics and Applications, 18(1):86–92, January/ February 1998. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

computer.org/cga/cg1998/
g1086abs.htm;   http:/
/dlib.computer.org/cg/
books/cg1998/pdf/g1086.
pdf.

Glassner:1998:AGNd

[Gla98f] Andrews Glassner. Andrew Glassner’s notebook: Penrose tiling. IEEE Computer Graphics and Applications, 18(4):78–86, July/ August 1998. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Glassner:1999:AGNc

[Gla99a]  Andrew Glassner.  An-

drew Glassner’s notebook: An open and shut case. IEEE Computer Graphics and Applications, 19(3):82– 92, May/June 1999. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http:

                                        //dlib.computer.org/cg/
                                        books/cg1999/pdf/g3082.
                                        pdf.

Glassner:1999:AGNf

[Gla99b] Andrew Glassner. Andrew Glassner’s notebook: Celtic knots, part 2. IEEE Computer Graphics and Applications, 19(6):82–86, November/December 1999. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http:

                                        //dlib.computer.org/cg/
                                        books/cg1999/pdf/g6082.
                                        pdf.

Glassner:1999:AGNe

[Gla99c] Andrew Glassner. Andrew Glassner’s notebook: Celtic knotwork, part 1. IEEE Computer Graphics and Applications, 19(5):78–84, September/October 1999.

CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL

http://dlib.computer.org/
cg/books/cg1999/pdf/g5078.
pdf.

Glassner:1999:AGNa

[Gla99d]  Andrew Glassner.  An-

drew Glassner’s notebook: Fourier polygons. IEEE Computer Graphics and Applications, 19(1):84–91, January/February 1999. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http:

//dlib.computer.org/cg/
books/cg1999/pdf/g1084.
pdf.

Glassner:1999:AGNd

[Gla99e] Andrew Glassner. Andrew Glassner’s notebook: O say, can you see? IEEE Computer Graphics and Applications, 19(4):88–93, July/ August 1999. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://dlib.
computer.org/cg/books/
cg1999/pdf/g4088.pdf.

Glassner:1999:AGNb

[Gla99f] Andrew Glassner. Andrew Glassner’s notebook: String crossings. IEEE Computer Graphics and Applications, 19(2):88–95, March/ April 1999. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL http://dlib.
                                        computer.org/cg/books/
                                        cg1999/pdf/g2088.pdf.

Glassner:2000:AGNa

[Gla00a] Andrew Glassner. Andrew Glassner’s notebook: Celtic knots, part 3. IEEE Computer Graphics and Applications, 20(1):70–75, January/ February 2000. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL http://dlib.
                                        computer.org/cg/books/
                                        cg2000/pdf/g1070.pdf.

Glassner:2000:AGNd

[Gla00b] Andrew Glassner. Andrew Glassner’s notebook: Hierarchical textures. IEEE Computer Graphics and Applications, 20(4):104–109, July/August 2000. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL http://dlib.
                                        computer.org/cg/books/
                                        cg2000/pdf/g4104.pdf.

Glassner:2000:AGNe

[Gla00c] Andrew Glassner. Andrew Glassner’s notebook: Soap bubbles: Part 1. IEEE Computer Graphics and Applications, 20(5):76–84, September/October 2000. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL http://dlib.
                                        computer.org/cg/books/
                                        cg2000/pdf/g5076.pdf.

Glassner:2000:AGNf

[Gla00d] Andrew Glassner. Andrew Glassner’s notebook: Soap bubbles: Part 2. IEEE Computer Graphics and Applications, 20(6):99–109, November/December 2000. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http:

//dlib.computer.org/cg/
books/cg2000/pdf/g6099.
pdf.

Glassner:2000:AGNb

[Gla00e] Andrew Glassner. Andrew Glassner’s notebook: The digital ceraunoscope: Synthetic thunder and lightning, part 1. IEEE Computer Graphics and Applications, 20(2):89–93, March/ April 2000. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://dlib.
computer.org/cg/books/
cg2000/pdf/g2089.pdf.

Glassner:2000:AGNc

[Gla00f] Andrew Glassner. Andrew Glassner’s notebook: The digital ceraunoscope: Synthetic thunder and lightning, part 2. IEEE Computer Graphics and Applications, 20(3):92–96, May/ June 2000. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://dlib.
computer.org/cg/books/
cg2000/pdf/g3092.pdf.

Glassner:2000:MF

[Gla00g] Andrew Glassner. Manned flight. IEEE Computer Graphics and Applications, 20(1):48–49, January/February 2000. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL

                                        http://dlib.computer.org/
                                        cg/books/cg2000/pdf/g1048.
                                        pdf.

Glassner:2001:AGNc

[Gla01a]  Andrew Glassner.  An-

drew Glassner’s notebook: a change of scene. IEEE Computer Graphics and Applications, 21(3):86–92, May/ June 2001. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

                                        computer.org/cga/cg2001/
                                        extras/g3086x1.avi;   http:
                                        //computer.org/cga/cg2001/
                                        extras/g3086x2.avi;   http:
                                        //computer.org/cga/cg2001/
                                        extras/g3086x3.avi;   http:
                                        //dlib.computer.org/cg/
                                        books/cg2001/pdf/g3086.
                                        pdf.

Glassner:2001:AGNa

[Gla01b]  Andrew Glassner.  An-

drew Glassner’s notebook: Fill ’er up! IEEE Computer Graphics and Applications, 21(1):78–85, January/ February 2001. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL http://dlib.

computer.org/cg/books/
cg2001/pdf/g1078.pdf.

Glassner:2001:AGNd

[Gla01c] Andrew Glassner. Andrew Glassner’s notebook: Quantum computing, part 1. IEEE Computer Graphics and Applications, 21(4):84– 92, July/August 2001. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http:

//dlib.computer.org/cg/
books/cg2001/pdf/g4084.
pdf.

Glassner:2001:AGNe

[Gla01d]  Andrew Glassner.  An-

drew Glassner’s notebook: Quantum computing, part

2. IEEE Computer Graphics and Applications, 21(5): 86–95, September/October 2001. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL

http://dlib.computer.org/
cg/books/cg2001/g5086abs.
htm;   http://dlib.computer.
org/cg/books/cg2001/pdf/
g5086.pdf.

Glassner:2001:AGNf

[Gla01e]  Andrew Glassner.  An-

drew Glassner’s notebook: Quantum computing, part

3. IEEE Computer Graphics and Applications, 21(6): 72–82, November/December 2001. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL

                                        http://dlib.computer.org/
                                        cg/books/cg2001/g6072abs.
                                        htm;   http://dlib.computer.
                                        org/cg/books/cg2001/pdf/
                                        g6072.pdf.

Glassner:2001:AGNb

[Gla01f] Andrew Glassner. Andrew Glassner’s notebook: Tricks of the trade. IEEE Computer Graphics and Applications, 21(2):80–87, March/ April 2001. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL http://dlib.
                                        computer.org/cg/books/
                                        cg2001/pdf/g2080.pdf.

Glassnar:2002:AGN

[Gla02a]  Andrew Glassnar.  An-

drew Glassner’s notebook: Digital weaving, part 1. IEEE Computer Graphics and Applications, 22 (6):108–118, November/ December 2002. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL http://dlib.
                                        computer.org/cg/books/
                                        cg2002/pdf/g6108.pdf.

Glassner:2002:AGNd

[Gla02b] Andrew Glassner. Andrew Glassner’s notebook: Duck! IEEE Computer Graphics and Applications, 22(4):88– 97, July/August 2002. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

                                        computer.org/cga/cg2002/

g4088abs.htm;   http:/
/dlib.computer.org/cg/
books/cg2002/pdf/g4088.
pdf.

Glassner:2002:AGNe

[Gla02c] Andrew Glassner. Andrew Glassner’s notebook: Getting the picture. IEEE Computer Graphics and Applications, 22(5):76–85, September/October 2002. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

computer.org/cga/cg2002/
extras/g5076.htm;   http:
//dlib.computer.org/cg/
books/cg2002/pdf/g5076.
pdf.

Glassner:2002:AGNb

[Gla02d] Andrew Glassner. Andrew Glassner’s notebook: Interactive pop-up card design, part 2. IEEE Computer Graphics and Applications, 22(2):74–85, March/ April 2002. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

computer.org/cga/cg2002/
g2074abs.htm;   http:/
/dlib.computer.org/cg/
books/cg2002/pdf/g2074.
pdf.

Glassner:2002:AGNa

[Gla02e] Andrew Glassner. Andrew Glassner’s notebook: Interactive pop-up card design, Part I. IEEE Computer Graphics and Applications, 22(1):79–86, January/ February 2002. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://dlib.
computer.org/cg/books/
cg2002/g1079abs.htm;
http://dlib.computer.org/
cg/books/cg2002/pdf/g1079.
pdf.

Glassner:2002:AGNc

[Gla02f]  Andrew Glassner.  An-

drew Glassner’s notebook: Putting the pieces together. IEEE Computer Graphics and Applications, 22(3):76– 86, May/June 2002. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

computer.org/cga/cg2002/
g3076abs.htm;   http:/
/dlib.computer.org/cg/
books/cg2002/pdf/g3076.
pdf.

Glassner:2003:AGNa

[Gla03a] Andrew Glassner. Andrew Glassner’s notebook: Digital weaving, part 2. IEEE Computer Graphics and Applications, 23(1):77–90, January/February 2003. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http:

//dlib.computer.org/cg/
books/cg2003/pdf/g1077.
pdf.

Glassner:2003:AGNb

[Gla03b]  Andrew Glassner.  An-

drew Glassner’s notebook: Digital weaving, part 3. IEEE Computer Graphics and Applications, 23(2):80– 89, March/April 2003. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http:

//dlib.computer.org/cg/
books/cg2003/pdf/g2080.
pdf.

Glassner:2003:AGNe

[Gla03c]  Andrew Glassner.  An-

drew Glassner’s notebook: DMorph. IEEE Computer Graphics and Applications, 23(5):78–83, September/October 2003. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http:

//csdl.computer.org/dl/
mags/cg/2003/05/g5078.
pdf.

Glassner:2003:AGNf

[Gla03d]  Andrew Glassner.  An-

drew Glassner’s notebook: Everyday computer graphics. IEEE Computer Graphics and Applications, 23(6):76–82, November/ December 2003. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http:

//csdl.computer.org/dl/
mags/cg/2003/06/g6076.
pdf.

Glassner:2003:AGNc

[Gla03e]  Andrew Glassner.  An-

drew Glassner’s notebook: Image search and replace. IEEE Computer Graphics and Applications, 23(3):80– 88, May/June 2003. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http:

//dlib.computer.org/cg/
books/cg2003/pdf/g3080.
pdf.

Glassner:2003:AGNd

[Gla03f] Andrew Glassner. Andrew Glassner’s notebook: Venn and now. IEEE Computer Graphics and Applications, 23(4):82–95, July/ August 2003. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http:/

/csdl.computer.org/dl/
mags/cg/2003/04/g4082.
htm;   http://csdl.computer.
org/dl/mags/cg/2003/04/
g4082.pdf.

Glassner:2004:AGNa

[Gla04a]  Andrew Glassner.  An-

drew Glassner’s notebook: About face. IEEE Computer Graphics and Applications, 24(1):86–91, January/February 2004. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http:

//csdl.computer.org/dl/
mags/cg/2004/01/g1086.
pdf.

Glassner:2004:AGNb

[Gla04b]  Andrew Glassner.  An-

drew Glassner’s notebook: Around and around. IEEE Computer Graphics and Applications, 24(2):98–108, March/April 2004. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http:

//csdl.computer.org/dl/
mags/cg/2004/02/g2098.
htm;   http://csdl.computer.
org/dl/mags/cg/2004/02/
g2098.pdf.

Glassner:2004:AGNe

[Gla04c] Andrew Glassner. Andrew Glassner’s notebook: Crop art, part 1. IEEE Computer Graphics and Applications, 24(5):86–99, September/October 2004. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http:

//csdl.computer.org/dl/
mags/cg/2004/05/g5086.
htm;   http://csdl.computer.
org/dl/mags/cg/2004/05/
g5086.pdf.

Glassner:2004:AGNf

[Gla04d] Andrew Glassner. Andrew Glassner’s notebook: Crop art, part 2. IEEE Computer Graphics and Applications, 24(6):78–85, November/December 2004. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http:

//csdl.computer.org/dl/

                                        mags/cg/2004/06/g6078.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/cg/2004/06/
                                        g6078.pdf.

Glassner:2004:AGNc

[Gla04e] Andrew Glassner. Andrew Glassner’s notebook: Digital cubism. IEEE Computer Graphics and Applications, 24(3):82–90, May/ June 2004. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http:/

                                        /csdl.computer.org/dl/
                                        mags/cg/2004/03/g3082.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/cg/2004/03/
                                        g3082.pdf.

Glassner:2004:AGNd

[Gla04f]  Andrew Glassner.  An-

drew Glassner’s notebook: Digital cubism, part 2. IEEE Computer Graphics and Applications, 24(4):84– 95, July/August 2004. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http:

                                        //csdl.computer.org/dl/
                                        mags/cg/2004/04/g4084.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/cg/2004/04/
                                        g4084.pdf.

Glassner:2005:AGNa

[Gla05a] Andrew Glassner. Andrew Glassner’s notebook: Crop art, part 3. IEEE Computer Graphics and Applications, 25(1):88–91, January/ February 2005. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Glassner:2005:AGNb

[Gla05b]  Andrew Glassner.  An-

drew Glassner’s notebook: Speed limit 55. IEEE Computer Graphics and Applications, 25(2):96–106, March/ April 2005. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Georgiev:2011:UFP

[GLC11] Todor Georgiev, Andrew Lumsdaine, and Georgi Chunev. Using focused plenoptic cameras for rich image capture. IEEE Computer Graphics and Applications, 31(1):62–73, January/ February 2011. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Goodall:2006:FCN

[GLRK06] John R. Goodall, Wayne G.

Lutters, Penny Rheingans, and Anita Komlodi. Focusing on context in network traffic analysis. IEEE Computer Graphics and Applications, 26(2):72–80, March/ April 2006. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Gupta:1987:PCD

[GM87] Satish Gupta and Daniel H. McCabe. Personal computer displays. IEEE Computer Graphics and Applications, 7(10):17–23, October 1987. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Grimm:2012:BAD

[GM12] Cindy Grimm and G. Elisabeta Marai. Biomedical applications: From data capture to modeling. IEEE Computer Graphics and Applications, 32(5):20–21, September/October 2012. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Gaither:1997:PVV

[GMNF97] Kelly Gaither, Robert Moorhead, Scott Nations, and Dan Fox. Projects in VR: Visualizing ocean circulation models through virtual environments. IEEE Computer Graphics and Applications, 17(1):16–19, January/ February 1997. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Gillmann:2023:TUE

[GMR^{+}23] Christina Gillmann, Robin Georg Claus Maack, Felix Raith, Juan F. Pérez, and Gerik Scheuermann. A taxonomy of uncertainty events in visual analytics. IEEE Computer Graphics and Applications, 43(5):62–71, September/October 2023. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Goldfeather:1989:NRC

[GMTF89] Jack Goldfeather, Steven Molnar, Greg Turk, and Henry Fuchs. Near real-time CSG rendering using tree normalization and geometric pruning. IEEE Computer Graphics and Applications, 9(3):20–28, May/June 1989. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Gardner:1983:IGD

[GN83] Thomas N. Gardner and H. Roice Nelson, Jr. Interactive graphics developments in energy exploration. IEEE Computer Graphics and Applications, 3(2):33–38, 40, 42–44, March/April 1983. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Goebel:1996:PVI

[Goe96] Martin Goebel. Projects in VR: industrial applications of VEs. IEEE Computer Graphics and Applications, 16(1):10–13, January/ February 1996. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Goldman:1983:QR

[Gol83a] Ronald N. Goldman. Quadrics of revolution. IEEE Computer Graphics and Applications, 3(2):68–71, 73–76, March/April 1983. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Goldman:1983:TAC

[Gol83b] Ronald N. Goldman. Two approaches to a computer model for quadric surfaces. IEEE Computer Graphics and Applications, 3(6):21– 24, September 1983. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Goldman:1983:UBF

[Gol83c] Ronald N. Goldman. An urnful of blending functions. IEEE Computer Graphics and Applications, 3(7):49– 54, October 1983. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Goldwasser:1984:GOD

[Gol84] Samuel M. Goldwasser. Generalized object display processor architecture. IEEE Computer Graphics and Applications, 4(10):43– 55, October 1984. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Goldman:1986:UMB

[Gol86] R. Goldman. Urn models and beta-splines. IEEE Computer Graphics and Applications, 6(2):57–64, February 1986. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Goldman:2000:ASC

[Gol00] Ron Goldman. The ambient spaces of computer graphics and geometric modeling. IEEE Computer Graphics and Applications, 20(2):76– 84, March/April 2000. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

computer.org/cga/cg2000/
g2076abs.htm;   http:/
/dlib.computer.org/cg/
books/cg2000/pdf/g2076.
pdf.

Goldman:2001:BAL

[Gol01] Ron Goldman. Baseball arithmetic and the laws of pseudoperspective. IEEE Computer Graphics and Applications, 21(2):70–78, March/April 2001. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://dlib.
computer.org/cg/books/
cg2001/pdf/g2070.pdf.

Goldman:2003:DLT

[Gol03] Ron Goldman. Deriving linear transformations in three dimensions. IEEE Computer Graphics and Applications, 23(3):66–71, May/ June 2003. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://dlib.
computer.org/cg/books/
cg2003/pdf/g3066.pdf;
http://www.computer.org/
cga/cg2003/g3066abs.htm.

Goldman:2012:DMB

[Gol12]  Glenn Goldman.  Digi-

tal media and the beginning designer. IEEE Computer Graphics and Applications, 32(2):14–21, March/ April 2012. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Goodfellow:1986:WWH

[Goo86] Michael J. Goodfellow. WHIM — window handler and input manager. IEEE Computer Graphics and Applications, 6(5):46–52, May/ June 1986. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Grubert:2018:OFV

[GOPK18] Jens Grubert, Eyal Ofek, Michel Pahud, and Per Ola Kristensson. The office of the future: Virtual, portable, and global. IEEE Computer Graphics and Applications, 38(6):125–133, ???? 2018. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL https:

//ieeexplore.ieee.org/
document/8617763/.

Gordon:1983:OCS

[Gor83] William J. Gordon. An operator calculus for surface and volume modeling. IEEE Computer Graphics and Applications, 3(7):18–22, October 1983. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Gore:1996:OS

[Gor96] Al Gore. Opening statement. IEEE Computer Graphics and Applications, 16(2):60–??, March/April 1996. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Goss:1990:RTP

[Gos90] Michael E. Goss. A real time particle system for display of ship wakes. IEEE Computer Graphics and Applications, 10(3):30–35, May/ June 1990. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Gott:1982:CGQ

[Got82]  Virgil W. Gott.  Com-

puter graphics for quality assurance. IEEE Computer Graphics and Applications, 2(1):39–41, 43–44, January/ February 1982. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Gotsis:2009:GVR

[Got09] Marientina Gotsis. Games, virtual reality, and the pursuit of happiness. IEEE Computer Graphics and Applications, 29(5):14–19, September/October 2009. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Green:1989:ECM

[GP89] Stuart A. Green and Derek J. Paddon. Exploiting coherence for multiprocessor ray tracing. IEEE Computer Graphics and Applications, 9(6):12–26, November/December 1989. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Guenter:1990:CAL

[GP90] Brian Guenter and Richard Parent. Computing the arc length of parametric curves. IEEE Computer Graphics and Applications, 10(3):72– 78, May/June 1990. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Grimes:1991:GEI

[GP91a] Jack Grimes and Mike Potel. Guest Editor’s introduction: Multimedia — it’s actually useful! IEEE Computer Graphics and Applications, 11(4):24–25, July/ August 1991. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Grimes:1991:WIM

[GP91b] Jack Grimes and Mike Potel. What is multimedia? IEEE Computer Graphics and Applications, 11(1):49– 52, January/February 1991. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Giertsen:1993:PVR

[GP93a] Christopher Giertsen and Johnny Petersen. Parallel volume rendering on a network of workstations. IEEE Computer Graphics and Applications, 13(6):16–23, November/December 1993. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Grimes:1993:GEI

[GP93b] Jack Grimes and Mike Potel. Guest Editors’ introduction: Amazing desktop applications. IEEE Computer Graphics and Applications, 13(3):14–??, May/ June 1993. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Garcia:2017:NIR

[GPA^{+}17] Juan C. Garcia, Bruno Patrao, Luis Almeida, Javier Perez, Paulo Menezes, Jorge Dias, and Pedro J. Sanz. A natural interface for remote operation of underwater robots. IEEE Computer Graphics and Applications, 37(1):34–43, January/ February 2017. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2017/01/mcg2017010034-
abs.html.

Gordon:1994:FPS

[GPR94] Dan Gordon, Michael A. Peterson, and R. Anthony Reynolds. Fast polygon scan conversion with medical applications. IEEE Computer Graphics and Applications, 14(6):20–27, November/December 1994. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Gillmann:2021:VMD

[GPS^{+}21] Christina Gillmann, Lucas Peter, Carlo Schmidt, Dorothee Saur, and Gerik Scheuermann. Visualizing multimodal deep learning for lesion prediction. IEEE Computer Graphics and Applications, 41(5):90–98, September/October 2021. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Gillmann:2019:HIS

[GPW^{+}19] C. Gillmann, T. Post, T. Wischgoll, H. Hagen, and R. Maciejewski. Hierarchical image semantics using probabilistic path propagations for biomedical research. IEEE Computer Graphics and Applications, 39(6):86–101, November/ December 2019. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Goldman:1994:CW

[GR94] J. Goldman and T. M.

Roy. The cosmic worm. IEEE Computer Graphics and Applications, 14(4): 12–14, July/August 1994. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Goldwasser:1985:PWR

[GRB^{+}85] Samuel M. Goldwasser, R. Anthony Reynolds, Ted Bapty, David Baraff, John Summers, David A. Talton, and Ed Walsh. Physician’s workstation with real-time performance. IEEE Computer Graphics and Applications, 5(12):44–57, December 1985. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Green:1984:TCH

[Gre84] Paul Green. Teaching a course on human factors and computer systems. IEEE Computer Graphics and Applications, 4(12):43–45, December 1984. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Green:1986:GI

[Gre86a] Mark Green. Graphics Interface 86. IEEE Computer Graphics and Applications, 6(11):11–67, November/December 1986. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Greene:1986:EMO

[Gre86b] Ned Greene. Environment mapping and other applications of world projection. IEEE Computer Graphics and Applications, 6(11): 21–29, November/December 1986. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Greenberg:1991:MAS

[Gre91a] Donald P. Greenberg. More accurate simulations at faster rates. IEEE Computer Graphics and Applications, 11(1):23–29, January/ February 1991. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Greenberg:1991:RTS

[Gre91b] Donald P. Greenberg. A ray tracing simulation of a radiosity simulation. IEEE Computer Graphics and Applications, 11(1):6–7, January/February 1991. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Grotch:1983:TSG

[Gro83] Stanley L. Grotch. Threedimensional and stereoscopic graphics for scientific data display and analysis. IEEE Computer Graphics and Applications, 3(8): 31–43, November/December 1983. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Gross:2006:GP

[Gro06] Markus Gross. Getting to the point… ? IEEE Computer Graphics and Applications, 26(5):96–99, September/October 2006. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://bell.
computer.org/dlcomments/

.

Gleicher:2023:PSD

[GRvL^{+}23] Michael Gleicher, Maria Riveiro, Tatiana von Landesberger, Oliver Deussen, Remco Chang, and Christina Gillman. A problem space for designing visualizations. IEEE Computer Graphics and Applications, 43(4): 111–120, July/August 2023. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Goldiez:1999:PVR

[GRW99] Brian Goldiez, Rodney Rogers, and Pam Woodard. Projects in VR: Real-time visual simulation on PCs. IEEE Computer Graphics and Applications, 19(1):11– 15, January/February 1999. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL

http://dlib.computer.org/
cg/books/cg1999/pdf/g1011.
pdf.

Granholm:1987:CDC

[GRWN87] John W. Granholm, Douglas D. Robertson, Peter S. Walker, and Philip C. Nelson. Computer design of custom femoral stem prostheses. IEEE Computer Graphics and Applications, 7(2):26–35, February 1987. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Gelberg:1987:SGE

[GS87a] Lawrence Gelberg and Thomas Stephenson. Supercomputing and graphics in the earth and planetary sciences. IEEE Computer Graphics and Applications, 7(7):26–33, July/August 1987. CODEN ICGADZ.

ISSN 0272-1716 (print), 1558-1756 (electronic).

Goldsmith:1987:ACO

[GS87b] Jeffrey Goldsmith and John Salmon. Automatic creation of object hierarchies for ray tracing. IEEE Computer Graphics and Applications, 7(5):14–20, May/June 1987. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Green:1988:IAL

[GS88] Mark Green and Hanqiu Sun. Interactive animation: a language and system for procedural modeling and motion. IEEE Computer Graphics and Applications, 8(6):52–64, November/December 1988. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Greiner:1994:MTB

[GS94] Gunther Greiner and Hans-Peter Seidel. Modeling with triangular B-splines. IEEE Computer Graphics and Applications, 14(2): 56–60, March/April 1994. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Ganovelli:2012:OCM

[GS12] Fabio Ganovelli and Roberto Scopigno. OCME: Out-of-Core Mesh Editing made practical. IEEE Computer Graphics and Applications, 32(3):46–58, May/ June 2012. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Gonzalez:2008:VBA

[GSF08] Francisco González, Mateu Sbert, and Miquel Feixas. Viewpoint-based ambient occlusion. IEEE Computer Graphics and Applications, 28(2):44–51, March/ April 2008. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Gunther:2016:VVC

[GSFT16] Tobias Gunther, Maik Schulze, Anke Friederici, and Holger Theisel. Visualizing volcanic clouds in the atmosphere and their impact on air traffic. IEEE Computer Graphics and Applications, 36(3):36–47, May/June 2016. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2016/03/mcg2016030036-
abs.html.

George:1990:RRD

[GSG90] David W. George, François X. Sillion, and Donald P. Greenberg. Radiosity redistribution for dynamic environments. IEEE Computer Graphics and Applications, 10(4):26–34, July/ August 1990. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Gillmann:2021:TOC

[GSG^{+}21] Christina Gillmann, Noeska N.

Smit, Eduard Gröller, Bernhard Preim, Anna Vilanova, and Thomas Wischgoll. Ten open challenges in medical visualization. IEEE Computer Graphics and Applications, 41(5):7–15, September/October 2021. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Greger:1998:IV

[GSHG98] Gene Greger, Peter Shirley, Philip M. Hubbard, and Donald P. Greenberg. The irradiance volume. IEEE Computer Graphics and Applications, 18(2):32–43, March/April 1998. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

computer.org/cga/cg1998/
g2032abs.htm;   http:/
/dlib.computer.org/cg/
books/cg1998/pdf/g2032.
pdf.

Georgel:2009:NTV

[GSN09] Pierre Fite Georgel, Pierre Schroeder, and Nassir Navab. Navigation tools for viewing augmented CAD models. IEEE Computer Graphics and Applications, 29(6): 65–73, November/December 2009. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Gillet:2005:TAI

[GSSO05] Alexandre Gillet, Michel Sanner, Daniel Stoffler, and Arthur Olson. Tangible augmented interfaces for structural molecular biology. IEEE Computer Graphics and Applications, 25(2): 13–17, March/April 2005. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Gillmann:2023:VW

[GSW23] Christina Gillmann, Johanna Schmidt, and Daniel Wiegreffe. Visualization in the wild. IEEE Computer Graphics and Applications, 43(6):46–49, November/December 2023. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Gillmann:2024:IVA

[GSW24] Christina Gillmann, Johanna Schmidt, and Daniel Wiegreffe. Interactive visualization in applications. IEEE Computer Graphics and Applications, 44(1):10– 12, January/February 2024. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Gupta:1993:VSM

[GT93] Suvajit Gupta and Joshua U.

Turner. Variational solid modeling for tolerance analysis. IEEE Computer Graphics and Applications, 13(3):64–74, May/June 1993. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Gueziec:1999:FSG

[GTHL99] André Guéziec, Gabriel Taubin, Bill Horn, and Francis Lazarus. A framework for streaming geometry in VRML. IEEE Computer Graphics and Applications, 19(2):68–78, March/ April 1999. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

computer.org/cga/cg1999/
g2068abs.htm;   http:/
/dlib.computer.org/cg/
books/cg1999/pdf/g2068.
pdf.

Goodman:1986:MSP

[GU86] T. Goodman and K. Unsworth.

Manipulating shape and producing geometric continuity in beta-spline surfaces. IEEE Computer Graphics and Applications, 6(2):50– 56, February 1986. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Gattullo:2015:LIA

[GUFG15] Michele Gattullo, Antonio E. Uva, Michele Fiorentino, and Joseph L. Gabbard. Legibility in industrial AR: Text style, color coding, and illuminance. IEEE Computer Graphics and Applications, 35(2):52–61, March/ April 2015. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL http://www.
                                        computer.org/csdl/mags/
                                        cg/2015/02/mcg2015020052-
                                        abs.html.

Gunther:2020:VTI

[Gün20]  T. Günther.  Visibility,

topology, and inertia: New methods in flow visualization. IEEE Computer Graphics and Applications, 40(2):103–111, March/April 2020. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Gust:1983:DEB

[Gus83] Raymond K. Gust. Designing effective business graphics. IEEE Computer Graphics and Applications, 3(4): 33–36, July/August 1983. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

George:1981:DES

[GV81] James E. George and Anders Vinberg. The display of engineering and scientific data. IEEE Computer Graphics and Applications, 1(3):49–54, July/ August 1981. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Giloth:1985:PP

[GV85] Copper Giloth and Jane Veeder. The paint problem. IEEE Computer Graphics and Applications, 5(7):66– 75, July/August 1985. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Guttag:1986:RVG

[GVA86] Karl M. Guttag, Jerry R.

Van Aken, and Michael D. Asal. Requirements for a VLSI graphics processor. IEEE Computer Graphics and Applications, 6(1):32– 47, January/February 1986. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Glassner:1998:GEI

[GW98] Andrew Glassner and Turner Whitted. Guest Editors’ introduction. IEEE Computer Graphics and Applications, 18(2):20–21, March/ April 1998. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

computer.org/cga/cg1998/
g2022abs.htm;   http:/
/dlib.computer.org/cg/

books/cg1998/pdf/g2020.
pdf.

Gooch:2008:CA

[GW08] Bruce Gooch and Brian Wyvill. Computational aesthetics. IEEE Computer Graphics and Applications, 28(2):23, March/ April 2008. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cg/2008/02/mcg2008020023.
pdf.

Gonzalez:1984:EEP

[GWA84] J. Camacho Gonzalez, M. H. Williams, and I. E. Aitchison. Evaluation of the effectiveness of Prolog for a CAD application. IEEE Computer Graphics and Applications, 4(3):67–75, March/ April 1984. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Gelband:1998:TET

[GWF98] Patrice Gelband, Sid Weber, and Susan Fryer. Taking a 2D educational title into 3D. IEEE Computer Graphics and Applications, 18(1):54–57, January/ February 1998. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

computer.org/cga/cg1998/
g1054abs.htm;   http:/

/dlib.computer.org/cg/
books/cg1998/pdf/g1054.
pdf.

Goodnight:2005:CPG

[GWH05] Nolan Goodnight, Rui Wang, and Greg Humphreys. Computation on programmable graphics hardware. IEEE Computer Graphics and Applications, 25(5):12–15, September/October 2005. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Gargantini:1986:VTV

[GWW86] Irene A. Gargantini, Timothy R. Walsh, and Oi-Lun Wu. Viewing transformations of voxel-based objects via linear octrees. IEEE Computer Graphics and Applications, 6(10):12–21, October 1986. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Gong:2015:APQ

[GXMW15] Wenyong Gong, Xiaohua Xie, Rui Ma, and Tieru Wu. Angle-preserving quadrilateral mesh parameterization. IEEE Computer Graphics and Applications, 35(6):51–59, November/ December 2015. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://www.
computer.org/csdl/mags/

cg/2015/06/mcg2015060051-
abs.html.

Guttman:1997:ACF

[GZM97] Michael A. Guttman, Elias A. Zerhouni, and Elliot R. McVeigh. Analysis of cardiac function from MR images. IEEE Computer Graphics and Applications, 17(1):30–38, January/ February 1997. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Gossard:1988:RDT

[GZS88] David C. Gossard, Robert P. Zuffante, and Hiroshi Sakurai. Representing dimensions, tolerances, and features in MCAE systems. IEEE Computer Graphics and Applications, 8(2):51– 59, March/April 1988. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http:/

/www.computer.org/csdl/
mags/cg/02/00000503-abs.
html;   http://www.computer.
org/csdl/mags/cg/02/mcg1988020051-
abs.html.

Hodges:2001:TPP

[HAB^{+}01] Larry F. Hodges, Page Anderson, Grigore C. Burdea, Hunter G. Hoffman, and Barbara O. Rothbaum. Treating psychological and physical disorders with VR. IEEE Computer Graphics and Applications, 21(6):25–33, November/ December 2001. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://dlib.
computer.org/cg/books/
cg2001/g6025abs.htm;
http://dlib.computer.org/
cg/books/cg2001/pdf/g6025.
pdf.

Hanna:1983:IPS

[HAG83] Samir L. Hanna, John F.

Abel, and Donald P. Greenberg. Intersection of parametric surfaces by means of look-up tables. IEEE Computer Graphics and Applications, 3(7):39–48, October 1983. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Haggerty:1990:CDC

[Hag90a]  Michael Haggerty.  Col-

ors from the dark continent. IEEE Computer Graphics and Applications, 10(3):8–10, May/June 1990. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Haggerty:1990:EWA

[Hag90b] Michael Haggerty. Exploring the world in the art of Matt Mullican. IEEE Computer Graphics and Applications, 10(3):3–7, May/ June 1990. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Haggerty:1990:LAC

[Hag90c] Michael Haggerty. Looking at clouds from both sides now. IEEE Computer Graphics and Applications, 10(5):5–7, September 1990. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Haggerty:1990:RHO

[Hag90d]  Michael Haggerty.  Re-

making history one pixel at a time. IEEE Computer Graphics and Applications, 10(4):3–6, July/ August 1990. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Haggerty:1991:AAR

[Hag91a] Michael Haggerty. The art of artificial reality. IEEE Computer Graphics and Applications, 11(1):8–14, January/ February 1991. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Haggerty:1991:EE

[Hag91b] Michael Haggerty. Evolution by esthetics. IEEE Computer Graphics and Applications, 11(2):5–9, March/ April 1991. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Haggerty:1994:DDT

[Hag94]  Michael Haggerty.  Dis-

plays on display: Translating mathematics and icons into art. IEEE Computer Graphics and Applications, 14(1):7–9, January/ February 1994. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Henstrom:2023:THV

[HAH^{+}23] Jordan A. Henstrom, Jonathan M.

Alexander, Jackson T. Hart, Ian N. Backus, and Raffaele De Amicis. Toward a holistic virtual environment for graph visualization and analysis. IEEE Computer Graphics and Applications, 43(3):48–53, May/ June 2023. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Haines:1987:PSG

[Hai87] Eric A. Haines. A proposal for standard graphics environments. IEEE Computer Graphics and Applications, 7(11):3–5, November/ December 1987. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Haines:2006:ITI

[Hai06] Eric Haines. An introductory tour of interactive rendering. IEEE Computer Graphics and Applications, 26(1):76–87, January/ February 2006. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Hall:1987:PIP

[Hal87] Tom Hall. Productivity issues in PC-CAD graphics. IEEE Computer Graphics and Applications, 7(10):13– 16, October 1987. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Hall:1999:CSC

[Hal99]  Roy Hall.  Comparing

spectral color computation methods. IEEE Computer Graphics and Applications, 19(4):36–45, July/ August 1999. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

computer.org/cga/cg1999/
g4036abs.htm;   http:/
/dlib.computer.org/cg/
books/cg1999/pdf/g4036.
pdf.

Hanson:2000:ABF

[Han00a] Mary Hanson. Applications: Beyond first aid: Emergency response teams turn to graphics. IEEE Computer Graphics and Applications, 20(6):12–18, November/December 2000. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http:

//dlib.computer.org/cg/
books/cg2000/pdf/g6012.
pdf.

Hanson:2000:AFB

[Han00b] Mary Hanson. Applications:

Fingerprint-based forensics identify Argentina’s desaparecidos. IEEE Computer Graphics and Applications, 20(5):7–10, September/October 2000. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://dlib.
computer.org/cg/books/
cg2000/pdf/g5007.pdf.

Hanson:2001:ASS

[Han01] Mary Hanson. Applications: Something to smile about: 3D graphics are revolutionizing oral health care. IEEE Computer Graphics and Applications, 21(4):14– 20, July/August 2001. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http:

//dlib.computer.org/cg/
books/cg2001/pdf/g4014.
pdf.

Hanson:2014:PSF

[Han14] Andrew J. Hanson. Putting science first: Distinguishing visualizations from pretty pictures. IEEE Computer Graphics and Applications, 34(4):63–69, July/August 2014. CODEN ICGADZ.

ISSN 0272-1716 (print), 1558-1756 (electronic).

Hart:1996:FIC

[Har96] John C. Hart. Fractal image compression and recurrent iterated function systems. IEEE Computer Graphics and Applications, 16(4): 25–33, July/August 1996. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Harrison:2000:BFR

[Har00] Beverly L. Harrison. Ebooks and the future of reading. IEEE Computer Graphics and Applications, 20(3):32–39, May/ June 2000. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

                                        computer.org/cga/cg2000/
                                        g3032abs.htm;   http:/
                                        /dlib.computer.org/cg/
                                        books/cg2000/pdf/g3032.
                                        pdf.

Hays:1995:ACA

[Hay95a] Nancy Hays. About the cover: Art under observation. IEEE Computer Graphics and Applications, 15(4):4–6, July/ August 1995. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Hays:1995:NEA

[Hay95b] Nancy Hays. In the news:

Excellence awards honor top picks of 1994; Silicon Graphics buys 3D modelers Wavefront and Alias; judge rejects government settlement with Microsoft; Microsoft joins dream team; software sales zoom in Western Europe and Latin America; JPL develops telerobotic calibration; 3D TV with no glasses; exhibit shows off robotic art. IEEE Computer Graphics and Applications, 15 (3):85–88, May/June 1995. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Hayes:1996:NSB

[Hay96a]  Nancy Hayes.  In the

news: Siggraph 96 brings the bayou inside; the enemy of my enemy is my friend; Communications Decency Act heads to Supreme Court; new television services go both ways. IEEE Computer Graphics and Applications, 16(5):90–91, September 1996. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Hays:1996:ASM

[Hay96b] Nancy Hays. Abstracts for solid modeling articles in March 1996 TVCG — IEEE Transactions on Visualization and Computer Graphics. IEEE Computer Graphics and Applications, 16(2): 16–17, March/April 1996. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Hays:1996:BLR

[Hay96c] Nancy Hays. Bell Labs researchers develop browsable TV. IEEE Computer Graphics and Applications, 16(2):90, March/April 1996. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Hays:1996:NVP

[Hay96d]  Nancy Hays.  In the

news: Visualization 95 provides tools, techniques; Comdex considers strippeddown PCs for Web; InfoVis 95 draws a crowd; VEs speed Bosnia peace negotiation. IEEE Computer Graphics and Applications, 16(1):88– 90, January/February 1996. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Hays:1996:IGR

[Hay96e] Nancy Hays. Interactive games remedy learning disabilities. IEEE Computer Graphics and Applications, 16(2):90, March/April 1996. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Hays:1996:MED

[Hay96f] Nancy Hays. Message from the Editors: Developing IEEE CG&A on the Web. IEEE Computer Graphics and Applications, 16(4):6, July/August 1996. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Hays:1997:ACE

[Hay97]  Nancy Hays.  About

the cover — expert user/ artist guides software development. IEEE Computer Graphics and Applications, 17(1):6–7, January/ February 1997. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Haruyama:1984:USM

[HB84] Shinichiro Haruyama and Brian A. Barsky. Using stochastic modeling for texture generation. IEEE Computer Graphics and Applications, 4(3):7–19, March/ April 1984. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Huang:1997:SNI

[HB97] Jianhua Huang and Ebad Banissi. Short note: An improved parallel circledrawing algorithm. IEEE Computer Graphics and Applications, 17(1):40–41, January/February 1997. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Harrower:2006:MOM

[HB06] Mark Harrower and Matt

Bloch.  MapShaper.org:

a map generalization Web service. IEEE Computer Graphics and Applications, 26(4):22–27, July/August 2006. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Hile:2008:POI

[HB08] Harlan Hile and Gaetano Borriello. Positioning and orientation in indoor environments using camera phones. IEEE Computer Graphics and Applications, 28(4):32–39, July/August 2008. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Hjelmervik:2016:IEB

[HB16] Jon M. Hjelmervik and Oliver J. D. Barrowclough. Interactive exploration of big scientific data: New representations and techniques. IEEE Computer Graphics and Applications, 36(3):6–9, May/June 2016. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2016/03/mcg2016030006-
abs.html.

Hein:2020:TBT

[HBB20] P. Hein, M. Bernhagen, and A. C. Bullinger. Two is better than one. improved attention guiding in AR by combining techniques. IEEE Computer Graphics and Applications, 40(5):57–66, September/October 2020. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Hersch:1995:PTG

[HBBG95] Roger D. Hersch, Claude Bétrisey, Justin Bur, and André Gürtler. Perceptually tuned generation of grayscale fonts. IEEE Computer Graphics and Applications, 15(6):78–89, November/December 1995. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Hua:2004:SSS

[HBG04] Hong Hua, Leonard D.

Brown, and Chunyu Gao. Scape: Supporting stereoscopic collaboration in augmented and projective environments. IEEE Computer Graphics and Applications, 24(1):66–75, January/ February 2004. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cg/2004/01/g1066abs.htm;
 http://csdl.computer.

                                        org/dl/mags/cg/2004/01/
                                        g1066.pdf.

Huang:2007:IIR

[HBP^{+}07] Jingshu Huang, Brian Bue, Avin Pattath, David S. Ebert, and Krystal M. Thomas. Interactive illustrative rendering on mobile devices. IEEE Computer Graphics and Applications, 27(3):48–56, May/ June 2007. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Hohne:1992:VAA

[HBR^{+}92a] Karl Heinz Hohne, Michael Bomans, Martin Reimer, Rainer Schubert, and Ulf Tiede. A volume-based anatomical atlas. IEEE Computer Graphics and Applications, 12(7):72–77, July/August 1992. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Hohne:1992:CVB

[HBR^{+}92b] Karl Heinz Hohne, Michael Bomans, Martin Rimer, Rainer Schubert, Ulf Tiede, and Werner Lierse. Correction to ‘A volume-based anatomical atlas’ (July 1992, 72–78). IEEE Computer Graphics and Applications, 12(6):89–??, November/December 1992. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Herrell:1995:HPE

[HBW95] Russ Herrell, Joe Baldwin, and Chris Wilcox. High-quality polygon edging. IEEE Computer Graphics and Applications, 15(4): 68–74, July/August 1995. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Hess:2014:VEP

[HBW^{+}14] Martin Hess, Sebastian Bremm, Stephanie Weissgraeber, Kay Hamacher, Michael Goesele, Josef Wiemeyer, and Tatiana von Landesberger. Visual exploration of parameter influence on phylogenetic trees. IEEE Computer Graphics and Applications, 34(2): 48–56, March/April 2014. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Hollock:2024:VAS

[HBW^{+}24] David C. Hollock, Nicholas J.

Brunsink, Austin B. Whittaker, Andrew Lawson, Toni B. Pence, Brittany Morago, Elham Ebrahimi, James Stocker, Amelia Moody, and Amy Taylor. Virtual access to STEM careers: In the field experiments. IEEE Computer Graphics and Applications, 44(2):73–80, March/ April 2024. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Hanrahan:2021:DR

[HC21] Pat Hanrahan and Edwin Catmull. The design of RenderMan. IEEE Computer Graphics and Applications, 41(4):7–15, July/ August 2021. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Huang:2005:TPV

[HCC05] Chun-Rong Huang, Chu-Song Chen, and Pau-Choo Chung. Tangible photorealistic virtual museum. IEEE Computer Graphics and Applications, 25(1):15– 17, January/February 2005. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Han:2020:CVT

[HCL^{+}20] P. Han, Y. Chen, I. Liu, Y. Jang, L. Tsai, A. Chang, and Y. Hung. A compelling virtual tour of the Dunhuang Cave with an immersive head-mounted display. IEEE Computer Graphics and Applications, 40(1):40– 55, January/February 2020. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Chi:1998:PIV

[hCRBK98] Ed Huai hsin Chi, John Riedl, Phillip Barry, and Joseph Konstan. Principles for information visualization spreadsheets. IEEE Computer Graphics and Applications, 18(4):30–38, July/ August 1998. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

computer.org/cga/cg1998/
g4030abs.htm.

Hinrichs:2013:IPD

[HCV^{+}13] Uta Hinrichs, Sheelagh Carpendale, Nina Valkanova, Kai Kuikkaniemi, Giulio Jacucci, and Andrew Vande Moere. Interactive public displays. IEEE Computer Graphics and Applications, 33(2): 25–27, March/April 2013. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Herman:2001:VVM

[HD01] Ivan Herman and David Duke. Visualization viewpoints: Minimal graphics. IEEE Computer Graphics and Applications, 21(6):18–21, November/ December 2001. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://dlib.
computer.org/cg/books/
cg2001/g6018abs.htm;

                                        http://dlib.computer.org/
                                        cg/books/cg2001/pdf/g6018.
                                        pdf.

Hersch:1992:MRP

[HDF^{+}92] Roger D. Hersch, Pierre Demartines, Doris G. Fridman, Jean-Marc Aberle, and Peter Pisan. Multiprocessor raster plotting. IEEE Computer Graphics and Applications, 12(4):79–??, July/ August 1992. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Hardwick:1995:CED

[HDKS95] Martin Hardwick, Blair R. Downie, Mike Kutcher, and David L. Spooner. Concurrent engineering with delta files. IEEE Computer Graphics and Applications, 15(1):62–68, January/ February 1995. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Hachet:2013:TBI

[HdlRL^{+}13] Martin Hachet, Jean-Baptiste de la Rivière, Jérémy Laviole, Aurélie Cohé, and Sébastien Cursan. Touchbased interfaces for interacting with 3D content in public exhibitions. IEEE Computer Graphics and Applications, 33(2):80–85, March/ April 2013. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Herbison-Evans:1982:RAH

[HE82] Don Herbison-Evans. Realtime animation of human figure drawings with hidden lines omitted. IEEE Computer Graphics and Applications, 2(9):27–33, November/December 1982. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Healey:2002:PPS

[HE02] Christopher G. Healey and James T. Enns. Perception and painting: a search for effective, engaging visualizations. IEEE Computer Graphics and Applications, 22(2):10–15, March/ April 2002. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

computer.org/cga/cg2002/
g2010abs.htm;   http://
computer.org/cga/cg2002/
pdf/g2010.pdf;   http:
//dlib.computer.org/cg/
books/cg2002/pdf/g2010.
pdf.

Heckbert:1986:MME

[Hec86a] Paul Heckbert. Making the magic egg: a personal account. IEEE Computer Graphics and Applications, 6(6):3–8, June 1986. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Heckbert:1986:STM

[Hec86b] Paul S. Heckbert. Survey of texture mapping. IEEE Computer Graphics and Applications, 6(11):56–67, November/December 1986. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Hedelman:1984:DFA

[Hed84] Harold Hedelman. A data flow approach to procedural modeling. IEEE Computer Graphics and Applications, 4(1):16–26, January/ February 1984. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Heermann:1999:FGA

[Hee99] Philip D. Heermann. Firstgeneration ASCI production visualization environment. IEEE Computer Graphics and Applications, 19(5):66–71, September/ October 1999. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

                                        computer.org/cga/cg1999/
                                        g5066abs.htm;   http:/
                                        /dlib.computer.org/cg/
                                        books/cg1999/pdf/g5066.
                                        pdf.

Heisserman:1994:GGD

[Hei94] Jeff Heisserman. Generative geometric design. IEEE Computer Graphics and Applications, 14(2):37–45, March/April 1994. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Heisig:2003:TWV

[Hei03] Stephen Heisig. Treemaps for workload visualization. IEEE Computer Graphics and Applications, 23(2):60– 67, March/April 2003. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http:

//dlib.computer.org/cg/
books/cg2003/pdf/g2060.
pdf;   http://www.computer.
org/cga/cg2003/g2060abs.
htm.

Heide:2018:TCI

[Hei18] Felix Heide. Transient convolutional imaging. IEEE Computer Graphics and Applications, 38(6):106–117, ???? 2018. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL https:

//ieeexplore.ieee.org/
document/8480632/.

Hersch:1986:DCF

[Her86] Roger D. Hersch. Descriptive contour fill of partly degenerated shapes. IEEE Computer Graphics and Applications, 6(7):61–70, July/ August 1986. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Hertzog:1996:ECM

[Her96] Bertram Hertzog. Editorin-Chief’s message: a new column, a new year — and more challenges. IEEE Computer Graphics and Applications, 16(1):6, January/ February 1996. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Hertzmann:2003:SSB

[Her03] Aaron Hertzmann. A survey of stroke-based rendering. IEEE Computer Graphics and Applications, 23(4):70–81, July/ August 2003. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cg/2003/04/g4070abs.htm;
 http://csdl.computer.
org/dl/mags/cg/2003/04/
g4070.htm;   http://csdl.
computer.org/dl/mags/cg/
2003/04/g4070.pdf.

Hertzmann:2023:CPU

[Her23] Aaron Hertzmann. The curse of performative user studies. IEEE Computer Graphics and Applications, 43(6):112–116, November/ December 2023. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Hertzmann:2024:NIS

[Her24] Aaron Hertzmann. New insights in smooth occluding contours for nonphotorealistic rendering. IEEE Computer Graphics and Applications, 44(1):76–85, January/ February 2024. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Huber:2017:CAR

[HEW17] Markus Huber, Bernhard Eberhardt, and Daniel Weiskopf. Cloth animation retrieval using a motionshape signature. IEEE Computer Graphics and Applications, 37(6):52–64, November/December 2017. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL

https://www.computer.org/
csdl/mags/cg/2017/06/mcg2017060052-
abs.html.

Havemann:2007:SRC

[HF07] Sven Havemann and Dieter W. Fellner. Seven research challenges of generalized 3D documents. IEEE Computer Graphics and Applications, 27(3):70–76, May/June 2007. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Hachisu:2018:SRA

[HF18] Taku Hachisu and Masaaki Fukumoto. SpiroSurface: A repulsive and attractive force display for interactive tabletops using a pneumatic system. IEEE Computer Graphics and Applications, 38(4):54–70, July/ August 2018. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        cg/2018/04/mcg2018040054-
                                        abs.html.

Han:2013:VBS

[HFZ^{+}13] Xiaoguang Han, Hongbo Fu, Hanlin Zheng, Ligang Liu, and Jue Wang. A video-based system for hand-driven stop-motion animation. IEEE Computer Graphics and Applications, 33(6):70–81, November/ December 2013. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Hall:1983:TRI

[HG83] Roy A. Hall and Donald P. Greenberg. A testbed for realistic image synthesis. IEEE Computer Graphics and Applications, 3(8): 10–20, November/December 1983. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Haines:1986:LBS

[HG86] E. A. Haines and D. P.

Greenberg. The light buffer: a shadow testing accelerator. IEEE Computer Graphics and Applications, 6(9):6–16, September 1986. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Holzschuch:2013:DMS

[HG13] Nicolas Holzschuch and Jean-Dominique Gascuel. Double- and multiple-scattering effects in translucent materials. IEEE Computer Graphics and Applications, 33 (3):66–76, May/June 2013. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Holmquist:2004:BIE

[HGK^{+}04] Lars Erik Holmquist, Hans-Werner Gellersen, Gerd Kortuem, Albrecht Schmidt, Martin Strohbach, Stavros Antifakos, Florian Michahelles, Bernt Schiele, Michael Beigl, and Ramia Mazé. Building intelligent environments with Smart-Its. IEEE Computer Graphics and Applications, 24(1):56–64, January/February 2004. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

csdl.computer.org/comp/
mags/cg/2004/01/g1056abs.
htm;   http://csdl.computer.

                                        org/dl/mags/cg/2004/01/
                                        g1056.pdf.

Hu:2010:VGC

[HGK10] Yifan Hu, Emden R. Gansner, and Stephen Kobourov. Visualizing graphs and clusters as maps. IEEE Computer Graphics and Applications, 30(6): 54–66, November/December 2010. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Hattab:2017:DS

[HGMT17] Ammar Hattab, Ian Gonsher, Daniel Moreno, and Gabriel Taubin. Differential 3D scanning. IEEE Computer Graphics and Applications, 37(3):43–51, May/ June 2017. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        cg/2017/03/mcg2017030043-
                                        abs.html.

Hatfield:1982:GST

[HH82] L. Hatfield and B. Herzog. Graphics software — from techniques to principles. IEEE Computer Graphics and Applications, 2(1):59– 80, January/February 1982. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Helman:1991:VVF

[HH91] James L. Helman and Lambertus Hesselink. Visualizing vector field topology in fluid flows. IEEE Computer Graphics and Applications, 11(3):36–46, May/ June 1991. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Haines:1992:CGH

[HH92a] Kenneth Haines and Debby Haines. Computer graphics for holography. IEEE Computer Graphics and Applications, 12(1):37–46, January/ February 1992. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Hanson:1992:IFD

[HH92b] Andrew J. Hanson and Pheng A. Heng. Illuminating the fourth dimension. IEEE Computer Graphics and Applications, 12(4): 54–??, July/August 1992. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Hirota:1995:PFF

[HH95] Koichi Hirota and Michitaka Hirose. Providing force feedback in virtual environments. IEEE Computer Graphics and Applications, 15(5):22–30, September 1995. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Hu:2001:PFB

[HH01] Changyuan Hu and Roger D. Hersch. Parameterizable fonts based on shape components. IEEE Computer Graphics and Applications, 21(3):70–85, May/ June 2001. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

                                        computer.org/cga/cg2001/
                                        g3070abs.htm;   http:/
                                        /dlib.computer.org/cg/
                                        books/cg2001/pdf/g3070.
                                        pdf.

Haroz:2008:SDB

[HH08] Steve Haroz and Katrin Heitmann. Seeing the difference between cosmological simulations. IEEE Computer Graphics and Applications, 28(5):37–45, September/October 2008. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Hogan:2020:DP

[HHH^{+}20] T. Hogan, U. Hinrichs, S. Huron, J. Alexander, and Y. Jansen. Data physicalization. IEEE Computer Graphics and Applications, 40(6):21–24, November/December 2020. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Hogan:2021:DPP

[HHH^{+}21] T. Hogan, U. Hinrichs, S. Huron, J. Alexander, and Y. Jansen. Data physicalization Part II. IEEE Computer Graphics and Applications, 41(1):63–64, January/ February 2021. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Hoffmann:1989:RSO

[HHK89] Christoph M. Hoffmann, John E. Hopcroft, and Michael J. Karasick. Robust set operations on polyhedral solids. IEEE Computer Graphics and Applications, 9(6):50–59, November/December 1989. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Holden:2017:FNS

[HHKK17] Daniel Holden, Ikhsanul Habibie, Ikuo Kusajima, and Taku Komura. Fast neural style transfer for motion data. IEEE Computer Graphics and Applications, 37(4):42–49, July/ August 2017. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2017/04/mcg2017040042-
abs.html.

Han:2014:IDP

[HHLL14] Jaehyun Han, Seongkook Heo, Hyong-Euk Lee, and Geehyuk Lee. The IrPen: A 6-DOF pen for interaction with tablet computers. IEEE Computer Graphics and Applications, 34 (3):22–29, May/June 2014. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Hagen:1992:SIA

[HHS^{+}92] Hans Hagen, Stefanie Hahmann, Thomas Schreiber, Yasuo Nakajima, and Burkard Wordenweber. Surface interrogation algorithms. IEEE Computer Graphics and Applications, 12(5):53–??, September 1992. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Hansen:2020:VPS

[HHY20] C. Hansen, I. Hotz, and A. Ynnerman. Visualization in public spaces. IEEE Computer Graphics and Applications, 40(2):16–17, March/ April 2020. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Hibbard:2004:TFP

[Hib04]  Bill Hibbard.  The top

five problems that motivated my work. IEEE Computer Graphics and Applications, 24(6):9–13, November/December 2004. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http:

//csdl.computer.org/dl/
mags/cg/2004/06/g6009.
htm;   http://csdl.computer.
org/dl/mags/cg/2004/06/
g6009.pdf.

Hillyard:1982:BGS

[Hil82] R. Hillyard. The build group of solid modelers. IEEE Computer Graphics and Applications, 2(3):43– 57, March/April 1982. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Hoxie:1998:PVD

[HILW98] Sue Hoxie, Gil Irizarry, Ben Lubetsky, and Darren Wetzel. Projects in VR: Developments in standards for networked virtual reality. IEEE Computer Graphics and Applications, 18(2):6–9, March/April 1998. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://dlib.
computer.org/cg/books/
cg1998/pdf/g2006.pdf.

Hansen:2003:GEI

[HJ03] Charles Hansen and Chris Johnson. Guest Editors’ introduction: Graphics applications for Grid computing. IEEE Computer Graphics and Applications, 23(2):20– 21, March/April 2003. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http:

                                        //dlib.computer.org/cg/
                                        books/cg2003/pdf/g2020.
                                        pdf.

Hong:2006:FVP

[HJCW06] Min Hong, Sunhwa Jung, Min-Hyung Choi, and Samuel W. J. Welch. Fast volume preservation for a massspring system. IEEE Computer Graphics and Applications, 26(5):83–91, September/October 2006. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Hobona:2006:WBV

[HJF06] Gobe Hobona, Philip James, and David Fairbairn. Webbased visualization of 3D geospatial data using Java3D. IEEE Computer Graphics and Applications, 26(4): 28–33, July/August 2006. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Henderson:1986:CGM

[HJO86] Lofton Henderson, Margaret Journey, and Chris Osland. The computer graphics metafile. IEEE Computer Graphics and Applications, 6(8):24–32, August 1986. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Hereld:2000:TIB

[HJS00] Mark Hereld, Ivan R. Judson, and Rick L. Stevens. Tutorial: Introduction to building projection-based tiled display systems. IEEE Computer Graphics and Applications, 20(4):22–28, July/August 2000. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

computer.org/cga/cg2000/
g4022abs.htm;   http:/
/dlib.computer.org/cg/
books/cg2000/pdf/g4022.
pdf.

Heiter:2022:ERL

[HKDL22] Edith Heiter, Bo Kang, Tijl De Bie, and Jefrey Lijffijt. Evaluating representation learning and graph layout methods for visualization. IEEE Computer Graphics and Applications, 42(3):19–28, May/ June 2022. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Kim:2017:FBA

[hKKSK17] Yun hyeong Kim, Taesoo Kwon, Daeun Song, and Young J. Kim. Full-body animation of human locomotion in reduced gravity using physics-based control. IEEE Computer Graphics and Applications, 37(6):28–39, November/ December 2017. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2017/06/mcg2017060028-
abs.html.

Hwang:2017:PBA

[HKSK17] Jaepyung Hwang, Kwanguk Kim, Il Hong Suh, and Taesoo Kwon. Performancebased animation using constraints for virtual object manipulation. IEEE Computer Graphics and Applications, 37(4):95–102, July/ August 2017. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2017/04/mcg2017040095-
abs.html.

Hinneburg:1999:HEV

[HKW99] Alexander Hinneburg, Daniel A. Keim, and Markus Wawryniuk. HD-Eye: Visual mining of high-dimensional data. IEEE Computer Graphics and Applications, 19(5):22–31, September/ October 1999. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

computer.org/cga/cg1999/
g5022abs.htm;   http:/
/dlib.computer.org/cg/
books/cg1999/pdf/g5022.
pdf.

Huang:2007:DSL

[HLB^{+}07] Jian Huang, Huadong Liu, Micah Beck, Andrew Gaston, Jinzhu Gao, and Terry Moore. Dynamic sharing of large-scale visualization. IEEE Computer Graphics and Applications, 27(1):20– 25, January/February 2007. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Huang:2011:CAD

[HLC11] Chung-Neng Huang, Ming-Yih Lee, and Chong-Ching Chang. Computer-aided design and manufacturing of customized insoles. IEEE Computer Graphics and Applications, 31(2):74–79, March/April 2011. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Hillaire:2008:DFB

[HLCC08] Sébastien Hillaire, Anatole Lécuyer, Rémi Cozot, and Géry Casiez. Depth-of-field blur effects for first-person navigation in virtual environments. IEEE Computer Graphics and Applications, 28(6):47–55, November/December 2008. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Hopf:2004:HSS

[HLE04] Matthias Hopf, Michael Luttenberger, and Thomas Ertl. Hierarchical splatting of scattered 4D data. IEEE Computer Graphics and Applications, 24(4):64– 72, July/August 2004. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http:

//csdl.computer.org/dl/
mags/cg/2004/04/g4064.
htm;   http://csdl.computer.
org/dl/mags/cg/2004/04/
g4064.pdf.

Hu:2013:SBN

[HLW13] Xianlin Hu, Aidong Lu, and Xintao Wu. Spectrumbased network visualization for topology analysis. IEEE Computer Graphics and Applications, 33(1):58– 68, January/February 2013. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Hodges:1985:SAT

[HM85] Larry F. Hodges and David F. McAllister. Stereo and alternating-pair techniques for display of computergenerated images. IEEE Computer Graphics and Applications, 5(9):38–45, September 1985. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Hackstadt:1995:VBV

[HM95] S. T. Hackstadt and A. D. Malony. Visualization blackboard: Visualizing parallel programs and performance. IEEE Computer Graphics and Applications, 15(4): 12–14, July/August 1995. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Hurtienne:2020:MFV

[HMC^{+}20] J. Hurtienne, F. Maas, A. Carolus, D. Reinhardt, C. Baur, and C. Wienrich. Move find: The value of kinaesthetic experience in a casual data representation. IEEE Computer Graphics and Applications, 40(6): 61–75, November/December 2020. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Hentschel:2018:AV

[HMHM18] Bernd Hentschel, Miriah Meyer, Hans Hagen, and Ross Maciejewski. Applied visualization. IEEE Computer Graphics and Applications, 38(3):30–32, May/ June 2018. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2018/03/mcg2018030030.
html.

Hahn:1983:CFP

[HML83] G. J. Hahn, C. B. Morgan, and W. E. Lorensen. Color face plots for displaying product performance. IEEE Computer Graphics and Applications, 3(1):23–26, 28– 29, January/February 1983. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Huang:2006:SSF

[HMRS06] Elaine M. Huang, Elizabeth D. Mynatt, Daniel M. Russell, and Alison E. Sue. Secrets to success and fatal flaws: The design of largedisplay groupware. IEEE Computer Graphics and Applications, 26(1):37–45, January/February 2006. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Huitric:1985:BSS

[HN85] Herve Huitric and Monique Nahas. B-spline surfaces: a tool for computer painting. IEEE Computer Graphics and Applications, 5(3): 39–47, March/April 1985. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Hofsetz:2004:IBR

[HNC^{+}04] Christian Hofsetz, Kim Ng, George Chen, Peter McGuinness, Nelson Max, and Yang Liu. Imagebased rendering of range data with estimated depth uncertainty. IEEE Computer Graphics and Applications, 24(4):34–42, July/ August 2004. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http:/

/csdl.computer.org/dl/
mags/cg/2004/04/g4034.
htm;   http://csdl.computer.
org/dl/mags/cg/2004/04/
g4034.pdf.

Herman:1991:VOM

[HO91] Gabor T. Herman and Dewey Odhner. Visualization optimization by multiple processors. IEEE Computer Graphics and Applications, 11(6):13–15, November/December 1991. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Hobbs:1981:CGD

[Hob81] L. C. Hobbs. Computer graphics display hardware. IEEE Computer Graphics and Applications, 1(1):25– 28, 30–32, 34, 36, 38– 39, January/February 1981. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Hodges:1992:TTS

[Hod92] Larry F. Hodges. Tutorial:

Time-multiplexed stereoscopic computer graphics. IEEE Computer Graphics and Applications, 12(2): 20–30, March/April 1992. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Hoetzlein:2012:GPR

[Hoe12] Rama C. Hoetzlein. Graphics performance in rich Internet applications. IEEE Computer Graphics and Applications, 32(5):98–104, September/October 2012. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Hoffmann:1993:ICS

[Hof93] Christoph M. Hoffmann. Implicit curves and surfaces in CAGD. IEEE Computer Graphics and Applications, 13(1):79–88, January/ February 1993. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Hope:1983:SB

[Hop83] Thomas W. Hope. The slide business. IEEE Computer Graphics and Applications, 3(4):55–56, 58, 59, July/August 1983. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Hornung:1984:MSV

[Hor84]  Christoph Hornung.  A

method for solving the visibility problem. IEEE Computer Graphics and Applications, 4(7):26–33, July/ August 1984. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Howard:2000:ACF

[How00] Caleb Howard. About the cover: Finding the secrets of (artificial) life. IEEE Computer Graphics and Applications, 20(1):4–5, January/ February 2000. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://dlib.
computer.org/cg/books/
cg2000/pdf/g1004.pdf.

Hu:1991:FPU

[HP91] Jianying Hu and Theo Pavlidis. Function plotting using conic splines. IEEE Computer Graphics and Applications, 11(1):89– 94, January/February 1991. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Hachet:2008:ECM

[HPG08] Martin Hachet, Joachim Pouderoux, and Pascal Guitton. 3D elastic control for mobile devices. IEEE Computer Graphics and Applications, 28(4):58–62, July/ August 2008. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Host:2020:NPE

[HPS20] G. Höst, K. Palmerius, and K. Schönborn. Nano for the public: An exploranation perspective. IEEE Computer Graphics and Applications, 40(2):32–42, March/ April 2020. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Hesselink:1994:RIV

[HPvW94] L. Hesselink, F. H. Post, and J. J. van Wijk. Research issues in vector and tensor field visualization. IEEE Computer Graphics and Applications, 14(2): 76–79, March/April 1994. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Han:2021:VAA

[HPX^{+}21] Dongming Han, Jiacheng Pan, Cong Xie, Xiaodong Zhao, Xiaonan Luo, and Wei Chen. A visual analytics approach for structural differences among graphs via deep learning. IEEE Computer Graphics and Applications, 41(5):18–31, September/October 2021. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Han:1998:FRC

[HR98] JungHyun Han and Aristides A. G. Requicha. Feature recognition from CAD models. IEEE Computer Graphics and Applications, 18(2):80–94, March/ April 1998. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

computer.org/cga/cg1998/
g2080abs.htm;   http:/
/dlib.computer.org/cg/
books/cg1998/pdf/g2080.
pdf.

Hincapie-Ramos:2016:GAI

[HROIK16] Juan David Hincapie-Ramos, Kasim Ozacar, Pourang P. Irani, and Yoshifumi Kitamura. GyroWand: An approach to IMU-based raycasting for augmented reality. IEEE Computer Graphics and Applications, 36(2):90–96, March/ April 2016. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2016/02/mcg2016020090-
abs.html.

Hardwick:1987:CSD

[HS87] Martin Hardwick and David L.

Spooner. Comparison of some data models for engineering objects. IEEE Computer Graphics and Applications, 7(3):56–66, March/ April 1987. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Hanson:2001:STB

[HS01] Mary Hanson and Emmanouil Skoufos. Speeding things up: Bicycle racing uses computers before reaching the starting line. IEEE Computer Graphics and Applications, 21(5):14–19, September/October 2001. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://dlib.
computer.org/cg/books/
cg2001/g5014abs.htm;
http://dlib.computer.org/
cg/books/cg2001/pdf/g5014.
pdf.

Hale:2004:DHD

[HS04] Kelly S. Hale and Kay M. Stanney. Deriving haptic design guidelines from human physiological, psychophysical, and neurological foundations. IEEE Computer Graphics and Applications, 24(2):33–39, March/ April 2004. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cg/2004/02/g2033abs.htm;
 http://csdl.computer.
org/dl/mags/cg/2004/02/
g2033.htm;   http://csdl.
computer.org/dl/mags/cg/
2004/02/g2033.pdf.

House:1998:VB

[HSAKD98] Donald H. House, Greg S. Schmidt, Scott A. Arvin, and Midori Kitagawa-DeLeon. Visualization blackboard. IEEE Computer Graphics and Applications, 18(1):12– 14, January/February 1998. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL

http://computer.org/cga/
cg1998/g1012abs.htm;
http://dlib.computer.org/
cg/books/cg1998/pdf/g1012.
pdf.

Hughes:2005:MRE

[HSHS05] Charles E. Hughes, Christopher B. Stapleton, Darin E. Hughes, and Eileen M. Smith. Mixed reality in education, entertainment, and training. IEEE Computer Graphics and Applications, 25(6):24–30, November/December 2005. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Hsu:2002:AMT

[HSK02] Steve Hsu, Harpreet S.

Sawhney, and Rakesh Kumar. Automated mosaics via topology inference. IEEE Computer Graphics and Applications, 22(2):44–54, March/April 2002. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

computer.org/cga/cg2002/
g2044abs.htm;   http:/
/dlib.computer.org/cg/
books/cg2002/pdf/g2044.
pdf.

Hsu:1998:CBF

[HT98] Kun Lung Hsu and Der Min Tsay. Corner blending of free-form N-sided holes. IEEE Computer Graphics and Applications, 18(1):72– 78, January/February 1998. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL

http://computer.org/cga/
cg1998/g1072abs.htm;
http://dlib.computer.org/
cg/books/cg1998/pdf/g1072.
pdf.

Hetzler:2004:AEU

[HT04] Elizabeth Hetzler and Alan Turner. Analysis experiences using information visualization. IEEE Computer Graphics and Applications, 24(5):22–26, September/October 2004. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http:

//csdl.computer.org/dl/
mags/cg/2004/05/g5022.
htm;   http://csdl.computer.
org/dl/mags/cg/2004/05/
g5022.pdf.

Han:2019:FFR

[HTZ^{+}19] J. Han, J. Tao, H. Zheng, H. Guo, D. Z. Chen, and C. Wang. Flow field reduction via reconstructing vector data from 3-D streamlines using deep learning. IEEE Computer Graphics and Applications, 39(4): 54–67, July/August 2019. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Herman:1983:DDI

[HU83] Gabor T. Herman and Jayaram K. Udupa. Display of 3-D digital images: Computational foundations and medical applications. IEEE Computer Graphics and Applications, 3(5):39–46, August 1983. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Huang:2011:BIE

[Hua11] Mary Huang. Bold, italic, emphatic — possibilities for interactive type. IEEE Computer Graphics and Applications, 31(3):8–13, May/ June 2011. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Huang:2021:MHV

[Hua21] K. T. Huang. Mapping the hazard: Visual analysis of flood impact on urban mobility. IEEE Computer Graphics and Applications, 41(1):26–34, January/ February 2021. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Huelskoetter:1983:ECP

[Hue83] Wayne R. Huelskoetter. The evolution of computerized presentation graphics at DI-COMED. IEEE Computer Graphics and Applications, 3(4):15–18, 22, 23, July/ August 1983. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Hursh:1982:SMC

[Hur82] Ronald E. Hursh. Starting and maintaining a computer-aided design system. IEEE Computer Graphics and Applications, 2(1):15–18, 20–22, January/ February 1982. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Haworth:2021:GCD

[HUS^{+}21] Brandon Haworth, Muhammad Usman, Davide Schaumann, Nilay Chakraborty, Glen Berseth, Petros Faloutsos, and Mubbasir Kapadia. Gamification of crowddriven environment design. IEEE Computer Graphics and Applications, 41(4): 107–117, July/August 2021. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Haber:1982:PCC

[HW82a] Ralph Norman Haber and Leland Wilkinson. Perceptual components of computer displays. IEEE Computer Graphics and Applications, 2(3):23–26, 28, 30, 32–35, May/June 1982. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Hopgood:1982:PAR

[HW82b] F. R. A. Hopgood and Robert W. Witty. Perq and advanced raster graphics workstations. IEEE Computer Graphics and Applications, 2(7):9–10, 12, 14–15, September 1982. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Hall:1990:API

[HW90] Mark Hall and Joe Warren. Adaptive polygonalization of implicitly defined surfaces. IEEE Computer Graphics and Applications, 10(6):33–42, November/December 1990. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Hsu:1995:SDA

[HW95] Siu-Chi Hsu and Tien-Tsin Wong. Simulating dust accumulation. IEEE Computer Graphics and Applications, 15(1):18–22, January/ February 1995. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Hull:2018:DTF

[HW18] Carmen Hull Hull and Wesley Willett. Data tectonics: A framework for building physical and immersive data representations. IEEE Computer Graphics and Applications, 38(5):11–17, September/October 2018. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2018/05/mcg2018050011-
abs.html.

Hodges:1996:VCF

[HWK^{+}96] Larry F. Hodges, Benjamin A. Watson, G. Drew Kessler, Barbara O. Rothbaum, and Dan Opdyke. Virtually conquering fear of flying. IEEE Computer Graphics and Applications, 16(6):42–49, November/December 1996. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Haleem:2019:ERF

[HWP^{+}19] H. Haleem, Y. Wang, A. Puri, S. Wadhwa, and H. Qu. Evaluating the readability of force directed graph layouts: A deep learning approach. IEEE Computer Graphics and Applications, 39(4):40–53, July/ August 2019. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Han:2016:HMR

[HWY16] Xiaoguang Han, Kwan-Yee K. Wong, and Yizhou Yu. 3D human model reconstruction from sparse uncalibrated views. IEEE Computer Graphics and Applications, 36(6):46–56, November/December 2016. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL https:/

/www.computer.org/csdl/
mags/cg/2016/06/mcg2016060046-
abs.html.

Han:2022:NSV

[HYL^{+}22] Xiaoyang Han, Xiaomin Yu, Guan Li, Jun Liu, Ying Zhao, and Guihua Shan. Narrative In Situ visual analysis for large-scale ocean eddy evolution. IEEE Computer Graphics and Applications, 42(3):65–73, May/ June 2022. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Hu:2003:ALS

[HYN03] Jinhui Hu, Suya You, and Ulrich Neumann. Approaches to large-scale urban modeling. IEEE Computer Graphics and Applications, 23(6):62–69, November/December 2003. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http:

//csdl.computer.org/dl/
mags/cg/2003/06/g6062.
pdf;   http://csdl.computer.
org/dl/mags/cg/2003/06/
g6062abs.htm.

Hertzmann:2011:PBC

[HZ11] Aaron Hertzmann and Victor Zordan. Physics-based characters. IEEE Computer Graphics and Applications, 31(4):20–21, July/ August 2011. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Herman:1992:SI

[HZB92] Gabor T. Herman, Jingsheng Zheng, and Carolyn A. Bucholtz. Shapebased interpolation. IEEE Computer Graphics and Applications, 12(3):69–79, May/June 1992. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Harless:2003:VCI

[HZHD03] William G. Harless, Marcia A. Zier, Michael G. Harless, and Robert C. Duncan. Virtual conversations: An interface to knowledge. IEEE Computer Graphics and Applications, 23(5):46–52, September/October 2003. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http:

                                        //csdl.computer.org/dl/
                                        mags/cg/2003/05/g5046.
                                        pdf;   http://csdl.computer.
                                        org/dl/mags/cg/2003/05/
                                        g5046abs.htm.

Anjyo:2003:SHC

[iAH03] Ken ichi Anjyo and Katsuaki Hiramitsu. Stylized highlights for cartoon rendering and animation. IEEE Computer Graphics and Applications, 23(4):54– 61, July/August 2003. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

csdl.computer.org/comp/
mags/cg/2003/04/extras/
g4054.htm;   http://csdl.
computer.org/comp/mags/
cg/2003/04/g4054abs.htm;
 http://csdl.computer.
org/dl/mags/cg/2003/04/
g4054.htm;   http://csdl.
computer.org/dl/mags/cg/
2003/04/g4054.pdf.

Isenberg:2009:CCR

[IBH^{+}09] Petra Isenberg, Anastasia Bezerianos, Nathalie Henry, Sheelagh Carpendale, and Jean-Daniel Fekete. Co-CoNutTrix: Collaborative retrofitting for information visualization. IEEE Computer Graphics and Applications, 29(5):44–57, September/October 2009. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Ivanov:2019:WAD

[IDJW19] A. Ivanov, K. Danyluk, C. Jacob, and W. Willett. A walk among the data. IEEE Computer Graphics and Applications, 39(3):19– 28, May/June 2019. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Isenberg:2003:DGS

[IFH^{+}03] Tobias Isenberg, Bert Freudenberg, Nick Halper, Stefan Schlechtweg, and Thomas Strothotte. A developer’s guide to silhouette algorithms for polygonal models. IEEE Computer Graphics and Applications, 23(4):28–37, July/ August 2003. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL http://csdl.
                                        computer.org/comp/mags/
                                        cg/2003/04/g4028abs.htm;
                                         http://csdl.computer.
                                        org/dl/mags/cg/2003/04/
                                        g4028.htm;   http://csdl.
                                        computer.org/dl/mags/cg/
                                        2003/04/g4028.pdf.

Interrante:1998:VF

[IG98] Victoria Interrante and Chester Grosch. Visualizing 3D flow. IEEE Computer Graphics and Applications, 18(4):49–53, July/ August 1998. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

                                        computer.org/cga/cg1998/
                                        g4049abs.htm.

Igarashi:2011:DDE

[Iga11] Yuki Igarashi. Deco: a design editor for rhinestone decorations. IEEE Computer Graphics and Applications, 31(5):90–94, September/October 2011. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Igarashi:2019:BIM

[Iga19] Y. Igarashi. BandWeavy:

Interactive modeling for craft band design. IEEE Computer Graphics and Applications, 39(5):96–103, September/October 2019. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Interrante:2018:VAR

[IHL18] Victoria Interrante, Tobias Hollerer, and Anatole Lecuyer. Virtual and augmented reality. IEEE Computer Graphics and Applications, 38(2):28–30, March/ April 2018. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2018/02/mcg2018020028.
html.

Igarashi:2010:HDE

[II10] Yuki Igarashi and Takeo Igarashi. Holly: a drawing editor for designing stencils. IEEE Computer Graphics and Applications, 30 (4):8–14, July/August 2010. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Igarashi:2015:EAI

[II15] Takeo Igarashi and Masahiko Inami. Exploration of alternative interaction techniques for robotic systems. IEEE Computer Graphics and Applications, 35(3):33– 41, May/June 2015. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http:/

/www.computer.org/csdl/
mags/cg/2015/03/mcg2015030033-
abs.html.

Ikeda:2022:PSD

[II22] Yuki Ikeda and Yuki Igarashi. Podiy: a system for design and production of pouches by novices. IEEE Computer Graphics and Applications, 42(2):81–88, March/ April 2022. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Isenberg:2013:DVI

[IIH^{+}13] Petra Isenberg, Tobias Isenberg, Tobias Hesselmann, Bongshin Lee, Ulrich von Zadow, and Anthony Tang. Data visualization on interactive surfaces: A research agenda. IEEE Computer Graphics and Applications, 33(2):16–24, March/ April 2013. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Itoh:2015:KNS

[IK15] Takayuki Itoh and Karsten Klein. Key-node-separated graph clustering and layouts for human relationship graph visualization. IEEE Computer Graphics and Applications, 35(6):30–40, November/December 2015. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL

http://www.computer.org/
csdl/mags/cg/2015/06/mcg2015060030-
abs.html.

Ikedo:1984:HTC

[Ike84] Tsuneo Ikedo. High-speed techniques for a 3-D color graphics terminal. IEEE Computer Graphics and Applications, 4(5):46–58, May/ June 1984. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Iizuka:2012:EMS

[IKMF12] Satoshi Iizuka, Yoshihiro Kanamori, Jun Mitani, and Yukio Fukui. Efficiently modeling 3D scenes from a single image. IEEE Computer Graphics and Applications, 32(6):18–25, November/December 2012. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Iones:2003:FRL

[IKSZ03] Andrey Iones, Anton Krupkin, Mateu Sbert, and Sergey Zhukov. Fast, realistic lighting for video games. IEEE Computer Graphics and Applications, 23(3):54– 64, May/June 2003. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http:

//dlib.computer.org/cg/
books/cg2003/pdf/g3054.
pdf;   http://www.computer.
org/cga/cg2003/g3054abs.
htm.

Ida:1983:SEN

[IL83] Nathan Ida and William Lord. Simulating electromagnetic NDT probe fields. IEEE Computer Graphics and Applications, 3(3):21– 28, May/June 1983. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Isenburg:2010:PSG

[ILC10] Martin Isenburg, Peter Lindstrom, and Hank Childs. Parallel and streaming generation of ghost data for structured grids. IEEE Computer Graphics and Applications, 30(3):32–44, May/June 2010. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Ikedo:1998:TSR

[IM98] Tsuneo Ikedo and Jianhua Ma. The Truga001: a scalable rendering processor. IEEE Computer Graphics and Applications, 18(2):59– 79, March/April 1998. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

computer.org/cga/cg1998/
g2059abs.htm;   http:/
/dlib.computer.org/cg/
books/cg1998/pdf/g2059.
pdf.

Igarashi:2011:DCA

[IM11] Takeo Igarashi and Radomir Mech. Digital-content authoring. IEEE Computer Graphics and Applications, 31(6):16–17, November/ December 2011. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Igarashi:2014:WIC

[IM14] Yuki Igarashi and Jun Mitani. Weavy: Interactive card-weaving design and construction. IEEE Computer Graphics and Applications, 34(4):22–29, July/ August 2014. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Itoh:2003:GOT

[IMS03] Takayuki Itoh, Kazunori Miyata, and Kenji Shimada. Generating organic textures with controlled anisotropy and directionality. IEEE Computer Graphics and Applications, 23(3):38–45, May/June 2003. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL http://dlib.
                                        computer.org/cg/books/
                                        cg2003/pdf/g3038.pdf;
                                        http://www.computer.org/
                                        cga/cg2003/g3038abs.htm.

Interrante:2000:VVH

[Int00] Victoria Interrante. Visualization viewpoints: Harnessing natural textures for multivariate visualization. IEEE Computer Graphics and Applications, 20(6):6–11, November/December 2000. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http:

                                        //dlib.computer.org/cg/
                                        books/cg2000/pdf/g6006.
                                        pdf.

Inui:2007:UGA

[IO07] Masatomo Inui and Atsushi Ohta. Using a GPU to accelerate die and mold fabrication. IEEE Computer Graphics and Applications, 27(1):82–88, January/ February 2007. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Illert:1992:GIO

[IP92] Chris Illert and Clifford A. Pickover. Generating irregularly oscillating fossil seashells. IEEE Computer Graphics and Applications, 12(3):18–22, May/ June 1992. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Illert:1995:NAS

[IS95] Chris Illert and Dave Sims.

In the news: Australian supercomputer graphics exhibition; VR World 95: The new element in VR is people. IEEE Computer Graphics and Applications, 15(4): 89–91, July/August 1995. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Ingulfsen:2022:NGV

[ISMGG22] Nicholas Ingulfsen, Simone Schaub-Meyer, Markus Gross, and Tobias Günther. News globe: Visualization of geolocalized news articles. IEEE Computer Graphics and Applications, 42(4): 40–51, July/August 2022. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Itoh:2006:HVN

[ITSK06] Takayuki Itoh, Hiroki Takakura, Atsushi Sawada, and Koji Koyamada. Hierarchical visualization of network intrusion detection data. IEEE Computer Graphics and Applications, 26(2):40–47, March/April 2006. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Irani:2001:UPS

[ITW01] Pourang Irani, Maureen Tingley, and Colin Ware. Using perceptual syntax to enhance semantic content in diagrams. IEEE Computer Graphics and Applications, 21(5):76–85, September/October 2001. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://dlib.
computer.org/cg/books/
cg2001/g5076abs.htm;
http://dlib.computer.org/
cg/books/cg2001/pdf/g5076.
pdf.

Iwata:1999:TTR

[Iwa99] Hiroo Iwata. The torus treadmill: Realizing locomotion in VEs. IEEE Computer Graphics and Applications, 19(6):30–35, November/December 1999. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

computer.org/cga/cg1999/
g6030abs.htm;   http:/
/dlib.computer.org/cg/
books/cg1999/pdf/g6030.
pdf.

Ikei:1997:VTD

[IWF97] Yasushi Ikei, Kazufumi Wakamatsu, and Shuichi Fukuda. Vibratory tactile display of image-based textures. IEEE Computer Graphics and Applications, 17(6):53–61, November/December 1997. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

computer.org/cga/cg1997/
g6053abs.htm.

Iwata:2005:C

[IYFN05] Hiroo Iwata, Hiroaki Yano, Hiroyuki Fukushima, and Haruo Noma. CirculaFloor. IEEE Computer Graphics and Applications, 25(1):64– 67, January/February 2005. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Igarashi:2007:GEI

[IZ07] Takeo Igarashi and Bob Zeleznik. Guest Editors’ introduction: Sketch-based interaction. IEEE Computer Graphics and Applications, 27(1):26–27, January/ February 2007. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cg/2007/01/g1026.pdf.

Juurlink:2008:GFL

[JAC^{+}08] Ben Juurlink, Iosif Antochi, Dan Crisu, Sorin Cotofana, and Stamatis Vassiliadis. GRAAL: a framework for low-power 3D graphics accelerators. IEEE Computer Graphics and Applications, 28(4):63–73, July/ August 2008. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Jacobson:2015:BLS

[Jac15] Alec Jacobson. Breathing life into shapes. IEEE Computer Graphics and Applications, 35(5):92–100, September/October 2015. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL http://www.
                                        computer.org/csdl/mags/
                                        cg/2015/05/mcg2015050092-
                                        abs.html.

Jaenichen:2017:VCC

[Jae17] Claudine Jaenichen. Visual communication and cognition in everyday decisionmaking. IEEE Computer Graphics and Applications, 37(6):10–18, November/ December 2017. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        cg/2017/06/mcg2017060010-
                                        abs.html.

Jain:2000:RR

[Jai00] Ramesh Jain. Real reality. IEEE Computer Graphics and Applications, 20(1):40– 41, January/February 2000. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL

                                        http://dlib.computer.org/
                                        cg/books/cg2000/pdf/g1040.
                                        pdf.

James:1987:TGP

[Jam87] J. Michael James. The third generation of PC graphics controllers. IEEE Computer Graphics and Applications, 7(10):24–27, October 1987. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Juan:2005:UAR

[JAM^{+}05] M. Carmen Juan, Mariano Alcañiz, Carlos Monserrat, Cristina Botella, Rosa M. Baños, and Belen Guerrero. Using augmented reality to treat phobias. IEEE Computer Graphics and Applications, 25(6):31–37, November/December 2005. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Joshi:2007:EDM

[JAPF07] Neel Joshi, Wojciech Matusik Shai Avidan, Hanspeter Pfister, and William T. Freeman. Exploring defocus matting: Nonparametric acceleration, superresolution, and off-center matting. IEEE Computer Graphics and Applications, 27(2):43–52, March/April 2007. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Jarett:1983:UCP

[Jar83] Irwin M. Jarett. Using charts to present international trade data. IEEE Computer Graphics and Applications, 3(9):55–61, December 1983. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Jin:1997:TBC

[JB97] Liqun Jin and David C. Banks. TennisViewer: a browser for competition trees. IEEE Computer Graphics and Applications, 17(4):63–65, July/August 1997. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Julier:2002:GEI

[JB02] Simon Julier and Gary Bishop. Guest Editors’ introduction: Tracking: How hard can it be? IEEE Computer Graphics and Applications, 22(6):22–23, November/December 2002. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http:

//dlib.computer.org/cg/
books/cg2002/pdf/g6022.
pdf.

Julier:2002:IFM

[JBBL02] Simon Julier, Yohan Baillot, Dennis Brown, and Marco Lanzagorta. Information filtering for mobile augmented reality. IEEE Computer Graphics and Applications, 22(5):12–15, September/October 2002. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://dlib.
computer.org/cg/books/
cg2002/pdf/g5012.pdf.

Janzen:2023:JPE

[JBCS23] Kyle Janzen, Chris Burke, Bruce D. Campbell, and Francesca Samsel. JNZN-BRK: Physical experiments in light, modulation, and substrate. IEEE Computer Graphics and Applications, 43(6):126–132, November/ December 2023. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Jena:2021:NBU

[JBD^{+}21] A. Jena, M. Butler, T. Dwyer, K. Ellis, U. Engelke, R. Kirkham, K. Marriott, C. Paris, and V. Rajamanickam. The next billion users of visualization. IEEE Computer Graphics and Applications, 41(2):8–16, March/ April 2021. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Jeong:2010:SNI

[JBH^{+}10] Won-Ki Jeong, Johanna Beyer, Markus Hadwiger, Rusty Blue, Charles Law, Amelio Vázquez-Reina, R. Clay Reid, Jeff Lichtman, and Hanspeter Pfister. Ssecrett and NeuroTrace: Interactive visualization and analysis tools for large-scale neuroscience data sets. IEEE Computer Graphics and Applications, 30(3):58–70, May/June 2010. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Jin:2007:CGM

[JCM07] Xiaogang Jin, Shaochun Chen, and Xiaoyang Mao. Computer-generated marbling textures: a GPUbased design system. IEEE Computer Graphics and Applications, 27(2):78–84, March/April 2007. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL http://csdl.
                                        computer.org/comp/mags/
                                        cg/2007/extras/g2078x1.
                                        avi.

Jerard:1989:MDE

[JDH^{+}89] Robert B. Jerard, Robert L. Drysdale, Kenneth E. Hauck, Barry Schaudt, and John Magewick. Methods for detecting errors in numerically controlled machining of sculptured surfaces. IEEE Computer Graphics and Applications, 9(1):26– 39, January/February 1989. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Jones:2004:NIP

[JDZ04] Thouis R. Jones, Frédo Durand, and Matthias Zwicker. Normal improvement for point rendering. IEEE Computer Graphics and Applications, 24(4):53–56, July/ August 2004. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http:/

/csdl.computer.org/dl/
mags/cg/2004/04/g4053.
htm;   http://csdl.computer.
org/dl/mags/cg/2004/04/
g4053.pdf.

Jerald:2018:HCV

[Jer18a]  Jason Jerald.  Human-

centered VR design: Five essentials every engineer needs to know. IEEE Computer Graphics and Applications, 38(2):15–21, March/ April 2018. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2018/02/mcg2018020015-
abs.html.

Jerald:2018:TSI

[Jer18b] Jason Jerald. A taxonomy of spatial interaction patterns and techniques. IEEE Computer Graphics and Applications, 38(1):11–19, January/ February 2018. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2018/01/mcg2018010011-
abs.html.

Johnson:1999:FSC

[JF99] Garrett M. Johnson and Mark D. Fairchild. Fullspectral color calculations in realistic image synthesis. IEEE Computer Graphics and Applications, 19(4):47– 53, July/August 1999. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

computer.org/cga/cg1999/
g4047abs.htm;   http:/
/dlib.computer.org/cg/
books/cg1999/pdf/g4047.
pdf.

Jayaram:1999:VVA

[JJW^{+}99] Sankar Jayaram, Uma Jayaram, Yong Wang, Hrishikesh Tirumali, Kevin Lyons, and Peter Hart. VADE: a virtual assembly design environment. IEEE Computer Graphics and Applications, 19(6):44–50, November/December 1999. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

computer.org/cga/cg1998/
g6044abs.htm;   http://
computer.org/cga/cg1999/
g6044xabs.htm;   http:
//dlib.computer.org/cg/
books/cg1999/pdf/g6044.
pdf.

Jerabkova:2009:SCD

[JK09] Lenka Jeřábková and Torsten Kuhlen. Stable cutting of deformable objects in virtual environments using XFEM. IEEE Computer Graphics and Applications, 29(2):61–71, March/ April 2009. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Jankun-Kelly:2003:DWB

[JKKM^{+}03] T. J. Jankun-Kelly, Oliver Kreylos, Kwan-Liu Ma, Bernd Hamann, Kenneth I. Joy, John Shalf, and E. Wes Bethel. Deploying Webbased visual exploration tools on the Grid. IEEE Computer Graphics and Applications, 23(2):40–50, March/April 2003. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://dlib.
computer.org/cg/books/
cg2003/pdf/g2040.pdf;
http://www.computer.org/
cga/cg2003/g2040abs.htm.

Jin:2023:EDS

[JKM23] Yu Jin, Laura Koesten, and Torsten Möller. Exploring the design space of three criteria decision making. IEEE Computer Graphics and Applications, 43(5): 26–38, September/October 2023. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Johnson:1998:PVM

[JLC98] Andy Johnson, Jason Leigh, and Jim Costigan. Projects in VR: Multiway teleimmersion at Supercomputing 97. IEEE Computer Graphics and Applications, 18(4):6–9, July/ August 1998. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Johnson:2002:VH

[JLC^{+}02] Andrew Johnson, Jason Leigh, Bryan Carter, Jim Sosnoski, and Steve Jones. Virtual Harlem. IEEE Computer Graphics and Applications, 22(5):61–67, September/October 2002. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL http://dlib.
                                        computer.org/cg/books/
                                        cg2002/pdf/g5061.pdf;
                                        http://www.computer.org/
                                        cga/cg2002/g5061abs.htm.

Jeong:2010:UCV

[JLJ^{+}10] Byungil Jeong, Jason Leigh, Andrew Johnson, Luc Renambot, Maxine Brown, Ratko Jagodic, Sungwon Nam, and Hyejung Hur. Ultrascale collaborative visualization using a display-rich global cyberinfrastructure. IEEE Computer Graphics and Applications, 30(3):71– 83, May/June 2010. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Johnson:2006:GSV

[JLMV06] Andrew Johnson, Jason Leigh, Paul Morin, and Peter Van Keken. GeoWall:

Stereoscopic visualization for geoscience research and education. IEEE Computer Graphics and Applications, 26(6):10–14, November/December 2006. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Jo:2015:WEW

[JLS15] Jaemin Jo, Bongshin Lee, and Jinwook Seo. Wordle-Plus: Expanding Wordle’s use through natural interaction and animation. IEEE Computer Graphics and Applications, 35(6):20–28, November/ December 2015. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
cg/2015/06/mcg2015060020-
abs.html.

Jacobs:2008:AHD

[JLW08] Katrien Jacobs, Celine Loscos, and Greg Ward. Automatic high-dynamic range image generation for dynamic scenes. IEEE Computer Graphics and Applications, 28(2):84–93, March/ April 2008. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Johannsen:1995:AOM

[JM95] Andreas Johannsen and Robert J. Moorhead II.

AGP: ocean model flow visualization. IEEE Computer Graphics and Applications, 15(4):28–33, July/ August 1995. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Johnson:2002:AES

[JMC^{+}02] Andrew Johnson, Thomas Moher, Yong-Joo Cho, Ya Ju Lin, Dave Haas, and Janet Kim. Augmenting elementary school education with VR. IEEE Computer Graphics and Applications, 22(2):6–9, March/ April 2002. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

computer.org/cga/cg2002/
g2006abs.htm;   http://
computer.org/cga/cg2002/
pdf/g2006.pdf;   http:
//dlib.computer.org/cg/
books/cg2002/pdf/g2006.
pdf.

Johnson:1999:REP

[JMOG99] Andrew Johnson, Thomas Moher, Stellan Ohlsson, and Mark Gillingham. The Round Earth Project — collaborative VR for conceptual learning. IEEE Computer Graphics and Applications, 19(6):60–69, November/December 1999. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

computer.org/cga/cg1999/
g6060abs.htm;   http:/
/dlib.computer.org/cg/
books/cg1999/pdf/g6060.
pdf.

Joslin:2001:SAN

[JMT^{+}01] Chris Joslin, Tom Molet, Nadia Magnenat Thalmann, Joaquim Esmerado, Daniel Thalmann, Ian Palmer, Nicholas Chilton, and Rae Earnshaw. Sharing attractions on the Net with VPark. IEEE Computer Graphics and Applications, 21(1):61–71, January/ February 2001. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

computer.org/cga/cg2001/
g1061abs.htm;   http:/
/dlib.computer.org/cg/
books/cg2001/pdf/g1061.
pdf.

Joel:2013:NSR

[Joe13] William J. Joel. The need to start research education early. IEEE Computer Graphics and Applications, 33(5):79–81, September/October 2013. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Johansson:2017:VVT

[JOG^{+}17] Jimmy Johansson, Tomasz Opach, Erik Glaas, Tina-Simone Neset, Carlo Navarra, Bjorn-Ola Linner, and Jan Ketil Rod. VisAdapt: A visualization tool to support climate change adaptation. IEEE Computer Graphics and Applications, 37(2):54–65, March/ April 2017. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2017/02/mcg2017020054-
abs.html.

Johnson:2004:TSV

[Joh04] Chris Johnson. Top scientific visualization research problems. IEEE Computer Graphics and Applications, 24(4):13–17, July/ August 2004. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http:/

/csdl.computer.org/dl/
mags/cg/2004/04/g4013.
htm;   http://csdl.computer.
org/dl/mags/cg/2004/04/
g4013.pdf.

Jones:1989:CNP

[Jon89] Christopher B. Jones. Cartographic name placement with Prolog. IEEE Computer Graphics and Applications, 9(5):36–47, September 1989. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Jones:2007:GGO

[Jon07] Michael T. Jones. Google’s geospatial organizing principle. IEEE Computer Graphics and Applications, 27 (4):8–13, July/August 2007. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Jager:1993:BMI

[JOnH93] Michael Jager, Utz Osterfeld, Hans-Josef Ackerman n, and Christoph Hornung. Building a multimedia ISDN PC. IEEE Computer Graphics and Applications, 13 (5):24–33, September 1993. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Joshi:2011:CBS

[Jos11] Pushkar Joshi. Curve-based shape modeling — a tutorial. IEEE Computer Graphics and Applications, 31(6): 18–23, November/December 2011. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Jos:2013:PPP

[JPSM13] Rui José, Hölder Pinto, Bruno Silva, and Ana Melro. Pins and posters: Paradigms for content publication on situated displays. IEEE Computer Graphics and Applications, 33(2): 64–72, March/April 2013. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Junior:2017:AAL

[JRM17] Amilcar Soares Junior, Chiara Renso, and Stan Matwin. ANALYTiC: An active learning system for trajectory classification. IEEE Computer Graphics and Applications, 37(5):28–39, September/October 2017. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2017/05/mcg2017050028-
abs.html.

Johnson:2003:NSV

[JS03] Chris R. Johnson and Allen R. Sanderson. A next step: Visualizing errors and uncertainty. IEEE Computer Graphics and Applications, 23(5):6–10, September/October 2003. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http:

//csdl.computer.org/dl/
mags/cg/2003/05/g5006.
pdf.

J:2010:VAF

[JS10] Heike Jänicke and Gerik Scheuermann. Visual analysis of flow features using information theory. IEEE Computer Graphics and Applications, 30(1):40–49, January/February 2010. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Jiang:2021:DFA

[JSDL21] Qi Jiang, Guodao Sun, Yue Dong, and Ronghua Liang. DT2VIS: a Focus+Context answer generation system to facilitate visual exploration of tabular data. IEEE Computer Graphics and Applications, 41(5):45–56, September/October 2021. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Jain:2016:PMS

[JSH16] Eakta Jain, Yaser Sheikh, and Jessica Hodgins. Predicting moves-on-stills for comic art using Viewer gaze data. IEEE Computer Graphics and Applications, 36(4):34–45, July/ August 2016. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2016/04/mcg2016040034-
abs.html.

Jin:2001:PVA

[JW01] Li Jin and Zhigang Wen.

Projects in VR: Adorning VRML worlds with environmental aspect. IEEE Computer Graphics and Applications, 21(1):6–9, January/ February 2001. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://dlib.
computer.org/cg/books/
cg2001/pdf/g1006.pdf.

Ji:2016:EBI

[JW16] Haru Ji and Graham Wakefield. Endogenous biologically inspired art of complex systems. IEEE Computer Graphics and Applications, 36(1):16–21, January/ February 2016. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL http://www.
                                        computer.org/csdl/mags/
                                        cg/2016/01/mcg2016010016-
                                        abs.html.

Jiao:2012:UTV

[JWHW12] Shaohui Jiao, Wen Wu, Pheng-Ann Heng, and Enhua Wu. Using time-varying texels to simulate withering grassland. IEEE Computer Graphics and Applications, 32(1):78–86, January/ February 2012. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Jadoon:2018:IPM

[JWL^{+}18] Aamir Khan Jadoon, Chenming Wu, Yong-Jin Liu, Ying He, and Charlie C. L. Wang. Interactive partitioning of 3D models into printable parts. IEEE Computer Graphics and Applications, 38(4):38–53, July/ August 2018. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/

cg/2018/04/mcg2018040038-
abs.html.

Jimenez:2010:RTR

[JWSG10] Jorge Jimenez, David Whelan, Veronica Sundstedt, and Diego Gutierrez. Realtime realistic skin translucency. IEEE Computer Graphics and Applications, 30(4):32–41, July/August 2010. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Jin:2008:VST

[JXW^{+}08] Xiaogang Jin, Jiayi Xu, Charlie C. L. Wang, Shengsheng Huang, and Jun Zhang. VR software and technology: Interactive control of large-crowd navigation in virtual environments using vector fields. IEEE Computer Graphics and Applications, 28(6):37–46, November/December 2008. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Kahaner:1994:JAV

[Kah94] David Kahaner. Japanese activities in virtual reality. IEEE Computer Graphics and Applications, 14(1):75– 78, January/February 1994. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Kalay:1987:WIG

[Kal87] Yehuda E. Kalay. Worldview: An integrated geometricmodeling/drafting systems. IEEE Computer Graphics and Applications, 7(2):36– 46, February 1987. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Kankaanpaa:1988:FFP

[Kan88] Arto I. Kankaanpaa. FIDS — a Flat-panel Interactive Display System. IEEE Computer Graphics and Applications, 8(2):71–82, March/ April 1988. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
cg/02/00000505-abs.html;
 http://www.computer.org/
csdl/mags/cg/02/mcg1988020071-
abs.html.

Kanyuk:2009:BSF

[Kan09] Paul Kanyuk. Brain springs: Fast physics for large crowds in WALLoE. IEEE Computer Graphics and Applications, 29(4):19–25, July/ August 2009. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Kasik:2000:VFC

[Kas00] David J. Kasik. Viewing the future of CAD. IEEE Computer Graphics and Applications, 20(1):34–35, January/ February 2000. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://dlib.
computer.org/cg/books/
cg2000/pdf/g1034.pdf.

Kasik:2008:TP

[Kas08]  David J. Kasik.  Tools

and products. IEEE Computer Graphics and Applications, 28(1):79–80, January/ February 2008. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://bell.
computer.org/dlcomments/
;   http://csdl.computer.
org/comp/mags/cg/2008/
01/mcg2008010079.pdf.

Kasik:2010:AGTa

[Kas10a] David J. Kasik. Advanced graphics technology. IEEE Computer Graphics and Applications, 30(2):95–96, March/April 2010. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Kasik:2010:AGTb

[Kas10b] David J. Kasik. Advanced graphics technology. IEEE Computer Graphics and Applications, 30(3):96, c3, May/June 2010. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Kasik:2011:AGTa

[Kas11a] David J. Kasik. Advanced graphics technology. IEEE Computer Graphics and Applications, 31(3):96, c3, May/June 2011. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Kasik:2011:AGTb

[Kas11b] David J. Kasik. Advanced graphics technology. IEEE Computer Graphics and Applications, 31(6):c3, November/December 2011. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Kasik:2011:TWC

[Kas11c] David J. Kasik. The third wave in computer graphics and interactive techniques. IEEE Computer Graphics and Applications, 31(4): 89–93, July/August 2011. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Kasik:2012:HS

[Kas12a] David J. Kasik. Hands on at SIGGRAPH 2012. IEEE Computer Graphics and Applications, 32(6):95–96, November/December 2012. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Kasik:2012:LHS

[Kas12b] David J. Kasik. Looking at the hard stuff. IEEE Computer Graphics and Applications, 32(3):94–95, May/ June 2012. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Kasik:2012:TP

[Kas12c]  David J. Kasik.  Tools

and products. IEEE Computer Graphics and Applications, 32(4):95–96, July/ August 2012. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Kasik:2013:SHE

[Kas13] David J. Kasik. Shrinking hardware, expanding interfaces. IEEE Computer Graphics and Applications, 33(1):87–88, January/ February 2013. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Kasik:2017:MJS

[Kas17]  David J. Kasik.  In

memoriam: John Staudhammer. IEEE Computer Graphics and Applications, 37(3):14–15, May/ June 2017. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2017/03/mcg2017030014.
html.

Korein:1982:TGG

[KB82] James U. Korein and Norman I. Badler. Techniques for generating the goaldirected motion of articulated structures. IEEE Computer Graphics and Applications, 2(9):71–74, 76– 81, November/December 1982. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Kaufman:1988:MPA

[KB88] Arie Kaufman and Reuven Bakalash. Memory and processing architecture for 3D voxel-based imagery. IEEE Computer Graphics and Applications, 8(6):10–23, November/December 1988. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Konstantinides:1992:MAI

[KB92] Konstantinos Konstantinides and Vasudev Bhaskaran. Monolithic architectures for image processing and compression. IEEE Computer Graphics and Applications, 12(6):75–86, November/December 1992. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Ko:1996:AHL

[KB96] Hyeongseok Ko and Norman I. Badler. Animating human locomotion with inverse dynamics. IEEE Computer Graphics and Applications, 16(2):50–59, March/ April 1996. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Karahalios:2009:SMS

[KB09] Karrie Karahalios and Tony Bergstrom. Social mirrors as social signals: Transforming audio into graphics. IEEE Computer Graphics and Applications, 29(5): 22–32, September/October 2009. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Koppelhuber:2014:LTF

[KB14] Alexander Koppelhuber and Oliver Bimber. LumiCon-Sense: A transparent, flexible, and scalable thin-film sensor. IEEE Computer Graphics and Applications, 34(5):98–102, September/ October 2014. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://csdl.
computer.org/csdl/mags/
cg/2014/05/mcg2014050098-
abs.html.

Kasik:2005:TCC

[KBF05] David J. Kasik, William Buxton, and David R. Ferguson. Ten CAD challenges. IEEE Computer Graphics and Applications, 25(2): 81–92, March/April 2005. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Knuth:2017:DW

[KBGK17] Martin Knuth, Jan Bender, Michael Goesele, and Arjan Kuijper. Deferred warping. IEEE Computer Graphics and Applications, 37(6):76–87, November/ December 2017. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2017/06/mcg2017060076-
abs.html.

Kandogan:2014:DIW

[KBHP14] Eser Kandogan, Aruna Balakrishnan, Eben M. Haber, and Jeffrey S. Pierce. From data to insight: Work practices of analysts in the enterprise. IEEE Computer Graphics and Applications, 34(5):42–50, September/October 2014. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://csdl.
computer.org/csdl/mags/
cg/2014/05/mcg2014050042-
abs.html.

Koelsch:2006:MIW

[KBHT06] Mathias Koelsch, Ryan Bane, Tobias Hoellerer, and Matthew Turk. Multimodal interaction with a wearable augmented reality system. IEEE Computer Graphics and Applications, 26(3):62–71, May/ June 2006. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Kang:2005:VFM

[KC05] Kai Kang and Rene Chevray.

Visualization of fluid mixing in microchannels. IEEE Computer Graphics and Applications, 25(6):16–20, November/December 2005. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Kaya:2023:EIC

[KCB^{+}23] Furkan Kaya, Elif Celik, Anil Ufuk Batmaz, Aunnoy K. Mutasim, and Wolfgang Stuerzlinger. Evaluation of an immersive COVID-19 data visualization. IEEE Computer Graphics and Applications, 43(1):76–83, January/ February 2023. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Kurc:2001:VLD

[KÇC^{+}01] Tahsin Kurc, Ümit Çatalyürek, Chialin Chang, Alan Sussman, and Joel Saltz. Visualization of large data sets with the Active Data Repository. IEEE Computer Graphics and Applications, 21(4):24–33, July/ August 2001. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

computer.org/cga/cg2001/

                                        g4024abs.htm;   http:/
                                        /dlib.computer.org/cg/
                                        books/cg2001/pdf/g4024.
                                        pdf.

Koumaditis:2020:EVV

[KCMK20] K. Koumaditis, F. Chinello, P. Mitkidis, and S. Karg. Effectiveness of virtual versus physical training: The case of assembly tasks, trainer’s verbal assistance, and task complexity. IEEE Computer Graphics and Applications, 40(5):41–56, September/October 2020. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Karduni:2017:USE

[KCW^{+}17] Alireza Karduni, Isaac Cho, Ginette Wessel, William Ribarsky, Eric Sauda, and Wenwen Dou. Urban Space Explorer: A visual analytics system for urban planning. IEEE Computer Graphics and Applications, 37(5):50–60, September/October 2017. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        cg/2017/05/mcg2017050050-
                                        abs.html.

Kasik:2020:IGI

[KD20] D. J. Kasik and J. C. Dill. Interactive graphics in industry: The early days. IEEE Computer Graphics and Applications, 40(5): 89–99, September/October 2020. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Kosminsky:2021:SVR

[Kd21] D. Kosminsky and D. T. de Oliveira. Slave voyages: Reflections on data sculptures. IEEE Computer Graphics and Applications, 41(1):65– 73, January/February 2021. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Kallman:1995:VTM

[KDH^{+}95a] Jeffrey S. Kallman, Anthony J. De Groot, Carol G. Hoover, William G. Hoover, Susanne M. Lee, and Frederick Wooten. Visualization techniques for molecular dynamics. IEEE Computer Graphics and Applications, 15(6):72–77, November/December 1995. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Kalvin:1995:RHF

[KDH95b] A. D. Kalvin, D. Dean, and J.-J. Hublin. Reconstruction of human fossils. IEEE Computer Graphics and Applications, 15(1):12– 15, January/February 1995. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Kosara:2008:VC

[KDHL08] Robert Kosara, Fritz Drury, Lars Erik Holmquist, and David H. Laidlaw. Visualization criticism. IEEE Computer Graphics and Applications, 28(3):13–15, May/June 2008. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Kyto:2017:ICP

[KDMH17] Mikko Kyto, Krupakar Dhinakaran, Aki Martikainen, and Perttu Hamalainen. Improving 3D character posing with a gestural interface. IEEE Computer Graphics and Applications, 37(1):70–78, January/ February 2017. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        cg/2017/01/mcg2017010070-
                                        abs.html.

Keefe:2010:IVI

[Kee10] Daniel F. Keefe. Integrating visualization and interaction research to improve scientific workflows. IEEE Computer Graphics and Applications, 30(2):8–13, March/ April 2010. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Kelman:1984:DWA

[Kel84] Barry I. Kelman. Distributed workstation architecture: the convergent cluster. IEEE Computer Graphics and Applications, 4(4): 36–38, April 1984. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Kennicott:1981:ITC

[Ken81] P. R. Kennicott. IGES and the translation of CAD/ CAM data bases. IEEE Computer Graphics and Applications, 1(1):??, January/ February 1981. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Kenwright:2020:TMS

[Ken20] B. Kenwright. There’s more to sound than meets the ear: Sound in interactive environments. IEEE Computer Graphics and Applications, 40(4):62–70, July/ August 2020. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Kergosien:1991:GSS

[Ker91] Yannick L. Kergosien. Generic sign systems in medical imaging. IEEE Computer Graphics and Applications, 11(5):46–65, September 1991. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Kerren:2013:IVC

[Ker13] Andreas Kerren. Information visualization courses for students with a computer science background. IEEE Computer Graphics and Applications, 33(2): 12–15, March/April 2013. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Krueger:1994:VBR

[KF94] W. Krueger and B. Froehlich. Visualization blackboard: The Responsive Workbench (virtual work environment). IEEE Computer Graphics and Applications, 14(3):12– 15, May/June 1994. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Kurtenbach:2005:GEI

[KF05] Gordon Kurtenbach and George Fitzmaurice. Guest Editors’ introduction: Applications of large displays. IEEE Computer Graphics and Applications, 25(4):22– 23, July/August 2005. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

                                        csdl.computer.org/comp/
                                        mags/cg/2005/04/g4022.
                                        pdf.

Kinkeldey:2022:BEV

[KFBI22] Christoph Kinkeldey, Jean-Daniel Fekete, Tanja Blascheck, and Petra Isenberg. Bit-Conduite: Exploratory visual analysis of entity activity on the Bitcoin network. IEEE Computer Graphics and Applications, 42(1):84– 94, January/February 2022. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Karer:2018:DEV

[KFH^{+}18] Benjamin Karer, Alina Freund, Michael Horst, Inga Scheler, Thomas Kossurok, and Franz-Josef Brandt. Designing effective visual interactive systems despite sparse availability of domain information. IEEE Computer Graphics and Applications, 38(5):54–69, September/October 2018. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2018/05/mcg2018050054-
abs.html.

Krejov:2014:MIE

[KGB14] Philip Krejov, Andrew Gilbert, and Richard Bowden. A multitouchless interface: Expanding user interaction. IEEE Computer Graphics and Applications, 34(3):40–48, May/ June 2014. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Kergosien:1994:BCV

[KGK94] Yannick L. Kergosien, Hironobu Gotoda, and Tosiyasu L. Kunii. Bending and creasing virtual paper. IEEE Computer Graphics and Applications, 14(1):40–48, January/ February 1994. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Kriz:1995:VBV

[KGM95] R. D. Kriz, E. H. Glaessgen, and J. D. MacRae. Visualization blackboard: Visualizing gradients in composite design and fabrication. IEEE Computer Graphics and Applications, 15(6): 10–13, November/December 1995. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Krone:2018:YMP

[KGR^{+}18] Michael Krone, Sebastian Grottel, Guido Reina, Christoph Muller, and Thomas Ertl. 10 years of MegaMol: The pain and gain of creating your own visualization framework. IEEE Computer Graphics and Applications, 38(1):109–114, January/ February 2018. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        cg/2018/01/mcg2018010109-
                                        abs.html.

Kenwright:1998:AVC

[KH98] David N. Kenwright and Robert Haimes. Automatic: Vortex core detection. IEEE Computer Graphics and Applications, 18(4):70–74, July/August 1998. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

computer.org/cga/cg1998/
g4070abs.htm.

Kuhlen:2014:QVC

[KH14] Torsten Wolfgang Kuhlen and Bernd Hentschel. Quo vadis CAVE: Does immersive visualization still matter? IEEE Computer Graphics and Applications, 34(5):14–21, September/ October 2014. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://csdl.
computer.org/csdl/mags/
cg/2014/05/mcg2014050014-
abs.html.

Hsieh:2013:VCA

[kHHBEH13] Chun ko Hsieh, Yi-Ping Hung, Moshe Ben-Ezra, and Hsin-Fang Hsieh. Viewing Chinese art on an interactive tabletop. IEEE Computer Graphics and Applications, 33(3):16–21, May/ June 2013. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Kuchar:2006:IIA

[KHHP06] Olga A. Kuchar, Thomas J. Hoeft, Susan Havre, and Kenneth A. Perrine. Isn’t it about time? IEEE Computer Graphics and Applications, 26(3):80–83, May/ June 2006. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Kosara:2003:USW

[KHI^{+}03] Robert Kosara, Christopher G. Healey, Victoria Interrante, David H. Laidlaw, and Colin Ware. User studies: Why, how, and when? IEEE Computer Graphics and Applications, 23(4):20– 25, July/August 2003. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http:

//csdl.computer.org/dl/
mags/cg/2003/04/g4020.
htm;   http://csdl.computer.
org/dl/mags/cg/2003/04/
g4020.pdf.

Kaufman:1994:RIV

[KHK^{+}94] A. Kaufman, K. H. Hohne, W. Kruger, L. Rosenblum, and P. Schroder. Research issues in volume visualization. IEEE Computer Graphics and Applications, 14(2):63–67, March/ April 1994. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Kendall:2011:TGL

[KHP^{+}11] Wesley Kendall, Jian Huang, Tom Peterka, Robert Latham, and Robert Ross. Toward a general I/O layer for parallel-visualization applications. IEEE Computer Graphics and Applications, 31(6):6–10, November/December 2011. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Kendall:2012:GQF

[KHP12] Wesley Kendall, Jian Huang, and Tom Peterka. Geometric quantification of features in large flow fields. IEEE Computer Graphics and Applications, 32(4): 46–54, July/August 2012. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Kuschfeldt:1998:EVC

[KHSE98] Sven Kuschfeldt, Michael Holzner, Ove Sommer, and Thomas Ertl. Efficient visualization of crash-worthiness simulations. IEEE Computer Graphics and Applications, 18(4):60–65, July/ August 1998. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

computer.org/cga/cg1998/
g4060abs.htm.

Kimura:1984:GDS

[Kim84] Fumihiko Kimura. GEOMAP-III: Designing solids with free-form surfaces. IEEE Computer Graphics and Applications, 4(6):58–72, June 1984. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Kakehi:2005:LTI

[KIN^{+}05] Yasuaki Kakehi, Makoto Iida, Takeshi Naemura, Yoshinari Shirai, Mitsunori Matsushita, and Takeshi Ohguro. Lumisight table: An interactive viewdependent tabletop display. IEEE Computer Graphics and Applications, 25(1):48– 53, January/February 2005. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Katzourin:2006:SIG

[KIQ^{+}06] Michael Katzourin, Daniel Ignatoff, Lincoln Quirk, Joseph J. LaViola, Jr., and Odest Chadwicke Jenkins. Swordplay: Innovating game development through VR. IEEE Computer Graphics and Applications, 26(6): 15–19, November/December 2006. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Kirk:2021:RCP

[Kir21] A. Kirk. A recipe of capabilities for pursuing expertise in data visualization:

A practitioner’s perspective. IEEE Computer Graphics and Applications, 41(1):58– 62, January/February 2021. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Keefe:2018:WRS

[KJA^{+}18] Daniel F. Keefe, Seth Johnson, Ross Altheimer, Deuk-Geun Hong, Robert Hunter, Andrea J. Johnson, Maura Rockcastle, Mark Swackhamer, and Aaron Wittkamper. Weather report: A site-specific artwork interweaving human experiences and scientific data physicalization. IEEE Computer Graphics and Applications, 38(4):10–16, July/ August 2018. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2018/04/mcg2018040010.
html.

Kauer:2018:MVD

[KJR^{+}18] Tobias Kauer, Sagar Joglekar, Miriam Redi, Luca Maria Aiello, and Daniele Quercia. Mapping and visualizing deep-learning urban beautification. IEEE Computer Graphics and Applications, 38(5):70–83, September/October 2018. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/

                                        cg/2018/05/mcg2018050070-
                                        abs.html.

Keim:1994:VDE

[KK94] Daniel A. Keim and Hans-Peter Kriegel. VisDB: Database exploration using multidimensional visualization. IEEE Computer Graphics and Applications, 14(5):40–49, September 1994. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Kanev:2002:IDF

[KK02] Kamen Kanev and Shigeo Kimura. Integrating dynamic full-body motion devices in interactive 3D entertainment. IEEE Computer Graphics and Applications, 22(4):76–86, July/ August 2002. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

                                        computer.org/cga/cg2002/
                                        g4076abs.htm;   http:/
                                        /dlib.computer.org/cg/
                                        books/cg2002/pdf/g4076.
                                        pdf.

Kitagawa:1984:FFB

[KKAM84] Hiroyuki Kitagawa, Tosiyasu L. Kunii, Motoei Azuma, and Sohei Misaki. Formgraphics: a form-based graphics architecture providing a database workbench. IEEE Computer Graphics and Applications, 4(6):38–54, 56, June 1984.

CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Kraus:2021:VIV

[KKF^{+}21] Matthias Kraus, Karsten Klein, Johannes Fuchs, Daniel A. Keim, Falk Schreiber, and Michael Sedlmair. The value of immersive visualization. IEEE Computer Graphics and Applications, 41(4):125–132, July/August 2021. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Kostakos:2013:MPD

[KKG^{+}13] Vassilis Kostakos, Hannu Kukka, Jorge Goncalves, Nikolaos Tselios, and Timo Ojala. Multipurpose public displays: How shortcut menus affect usage. IEEE Computer Graphics and Applications, 33(2):56–63, March/April 2013. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Kramer:1990:GAO

[KKGH90] David M. Kramer, Leon Kaufman, Ricardo J. Guzman, and Christine Hawryszko. A general algorithm for oblique image reconstruction. IEEE Computer Graphics and Applications, 10(2):62–65, March/April 1990. CODEN ICGADZ.

ISSN 0272-1716 (print), 1558-1756 (electronic).

Kumar:1996:IDL

[KKM96] Subodh Kumar, Shankar Krishnan, and Dinesh Manocha. Projects in VR: Interactive display of large solid models for walkthroughs. IEEE Computer Graphics and Applications, 16(2): 9–11, March/April 1996. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Kajimoto:2004:SES

[KKTI04] Hiroyuki Kajimoto, Naoki Kawakami, Susumu Tachi, and Masahiko Inami. Smart-Touch: Electric skin to touch the untouchable. IEEE Computer Graphics and Applications, 24(1):36– 43, January/February 2004. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL

http://csdl.computer.org/
comp/mags/cg/2004/01/g1036abs.
htm;   http://csdl.computer.
org/dl/mags/cg/2004/01/
g1036.pdf.

Klosowski:2002:PVD

[KKV^{+}02] James T. Klosowski, Peter D. Kirchner, Julia Valuyeva, Greg Abram, Christopher J. Morris, Robert H. Wolfe, and Thomas Jackman. Projects in VR: Deep view: High-resolution reality. IEEE Computer Graphics and Applications, 22(3): 12–15, May/June 2002. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

computer.org/cga/cg2002/
g3012abs.htm;   http:/
/dlib.computer.org/cg/
books/cg2002/pdf/g3012.
pdf.

Keefe:2005:ACD

[KKVL05] Daniel F. Keefe, David B.

Karelitz, Eileen L. Vote, and David H. Laidlaw. Artistic collaboration in designing VR visualizations. IEEE Computer Graphics and Applications, 25(2):18–23, March/April 2005. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Kyung:2009:UPH

[KL09] Ki-Uk Kyung and Jun-Young Lee. Ubi-Pen: a haptic interface with texture and vibrotactile display. IEEE Computer Graphics and Applications, 29(1):56– 64, January/February 2009. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Kouril:2018:SUE

[KL18] Pavel Kouril and Fotis Liarokapis. Simulation of underwater excavation using dredging procedures. IEEE Computer Graphics and Applications, 38(2):103–111, March/April 2018. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        cg/2018/02/mcg2018020103-
                                        abs.html.

Klasky:1989:CAV

[Kla89] Ronald S. Klasky. Computer animation for visualizing terrain data. IEEE Computer Graphics and Applications, 9(3):12–13, May/ June 1989. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Klassen:1993:IAA

[Kla93] R. Victor Klassen. Increasing the apparent addressability of supersampling grids. IEEE Computer Graphics and Applications, 13(5):74–77, September 1993. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Komura:2015:VRS

[KLL^{+}15] Taku Komura, Rynson W. H. Lau, Ming C. Lin, Aditi Majumder, Dinesh Manocha, and Wei Wei Xu. Virtual reality software and technology. IEEE Computer Graphics and Applications, 35(5):20–21, September/October 2015. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
cg/2015/05/mcg2015050020.
html.

Kang:2012:SLD

[KLR12] Nahyup Kang, Min Woo Lee, and Taehyun Rhee. Simulating liver deformation during respiration using sparse local features. IEEE Computer Graphics and Applications, 32(5): 29–38, September/October 2012. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Koivunen:1988:HIA

[KM88] Marja-Riitta Koivunen and Martti Mantyla. HutWindows: an improved architecture for a user interface management system. IEEE Computer Graphics and Applications, 8(1):43– 52, January/February 1988. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Kao:2000:GEI

[KM00] David L. Kao and Kwan-Liu Ma. Guest Editors’ introduction: The life cycle of a visualization case study. IEEE Computer Graphics and Applications, 20(5):29–31, September/ October 2000. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://dlib.

                                        computer.org/cg/books/
                                        cg2000/pdf/g5029.pdf.

Kirby:2013:VCW

[KM13] Robert M. Kirby and Miriah Meyer. Visualization collaborations: What works and why. IEEE Computer Graphics and Applications, 33(6):82–88, November/December 2013. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Kasap:2009:MTR

[KMCMT09] Zerrin Kasap, Maher Ben Moussa, Parag Chaudhuri, and Nadia Magnenat-Thalmann. Making them remember—emotional virtual characters with memory. IEEE Computer Graphics and Applications, 29(2): 20–29, March/April 2009. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Kosara:2002:FTL

[KMH02] Robert Kosara, Silvia Miksch, and Helwig Hauser. Focus + context taken literally. IEEE Computer Graphics and Applications, 22(1):22–29, January/February 2002. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http:

                                        //dlib.computer.org/cg/
                                        books/cg2002/g1022abs.
                                        htm;   http://dlib.computer.

org/cg/books/cg2002/pdf/
g1022.pdf.

Kim:2018:TBH

[KmKK^{+}18] Daehwan Kim, Hye mi Kim, Hang-Kee Kim, Su-Ran Park, Ki-Suk Lee, and Ki-Hong Kim. ThunderPunch: A bare-hand, gesture-based, large interactive display interface with upper-body-part detection in a top view. IEEE Computer Graphics and Applications, 38(5):100–111, September/ October 2018. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2018/05/mcg2018050100-
abs.html.

Kniss:2001:ITB

[KMM^{+}01] Joe Kniss, Patrick Mc-Cormick, Allen McPherson, James Ahrens, Jamie Painter, Alan Keahey, and Charles Hansen. Interactive texture-based volume rendering for large data sets. IEEE Computer Graphics and Applications, 21(4):52– 61, July/August 2001. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

computer.org/cga/cg2001/
g4052abs.htm;   http:/
/dlib.computer.org/cg/
books/cg2001/pdf/g4052.
pdf.

Kasik:2007:GEI

[KMS07] David J. Kasik, Dinesh Manocha, and Philipp Slusallek. Guest Editors’ introduction: Real-time interaction with complex models. IEEE Computer Graphics and Applications, 27(6):17–19, November/ December 2007. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL http://csdl.
                                        computer.org/comp/mags/
                                        cg/2007/06/mcg2007060017.
                                        pdf.

Kalra:1998:RTA

[KMTM^{+}98] Prem Kalra, Nadia Magnenat-Thalmann, Laurent Moccozet, Gael Sannier, Amaury Aubel, and Daniel Thalmann. Real-time animation of realistic virtual humans. IEEE Computer Graphics and Applications, 18(5):42–57, September/ October 1998. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

                                        computer.org/cga/cg1998/
                                        g5042abs.htm;   http:/
                                        /dlib.computer.org/cg/
                                        books/cg1998/pdf/g5042.
                                        pdf.

Kanade:2007:VRP

[KN07] Takeo Kanade and P. J. Narayanan. Virtualized reality: Perspectives on 4D digitization of dynamic events. IEEE Computer Graphics and Applications, 27(3):32– 40, May/June 2007. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Kishore:2018:BNS

[KND^{+}18] Sameer Kishore, Xavi Navarro, Eva Dominguez, Nonny De La Pena, and Mel Slater. Beaming into the news: A system for and case study of tele-immersive journalism. IEEE Computer Graphics and Applications, 38(2):89–101, March/ April 2018. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2018/02/mcg2018020089-
abs.html.

Koutsofios:1999:VBV

[KNK99] Eleftherios E. Koutsofios, Stephen C. North, and Daniel A. Keim. Visualization blackboard: Visualizing large telecommunication data sets. IEEE Computer Graphics and Applications, 19(3):16–19, May/ June 1999. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://dlib.
computer.org/cg/books/
cg1999/pdf/g3016.pdf.

Kaufman:1993:GEI

[KNR93] Arie E. Kaufman, Gregory M. Nielson, and Lawrence J. Rosenblum. Guest Editor’s introduction: The visualization revolution. IEEE Computer Graphics and Applications, 13(4):16–??, July/August 1993. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Kohlhammer:2012:TVP

[KNRB12] Jorn Kohlhammer, Kawa Nazemi, Tobias Ruppert, and Dirk Burkhardt. Toward visualization in policy modeling. IEEE Computer Graphics and Applications, 32(5):84–89, September/October 2012. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Kochhar:1994:CPH

[Koc94] Sandeep Kochhar. CCAD: a paradigm for humancomputer cooperation in design. IEEE Computer Graphics and Applications, 14(3):54–65, May/ June 1994. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Koh:1993:VSM

[Koh93] Eng-Kiat Koh. Visualization blackboard: Visualizing the stock market crash. IEEE Computer Graphics and Applications, 13(2): 14–16, March/April 1993. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Koning:2012:TCA

[Kon12] Wobbe F. Koning. Teaching 3D computer animation to illustrators: The instructor as translator and technical director. IEEE Computer Graphics and Applications, 32(5):81–83, September/October 2012. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Koprowski:2003:ADR

[Kop03] Gene J. Koprowski. An animated day at the races. IEEE Computer Graphics and Applications, 23(5): 19–23, September/October 2003. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL

http://csdl.computer.org/
dl/mags/cg/2003/05/g5019.
pdf.

Kornfeld:1987:IPD

[Kor87] Cary D. Kornfeld. The image prism: a device for rotating and mirroring bitmap images. IEEE Computer Graphics and Applications, 7(5):21–30, May/June 1987. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Kosara:2015:TDK

[Kos15] Robert Kosara. Tapestry: A different kind of conference on storytelling with data. IEEE Computer Graphics and Applications, 35(2):12– 14, March/April 2015. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http:/

/www.computer.org/csdl/
mags/cg/2015/02/mcg2015020012-
abs.html.

Kosara:2016:POV

[Kos16] Robert Kosara. Presentationoriented visualization techniques. IEEE Computer Graphics and Applications, 36(1):80–85, January/ February 2016. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
cg/2016/01/mcg2016010080-
abs.html.

Kosara:2022:MTM

[Kos22] Robert Kosara. More than meets the eye: a closer look at encodings in visualization. IEEE Computer Graphics and Applications, 42(2):110–114, March/April 2022. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Kosara:2023:NDA

[Kos23] Robert Kosara. Notebooks for data analysis and visualization: Moving beyond the data. IEEE Computer Graphics and Applications, 43(1):91–96, January/ February 2023. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Kozak:2010:ISD

[Koz10] Marcin Kozak. Improved scatterplot design. IEEE Computer Graphics and Applications, 30(6):3–7, November/December 2010. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Kozak:2012:WSF

[Koz12] Marcin Kozak. Watch out for superman: First visualize, then analyze. IEEE Computer Graphics and Applications, 32(3):6–9, May/ June 2012. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Kim:2011:DCS

[KP11] Junggon Kim and Nancy S.

Pollard. Direct control of simulated nonhuman characters. IEEE Computer Graphics and Applications, 31(4):56–65, July/August 2011. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Kim:2012:VME

[KP12] Jinah Kim and Jinah Park.

Visualizing marine environmental changes to the Saemangeum coast. IEEE Computer Graphics and Applications, 32(6):82–87, November/December 2012. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Kim:2014:OMU

[KP14] Taeho Kim and Jinah Park. 3D object manipulation using virtual handles with a grabbing metaphor. IEEE Computer Graphics and Applications, 34(3):30–38, May/June 2014. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Kim:2018:CHE

[KP18] Seokyeol Kim and Jinah Park. Collaborative haptic exploration of dynamic remote environments. IEEE Computer Graphics and Applications, 38(5):84–99, September/ October 2018. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2018/05/mcg2018050084-
abs.html.

Kulshreshth:2017:EGE

[KPL17] Arun Kulshreshth, Kevin Pfeil, and Joseph J. LaViola. Enhancing the gaming experience using 3D spatial user interface technologies. IEEE Computer Graphics and Applications, 37(3):16–23, May/June 2017. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2017/03/mcg2017030016-
abs.html.

Keim:2005:MAB

[KPN05] Daniel A. Keim, Christian Panse, and Stephen C. North. Medial-axis-based cartograms. IEEE Computer Graphics and Applications, 25(3):60–68, May/ June 2005. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Keim:2004:VDM

[KPSN04] Daniel A. Keim, Christian Panse, Mike Sips, and Stephen C. North. Visual data mining in large geospatial point sets. IEEE Computer Graphics and Applications, 24(5):36–44, September/October 2004. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http:

//csdl.computer.org/dl/
mags/cg/2004/05/g5036.
htm;   http://csdl.computer.
org/dl/mags/cg/2004/05/
g5036.pdf.

Kroszynski:1989:GDT

[KPTS89] Uri I. Kroszynski, Bjarne Palstroem, Erik Trostmann, and Ernst G. Schlechtendahl. Geometric data transfer between CAD systems: Solid models. IEEE Computer Graphics and Applications, 9(5):56–71 (or 57–71??), September 1989. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Keim:2013:BDV

[KQM13] Daniel Keim, Huamin Qu, and Kwan-Liu Ma. Big-data visualization. IEEE Computer Graphics and Applications, 33(4):20–21, July/ August 2013. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Kunii:1986:CGT

[KR86] Tosiyasu L. Kunii and David F. Rogers. Computer Graphics Tokyo 85. IEEE Computer Graphics and Applications, 6(4):14– 58, April 1986. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Kraak:2006:BG

[Kra06] Menno Jan Kraak. Beyond geovisualization. IEEE Computer Graphics and Applications, 26(4):6–9, July/ August 2006. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Kuhnert:2011:VPS

[KRB11] Tom Kuhnert, Stephan Rusdorf, and Guido Brunnett. Virtual prototyping of shoes. IEEE Computer Graphics and Applications, 31(5):30–42, September/October 2011. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Kretzschmar:2022:SVI

[KRG^{+}22] Vanessa Kretzschmar, Allan Rocha, Fabian Günther, Markus Stommel, and Gerik Scheuermann. Stress visualization for interface optimization of a hybrid component using surface tensor spines. IEEE Computer Graphics and Applications, 42(2):45–55, March/ April 2022. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Krikke:2000:AAM

[Kri00] Jan Krikke. Applications:

Axonometry: a matter of perspective. IEEE Computer Graphics and Applications, 20(4):7–11, July/ August 2000. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://dlib.
computer.org/cg/books/
cg2000/pdf/g4007.pdf.

Krikke:2001:AGA

[Kri01] Jan Krikke. Applications:

Graphics applications over the Wireless Web: Japan sets the pace. IEEE Computer Graphics and Applications, 21(3):9–15, May/ June 2001. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL http://dlib.
                                        computer.org/cg/books/
                                        cg2001/pdf/g3009.pdf.

Krikke:2003:SRJ

[Kri03]  Jan Krikke.  Samurai

Romanesque, J2ME, and the Battle for Mobile Cyberspace. IEEE Computer Graphics and Applications, 23(1):16–23, January/ February 2003. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL http://dlib.
                                        computer.org/cg/books/
                                        cg2003/pdf/g1016.pdf.

Krikke:2004:SVT

[Kri04]  Jan Krikke.  Streaming

video transforms the media industry. IEEE Computer Graphics and Applications, 24(4):6–12, July/ August 2004. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL http://csdl.
                                        computer.org/comp/mags/
                                        cg/2004/04/g4006.pdf;
                                        http://csdl.computer.org/
                                        dl/mags/cg/2004/04/g4006.
                                        htm.

Krikke:2005:NRT

[Kri05] Jan Krikke. Near real-time tsunami computer simulations within reach. IEEE Computer Graphics and Applications, 25(5):16–21, September/October 2005. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Krikke:2006:CGA

[Kri06]  Jan Krikke.  Computer

graphics advances the art of anime. IEEE Computer Graphics and Applications, 26(3):14–19, May/ June 2006. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Kolojejchick:1997:IAT

[KRL97] John Kolojejchick, Steven F.

Roth, and Peter Lucas. Information appliances and tools in Visage. IEEE Computer Graphics and Applications, 17(4):32–41, July/ August 1997. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Kroeker:2002:AGS

[Kro02] Kirk L. Kroeker. Applications: Graphics and security: Exploring visual biometrics. IEEE Computer Graphics and Applications, 22(4):16–21, July/ August 2002. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

computer.org/cga/cg2002/
g4016abs.htm;   http:/
/dlib.computer.org/cg/

                                        books/cg2002/pdf/g4016.
                                        pdf.

Kroeker:2004:SDN

[Kro04] Kirk L. Kroeker. Seeing data: New methods for understanding information. IEEE Computer Graphics and Applications, 24(3):6– 12, May/June 2004. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http:

                                        //csdl.computer.org/dl/
                                        mags/cg/2004/03/g3006.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/cg/2004/03/
                                        g3006.pdf.

Krueger:1995:AVR

[Kru95] M. W. Krueger. Automating virtual reality. IEEE Computer Graphics and Applications, 15(1):9–11, January/ February 1995. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Kunii:1994:RIM

[KS94] T. L. Kunii and Y. Shinagawa. Research issues in modeling complex object shapes. IEEE Computer Graphics and Applications, 14(2):80–83, March/ April 1994. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Klimansezewski:1997:FRT

[KS97] Kryzsztof S. Klimansezewski and Thomas W. Sederberg. Faster ray tracing using adaptive grids. IEEE Computer Graphics and Applications, 17(1):42–51, January/ February 1997. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Kamen:1998:TRU

[KS98] Yakov Kamen and Leon Shirman. Triangle rendering using adaptive subdivision. IEEE Computer Graphics and Applications, 18(2):95– 103, March/April 1998. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

computer.org/cga/cg1998/
g2095abs.htm;   http:/
/dlib.computer.org/cg/
books/cg1998/pdf/g2095.
pdf.

Karkanis:2001:CDT

[KS01] Tasso Karkanis and A. James Stewart. Curvature-dependent triangulation of implicit surfaces. IEEE Computer Graphics and Applications, 21(2):60–69, March/ April 2001. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

computer.org/cga/cg2001/
g2060abs.htm;   http:/
/dlib.computer.org/cg/

                                        books/cg2001/pdf/g2060.
                                        pdf.

Kara:2007:SBS

[KS07] Levent Burak Kara and Kenji Shimada. Sketchbased 3D-shape creation for industrial styling design. IEEE Computer Graphics and Applications, 27(1):60– 71, January/February 2007. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Kirby:2008:NVV

[KS08] Robert M. Kirby and Cláudio T. Silva. The need for verifiable visualization. IEEE Computer Graphics and Applications, 28(5): 78–83, September/October 2008. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Kerren:2014:WII

[KS14] Andreas Kerren and Falk Schreiber. Why integrate InfoVis and SciVis?: An example from systems biology. IEEE Computer Graphics and Applications, 34(6):69–73, November/ December 2014. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL http://csdl.
                                        computer.org/csdl/mags/
                                        cg/2014/06/mcg2014060069-
                                        abs.html.

Kaiyrbekov:2020:DSB

[KS20] K. Kaiyrbekov and M. Sezgin. Deep stroke-based sketched symbol reconstruction and segmentation. IEEE Computer Graphics and Applications, 40(1): 112–126, January/February 2020. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Kurzyna:2022:BPM

[Ks22] Marcin Kurzyna and Tomasz Kwapi ski. Braid Plot: a mixed palette plotting method as an extension of contour plot. IEEE Computer Graphics and Applications, 42(1): 95–104, January/February 2022. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Kim:2004:HRT

[KSD04] Laehyun Kim, Gaurav S.

Sukhatme, and Mathieu Desbrun. A haptic-rendering technique based on hybrid surface representation. IEEE Computer Graphics and Applications, 24(2):66– 75, March/April 2004. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

csdl.computer.org/comp/
mags/cg/2004/02/g2066abs.
htm;   http://csdl.computer.
org/dl/mags/cg/2004/02/
g2066.htm;   http://csdl.
computer.org/dl/mags/cg/
2004/02/g2066.pdf.

Kanai:2000:MAT

[KSK00] Takashi Kanai, Hiromasa Suzuki, and Fumihiko Kimura. Metamorphosis of arbitrary triangular meshes. IEEE Computer Graphics and Applications, 20(2):62–75, March/April 2000. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

                                        computer.org/cga/cg2000/
                                        g2062abs.htm;   http:/
                                        /dlib.computer.org/cg/
                                        books/cg2000/pdf/g2062.
                                        pdf.

Keck:2021:DFA

[KSK21] Mandy Keck, Elena Stoll, and Dietrich Kammer. A didactic framework for analyzing learning activities to design InfoVis courses. IEEE Computer Graphics and Applications, 41(6):80–90, November/December 2021. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Karima:1985:PDC

[KSM85] Medhat Karima, Kuldip S. Sadhal, and Tim O. Mc-Neil. From paper drawings to computer-aided design. IEEE Computer Graphics and Applications, 5(2):27– 39, February 1985. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Kapadia:2011:BAF

[KSRF11] Mubbasir Kapadia, Shawn Singh, Glenn Reinman, and Petros Faloutsos. A behavior-authoring framework for multiactor simulations. IEEE Computer Graphics and Applications, 31(6):45–55, November/December 2011. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Klein:2013:XXC

[KSRS13] Felix Klein, Kristian Sons, Dmitri Rubinstein, and Philipp Slusallek. XML3D and Xflow: Combining declarative 3D for the Web with generic data flows. IEEE Computer Graphics and Applications, 33(5): 38–47, September/October 2013. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Kim:2007:FSL

[KSSL07] Theodore Kim, Jason Sewall, Avneesh Sud, and Ming C. Lin. Fast simulation of Laplacian growth. IEEE Computer Graphics and Applications, 27(2): 68–76, March/April 2007. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Kunii:1985:GTB

[KSY85] Tosiyasu L. Kunii, Toshiaki Satoh, and Kazunori Yamaguchi. Generation of topological boundary representations from octree encoding. IEEE Computer Graphics and Applications, 5(3): 29–38, March/April 1985. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Kalvin:1996:SPM

[KT96] Alan D. Kalvin and Russell H. Taylor. Superfaces: Polygonal mesh simplification with bounded error. IEEE Computer Graphics and Applications, 16 (3):64–77, May/June 1996. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Kasik:2002:EGD

[KTA^{+}02] David J. Kasik, James J. Troy, Stephen R. Amorosi, Marie O. Murray, and Shankar N. Swamy. Evaluating graphics displays for complex 3D models. IEEE Computer Graphics and Applications, 22(3):56–64, May/June 2002. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

computer.org/cga/cg2002/
g3056abs.htm;   http:/
/dlib.computer.org/cg/
books/cg2002/pdf/g3056.
pdf.

Kadlec:2010:KAV

[KTD10] Benjamin J. Kadlec, Henry M.

Tufo, and Geoffrey A. Dorn. Knowledge-assisted visualization and segmentation of geologic features. IEEE Computer Graphics and Applications, 30(1):30–39, January/February 2010. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Kubo:1984:AGI

[Kub84] Sachio Kubo. Alis: a geographical information system for urban research. IEEE Computer Graphics and Applications, 4(5):68– 76, May/June 1984. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Kulik:2009:BRM

[Kul09] Alexander Kulik. Building on realism and magic for designing 3D interaction techniques. IEEE Computer Graphics and Applications, 29(6):22–33, November/December 2009. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Kunii:2000:DC

[Kun00] Tosiyasu L. Kunii. Discovering cyberworlds. IEEE Computer Graphics and Applications, 20(1):64–65, January/February 2000. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http:

//dlib.computer.org/cg/
books/cg2000/pdf/g1064.
pdf.

Kwon:2017:SSV

[KVHD17] Bum Chul Kwon, Janu Verma, Peter J. Haas, and Cagatay Demiralp. Sampling for scalable visual analytics. IEEE Computer Graphics and Applications, 37(1):100–108, January/ February 2017. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2017/01/mcg2017010100-
abs.html.

Kamiyama:2005:VBS

[KVM^{+}05] Kazuto Kamiyama, Kevin Vlack, Terukazu Mizota, Hiroyuki Kajimoto, Naoki Kawakami, and Susumu Tachi. Vision-based sensor for real-time measuring of surface traction fields. IEEE Computer Graphics and Applications, 25(1):68– 75, January/February 2005. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Komerska:2004:HSS

[KW04] Rick Komerska and Colin Ware. Haptic state-surface interactions. IEEE Computer Graphics and Applications, 24(6):52–59, November/December 2004. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http:

//csdl.computer.org/dl/
mags/cg/2004/06/g6052.
htm;   http://csdl.computer.
org/dl/mags/cg/2004/06/
g6052.pdf.

Koehler:2010:KAR

[KW10] Christopher Koehler and Thomas Wischgoll. Knowledgeassisted reconstruction of the human rib cage and lungs. IEEE Computer Graphics and Applications, 30(1):17–29, January/February 2010. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Kurzhals:2015:ETP

[KW15] Kuno Kurzhals and Daniel Weiskopf. Eye tracking for personal visual analytics. IEEE Computer Graphics and Applications, 35(4):64–72, July/ August 2015. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
cg/2015/04/mcg2015040064-
abs.html.

Kasik:2023:BCA

[KWJ23] David J. Kasik, Mary C.

Whitton, and Chris R. Johnson. The big 50: Celebrating 50 ACM SIG-GRAPH conferences. IEEE Computer Graphics and Applications, 43(4):12–80, July/August 2023. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Karyda:2021:NPS

[KWK21] M. Karyda, D. Wilde, and M. G. Kjærsgaard. Narrative physicalization: Supporting interactive engagement with personal data. IEEE Computer Graphics and Applications, 41(1):74– 86, January/February 2021. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Kurzion:1997:ISD

[KY97] Yair Kurzion and Roni Yagel. Interactive space deformation with hardwareassisted rendering. IEEE Computer Graphics and Applications, 17(5):66–77, September/October 1997. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL

                                        http://dlib.computer.org/
                                        cg/books/cg1997/pdf/g5066.
                                        pdf.

Kang:2024:IAV

[KYK^{+}24] Hosan Kang, Jinseong Yang, Beom-Seok Ko, Bo-Seong Kim, Oh-Young Song, and Soo-Mi Choi. Integrated augmented and virtual reality technologies for realistic fire drill training. IEEE Computer Graphics and Applications, 44(2): 89–99, March/April 2024.

CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Koppal:2011:VCE

[KZC^{+}11] Sanjeev J. Koppal, C. Lawrence Zitnick, Michael F. Cohen, Sing Bing Kang, Bryan Ressler, and Alex Colburn. A viewer-centric editor for 3D movies. IEEE Computer Graphics and Applications, 31(1):20–35, January/ February 2011. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Liarokapis:2020:CLS

[LA20a] F. Liarokapis and E. F. Anderson. Collaborating and learning in shared virtual environments. IEEE Computer Graphics and Applications, 40(5):8–9, September/October 2020. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Liarokapis:2020:GVE

[LA20b] F. Liarokapis and E. F. Anderson. Graphics and virtual environments for serious games. IEEE Computer Graphics and Applications, 40(4):8–9, July/ August 2020. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Lira:2014:VAS

[LAC^{+}14] Wallace P. Lira, Ronnie Alves, Jean M. R. Costa, Gustavo Pessin, Lilyan Galvao, Ana C. Cardoso, and Cleidson R. B. de Souza. A visual-analytics system for railway safety management. IEEE Computer Graphics and Applications, 34(5):52–57, September/ October 2014. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://csdl.
computer.org/csdl/mags/
cg/2014/05/mcg2014050052-
abs.html.

Laidlaw:2001:VVL

[Lai01] David H. Laidlaw. Visualization viewpoints: Loose, artistic “Textures” for visualization. IEEE Computer Graphics and Applications, 21(2):6–9, March/ April 2001. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://dlib.
computer.org/cg/books/
cg2001/pdf/g2006.pdf.

Lext:2001:BAR

[LAM01] Jonas Lext, Ulf Assarsson, and Tomas Möller. A benchmark for animated ray tracing. IEEE Computer Graphics and Applications, 21(2):22–30, March/ April 2001. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

computer.org/cga/cg2001/
g2022abs.htm;   http:/
/dlib.computer.org/cg/
books/cg2001/pdf/g2022.
pdf.

Laramee:2010:UVD

[Lar10] Robert Laramee. Using visualization to debug visualization software. IEEE Computer Graphics and Applications, 30(6):67–73, November/December 2010. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Laramee:2011:HRV

[Lar11] Robert S. Laramee. How to read a visualization research paper: Extracting the essentials. IEEE Computer Graphics and Applications, 31(3):78–82, May/ June 2011. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Latta:1998:TOE

[Lat98] John N. Latta. Technologybased out-of-home entertainment. IEEE Computer Graphics and Applications, 18(1):24–28, January/ February 1998. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

computer.org/cga/cg1998/
g1024abs.htm;   http:/
/dlib.computer.org/cg/

                                        books/cg1998/pdf/g1024.
                                        pdf.

Lavey:1983:LPC

[Lav83] Warren G. Lavey. Legal protection of computer graphics software. IEEE Computer Graphics and Applications, 3(1):11–16, 18, 20– 21, January/February 1983. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

LaViola:2007:AMS

[LaV07] Joseph J. LaViola, Jr. Advances in mathematical sketching: Moving toward the paradigm’s full potential. IEEE Computer Graphics and Applications, 27(1): 38–48, January/February 2007. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

LaViola:2008:BVS

[LaV08] Joseph J. LaViola, Jr. Bringing VR and spatial 3D interaction to the masses through video games. IEEE Computer Graphics and Applications, 28(5):10–15, September/October 2008. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Lawton:1999:ACA

[Law99]  Mark Lawton.  About

the cover: Advancing 3D through VRML on the Web. IEEE Computer Graphics and Applications, 19(2):4–5, March/April 1999. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://dlib.
computer.org/cg/books/
cg1999/pdf/g2004.pdf.

Lazear:1987:PHD

[Laz87] Tom Lazear. PC hardware developments and CAD software. IEEE Computer Graphics and Applications, 7(10):28–31, October 1987. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Lutz:2020:ELM

[LB20] E. Lutz and A. Bares.

Eleanor Lutz: Making art from public data. IEEE Computer Graphics and Applications, 40(1):8–13, January/February 2020. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Lavender:1992:VDS

[LBD^{+}92] David Lavender, Adrian Bowyer, James Davenport, Andrew Wallis, and John Woodwark. Voronoi diagrams of set-theoretic solid models. IEEE Computer Graphics and Applications, 12(5):69–??, September 1992. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Lee:2019:IA

[LBDM19] B. Lee, B. Bach, T. Dwyer, and K. Marriott. Immersive analytics. IEEE Computer Graphics and Applications, 39(3):16–18, May/ June 2019. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Lefohn:2003:OAH

[LBS^{+}03] Aaron Lefohn, Brian Budge, Peter Shirley, Richard Caruso, and Erik Reinhard. An ocularist’s approach to human iris synthesis. IEEE Computer Graphics and Applications, 23(6):70–75, November/December 2003. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http:

                                        //csdl.computer.org/dl/
                                        mags/cg/2003/06/g6070.
                                        pdf;   http://csdl.computer.
                                        org/dl/mags/cg/2003/06/
                                        g6070abs.htm.

Lippman:1985:CWP

[LBSS85] Andrew Lippman, Walter Bender, Gitta Solomon, and Mitsuo Saito. Color word processing. IEEE Computer Graphics and Applications, 5(6):41–46, June 1985. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Li:2017:CDA

[LBW^{+}17] Guanbin Li, Sai Bi, Jue Wang, Yingqing Xu, and Yizhou Yu. ColorSketch: A drawing assistant for generating color sketches from photos. IEEE Computer Graphics and Applications, 37(3):70–81, May/ June 2017. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2017/03/mcg2017030070-
abs.html.

Lamb:2016:VRG

[LC16] Peter Lamb and Hayden Croft. Visualizing rugby game styles using selforganizing maps. IEEE Computer Graphics and Applications, 36(6):11–15, November/December 2016. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL

https://www.computer.org/
csdl/mags/cg/2016/06/mcg2016060011-
abs.html.

Lin:2023:BOC

[LCB^{+}23] Tica Lin, Zhutian Chen, Johanna Beyer, Yingcai Wu, Hanspeter Pfister, and Yalong Yang. The ball is in our court: Conducting visualization research with sports experts. IEEE Computer Graphics and Applications, 43(1):84–90, January/ February 2023. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Liu:1996:DCN

[LCC^{+}96] Pei-Wen Liu, Lih-Shyang Chen, Su-Chou Chen, Jong-Ping Chen, Fang-Yi Lin, and Shy-Shang Hwang. Distributed computing: New power for scientific visualization. IEEE Computer Graphics and Applications, 16(3):42–51, May/ June 1996. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Li:2000:BUS

[LCC^{+}00] Kai Li, Han Chen, Yuqun Chen, Douglas W. Clark, Perry Cook, Stefanos Damianakis, Georg Essl, Adam Finkelstein, Thomas Funkhouser, Timothy Housel, Allison Klein, Zhiyan Liu, Emil Praun, Rudrajit Samanta, Ben Shedd, Jaswinder Pal Singh, George Tzanetakis, and Jiannan Zheng. Building and using a scalable display wall system. IEEE Computer Graphics and Applications, 20(4):29–37, July/August 2000. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

                                        computer.org/cga/cg2000/
                                        g4029abs.htm;   http:/
                                        /dlib.computer.org/cg/
                                        books/cg2000/pdf/g4029.
                                        pdf.

Lores:1983:EGS

[LCG83] M. E. Lores, S. H. Chasen, and J. M. Garner. Evaluation of 3-D graphics software: a case study. IEEE Computer Graphics and Applications, 3(8):73–77, November/December 1983. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Lee:2002:NFB

[LCG^{+}02] John Peter Lee, Daniel Carr, Georges Grinstein, John Kinney, and Jeffrey Saffer. The next frontier for bio- and cheminformatics visualization. IEEE Computer Graphics and Applications, 22(5):6–11, September/October 2002. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://dlib.
computer.org/cg/books/
cg2002/pdf/g5006.pdf.

Lee:2020:RBA

[LCI^{+}20] B. Lee, E. K. Choe, P. Isenberg, K. Marriott, and J. Stasko. Reaching broader audiences with data visualization. IEEE Computer Graphics and Applications, 40(2):82–90, March/ April 2020. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Lok:2014:MRH

[LCR^{+}14] Benjamin Lok, Joon Hao Chuah, Andrew Robb, Andrew Cordar, Samsun Lampotang, Adam Wendling, and Casey White. Mixedreality humans for team training. IEEE Computer Graphics and Applications, 34(3):72–75, May/ June 2014. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Laurenzo:2021:TLP

[LCS21] Tomás Laurenzo, Bruce D. Campbell, and Francesca Samsel. Tomás Laurenzo: Pushing back on colonization. IEEE Computer Graphics and Applications, 41(4):118–124, July/August 2021. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Liu:2012:ICU

[LCW12] Shengjun Liu, Kwan-Chung Chan, and Charlie C. L. Wang. Iterative consolidation of unorganized point clouds. IEEE Computer Graphics and Applications, 32(3):70–83, May/ June 2012. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Lehner:1997:PVD

[LD97] Valerie D. Lehner and Thomas A. DeFanti. Projects in VR: Distributed virtual reality: Supporting remote collaboration in vehicle design. IEEE Computer Graphics and Applications, 17(2):13–17, March/ April 1997. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Lintermann:1999:IMP

[LD99] Bernd Lintermann and Oliver Deussen. Interactive modeling of plants. IEEE Computer Graphics and Applications, 19(1):56–65, January/February 1999. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

computer.org/cga/cg1999/
g1056abs.htm;   http:/
/dlib.computer.org/cg/
books/cg1999/pdf/g1056.
pdf.

Li:2008:OBB

[LD08] Qing Li and Zhigang Deng.

Orthogonal-blendshape-based editing system for facial motion capture data. IEEE Computer Graphics and Applications, 28(6):76–82, November/December 2008. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Lear:1996:EAF

[Lea96a] Anne C. Lear. 1995 excellence awards favor 3D products; display technology matches laser prints on screen; on-line cancer research database adapts to Web; bringing the office home from the office; VRML group adopts moving worlds proposal for V.2.0; MPEG-4 goal: Wireless images. IEEE Computer Graphics and Applications, 16(3):89–92, May/ June 1996. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Lear:1996:ACE

[Lea96b] Anne C. Lear. About the cover: Exploring synthetic possibilities of virtual space. IEEE Computer Graphics and Applications, 16(2):4–5, March/April 1996. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Lear:1996:ACF

[Lea96c] Anne C. Lear. About the cover: From fiber to pixels to fiber again. IEEE Computer Graphics and Applications, 16(5):4–5, September 1996. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Lear:1996:NPG

[Lea96d] Anne C. Lear. In the news: Portable graphics are free, again; compression technique benefits medical images; 3D in a cube; TCP incompatibility plagues satellites; flat displays in production. IEEE Computer Graphics and Applications, 16(6):76–78, November/ December 1996. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Lear:1996:TRI

[Lea96e] Anne C. Lear. Telemedicine reception is still fuzzy, despite advances. IEEE Computer Graphics and Applications, 16(2):91, March/ April 1996. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Lear:1997:ADO

[Lea97a] Anne C. Lear. Applications: Digital orthophotography: Mapping with pictures. IEEE Computer Graphics and Applications, 17(5):12–14, September/ October 1997. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://dlib.
computer.org/cg/books/
cg1997/pdf/g5012.pdf.

Lear:1997:AVR

[Lea97b] Anne C. Lear. Applications:

Virtual reality provides real therapy. IEEE Computer Graphics and Applications, 17(4):16–20, July/August 1997. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Lear:1997:NDT

[Lea97c] Anne C. Lear. In the news — digital TV standard set, Visualization 96, InfoVis 96. IEEE Computer Graphics and Applications, 17(1):85– 87, January/February 1997. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Lear:1997:NIE

[Lea97d] Anne C. Lear. In the news: 1996 Industry Excellence awards; VRAIS 97: Bringing more substance to VR. IEEE Computer Graphics and Applications, 17 (3):83–84, May/June 1997. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Lecuyer:2017:PSV

[Lec17] Anatole Lecuyer. Playing with senses in VR: Alternate perceptions combining vision and touch. IEEE Computer Graphics and Applications, 37(1):20–26, January/ February 2017. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        cg/2017/01/mcg2017010020-
                                        abs.html.

Lee:1999:SOB

[Lee99]  Jintae Lee.  Simulating

oriental black-ink painting. IEEE Computer Graphics and Applications, 19(3):74– 81, May/June 1999. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

computer.org/cga/cg1999/
g3074abs.htm;   http:/
/dlib.computer.org/cg/
books/cg1999/pdf/g3074.
pdf.

Lee:2008:RRO

[Lee08]  Jehee Lee.  Represent-

ing rotations and orientations in geometric computing. IEEE Computer Graphics and Applications, 28(2): 75–83, March/April 2008. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Leray:1982:CSS

[Ler82] P. Leray. The CIG system of synthetic image generation. IEEE Computer Graphics and Applications, 2(7):89– 92, July/August 1982. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Levoy:1988:DSV

[Lev88] Marc Levoy. Display of surfaces from volume data. IEEE Computer Graphics and Applications, 8(3):29– 37, May/June 1988. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). See corrigendum [Lev00, WMG00].

Levoy:1990:HRT

[Lev90] Marc Levoy. A hybrid ray tracer for rendering polygon and volume data. IEEE Computer Graphics and Applications, 10(2):33–40, March/April 1990. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Levoy:2000:LEE

[Lev00]  Marc Levoy.  Letter to

the editor: Error in volume rendering paper was in exposition only. IEEE Computer Graphics and Applications, 20(4):6, July/ August 2000. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http:/

                                        /dlib.computer.org/cg/
                                        books/cg2000/pdf/g4006.

pdf. See [Lev88, WMG00].

Levoy:2010:EPC

[Lev10] Marc Levoy. Experimental platforms for computational photography. IEEE Computer Graphics and Applications, 30(5):81–87, September/October 2010. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Lee:1984:SFI

[LF84] Randy B. Lee and David A. Fredricks. Special feature: Intersection of parametric surfaces and a plane. IEEE Computer Graphics and Applications, 4(8):48–51, August 1984. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Lee:1987:MUC

[LF87] Yung-Chia Lee and King-Sun Fu. Machine understanding of CSG: Extraction and unification of manufacturing features. IEEE Computer Graphics and Applications, 7(1):20–32, January/ February 1987. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Levkowitz:2018:RGG

[LFE^{+}18] Haim Levkowitz, John T.

Fallon, Jose L. Encarnacao, Catherine Plaisant, Jean Scholtz, Mark Whiting, Kris Cook, Daniel Keim, and Theresa-Marie Rhyne. Remembering Georges Grinstein. IEEE Computer Graphics and Applications, 38(3):150–156, May/June 2018. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2018/03/mcg2018030150.
html.

Lockton:2020:TTL

[LFFB20] D. Lockton, L. Forlano, J. Fass, and L. Brawley. Thinking with things: Landscapes, connections, and performances as modes of building shared understanding. IEEE Computer Graphics and Applications, 40(6): 38–50, November/December 2020. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Lopez-Fernandez:2023:VRS

[LFMGPG23] Daniel López-Fernández, Jesús Mayor, Marta García-Pérez, and Aldo Gordillo. Are virtual reality serious video games more effective than web video games? IEEE Computer Graphics and Applications, 43(2): 32–42, March/April 2023. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Leithinger:2015:SDS

[LFOI15] Daniel Leithinger, Sean Follmer, Alex Olwal, and Hiroshi Ishii. Shape displays: Spatial interaction with dynamic physical form. IEEE Computer Graphics and Applications, 35(5):5–11, September/October 2015. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL http://www.
                                        computer.org/csdl/mags/
                                        cg/2015/05/mcg2015050005-
                                        abs.html.

Long:2001:ESV

[LFPS01] Fuhui Long, Dagan Feng, Hanchuan Peng, and Wan-Chi Siu. Extracting semantic video objects. IEEE Computer Graphics and Applications, 21(1):48–55, January/February 2001. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

computer.org/cga/cg2001/
g1048abs.htm;   http:/
/dlib.computer.org/cg/
books/cg2001/pdf/g1048.
pdf.

Li:2019:FSS

[LFT19] Lei Li, Hongbo Fu, and Chiew-Lan Tai. Fast sketch segmentation and labeling with deep learning. IEEE Computer Graphics and Applications, 39(2):38–51, March/April 2019. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL https:

//ieeexplore.ieee.org/
document/8565976/.

Lamotte:1997:VCV

[LFV^{+}97] Wim Lamotte, Eddy Flerackers, Frank Van Reeth, Rae Earnshaw, and Joao Mena De Matos. Visenet: Collaborative 3D visualization and VR over ATM networks. IEEE Computer Graphics and Applications, 17(2):66–75, March/ April 1997. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Lamata:2006:CFL

[LGK^{+}06] Pablo Lamata, Enrique J. Gómez, Fernando Bello Roger L. Kneebone, Rajesh Aggarwal, and Félix Lamata. Conceptual framework for laparoscopic VR simulators. IEEE Computer Graphics and Applications, 26(6):69–79, November/December 2006. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Lenz:1986:DDV

[LGLD86] Reiner Lenz, Bjorn Gudmundsson, Bjorn Lindskog, and Per E. Danielsson. Display of density volumes. IEEE Computer Graphics and Applications, 6(7):20– 29, July/August 1986. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Lecuyer:2013:TAV

[LGM13] Anatole Lecuyer, Laurent George, and Maud Marchal. Toward adaptive VR simulators combining visual, haptic, and brain-computer interfaces. IEEE Computer Graphics and Applications, 33(5):18–23, September/October 2013. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Li:2012:CTS

[LGW^{+}12] Shun Li, Jixiang Guo, Qiong Wang, Qiang Meng, Yim-Pan Chui, Jing Qin, and Pheng-Ann Heng. A catheterization-training simulator based on a fast multigrid solver. IEEE Computer Graphics and Applications, 32(6):56–70, November/December 2012. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Lee:2004:BSA

[LGZ04] Alison Lee, Andreas Girgensohn, and Jun Zhang. Browsers to support awareness and social interaction. IEEE Computer Graphics and Applications, 24(5):66–75, September/ October 2004. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http:

//csdl.computer.org/dl/
mags/cg/2004/05/g5066.
htm;   http://csdl.computer.
org/dl/mags/cg/2004/05/
g5066.pdf.

Levkowitz:1992:CSI

[LH92] Haim Levkowitz and Gabor T. Herman. Color scales for image data. IEEE Computer Graphics and Applications, 12(1):72–80, January/ February 1992. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Leistikow:2020:AEV

[LHFL20] S. Leistikow, K. Huesmann, A. Fofonov, and L. Linsen. Aggregated ensemble views for deep water asteroid impact simulations. IEEE Computer Graphics and Applications, 40(1):72– 81, January/February 2020. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Lotteraner:2023:CII

[LHM23] Laura Lotteraner, Thilo Hofmann, and Torsten Möller. The challenge of interdisciplinarity at the intersection of groundwater management and visualization research. IEEE Computer Graphics and Applications, 43(6):50–63, November/December 2023. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Liu:2021:EAS

[LHZ^{+}21] Xiaojun Liu, Changyan He, Hantao Zhao, Jinyuan Jia, and Chang Liu. ExteriorTag: Automatic semantic annotation of BIM building exterior via voxel index analysis. IEEE Computer Graphics and Applications, 41(3):48–58, May/ June 2021. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Li:2022:RCS

[Li22a] Minchen Li. Reliable contact simulation with IPC. IEEE Computer Graphics and Applications, 42(3): 108–114, May/June 2022. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Li:2022:DVC

[Li22b] Tzu-Mao Li. Differentiable visual computing: Challenges and opportunities. IEEE Computer Graphics and Applications, 42(2): 101–109, March/April 2022. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Lewis:2010:DMB

[LiA10] J. P. Lewis and Ken ichi Anjyo. Direct manipulation blendshapes. IEEE Computer Graphics and Applications, 30(4):42–50, July/ August 2010. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Lichten:1984:CDA

[Lic84] Larry Lichten. Computeraided design applications on microcomputers. IEEE Computer Graphics and Applications, 4(10):25–28, October 1984. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Lillehagen:1981:CWS

[Lil81] Frank M. Lillehagen. CAD/ CAM work stations for man-model communication. IEEE Computer Graphics and Applications, 1(3):17– ??, July/August 1981. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Lin:1999:PWM

[Lin99] Y. Arthur Lin. Parametric wafer map visualization. IEEE Computer Graphics and Applications, 19(4):14– 17, July/August 1999. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http:

                                        //dlib.computer.org/cg/
                                        books/cg1999/pdf/g4014.
                                        pdf.

Lee:2019:BID

[LIS^{+}19] B. Lee, K. Isaacs, D. A. Szafir, G. E. Marai, C. Turkay, M. Tory, S. Carpendale, and A. Endert. Broadening intellectual diversity in visualization research papers. IEEE Computer Graphics and Applications, 39(4): 78–85, July/August 2019. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Livingston:2005:EHF

[Liv05] Mark A. Livingston. Evaluating human factors in augmented reality systems.

IEEE Computer Graphics and Applications, 25(6):6–9, November/December 2005. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Leigh:1996:STC

[LJ96] Jason Leigh and Andrew E.

Johnson. Supporting transcontinental collaborative work in persistent virtual environments. IEEE Computer Graphics and Applications, 16(4):47–51, July/ August 1996. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

LaViola:2015:NUI

[LJ15] Joseph J. LaViola and Odest Chadwicke Jenkins. Natural user interfaces for adjustable autonomy in robot control. IEEE Computer Graphics and Applications, 35(3):20–21, May/ June 2015. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
cg/2015/03/mcg2015030020.
html.

Limper:2013:FPL

[LJB^{+}13] Max Limper, Yvonne Jung, Johannes Behr, Timo Sturm, Tobias Franke, Karsten Schwenk, and Arjan Kuijper. Fast, progressive loading of binary-encoded declarative-3D Web content. IEEE Computer Graphics and Applications, 33(5): 26–36, September/October 2013. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Lu:2012:MM

[LJJ^{+}12] Shufang Lu, Aubrey Jaffer, Xiaogang Jin, Hanli Zhao, and Xiaoyang Mao. Mathematical marbling. IEEE Computer Graphics and Applications, 32(6):26–35, November/December 2012. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Lu:2017:SMM

[LJJ^{+}17] Shufang Lu, Xiaogang Jin, Aubrey Jaffer, Fei Gao, and Xiaoyang Mao. Solid mathematical marbling. IEEE Computer Graphics and Applications, 37(2):90–98, March/April 2017. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        cg/2017/02/mcg2017020090-
                                        abs.html.

Liu:1996:AGI

[LJR96] Ning Liu, Houzhi Jin, and Alyn P. Rockwood. Antialiasing by Gaussian integration. IEEE Computer Graphics and Applications, 16(3):58–63, May/ June 1996. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Liewald:1982:IDT

[LK82] Michael H. Liewald and Philip R. Kennicott. Intersystem data transfer via IGES. IEEE Computer Graphics and Applications, 2(3):55–58, 60–63, May/ June 1982. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Lee:1995:MAH

[LK95a] Jintae Lee and Tosiyasu L.

Kunii. Model-based analysis of hand posture. IEEE Computer Graphics and Applications, 15(5):77–86, September 1995. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Loftin:1995:THS

[LK95b] R. Bowen Loftin and Patrick J.

Kenney. Training the Hubble space telescope flight team. IEEE Computer Graphics and Applications, 15(5):31–37, September 1995. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Lau:2003:GEI

[LK03] Rynson W. H. Lau and Tosiyasu L. Kunii. Guest Editors’ introduction: Web graphics. IEEE Computer Graphics and Applications, 23(1):26–27, January/ February 2003. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://dlib.
computer.org/cg/books/
cg2003/pdf/g1026.pdf.

Lockwood:2008:UVH

[LK08] Dave Lockwood and Erik Kruger. Using VR for human development in Africa. IEEE Computer Graphics and Applications, 28(3):99– 103, May/June 2008. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Lassagne:2019:ESF

[LKPM19] Antoine Lassagne, Andras Kemeny, Javier Posselt, and Frédéric Merienne. Evaluation of spatial filtering algorithms for visual interactions in CAVEs. IEEE Computer Graphics and Applications, 39(1):53–63, January/February 2019. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL https:

//ieeexplore.ieee.org/
document/8663524/.

Lee:1996:NSI

[LKS96] Craig A. Lee, Carl Kesselman, and Stephen Schwab. Near-real-time satellite image processing: Metacomputing in CC++. IEEE Computer Graphics and Applications, 16(4):79–84, July/August 1996. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Lathrop:1990:ARS

[LKV90] Olin Lathrop, David Kirk, and Doug Voorhies. Accurate rendering by subpixel addressing. IEEE Computer Graphics and Applications, 10(5):45–53, September 1990. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Lidal:2007:DIO

[LLG^{+}07] Endre M. Lidal, Tor Langeland, Christopher Giertsen, Jens Grimsgaard, and Rolf Helland. A decade of increased oil recovery in virtual reality. IEEE Computer Graphics and Applications, 27(6):94–97, November/December 2007. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Lien:1984:SMC

[lLK84] Sheue ling Lien and James T.

Kajiya. Symbolic method for calculating the integral properties of arbitrary nonconvex polyhedra. IEEE Computer Graphics and Applications, 4(10):35–41, October 1984. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Lau:2003:EWG

[LLK^{+}03] Rynson W. H. Lau, Frederick Li, Tosiyasu L. Kunii, Baining Guo, Bo Zhang, Nadia Magnenat-Thalmann, Sumedha Kshirsagar, Daniel Thalmann, and Mario Gutierrez. Emerging Web graphics standards and technologies. IEEE Computer Graphics and Applications, 23(1):66–75, January/ February 2003. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL http://dlib.
                                        computer.org/cg/books/
                                        cg2003/pdf/g1066.pdf;
                                        http://www.computer.org/
                                        cga/cg2003/g1066abs.htm.

Liang:2013:OSF

[LLL^{+}13] Yun Liang, Yong-Jin Liu, Xiao-Nan Luo, Lexing Xie, and Xiaolan Fu. Optimalscaling-factor assignment for patch-wise image retargeting. IEEE Computer Graphics and Applications, 33(5): 68–78, September/October 2013. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Lavoue:2016:QAP

[LLML16] Guillaume Lavoue, Hantao Liu, Karol Myszkowski, and Weisi Lin. Quality assessment and perception in computer graphics. IEEE Computer Graphics and Applications, 36(4):21–22, July/ August 2016. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2016/04/mcg2016040021.
html.

Linsen:2009:VVL

[LLR09] Lars Linsen, Tran Van Long, and Paul Rosenthal. Visualization viewpoints: Linking multidimensional feature space cluster visualization to multifield surface extraction. IEEE Computer Graphics and Applications, 29(3):85–89, May/ June 2009. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Li:2023:LOS

[LLY23] Hao Li, Changjiang Liu, and Yang Yang. Layer-Net: a one-step layered network for semantic segmentation at night. IEEE Computer Graphics and Applications, 43(6):9–21, November/December 2023. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Liu:2013:HBT

[LLZ13] Guanyang Liu, Keke Lu, and Yuru Zhang. Hapticbased training for tank gunnery using decoupled motion control. IEEE Computer Graphics and Applications, 33(2):73–79, March/ April 2013. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Lance:2010:EGM

[LM10] Brent Lance and Stacy Marsella. The expressive gaze model: Using gaze to express emotion. IEEE Computer Graphics and Applications, 30(4):62–73, July/August 2010. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Leizea:2017:RTV

[LMA^{+}17] Ibai Leizea, Ainitze Mendizabal, Hugo Alvarez, Iker Aguinaga, Diego Borro, and Emilio Sanchez. Realtime visual tracking of deformable objects in robotassisted surgery. IEEE Computer Graphics and Applications, 37(1):56–68, January/February 2017. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL https:/

                                        /www.computer.org/csdl/
                                        mags/cg/2017/01/mcg2017010056-
                                        abs.html.

Lounsbery:1992:PSI

[LMD92] Michael Lounsbery, Stephen Mann, and Tony DeRose. Parametric surface interpolation. IEEE Computer Graphics and Applications, 12(5):45–??, September 1992. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Leith:1989:CGC

[LMP89] ArDean Leith, Michael Marko, and Donald Parsons. Computer graphics for cellular reconstruction. IEEE Computer Graphics and Applications, 9(5):16– 23, September 1989. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Lalioti:2001:PVN

[LMPW01] Vali Lalioti, Andiers Malan, James Pun, and Juergen Wind. Projects in VR: Ndebele painting in VR. IEEE Computer Graphics and Applications, 21(2):10– 13, March/April 2001. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http:

//dlib.computer.org/cg/
books/cg2001/pdf/g2010.
pdf.

Legg:2017:GVF

[LMWC17] Philip A. Legg, Eamonn Maguire, Simon Walton, and Min Chen. Glyph visualization: A fail-safe design scheme based on quasi-Hamming distances. IEEE Computer Graphics and Applications, 37(2):31–41, March/April 2017. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        cg/2017/02/mcg2017020031-
                                        abs.html.

Liu:2011:EPT

[LMZ11] Yong-Jin Liu, Cui-Xia Ma, and Dong-Liang Zhang. EasyToy: Plush toy design using editable sketching curves. IEEE Computer Graphics and Applications, 31(2):49–57, March/ April 2011. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Latta:1994:CVR

[LO94] John N. Latta and David J. Oberg. A conceptual virtual reality model. IEEE Computer Graphics and Applications, 14(1):23–29, January/ February 1994. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Lin:2005:SPH

[LO05] Ming C. Lin and Miguel A. Otaduy. Sensation-preserving haptic rendering. IEEE Computer Graphics and Applications, 25(4):8–11, July/ August 2005. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Lin:2008:GEI

[LOB08] Ming C. Lin, Miguel A.

Otaduy, and Ronan Boulic. Guest Editors’ introduction: Virtual reality software and technology. IEEE Computer Graphics and Applications, 28(6):18–19, November/December 2008. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Lage:2016:SDE

[LOC^{+}16] Marcos Lage, Jorge Piazentin Ono, Daniel Cervone, Justin Chiang, Carlos Dietrich, and Claudio T. Silva. StatCast Dashboard: Exploration of spatiotemporal baseball data. IEEE Computer Graphics and Applications, 36(5):28–37, September/ October 2016. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2016/05/mcg2016050028-
abs.html.

Lodding:1983:II

[Lod83] Kenneth N. Lodding. Iconic interfacing. IEEE Computer Graphics and Applications, 3(2):11–12, 14–16, 18–20, March/April 1983. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Lok:2006:TCS

[Lok06] Benjamin Lok. Teaching communication skills with virtual humans. IEEE Computer Graphics and Applications, 26(3):10–13, May/ June 2006. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Lorensen:2020:HMC

[Lor20]  W. E. Lorensen.  His-

tory of the marching cubes algorithm. IEEE Computer Graphics and Applications, 40(2):8–15, March/ April 2020. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Love:2005:VSM

[LPK05] Alison L. Love, Alex Pang, and David L. Kao. Visualizing spatial multivalue data. IEEE Computer Graphics and Applications, 25(3):69– 79, May/June 2005. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Lawrence:2004:SVH

[LPLN04] Dale A. Lawrence, Lucy Y. Pao, Christopher D. Lee, and Roman Y. Novoselov. Synergistic visual/haptic rendering modes for scientific visualization. IEEE Computer Graphics and Applications, 24(6):22–30, November/December 2004. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL

http://csdl.computer.org/
dl/mags/cg/2004/06/g6022.
htm;   http://csdl.computer.
org/dl/mags/cg/2004/06/
g6022.pdf.

Lopez-Pernas:2022:EDV

[LPMAA^{+}22] Sonsoles López-Pernas, Andres Munoz-Arcentales, Carlos Aparicio, Enrique Barra, Aldo Gordillo, Joaquín Salvachúa, and Juan Quemada. Educational data virtual lab: Connecting the dots between data visualization and analysis. IEEE Computer Graphics and Applications, 42(5):76–83, September/October 2022. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Linares-Pellicer:2011:CGD

[LPMEPCM11] Jordi Linares-Pellicer, Pau Mico, Javier Esparza-Peidro, and Empar Carrasquer-Moya. Computer graphics: From desktop to mobile and Web. IEEE Computer Graphics and Applications, 31(4):94–96, c3, July/ August 2011. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Linares-Pellicer:2018:FGD

[LPOLID18] Jordi Linares-Pellicer, Jorge Orta-Lopez, and Juan Izquierdo-Domenech. Flipping game development. IEEE Computer Graphics and Applications, 38(6):118– 124, ???? 2018. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL https:

                                        //ieeexplore.ieee.org/
                                        document/8494833/.

Levine:2012:MPM

[LPZ12] Joshua A. Levine, Rasmus R. Paulsen, and Yongjie Zhang. Mesh processing in medical-image analysis — a tutorial. IEEE Computer Graphics and Applications, 32(5):22–28, September/October 2012. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Lindstrom:2018:SIV

[LQ18] Peter Lindstrom and Huamin Qu. Special issue: Visualization connections. IEEE Computer Graphics and Applications, 38(6):24–25, ???? 2018. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL https:

                                        //ieeexplore.ieee.org/
                                        document/8617762/.

Lee:1990:NUN

[LR90] Mark E. Lee and Richard A. Redner. A note on the use of nonlinear filtering in computer graphics. IEEE Computer Graphics and Applications, 10(3):23–29, May/ June 1990. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Loscos:2014:NBT

[LR14] Celine Loscos and Holly Rushmeier. The next big thing [Guest Editors’ introduction]. IEEE Computer Graphics and Applications, 34(6):24–25, November/ December 2014. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://csdl.
computer.org/csdl/mags/
cg/2014/06/mcg2014060024.
html.

Losev:2022:EDD

[LRCT22] Tatiana Losev, Justin Raynor, Sheelagh Carpendale, and Melanie Tory. Embracing disciplinary diversity in visualization. IEEE Computer Graphics and Applications, 42(6):64–71, November/December 2022. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Lee:2015:MTT

[LRIC15] Bongshin Lee, Nathalie Henry Riche, Petra Isenberg, and Sheelagh Carpendale. More than telling a story: Transforming data into visually shared stories. IEEE Computer Graphics and Applications, 35(5):84–90, September/October 2015. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
cg/2015/05/mcg2015050084-
abs.html.

Lindeman:2009:PVP

[LRS09] Robert W. Lindeman, Dirk Reiners, and Anthony Steed. Projects in VR: Practicing what we preach: IEEE VR 2009 virtual program committee meeting. IEEE Computer Graphics and Applications, 29(2):80–83, March/ April 2009. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Livnat:2012:EVA

[LRS12] Yarden Livnat, Theresa-Marie Rhyne, and Matthew H. Samore. Epinome: a visualanalytics workbench for epidemiology data. IEEE Computer Graphics and Applications, 32(2):89–95, March/ April 2012. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Lo:2023:SVW

[LRZ23] Wei Hong Lo, Holger Regenbrecht, and Stefanie Zollmann. Sports visualization in the wild: The impact of technical factors on user experience in augmented reality sports spectating. IEEE Computer Graphics and Applications, 43(6):64–74, November/December 2023. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Lehar:1984:HMC

[LS84] A. F. Lehar and R. J.

Stevens. High-speed manipulation of the color chromaticity of digital images. IEEE Computer Graphics and Applications, 4(2):34– 39, February 1984. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Lansdown:1995:ERR

[LS95] John Lansdown and Simon Schofield. Expressive rendering: a review of nonphotorealistic techniques. IEEE Computer Graphics and Applications, 15(3):29–37, May/June 1995. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Lin:2004:HRB

[LS04a] Ming Lin and Kenneth Salisbury. Haptic rendering— beyond visual computing. IEEE Computer Graphics and Applications, 24(2):22– 23, March/April 2004. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

csdl.computer.org/comp/
mags/cg/2004/02/g2022.

pdf;   http://csdl.computer.
org/dl/mags/cg/2004/02/
g2022.htm.

Lin:2004:GEI

[LS04b] Ming C. Lin and Kenneth Salisbury. Guest Editors’ introduction: Haptic rendering—Touch-Enabled interfaces. IEEE Computer Graphics and Applications, 24(6):20–21, November/ December 2004. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cg/2004/06/g6020.pdf;
http://csdl.computer.org/
dl/mags/cg/2004/06/g6020.
htm.

Lamberti:2008:EGR

[LS08] Fabrizio Lamberti and Andrea Sanna. Extensible GUIs for remote application control on mobile devices. IEEE Computer Graphics and Applications, 28(4): 50–57, July/August 2008. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Liu:2021:GMP

[LS21] Ligang Liu and Scott Schaefer. Geometric modeling and processing. IEEE Computer Graphics and Applications, 41(3):18–19, May/ June 2021. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Liu:2018:ADD

[LSB18] Li Liu, Deborah Silver, and Karen Bemis. Applicationdriven design: Help students understand employment and see the “Big Picture”. IEEE Computer Graphics and Applications, 38(3):90–105, May/June 2018. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2018/03/mcg2018030090-
abs.html.

Liu:2021:VAI

[LSB21] L. Liu, D. Silver, and K. Bemis. Visualizing acoustic imaging of hydrothermal plumes on the seafloor. IEEE Computer Graphics and Applications, 41(2): 63–75, March/April 2021. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Luo:2017:ISS

[LSC^{+}17] Wei Luo, Michael Steptoe, Zheng Chang, Robert Link, Leon Clarke, and Ross Maciejewski. Impact of spatial scales on the intercomparison of climate scenarios. IEEE Computer Graphics and Applications, 37(5):40–49, September/October 2017. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/

                                        cg/2017/05/mcg2017050040-
                                        abs.html.

Loefer:1983:IBC

[LSFW83] Gene R. Loefer, David E. Schmieder, William M. Finlay, and Marshall R. Weathersby. An infrared background clutter model using 3-D computer graphics. IEEE Computer Graphics and Applications, 3(2): 55–62, 64–66, March/April 1983. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Loftin:2004:TPO

[LSMC04] R. Bowen Loftin, Mark W. Scerbo, Frederic D. McKenzie, and Jean M. Catanzaro. Training in peacekeeping operations using virtual environments. IEEE Computer Graphics and Applications, 24(4):18–21, July/ August 2004. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http:/

                                        /csdl.computer.org/dl/
                                        mags/cg/2004/04/g4018.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/cg/2004/04/
                                        g4018.pdf.

Lin:2012:SCF

[LSQ^{+}12] Yuxu Lin, Mingli Song, Dao Thi Phuong Quynh, Ying He, and Chun Chen. Sparse coding for flexible, robust 3D facial-expression synthesis. IEEE Computer Graphics and Applications, 32(2):

76–88, March/April 2012. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Leibe:2000:TSI

[LSR^{+}00] Bastian Leibe, Thad Starner, William Ribarsky, Zachary Wartell, David Krum, Justin Weeks, Bradley Singletary, and Larry Hodges. Toward spontaneous interaction with the perceptive workbench. IEEE Computer Graphics and Applications, 20(6):54–65, November/December 2000. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

computer.org/cga/cg2000/
g6054abs.htm;   http:/
/dlib.computer.org/cg/
books/cg2000/pdf/g6054.
pdf.

Lokki:2002:CIV

[LSV^{+}02] Tapio Lokki, Lauri Savioja, Riitta Väänänen, Jyri Huopaniemi, and Tapio Takala. Creating interactive virtual auditory environments. IEEE Computer Graphics and Applications, 22(4):49–57, July/ August 2002. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

computer.org/cga/cg2002/
g4049abs.htm;   http:/
/dlib.computer.org/cg/
books/cg2002/pdf/g4049.
pdf.

Luccioni:2021:UAI

[LSVB21] A. Luccioni, V. Schmidt, V. Vardanyan, and Y. Bengio. Using artificial intelligence to visualize the impacts of climate change. IEEE Computer Graphics and Applications, 41(1):8– 14, January/February 2021. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Losada:2016:BBV

[LTB16] Antonio G. Losada, Roberto Theron, and Alejandro Benito. BKViz: A basketball visual analysis tool. IEEE Computer Graphics and Applications, 36(6):58–68, November/ December 2016. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        cg/2016/06/mcg2016060058-
                                        abs.html.

Latif:2022:TRA

[LTB22] Shahid Latif, Hagen Tarner, and Fabian Beck. Talking realities: Audio guides in virtual reality visualizations. IEEE Computer Graphics and Applications, 42(1):73– 83, January/February 2022. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Le:2016:CHV

[LTD16] Thai Le, Hilaire J. Thompson, and George Demiris. A comparison of health visualization evaluation techniques with older adults. IEEE Computer Graphics and Applications, 36(4):67– 77, July/August 2016. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL https:/

/www.computer.org/csdl/
mags/cg/2016/04/mcg2016040067-
abs.html.

Lischinski:1992:DMA

[LTG92] Dani Lischinski, Filippo Tampieri, and Donald P. Greenberg. Discontinuity meshing for accurate radiosity. IEEE Computer Graphics and Applications, 12(6): 25–39, November/December 1992. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Laramee:2018:VSC

[LTJ18] Robert S. Laramee, Cagatay Turkay, and Alark Joshi. Visualization for smart city applications. IEEE Computer Graphics and Applications, 38(5):36–37, September/October 2018. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2018/05/mcg2018050036.
html.

Lee:2013:FTV

[LTS^{+}13] Teng-Yok Lee, Xin Tong, Han-Wei Shen, Pak Chung Wong, Samson Hagos, and L. Ruby Leung. Feature tracking and visualization of the Madden–Julian oscillation in climate simulation. IEEE Computer Graphics and Applications, 33(4): 29–37, July/August 2013. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Loke:1992:RFD

[LTSE92] Teng-See Loke, Daniel Tan, Hock-Soon Seah, and Meng-Hwa Er. Rendering fireworks displays. IEEE Computer Graphics and Applications, 12(3):33–43, May/ June 1992. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Luebke:2001:DSP

[Lue01] David P. Luebke. A developer’s survey of polygonal simplification algorithms. IEEE Computer Graphics and Applications, 21(3):24– 35, May/June 2001. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

                                        computer.org/cga/cg2001/
                                        g3024abs.htm;   http:/
                                        /dlib.computer.org/cg/
                                        books/cg2001/pdf/g3024.
                                        pdf.

Liu:2023:GIM

[LV23] Liqun Liu and Romain Vuillemot. A generic interactive membership function for categorization of quantities. IEEE Computer Graphics and Applications, 43(5):39–48, September/October 2023. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Lamousin:1994:NFD

[LW94] Henry J. Lamousin and Warren N. Waggenspack, Jr. NURBS-based free-form deformations. IEEE Computer Graphics and Applications, 14(6):59–65, November/December 1994. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Leung:2013:CSS

[LW13] Yuen-Shan Leung and Charlie C. L. Wang. Conservative sampling of solids in image space. IEEE Computer Graphics and Applications, 33(1):32–43, January/ February 2013. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Lu:2014:BIS

[LWM14] Yafeng Lu, Feng Wang, and Ross Maciejewski. Business intelligence from social media: A study from the VAST Box Office Challenge. IEEE Computer Graphics and Applications, 34(5):58–69, September/ October 2014. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://csdl.
computer.org/csdl/mags/
cg/2014/05/mcg2014050058-
abs.html.

Lee:1998:PMA

[LWS98] Seungyong Lee, George Wolberg, and Sung Yong Shin. Polymorph: Morphing among multiple images. IEEE Computer Graphics and Applications, 18(1):58– 71, January/February 1998. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL

http://computer.org/cga/
cg1998/g1058abs.htm;
http://dlib.computer.org/
cg/books/cg1998/pdf/g1058.
pdf.

Lu:2020:ICT

[LWY^{+}20] J. Lu, J. Wang, H. Ye, Y. Gu, Z. Ding, M. Xu, and W. Chen. Illustrating changes in time-series data with data video. IEEE Computer Graphics and Applications, 40(2):18–31, March/ April 2020. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Li:2018:CHV

[LWZ^{+}18] Mingran Li, Wenjie Wu, Junhan Zhao, Keyuan Zhou, David Perkis, Timothy N. Bond, Kevin Mumford, David Hummels, and Yingjie Victor Chen. CareerVis: Hierarchical visualization of career pathway data. IEEE Computer Graphics and Applications, 38(6):96–105, ???? 2018. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL https:

//ieeexplore.ieee.org/
document/8617733/.

Li:2019:PAN

[LWZY19] Z. Li, M. Wu, J. Zheng, and H. Yu. Perceptual adversarial networks with a feature pyramid for image translation. IEEE Computer Graphics and Applications, 39(4):68–77, July/ August 2019. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Luo:2004:PAH

[LX04] Qi Luo and Jing Xiao.

Physically accurate haptic rendering with dynamic effects. IEEE Computer Graphics and Applications, 24(6):60–69, November/December 2004. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http:

//csdl.computer.org/dl/
mags/cg/2004/06/g6060.
htm;   http://csdl.computer.
org/dl/mags/cg/2004/06/
g6060.pdf.

Lorensen:1992:GGV

[LY92] William E. Lorensen and Boris Yamrom. Golf green visualization. IEEE Computer Graphics and Applications, 12(4):35–??, July/ August 1992. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Lin:2002:ERF

[LYO02] I-Chen Lin, Jeng-Sheng Yeh, and Ming Ouhyoung. Extracting realistic 3D facial animation parameters from multiview video clips. IEEE Computer Graphics and Applications, 22(6):72–80, November/ December 2002. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL http://dlib.
                                        computer.org/cg/books/
                                        cg2002/pdf/g6072.pdf;
                                        http://www.computer.org/
                                        cga/cg2002/g6072abs.htm.

Li:2014:ISV

[LZF^{+}14] Qihang Li, Gabriel Zachmann, David Feng, Kun Huang, and Raghu Machiraju. 2013 IEEE Scientific Visualization Contest winner: Observing genomics and phenotypical patterns in the developing mouse brain. IEEE Computer Graphics and Applications, 34(5):88–97, September/October 2014. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://csdl.
computer.org/csdl/mags/
cg/2014/05/mcg2014050088.
html.

Liu:2021:CEE

[LZR^{+}21] Na Liu, Dan Zhang, Xudong Ru, Haichuan Zhao, Xingce Wang, and Zhongke Wu. Caricature expression extrapolation based on Kendall shape space theory. IEEE Computer Graphics and Applications, 41(3):71–84, May/June 2021. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Liu:2004:IBS

[LZS04] Zicheng Liu, Zhengyou Zhang, and Ying Shan. Image-based surface detail transfer. IEEE Computer Graphics and Applications, 24(3):30–35, May/ June 2004. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cg/2004/03/g3030abs.htm;
http://csdl.computer.
org/dl/mags/cg/2004/03/
g3030.htm;   http://csdl.
computer.org/dl/mags/cg/
2004/03/g3030.pdf.

McCormick:1998:VBV

[MA98] Patrick S. McCormick and James P. Ahrens. Visualization blackboard: Visualization of wildfire simulations.

IEEE Computer Graphics and Applications, 18(2):17– 19, March/April 1998. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

computer.org/cga/cg1998/
extras/g2019x.htm;   http:
//dlib.computer.org/cg/
books/cg1998/pdf/g2017.
pdf.

Ma:2000:VVV

[Ma00] Kwan-Liu Ma. Visualization viewpoints: Visualizing visualizations: User interfaces for managing and exploring scientific visualization data. IEEE Computer Graphics and Applications, 20(5):16–19, September/October 2000. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://dlib.
computer.org/cg/books/
cg2000/pdf/g5016.pdf.

Ma:2001:GEI

[Ma01] Kwan-Liu Ma. Guest Editor’s introduction: Largescale data visualization. IEEE Computer Graphics and Applications, 21(4):22– 23, July/August 2001. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http:

//dlib.computer.org/cg/
books/cg2001/pdf/g4022.
pdf.

Ma:2006:GEI

[Ma06] Kwan-Liu Ma. Guest Editor’s introduction: Visualization for cybersecurity. IEEE Computer Graphics and Applications, 26(2):26– 27, March/April 2006. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

csdl.computer.org/comp/
mags/cg/2006/02/g2026.
pdf.

Ma:2007:MLB

[Ma07] Kwan-Liu Ma. Machine learning to boost the next generation of visualization technology. IEEE Computer Graphics and Applications, 27(5):6–9, September/October 2007. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Ma:2009:SVE

[Ma09]  Kwan-Liu Ma.  In situ

visualization at extreme scale: Challenges and opportunities. IEEE Computer Graphics and Applications, 29(6):14–19, November/December 2009. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Ma:2023:PVR

[Ma23] Kwan-Liu Ma. Pushing visualization research frontiers: Essential topics not addressed by machine learning. IEEE Computer Graphics and Applications, 43(1): 97–102, January/February 2023. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Mabry:1982:GNC

[Mab82] J. H. Mabry. Graphics in numerical control — the user’s challenge. IEEE Computer Graphics and Applications, 2(1):45–49, January/ February 1982. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Madhavan:2005:DRE

[MAB05] Krishna P. C. Madhavan, Laura L. Arns, and Gary R. Bertoline. A distributed rendering environment for teaching animation and scientific visualization. IEEE Computer Graphics and Applications, 25(5):32–38, September/October 2005. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Machover:1981:GSI

[Mac81] Carl Machover. A guide to sources of information about computer graphics. IEEE Computer Graphics and Applications, 1(1):73– 75, 78, 80–85, January/ February 1981. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Machover:1983:UGS

[Mac83] Carl Machover. Updated guide to sources of information about computer graphics. IEEE Computer Graphics and Applications, 3(1): 49–59, January/February 1983. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Machover:1985:BSI

[Mac85]  Carl Machover.  Back-

ground and source information about computer graphics. IEEE Computer Graphics and Applications, 5(1): 68–81, January/February 1985. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Maciejewski:1990:DIE

[Mac90] Anthony A. Maciejewski. Dealing with the ill-conditioned equations of motion for articulated figures. IEEE Computer Graphics and Applications, 10(3):63–71, May/ June 1990. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Machover:1991:BCG

[Mac91] Carl Machover. The business of computer graphics. IEEE Computer Graphics and Applications, 11(1):35– 41, January/February 1991. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Machover:1994:FDC

[Mac94]  Carl Machover.  Four

decades of computer graphics. IEEE Computer Graphics and Applications, 14(6): 14–19, November/December 1994. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Machover:1996:THP

[Mac96] Carl Machover. Top 10 hardware products for 1995. IEEE Computer Graphics and Applications, 16(1): 91, January/February 1996. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Macedonia:1997:NCM

[Mac97a] Michael R. Macedonia. In the news: Carl Machover Symposium; CG&A winds the bronze; new 3D technologies take shape; Xrays go digital; tech notes; throwing in the chips; State of the Art opens; business briefs. IEEE Computer Graphics and Applications, 17(4):90–92, July/ August 1997. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Machover:1997:THP

[Mac97b] Carl Machover. The top 10 hardware products of 1996. IEEE Computer Graphics and Applications, 17(1): 88, January/February 1997. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Machover:1998:TTH

[Mac98]  Carl Machover.  Top

ten hardware products of 1997. IEEE Computer Graphics and Applications, 18(1):93, January/ February 1998. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

computer.org/cga/cg1998/
g1094abs.htm;   http:/
/dlib.computer.org/cg/
books/cg1998/pdf/g1094.
pdf.

MacDonald:1999:TUC

[Mac99a] Lindsay W. MacDonald. Tutorial: Using color effectively in computer graphics. IEEE Computer Graphics and Applications, 19(4):20– 35, July/August 1999. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

computer.org/cga/cg1999/
g4020abs.htm;   http:/
/dlib.computer.org/cg/
books/cg1999/pdf/g4020.
pdf.

Machover:1999:THN

[Mac99b] Carl Machover. Top 10 hardware: New products of 1998. IEEE Computer Graphics and Applications, 19(1):92, January/ February 1999. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL http://dlib.
                                        computer.org/cg/books/
                                        cg1999/pdf/g1092.pdf.

Machover:2000:BCG

[Mac00a] Carl Machover. The business of computer graphics. IEEE Computer Graphics and Applications, 20(1):44– 45, January/February 2000. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL

                                        http://dlib.computer.org/
                                        cg/books/cg2000/pdf/g1044.
                                        pdf.

Mackinlay:2000:OIV

[Mac00b] Jock D. Mackinlay. Opportunities for information visualization. IEEE Computer Graphics and Applications, 20(1):22–23, January/ February 2000. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL http://dlib.
                                        computer.org/cg/books/
                                        cg2000/pdf/g1022.pdf.

Machover:2001:TTN

[Mac01] Carl Machover. Top 10 hardware products of 2000.

IEEE Computer Graphics and Applications, 21(1): 93, January/February 2001. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL

http://dlib.computer.org/
cg/books/cg2001/pdf/g1093.
pdf.

Machover:2002:TNP

[Mac02] Carl Machover. 2001 top 10 new products. IEEE Computer Graphics and Applications, 22(1):87–88, January/ February 2002. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://dlib.
computer.org/cg/books/
cg2002/g1087abs.htm;
http://dlib.computer.org/
cg/books/cg2002/pdf/g1087.
pdf.

Machover:2003:THS

[Mac03] Carl Machover. Top 10 hardware and software products of 2002. IEEE Computer Graphics and Applications, 23(1):92–93, January/ February 2003. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://dlib.
computer.org/cg/books/
cg2003/pdf/g1092.pdf.

Machover:2004:THP

[Mac04] Carl Machover. Top hardware products of 2003. IEEE Computer Graphics and Applications, 24(1):92, January/February 2004. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http:

                                        //csdl.computer.org/dl/
                                        mags/cg/2004/01/g1092.
                                        pdf.

Macedonia:2009:GSV

[Mac09] Mike Macedonia. Graphically speaking: Virtual worlds: a new reality for treating post-traumatic stress disorder. IEEE Computer Graphics and Applications, 29(1):86–88, January/ February 2009. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Macedonia:2013:W

[Mac13] Michael Macedonia. 3D for the Web. IEEE Computer Graphics and Applications, 33(5):24–25, September/October 2013. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Makkuni:1987:GRP

[Mak87] Ranjit Makkuni. A gestural representation of the process of composing Chinese temples. IEEE Computer Graphics and Applications, 7(12):45–61, December 1987. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Mammen:1989:TAA

[Mam89] Abraham Mammen. Transparency and antialiasing algorithms implemented with the virtual pixel maps technique. IEEE Computer Graphics and Applications, 9(4):43–55, July/ August 1989. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Manocha:1994:SSP

[Man94]  Dinesh Manocha.  Solv-

ing systems of polynomial equations. IEEE Computer Graphics and Applications, 14(2):46–55, March/ April 1994. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Marcus:1982:DFI

[Mar82]  Aaron Marcus.  Design-

ing the face of an interface. IEEE Computer Graphics and Applications, 2(1):23–26, 28–29, January/ February 1982. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Marcus:1983:SFG

[Mar83] Aaron Marcus. Special feature: Graphic design for computer graphics. IEEE Computer Graphics and Applications, 3(4):63–70, July/ August 1983. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Marcus:1984:CII

[Mar84] Aaron Marcus. Corporate identity for iconic interface design: The graphic design perspective. IEEE Computer Graphics and Applications, 4(12):24–32, December 1984. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Marcos:2007:DAW

[Mar07] Adérito Fernandes Marcos. Digital art: When artistic and cultural muse merges with computer technology. IEEE Computer Graphics and Applications, 27(5):98–103, September/ October 2007. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Massie:1998:TGM

[Mas98] Thomas Massie. A tangible goal for 3D modeling. IEEE Computer Graphics and Applications, 18(3):62– 65, May/June 1998. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

computer.org/cga/cg1998/
g3062abs.htm;   http:/
/dlib.computer.org/cg/
books/cg1998/pdf/g3062.
pdf.

Masia:2018:CID

[Mas18] Belen Masia. Computational imaging and displays: Capturing and displaying richer representations of the world. IEEE Computer Graphics and Applications, 38(2):112–120, March/April 2018. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2018/02/mcg2018020112-
abs.html.

Mathews:1996:VB

[Mat96] G. Jason Mathews. Visualization blackboard. IEEE Computer Graphics and Applications, 16(6):6–9, November/December 1996. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Max:1983:CRM

[Max83] Nelson L. Max. Computer representation of molecular surfaces. IEEE Computer Graphics and Applications, 3(5):21–29, August 1983. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Max:1990:ASD

[Max90] Nelson L. Max. Antialiasing scan-line data. IEEE Computer Graphics and Applications, 10(1):18–30, January/ February 1990. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Mayo:1983:DAL

[May83] John S. Mayo. Design automation: Lessons of the past, challenges for the future. IEEE Computer Graphics and Applications, 3(6):13–18, September 1983. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Monheit:1991:KMH

[MB91] Gary Monheit and Norman I. Badler. A kinematic model of the human spine and torso. IEEE Computer Graphics and Applications, 11(2):29–38, March/ April 1991. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Max:1994:BSV

[MB94] Nelson L. Max and Barry G. Becker. Bump shading for volume textures. IEEE Computer Graphics and Applications, 14(4):18–20, July/August 1994. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

McGrath:2005:VLS

[MB05] Michael B. McGrath and Judith R. Brown. Visual learning for science and engineering. IEEE Computer Graphics and Applications, 25(5):56–63, September/October 2005. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Machover:1995:GEI

[MBB95] Carl Machover, Jeff Bangert, and Colette Bangert. Guest Editor’s introduction. IEEE Computer Graphics and Applications, 15(3):19–??, May/June 1995. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Magro:2020:CBD

[MBSD20] M. Magro, K. Bugeja, S. Spina, and K. Debattista. Cloud-based dynamic GI for shared VR experiences. IEEE Computer Graphics and Applications, 40(5):10–25, September/October 2020. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Morel:2017:TMR

[MBV17] Jules Morel, Alexandra Bac, and Cedric Vega. Terrain model reconstruction from terrestrial LiDAR data using radial basis functions. IEEE Computer Graphics and Applications, 37(5):72–84, September/ October 2017. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.

                                        computer.org/csdl/mags/
                                        cg/2017/05/mcg2017050072-
                                        abs.html.

MacDonald:1985:ACG

[MC85] Connie L. MacDonald and Ian K. Crain. Applied computer graphics in a geographic information system: Problems and successes. IEEE Computer Graphics and Applications, 5(10):34–39, October 1985. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

McClearly:1983:EG

[McC83] G. F. McClearly, Jr. An effective graphic ‘vocabulary’. IEEE Computer Graphics and Applications, 3(2):46– 53, March/April 1983. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Molnar:1994:SCP

[MCEF94] Steven Molnar, Michael Cox, David Ellsworth, and Henry Fuchs. A sorting classification of parallel rendering. IEEE Computer Graphics and Applications, 14(4): 23–32, July/August 1994. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

McGuire:1993:OSE

[McG93] Frank McGuire. The origins of sculpture: Evolutionary 3D design. IEEE Computer Graphics and Applications, 13(1):9–11, January/ February 1993. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Masia:2021:IDS

[MCGS21] Belen Masia, Javier Camon, Diego Gutierrez, and Ana Serrano. Influence of directional sound cues on users exploration across 360^{∘} movie cuts. IEEE Computer Graphics and Applications, 41(4):64–75, July/August 2021. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

McLaughlin:1983:SPI

[McL83] Harry W. McLaughlin.

Shape-preserving planar interpolation: an algorithm. IEEE Computer Graphics and Applications, 3(3):58– 62, 64–67, May/June 1983. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Morris:2006:MGD

[MCP^{+}06] Meredith Ringel Morris, Anthony Cassanego, Andreas Paepcke, Terry Winograd, Ann Marie Piper, and Anqi Huang. Mediating group dynamics through tabletop interface design. IEEE Computer Graphics and Applications, 26(5):65–73, September/October 2006. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Miebach:2022:NMS

[MCS22] Nathalie Miebach, Bruce D. Campbell, and Francesca Samsel. Nathalie Miebach: Sculpted data infused with craftsmanship. IEEE Computer Graphics and Applications, 42(1):7–16, January/ February 2022. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Magnor:2021:RV

[MCSH21] Marcus Magnor, Susana Castillo, and Alexander Sorkine-Hornung. Real VR. IEEE Computer Graphics and Applications, 41(4): 27–28, July/August 2021. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Max:1993:VCM

[MCW93] Nelson Max, Roger Crawfis, and Dean Williams. Visualization for climate modeling. IEEE Computer Graphics and Applications, 13(4): 34–40, July/August 1993. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Machover:1995:NP

[MD95a] Carl Machover and John Dill. New products. IEEE Computer Graphics and Applications, 15(5):93–95, July/August 1995. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Monro:1995:RAD

[MD95b] Donald M. Monro and Frank Dudbridge. Rendering algorithms for deterministic fractals. IEEE Computer Graphics and Applications, 15(1):32–41, January/ February 1995. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Machover:1996:NPa

[MD96a] Carl Machover and John Dill. New products. IEEE Computer Graphics and Applications, 16(5):92–95, September 1996. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Machover:1996:NPb

[MD96b] Carl Machover and John Dill. New products. IEEE Computer Graphics and Applications, 16(6):79–84, November/December 1996. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Machover:1996:NPI

[MD96c] Carl Machover and John Dill. New products: Including the top 10 hardware/ software products for 1995.

IEEE Computer Graphics and Applications, 16(1):93– 95, January/February 1996. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Machover:1997:NPa

[MD97a] Carl Machover and John Dill. New products. IEEE Computer Graphics and Applications, 17(5):85–87, September/October 1997. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL

http://dlib.computer.org/
cg/books/cg1997/pdf/g5085.
pdf.

Machover:1997:NPb

[MD97b] Carl Machover and John Dill. New products. IEEE Computer Graphics and Applications, 17 (6):104–109, November/ December 1997. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://dlib.
computer.org/cg/books/
cg1997/pdf/g6104.pdf.

Machover:1998:NPa

[MD98a] Carl Machover and John Dill. New products. IEEE Computer Graphics and Applications, 18(1):95–96, January/February 1998. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

computer.org/cga/cg1998/

g1095abs.htm;   http:/
/dlib.computer.org/cg/
books/cg1998/pdf/g1095.
pdf.

Machover:1998:NPb

[MD98b] Carl Machover and John Dill. New products. IEEE Computer Graphics and Applications, 18(2):109–111, March/April 1998. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://dlib.
computer.org/cg/books/
cg1998/pdf/g2109.pdf.

Machover:1998:NPc

[MD98c] Carl Machover and John Dill. New products. IEEE Computer Graphics and Applications, 18(3):91–95, May/June 1998. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://dlib.
computer.org/cg/books/
cg1998/pdf/g3091.pdf.

Machover:1998:NPd

[MD98d] Carl Machover and John Dill. New products. IEEE Computer Graphics and Applications, 18(4):92–95, July/August 1998. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Machover:1998:NPe

[MD98e] Carl Machover and John Dill. New products. IEEE Computer Graphics and Applications, 18(5):90–95, September/October 1998. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL

http://dlib.computer.org/
cg/books/cg1998/pdf/g5090.
pdf.

Machover:1998:NPf

[MD98f] Carl Machover and John Dill. New products. IEEE Computer Graphics and Applications, 18(6):81–85, November/December 1998. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL

http://dlib.computer.org/
cg/books/cg1998/pdf/g6081.
pdf.

Machover:1999:NPa

[MD99a] Carl Machover and John Dill. New products. IEEE Computer Graphics and Applications, 19(1):94–95, January/February 1999. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http:

//dlib.computer.org/cg/
books/cg1999/pdf/g1094.
pdf.

Machover:1999:NPb

[MD99b] Carl Machover and John Dill. New products. IEEE Computer Graphics and Applications, 19(4):94–96, July/August 1999. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://dlib.
computer.org/cg/books/
cg1999/pdf/g4094.pdf.

Machover:1999:NPc

[MD99c] Carl Machover and John Dill. New products. IEEE Computer Graphics and Applications, 19(6):93–99, November/December 1999. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL

http://dlib.computer.org/
cg/books/cg1999/pdf/g6093.
pdf.

Machover:1999:NPHc

[MD99d] Carl Machover and John Dill. New products: Hardware: Board handles volume 3D in real time; TDZ 2000 adds graphics accelerator; Evans & Sutherland simulates reality; Dukane makes compact SVGA projector; digital video camera for the Internet. software: Inscribe 3.0 provides digitizing interface; collaboration through E-mail; Okino offers animation translators; Design Suite targets CAD/CAM; 3D applications for the office; Thomas Publishing releases CAD-Blocks 2.0; visualization enhances spreadsheets; Altia updates simulation graphics software; Framework provides Web support to projects; view markup via the Web. IEEE Computer Graphics and Applications, 19(5):85–88, September/October 1999. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://dlib.
computer.org/cg/books/
cg1999/pdf/g5085.pdf.

Machover:1999:NPHb

[MD99e] Carl Machover and John Dill. New products: Hardware: Displays: Touchscreen resists scratches; Proview introduces 17-inch color monitor; Dimension releases 3D display; transportable virtual display available. processors: Lowpriced engine renders animation; 3Dlabs offers graphics processor. digitizers: 3D digitizer generates objects fast. workstations: Work module for high-tech professionals. software: Life cycle management: Windchill 2.0 manages life cycles. visualization: Data modeling for river network analysis; software for visual simulation released; AEA releases CFX-5 for CFD analysis. animation: Cinema 4D Go entry-level version out. modeling: Modeling tool for 3D offered; 3D Systems announces 3D lightyear. CAD/CAM: Framework releases model Web publisher; Thomas releases updated CAD titles; Internet publishing demonstration launched; Vision Numeric offers 3D reproduction; Cadcentre launches VECnt; Cadcentre announces Review LE. virtual reality: Virtual reality for business presentations. IEEE Computer Graphics and Applications, 19(3):99–103, May/ June 1999. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://dlib.
computer.org/cg/books/
cg1999/pdf/g3099.pdf.

Machover:1999:NPHa

[MD99f] Carl Machover and John Dill. New products: Hardware: Peripherals: Mita offers desktop color laser printer; Proview announces low priced monitors; digital flat-panel solution pack shipped; compact projector from Dukane. systems: PC real-time 3D simulation introduced. boards: Mitsubishi vg500 does volume rendering. software: Web tools: 3DML released by Flatland; Sun delivers Java 2; doc-to-net introduced by Skyline Tools; software makes Web photo editing easy; JustEdit speeds up Web edits. authoring tools: Director 7 and Dreamweaver introduced; Famous offers animation solutions. CAD/CAM: 3D Builder-Pro shipped; Toolbox/SM available from Cimlogic; AutoPlant 97 V1.10 plant design tool; productivity booster for DataCad; enhancements to dv/MockUp functionality. visualization tools: TerraVista Version 2.0 announced. color: Color matching system by Color Solutions. IEEE Computer Graphics and Applications, 19(2):96–99, March/ April 1999. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL http://dlib.
                                        computer.org/cg/books/
                                        cg1999/pdf/g2096.pdf.

Machover:2000:TTN

[MD00a] Carl Machover and John Dill. 1999 top ten new products. IEEE Computer Graphics and Applications, 20(1):76–77, January/ February 2000. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL http://dlib.
                                        computer.org/cg/books/
                                        cg2000/pdf/g1076.pdf.

Machover:2000:NPa

[MD00b] Carl Machover and John Dill. New products. IEEE Computer Graphics and Applications, 20(1):78–79, January/February 2000. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http:

                                        //dlib.computer.org/cg/
                                        books/cg2000/pdf/g1078.
                                        pdf.

Machover:2000:NPb

[MD00c] Carl Machover and John Dill. New products. IEEE Computer Graphics and Applications, 20(2):94–96, March/April 2000. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://dlib.
computer.org/cg/books/
cg2000/pdf/g2094.pdf.

Machover:2000:NPc

[MD00d] Carl Machover and John Dill. New products. IEEE Computer Graphics and Applications, 20(4):110–112, July/August 2000. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://dlib.
computer.org/cg/books/
cg2000/pdf/g4110.pdf.

Machover:2000:NPd

[MD00e] Carl Machover and John Dill. New products. IEEE Computer Graphics and Applications, 20(5):85–88, September/October 2000. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL

http://dlib.computer.org/
cg/books/cg2000/pdf/g5085.
pdf.

Machover:2000:NPH

[MD00f] Carl Machover and John Dill. New products: Hardware: Modeling system for office environment; smart fabric control surface support; head tracker enables VR; programmable 3D/2D input device; digital projectors; portable 3D scene digitizer; slim-line panel PC. software: Visual effects upgrade; mobile CAD; 3D development tool; visual interpretation of financial data; software development for Linux; feature recognition program; Altia Java connection. IEEE Computer Graphics and Applications, 20(6):110–112, November/ December 2000. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://dlib.
computer.org/cg/books/
cg2000/pdf/g6110.pdf.

Machover:2001:NPa

[MD01a] Carl Machover and John Dill. New products. IEEE Computer Graphics and Applications, 21(1):95–96, January/February 2001. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http:

//dlib.computer.org/cg/
books/cg2001/pdf/g1095.
pdf.

Machover:2001:NPb

[MD01b] Carl Machover and John Dill. New products. IEEE Computer Graphics and Applications, 21(2):93–96, March/April 2001. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://dlib.
computer.org/cg/books/
cg2001/pdf/g2093.pdf.

Machover:2001:NPc

[MD01c] Carl Machover and John Dill. New products. IEEE Computer Graphics and Applications, 21(3):93–96, May/June 2001. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://dlib.
computer.org/cg/books/
cg2001/pdf/g3093.pdf.

Machover:2001:NPd

[MD01d] Carl Machover and John Dill. New products. IEEE Computer Graphics and Applications, 21(4):93–96, July/August 2001. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://dlib.
computer.org/cg/books/
cg2001/pdf/g4093.pdf.

Machover:2001:NPe

[MD01e] Carl Machover and John Dill. New products. IEEE Computer Graphics and Applications, 21(5):96, C3, September/October 2001. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Machover:2001:NPf

[MD01f] Carl Machover and John Dill. New products. IEEE Computer Graphics and Applications, 21(6):83–86, November/December 2001. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL

http://dlib.computer.org/
cg/books/cg2001/g6083abs.
htm;   http://dlib.computer.
org/cg/books/cg2001/pdf/
g6083.pdf.

Machover:2002:NPa

[MD02a] Carl Machover and John Dill. New products. IEEE Computer Graphics and Applications, 22(2):92–96, March/April 2002. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

computer.org/cga/cg2002/
g2092abs.htm;   http:/
/dlib.computer.org/cg/
books/cg2002/pdf/g2092.
pdf.

Machover:2002:NPb

[MD02b] Carl Machover and John Dill. New products. IEEE Computer Graphics and Applications, 22(3):87–88, May/June 2002. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

computer.org/cga/cg2002/
g3087abs.htm;   http:/
/dlib.computer.org/cg/
books/cg2002/pdf/g3087.
pdf.

Machover:2002:NPc

[MD02c] Carl Machover and John Dill. New products. IEEE Computer Graphics and Applications, 22(4):103–104, c3, July/August 2002. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

computer.org/cga/cg2002/
g4103abs.htm;   http:/
/dlib.computer.org/cg/
books/cg2002/pdf/g4103.
pdf.

Machover:2002:NPd

[MD02d] Carl Machover and John Dill. New products. IEEE Computer Graphics and Applications, 22 (5):86–88, c3, September/ October 2002. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://dlib.
computer.org/cg/books/
cg2002/pdf/g5086.pdf.

Mann:2002:GAC

[MD02e] Stephen Mann and Leo Dorst. Geometric algebra: a computational framework for geometrical applications (Part 2). IEEE Computer Graphics and Applications, 22(4):58–67, July/ August 2002. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

computer.org/cga/cg2002/
g4058abs.htm;   http:/
/dlib.computer.org/cg/
books/cg2002/pdf/g4058.
pdf.

Meyer:2018:RRA

[MD18] Miriah Meyer and Jason Dykes. Reflection on reflection in applied visualization research. IEEE Computer Graphics and Applications, 38(6):9–16, ???? 2018. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL https:

                                        //ieeexplore.ieee.org/
                                        document/8617760/.

Machover:2002:NPe

[ME02] Carl Machover and Miguel Encarnação. New products. IEEE Computer Graphics and Applications, 22(6):132–135, November/ December 2002. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL http://dlib.
                                        computer.org/cg/books/
                                        cg2002/pdf/g6132.pdf.

Machover:2003:NPa

[ME03a] Carl Machover and Miguel Encarnação. New products. IEEE Computer Graphics and Applications, 23(1):94– 96, January/February 2003. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL

                                        http://dlib.computer.org/
                                        cg/books/cg2003/pdf/g1094.
                                        pdf.

Machover:2003:NPb

[ME03b] Carl Machover and Miguel Encarnação. New products. IEEE Computer Graphics and Applications, 23(2):92– 95, March/April 2003. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http:

//dlib.computer.org/cg/
books/cg2003/pdf/g2092.
pdf.

Machover:2003:NPc

[ME03c] Carl Machover and Miguel Encarnação. New products. IEEE Computer Graphics and Applications, 23(3):92– 95, May/June 2003. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http:

//dlib.computer.org/cg/
books/cg2003/pdf/g3092.
pdf.

Machover:2003:NPd

[ME03d] Carl Machover and Miguel Encarnação. New products. IEEE Computer Graphics and Applications, 23(4):102– 104, July/August 2003. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http:

//csdl.computer.org/dl/
mags/cg/2003/04/g4102.
htm;   http://csdl.computer.
org/dl/mags/cg/2003/04/
g4102.pdf.

Machover:2003:NPe

[ME03e] Carl Machover and Miguel Encarnação. New products. IEEE Computer Graphics and Applications, 23(5):94–96, September/ October 2003. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http:

                                        //csdl.computer.org/dl/
                                        mags/cg/2003/05/g5094.
                                        pdf.

Machover:2003:NPf

[ME03f] Carl Machover and Miguel Encarnação. New products. IEEE Computer Graphics and Applications, 23(6):98–101, November/ December 2003. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http:

                                        //csdl.computer.org/dl/
                                        mags/cg/2003/06/g6098.
                                        pdf.

Machover:2004:NP

[ME04] Carl Machover and Miguel Encarnaçao. New products. IEEE Computer Graphics and Applications, 24(1):94– 95, January/February 2004. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL

                                        http://csdl.computer.org/
                                        dl/mags/cg/2004/01/g1094.
                                        pdf.

McGraw:2022:HRI

[ME22] Tim McGraw and Michael Eddy. Hybrid rendering for interactive visualization of mantle convection. IEEE Computer Graphics and Applications, 42(6):96–106, November/December 2022.

CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Muller:1993:GCV

[MED^{+}93] Jan-Peter Muller, Philip Eales, Tim Day, Lee Kellgren, Athula Mandanayake, Andrew Newton, David Rees, Sam Richards, Kevin Tildsley, Gunter Schreier, Heidelotte Craubner, Hilko Hoffmann, Robert Meisner, Peter Schickl, and Angelika Schnagl. Global change video: Visualization freeze-frames. IEEE Computer Graphics and Applications, 13(3):11–13, May/ June 1993. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Meier:1985:BPT

[Mei85] Barbara Meier. BUCOLIC:

a program for teaching color theory to art students. IEEE Computer Graphics and Applications, 5(7):57–65, July/ August 1985. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Menon:1994:CSR

[Men94] Jai P. Menon. Constructive shell representations for freeform surfaces and solids. IEEE Computer Graphics and Applications, 14(2): 24–36, March/April 1994. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Machover:1998:GEG

[MEW98] Carl Machover, Nick England, and Turner Whitted. Graphics in entertainment: Guest Editors’ introduction. IEEE Computer Graphics and Applications, 18(1):22–23, January/ February 1998. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

                                        computer.org/cga/cg1998/
                                        g1022abs.htm;   http:/
                                        /dlib.computer.org/cg/
                                        books/cg1998/pdf/g1022.
                                        pdf.

Mendelbaum:1989:CSG

[MF89] H. G. Mendelbaum and David Finkelman. CASDA: Synthesized graphic design of real-time systems. IEEE Computer Graphics and Applications, 9(1):40–46, January/February 1989. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Manghisi:2017:ESS

[MFG^{+}17] Vito M. Manghisi, Michele Fiorentino, Michele Gattullo, Antonio Boccaccio, Vitoantonio Bevilacqua, Giuseppe L. Cascella, Michele Dassisti, and Antonio E. Uva. Experiencing the sights, smells, sounds, and climate of Southern Italy in VR. IEEE Computer Graphics and Applications, 37(6):19–25, November/ December 2017. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2017/06/mcg2017060019-
abs.html.

Ma:2008:VRG

[MFL08] Kwan-Liu Ma, Issei Fujishiro, and Hua Li. Visualization research is growing and expanding. IEEE Computer Graphics and Applications, 28(5):22–23, September/October 2008. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cg/2008/05/mcg2008050022.
pdf.

Mcloughlin:2016:VSP

[MFM^{+}16] Leigh Mcloughlin, Oleg Fryazinov, Mark Moseley, Mathieu Sanchez, Valery Adzhiev, Peter Comninos, and Alexander Pasko. Virtual sculpting and 3D printing for young people with disabilities. IEEE Computer Graphics and Applications, 36(1):22–28, January/ February 2016. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
cg/2016/01/mcg2016010022-
abs.html.

Max:1988:SHM

[MG88a] Nelson L. Max and Elizabeth D. Getzoff. Spherical harmonic molecular surfaces. IEEE Computer Graphics and Applications, 8(4):42–50, July/ August 1988. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Meyer:1988:CVC

[MG88b] Gary W. Meyer and Donald P. Greenberg. Colordefective vision and computer graphics displays. IEEE Computer Graphics and Applications, 8(5): 28–40, September/October 1988. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Miller:1992:UTB

[MG92] James R. Miller and Ronald N. Goldman. Using tangent balls to find plane sections of natural quadrics. IEEE Computer Graphics and Applications, 12(2): 68–82, March/April 1992. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Miller:2004:IVP

[MG04] James R. Miller and Estela A. Gavosto. The immersive visualization probe for exploring n-dimensional spaces. IEEE Computer Graphics and Applications, 24(1):76–85, January/ February 2004. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cg/2004/01/g1076abs.htm;
http://csdl.computer.
org/dl/mags/cg/2004/01/
g1076.pdf.

Michalski:2014:DEM

[MG14] Rafal Michalski and Jerzy Grobelny. Designing emergencymedical-service helicopter interiors using virtual manikins. IEEE Computer Graphics and Applications, 34(2): 16–23, March/April 2014. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Mansoor:2021:VAS

[MGA^{+}21] Hamid Mansoor, Walter Gerych, Abdulaziz Alajaji, Luke Buquicchio, Kavin Chandrasekaran, Emmanuel Agu, and Elke A. Rundensteiner. Visual analytics of smartphone-sensed human behavior and health. IEEE Computer Graphics and Applications, 41(3):96– 104, May/June 2021. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Miles:2020:STS

[MGC20] H. C. Miles, M. D. Gunn, and A. J. Coates. Seeing through the science eyes of the ExoMars Rover. IEEE Computer Graphics and Applications, 40(2):71–81, March/April 2020. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Mueller:2011:CCM

[MGN^{+}11] Klaus Mueller, Supriya Garg, Julia EunJu Nam, Tamara Berg, and Kevin T. McDonnell. Can computers master the art of communication?: a focus on visual analytics. IEEE Computer Graphics and Applications, 31(3):14–21, May/ June 2011. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Machiraju:2014:VAB

[MGO14] Raghu Machiraju, Carsten Gorg, and Arthur Olson. Visual analytics for biological data. IEEE Computer Graphics and Applications, 34(2):24–25, March/ April 2014. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Martinez:2016:IVG

[MGO^{+}16] Jonatan Martinez, Arturo Garcia, Miguel Oliver, Jose Pascual Molina, and Pascual Gonzalez. Identifying virtual 3D geometric shapes with a vibrotactile glove. IEEE Computer Graphics and Applications, 36(1):42–51, January/ February 2016. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
cg/2016/01/mcg2016010042-
abs.html.

MacEachren:2004:GKC

[MGP^{+}04] Alan M. MacEachren, Mark Gahegan, William Pike, Isaac Brewer, Guoray Cai, Eugene Lengerich, and Frank Hardisty. Geovisualization for knowledge construction and decision support. IEEE Computer Graphics and Applications, 24(1):13–17, January/February 2004. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http:

//csdl.computer.org/dl/
mags/cg/2004/01/g1013.
pdf.

Meuschke:2018:MCA

[MGW^{+}18] Monique Meuschke, Tobias Gunther, Ralph Wickenhofer, Markus Gross, Bernhard Preim, and Kai Lawonn. Management of cerebral aneurysm descriptors based on an automatic ostium extraction. IEEE Computer Graphics and Applications, 38(3):58–72, May/June 2018. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        cg/2018/03/mcg2018030058-
                                        abs.html.

Mahvash:2004:HFH

[MH04] Mohsen Mahvash and Vincent Hayward. Highfidelity haptic synthesis of contact with deformable bodies. IEEE Computer Graphics and Applications, 24(2):48–55, March/ April 2004. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL http://csdl.
                                        computer.org/comp/mags/
                                        cg/2004/02/g2048abs.htm;
                                         http://csdl.computer.
                                        org/dl/mags/cg/2004/02/
                                        g2048.htm;   http://csdl.
                                        computer.org/dl/mags/cg/
                                        2004/02/g2048.pdf.

Miller:1998:FTC

[MHC98] Gavin Miller, Mark Halstead, and Michael Clifton. On-the-fly texture computation for real-time surface shading. IEEE Computer Graphics and Applications, 18(2):44–58, March/ April 1998. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

                                        computer.org/cga/cg1998/
                                        g2044abs.htm;   http:/
                                        /dlib.computer.org/cg/
                                        books/cg1998/pdf/g2044.
                                        pdf.

Maciejewski:2009:GSS

[MHR^{+}09] Ross Maciejewski, Ryan Hafen, Stephen Rudolph, George Tebbetts, William S. Cleveland, Shaun J. Grannis, and David S. Ebert. Generating synthetic syndromicsurveillance data for evaluating visual-analytics techniques. IEEE Computer Graphics and Applications, 29(3):18–28, May/ June 2009. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Matzen:2023:NVR

[MHTD23] Laura E. Matzen, Breannan C. Howell, Michael C. S. Trumbo, and Kristin M. Divis. Numerical and visual representations of uncertainty lead to different patterns of decision making. IEEE Computer Graphics and Applications, 43(5): 72–82, September/October 2023. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Maciejewski:2008:MSA

[MIA^{+}08] Ross Maciejewski, Tobias Isenberg, William M. Andrews, David S. Ebert, Mario Costa Sousa, and Wei Chen. Measuring stipple aesthetics in hand-drawn and computer-generated images. IEEE Computer Graphics and Applications, 28(2):62–74, March/April 2008. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Mynatt:2000:DAW

[MIEL00] Elizabeth D. Mynatt, Takeo Igarashi, W. Keith Edwards, and Anthony LaMarca. Designing an augmented writing surface. IEEE Computer Graphics and Applications, 20(4):55–61, July/ August 2000. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

                                        computer.org/cga/cg2000/
                                        g4055abs.htm;   http:/
                                        /dlib.computer.org/cg/
                                        books/cg2000/pdf/g4055.
                                        pdf.

Miller:1986:SSS

[Mil86] James R. Miller. Sculptured surfaces in solid models: Issues and alternative approaches. IEEE Computer Graphics and Applications, 6(12):37–48, December 1986. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Miller:1988:NP

[Mil88a]  Gavin Miller.  Natural

phenomena. IEEE Computer Graphics and Applications, 8(5):4–7, September/October 1988. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Miller:1988:AQS

[Mil88b] James R. Miller. Analysis of quadric-surface based solid models. IEEE Computer Graphics and Applications, 8(1):28–42, January/ February 1988. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Miller:1989:AIS

[Mil89] James R. Miller. Architectural issues in solid modelers. IEEE Computer Graphics and Applications, 9(5):72–87, September 1989. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Miller:1993:IBE

[Mil93] James R. Miller. Incremental boundary evaluation using inference of edge classifications. IEEE Computer Graphics and Applications, 13(1):71–78, January/ February 1993. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Miller:1999:AVG

[Mil99a] James R. Miller. Applications of vector geometry for robustness and speed. IEEE Computer Graphics and Applications, 19(4):68– 73, July/August 1999. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

computer.org/cga/cg1999/
g4068abs.htm;   http:/
/dlib.computer.org/cg/
books/cg1999/pdf/g4068.
pdf.

Miller:1999:VGC

[Mil99b] James R. Miller. Vector geometry for computer graphics. IEEE Computer Graphics and Applications, 19(3): 66–73, May/June 1999. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

computer.org/cga/cg1999/
g3066abs.htm;   http:/
/dlib.computer.org/cg/
books/cg1999/pdf/g3066.
pdf.

Mills:2003:DTS

[Mil03] Chad Mills. 3D design tools speed NASA Space Shuttle work. IEEE Computer Graphics and Applications, 23(3):14–19, May/ June 2003. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://dlib.
computer.org/cg/books/
cg2003/pdf/g3014.pdf.

Miller:2007:ABV

[Mil07] James R. Miller. Attribute blocks: Visualizing multiple continuously defined attributes. IEEE Computer Graphics and Applications, 27(3):57–69, May/ June 2007. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Miller:2012:RTV

[Mil12] James R. Miller. Real-time visualization of domain coverage by dynamically moving sensors. IEEE Computer Graphics and Applications, 32(4):8–13, July/ August 2012. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Mol:2008:UGE

[MJC08] Antônio Carlos A. Mól, Carlos Alexandre F. Jorge, and Pedro M. Couto. Using a game engine for VR simulations in evacuation planning. IEEE Computer Graphics and Applications, 28(3):6–12, May/June 2008. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

MacLeod:1993:VBV

[MJM93] Robert S. MacLeod, Christopher R. Johnson, and Mike A. Matheson. Visualization blackboard: Visualizing bioelectric fields. IEEE Computer Graphics and Applications, 13(4): 10–12, July/August 1993. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Munzner:2006:NNV

[MJM^{+}06] Tamara Munzner, Chris Johnson, Robert Moorhead, Hanspeter Pfister, Penny Rheingans, and Terry S. Yoo. NIH–NSF Visualization Research Challenges Report summary. IEEE Computer Graphics and Applications, 26(2):20–24, March/April 2006. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Mahyar:2020:DTS

[MJS20] N. Mahyar, M. Jasim, and A. Sarvghad. Designing technology for sociotechnical problems: Challenges and considerations. IEEE Computer Graphics and Applications, 40(6):76–87, November/December 2020. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Mudur:1984:IMP

[MK84] S. P. Mudur and P. A. Koparkar. Interval methods for processing geometric objects. IEEE Computer Graphics and Applications, 4(2):7–17, February 1984. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Machover:2007:TPa

[MK07a] Carl Machover and David J. Kasik. Tools and products. IEEE Computer Graphics and Applications, 27(4):96, 94–95, July/ August 2007. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://bell.
computer.org/dlcomments/
;   http://csdl.computer.
org/comp/mags/cg/2007/
04/g4096.pdf.

Machover:2007:TPb

[MK07b] Carl Machover and David J.

Kasik. Tools and products. IEEE Computer Graphics and Applications, 27(5):104, 89–91, September/October 2007. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://bell.
computer.org/dlcomments/
;   http://csdl.computer.
org/comp/mags/cg/2007/
05/g5104.pdf.

Machover:2007:TPc

[MK07c] Carl Machover and David J.

Kasik. Tools and products. IEEE Computer Graphics and Applications, 27 (6):100, 98–99, November/ December 2007. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cg/2007/06/mcg2007060100.
pdf.

Mao:1987:HRG

[MKFN87] Xiaoyang Mao, Tosiyasu L.

Kunii, Issei Fujishiro, and Tsukasa Noma. Hierarchical representations of 2D/3D gray-scale images and their 2D/3D two-way conversion. IEEE Computer Graphics and Applications, 7(12):37– 44, December 1987. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Mashal:2020:DYF

[MKH20] S. Mashal, M. Kranz, and G. Hoelzl. Do you feel like flying? A study of flying perception in virtual reality for future game development. IEEE Computer Graphics and Applications, 40(4):51–61, July/ August 2020. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Moezzi:1996:RMV

[MKKJ96] Saied Moezzi, Arun Katkere, Don Y. Kuramura, and Ramesh Jain. Reality modeling and visualization from multiple video sequences. IEEE Computer Graphics and Applications, 16(6): 58–63, November/December 1996. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Mostafawy:1997:VBV

[MKL97] Sina Mostafawy, Omid Kermani, and Holger Lubatschowski. Visualization blackboard: Virtual eye: Retinal image visualization of the human eye. IEEE Computer Graphics and Applications, 17(1):8– 12, January/February 1997. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

MacIntyre:2005:GEI

[ML05] Blair MacIntyre and Mark A.

Livingston. Guest Editors’ introduction: Moving mixed reality into the real world. IEEE Computer Graphics and Applications, 25(6):22–23, November/December 2005. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

csdl.computer.org/comp/
mags/cg/2005/06/g6022.
pdf.

Macedonia:2016:CGD

[ML16] Mike Macedonia and Mark A.

Livingston. Computer graphics for defense applications. IEEE Computer Graphics and Applications, 36(6):16–17, November/ December 2016. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2016/06/mcg2016060016.
html.

Ma:2012:SSU

[MLF^{+}12] Kwan-Liu Ma, Isaac Liao, Jennifer Frazier, Helwig Hauser, and Helen-Nicole Kostis. Scientific storytelling using visualization. IEEE Computer Graphics and Applications, 32(1):12– 19, January/February 2012. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Marai:2019:IAL

[MLJ19] G. E. Marai, J. Leigh, and A. Johnson. Immersive analytics lessons from the Electronic Visualization Laboratory: A 25-year perspective. IEEE Computer Graphics and Applications, 39(3):54–66, May/ June 2019. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Musgrave:1989:NEM

[MM89] F. Kenton Musgrave and Benoit B. Mandelbrot. Natura ex machina. IEEE Computer Graphics and Applications, 9(1):4–7, January/February 1989. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Mustafa:2014:EMW

[MM14] Maryam Mustafa and Marcus Magnor. ElectroEncephaloGraphics: Making waves in computer graphics research. IEEE Computer Graphics and Applications, 34(6):46–56, November/ December 2014. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://csdl.
computer.org/csdl/mags/
cg/2014/06/mcg2014060046-
abs.html.

Morago:2021:EFE

[MM21] B. Morago and D. Monahan. Enhancing film education learning outcomes with virtual experiences. IEEE Computer Graphics and Applications, 41(2): 99–105, March/April 2021. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Murugesan:2019:DVI

[MMD^{+}19] S. Murugesan, S. Malik, F. Du, E. Koh, and T. M. Lai. DeepCompare: Visual and interactive comparison of deep learning model performance. IEEE Computer Graphics and Applications, 39(5):47–59, September/October 2019. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Merillou:2012:SHS

[MMGG12] Nicolas Merillou, Stephane Merillou, Eric Galin, and Djamchid Ghazanfarpour. Simulating how salt decay ages buildings. IEEE Computer Graphics and Applications, 32(2):44–54, March/ April 2012. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Molnar:2015:UVD

[MML^{+}15] Andreea Molnar, Ann F. McKenna, Qing Liu, Mihaela Vorvoreanu, and Krishna Madhavan. Using visualization to derive insights from research funding portfolios. IEEE Computer Graphics and Applications, 35(3):91–c3, May/ June 2015. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL http://www.
                                        computer.org/csdl/mags/
                                        cg/2015/03/mcg2015030091-
                                        abs.html.

Mathis:2021:EDS

[MMMM21] B. Mathis, Y. Ma, M. Mancenido, and R. Maciejewski. Exploring the design space of Sankey diagrams for the food–energy–water nexus. IEEE Computer Graphics and Applications, 41(2): 25–34, March/April 2021. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Ma:2021:VAS

[MMN^{+}21] Yuxin Ma, Prannoy Chandra Pydi Medini, Jake R. Nelson, Ran Wei, Tony H. Grubesic, Jorge A. Sefair, and Ross Maciejewski. A visual analytics system for oil spill response and recovery. IEEE Computer Graphics and Applications, 41(6):91–100, November/ December 2021. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

McGuire:2007:OST

[MMP^{+}07] Morgan McGuire, Wojciech Matusik, Hanspeter Pfister, Billy Chen, John F. Hughes, and Shree K. Nayar. Optical splitting trees for highprecision monocular imaging. IEEE Computer Graphics and Applications, 27(2): 32–42, March/April 2007. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Mitasova:2006:RTL

[MMR^{+}06] Helena Mitasova, Lubos Mitas, Carlo Ratti, Hiroshi Ishii, Jason Alonso, and Russell S. Harmon. Realtime landscape model interaction using a tangible geospatial modeling environment. IEEE Computer Graphics and Applications, 26(4):55–63, July/ August 2006. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Meinzer:1991:HRT

[MMS^{+}91] Hans-Peter Meinzer, Kirsten Meetz, Dinu Scheppelmann, Uwe Engelmann, and Hans Jurgen Baur. The Heidelberg ray tracing model.

IEEE Computer Graphics and Applications, 11(6): 34–43, November/December 1991. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Menon:1994:MPS

[MMZ94] Jai Menon, Richard J. Marisa, and Jovan Zagajac. More powerful solid modeling through ray representations. IEEE Computer Graphics and Applications, 14(3):22–35, May/ June 1994. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Mehl:1984:VSG

[MN84] Max E. Mehl and Stefan J. Noll. A VLSI support for GKS. IEEE Computer Graphics and Applications, 4(8):52–55, August 1984. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Macedonia:1997:ATR

[MN97] Michael R. Macedonia and Stefan Noll. An application: a transatlantic research and development environment. IEEE Computer Graphics and Applications, 17(2):76–82, March/ April 1997. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Morales:2022:BIV

[MNC^{+}22] Leonardo Morales, Pablo Navarro, Celia Cintas, Rolando González-José, Virginia Ramallo, and Claudio Delrieux. Bulsarapp: Interactive visual analysis for surname trend exploration. IEEE Computer Graphics and Applications, 42(4): 28–39, July/August 2022. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Maguire:1991:GAM

[MNR^{+}91] Gerald Q. Maguire, Jr., Marilyn E. Noz, Henry Rusinek, Judith Jaeger, Elissa L. Kramer, Joseph J. Sanger, and Gwenn Smith. Graphics applied to medical image registration. IEEE Computer Graphics and Applications, 11(2):20–28, March/April 1991. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Monks:2000:AGB

[MOD00] Michael Monks, Byong Mok Oh, and Julie Dorsey. Audioptimization: Goal-based acoustic design. IEEE Computer Graphics and Applications, 20(3):76–91, May/ June 2000. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

computer.org/cga/cg2000/
g3076abs.htm;   http:/

                                        /dlib.computer.org/cg/
                                        books/cg2000/pdf/g3076.
                                        pdf.

Moher:1985:VCD

[Moh85] Thomas G. Moher. Videocassette course development using microcomputer graphics. IEEE Computer Graphics and Applications, 5(6): 34–40, June 1985. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Moller:2018:BSR

[Mol18a] Torsten Moller. A breezy summer read. IEEE Computer Graphics and Applications, 38(4):4, July/ August 2018. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        cg/2018/04/mcg2018040004.
                                        html.

Moller:2018:BBB

[Mol18b] Torsten Moller. Building bridges between research and applications. IEEE Computer Graphics and Applications, 38(1):4–7, January/February 2018. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL https:/

                                        /www.computer.org/csdl/
                                        mags/cg/2018/01/mcg2018010004.
                                        html.

Moller:2018:DEB

[Mol18c] Torsten Moller. A dynamic editorial board. IEEE Computer Graphics and Applications, 38(5):4–6, September/October 2018. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2018/05/mcg2018050004.
html.

Moller:2019:CC

[Möl19] Torsten Möller. Changes at CG&A. IEEE Computer Graphics and Applications, 39(1):5–8, January/February 2019. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL https:

//ieeexplore.ieee.org/
document/8663652/.

Moller:2020:O

[Möl20] T. Möller. Origins. IEEE Computer Graphics and Applications, 40(1):14–19, January/February 2020. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Moller:2022:IO

[Möl22]  Torsten Möller.  Incom-

ing and outgoing. IEEE Computer Graphics and Applications, 42(6):123–124, November/December 2022. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Mamrak:1989:AGT

[MOP89] Sandra A. Mamrak, Conleth S. O’Connell, and Richard E. Parent. The automatic generation of translation software for graphic objects. IEEE Computer Graphics and Applications, 9(6):34–42 (or 34–43??), November/December 1989. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Morie:1998:CTE

[Mor98] Jacquelyn Ford Morie. CGI training for the entertainment: Film industry. IEEE Computer Graphics and Applications, 18(1):30–37, January/February 1998. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

                                        computer.org/cga/cg1998/
                                        g1030abs.htm;   http:/
                                        /dlib.computer.org/cg/
                                        books/cg1998/pdf/g1030.
                                        pdf.

Morrison:2005:CBI

[Mor05] Gerald D. Morrison. A camera-based input device for large interactive displays. IEEE Computer Graphics and Applications, 25(4): 52–57, July/August 2005. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Moreland:2016:TSV

[Mor16] Kenneth Moreland. The tensions of in situ visualization. IEEE Computer Graphics and Applications, 36(2):5–9, March/April 2016. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2016/02/mcg2016020005-
abs.html.

Mirolo:1989:SMS

[MP89] Claudio Mirolo and Enrico Pagello. A solid modeling system for robot action planning. IEEE Computer Graphics and Applications, 9(1):55–69, January/ February 1989. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Ma:1999:GEI

[MP99] Kwan-Liu Ma and Frits H.

Post. Guest Editors’ introduction: Visualization case studies. IEEE Computer Graphics and Applications, 19(5):48–49, September/October 1999. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://dlib.
computer.org/cg/books/
cg1999/pdf/g5048.pdf.

Ma:2001:MPS

[MP01] Kwan-Liu Ma and Steven Parker. Massively parallel software rendering for visualizing large-scale data sets. IEEE Computer Graphics and Applications, 21(4):72– 83, July/August 2001. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

computer.org/cga/cg2001/
g4072abs.htm;   http:/
/dlib.computer.org/cg/
books/cg2001/pdf/g4072.
pdf.

Musse:2023:FPC

[MP23] Soraia Raupp Musse and Nuria Pelechano. Fighting pandemics with computer graphics and applications. IEEE Computer Graphics and Applications, 43(1):37– 38, January/February 2023. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Martin:2014:AMT

[MPDR14] Iain Martin, Steve Parkes, Martin Dunstan, and Nick Rowell. Asteroid modeling for testing spacecraft approach and landing. IEEE Computer Graphics and Applications, 34(4):52–62, July/August 2014. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Ma:1994:PVR

[MPHK94] Kwan-Liu Ma, James S. Painter, Charles D. Hansen, and Michael F. Krogh. Parallel volume rendering using binary-swap compositing. IEEE Computer Graphics and Applications, 14(4): 59–68, July/August 1994. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

MacDonald:1981:TDM

[MPP81] John B. MacDonald, Mary Kay Podlecki, and Milt J. Pappas. Technical documentation by magic. IEEE Computer Graphics and Applications, 1(2):27–30, 32, 34, 36, April 1981. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Melgare:2022:IES

[MQM22] Julia Kubiak Melgare, Rossana Baptista Queiroz, and Soraia Raupp Musse. Investigating emotion style in human faces using clustering methods. IEEE Computer Graphics and Applications, 42(4):52–71, July/ August 2022. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

McCormick:1999:VVV

[MQR99] Patrick S. McCormick, Ji Qiang, and Robert D. Ryne. Visualization viewpoints: Visualizing highresolution accelerator physics. IEEE Computer Graphics and Applications, 19(5):

11–13, September/October 1999. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL

                                        http://computer.org/cga/
                                        cg1999/g5011abs.htm;
                                        http://dlib.computer.org/
                                        cg/books/cg1999/pdf/g5011.
                                        pdf.

Montani:1987:VRH

[MR87] Claudio Montani and Michele Re. Vector and raster hidden-surface removal using parallel connected stripes. IEEE Computer Graphics and Applications, 7(7):14– 23, July/August 1987. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Macedonia:1998:PVD

[MR98] Michael R. Macedonia and Lawrence J. Rosenblum. Projects in VR: Designing animal habitats within an immersive VE. IEEE Computer Graphics and Applications, 18(5):9–13, September/October 1998. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL http://dlib.
                                        computer.org/cg/books/
                                        cg1998/pdf/g5009.pdf.

Matsuba:1999:BTA

[MR99] Stephen N. Matsuba and Bernie Roehl. “Bottom, Thou Art Translated”: The making of VRML Dream. IEEE Computer Graphics and Applications, 19(2):45– 51, March/April 1999. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

computer.org/cga/cg1999/
g2045abs.htm;   http:/
/dlib.computer.org/cg/
books/cg1999/pdf/g2045.
pdf.

Madanagopal:2019:APP

[MRB19] K. Madanagopal, E. D. Ragan, and P. Benjamin. Analytic provenance in practice: The role of provenance in real-world visualization and data analysis environments. IEEE Computer Graphics and Applications, 39(6):30–45, November/December 2019. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Muller:2012:LSV

[MRBW12] Christoph Muller, Guido Reina, Michael Burch, and Daniel Weiskopf. Largescale visualization projects for teaching software engineering. IEEE Computer Graphics and Applications, 32(4):14–19, July/ August 2012. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Mazalek:2006:TEA

[MRD06] Ali Mazalek, Matthew Reynolds, and Glorianna Davenport. TViews: An extensible architecture for multiuser digital media tables. IEEE Computer Graphics and Applications, 26(5):47–55, September/ October 2006. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

MacIntyre:2013:WGA

[MRL13] Blair MacIntyre, Hafez Rouzati, and Martin Lechner. Walled gardens: Apps and data as barriers to augmenting reality. IEEE Computer Graphics and Applications, 33(3):77–81, May/ June 2013. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Mantyla:1982:GSM

[MS82] Martti Mantyla and Reijo Sulonen. GWB: a solid modeler with Euler operators. IEEE Computer Graphics and Applications, 2(7):17–??, September 1982. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

McCormack:2024:JMA

[MS24] Jon McCormack and Francesca Samsel. Jon McCormack: Art infused with [artificial] intelligence. IEEE Computer Graphics and Applications, 44(2):46–54, March/ April 2024. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

McCarty:1994:VCD

[MSA^{+}94] W. Dean McCarty, Steven Sheasby, Philip Amburn, Martin R. Stytz, and Chip Switzer. A virtual cockpit for a distributed interactive simulation. IEEE Computer Graphics and Applications, 14(1):49–54, January/ February 1994. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Morris:2006:VSB

[MSB^{+}06] Dan Morris, Christopher Sewell, Federico Barbagli, Kenneth Salisbury, Nikolas H. Blevins, and Sabine Girod. Visuohaptic simulation of bone surgery for training and evaluation. IEEE Computer Graphics and Applications, 26(6): 48–57, November/December 2006. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Marchant:1989:SBE

[MSC89] Gary LaMont Marchant, Michael B. Stephenson, and Troy Crowfoot. A set of benchmarks for evaluating engineering workstations. IEEE Computer Graphics and Applications, 9(3):29–33, May/June 1989. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Marques:2023:HEI

[MSDS23] Bernardo Marques, Samuel Silva, Paulo Dias, and Beatriz Sousa Santos. How to evaluate if collaborative augmented reality speaks to its users. IEEE Computer Graphics and Applications, 43(5):107–113, September/ October 2023. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Mossman:2023:MAR

[MSED23] Christopher Mossman, Faramarz F. Samavati, Katayoon Etemad, and Peter Dawson. Mobile augmented reality for adding detailed multimedia content to historical physicalizations. IEEE Computer Graphics and Applications, 43(3):71–83, May/June 2023. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Marini:2007:SSP

[MSF07] Simone Marini, Michela Spagnuolo, and Bianca Falcidieno. Structural shape prototypes for the automatic classification of 3D objects. IEEE Computer Graphics and Applications, 27(4): 28–37, July/August 2007. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Mendez:2008:GSM

[MSH^{+}08] Erick Mendez, Gerhard Schall, Sven Havemann, Sebastian Junghanns, Dieter Fellner, and Dieter Schmalstieg. Generating semantic 3D models of underground infrastructure. IEEE Computer Graphics and Applications, 28(3):48–57, May/ June 2008. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Meier:2004:ICP

[MSK04] Barbara J. Meier, Anne Morgan Spalter, and David B. Karelitz. Interactive color palette tools. IEEE Computer Graphics and Applications, 24(3):64–72, May/ June 2004. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cg/2004/03/g3064abs.htm;
http://csdl.computer.
org/dl/mags/cg/2004/03/
g3064.htm;   http://csdl.
computer.org/dl/mags/cg/
2004/03/g3064.pdf.

Meyer-Spradow:2009:VRP

[MSRMH09] Jennis Meyer-Spradow, Timo Ropinski, Jörg Mensmann, and Klaus Hinrichs. Voreen: a rapid-prototyping environment for ray-casting-based volume visualizations. IEEE Computer Graphics and Applications, 29(6): 6–13, November/December 2009. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Moreland:2016:VMA

[MSU^{+}16] Kenneth Moreland, Christopher Sewell, William Usher, Li ta Lo, Jeremy Meredith, David Pugmire, James Kress, Hendrik Schroots, Kwan-Liu Ma, Hank Childs, Matthew Larsen, Chun-Ming Chen, Robert Maynard, and Berk Geveci. VTK-m: Accelerating the visualization toolkit for massively threaded architectures. IEEE Computer Graphics and Applications, 36(3):48–58, May/ June 2016. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2016/03/mcg2016030048-
abs.html.

Marcolin:2021:AVR

[MSU^{+}21] Federica Marcolin, Giulia Wally Scurati, Luca Ulrich, Francesca Nonis, Enrico Vezzetti, Nicolò Dozio, and Francesco Ferrise. Affective virtual reality: How to design artificial experiences impacting human emotions. IEEE Computer Graphics and Applications, 41(6):171–178, November/ December 2021. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

MacInnes:2010:VCE

[MSW10] Joseph MacInnes, Stephanie Santosa, and William Wright. Visual classification: Expert knowledge guides machine learning. IEEE Computer Graphics and Applications, 30(1):8–14, January/ February 2010. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Marner:2014:SUI

[MSWT14] Michael R. Marner, Ross T.

Smith, James A. Walsh, and Bruce H. Thomas. Spatial user interfaces for large-scale projector-based augmented reality. IEEE Computer Graphics and Applications, 34(6):74–82, November/ December 2014. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://csdl.
computer.org/csdl/mags/
cg/2014/06/mcg2014060074.
html.

Machover:1994:GEI

[MT94] Carl Machover and Steve E.

Tice. Guest Editors’ introduction: Virtual reality. IEEE Computer Graphics and Applications, 14(1):15– 16, January/February 1994. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Mohan:2007:ESS

[MTC07] Ankit Mohan, Jack Tumblin, and Prasun Choudhury. Editing soft shadows in a digital photograph. IEEE Computer Graphics and Applications, 27(2): 23–31, March/April 2007. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Menin:2018:AVT

[MTN18] Aline Menin, Rafael Torchelsen, and Luciana Nedel. An analysis of VR technology used in immersive simulations with a serious game perspective. IEEE Computer Graphics and Applications, 38(2):57–73, March/ April 2018. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        cg/2018/02/mcg2018020057-
                                        abs.html.

Mihalisin:1991:VMF

[MTS91] Ted Mihalisin, John Timlin, and John Schwegler. Visualizing multivariate functions, data, and distributions. IEEE Computer Graphics and Applications, 11(3):28– 35, May/June 1991. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Miyazaki:2020:AGT

[MTS^{+}20] T. Miyazaki, T. Tsuchiya, Y. Sugaya, S. Omachi, M. Iwamura, S. Uchida, and K. Kise. Automatic generation of typographic font from small font subset. IEEE Computer Graphics and Applications, 40(1): 99–111, January/February 2020. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Magnenat-Thalmann:1983:UHG

[MTT83] Nadia Magnenat-Thalmann and Daniel Thalmann. The use of high-level 3-D graphical types in the Mira animation system. IEEE Computer Graphics and Applications, 3(9):9–16, December 1983. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Magnenat-Thalmann:1985:GI

[MTT85a] Nadia Magnenat-Thalmann and Daniel Thalmann. Graphics Interface ’85. IEEE Computer Graphics and Applications, 5(10):22– 64, October 1985. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Magnenat-Thalmann:1985:IBC

[MTT85b] Nadia Magnenat-Thalmann and Daniel Thalmann. An indexed bibliography on computer animation. IEEE Computer Graphics and Applications, 5(7):76–86, July/ August 1985. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Magnenat-Thalmann:1985:TCA

[MTT85c] Nadia Magnenat-Thalmann and Daniel Thalmann. Three-dimensional computer animation: More an evolution than a motion problem. IEEE Computer Graphics and Applications, 5(10):47–57, October 1985. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Magnenat-Thalmann:1986:SCE

[MTT86] N. Magnenat-Thalmann and D. Thalmann. Special cinematographic effects with virtual movie cameras. IEEE Computer Graphics and Applications, 6(4):43–50, April 1986. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Magnenat-Thalmann:1987:IBI

[MTT87a] N. Magnenat-Thalmann and D. Thalmann. An indexed bibliography on image synthesis. IEEE Computer Graphics and Applications, 7(8):27–38, August 1987. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Magnenat-Thalmann:1987:DSA

[MTT87b] Nadia Magnenat-Thalmann and Daniel Thalmann. Direction of synthetic actors in the film Rendez-Vous à Montréal. IEEE Computer Graphics and Applications, 7(12):9–19, December 1987. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Magnenat-Thalmann:1991:CMA

[MTT91] Nadia Magnenat-Thalmann and Daniel Thalmann. Complex models for animating synthetic actors. IEEE Computer Graphics and Applications, 11(5):32–44, September 1991. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Magnenat-Thalmann:1985:MED

[MTTF85] Nadia Magnenat-Thalmann, Daniel Thalmann, and Mario Fortin. MIRANIM: An extensible directororiented system for the animation of realistic images. IEEE Computer Graphics and Applications, 5(3):61– 73, March/April 1985. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Mueller:2018:RIT

[Mue18]  Stefanie Mueller.  Re-

thinking interaction techniques for personal fabrication. IEEE Computer Graphics and Applications, 38(5):18–25, September/ October 2018. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2018/05/mcg2018050018-
abs.html.

Munzner:1998:ELG

[Mun98] Tamara Munzner. Exploring large graphs in 3D hyperbolic space. IEEE Computer Graphics and Applications, 18(4):18–23, July/ August 1998. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

computer.org/cga/cg1998/
g4018abs.htm.

Munzner:2002:GEI

[Mun02] Tamara Munzner. Guest Editor’s introduction: Information visualization. IEEE Computer Graphics and Applications, 22(1):20–21, January/February 2002. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http:

//dlib.computer.org/cg/
books/cg2002/g1020abs.
htm;   http://dlib.computer.
org/cg/books/cg2002/pdf/
g1020.pdf.

Marrinan:2010:UIM

[MUN^{+}10] Thomas Marrinan, Timothy Urness, Charles Nelson, Kory Kreimeyer, and Jordan Mirocha. Understanding and interpreting multivalued astronomical data. IEEE Computer Graphics and Applications, 30(5): 12–17, September/October 2010. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Murch:1984:PPE

[Mur84a] Gerald M. Murch. Physiological principles for the effective use of color. IEEE Computer Graphics and Applications, 4(11):49–54, November/December 1984. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Murray:1984:VDT

[Mur84b] William E. Murray. Video display terminals: Radiation issues. IEEE Computer Graphics and Applications, 4(4):41–44, April 1984. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Muraki:1993:VDW

[Mur93] Shigeru Muraki. Volume data and wavelet transforms. IEEE Computer Graphics and Applications, 13(4):50–56, July/August 1993. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Musgrave:1990:NEM

[Mus90a] F. Kenton Musgrave. Natura ex machina II. IEEE Computer Graphics and Applications, 10(6):5–7, November/ December 1990. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Musgrave:1990:NRT

[Mus90b] F. Kenton Musgrave. A note on ray tracing mirages. IEEE Computer Graphics and Applications, 10(6): 10–12, November/December 1990. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Musgrave:1999:ACT

[Mus99] F. Kenton Musgrave. About the cover: Towards a synthetic universe. IEEE Computer Graphics and Applications, 19(6):4–5, November/ December 1999. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL http://dlib.
                                        computer.org/cg/books/
                                        cg1999/extras/g6004x1.
                                        mpg;   http://dlib.computer.
                                        org/cg/books/cg1999/pdf/
                                        g6004.pdf.

Moran:1993:VSS

[MV93] Christopher J. Moran and Guy Vezina. Visualizing soil surfaces and crop residues. IEEE Computer Graphics and Applications, 13(2):

40–47, March/April 1993. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Maupu:2005:SIM

[MVWB05] Damien Maupu, Mark H.

Van Horn, Susan Weeks, and Elizabeth Bullitt. 3D stereo interactive medical visualization. IEEE Computer Graphics and Applications, 25(5):67–71, September/October 2005. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Mirante:1982:RSA

[MW82] Anthony Mirante and Nicholas Weingarten. The radial sweep algorithm for constructing triangulated irregular networks. IEEE Computer Graphics and Applications, 2(3):11–13, 15–21, May/June 1982. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Mastin:1987:FSO

[MWM87] Gary A. Mastin, Peter A.

Watterberg, and John F. Mareda. Fourier synthesis of ocean scenes. IEEE Computer Graphics and Applications, 7(3):16–23, March/ April 1987. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Meng:2015:VFG

[MWY^{+}15] Zhang Meng, Si Weixin, Qian Yinling, Sun Hanqiu, Qin Jing, and Pheng-Ann Heng. Vortex filaments in grids for scalable, fine smoke simulation. IEEE Computer Graphics and Applications, 35(6):60–68, November/ December 2015. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL http://www.
                                        computer.org/csdl/mags/
                                        cg/2015/06/mcg2015060060-
                                        abs.html.

Myers:1982:IPS

[Mye82] Ware Myers. An industrial perspective on solid modeling. IEEE Computer Graphics and Applications, 2(2): 86–97, March/April 1982. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Myers:1984:UIS

[Mye84a] Brad A. Myers. The user interface for SAPPHIRE. IEEE Computer Graphics and Applications, 4(12):13– 23, December 1984. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Myers:1984:EVD

[Mye84b] Ware Myers. Ergonomics of video display terminals. IEEE Computer Graphics and Applications, 4(1):7– 15, January/February 1984.

CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Myers:1984:SGD

[Mye84c] Ware Myers. Staking out the graphics display pipeline. IEEE Computer Graphics and Applications, 4(7):60– 65, July/August 1984. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Myers:1985:CGN

[Mye85] Ware Myers. Computer graphics: the next 20 years. IEEE Computer Graphics and Applications, 5(8):69– 76, August 1985. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Myers:1987:CIT

[Mye87] Brad A. Myers. Creating interaction techniques by demonstration. IEEE Computer Graphics and Applications, 7(9):51–60, September 1987. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Myers:1988:TWM

[Mye88] Brad A. Myers. A taxonomy of window manager user interfaces. IEEE Computer Graphics and Applications, 8(5):65–84, September/October 1988. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Myklebust:1988:GEI

[Myk88] Arvid Myklebust. Guest Editor’s introduction: Mechanical computer-aided engineering. IEEE Computer Graphics and Applications, 8(2):24–25, March/ April 1988. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
cg/02/mcg1988020024.pdf.

Mortensen:2008:RTG

[MYK^{+}08] Jesper Mortensen, Insu Yu, Pankaj Khanna, Franco Tecchia, Bernhard Spanlang, Giuseppe Marino, and Mel Slater. Real-time global illumination for VR applications. IEEE Computer Graphics and Applications, 28(6):56–64, November/December 2008. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Mynatt:2008:UHP

[Myn08]  Elizabeth Mynatt.  Un-

locking human potential through technical innovation. IEEE Computer Graphics and Applications, 28(2):104, 100–103, March/ April 2008. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://bell.

computer.org/dlcomments/

.

Maim:2009:YAR

[MYPT09] Jonathan Maïm, Barbara Yersin, Julien Pettré, and Daniel Thalmann. YaQ: An architecture for real-time navigation and rendering of varied crowds. IEEE Computer Graphics and Applications, 29(4):44–53, July/ August 2009. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Ma:2009:UCI

[MYT09] Jonathan Maïm, Barbara Yersin, and Daniel Thalmann. Unique character instances for crowds. IEEE Computer Graphics and Applications, 29(6):82–90, November/December 2009. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Moller:2022:PVI

[MZ22] Torsten Möller and Richard Zhang. Perception, visual inference, and exploratory visualization. IEEE Computer Graphics and Applications, 42(1):17–19, January/ February 2022. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Ma:2021:CSD

[MZC^{+}21] C. Ma, Y. Zhao, A. Curtis, F. Kamw, S. AL-Dohuki, J. Yang, S. Jamonnak, and I. Ali. CLEVis: A semantic driven visual analytics system for community level events. IEEE Computer Graphics and Applications, 41(2):49–62, March/ April 2021. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Miao:2021:EES

[MZL^{+}21] Yongwei Miao, Lei Zhang, Jiazong Liu, Jinrong Wang, and Fuchang Liu. An end-to-end shape-preserving point completion network. IEEE Computer Graphics and Applications, 41(3):20– 33, May/June 2021. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Macedonia:1995:ERM

[MZP^{+}95] Michael R. Macedonia, Michael J. Zyda, David R. Pratt, Donald P. Brutzman, and Paul T. Barham. Exploiting reality with multicast groups. IEEE Computer Graphics and Applications, 15(5):38–45, September 1995. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Nowak:2023:IVO

[NAB^{+}23] Stan Nowak, Bon Adriel Aseniero, Lyn Bartram, Tovi Grossman, George Fitzmaurice, and Justin Matejka. Identifying visualization opportunities to help architects manage the complexity of building codes. IEEE Computer Graphics and Applications, 43(6): 75–86, November/December 2023. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Nadeau:1999:TBV

[Nad99] David R. Nadeau. Tutorial: Building virtual worlds with VRML. IEEE Computer Graphics and Applications, 19(2):18–29, March/ April 1999. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

computer.org/cga/cg1999/
g2018abs.htm;   http:/
/dlib.computer.org/cg/
books/cg1999/pdf/g2018.
pdf.

Nakamura:1984:TDD

[Nak84] Shohei Nakamura. Threedimensional digital display of ultrasonograms. IEEE Computer Graphics and Applications, 4(5):36–45, May/ June 1984. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Naps:2005:JSA

[Nap05] Thomas L. Naps. JHAVÉ:

Supporting algorithm visualization. IEEE Computer Graphics and Applications, 25(5):49–55, September/October 2005. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Navab:2004:DKA

[Nav04]  Nassir Navab.  Devel-

oping killer apps for industrial augmented reality. IEEE Computer Graphics and Applications, 24(3):16– 20, May/June 2004. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http:

                                        //csdl.computer.org/dl/
                                        mags/cg/2004/03/g3016.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/cg/2004/03/
                                        g3016.pdf.

Ning:1993:EIS

[NB93] Paul Ning and Jules Bloomenthal. An evaluation of implicit surface tilers. IEEE Computer Graphics and Applications, 13(6):33–41, November/December 1993. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Ni:1994:PEB

[NB94] Xiujun Ni and M. Susan Bloor. Performance evaluation of boundary data structures. IEEE Computer Graphics and Applications, 14(6):66–77, November/December 1994. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Nesbitt:2004:FTP

[NB04] Keith V. Nesbitt and Stephen Barrass. Finding trading patterns in stock market data. IEEE Computer Graphics and Applications, 24(5):45–55, September/October 2004. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http:

//csdl.computer.org/dl/
mags/cg/2004/05/g5045.
htm;   http://csdl.computer.
org/dl/mags/cg/2004/05/
g5045.pdf.

Nielson:1994:RID

[NBG^{+}94] G. M. Nielson, P. Brunet, M. Gross, H. Hagen, and S. V. Klimenko. Research issues in data modeling for scientific visualization. IEEE Computer Graphics and Applications, 14(2): 70–73, March/April 1994. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Noimark:2003:SSM

[NCO03] Yuval Noimark and Daniel Cohen-Or. Streaming scenes to MPEG-4 video-enabled devices. IEEE Computer Graphics and Applications, 23(1):58–64, January/ February 2003. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL http://dlib.
                                        computer.org/cg/books/
                                        cg2003/pdf/g1058.pdf;
                                        http://www.computer.org/
                                        cga/cg2003/g1058abs.htm.

Ni:2011:VRS

[NCQ^{+}11] Dong Ni, Wing Yin Chan, Jing Qin, Yim Pan Chui, Yingge Qu, Simon S. M. Ho, and Pheng Ann Heng. A virtual reality simulator for ultrasound-guided biopsy training. IEEE Computer Graphics and Applications, 31(2):36–48, March/ April 2011. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Nienhaus:2005:DDU

[ND05] Marc Nienhaus and Jurgen Dollner. Depicting dynamics using principles of visual art and narrations. IEEE Computer Graphics and Applications, 25(3):40–51, May/ June 2005. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Nedel:2016:UIV

[NdSM^{+}16] Luciana Nedel, Vinicius Costa de Souza, Aline Menin, Lucia Sebben, Jackson Oliveira, Frederico Faria, and Anderson Maciel. Using immersive virtual reality to reduce work accidents in developing countries. IEEE Computer Graphics and Applications, 36(2):36–46, March/April 2016. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2016/02/mcg2016020036-
abs.html.

Nelson:1993:VUD

[NE93] Thomas R. Nelson and T. Todd Elvins. Visualization of 3D ultrasound data. IEEE Computer Graphics and Applications, 13(6): 50–57, November/December 1993. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Neal:1986:TCG

[Nea86] Margaret Neal. Tour of computer graphics in Japan. IEEE Computer Graphics and Applications, 6(6):11– 20, June 1986. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Neal:1988:TBS

[Nea88] Margaret Neal. A tool for building sculptures. IEEE Computer Graphics and Applications, 8(2):4–7, March/ April 1988. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
cg/02/mcg1988020004-abs.
html.

Neal:1989:CMA

[Nea89a] Margaret Neal. Computer museum acquires the interactive image. IEEE Computer Graphics and Applications, 9(2):12–18, 81, March/April 1989. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Neal:1989:IAF

[Nea89b] Margaret Neal. An interview with Akira Fujimoto. IEEE Computer Graphics and Applications, 9(3):4–10, May/June 1989. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Neal:1989:KIS

[Nea89c] Margaret Neal. Keep it simple. IEEE Computer Graphics and Applications, 9(5):3– 5, September 1989. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Neumann:1994:CCP

[Neu94] Ulrich Neumann. Communication costs for parallel volume-rendering algorithms. IEEE Computer Graphics and Applications, 14(4):49–58, July/ August 1994. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Ney:1991:ETM

[NF91] Derek R. Ney and Elliot K.

Fishman. Editing tools for 3D medical imaging. IEEE Computer Graphics and Applications, 11(6):63–71, November/December 1991. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Nataneli:2011:BSR

[NF11] Gabriele Nataneli and Petros Faloutsos. Bringing sketch recognition into your hands. IEEE Computer Graphics and Applications, 31(3):32–41, May/ June 2011. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Nielson:1991:VMS

[NFHL91] Gregory M. Nielson, Thomas A.

Foley, Bernd Hamann, and David Lane. Visualizing and modeling scattered multivariate data. IEEE Computer Graphics and Applications, 11(3):47–55, May/ June 1991. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Ney:1990:VRC

[NFMD90] Derek R. Ney, Elliot K.

Fishman, Donna Magid, and Robert A. Drebin. Volumetric rendering of computed tomography data: Principles and techniques. IEEE Computer Graphics and Applications, 10(2):24–32, March/April 1990. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Ng:1996:CGT

[NG96] Hing N. Ng and Richard L. Grimsdale. Computer graphics techniques for modeling cloth. IEEE Computer Graphics and Applications, 16(5):28–41, September 1996. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Nickel:1984:IW

[Nic84] Randy Nickel. The IRIS workstation. IEEE Computer Graphics and Applications, 4(8):30–34, August 1984. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Nielson:1986:RNI

[Nie86] G. Nielson. A rectangular nu-spline for interactive surface design. IEEE Computer Graphics and Applications, 6(2):35–41, February 1986. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Nielson:1992:GEI

[Nie92] Gregory M. Nielson. Guest Editors’ introduction: Computeraided geometric design.

IEEE Computer Graphics and Applications, 12(5):29– ??, September 1992. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Nielson:1993:GEI

[Nie93a] Gregory M. Nielson. Guest Editor’s introduction CAGD’s top ten: What to watch. IEEE Computer Graphics and Applications, 13(1):35– 37, January/February 1993. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Nielson:1993:SDM

[Nie93b] Gregory M. Nielson. Scattered data modeling. IEEE Computer Graphics and Applications, 13(1):60–70, January/February 1993. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Nielsen:2007:DCP

[Nie07] Frank Nielsen. The digital chameleon principle: Computing invisibility by rendering transparency. IEEE Computer Graphics and Applications, 27(1):90–96, January/February 2007. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Nakamae:1989:CIA

[NINT89] Eihachiro Nakamae, Takao Ishizaki, Tomoyuki Nishita, and Shinichi Takita. Compositing 3D images with antialiasing and various shading effects. IEEE Computer Graphics and Applications, 9(2):21–29, March/ April 1989. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Noma:1985:AEA

[NK85] Tsukasa Noma and Tosiyasu L. Kunii. Animengine: An engineering animation system. IEEE Computer Graphics and Applications, 5(10):24– 33, October 1985. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Nielson:1994:VG

[NK94] Gregory M. Nielson and Arie E. Kaufman. Visualization graduates. IEEE Computer Graphics and Applications, 14(5):17–18, September 1994. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Nam:2014:MPS

[NK14] Giljoo Nam and Min H. Kim. Multispectral photometric stereo for acquiring high-fidelity surface normals. IEEE Computer Graphics and Applications, 34(6):57–68, November/ December 2014. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://csdl.
computer.org/csdl/mags/
cg/2014/06/mcg2014060057-
abs.html.

Nelson:1984:AAA

[NL84] D. L. Nelson and P. J.

Leach. The architecture and applications of the Apollo Domain. IEEE Computer Graphics and Applications, 4(4):58–66, April 1984. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Nyland:2001:VRW

[NL01] Lars Nyland and Anselmo Lastra. Visualizing the real world. IEEE Computer Graphics and Applications, 21(5):6–10, September/October 2001. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://dlib.
computer.org/cg/books/
cg2001/g5006abs.htm;
http://dlib.computer.org/
cg/books/cg2001/pdf/g5006.
pdf.

Nixon:2002:FBS

[NL02] Daniel Nixon and Richard Lobb. A fluid-based softobject model. IEEE Computer Graphics and Applications, 22(4):68–75, July/ August 2002. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

                                        computer.org/cga/cg2002/
                                        g4068abs.htm;   http:/
                                        /dlib.computer.org/cg/
                                        books/cg2002/pdf/g4068.
                                        pdf.

Nieuwenhuizen:2009:IDG

[NLvLM09] Karin Nieuwenhuizen, Lei Liu, Robert van Liere, and Jean-Bernard Martens. Insights from dividing 3D goaldirected movements into meaningful phases. IEEE Computer Graphics and Applications, 29(6):44–53, November/December 2009. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Neophytou:2008:CSC

[NM08] Neophytos Neophytou and Klaus Mueller. Color-space CAD: Direct gamut editing in 3D. IEEE Computer Graphics and Applications, 28(3):88–98, May/ June 2008. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Nelson:2006:DOB

[NMRI06] Mark J. Nelson, Michael Mateas, David L. Roberts, and Charles L. Isbell, Jr. Declarative optimizationbased drama management in interactive fiction. IEEE Computer Graphics and Applications, 26(3):32–41, May/June 2006. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Nakamoto:2008:CIO

[NOK^{+}08] Takamichi Nakamoto, Shigeki Otaguro, Masashi Kinoshita, Masahiko Nagahama, Keita Ohinishi, and Taro Ishida. Cooking up an interactive olfactory game display. IEEE Computer Graphics and Applications, 28(1):75– 78, January/February 2008. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

North:2006:TMV

[Nor06] Chris North. Toward measuring visualization insight. IEEE Computer Graphics and Applications, 26(3):6–9, May/June 2006. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Neff:2017:ANV

[NP17] Michael Neff and Catherine Pelachaud. Animation of natural virtual characters. IEEE Computer Graphics and Applications, 37(4):14–16, July/ August 2017. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2017/04/mcg2017040014.
html.

Nilsson:2018:YRR

[NPB^{+}18] Niels Christian Nilsson, Tabitha Peck, Gerd Bruder, Eri Hodgson, Stefania Serafin, Mary Whitton, Frank Steinicke, and Evan Suma Rosenberg. 15 years of research on redirected walking in immersive virtual environments. IEEE Computer Graphics and Applications, 38(2):44–56, March/ April 2018. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2018/02/mcg2018020044-
abs.html.

Nakasone:2009:ACV

[NPH^{+}09] Arturo Nakasone, Helmut Prendinger, Simon Holland, Piet Hut, Jun Makino, and Ken Miura. AstroSim: Collaborative visualization of an astrophysics simulation in second life. IEEE Computer Graphics and Applications, 29(5):69–81, September/October 2009. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Norton:2003:EVD

[NR03] Alan Norton and Alyn Rockwood. Enabling viewdependent progressive volume visualization on the Grid. IEEE Computer Graphics and Applications, 23(2):22–31, March/April 2003. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL

http://dlib.computer.org/
cg/books/cg2003/pdf/g2022.
pdf;   http://www.computer.
org/cga/cg2003/g2022abs.
htm.

Nourani:2022:DVD

[NRH^{+}22] Mahsan Nourani, Chiradeep Roy, Donald R. Honeycutt, Eric D. Ragan, and Vibhav Gogate. DETOXER: a visual debugging tool with multiscope explanations for temporal multilabel classification. IEEE Computer Graphics and Applications, 42(6):37–46, November/December 2022. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Nuernberger:2020:UWO

[NTB^{+}20] B. Nuernberger, R. Tapella, S. Berndt, S. Y. Kim, and S. Samochina. Under water to outer space: Augmented reality for astronauts and beyond. IEEE Computer Graphics and Applications, 40(1):82–89, January/February 2020. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Naemura:2002:RTV

[NTH02] Takeshi Naemura, Junji Tago, and Hiroshi Harashima. Real-time videobased modeling and rendering of 3D scenes. IEEE Computer Graphics and Applications, 22(2):66–73, March/April 2002. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

computer.org/cga/cg2002/
g2066abs.htm;   http:/
/dlib.computer.org/cg/
books/cg2002/pdf/g2066.
pdf.

Nara:2001:SAW

[NTM^{+}01] Takaaki Nara, Masaya Takasaki, Taro Maeda, Toshiro Higuchi, Shigeru Ando, and Susumu Tachi. Surface acoustic wave tactile display. IEEE Computer Graphics and Applications, 21(6):56–63, November/ December 2001. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://dlib.
computer.org/cg/books/
cg2001/g6056abs.htm;
http://dlib.computer.org/
cg/books/cg2001/pdf/g6056.
pdf.

Navab:2007:AWD

[NTS^{+}07] Nassir Navab, Joerg Traub, Tobias Sielhorst, Marco Feuerstein, and Christoph Bichlmeier. Action- and workflow-driven augmented reality for computer-aided medical procedures. IEEE Computer Graphics and Applications, 27(5):10–14, September/October 2007. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Nurminen:2008:MCM

[Nur08] Antti Nurminen. Mobile 3D city maps. IEEE Computer Graphics and Applications, 28(4):20–31, July/ August 2008. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Nielson:1992:IAK

[NVC92] Gregory M. Nielson, Keith Voegele, and Brian Collins. Introduction to “An Annotated and Keyworded Bibliography for Scientific Visualization”. IEEE Computer Graphics and Applications, 12(4):23–??, July/ August 1992. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Noma:2000:DVH

[NZB00] Tsukasa Noma, Liwei Zhao, and Norman I. Badler. Design of a virtual human presenter. IEEE Computer Graphics and Applications, 20(4):79–85, July/ August 2000. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

computer.org/cga/cg2000/
g4079abs.htm;   http:/
/dlib.computer.org/cg/

                                        books/cg2000/pdf/g4079.
                                        pdf.

Nilsson:2021:PVR

[NZS21] Niels Christian Nilsson, André Zenner, and Adalberto L. Simeone. Propping up virtual reality with haptic proxies. IEEE Computer Graphics and Applications, 41(5):104–112, September/ October 2021. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Ota:1985:AFP

[OATO85] Yoshimi Ota, Hiroshi Arai, Shinji Tokumasu, and Toshio Ochi. An automated finite polygon division method for 3-D objects. IEEE Computer Graphics and Applications, 5(4):60–70, April 1985. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Olsen:1984:CUI

[OBE^{+}84] Dan R. Olsen, Jr., William Buxton, Roger Ehrich, David J. Kasik, James R. Rhyne, and John Sibert. A context for user interface management. IEEE Computer Graphics and Applications, 4(12):33–42, December 1984. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Oliva:2023:MNI

[OBG^{+}23] Ramon Oliva, Alejandro Beacco, Jaime Gallego, Raul Gallego Abellan, and Mel Slater. The making of a newspaper interview in virtual reality: Realistic avatars, philosophy, and sushi. IEEE Computer Graphics and Applications, 43(6):117–125, November/ December 2023. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Olaizola:2023:VAP

[OBO^{+}23] Igor García Olaizola, Jan Lukas Bruse, Juan Odriozola, Arkaitz Artetxe, David Velásquez, Marco Quartulli, and Jorge Posada. Visual analytics platform for centralized COVID-19 digital contact tracing. IEEE Computer Graphics and Applications, 43(1):53–64, January/ February 2023. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Ouyang:2014:BMPb

[OC14] Peichang Ouyang and Kwokwai Chung. Beautiful math, part 3: Hyperbolic aesthetic patterns based on conformal mappings. IEEE Computer Graphics and Applications, 34(2):72–79, March/ April 2014. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Offenhuber:2020:DOC

[OCS20] D. Offenhuber, B. Campbell, and F. Samsel. Dietmar Offenhuber: Collaboration via the many traces our data sets leave behind. IEEE Computer Graphics and Applications, 40(5):120–126, September/October 2020. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Olsen:1988:MEG

[OD88] Dan R. Olsen, Jr. and John R. Dance. Macros by example in a graphical UIMS. IEEE Computer Graphics and Applications, 8(1):68–78, January/ February 1988. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Owen:2013:HVC

[ODE^{+}13] G. Scott Owen, Gitta Domik, David S. Ebert, Jorn Kohlhammer, Holly Rushmeier, Beatriz Sousa Santos, and Daniel Weiskopf. How visualization courses have changed over the past 10 years. IEEE Computer Graphics and Applications, 33(4):14–19, July/ August 2013. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Ouyang:2015:BMPa

[ODW15] Peichang Ouyang, Feng Ding, and Xinchang Wang. Beautiful math, part 4: Polygonal aesthetic patterns based on the Schwarz– Christoffel mapping. IEEE Computer Graphics and Applications, 35(4):22–25, July/August 2015. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
cg/2015/04/mcg2015040022-
abs.html.

Ouyang:2014:BMPa

[OF14] Peichang Ouyang and Robert W.

Fathauer. Beautiful math, part 2: Aesthetic patterns based on fractal tilings. IEEE Computer Graphics and Applications, 34(1):68– 76, January/February 2014. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Offenhuber:2020:WWT

[Off20]  D. Offenhuber.  What

we talk about when we talk about data physicality. IEEE Computer Graphics and Applications, 40(6): 25–37, November/December 2020. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Olson:1991:FVP

[OG91] Arthur J. Olson and David S.

Goodsell. A functional view of proteins. IEEE Computer Graphics and Applications, 11(1):15–17, January/ February 1991. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Oian:1982:TSC

[Oia82] Jorn Oian. Trends in Scandinavian CAD development. IEEE Computer Graphics and Applications, 2(3):51– 56, 58, July/August 1982. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Obermaier:2014:FCE

[OJ14] Harald Obermaier and Kenneth I. Joy. Future challenges for ensemble visualization. IEEE Computer Graphics and Applications, 34(3):8–11, May/June 2014. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Otori:2009:TSM

[OK09] Hirofumi Otori and Shigeru Kuriyama. Texture synthesis for mobile data communications. IEEE Computer Graphics and Applications, 29(6):74–81, November/December 2009. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Otto:2012:IVC

[OKET12] Mathias Otto, Alexander Kuhn, Wito Engelke, and Holger Theisel. 2011 IEEE Visualization Contest winner: Visualizing unsteady vortical behavior of a centrifugal pump. IEEE Computer Graphics and Applications, 32(5):12–19, September/October 2012. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Ou:2013:EPR

[OKKP13] Jiawei Ou, Ondrej Karlik, Jaroslav Krivanek, and Fabio Pellacini. Evaluating progressive-rendering algorithms in appearance design tasks. IEEE Computer Graphics and Applications, 33(6):58–68, November/December 2013. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Okino:1984:EDA

[OKM84] Norio Okino, Yukinori Kakazu, and Masamichi Morimoto. Extended depthbuffer algorithms for hiddensurface visualization. IEEE Computer Graphics and Applications, 4(5):79–84, 86– 88, May/June 1984. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Onoue:2017:LGD

[OKS^{+}17] Yosuke Onoue, Nobuyuki Kukimoto, Naohisa Sakamoto, Kazuo Misue, and Koji Koyamada. Layered graph drawing for visualizing evaluation structures. IEEE Computer Graphics and Applications, 37(2):20–30, March/April 2017. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        cg/2017/02/mcg2017020020-
                                        abs.html.

Olsen:1986:ETU

[Ols86] Dan R. Olsen, Jr. Editing templates: a user interface generation tool. IEEE Computer Graphics and Applications, 6(11):40–45, November/December 1986. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Okabe:2020:IVC

[ONDA20] M. Okabe, K. Noda, Y. Dobashi, and K. Anjyo. Interactive video completion. IEEE Computer Graphics and Applications, 40(1): 127–139, January/February 2020. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Ochiai:2014:CMT

[OOHR14] Yoichi Ochiai, Alexis Oyama, Takayuki Hoshi, and Jun Rekimoto. The colloidal metamorphosis: Time division multiplexing of the reflectance state. IEEE Computer Graphics and Applications, 34(4):42–51, July/ August 2014. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

OSullivan:2011:CPF

[O’S11] Carol O’Sullivan. CG&A — past and future themes. IEEE Computer Graphics and Applications, 31(4):5–6, July/August 2011. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Olsen:2011:NMI

[OSJ11] Luke Olsen, Faramarz Samavati, and Joaquim Jorge. NaturaSketch: Modeling from images and natural sketches. IEEE Computer Graphics and Applications, 31(6):24–34, November/December 2011. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Oka:2002:RTF

[OSK02] Kenji Oka, Yoichi Sato, and Hideki Koike. Real-time fingertip tracking and gesture recognition. IEEE Computer Graphics and Applications, 22(6):64–71, November/December 2002. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http:

                                        //dlib.computer.org/cg/
                                        books/cg2002/pdf/g6064.
                                        pdf;   http://www.computer.
                                        org/cga/cg2002/g6064abs.
                                        htm.

Ofek:1997:MTI

[OSRW97] Eyal Ofek, Erez Shilat, Ari Rappoport, and Michael Werman. Multiresolution textures from image sequences. IEEE Computer Graphics and Applications, 17(2):18–29, March/April 1997. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Oviatt:2003:ARM

[Ovi03] Sharon Oviatt. Advances in robust multimodal interface design. IEEE Computer Graphics and Applications, 23(5):62–68, September/October 2003. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http:

                                        //csdl.computer.org/dl/
                                        mags/cg/2003/05/g5062.
                                        pdf;   http://csdl.computer.
                                        org/dl/mags/cg/2003/05/
                                        g5062abs.htm.

Ouyang:2017:BMP

[OWZ17] Peichang Ouyang, Xinchang Wang, and Yongman Zhao. Beautiful math, part 6: Visualizing 4D regular polytopes using the kaleidoscope principle. IEEE Computer Graphics and Applications, 37(5):92–97, September/October 2017. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2017/05/mcg2017050092-
abs.html.

OBrien:2000:CAP

[OZH00] James F. O’Brien, Victor B.

Zordan, and Jessica K. Hodgins. Combining active and passive simulations for secondary motion. IEEE Computer Graphics and Applications, 20(4):86–96, July/ August 2000. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

computer.org/cga/cg2000/
g4086abs.htm;   http:/
/dlib.computer.org/cg/
books/cg2000/pdf/g4086.
pdf.

Ouyang:2015:BMPb

[OZH15] Peichang Ouyang, Weiguo Zhao, and Xuan Huang. Beautiful math, part 5: Colorful Archimedean tilings from dynamical systems. IEEE Computer Graphics and Applications, 35(6): 90–96, November/December 2015. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL

http://www.computer.org/
csdl/mags/cg/2015/06/mcg2015060090-
abs.html.

Oztireli:2015:MSG

[Ozt15] Cengiz Oztireli. Making sense of geometric data. IEEE Computer Graphics and Applications, 35(4):100– 106, July/August 2015. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http:/

                                        /www.computer.org/csdl/
                                        mags/cg/2015/04/mcg2015040100-
                                        abs.html.

Patchett:2018:OST

[PA18] John Patchett and James Ahrens. Optimizing scientist time through in situ visualization and analysis. IEEE Computer Graphics and Applications, 38(1):119–127, January/ February 2018. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        cg/2018/01/mcg2018010119-
                                        abs.html.

Phillips:1986:BFR

[PABS86] Richard L. Phillips, Daniel E. Atkins, Nancy A. Benovich, and Brian D. Schipper. A bridge from full-function to reduced-function workstations. IEEE Computer Graphics and Applications, 6(5):53–57, May/June 1986. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Pachet:2004:BCJ

[Pac04] François Pachet. Beyond the cybernetic jam fantasy: The continuator. IEEE Computer Graphics and Applications, 24(1):31–35, January/ February 2004. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cg/2004/01/g1031abs.htm;
 http://csdl.computer.
org/dl/mags/cg/2004/01/
g1031.pdf.

Pack:2010:VTC

[Pac10] Michael L. Pack. Visualization in transportation: Challenges and opportunities for everyone. IEEE Computer Graphics and Applications, 30(4):90–96, July/August 2010. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Paller:1981:IMP

[Pal81] Alan T. Paller. Improving management productivity with computer graphics. IEEE Computer Graphics and Applications, 1(4):9–14, 16, October 1981. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Palyka:1985:CAD

[Pal85] Duane M. Palyka. Computer/art — depolarization and unification. IEEE Computer Graphics and Applications, 5(7):46–56, July/ August 1985. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Palmer:1992:LMV

[Pal92] Thomas C. Palmer. A language for molecular visualization. IEEE Computer Graphics and Applications, 12(3):23–32, May/ June 1992. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Pang:1990:LAP

[Pan90]  Alex T. Pang.  Line-

drawing algorithms for parallel machines. IEEE Computer Graphics and Applications, 10(5):54–59, September 1990. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Pang:1994:SR

[Pan94] Alex Pang. Spray rendering. IEEE Computer Graphics and Applications, 14 (5):57–63, September 1994. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Panozzo:2015:DQR

[Pan15]  Daniele Panozzo.  De-

mystifying quadrilateral remeshing. IEEE Computer Graphics and Applications, 35(2):88–95, March/ April 2015. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
cg/2015/02/mcg2015020088-
abs.html.

Pape:1996:HVR

[Pap96] Dave Pape. A hardwareindependent virtual reality development system. IEEE Computer Graphics and Applications, 16(4):44–47, July/August 1996. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Paquette:1996:ASA

[Paq96] S. Paquette. Applications:

3D scanning in apparel design and human engineering. IEEE Computer Graphics and Applications, 16 (5):11–15, September 1996. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Parke:1982:PMF

[Par82] Frederic I. Parke. Parameterized models for facial animation. IEEE Computer Graphics and Applications, 2(9):61–64, 66–68, November/December 1982. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Paradiso:1998:GEI

[Par98] Joseph A. Paradiso. Guest Editor’s introduction: Getting the picture. IEEE Computer Graphics and Applications, 18(3):26–27, May/ June 1998. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL http://dlib.
                                        computer.org/cg/books/
                                        cg1998/pdf/g3026.pdf.

Pulli:2005:DGP

[PARV05] Kari Pulli, Tomi Aarnio, Kimmo Roimela, and Jani Vaarala. Designing graphics programming interfaces for mobile devices. IEEE Computer Graphics and Applications, 25(6):66–75, November/December 2005. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Patrikalakis:1993:SI

[Pat93] Nicholas M. Patrikalakis. Surface-to-surface intersections. IEEE Computer Graphics and Applications, 13(1):89–95, January/ February 1993. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Patow:2012:UFG

[Pat12]  Gustavo Patow.  User-

friendly graph editing for procedural modeling of buildings. IEEE Computer Graphics and Applications, 32(2):66–75, March/ April 2012. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Pavlidis:1985:SCR

[Pav85] Theo Pavlidis. Scan conversion of regions bounded by parabolic splines. IEEE Computer Graphics and Applications, 5(6):47–53, June 1985. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Plunkett:1985:VRA

[PB85] David J. Plunkett and Michael J. Bailey. The vectorization of a ray-tracing algorithm for improved execution speed. IEEE Computer Graphics and Applications, 5(8):52–60, August 1985. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Pueyo:1987:PSS

[PB87] X. Pueyo and P. Brunet.

A parametric space-based scan-line algorithm rendering bicubic surfaces. IEEE Computer Graphics and Applications, 7(11):17–25, November/December 1987. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Pimentel:1994:TYS

[PB94] Ken Pimentel and Brian Blau. Teaching your system to share. IEEE Computer Graphics and Applications, 14(1):60–65, January/ February 1994. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Pelechano:2006:MCT

[PB06] Nuria Pelechano and Norman I. Badler. Modeling crowd and trained leader behavior during building evacuation. IEEE Computer Graphics and Applications, 26(6):80–86, November/December 2006. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Pelachaud:2010:GEI

[PB10] Catherine Pelachaud and Tamy Boubekeur. Guest Editors’ introduction: Digital human faces: From creation to emotion. IEEE Computer Graphics and Applications, 30(4):18–19, July/ August 2010. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Potter:1996:EVE

[PBM^{+}96] Clinton Potter, Rachael Brady, Patrick Moran, Carl Gregory, Bridget Carragher, Nick Kisseberth, Joseph Lyding, and Jason Lindquist.

EVAC: a virtual environment for control of remote imaging instrumentation. IEEE Computer Graphics and Applications, 16(4): 62–66, July/August 1996. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Popescu:2014:FPC

[PBR^{+}14] Voicu Popescu, Bedrich Benes, Paul Rosen, Jian Cui, and Lili Wang. A flexible pinhole camera model for coherent nonuniform sampling. IEEE Computer Graphics and Applications, 34(4):30–41, July/August 2014. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Polys:2008:FSI

[PBSB08] Nicholas F. Polys, Don Brutzman, Anthony Steed, and Johannes Behr. Future standards for immersive VR: Report on the IEEE Virtual Reality 2007 Workshop. IEEE Computer Graphics and Applications, 28(2):94–99, March/ April 2008. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Perin:2016:UGC

[PBV16] Charles Perin, Jeremy Boy, and Frederic Vernier. Using gap charts to visualize the temporal evolution of ranks and scores. IEEE Computer Graphics and Applications, 36(5):38–49, September/October 2016. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2016/05/mcg2016050038-
abs.html.

Pratorius:2015:STF

[PBVH15] Manuel Pratorius, Ulrich Burgbacher, Dimitar Valkov, and Klaus Hinrichs. Sensing thumb-to-finger taps for symbolic input in VR/AR environments. IEEE Computer Graphics and Applications, 35(5):42–54, September/October 2015. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
cg/2015/05/mcg2015050042-
abs.html.

Perovich:2021:DCB

[PCG^{+}21] L. J. Perovich, P. Cai, A. Guo, K. Zimmerman, K. Paseman, D. Espinoza Silva, and J. G. Brody. Data clothing and BigBar-Chart: Designing physical data reports on indoor pollutants for individuals and communities. IEEE Computer Graphics and Applications, 41(1):87–98, January/ February 2021. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Proulx:2007:NGV

[PCH^{+}07] Pascale Proulx, Lynn Chien, Robert Harper, David Schroh, Thomas Kapler, David Jonker, and William Wright. nSpace and GeoTime: a VAST 2006 case study. IEEE Computer Graphics and Applications, 27(5): 46–56, September/October 2007. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Ponti:2016:PAU

[PCJ^{+}16] Moacir Ponti, Arthur A.

Chaves, Fabio R. Jorge, Gabriel B. P. Costa, Adimara Colturato, and Kalinka R. L. J. C. Branco. Precision agriculture: Using lowcost systems to acquire lowaltitude images. IEEE Computer Graphics and Applications, 36(4):14–20, July/ August 2016. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2016/04/mcg2016040014-
abs.html.

Pietroni:2010:STS

[PCOS10] Nico Pietroni, Paolo Cignoni, Miguel Otaduy, and Roberto Scopigno. Solid-texture synthesis: a survey. IEEE Computer Graphics and Applications, 30(4):74–89, July/ August 2010. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Parker:1995:VPT

[PCS95] S. E. Parker, J. C. Cummings, and R. Samtaney. Visualization of plasma turbulence. IEEE Computer Graphics and Applications, 15(2):7–10, March/ April 1995. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Pan:2009:VRD

[PCZ^{+}09] Zhigeng Pan, Wenzhi Chen, Mingmin Zhang, Jianfeng Liu, and Gangshan Wu. Virtual reality in the Digital Olympic Museum. IEEE Computer Graphics and Applications, 29(5):91–95, September/October 2009. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Paris:2009:ADP

[PD09] Sébastien Paris and Stéphane Donikian. Activity-driven populace: a cognitive approach to crowd simulation. IEEE Computer Graphics and Applications, 29(4): 34–43, July/August 2009. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Pirker:2021:PVR

[PD21] Johanna Pirker and Andreas Dengel. The potential of 360^{∘} virtual reality videos and real VR for education: a literature review. IEEE Computer Graphics and Applications, 41(4):76–89, July/ August 2021. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Peters:2009:MGP

[PE09] Christopher Peters and Cathy Ennis. Modeling groups of plausible virtual pedestrians. IEEE Computer Graphics and Applications, 29(4):54–63, July/ August 2009. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Peachy:1988:GIS

[Pea88] Darwin Peachy. Graphics interface 88, selected papers. IEEE Computer Graphics and Applications, 8(6):26– 70, November/December 1988. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Pea:1991:LTM

[Pea91]  Roy D. Pea.  Learning

through multimedia. IEEE Computer Graphics and Applications, 11(4):58–66, July/August 1991. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Peddie:2001:DMT

[Ped01] Jon Peddie. Digital media technology: Industry trends and developments. IEEE Computer Graphics and Applications, 21(1):27–31, January/February 2001. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

                                        computer.org/cga/cg2001/
                                        g1027abs.htm;   http:/
                                        /dlib.computer.org/cg/
                                        books/cg2001/pdf/g1027.
                                        pdf.

Poulton:1992:BFB

[PEMF92] John Poulton, John Eyles, Steven Molnar, and Henry Fuchs. Breaking the framebuffer bottleneck with logicenhanced memories. IEEE Computer Graphics and Applications, 12(6):65–74, November/December 1992. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Pendleton:2010:WAB

[Pen10] Chris Pendleton. The world according to Bing. IEEE Computer Graphics and Applications, 30(4):15–17, July/August 2010. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Perlin:2016:FRH

[Per16] Ken Perlin. Future reality: How emerging technologies will change language itself. IEEE Computer Graphics and Applications, 36(3):84–89, May/ June 2016. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2016/03/mcg2016030084-
abs.html.

Perin:2021:WSL

[Per21] Charles Perin. What students learn with personal data physicalization. IEEE Computer Graphics and Applications, 41(6):48–58, November/December 2021. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Peters:2012:CV

[Pet12]  Jorg Peters.  Changing

variables. IEEE Computer Graphics and Applications, 32(3):88–93, May/ June 2012. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Perlmutter:1983:CHB

[PF83] Robert J. Perlmutter and Stephen S. Friedland. Computergenerated holograms in biology and medicine. IEEE Computer Graphics and Applications, 3(5):47–51, August 1983. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Powers:1986:CGV

[PFA86] Thomas Powers, Andrea Frankel, and David Arnold. The computer graphics virtual device interface. IEEE Computer Graphics and Applications, 6(8):33–41, August 1986. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Paloc:2006:ORS

[PFB06] Céline Paloc, Alessandro Faraci, and Fernando Bello. Online remeshing for soft tissue simulation in surgical training. IEEE Computer Graphics and Applications, 26(6):24–34, November/December 2006. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Peterson:1995:ISD

[PFH95] Philip Peterson, F. David Fracchia, and Brian Hayden. Integrating spatial data display with virtual reconstruction. IEEE Computer Graphics and Applications, 15(4):40–46, July/ August 1995. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Pastor:2003:RTA

[PFS03] Oscar Meruvia Pastor, Bert Freudenberg, and Thomas Strothotte. Real-time animated stippling. IEEE Computer Graphics and Applications, 23(4):62–68, July/August 2003. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

csdl.computer.org/comp/
mags/cg/2003/04/extras/
g4062.htm;   http://csdl.
computer.org/comp/mags/
cg/2003/04/g4062abs.htm;
 http://csdl.computer.
org/dl/mags/cg/2003/04/
g4062.htm;   http://csdl.
computer.org/dl/mags/cg/
2003/04/g4062.pdf.

Pruyn:1993:ECG

[PG93] Peter W. Pruyn and Donald P. Greenberg. Exploring 3D computer graphics in cockpit avionics. IEEE Computer Graphics and Applications, 13(3):28–35, May/June 1993. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Pfister:2004:PBC

[PG04] Hanspeter Pfister and Markus Gross. Point-based computer graphics. IEEE Computer Graphics and Applications, 24(4):22–23, July/ August 2004. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cg/2004/04/g4022.pdf;
http://csdl.computer.org/

                                        dl/mags/cg/2004/04/g4022.
                                        htm.

Potter:2013:VUM

[PGA13] Kristin Potter, Samuel Gerber, and Erik W. Anderson. Visualization of uncertainty without a mean. IEEE Computer Graphics and Applications, 33(1):75–79, January/ February 2013. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Preston:2019:UAV

[PGM19] A. Preston, M. Gomov, and K. Ma. Uncertaintyaware visualization for analyzing heterogeneous wildfire detections. IEEE Computer Graphics and Applications, 39(5):72–82, September/October 2019. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Plaisant:2008:EVA

[PGS^{+}08] Catherine Plaisant, Georges Grinstein, Jean Scholtz, Mark Whiting, Theresa O’Connell, Sharon Laskowski, Lynn Chien, Annie Tat, William Wright, Carsten Görg, Zhicheng Liu, Neel Parekh, Kanupriya Singhal, and John Stasko. Evaluating visual analytics at the 2007 VAST Symposium contest. IEEE Computer Graphics and Applications, 28(2):12–21, March/ April 2008. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Plaisant:2009:GEI

[PGS09] Catherine Plaisant, Georges Grinstein, and Jean Scholtz. Guest Editors’ introduction: Visual-analytics evaluation. IEEE Computer Graphics and Applications, 29(3):16– 17, May/June 2009. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Panagiotidou:2020:DBM

[PGV20] G. Panagiotidou, S. Görücü, and A. Vande Moere. Data badges: Making an academic profile through a DIY wearable physicalization. IEEE Computer Graphics and Applications, 40(6): 51–60, November/December 2020. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Pavlidis:1998:UCX

[PH98] Theo Pavlidis and Kevin Hunter. Using color in the X Window System versus Microsoft Windows: Part 1. IEEE Computer Graphics and Applications, 18(6):64–73, November/December 1998. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

computer.org/cga/cg1998/

                                        g6064abs.htm;   http:/
                                        /dlib.computer.org/cg/
                                        books/cg1998/pdf/g6064.
                                        pdf.

Pavlidis:1999:UCX

[PH99] Theo Pavlidis and Kevin Hunter. Using color in the X Window System versus Microsoft Windows: Part 2. IEEE Computer Graphics and Applications, 19(1):75–83, January/ February 1999. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

                                        computer.org/cga/cg1999/
                                        g1075abs.htm;   http:/
                                        /dlib.computer.org/cg/
                                        books/cg1999/pdf/g1075.
                                        pdf.

Pedersen:2020:SAC

[PHAS20] K. Pedersen, V. Hulusic, P. Amelidis, and T. Slattery. Spatialized audio in a custom-built OpenGLbased ear training virtual environment. IEEE Computer Graphics and Applications, 40(5):67–81, September/October 2020. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Paier:2021:EBF

[PHE21] Wolfgang Paier, Anna Hilsmann, and Peter Eisert. Example-based facial animation of virtual reality avatars using autoregressive neural networks. IEEE Computer Graphics and Applications, 41(4): 52–63, July/August 2021. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Phillips:1991:IMV

[Phi91] Richard L. Phillips. An interpersonal multimedia visualization system. IEEE Computer Graphics and Applications, 11(3):20–27, May/June 1991. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Patro:2011:SSS

[PIBV11] Robert Patro, Cheuk Yiu Ip, Sujal Bista, and Amitabh Varshney. Social snapshot: a system for temporally coupled social photography. IEEE Computer Graphics and Applications, 31(1):74– 84, January/February 2011. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Pickover:1989:SRS

[Pic89] Clifford A. Pickover. A short recipe for seashell synthesis. IEEE Computer Graphics and Applications, 9(6):8–11, November/December 1989. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Pickover:1990:VCL

[Pic90] Clifford A. Pickover. Visualizing chaos: Lyapunov surfaces and volumes. IEEE Computer Graphics and Applications, 10(2):15–19, March/April 1990. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Picott:1992:ELA

[Pic92] Kevin P. Picott. Extensions of the linear and area lighting models. IEEE Computer Graphics and Applications, 12(2):31–38, March/ April 1992. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Pickover:1995:GET

[Pic95] Clifford A. Pickover. Generating extraterrestrial terrain. IEEE Computer Graphics and Applications, 15(2):18–21, March/April 1995. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Picard:2000:SE

[Pic00] Rosalind W. Picard. Synthetic emotion. IEEE Computer Graphics and Applications, 20(1):52–53, January/ February 2000. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL http://dlib.
                                        computer.org/cg/books/
                                        cg2000/pdf/g1052.pdf.

Piegl:1987:IDI

[Pie87a] L. Piegl. Interactive data interpolation by rational Bézier curves. IEEE Computer Graphics and Applications, 7(4):45–58, April 1987. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Piegl:1987:ICP

[Pie87b] Leslie Piegl. Infinite control points: a method for representing surfaces of revolution using boundary data. IEEE Computer Graphics and Applications, 7(3):45– 55, March/April 1987. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Piegl:1987:LDS

[Pie87c] Leslie Piegl. Less data for shapes. IEEE Computer Graphics and Applications, 7(8):48–49, August 1987. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Piegl:1991:NS

[Pie91] Les Piegl. On NURBS:

a survey. IEEE Computer Graphics and Applications, 11(1):55–71, January/ February 1991. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Piekarski:2006:MTM

[Pie06] Wayne Piekarski. 3D modeling with the Tinmith Mobile Outdoor Augmented Reality System. IEEE Computer Graphics and Applications, 26(1):14–17, January/ February 2006. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Pinedo:1991:WCM

[Pin91] David Pinedo. Window clipping methods in graphics accelerators. IEEE Computer Graphics and Applications, 11(3):75–84, May/ June 1991. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Pitot:1993:VP

[Pit93] Paul Pitot. The Voxar Project. IEEE Computer Graphics and Applications, 13(1):27–33, January/ February 1993. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Papathomas:1988:TAD

[PJC88] Thomas V. Papathomas, Bela Julesz, and Sarah E. Chodrow. True 3D animation for displaying VLSI modeling data (applications briefs). IEEE Computer Graphics and Applications, 8(1):6–9, January/February 1988. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Phillips:2024:TCA

[PJC24] Connor Phillips, Junfeng Jiao, and Emmalee Clubb. Testing the capability of AI art tools for urban design. IEEE Computer Graphics and Applications, 44(2): 37–45, March/April 2024. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Prevost:1996:VBV

[PJJF96] N. Prevost, R. E. Jennings, L. Jorgenson, and F. D. Fracehia. Visualization blackboard: Visualization in propositional logic. IEEE Computer Graphics and Applications, 16(2):6–8, March/April 1996. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Pulleyblank:1987:FVC

[PK87] Ronald W. Pulleyblank and John A. Kapenga. The feasibility of a VLSI chip for ray tracing bicubic patches. IEEE Computer Graphics and Applications, 7(3):33– 44, March/April 1987. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Pulli:2008:MG

[PK08] Kari Pulli and Scott Klemmer. Mobile graphics. IEEE Computer Graphics and Applications, 28(4):18–19, July/August 2008. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cg/2008/04/mcg2008040018.
pdf.

Pu:2007:NDC

[PKJ^{+}07] Jiantao Pu, Yagnanarayanan Kalyanaraman, Subramaniam Jayanti, Karthik Ramani, and Zygmunt Pizlo. Navigation and discovery in 3D CAD repositories. IEEE Computer Graphics and Applications, 27(4): 38–47, July/August 2007. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Park:2018:GAP

[PKM18] Ji Hwan Park, Arie Kaufman, and Klaus Mueller. Graphoto: Aesthetically pleasing charts for casual information visualization. IEEE Computer Graphics and Applications, 38(6):67– 82, ???? 2018. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL https:

//ieeexplore.ieee.org/
document/8617746/.

Plass:2022:UED

[PKN^{+}22] Markus Plass, Michaela Kargl, Patrick Nitsche, Emilian Jungwirth, Andreas Holzinger, and Heimo Müller. Understanding and explaining diagnostic paths: Toward augmented decision making. IEEE Computer Graphics and Applications, 42(6):47–57, November/December 2022. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Papaioannou:2001:VAA

[PKT01] Georgios Papaioannou, Evaggelia-Aggeliki Karabassi, and Theoharis Theoharis. Virtual archaeologist: Assembling the past. IEEE Computer Graphics and Applications, 21(2):53–59, March/ April 2001. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

computer.org/cga/cg2001/
g2053abs.htm;   http:/
/dlib.computer.org/cg/
books/cg2001/pdf/g2053.
pdf.

Perer:2019:VDS

[PL19] A. Perer and S. Liu. Visualization in data science. IEEE Computer Graphics and Applications, 39(5): 18–19, September/October 2019. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Pfister:2001:VVT

[PLB^{+}01] Hanspeter Pfister, Bill Lorensen, Chandrajit Baja, Gordon Kinklmann, Will Schroeder, Lisa Sobierajski Avila, Ken Martin, Raghu Machiraju, and Jinho Lee. Visualization viewpoints: The transfer function bake-off. IEEE Computer Graphics and Applications, 21(3):16–23, May/ June 2001. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL http://dlib.
                                        computer.org/cg/books/
                                        cg2001/pdf/g3016.pdf.

Papenhausen:2019:PVO

[PLM^{+}19] Eric Papenhausen, M. Harper Langston, Benoit Meister, Richard A. Lethin, and Klaus Mueller. PUMA-V: Optimizing parallel code performance through interactive visualization. IEEE Computer Graphics and Applications, 39(1):84–99, January/February 2019. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL https:

                                        //ieeexplore.ieee.org/
                                        document/8663646/.

Pang:2016:TFC

[PLS^{+}16] Xufang Pang, Rynson W. H. Lau, Zhan Song, Yangyan Li, and Shengfeng He. A tool-free calibration method for turntable-based 3D scanning systems. IEEE Computer Graphics and Applications, 36(1):52–61, January/ February 2016. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
cg/2016/01/mcg2016010052-
abs.html.

Pfaff:1982:TCI

[PM82] Guenther E. Pfaff and Gerd Maderlechner. Tools for configuring interactive pictureprocessing systems. IEEE Computer Graphics and Applications, 2(3):35–48, July/ August 1982. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Puk:1986:GTD

[PM86] Richard F. Puk and John I.

McConnell. GKS-3D: a three-dimensional extension to the Graphical Kernel System. IEEE Computer Graphics and Applications, 6(8):42–49, August 1986. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Pu:1997:VB

[PM97] Pearl Pu and George Melissargos. Visualization blackboard. IEEE Computer Graphics and Applications, 17(4):6–8, July/ August 1997. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Parker:2006:IPT

[PMI06] J. Karen Parker, Regan L. Mandryk, and Kori M. Inkpen. Integrating point and touch for interaction with digital tabletop displays. IEEE Computer Graphics and Applications, 26(5):28–35, September/ October 2006. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Peck:2021:DBC

[PMQ21] Tabitha C. Peck, Kyla A. McMullen, and John Quarles. DiVRsify: Break the cycle and develop VR for everyone. IEEE Computer Graphics and Applications, 41(6):133–142, November/ December 2021. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Peinado:2009:FBA

[PMR^{+}09] Manuel Peinado, Damien Maupu, Daniel Raunhardt, Daniel Meziat, Daniel Thalmann, and Ronan Boulic. Full-body avatar control with environment awareness. IEEE Computer Graphics and Applications, 29(3):62–75, May/June 2009. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Pair:2003:FCH

[PNPS03] Jarrell Pair, Ulrich Neumann, Diane Piepol, and Bill Swartout. Flat-World: Combining Hollywood set-design techniques with VR. IEEE Computer Graphics and Applications, 23(1):12–15, January/ February 2003. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://dlib.
computer.org/cg/books/
cg2003/pdf/g1012.pdf.

Phillips:1984:ALD

[PO84] Mark B. Phillips and Garrett M. Odell. An algorithm for locating and displaying the intersection of two arbitrary surfaces. IEEE Computer Graphics and Applications, 4(9):48–56, 58, September 1984. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Potel:1996:CGD

[Pot96] Michael J. Potel. Computer graphics and DNA sequencing. IEEE Computer Graphics and Applications, 16(6): 14–19, November/December 1996. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Potel:1998:AMS

[Pot98] Michael J. Potel. Applications: Motion sick in cyberspace. IEEE Computer Graphics and Applications, 18(1):16–21, January/ February 1998. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

computer.org/cga/cg1998/
g1016abs.htm;   http:/
/dlib.computer.org/cg/
books/cg1998/pdf/g1016.
pdf.

Potel:2000:A

[Pot00]  Mike Potel.  Applica-

tions 2000. IEEE Computer Graphics and Applications, 20(1):42–43, January/ February 2000. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://dlib.
computer.org/cg/books/
cg2000/pdf/g1042.pdf.

Potel:2004:DA

[Pot04] Mike Potel. A decade of applications. IEEE Computer Graphics and Applications, 24(6):14–19, November/ December 2004. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cg/2004/06/g6014.pdf;
http://csdl.computer.org/
comp/mags/cg/2004/07/g7toc.
htm;   http://csdl.computer.
org/dl/mags/cg/2004/06/
g6014.htm.

Potel:2012:CD

[Pot12] Mike Potel. CG&A’s departments. IEEE Computer Graphics and Applications, 32(2):6–7, March/ April 2012. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Pham:1995:CCI

[PP95] Binh Pham and Glen Pringle. Color correction for an image sequence. IEEE Computer Graphics and Applications, 15(3):38–42, May/June 1995. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Pang:1998:GEI

[PP98] Alex Pang and Hans-Georg Pagendarm. Guest Editors’ introduction: Visualization for everyone. IEEE Computer Graphics and Applications, 18(4):47–48, July/ August 1998. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Ponce:2011:MNB

[PP11] Ernesto Ponce and Daniel Ponce. Modeling neck and brain injuries in infants. IEEE Computer Graphics and Applications, 31(6): 90–96, November/December 2011. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Poelke:2012:DCC

[PP12] Konstantin Poelke and Konrad Polthier. Domain coloring of complex functions: An implementation-oriented introduction. IEEE Computer Graphics and Applications, 32(5):90–97, September/October 2012. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Ponce:2014:MHA

[PPA14] Ernesto Ponce, Daniel Ponce, and Max Andresen. Modeling heading in adult soccer players. IEEE Computer Graphics and Applications, 34(5):8–13, September/October 2014. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL http://csdl.
                                        computer.org/csdl/mags/
                                        cg/2014/05/mcg2014050008-
                                        abs.html.

Pasko:2005:BBF

[PPK05] Galina I. Pasko, Alexander A. Pasko, and Tosiyasu L. Kunii. Bounded blending for function-based shape modeling. IEEE Computer Graphics and Applications, 25(2): 36–45, March/April 2005. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Papadopoulos:2015:RDI

[PPKM15] Charilaos Papadopoulos, Kaloian Petkov, Arie E.

Kaufman, and Klaus Mueller. The reality deck — an immersive gigapixel display. IEEE Computer Graphics and Applications, 35(1):33– 45, January/February 2015. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL

http://csdl.computer.org/
csdl/mags/cg/2015/01/mcg2015010033-
abs.html.

Patel:2023:CSE

[PPVH23] Dixit Bharatkumar Patel, Yong Pei, Mitesh Vasoya, and Paul J. Hershberger. Computer-supported experiential learning-based tool for healthcare skills. IEEE Computer Graphics and Applications, 43(2):57–68, March/April 2023. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Pang:2012:FRD

[PQC^{+}12] Wai-Man Pang, Jing Qin, Michael Cohen, Pheng-Ann Heng, and Kup-Sze Choi. Fast rendering of diffusion curves with triangles. IEEE Computer Graphics and Applications, 32(4):68–78, July/August 2012. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Powell:2008:SSP

[PR08] Alexander Powell and Jarek Rossignac. ScrewBender:

Smoothing piecewise helical motions. IEEE Computer Graphics and Applications, 28(1):56–63, January/ February 2008. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Poulin:2012:MA

[PR12] Pierre Poulin and Holly Rushmeier. Material appearance. IEEE Computer Graphics and Applications, 32(2):22–23, March/ April 2012. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Pratt:1984:SMI

[Pra84a] M. J. Pratt. Solid modeling and the interface between design and manufacture. IEEE Computer Graphics and Applications, 4(7):52–56, 58–59, July/ August 1984. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Pratt:1984:SPI

[Pra84b] V. R. Pratt. Standards and performance issues in the workstation market. IEEE Computer Graphics and Applications, 3(4):71–76, April 1984. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Pearlman:2007:VDD

[PRdJ07] Jason Pearlman, Penny Rheingans, and Marie des Jardins. Visualizing diversity and depth over a set of objects. IEEE Computer Graphics and Applications, 27(5):35–45, September/October 2007. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Prince:2000:CAN

[Pri00] Patric D. Prince. Computer art in the new millennium. IEEE Computer Graphics and Applications, 20(1):26–27, January/ February 2000. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://dlib.
computer.org/cg/books/
cg2000/pdf/g1026.pdf.

Paul:2015:SFA

[PRN15] Celeste Lyn Paul, Randall Rohrer, and Bohdan Nebesh. A “Design First” approach to visualization innovation. IEEE Computer Graphics and Applications, 35(1):12–18, January/ February 2015. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://csdl.
computer.org/csdl/mags/
cg/2015/01/mcg2015010012-
abs.html.

Prusinkiewicz:1990:EBP

[Pru90] Przemyslaw Prusinkiewicz. Exploring the beauty of plants. IEEE Computer Graphics and Applications, 10(2):3–6, March/April 1990. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Putnam:1986:BON

[PS86] L. K. Putnam and P. A. Subrahmanyam. Boolean operations on N-Dimensional objects. IEEE Computer Graphics and Applications, 6(6):43–51, June 1986. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Prusinkiewicz:1988:KCA

[PS88] Przemyslaw Prusinkiewicz and Glen Sandness. Koch curves as attractors and repellers. IEEE Computer Graphics and Applications, 8(6):26–40, November/December 1988. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Perer:2009:ISV

[PS09] Adam Perer and Ben Shneiderman. Integrating statistics and visualization for exploratory power: From longterm case studies to design guidelines. IEEE Computer Graphics and Applications, 29(3):39–51, May/ June 2009. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Plaisant:2022:LTE

[PS22] Catherine Plaisant and Ben Shneiderman. Lightning and thunder: The early days of interactive information visualization at the University of Maryland. IEEE Computer Graphics and Applications, 42(4):103–113, July/ August 2022. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Parker:2009:ABC

[PSE^{+}09] James R. Parker, Nathan Sorenson, Nooshin Esmaeili, Ronan Sicre, Phillipa Gil, Vishal Kochlar, Lori Shyba, and John Heerema. Applications: The booze cruise: Impaired driving in virtual spaces. IEEE Computer Graphics and Applications, 29(2):6–10, March/ April 2009. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Papathomas:1987:SAV

[PSJ87] Thomas V. Papathomas, James A. Schiavone, and Bela Julesz. Stereo animation for very large data bases: Case study — meteorology. IEEE Computer Graphics and Applications, 7(9):18–27, September 1987. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Pang:2013:EVR

[PSX13] Xufang Pang, Zhan Song, and Wuyuan Xie. Extracting valley-ridge lines from point-cloud-based 3D fingerprint models. IEEE Computer Graphics and Applications, 33(4):73–81, July/ August 2013. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Piegl:1989:MRB

[PT89] Leslie Piegl and Wayne Tiller. A menagerie of rational B-spline circles. IEEE Computer Graphics and Applications, 9(5):48– 56, September 1989. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Payne:1990:SMB

[PT90] Bradley A. Payne and Arthur W. Toga. Surface mapping brain function on 3D models. IEEE Computer Graphics and Applications, 10(5):33–41, September 1990. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Payne:1992:DFM

[PT92] Bradley A. Payne and Arthur W. Toga. Distance field manipulation of surface models. IEEE Computer Graphics and Applications, 12(1):65–71, January/ February 1992. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Payne:1994:SRM

[PT94] Bradley A. Payne and Arthur W. Toga. Surface reconstruction by multiaxial triangulation. IEEE Computer Graphics and Applications, 14(6):28–35, November/December 1994. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Piegl:2000:RCP

[PT00] Les A. Piegl and Wayne Tiller. Reducing control points in surface interpolation. IEEE Computer Graphics and Applications, 20(5):70–74, September/October 2000. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

computer.org/cga/cg2000/
g5070abs.htm;   http:/
/dlib.computer.org/cg/
books/cg2000/pdf/g5070.
pdf.

Piegl:2003:ASR

[PT03] Les A. Piegl and Wayne Tiller. Approximating surfaces of revolution by nonrational B-splines. IEEE Computer Graphics and Applications, 23(3):46–52, May/ June 2003. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://dlib.
computer.org/cg/books/
cg2003/pdf/g3046.pdf;
http://www.computer.org/
cga/cg2003/g3046abs.htm.

Posada:2015:VCK

[PTB^{+}15] Jorge Posada, Carlos Toro, Inigo Barandiaran, David Oyarzun, Didier Stricker, Raffaele de Amicis, Eduardo B. Pinto, Peter Eisert, Jurgen Dollner, and Ivan Vallarino. Visual computing as a key enabling technology for industrie 4.0 and industrial Internet. IEEE Computer Graphics and Applications, 35(2):26–40, March/ April 2015. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
cg/2015/02/mcg2015020026-
abs.html.

Petrov:1998:ODB

[PTR^{+}98] Michael Petrov, Andrey Talapov, Timothy Robertson, Alexei Lebedev, Alexander Zhilyaev, and Leonid Polonskiy. Optical 3D digitizers: Bringing life to the virtual world. IEEE Computer Graphics and Applications, 18(3):28–37, May/ June 1998. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

computer.org/cga/cg1998/
g3028abs.htm;   http:/
/dlib.computer.org/cg/
books/cg1998/pdf/g3028.
pdf.

Patel:1996:ACS

[PVML96] Vikas V. Patel, Michael W.

Vannier, Jeffrey L. Marsh, and Lun-Jou Lo. Assessing craniofacial surgical simulation. IEEE Computer Graphics and Applications, 16(1):46–54, January/ February 1996. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Pollefeys:2003:RAF

[PVV^{+}03] Marc Pollefeys, Luc Van Gool, Maarten Vergauwen, Kurt Cornelis, Frank Verbiest, and Jan Tops. 3D recording for archaeological fieldwork. IEEE Computer Graphics and Applications, 23(3):20–27, May/ June 2003. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://dlib.
computer.org/cg/books/
cg2003/pdf/g3020.pdf;

                                        http://www.computer.org/
                                        cga/cg2003/g3020abs.htm.

Pretorius:2007:BSG

[PvW07] A. Johannes Pretorius and Jarke J. van Wijk. Bridging the semantic gap: Visualizing transition graphs with user-defined diagrams. IEEE Computer Graphics and Applications, 27(5): 58–66, September/October 2007. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Pang:1997:CVC

[PW97] Alex Pang and Craig Wittenbrink. Collaborative 3D visualization with CSpray. IEEE Computer Graphics and Applications, 17(2): 32–41, March/April 1997. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Potel:2014:VYA

[PW14] Mike Potel and Pak Chung Wong. Visualizing 20 years of applications. IEEE Computer Graphics and Applications, 34(6):6–11, November/December 2014. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

                                        csdl.computer.org/csdl/
                                        mags/cg/2014/06/mcg2014060006.
                                        html.

Pang:2004:ELV

[PWH04] Wai-Man Pang, Tien-Tsin Wong, and Pheng-Ann Heng. Estimating light vectors in real time. IEEE Computer Graphics and Applications, 24(3):36–43, May/June 2004. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cg/2004/03/g3036abs.htm;
 http://csdl.computer.
org/dl/mags/cg/2004/03/
g3036.htm;   http://csdl.
computer.org/dl/mags/cg/
2004/03/g3036.pdf.

Peleg:1997:IMM

[PWW97] Alex Peleg, Sam Wilkie, and Uri Weiser. Intel MMX for multimedia PCs. Communications of the ACM, 40(1):24–38, January 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See also Blinn’s comments [Bli97b] about MMX instruction set deficiencies.

Prince:2002:ARC

[PXC02] Simon Prince, Ke Xu, and Adrian Cheok. Augmented reality camera tracking with homographies. IEEE Computer Graphics and Applications, 22(6):39–45, November/December 2002. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http:

//dlib.computer.org/cg/
books/cg2002/pdf/g6039.
pdf;   http://www.computer.

                                        org/cga/cg2002/g6039abs.
                                        htm.

Posada:2018:GMT

[PZD^{+}18] Jorge Posada, Mikel Zorrilla, Ana Dominguez, Bruno Simoes, Peter Eisert, Didier Stricker, Jason Rambach, Jurgen Dollner, and Miguel Guevara. Graphics and media technologies for operators in Industry 4.0. IEEE Computer Graphics and Applications, 38(5):119–132, September/ October 2018. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        cg/2018/05/mcg2018050119-
                                        abs.html.

Qu:2015:VAM

[QC15] Huamin Qu and Qing Chen. Visual analytics for MOOC data. IEEE Computer Graphics and Applications, 35(6):69–75, November/ December 2015. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL http://www.
                                        computer.org/csdl/mags/
                                        cg/2015/06/mcg2015060069-
                                        abs.html.

Qin:2010:LBM

[QCP^{+}10] Jing Qin, Yim-Pan Chui, Wai-Man Pang, Kup-Sze Choi, and Pheng-Ann Heng. Learning blood management in orthopedic surgery through gameplay. IEEE Computer Graphics and Applications, 30(2):45–57, March/April 2010. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Qin:2002:ACS

[QNT02] Xueying Qin, Eihachiro Nakamae, and Katsumi Tadamura. Automatically compositing still images and landscape video sequences. IEEE Computer Graphics and Applications, 22(1):68–78, January/ February 2002. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://dlib.
computer.org/cg/books/
cg2002/extras/g1068x1.
mpg;   http://dlib.computer.
org/cg/books/cg2002/g1068abs.
htm;   http://dlib.computer.
org/cg/books/cg2002/pdf/
g1068.pdf.

Robertson:1993:GEI

[RA93] Philip K. Robertson and David J. Abel. Guest Editors’ introduction: Graphics and environmental decision making. IEEE Computer Graphics and Applications, 13(2):25–??, March/ April 1993. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Requicha:2006:CCN

[RA06] Ari Requicha and Dan Arbuckle. CAD/CAM for nanoscale self-assembly. IEEE Computer Graphics and Applications, 26(2): 88–91, March/April 2006. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Reinhard:2001:CTB

[RAGS01] Erik Reinhard, Michael Ashikhmin, Bruce Gooch, and Peter Shirley. Color transfer between images. IEEE Computer Graphics and Applications, 21(5): 34–41, September/October 2001. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL

                                        http://dlib.computer.org/
                                        cg/books/cg2001/g5034abs.
                                        htm;   http://dlib.computer.
                                        org/cg/books/cg2001/pdf/
                                        g5034.pdf.

Rao:1997:RRW

[Rao97] Ramana Rao. From research to real world with Z-GUI. IEEE Computer Graphics and Applications, 17(4): 71–73, July/August 1997. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Rafferty:1998:IAA

[RAPL98] Matthew M. Rafferty, Daniel G. Aliaga, Voicu Popescu, and Anselmo A. Lastra. Images for accelerating architectural walkthroughs. IEEE Computer Graphics and Applications, 18(6):38–45, November/December 1998. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

computer.org/cga/cg1998/
g6038abs.htm;   http:/
/dlib.computer.org/cg/
books/cg1998/pdf/g6038.
pdf.

Rosenblum:1992:V

[RB92] Lawrence J. Rosenblum and Bruce E. Brown. Visualization. IEEE Computer Graphics and Applications, 12(4):18–??, July/August 1992. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Rioux:1993:WLS

[RB93] Marc Rioux and Tony Bird.

White lasers, synced scan. IEEE Computer Graphics and Applications, 13(3):15– 17, May/June 1993. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Rodger:2000:CRP

[RB00] James C. Rodger and Roger A. Browse. Choosing rendering parameters for effective communication of 3D shape. IEEE Computer Graphics and Applications, 20(2):20–28, March/ April 2000. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

                                        computer.org/cga/cg2000/
                                        g2020abs.htm;   http:/
                                        /dlib.computer.org/cg/
                                        books/cg2000/pdf/g2020.
                                        pdf.

Reitinger:2006:LSP

[RBBS06] Bernhard Reitinger, Alexander Bornik, Reinhard Beichel, and Dieter Schmalstieg. Liver surgery planning using virtual reality. IEEE Computer Graphics and Applications, 26(6): 36–47, November/December 2006. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Rysavy:2014:DGB

[RBD14] Steven J. Rysavy, Dennis Bromley, and Valerie Daggett. DIVE: A graph-based visual-analytics framework for big data. IEEE Computer Graphics and Applications, 34(2): 26–37, March/April 2014. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Rossler:2008:DSG

[RBE08] Friedemann Rößler, Ralf P. Botchen, and Thomas Ertl. Dynamic shader generation for GPU-based multivolume ray casting. IEEE Computer Graphics and Applications, 28(5):66–77, September/October 2008. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Rosenblum:1995:GEI

[RBF95] Lawrence J. Rosenblum, Steve Bryson, and Steven K. Feiner. Guest Editors’ introduction: Virtual reality unbound. IEEE Computer Graphics and Applications, 15(5):19–21, September 1995. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Rouiller:2013:PSV

[RBK^{+}13] Olivier Rouiller, Bernd Bickel, Jan Kautz, Wojciech Matusik, and Marc Alexa. 3D-printing spatially varying BRDFs. IEEE Computer Graphics and Applications, 33(6):48–57, November/December 2013. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Reiser:2017:MTI

[RBMS17] Susan Reiser, Rebecca Bruce, Jackson Martin, and Brent Skidmore. Making together: An interdisciplinary, inter-institutional assistive-technology project. IEEE Computer Graphics and Applications, 37(5): 9–14, September/October 2017. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL

https://www.computer.org/

                                        csdl/mags/cg/2017/05/mcg2017050009-
                                        abs.html.

Ribarsky:1994:VBV

[RBOV94] William Ribarsky, Jay Bolter, Augusto Op den Bosch, and Ron Van Teylingen. Visualization blackboard: Visualization and analysis using virtual reality. IEEE Computer Graphics and Applications, 14(1):10– 12, January/February 1994. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Rhyne:1993:VED

[RBR^{+}93] Theresa Rhyne, Mark Bolstad, Penny Rheingans, Lynne Petterson, and Walter Shackelford. Visualizing environmental data at the EPA. IEEE Computer Graphics and Applications, 13(2):34–38, March/ April 1993. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Rober:2021:VCS

[RBS21] N. Röber, M. Böttinger, and B. Stevens. Visualization of climate science simulation data. IEEE Computer Graphics and Applications, 41(1):42–48, January/ February 2021. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Rosenblum:1998:GEI

[RBT98] Lawrence Rosenblum, Grigore Burdea, and Susumu Tachi. Guest Editors’ introduction: VR reborn. IEEE Computer Graphics and Applications, 18(6):21–23, November/December 1998. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL

http://dlib.computer.org/
cg/books/cg1998/pdf/g6021.
pdf.

Regenbrecht:2005:ARP

[RBW05] Holger Regenbrecht, Gregory Baratoff, and Wilhelm Wilke. Augmented reality projects in the automotive and aerospace industries. IEEE Computer Graphics and Applications, 25(6): 48–56, November/December 2005. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Richer:1985:GGI

[RC85] Mark H. Richer and William J.

Clancey. GUIDON-WATCH: a graphic interface for viewing a knowledge-based system. IEEE Computer Graphics and Applications, 5(11):51–64, November/ December 1985. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Rosli:2015:GPM

[RC15] Muhammad Hafiz Wan Rosli and Andres Cabrera. Gestalt principles in multimodal data representation. IEEE Computer Graphics and Applications, 35(2):80–87, March/ April 2015. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
cg/2015/02/mcg2015020080-
abs.html.

Rose:1998:VAM

[RCB98] Charles Rose, Michael F. Cohen, and Bobby Bodenheimer. Verbs and adverbs: Multidimensional motion interpolation. IEEE Computer Graphics and Applications, 18(5):32–41, September/October 1998. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

computer.org/cga/cg1998/
g5032abs.htm;   http:/
/dlib.computer.org/cg/
books/cg1998/pdf/g5032.
pdf.

Robertson:2005:LDU

[RCB^{+}05] George Robertson, Mary Czerwinski, Patrick Baudisch, Brian Meyers, Daniel Robbins, Greg Smith, and Desney Tan. The largedisplay user experience. IEEE Computer Graphics and Applications, 25(4):

44–51, July/August 2005. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Rew:1990:NIS

[RD90] Russ Rew and Glenn Davis.

NetCDF: An interface for scientific data access. IEEE Computer Graphics and Applications, 10(4):76–82, July/August 1990. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Rodrigues:2020:DCR

[RdCFP^{+}20] R. Rodrigues, P. d. C. Ferreira, R. Prada, P. Paulino, and A. M. V. Simão. Developing children’s regulation of learning in problemsolving with a serious game. IEEE Computer Graphics and Applications, 40(5): 26–40, September/October 2020. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Rosenblum:1997:PVS

[RDDT97] Larry Rosenblum, Jim Durbin, Robert Doyle, and David Tate. Projects in VR: Situational awareness using the Responsive Workbench. IEEE Computer Graphics and Applications, 17(4): 12–13, July/August 1997. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Rushmeier:2007:RNF

[RDDY07] Holly Rushmeier, Jason Dykes, John Dill, and Peter Yoon. Revisiting the need for formal education in visualization. IEEE Computer Graphics and Applications, 27(6):12–16, November/December 2007. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Rolland:2002:VID

[RDH^{+}02] Jannick Rolland, Larry Davis, Yonggang Ha, Catherine Meyer, Vesselin Shaoulov, Avni Akcay, Haocheng Zheng, Robert Banks, and Benjamin Del Vento. 3D visualization and imaging in distributed collaborative environments. IEEE Computer Graphics and Applications, 22(1):11–13, January/ February 2002. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://dlib.
computer.org/cg/books/
cg2002/g1011abs.htm;
http://dlib.computer.org/
cg/books/cg2002/pdf/g1011.
pdf.

Rosenblum:1996:PVS

[RDO^{+}96] Lawrence Rosenblum, Jim Durbin, Upul Obeysekare, Linda Sibert, David Tate, James Templeman, Jyoti Agrawal, Daniel Fasulo, Thomas Meyer, Greg Newton, and Amit Shalev.

Projects in VR: Shipboard VR: From damage control to design. IEEE Computer Graphics and Applications, 16(6):10–13, November/December 1996. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Ropinski:2010:AVI

[RDRS10] Timo Ropinski, Christian Doring, and Christof Rezk-Salama. Advanced volume illumination with unconstrained light source positioning. IEEE Computer Graphics and Applications, 30(6):29–41, November/December 2010. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Reddy:2001:POG

[Red01] Martin Reddy. Perceptually optimized 3D graphics. IEEE Computer Graphics and Applications, 21(5):68–75, September/ October 2001. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://dlib.
computer.org/cg/books/
cg2001/g5068abs.htm;
http://dlib.computer.org/
cg/books/cg2001/pdf/g5068.
pdf.

Reed:1984:GWE

[Ree84] Jon Reed. Graphics in workstations — emerging systems and solutions. IEEE Computer Graphics and Applications, 4(8):11–16, August 1984. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Ryall:2006:IDW

[REF^{+}06] Kathy Ryall, Alan Esenther, Clifton Forlines, Chia Shen, Sam Shipman, Meredith Ringel Morris, Katherine Everitt, and Frédéric D. Vernier. Identity-differentiating widgets for multiuser interactive surfaces. IEEE Computer Graphics and Applications, 26(5):56–64, September/October 2006. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Reisman:1991:DMA

[Rei91] Sorel Reisman. Developing multimedia applications. IEEE Computer Graphics and Applications, 11(4): 52–57, July/August 1991. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Reilly:2017:NHJ

[Rei17] Susan Reilly. The need to help journalists with data and information visualization. IEEE Computer Graphics and Applications, 37(2):8–10, March/ April 2017. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2017/02/mcg2017020008-
abs.html.

Reina:2023:CID

[Rei23] Guido Reina. Can image data facilitate reproducibility of graphics and visualizations? toward a trusted scientific practice. IEEE Computer Graphics and Applications, 43(2):89–100, March/ April 2023. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Rensink:2015:PUV

[Ren15] Ronald A. Rensink. Preparing undergraduates for visual analytics. IEEE Computer Graphics and Applications, 35(2):16–20, March/ April 2015. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
cg/2015/02/mcg2015020016-
abs.html.

Robertson:1994:RIF

[RET^{+}94] P. K. Robertson, R. A.

Earnshaw, D. Thalmann, M. Grave, J. Gallop, and E. M. De Jong. Research issues in the foundations of visualization. IEEE Computer Graphics and Applications, 14(2):73–76, March/ April 1994. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Rave:2022:MVA

[RFA^{+}22] Hennes Rave, Johannes Fincke, Steffen Averkamp, Beate Tangerding, Luca P. Wehrenberg, Tim Gerrits, Karim Huesmann, Simon Leistikow, and Lars Linsen. Multifaceted visual analysis of oceanographic simulation ensemble data. IEEE Computer Graphics and Applications, 42(4):80–88, July/ August 2022. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Reda:2013:VLH

[RFK^{+}13] Khairi Reda, Alessandro Febretti, Aaron Knoll, Jillian Aurisano, Jason Leigh, Andrew Johnson, Michael E. Papka, and Mark Hereld. Visualizing large, heterogeneous data in hybrid-reality environments. IEEE Computer Graphics and Applications, 33(4): 38–48, July/August 2013. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Rost:1989:PNT

[RFN89] Randi J. Rost, Jeffrey D. Friedberg, and Peter L. Nishimoto. PEX: a networktransparent 3D graphics system. IEEE Computer Graphics and Applications, 9(4):14–26, July/August 1989. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Rigau:2008:IAM

[RFS08] Jaume Rigau, Miquel Feixas, and Mateu Sbert. Informational aesthetics measures. IEEE Computer Graphics and Applications, 28(2): 24–34, March/April 2008. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Roth:1995:VVA

[RG95] Mitchell Roth and Rick Guritz. Visualization of volcanic ash clouds. IEEE Computer Graphics and Applications, 15(4):34–39, July/August 1995. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Rhodes:1983:CGI

[RGA^{+}83] Michael L. Rhodes, William V.

Glenn, Jr., Yu-Ming Azzawi, Robert S. Howland, and Don L. Hibbard. Computer graphics and an interactive stereotactic system for CT-aided neurosurgery. IEEE Computer Graphics and Applications, 3(5):31– 38, August 1983. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Royan:2007:NBV

[RGCB07] Jérôme Royan, Patrick Gioia, Romain Cavagna, and Christian Bouville. Network-based visualization of 3D landscapes and city models. IEEE Computer Graphics and Applications, 27(6):70–79, November/December 2007. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Rimensberger:2019:VCA

[RGG19] Noël Rimensberger, Markus Gross, and Tobias Günther. Visualization of clouds and atmospheric air flows. IEEE Computer Graphics and Applications, 39(1):12–25, January/February 2019. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL https:

                                        //ieeexplore.ieee.org/
                                        document/8663522/.

Ren:2006:IID

[RGL^{+}06] Pin Ren, Yan Gao, Zhichun Li, Yan Chen, and Benjamin Watson. IDGraphs: Intrusion detection and analysis using stream compositing. IEEE Computer Graphics and Applications, 26(2): 28–39, March/April 2006. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Raya:2021:VRA

[RGRLFM21] L. Raya, J. J. García-Rueda, D. López-Fernández, and J. Mayor. Virtual reality application for fostering interest in art. IEEE Computer Graphics and Applications, 41(2):106–113, March/April 2021. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Rasche:2005:DPR

[RGW05] Karl Rasche, Robert Geist, and James Westall. Detail preserving reproduction of color images for monochromats and dichromats. IEEE Computer Graphics and Applications, 25(3):22–30, May/June 2005. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Rushmeier:1995:VRP

[RHC95] Holly Rushmeier, Anthony Hamins, and Mun Young Choi. Volume rendering of pool fire data. IEEE Computer Graphics and Applications, 15(4):62–67, July/ August 1995. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Rheingans:2002:WTY

[Rhe02]  Penny Rheingans.  Are

we there yet? exploring with dynamic visualization. IEEE Computer Graphics and Applications, 22(1):6–10, January/ February 2002. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://dlib.
computer.org/cg/books/
cg2002/g1006abs.htm;
http://dlib.computer.org/
cg/books/cg2002/pdf/g1006.
pdf.

Rhodes:1990:CGM

[Rho90]  Michael Rhodes.  Com-

puter graphics in medicine. IEEE Computer Graphics and Applications, 10(2): 20–23, March/April 1990. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Rhodes:1991:CGM

[Rho91] Michael L. Rhodes. Computer graphics in medicine: The past decade. IEEE Computer Graphics and Applications, 11(1):52–54, January/February 1991. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Rhodes:1993:CGM

[Rho93] Michael L. Rhodes. Computer graphics in medicine: Guest Editor’s introduction. IEEE Computer Graphics and Applications, 13(6):49, November/December 1993. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Rhodes:1996:ECA

[Rho96] Michael L. Rhodes. An electronic CG&A: computer graphics education. IEEE Computer Graphics and Applications, 16(3):17–18, May/June 1996. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Rhodes:1997:CGM

[Rho97a] Michael L. Rhodes. Computer graphics and medicine: a complex partnership. IEEE Computer Graphics and Applications, 17(1):22– 28, January/February 1997. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Rhodes:1997:GEI

[Rho97b] Michael L. Rhodes. Guest Editor’s introduction: Simulation and modeling. IEEE Computer Graphics and Applications, 17(5):15, September/October 1997. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://dlib.
computer.org/cg/books/
cg1997/pdf/g5015.pdf.

Rhyne:2000:SVN

[Rhy00] Theresa-Marie Rhyne. Scientific visualization in the next millennium. IEEE Computer Graphics and Applications, 20(1):20–21, January/February 2000. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http:

//dlib.computer.org/cg/
books/cg2000/pdf/g1020.
pdf.

Rhyne:2003:DDB

[Rhy03] Theresa-Marie Rhyne. Does the difference between information and scientific visualization really matter? IEEE Computer Graphics and Applications, 23(3):6–8, May/ June 2003. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://dlib.
computer.org/cg/books/
cg2003/pdf/g3006.pdf.

Rhyne:2011:EVT

[Rhy11] Theresa-Marie Rhyne. Exploring visualization theory. IEEE Computer Graphics and Applications, 31(3):6–7, May/June 2011. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Ribarsky:1998:TTC

[Rib98]  William Ribarsky.  The

times they are a-changing: PC graphics moves in. IEEE Computer Graphics and Applications, 18(3):20–25, May/June 1998. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://dlib.
computer.org/cg/books/
cg1998/pdf/g3020.pdf.

Richards:1985:EES

[Ric85] Larry G. Richards. Engineering education: a status report on the CAD/CAM revolution. IEEE Computer Graphics and Applications, 5(2):19–25, February 1985. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Riche:2015:DDD

[Ric15] Nathalie Henry Riche. Datadriven discoveries: Pushing visualization research further. IEEE Computer Graphics and Applications, 35(3):42–43, May/ June 2015. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
cg/2015/03/mcg2015030042.
html.

Riesenfeld:1981:HCP

[Rie81] Richard F. Riesenfeld. Homogeneous coordinates and projective planes in computer graphics. IEEE Computer Graphics and Applications, 1(1):50–55, January/ February 1981. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Riekert:1993:EAO

[Rie93] Wolf-Fritz Riekert. Extracting area objects from raster image data. IEEE Computer Graphics and Applications, 13(2):68–73, March/ April 1993. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Riesenfeld:2022:DBS

[Rie22] Richard F. Riesenfeld. The development of B-splines for CAD. IEEE Computer Graphics and Applications, 42(2):90–100, March/ April 2022. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Ribarsky:1999:DVU

[RKJH99] William Ribarsky, Jochen Katz, Frank Jiang, and Aubrey Holland. Discovery visualization using fast clustering. IEEE Computer Graphics and Applications, 19(5):32–39, September/October 1999. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

                                        computer.org/cga/cg1999/
                                        g5032abs.htm;   http:/
                                        /dlib.computer.org/cg/
                                        books/cg1999/pdf/g5032.
                                        pdf.

Rhodes:1987:ACG

[RKRW87] Michael L. Rhodes, Yu-Ming Kuo, Stephen L. G. Rothman, and Charles Woznick. An application of computer graphics and networks to anatomic model and prosthesis manufacturing. IEEE Computer Graphics and Applications, 7(2):12– 25, February 1987. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Raghuvanshi:2007:PBS

[RL07] Nikunj Raghuvanshi and Ming C. Lin. Physically based sound synthesis for large-scale virtual environments. IEEE Computer Graphics and Applications, 27(1):14–18, January/ February 2007. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Rey:2024:DLT

[RLCI24] Bradley Rey, Bongshin Lee, Eun Kyoung Choe, and Pourang Irani. Databiting: Lightweight, transient, and insight rich exploration of personal data. IEEE Computer Graphics and Applications, 44(2):65–72, March/ April 2024. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Ribo:2002:HTO

[RLG^{+}02] Miguel Ribo, Peter Lang, Harald Ganster, Markus Brandner, Christoph Stock, and Axel Pinz. Hybrid tracking for outdoor augmented reality applications. IEEE Computer Graphics and Applications, 22(6):54–63, November/ December 2002. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://dlib.
computer.org/cg/books/
cg2002/pdf/g6054.pdf;
http://www.computer.org/
cga/cg2002/g6054abs.htm.

Reddy:1999:TIV

[RLIB99] Martin Reddy, Yvan Leclerc, Lee Iverson, and Nat Bletter. TerraVision II: Visualizing massive terrain databases in VRML. IEEE Computer Graphics and Applications, 19(2):30–38, March/April 1999. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

computer.org/cga/cg1999/
g2030abs.htm;   http:/
/dlib.computer.org/cg/
books/cg1999/pdf/g2030.
pdf.

Ren:2013:FGN

[RLOW13] Gang Ren, Chuan Li, Eamonn O’Neill, and Phil Willis. 3D freehand gestural navigation for interactive public displays. IEEE Computer Graphics and Applications, 33(2):47–55, March/ April 2013. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Rubel:2016:WSV

[RLV^{+}16] Oliver Rubel, Burlen Loring, Jean-Luc Vay, David P. Grote, Remi Lehe, Stepan Bulanov, Henri Vincenti, and E. Wes Bethel. WarpIV: In situ visualization and analysis of ion accelerator simulations. IEEE Computer Graphics and Applications, 36(3):22–35, May/ June 2016. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2016/03/mcg2016030022-
abs.html.

Raczkowsky:1989:SCR

[RM89] J. Raczkowsky and K. H.

Mittenbuehler. Simulation of cameras in robot applications. IEEE Computer Graphics and Applications, 9(1):16–25, January/ February 1989. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Ruprecht:1995:IWS

[RM95] Detlef Ruprecht and Heinrich Müller. Image warning with scattered data interpolation. IEEE Computer Graphics and Applications, 15(2):37–43, March/ April 1995. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Rhyne:2006:GEI

[RMD06] Theresa-Marie Rhyne, Alan MacEachren, and Jason Dykes. Guest Editors’ introduction: Exploring geovisualization. IEEE Computer Graphics and Applications, 26(4):20–21, July/ August 2006. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL http://csdl.
                                        computer.org/comp/mags/
                                        cg/2006/04/g4020.pdf.

Risch:1996:VEM

[RMDT96] John S. Risch, Richard A. May, Scott T. Dowson, and James J. Thomas. A virtual environment for multimedia intelligence data analysis. IEEE Computer Graphics and Applications, 16(6): 33–41, November/December 1996. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Reina:2014:IMO

[RME14] Guido Reina, Thomas Muller, and Thomas Ertl. Incorporating modern OpenGL into computer graphics education. IEEE Computer Graphics and Applications, 34(4):16–21, July/August 2014. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Raposo:2016:VCP

[RMG16] Alberto Raposo, Soraia Raupp Musse, and James Gain. Visual computing and the progress of developing countries. IEEE Computer Graphics and Applications, 36(2):22–23, March/April 2016. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL

https://www.computer.org/
csdl/mags/cg/2016/02/mcg2016020022.
html.

Robinson:1991:GWE

[RMM^{+}91] Ken Robinson, Mark Martin, Max Mehl, Allan Davison, Kieron Drake, and Mel Slater. Graphics workstations: A European perspective. IEEE Computer Graphics and Applications, 11(2):91–103, March/April 1991. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Raj:2016:ESA

[RMP^{+}16] Mukund Raj, Mahsa Mirzargar, J. Samuel Preston, Robert M. Kirby, and Ross T. Whitaker. Evaluating shape alignment via ensemble visualization. IEEE Computer Graphics and Applications, 36(3):60–71, May/ June 2016. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/

cg/2016/03/mcg2016030060-
abs.html.

Rojo:2019:VRT

[RMR^{+}19] Diego Rojo, Jesús Mayor, José Jesús García Rueda, Laura Raya, and Mike Potel. A virtual reality training application for adults with Asperger’s Syndrome. IEEE Computer Graphics and Applications, 39(2):104– 111, March/April 2019. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL https:

//ieeexplore.ieee.org/
document/8673005/.

Rossignac:1994:RIM

[RN94] J. R. Rossignac and M. Novak. Research issues in model-based visualization of complex data sets. IEEE Computer Graphics and Applications, 14(2):83–85, March/April 1994. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Rincon-Nigro:2014:GAI

[RNNTD14] Mario Rincon-Nigro, Nikhil V. Navkar, Nikolaos V. Tsekos, and Zhigang Deng. GPUaccelerated interactive visualization and planning of neurosurgical interventions. IEEE Computer Graphics and Applications, 34(1):22– 31, January/February 2014. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Robertson:1986:GCS

[RO86] Philip K. Robertson and John F. O’Callaghan. The generation of color sequences for univariate and bivariate mapping. IEEE Computer Graphics and Applications, 6(2):24–32, February 1986. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Renze:1996:GUD

[RO96] Kevin J. Renze and James H.

Oliver. Generalized unstructured decimation. IEEE Computer Graphics and Applications, 16(6):24–32, November/December 1996. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Robertson:1987:FPV

[Rob87] Philip K. Robertson. Fast perspective views of images using one-dimensional operations. IEEE Computer Graphics and Applications, 7(2):47–56, February 1987. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Robertson:1988:VCG

[Rob88] Philip K. Robertson. Visualizing color gamuts: a user interface for the effective use of perceptual color spaces in data displays. IEEE Computer Graphics and Applications, 8(5):50–64, September/October 1988. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Robertson:1989:STR

[Rob89] Philip K. Robertson. Spatial transformations for rapid scan-line surface shadowing. IEEE Computer Graphics and Applications, 9(2): 30–38, March/April 1989. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Robertson:1991:MCD

[Rob91] Philip K. Robertson. A methodology for choosing data representations. IEEE Computer Graphics and Applications, 11(3):56–67, May/June 1991. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Rockwood:1987:GST

[Roc87] Alyn P. Rockwood. A generalized scanning technique for display of parametrically defined surfaces. IEEE Computer Graphics and Applications, 7(8):15–26, August 1987. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Rodden:2014:ASV

[Rod14] Kerry Rodden. Applying a sunburst visualization to summarize user navigation sequences. IEEE Computer Graphics and Applications, 34(5):36–40, September/October 2014. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://csdl.
computer.org/csdl/mags/
cg/2014/05/mcg2014050036-
abs.html.

Rokita:1996:GDF

[Rok96] Przemyslaw Rokita. Generating depth-of-field effects in virtual reality applications. IEEE Computer Graphics and Applications, 16(2): 18–21, March/April 1996. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Rosenblum:1989:VOD

[Ros89a] Lawrence J. Rosenblum.

Visualizing oceanographic data. IEEE Computer Graphics and Applications, 9(3):14–19, May/June 1989. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Ross:1989:MGT

[Ros89b] Robert Ross. The map geometric transformation. IEEE Computer Graphics and Applications, 9(3):70– 75, May/June 1989. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Rosenblum:1994:RIS

[Ros94] Lawrence J. Rosenblum. Research issues in scientific visualization. IEEE Computer Graphics and Applications, 14(2):61–63, March/ April 1994. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Rosenblum:1997:PVA

[Ros97] Lawrence J. Rosenblum. Projects in VR: Applications of the Responsive Workbench. IEEE Computer Graphics and Applications, 17(4):10, July/ August 1997. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Rosenblum:2000:VAR

[Ros00a] Lawrence Rosenblum. Virtual and augmented reality 2020. IEEE Computer Graphics and Applications, 20(1):38–39, January/ February 2000. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL http://dlib.
                                        computer.org/cg/books/
                                        cg2000/pdf/g1038.pdf.

Rosendahl:2000:FDG

[Ros00b] Carl Rosendahl. Future directions of graphics in entertainment. IEEE Computer Graphics and Applications, 20(1):30–31, January/ February 2000. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://dlib.
computer.org/cg/books/
cg2000/pdf/g1030.pdf.

Rosenblum:2021:AGR

[Ros21] Lawrence J. Rosenblum.

Adventures of a government researcher. IEEE Computer Graphics and Applications, 41(6):187–197, November/ December 2021. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Rottensteiner:2003:AGH

[Rot03] Franz Rottensteiner. Automatic generation of highquality building models from lidar data. IEEE Computer Graphics and Applications, 23(6):42–50, November/December 2003. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http:

//csdl.computer.org/dl/
mags/cg/2003/06/g6042.
pdf;   http://csdl.computer.
org/dl/mags/cg/2003/06/
g6042abs.htm.

Roulier:1987:CGR

[Rou87]  John A. Roulier.  A

convexity-preserving grid refinement algorithm for interpolation of bivariate functions. IEEE Computer Graphics and Applications, 7(1):57–62, January/ February 1987. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Roussou:2009:PVV

[Rou09] Maria Roussou. Projects in VR: a VR playground for learning abstract mathematics concepts. IEEE Computer Graphics and Applications, 29(1):82–85, January/ February 2009. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Rowland:1995:MFA

[RP95] Duncan A. Rowland and David I. Perrett. Manipulating facial appearance through shape and color. IEEE Computer Graphics and Applications, 15(5):70– 76, September 1995. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Rosen:2011:NAI

[RPHW11] Paul Rosen, Voicu Popescu, Kyle Hayward, and Chris Wyman. Nonpinhole approximations for interactive rendering. IEEE Computer Graphics and Applications, 31(6):68–83, November/December 2011. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Ruhland:2017:DDA

[RPM17] Kerstin Ruhland, Mukta Prasad, and Rachel Mc-Donnell. Data-driven approach to synthesizing facial animation using motion capture. IEEE Computer Graphics and Applications, 37(4):30–41, July/ August 2017. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2017/04/mcg2017040030-
abs.html.

Rajabiyazdi:2021:CPH

[RPOC21] Fateme Rajabiyazdi, Charles Perin, Lora Oehlberg, and Sheelagh Carpendale. Communicating patient health data: a wicked problem. IEEE Computer Graphics and Applications, 41(6):179– 186, November/December 2021. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Reilly:1984:IVD

[RR84] Susan Smith Reilly and John W. Roach. Improved visual design for graphics display. IEEE Computer Graphics and Applications, 4(2):42–51, February 1984. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Rogers:1985:EGL

[RR85] David F. Rogers and Linda M.

Rybak. On an efficient general line-clipping algorithm. IEEE Computer Graphics and Applications, 5(1):82– 86, January/February 1985. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Rogers:1986:RDG

[RR86a] David F. Rogers and Stephen D. Rogers. A raster display graphics package for education. IEEE Computer Graphics and Applications, 6(4):51–58, April 1986. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Rossignac:1986:DDT

[RR86b] J. R. Rossignac and A. A. G. Requicha. Depth-buffering display techniques for constructive solid geometry. IEEE Computer Graphics and Applications, 6(9):29– 39, September 1986. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Requicha:1992:SMB

[RR92] Aristides A. G. Requicha and Jarek R. Rossignac. Solid modeling and beyond. IEEE Computer Graphics and Applications, 12(5):31– ??, September 1992. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Reiser:1995:ACU

[RR95] Hans Reiser and Beverly Reiser. Aesthetic considerations unique to interactive multimedia. IEEE Computer Graphics and Applications, 15(3):24–28, May/ June 1995. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Rhodes:1996:AST

[RR96] Michael L. Rhodes and Douglas D. Robertson. Applications in surgery and therapy: Guest Editors’ introduction. IEEE Computer Graphics and Applications, 16(1):28– 29, January/February 1996. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Ribarsky:2003:GEI

[RR03] Bill Ribarsky and Holly Rushmeier. Guest Editors’ introduction: 3D reconstruction and visualization. IEEE Computer Graphics and Applications, 23(6):20–21, November/December 2003. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http:

//csdl.computer.org/dl/
mags/cg/2003/06/g6020.
pdf.

Roberts:2014:VBD

[RRB^{+}14] Jonathan C. Roberts, Panagiotis D. Ritsos, Sriram Karthik Badam, Dominique Brodbeck, Jessie Kennedy, and Niklas Elmqvist. Visualization beyond the desktop — the next big thing. IEEE Computer Graphics and Applications, 34(6):26–34, November/ December 2014. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://csdl.
computer.org/csdl/mags/
cg/2014/06/mcg2014060026-
abs.html.

Raya:2023:DVR

[RRF^{+}23] Laura Raya, Juan Jesús Ruiz, Marc Fabián, Adrián Ron, Javier García, and Cristina Verdú. Development of a virtual reality tool for the treatment of pediatric patients in the ICU. IEEE Computer Graphics and Applications, 43(2): 69–77, March/April 2023. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Raposo:2016:SIV

[RRR^{+}16] Alberto Raposo, Luciano Reis, Vinicius Rodrigues, Rodrigo Pinheiro, Pablo Elias, and Renato Cherullo. A system for integrated visualization in oil exploration and production. IEEE Computer Graphics and Applications, 36(3):10–16, May/ June 2016. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2016/03/mcg2016030010-
abs.html.

Rodriguez:2015:VBS

[RRV^{+}15] Alejandro Rodriguez, Beatriz Rey, Ma Dolores Vara, Maja Wrzesien, Mariano Alcaniz, Rosa Ma Banos, and David Perez-Lopez. A VR-based serious game for studying emotional regulation in adolescents. IEEE Computer Graphics and Applications, 35(1):65–73, January/February 2015. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

csdl.computer.org/csdl/
mags/cg/2015/01/mcg2015010065-
abs.html.

Rohrer:1997:WIV

[RS97] Randall M. Rohrer and Edward Swing. Web-based information visualization. IEEE Computer Graphics and Applications, 17(4): 52–59, July/August 1997. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Robertson:1999:GEI

[RS99] Phil Robertson and Jürgen Shönhut. Guest Editors’ introduction: Color in computer graphics. IEEE Computer Graphics and Applications, 19(4):18–19, July/ August 1999. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL http://dlib.
                                        computer.org/cg/books/
                                        cg1999/pdf/g4018.pdf.

Riener:2014:HPB

[RS14] Andreas Riener and Andreas Sippl. Head-pose-based attention recognition on large public displays. IEEE Computer Graphics and Applications, 34(1):32– 41, January/February 2014. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Ritter:2001:VPN

[RSDP01] Flex Ritter, Thomas Strothotte, Oliver Dresden, and Bernhard Preim. Virtual 3D puzzles: a new method for exploring geometric models in VR. IEEE Computer Graphics and Applications, 21(5):11–13, September/October 2001. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL http://dlib.
                                        computer.org/cg/books/
                                        cg2001/g5011abs.htm;
                                        http://dlib.computer.org/
                                        cg/books/cg2001/pdf/g5011.
                                        pdf.

Rohrer:1999:SBV

[RSE99] Randall M. Rohrer, John L. Sibert, and David S. Ebert. A shape-based visual interface for text retrieval.

IEEE Computer Graphics and Applications, 19(5): 40–47, September/October 1999. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL

http://computer.org/cga/
cg1999/g5040abs.htm;
http://dlib.computer.org/
cg/books/cg1999/pdf/g5040.
pdf.

Riley:2006:VSN

[RSK^{+}06] Kirk Riley, Yuyan Song, Martin Kraus, David S. Ebert, and Jason J. Levit. Visualization of structured nonuniform grids. IEEE Computer Graphics and Applications, 26(1):46–55, January/February 2006. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Ryan:2019:TDV

[RSL^{+}19] Lindy Ryan, Deborah Silver, Robert S. Laramee, David Ebert, and Theresa-Marie Rhyne. Teaching data visualization as a skill. IEEE Computer Graphics and Applications, 39(2):95– 103, March/April 2019. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL https:

//ieeexplore.ieee.org/
document/8673006/.

Rakkolainen:2015:MUI

[RSP15] Ismo Rakkolainen, Antti Sand, and Karri Palovuori.

Midair user interfaces employing particle screens. IEEE Computer Graphics and Applications, 35(2):96– 102, March/April 2015. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http:/

                                        /www.computer.org/csdl/
                                        mags/cg/2015/02/mcg2015020096-
                                        abs.html.

Rogers:1983:SBS

[RSR83a] D. F. Rogers, S. G. Satterfield, and F. A. Rodriguez. Shiphulls, B-spline surfaces, and CAD/CAM. IEEE Computer Graphics and Applications, 3(12):37–45, December 1983. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Rogers:1983:SHB

[RSR83b] David F. Rogers, Steven G. Satterfield, and Francisco A. Rodriguez. Ship hulls, Bspline surfaces, and CAD/ CAM. IEEE Computer Graphics and Applications, 3(9):37–45, December 1983. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Raposo:2009:EIV

[RSS^{+}09] Alberto Raposo, Ismael Santos, Luciano Soares, Gustavo Wagner, Eduardo Corseuil, and Marcelo Gattass. Environ: Integrating VR and CAD in engineering projects. IEEE Computer Graphics and Applications, 29(6):91–95, November/December 2009. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Rosenblum:1996:GEI

[RSZ96] Lawrence Rosenblum, Sharon Stansfield, and Michael Zyda. Guest Editor’s introduction: On apples, oranges, and the interdisciplinary nature of VR. IEEE Computer Graphics and Applications, 16(6):20–22, November/December 1996. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Rushmeier:2020:ACH

[RSZ20] H. Rushmeier, F. Samsel, and J. Zhang. Art and cultural heritage. IEEE Computer Graphics and Applications, 40(3):17–18, May/ June 2020. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Ronfard:2007:IC

[RT07] Rémi Ronfard and Gabriel Taubin. Introducing 3D cinematography. IEEE Computer Graphics and Applications, 27(3):18–20, May/ June 2007. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Raskar:2005:HRW

[RTF^{+}05] Ramesh Raskar, Kar-Han Tan, Rogerio S. Feris, Matthew Turk, and James Kobler Jingyi Yu. Harnessing real-world depth edges with multiflash imaging. IEEE Computer Graphics and Applications, 25(1):32– 38, January/February 2005. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Rushmeier:2011:ISI

[Rus11] Holly Rushmeier. Introduction to the special issue: Camera culture. IEEE Computer Graphics and Applications, 31(1):7–8, January/ February 2011. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Requicha:1982:SMH

[RV82] A. A. G. Requicha and H. B. Voelcker. Solid modeling: a historical summary and contemporary assessment. IEEE Computer Graphics and Applications, 2(2):9– 16, 18–20, 22–24, March/ April 1982. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Requicha:1983:SMC

[RV83] A. A. G. Requicha and H. B. Voelcker. Solid modeling: Current status and research directions. IEEE Computer Graphics and Applications, 3(7):25–30, 32– 37, October 1983. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Reiners:2007:CLP

[RV07] Dirk Reiners and Gerrit Voss. Climbing longs peak: The steep road to the future of OpenGL. IEEE Computer Graphics and Applications, 27(4):84–89, July/ August 2007. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Rossignac:2010:RFS

[RV10] Jarek Rossignac and Abhishek Venkatesh. Ringing: Frugal subdivision of curves and surfaces. IEEE Computer Graphics and Applications, 30(2):22–33, March/ April 2010. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Roerdink:2021:MW

[RvW21] Jos Roerdink and Jarke van Wijk. Michel Westenberg (1973–2021). IEEE Computer Graphics and Applications, 41(6):198–199, November/December 2021. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Russ:1999:LSM

[RW99] Keith Russ and Andrew Wetherelt. Large-scale mine visualization using VRML. IEEE Computer Graphics and Applications, 19(2):39– 44, March/April 1999. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

computer.org/cga/cg1999/
g2039abs.htm;   http:/
/dlib.computer.org/cg/
books/cg1999/pdf/g2039.
pdf.

Repenning:2014:BMF

[RWB^{+}14] Alexander Repenning, David C. Webb, Catharine Brand, Fred Gluck, Ryan Grover, Susan Miller, Hilarie Nickerson, and Muyang Song. Beyond Minecraft: Facilitating computational thinking through modeling and programming in 3D. IEEE Computer Graphics and Applications, 34(3):68–71, May/June 2014. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Ribarsky:2002:PVU

[RWF02] William Ribarsky, Tony Wasilewski, and Nickolas Faust. Projects in VR: From urban terrain models to visible cities. IEEE Computer Graphics and Applications, 22(4):11–15, July/ August 2002. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

computer.org/cga/cg2002/
g4010abs.htm;   http:/
/dlib.computer.org/cg/
books/cg2002/pdf/g4010.
pdf.

Read:2009:APD

[RXB09] Huw Read, Konstantinos Xynos, and Andrew Blyth. Applications: Presenting DEViSE: Data exchange for visualizing security events. IEEE Computer Graphics and Applications, 29(3):6– 11, May/June 2009. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Riedl:2006:LSG

[RY06] Mark O. Riedl and R. Michael Young. From linear story generation to branching story graphs. IEEE Computer Graphics and Applications, 26(3):23–31, May/ June 2006. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Rosenfeld:2004:POB

[RZSP04] Dan Rosenfeld, Michael Zawadzki, Jeremi Sudol, and Ken Perlin. Physical objects as bidirectional user interface elements. IEEE Computer Graphics and Applications, 24(1):44–49, January/ February 2004. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL http://csdl.
                                        computer.org/comp/mags/
                                        cg/2004/01/g1044abs.htm;
                                         http://csdl.computer.
                                        org/dl/mags/cg/2004/01/
                                        g1044.pdf.

Sederberg:1985:SSP

[SA85] Thomas W. Sederberg and D. C. Anderson. Steiner surface patches. IEEE Computer Graphics and Applications, 5(5):23–36, May/ June 1985. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Schwarz:2023:LWR

[SAB^{+}23] Chris Schwarz, Omar Ahmad, Timothy Brown, John Gaspar, Gregory Wagner, and Daniel V. McGehee. The long and winding road: 25 years of the National Advanced Driving Simulator. IEEE Computer Graphics and Applications, 43(4): 121–128, July/August 2023. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Sadagic:2016:DCV

[Sad16]  Amela Sadagic.  Design

and choice of visual display solutions in the training domain. IEEE Computer Graphics and Applications, 36(6):18–25, November/ December 2016. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2016/06/mcg2016060018-
abs.html.

Schleich:2002:UNC

[SADC02] Jean-Marc Schleich, Claude Almange, Jean-Louis Dillenseger, and Jean-Louis Coatrieux. Understanding normal cardiac development using animated models. IEEE Computer Graphics and Applications, 22(1):14–19, January/ February 2002. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://dlib.
computer.org/cg/books/
cg2002/g1014abs.htm;
http://dlib.computer.org/
cg/books/cg2002/pdf/g1014.
pdf.

Shah:2023:WWB

[SADH23] Sachin Shah, Kamran Ali, Lisa Dieker, and Charles Hughes. WAVE: a Webbased platform for delivering knowledge-driven virtual experiences. IEEE Computer Graphics and Applications, 43(3):54–60, May/ June 2023. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Schroeder:2000:VVT

[SAH00] William J. Schroeder, Lisa S.

Avila, and William Hoffman. Visualizing with VTK: a tutorial. IEEE Computer Graphics and Applications, 20(5):20–27, September/October 2000. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL http://dlib.
                                        computer.org/cg/books/
                                        cg2000/pdf/g5020.pdf.

Suzuki:2013:VTU

[SAHK13] Taro Suzuki, Yoshiharu Amano, Takumi Hashizume, and Nobuaki Kubo. Vehicle teleoperation using 3D maps and GPS time synchronization. IEEE Computer Graphics and Applications, 33(5):82–88, September/October 2013. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Sandbank:2001:DTC

[San01] Charles P. Sandbank. Digital TV in the convergent environment. IEEE Computer Graphics and Applications, 21(1):32–36, January/ February 2001. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

                                        computer.org/cga/cg2001/
                                        g1032abs.htm;   http:/
                                        /dlib.computer.org/cg/
                                        books/cg2001/pdf/g1032.
                                        pdf.

Sobierajski:1995:VCA

[SAO^{+}95] Lisa M. Sobierajski, Ricardo S. Avila, Donald M.

O’Malley, Sidney Wang, and Arie E. Kaufman. Visualization of calcium activity in nerve cells. IEEE Computer Graphics and Applications, 15(4):55–61, July/ August 1995. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Sarraga:1982:CSA

[Sar82] Ramon F. Sarraga. Computation of surface areas in GMSOLID. IEEE Computer Graphics and Applications, 2(7):65–70, September 1982. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Sarraga:1990:CMS

[Sar90] Ramon F. Sarraga. Computer modeling of surfaces with arbitrary shapes. IEEE Computer Graphics and Applications, 10(2):67–77, March/April 1990. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Sawzin:1983:DCG

[Saw83] Stephen A. Sawzin. The design of a computer graphics training program. IEEE Computer Graphics and Applications, 3(8):53–57, November/December 1983. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Sawyer:2008:GSC

[Saw08] Ben Sawyer. Graphically speaking: From cells to Cell Processors: The integration of health and video games. IEEE Computer Graphics and Applications, 28(6): 83–85, November/December 2008. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Shectman:1981:EED

[SB81] Robert M. Shectman and Doug E. Bering. The electronics engineer’s design station. IEEE Computer Graphics and Applications, 1(2):15–23, April 1981. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Salesin:1986:TBG

[SB86a] David Salesin and Ronen Barzel. Two-bit graphics (alpha bit mapping). IEEE Computer Graphics and Applications, 6(6):36– 42, June 1986. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Sweeney:1986:RTF

[SB86b] Michael Sweeney and Richard H. Bartels. Ray tracing free-form B-spline surfaces. IEEE Computer Graphics and Applications, 6(2):41– 49, February 1986. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Schwarz:1997:ODG

[SB97] Adam J. Schwarz and Barry G. Blundell. Optimizing dot graphics for volumetric displays. IEEE Computer Graphics and Applications, 17(3):72–78, May/ June 1997. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Simon:2002:PEP

[SB02] Gilles Simon and Marie-Odile Berger. Pose estimation for planar structures. IEEE Computer Graphics and Applications, 22(6):46–53, November/ December 2002. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://dlib.
computer.org/cg/books/
cg2002/pdf/g6046.pdf;
http://www.computer.org/
cga/cg2002/g6046abs.htm.

Shalf:2003:GFV

[SB03a] John Shalf and E. Wes Bethel. The Grid and future visualization system architectures. IEEE Computer Graphics and Applications, 23(2):6–9, March/ April 2003. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://dlib.

computer.org/cg/books/
cg2003/pdf/g2006.pdf.

Suffern:2003:RII

[SB03b] Kevin G. Suffern and Ronald J. Balsys. Rendering the intersections of implicit surfaces. IEEE Computer Graphics and Applications, 23(5):70–77, September/October 2003. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http:

//csdl.computer.org/dl/
mags/cg/2003/05/g5070.
pdf;   http://csdl.computer.
org/dl/mags/cg/2003/05/
g5070abs.htm.

Sakas:2006:GEI

[SB06] Georgios Sakas and Ulrich Bockholt. Guest Editors’ introduction: Simulators and closed interaction loops. IEEE Computer Graphics and Applications, 26(6):22–23, November/December 2006. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

csdl.computer.org/comp/
mags/cg/2006/06/g6022.
pdf.

Steinicke:2013:UPI

[SB13] Frank Steinicke and Gerd Bruder. Using perceptual illusions for redirected walking. IEEE Computer Graphics and Applications, 33 (1):6–11, January/February 2013. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Schultz:2016:ILD

[SB16] Chris Schultz and Mike Bailey. Interacting with large 3D datasets on a mobile device. IEEE Computer Graphics and Applications, 36(5):19–23, September/October 2016. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2016/05/mcg2016050019-
abs.html.

Shellshear:2015:MSF

[SBC15] Evan Shellshear, Rolf Berlin, and Johan S. Carlson. Maximizing smart factory systems by incrementally updating point clouds. IEEE Computer Graphics and Applications, 35(2):62–69, March/April 2015. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
cg/2015/02/mcg2015020062-
abs.html.

Saffer:2004:VAP

[SBCvdS04] Jeffrey D. Saffer, Vicki L.

Burnett, Guang Chen, and Peter van der Spek. Visual analytics in the pharmaceutical industry. IEEE Computer Graphics and Applications, 24(5):10–15, September/October 2004. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http:

                                        //csdl.computer.org/dl/
                                        mags/cg/2004/05/g5010.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/cg/2004/05/
                                        g5010.pdf.

Scherzinger:2017:IEC

[SBD^{+}17] Aaron Scherzinger, Tobias Brix, Dominik Drees, Andreas Volker, Kiril Radkov, Niko Santalidis, Alexander Fieguth, and Klaus H. Hinrichs. Interactive exploration of cosmological dark-matter simulation data. IEEE Computer Graphics and Applications, 37(2):80–89, March/ April 2017. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        cg/2017/02/mcg2017020080-
                                        abs.html.

Stucki:1995:GEI

[SBE95] Peter Stucki, Jack Bresenham, and Rae Earnshaw. Guest Editors’ introduction: Computer graphics in RP technology. IEEE Computer Graphics and Applications, 15(6):17–19, November/December 1995. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Subramanian:1997:VBM

[SBHM97] Kalpathi R. Subramanian, David P. Bashor, William V. Hasty, and Sandra M. Merkel. Visualization blackboard: Multilevel visualization of spinal reflex circuit simulations. IEEE Computer Graphics and Applications, 17(3):6–10, May/ June 1997. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Stytz:2001:VS

[SBJ^{+}01] Martin R. Stytz, Sheila B.

Banks, Troy D. Johnson, John M. Lewis, and Scott A. Rothermel. The Virtual SpacePlane. IEEE Computer Graphics and Applications, 21(2):42–52, March/ April 2001. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

computer.org/cga/cg2001/
g2042abs.htm;   http:/
/dlib.computer.org/cg/
books/cg2001/pdf/g2042.
pdf.

Shuey:1986:PSD

[SBM86] David Shuey, David Bailey, and Thomas P. Morrissey. PHIGS: a standard, dynamic, interactive graphics interface. IEEE Computer Graphics and Applications, 6(8):50–57, August 1986. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Schmitz:2020:HFP

[SBMK20] P. Schmitz, T. Blut, C. Mattes, and L. Kobbelt. Highfidelity point-based rendering of large-scale 3-D scan datasets. IEEE Computer Graphics and Applications, 40(3):19–31, May/ June 2020. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Satylmys:2017:MLD

[SBRCD17] Pynar Satylmys, Thomas Bashford-Rogers, Alan Chalmers, and Kurt Debattista. A machine-learning-driven sky model. IEEE Computer Graphics and Applications, 37(1):80–91, January/February 2017. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL

                                        https://www.computer.org/
                                        csdl/mags/cg/2017/01/mcg2017010080-
                                        abs.html.

Stoiber:2010:MST

[SBS10] Nicolas Stoiber, Gaspard Breton, and Renaud Seguier. Modeling short-term dynamics and variability for realistic interactive facial animation. IEEE Computer Graphics and Applications, 30(4):51–61, July/ August 2010. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Stytz:1996:SBT

[SBSW96] Martin R. Stytz, Elizabeth G. Block, Brian B. Soltz, and Kirk Wilson. The synthetic BattleBridge: a tool for large-scale VEs. IEEE Computer Graphics and Applications, 16(1):16– 26, January/February 1996. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Sabella:1989:OAS

[SC89] Paolo Sabella and Ingrid Carlbom. An objectoriented approach to the solid modeling of empirical data. IEEE Computer Graphics and Applications, 9(5):24–35, September 1989. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Scott:2006:GEI

[SC06] Stacey D. Scott and Sheelagh Carpendale. Guest Editors’ introduction: Interacting with digital tabletops. IEEE Computer Graphics and Applications, 26(5):24–27, September/ October 2006. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cg/2006/05/g5024.pdf.

Sohn:2012:SSS

[SC12] Eisung Sohn and Yoon-Chul Choy. Sketch-n-Stretch:

Sketching animations using cutouts. IEEE Computer Graphics and Applications, 32(3):59–69, May/ June 2012. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Shelke:2021:ARV

[SC21] Yogesh Shelke and Chinmay Chakraborty. Augmented reality and virtual reality transforming spinal imaging landscape: a feasibility study. IEEE Computer Graphics and Applications, 41(3):124–138, May/ June 2021. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Salisbury:2004:HRI

[SCB04a] Kenneth Salisbury, Francois Conti, and Federico Barbagli. Haptic rendering: Introductory concepts. IEEE Computer Graphics and Applications, 24(2):24– 32, March/April 2004. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

csdl.computer.org/comp/
mags/cg/2004/02/g2024abs.
htm;   http://csdl.computer.
org/dl/mags/cg/2004/02/
g2024.htm;   http://csdl.
computer.org/dl/mags/cg/
2004/02/g2024.pdf.

Schmidt:2004:MVR

[SCB^{+}04b] Greg S. Schmidt, Sue-Ling Chen, Aaron N. Bryden, Mark A. Livingston, Lawrence J. Rosenblum, and Bryan R. Osborn. Multidimensional visual representations for underwater environmental uncertainty. IEEE Computer Graphics and Applications, 24(5):56–65, September/ October 2004. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http:

//csdl.computer.org/dl/
mags/cg/2004/05/g5056.
htm;   http://csdl.computer.
org/dl/mags/cg/2004/05/
g5056.pdf.

Schachter:1981:CIG

[Sch81] Bruce J. Schachter. Computer image generation for flight simulation. IEEE Computer Graphics and Applications, 1(4):29–32, 34– 38, October 1981. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Schonhut:1986:PGN

[Sch86] J. Schönhut. Are PHIGS and GKS necessarily incompatible? IEEE Computer Graphics and Applications, 6(7):51–3, July/ August 1986. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Schuette:1991:VBA

[Sch91] Lawrence C. Schuette. Visualization blackboard — acoustic holography. IEEE Computer Graphics and Applications, 11(4):12–13, July/August 1991. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Schichtel:1993:BSF

[Sch93a] Markus Schichtel. G^{2} blend surfaces and filling of Nsided holes. IEEE Computer Graphics and Applications, 13(5):68–73, September 1993. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Schroeder:1993:VBV

[Sch93b] Florian Schroeder. Visualization blackboard: Visualizing meteorological data for a lay audience. IEEE Computer Graphics and Applications, 13(5):12–14, September 1993. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Schumaker:1993:TC

[Sch93c] Larry L. Schumaker. Triangulations in CAGD. IEEE Computer Graphics and Applications, 13(1):47–52, January/February 1993. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Schreiner:1999:NVI

[Sch99] Karl Schreiner. Night vision: Infrared takes to the road. IEEE Computer Graphics and Applications, 19(5):6–10, September/October 1999. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://dlib.
computer.org/cg/books/
cg1999/pdf/g5006.pdf.

Schubert:2000:UFS

[Sch00] Richard Schubert. Using a flatbed scanner as a stereoscopic near-field camera. IEEE Computer Graphics and Applications, 20(2):38– 45, March/April 2000. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

computer.org/cga/cg2000/
g2038abs.htm;   http:/
/dlib.computer.org/cg/
books/cg2000/pdf/g2038.
pdf.

Scott:2005:SBM

[SCH05] Stacey D. Scott, M. Sheelagh T. Carpendale, and Stefan Habelski. Storage bins: Mobile storage for collaborative tabletop displays. IEEE Computer Graphics and Applications, 25(4): 58–65, July/August 2005. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Schreiner:2007:WWM

[Sch07] Keri Schreiner. Where we at? mobile phones bring GPS to the masses. IEEE Computer Graphics and Applications, 27(3):6–11, May/ June 2007. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Schreiner:2008:AUP

[Sch08a] Keri Schreiner. Applications: Uniting the paper and digital worlds. IEEE Computer Graphics and Applications, 28(6):6–10, November/December 2008. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Schreiner:2008:DGT

[Sch08b]  Keri Schreiner.  Digital

games target social change. IEEE Computer Graphics and Applications, 28(1):12– 17, January/February 2008. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Schulz:2011:TNT

[Sch11] Hans-Jorg Schulz. Treevis.net: a tree visualization reference. IEEE Computer Graphics and Applications, 31(6):11–15, November/ December 2011. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Schreck:2017:WFC

[Sch17] Tobias Schreck. What features can tell us about shape. IEEE Computer Graphics and Applications, 37(3):82–87, May/June 2017. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL

https://www.computer.org/
csdl/mags/cg/2017/03/mcg2017030082-
abs.html.

Schneider:2021:FSI

[Sch21] Bengt-Olaf Schneider. Fukudastyle image tiling. IEEE Computer Graphics and Applications, 41(5):138–142, September/October 2021. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Staudhammer:1990:ICC

[SCN90] J. Staudhammer, T. Culviner, and M. Neal. An interview with Charles Csuri. IEEE Computer Graphics and Applications, 10(1):5– 9, January/February 1990. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Shilkrot:2013:GPI

[SCOSL13] Roy Shilkrot, Daniel Cohen-Or, Ariel Shamir, and Ligang Liu. Garment personalization via identity transfer. IEEE Computer Graphics and Applications, 33(4): 62–72, July/August 2013. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Sedbon:2022:MSE

[SCS22] Michael Sedbon, Bruce D. Campbell, and Francesca Samsel. Michael Sedbon: Explorations in coupling artificial and natural systems. IEEE Computer Graphics and Applications, 42(6): 58–63, November/December 2022. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Selikoff:2023:NSE

[SCS23] Nathan Selikoff, Bruce D. Campbell, and Francesca Samsel. Nathan Selikoff: Explorations in higher dimensionality and complexity. IEEE Computer Graphics and Applications, 43 (5):122–130, September/ October 2023. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Swezey:1983:CSH

[SD83] Robert W. Swezey and Elaine G. Davis. A case study of human factors guidelines in computer graphics. IEEE Computer Graphics and Applications, 3(8):21–30, November/December 1983. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Sowizral:1999:PVJ

[SD99] Henry A. Sowizral and Michael F. Deering. Projects in VR: The Java 3D API and virtual reality. IEEE Computer Graphics and Applications, 19(3):12–15, May/ June 1999. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://dlib.
computer.org/cg/books/
cg1999/pdf/g3012.pdf.

Sezgin:2007:SIU

[SD07] Tevfik Metin Sezgin and Randall Davis. Sketch interpretation using multiscale models of temporal patterns. IEEE Computer Graphics and Applications, 27(1):28–37, January/ February 2007. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Sra:2022:SVT

[SDM22] Misha Sra, Valdemar Danry, and Pattie Maes. Situated VR: Toward a congruent hybrid reality without experiential artifacts. IEEE Computer Graphics and Applications, 42(3):7–18, May/ June 2022. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Sprinks:2020:MMC

[SDPW20] J. Sprinks, L. Dowthwaite, G. Priestnall, and J. Wardlaw. MarsCAPE: Mars communicated through an augmented, physical environment. IEEE Computer Graphics and Applications, 40(2):43–56, March/ April 2020. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Stollnitz:1995:WCGa

[SDS95a] Eric J. Stollnitz, Tony D. DeRose, and David H. Salesin. Wavelets for computer graphics: a primer, part 1. IEEE Computer Graphics and Applications, 15(3):76–84, May/ June 1995. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Stollnitz:1995:WCGb

[SDS95b] Eric J. Stollnitz, Tony D. DeRose, and David H. Salesin. Wavelets for computer graphics: a primer, part 2. IEEE Computer Graphics and Applications, 15(4):75–85, July/August 1995. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Santos:2011:IUS

[SDS^{+}11] Beatriz Sousa Santos, Paulo Dias, Samuel Silva, Carlos Ferreira, and Joaquim Madeira. Integrating user studies into computer graphicsrelated courses. IEEE Computer Graphics and Applications, 31(5):14–17, September/October 2011. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Spetsieris:1993:ICF

[SDT^{+}93] Phoebe G. Spetsieris, Vijay Dhawan, Shugo Takikawa, Donald Margouleff, and David Eidelberg. Imaging cerebral function. IEEE Computer Graphics and Applications, 13(1):15–26, January/February 1993. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Saket:2019:DID

[SE19] B. Saket and A. Endert.

Demonstrational interaction for data visualization. IEEE Computer Graphics and Applications, 39(3):67–72, May/June 2019. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Streeb:2021:WVA

[SEAKC21] D. Streeb, M. El-Assady, D. A. Keim, and M. Chen. Why visualize? Arguments for visual support in decision making. IEEE Computer Graphics and Applications, 41(2):17–22, March/ April 2021. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Sederberg:1990:STC

[Sed90a] Thomas W. Sederberg. Surfaces — techniques for cubic algebraic surfaces. IEEE Computer Graphics and Applications, 10(4):14–25, July/August 1990. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Sederberg:1990:TCA

[Sed90b] Thomas W. Sederberg. Techniques for cubic algebraic surfaces. IEEE Computer Graphics and Applications, 10(5):12–21, September 1990. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Seidel:1993:IPF

[Sei93] Hans-Peter Seidel. An introduction to polar forms. IEEE Computer Graphics and Applications, 13(1):38– 46, January/February 1993. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Senger:1999:VSL

[Sen99] Steven Senger. Visualizing and segmenting large volumetric data sets. IEEE Computer Graphics and Applications, 19(3):32–37, May/June 1999. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

computer.org/cga/cg1999/
g3032abs.htm;   http:/
/dlib.computer.org/cg/
books/cg1999/pdf/g3032.
pdf.

Svakhine:2005:IME

[SES05] Nikolai Svakhine, David S.

Ebert, and Don Stredney. Illustration motifs for effective medical volume illustration. IEEE Computer Graphics and Applications, 25(3):31–39, May/ June 2005. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Sederberg:1992:AIB

[SF92] Thomas W. Sederberg and Rida T. Farouki. Approximation by interval Bézier curves. IEEE Computer Graphics and Applications, 12(5):87–??, September 1992. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Scurati:2020:LFW

[SF20] G. W. Scurati and F. Ferrise. Looking into a future which hopefully will not become reality: How computer graphics can impact our behavior a study of the potential of VR. IEEE Computer Graphics and Applications, 40(5):82–88, September/October 2020. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Stork:2024:WSP

[SF24] André Stork and Chi-Wing Fu. Wildfire, snow, and, panoramic ray tracing. IEEE Computer Graphics and Applications, 44(1):38– 39, January/February 2024. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Sun:1999:DSC

[SFCD99] Yinlong Sun, F. David Fracchia, Thomas W. Calvert, and Mark S. Drew. Deriving spectra from colors and rendering light interference. IEEE Computer Graphics and Applications, 19(4):61– 67, July/August 1999. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

                                        computer.org/cga/cg1999/
                                        g4061abs.htm;   http:/
                                        /dlib.computer.org/cg/
                                        books/cg1999/pdf/g4061.
                                        pdf.

Santos:2016:UHE

[SFD16] Beatriz Sousa Santos, Beatriz Quintino Ferreira, and Paulo Dias. Using heuristic evaluation to foster visualization analysis and design skills. IEEE Computer Graphics and Applications, 36(1):86–90, January/ February 2016. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
cg/2016/01/mcg2016010086-
abs.html.

Schikore:2000:HRM

[SFF^{+}00] Daniel R. Schikore, Richard A.

Fischer, Randall Frank, Ross Gaunt, John Hobson, and Brad Whitlock. Highresolution multiprojector display walls. IEEE Computer Graphics and Applications, 20(4):38–44, July/ August 2000. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

computer.org/cga/cg2000/
g4038abs.htm;   http:/
/dlib.computer.org/cg/
books/cg2000/pdf/g4038.
pdf.

Sobel:2004:PFS

[SFL^{+}04] Jason S. Sobel, Andrew S.

Forsberg, David H. Laidlaw, Robert C. Zeleznik, Daniel F. Keefe, Igor Pivkin, George E. Karniadakis, Peter Richardson, and Sharon Swartz. Particle flurries: Synoptic 3D pulsatile flow visualization. IEEE Computer Graphics and Applications, 24(2):76–85, March/ April 2004. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cg/2004/02/g2076abs.htm;
 http://csdl.computer.

                                        org/dl/mags/cg/2004/02/
                                        g2076.htm;   http://csdl.
                                        computer.org/dl/mags/cg/
                                        2004/02/g2076.pdf.

Slater:2007:URP

[SFT^{+}07] Mel Slater, Antonio Frisoli, Franco Tecchia, Christoph Guger, Beau Lotto, Anthony Steed, Gert Pfurtscheller, Robert Leeb, Miriam Reiner, Maria V. Sanchez-Vives, Paul Verschure, and Ulysses Bernardet. Understanding and realizing presence in the Presenccia Project. IEEE Computer Graphics and Applications, 27(4): 90–93, July/August 2007. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Sederberg:1986:AGC

[SG86a] Thomas W. Sederberg and Ronald N. Goldman. Algebraic geometry for computeraided geometric design. IEEE Computer Graphics and Applications, 6(6):52– 59, June 1986. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Sparks:1986:LBC

[SG86b] Madeleine R. Sparks and Julian R. Gallop. Language bindings for computer graphics standards. IEEE Computer Graphics and Applications, 6(8):58–65, August 1986. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Sakurai:1990:RSF

[SG90] Hiroshi Sakurai and David C.

Gossard. Recognizing shape features in solid models. IEEE Computer Graphics and Applications, 10(5):22– 32, September 1990. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Schweppe:2011:LTV

[SG11] Marla Schweppe and Joe Geigel. Live theater on a virtual stage: Incorporating soft skills and teamwork in computer graphics education. IEEE Computer Graphics and Applications, 31(1):85–89, January/ February 2011. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Selzer:2021:HVQ

[SGC21] Matias Nicolás Selzer, María Luján Ganuza, and Silvia Mabel Castro. High visualquality scenes in low-cost virtual reality with collisions and irregular surfaces. IEEE Computer Graphics and Applications, 41(4): 40–51, July/August 2021. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Serafin:2018:SIV

[SGE^{+}18] Stefania Serafin, Michele Geronazzo, Cumhur Erkut, Niels C. Nilsson, and Rolf Nordahl. Sonic interactions in virtual reality: State of the art, current challenges, and future directions. IEEE Computer Graphics and Applications, 38(2):31–43, March/April 2018. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2018/02/mcg2018020031-
abs.html.

Su:2012:MIV

[SGEF^{+}12] Sara Su, Connor Gramazio, Daniela Extrum-Fernandez, Caitlin Crumm, Lenore J. Cowen, Matt Menke, and Megan Strait. Molli: Interactive visualization for exploratory protein analysis. IEEE Computer Graphics and Applications, 32(5): 62–69, September/October 2012. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Shah:2022:GAC

[SGG^{+}22] Harshil Shah, Sambit Ghadai, Dhruv Gamdha, Alex Schuster, Ivan Thomas, Nathan Greiner, and Adarsh Krishnamurthy. GPU-accelerated collision analysis of vehicles in a point cloud environment. IEEE Computer Graphics and Applications, 42(5):37–50, September/October 2022. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Spooner:1997:UVP

[SH97] David L. Spooner and Martin Hardwick. Using views for product data exchange. IEEE Computer Graphics and Applications, 17(5):58–65, September/October 1997. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://dlib.
computer.org/cg/books/
cg1997/pdf/g5058.pdf.

Stapleton:2003:IIT

[SH03] Christopher Stapleton and Charles Hughes. Interactive imagination: Tapping the emotions through interactive story for compelling simulations. IEEE Computer Graphics and Applications, 23(5):11–15, September/October 2003. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http:

//csdl.computer.org/dl/
mags/cg/2003/05/g5011.
pdf.

Stapleton:2006:BSC

[SH06] Christopher Stapleton and Charles E. Hughes. Believing is seeing: Cultivating radical media innovations. IEEE Computer Graphics and Applications, 26(1):88– 93, January/February 2006. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Starck:2007:SCP

[SH07] Jonathan Starck and Adrian Hilton. Surface capture for performance-based animation. IEEE Computer Graphics and Applications, 27(3):21–31, May/ June 2007. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Sheffi:1982:NCG

[She82] Yosef Sheffi. A note on computer graphics in network evacuation studies. IEEE Computer Graphics and Applications, 2(7):57–60, 62, September 1982. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Sheppard:2004:VBC

[She04] Laurel M. Sheppard. Virtual building for construction projects. IEEE Computer Graphics and Applications, 24(1):6–12, January/February 2004. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http:

                                        //csdl.computer.org/dl/
                                        mags/cg/2004/01/g1006.
                                        pdf.

Sheppard:2005:VSB

[She05] Laurel M. Sheppard. Virtual surgery brings back smiles. IEEE Computer Graphics and Applications, 25(1):6– 11, January/February 2005. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Sheppard:2006:VEV

[She06] Laurel M. Sheppard. Visual effects and video analysis lead to Olympics victories. IEEE Computer Graphics and Applications, 26(2):6–11, March/ April 2006. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cg/2006/extras/g2006x1.
pdf.

Shen:2022:DDS

[She22] I-Chao Shen. Data-driven sketch beautification with neural feature representation. IEEE Computer Graphics and Applications, 42(4):72–79, July/August 2022. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Shimomura:1983:MAG

[Shi83]  Takao Shimomura.  A

method for automatically generating business graphs. IEEE Computer Graphics and Applications, 3(6):55– 59, September 1983. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Shires:1986:NVG

[Shi86] Glen Shires. A new VLSI graphics coprocessor — the Intel 82786. IEEE Computer Graphics and Applications, 6(10):49–55, October 1986. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Shi:2017:GAC

[Shi17] Weili Shi. A generative approach to Chinese shanshui painting. IEEE Computer Graphics and Applications, 37(1):15–19, January/ February 2017. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2017/01/mcg2017010015-
abs.html.

Sinclair:2009:UVB

[SHMN09] Jeff Sinclair, Philip Hingston, Martin Masek, and Kazunori (Ken) Nosaka. Using a virtual body to aid in exergaming system development. IEEE Computer Graphics and Applications, 29(2): 39–48, March/April 2009. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Shneiderman:2003:VVW

[Shn03] Ben Shneiderman. Visualization viewpoints: Why not make interfaces better than 3D reality? IEEE Computer Graphics and Applications, 23(6):12–15, November/December 2003. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http:

//csdl.computer.org/dl/
mags/cg/2003/06/g6012.
pdf.

Shneiderman:2020:TAI

[Shn20] B. Shneiderman. Tribute to Alfred Inselberg (1936– 2019). IEEE Computer Graphics and Applications, 40(3):120–122, May/June 2020. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Scruggs:2021:BML

[SHS^{+}21] C. Scruggs, C. Henkel, C. Stolper, K. Cook, and R. J. Crouser. Blending machine learning and interaction design in Audio Explorer. IEEE Computer Graphics and Applications, 41(2):89–95, March/ April 2021. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Shugrina:2023:DIT

[Shu23] Maria Shugrina. Designing intelligent tools for creative people. IEEE Computer Graphics and Applications, 43(3):102–110, May/ June 2023. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Shigeyama:2023:PMS

[SHY^{+}23] Jotaro Shigeyama, Takeru Hashimoto, Shigeo Yoshida, Takuji Narumi, Tomohiro Tanikawa, and Michitaka Hirose. Presenting morphing shape illusion: Enhanced sense of morphing virtual object with weight shifting VR controller by computational perception model. IEEE Computer Graphics and Applications, 43(4): 81–89, July/August 2023. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Senay:1994:KSV

[SI94] Hikmet Senay and Eve Ignatius. A knowledgebased system for visualization design. IEEE Computer Graphics and Applications, 14(6):36–47, November/December 1994. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Silver:1995:OV

[Sil95] Deborah Silver. Objectoriented visualization. IEEE Computer Graphics and Applications, 15(3):54–62, May/June 1995. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Sims:1993:DAS

[Sim93a] Dave Sims. Desktop applications support your local sheriff. IEEE Computer Graphics and Applications, 13(4):14, July/August 1993. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Sims:1993:GDC

[Sim93b] Dave Sims. Graphic design on a cosmic scale. IEEE Computer Graphics and Applications, 13(3):81, May/ June 1993. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Sims:1993:PWL

[Sim93c] Dave Sims. The point where lines converge. IEEE Computer Graphics and Applications, 13(4):7–9, July/ August 1993. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Sims:1993:VET

[Sim93d] David Sims. Virtual evidence on trial. IEEE Computer Graphics and Applications, 13(2):11–13, March/ April 1993. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Sims:1994:BRO

[Sim94a] D. Sims. Biometric recognition: our hands, eyes, and faces give us away. IEEE Computer Graphics and Applications, 14(5):14– 15, September 1994. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Sims:1994:DF

[Sim94b] D. Sims. Decriminalizing the fingerprint. IEEE Computer Graphics and Applications, 14(4):15–16, July/ August 1994. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Sims:1994:NRA

[Sim94c] D. Sims. New realities in aircraft design and manufacture. IEEE Computer Graphics and Applications, 14(2):91, March/April 1994. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Sims:1994:MCE

[Sim94d]  Dave Sims.  Multime-

dia camp empowers disabled kids. IEEE Computer Graphics and Applications, 14(1):13–14, January/ February 1994. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Sims:1995:AAO

[Sim95a] D. Sims. Applications: At Oak Ridge, a car crash on the World Wide Web. IEEE Computer Graphics and Applications, 15(3):16– 18, May/June 1995. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Sims:1995:AGU

[Sim95b]  D. Sims.  Applications:

From the ground up: building a high-resolution seismic model. IEEE Computer Graphics and Applications, 15(4):15–17, July/ August 1995. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Sims:1995:AMA

[Sim95c]  D. Sims.  Applications:

Molecules at your fingertips. IEEE Computer Graphics and Applications, 15(6): 14–16, November/December 1995. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Sims:1995:AUA

[Sim95d] D. Sims. Applications: Undersea and in the air: VR offers a thrill a minute. IEEE Computer Graphics and Applications, 15(5):15– 18, September 1995. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Sims:1995:SHT

[Sim95e] D. Sims. See how they run: modeling evacuations in VR. IEEE Computer Graphics and Applications, 15(2): 11–13, March/April 1995. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Sims:1996:ACO

[Sim96a] Dave Sims. About the cover: Osmose: Is VR supposed to be this relaxing? IEEE Computer Graphics and Applications, 16(6):4–5, November/December 1996. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Sims:1996:APV

[Sim96b] Dave Sims. Applications: Putting the visible human to work. IEEE Computer Graphics and Applications, 16(1):14–15, January/ February 1996. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Sims:1996:WEA

[Sim96c] Dave Sims. WWW extends “Apprentice’s Assistant” to global medical research; recreating a lost village in 3D space. IEEE Computer Graphics and Applications, 16(3):14–16, May/ June 1996. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Sims:1997:AAM

[Sim97] Dave Sims. Applications archaeological models: Pretty pictures or research tools? IEEE Computer Graphics and Applications, 17(1):13– 15, January/February 1997. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Sing:2001:MAA

[Sin01a] Gary Sing. Mixing audio and animation: Contemporary synesthesia. IEEE Computer Graphics and Applications, 21(5):4–5, September/October 2001. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://dlib.
computer.org/cg/books/
cg2001/g5004abs.htm;
http://dlib.computer.org/
cg/books/cg2001/pdf/g5004.
pdf.

Singh:2001:ACA

[Sin01b] Gary Singh. About the cover: Animating mechanicals: Gears, Moebius bands, and sphericons. IEEE Computer Graphics and Applications, 21(6):4–5, November/ December 2001. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL http://dlib.
                                        computer.org/cg/books/
                                        cg2001/g6004abs.htm;
                                        http://dlib.computer.org/
                                        cg/books/cg2001/pdf/g6004.
                                        pdf.

Singh:2001:ACS

[Sin01c] Gary Singh. About the cover: Showing off in 3D. IEEE Computer Graphics and Applications, 21(2):4–5, March/April 2001. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL http://dlib.
                                        computer.org/cg/books/
                                        cg2001/pdf/g2004.pdf.

Singh:2001:ADB

[Sin01d] Gary Singh. Across disciplinary boundaries. IEEE Computer Graphics and Applications, 21(3):4–5, May/ June 2001. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Singh:2002:ACD

[Sin02a] Gary Singh. About the cover: Digital visions from the subatomic realm. IEEE Computer Graphics and Applications, 22(1):4–5, January/February 2002. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http:

                                        //dlib.computer.org/cg/
                                        books/cg2002/g1004abs.
                                        htm;   http://dlib.computer.

org/cg/books/cg2002/pdf/
g1004.pdf.

Singh:2002:ACA

[Sin02b] Gary Singh. About the cover: Into the abstract. IEEE Computer Graphics and Applications, 22(6):4–5, November/December 2002. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL

http://dlib.computer.org/
cg/books/cg2002/pdf/g6004.
pdf.

Singh:2002:ACSb

[Sin02c] Gary Singh. About the cover: Shedding light on his world. IEEE Computer Graphics and Applications, 22(5):4–5, September/October 2002. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://dlib.
computer.org/cg/books/
cg2002/pdf/g5004.pdf.

Singh:2002:ACSa

[Sin02d] Gary Singh. About the cover: Surrendering control. IEEE Computer Graphics and Applications, 22(4):4–5, July/August 2002. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

computer.org/cga/cg2002/
g4004abs.htm;   http:/
/dlib.computer.org/cg/
books/cg2002/pdf/g4004.
pdf.

Singh:2002:ACW

[Sin02e] Gary Singh. About the cover: Worldwide design. IEEE Computer Graphics and Applications, 22(2):4–5, March/April 2002. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

                                        computer.org/cga/cg2002/
                                        g2004abs.htm;   http:/
                                        /dlib.computer.org/cg/
                                        books/cg2002/pdf/g2004.
                                        pdf.

Singh:2002:IEC

[Sin02f] Gary Singh. If Escher had a computer. IEEE Computer Graphics and Applications, 22(3):4–5, May/ June 2002. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

                                        computer.org/cga/cg2002/
                                        g3004abs.htm;   http:/
                                        /dlib.computer.org/cg/
                                        books/cg2002/pdf/g3004.
                                        pdf.

Singh:2003:ACD

[Sin03a] Gary Singh. About the cover: Designed for life. IEEE Computer Graphics and Applications, 23(1):4– 5, January/February 2003. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL

                                        http://dlib.computer.org/
                                        cg/books/cg2003/pdf/g1004.
                                        pdf.

Singh:2003:ACI

[Sin03b] Gary Singh. About the cover: Improvising from life. IEEE Computer Graphics and Applications, 23(5):4– 5, September/October 2003. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL

http://csdl.computer.org/
dl/mags/cg/2003/05/g5004.
pdf.

Singh:2003:ACMb

[Sin03c] Gary Singh. About the cover: Modeling cities one segment at a time. IEEE Computer Graphics and Applications, 23(6):4–5, November/December 2003. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL

http://csdl.computer.org/
dl/mags/cg/2003/06/g6004.
pdf.

Singh:2003:ACMa

[Sin03d] Gary Singh. About the cover: Music to the eyes. IEEE Computer Graphics and Applications, 23(4):4–5, July/August 2003. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cg/2003/04/g4004.pdf;
http://csdl.computer.org/
dl/mags/cg/2003/04/g4004.
htm.

Singh:2003:ACU

[Sin03e] Gary Singh. About the cover: Unlocking the Grid. IEEE Computer Graphics and Applications, 23(2):4–5, March/April 2003. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL http://dlib.
                                        computer.org/cg/books/
                                        cg2003/pdf/g2004.pdf.

Singh:2003:ACV

[Sin03f] Gary Singh. About the cover: Vivid surrealism. IEEE Computer Graphics and Applications, 23(3):4–5, May/June 2003. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL http://dlib.
                                        computer.org/cg/books/
                                        cg2003/pdf/g3004.pdf.

Singh:2004:ACD

[Sin04a] Gary Singh. About the cover: Design by points. IEEE Computer Graphics and Applications, 24(4):4–5, July/August 2004. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL http://csdl.
                                        computer.org/comp/mags/
                                        cg/2004/04/g4004.pdf;
                                        http://csdl.computer.org/
                                        dl/mags/cg/2004/04/g4004.
                                        htm.

Singh:2004:BH

[Sin04b]  Gary Singh.  Brushing

into haptics. IEEE Computer Graphics and Applications, 24(2):4–5, March/ April 2004. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http:/

/csdl.computer.org/dl/
mags/cg/2004/02/g2004.
htm;   http://csdl.computer.
org/dl/mags/cg/2004/02/
g2004.pdf.

Singh:2004:BTP

[Sin04c]  Gary Singh.  Building

terrain from pebbles to mountains. IEEE Computer Graphics and Applications, 24(1):4–5, January/February 2004. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http:

//csdl.computer.org/dl/
mags/cg/2004/01/g1004.
pdf.

Singh:2004:HP

[Sin04d]  Gary Singh.  Haptic

painting. IEEE Computer Graphics and Applications, 24(6):4–5, November/ December 2004. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cg/2004/06/g6004.pdf;
http://csdl.computer.org/
dl/mags/cg/2004/06/g6004.
htm.

Singh:2004:TEA

[Sin04e]  Gary Singh.  Trial-and-

error artistry. IEEE Computer Graphics and Applications, 24(3):4–5, May/ June 2004. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http:/

                                        /csdl.computer.org/dl/
                                        mags/cg/2004/03/g3004.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/cg/2004/03/
                                        g3004.pdf.

Singh:2004:VF

[Sin04f] Gary Singh. Visual fusion. IEEE Computer Graphics and Applications, 24(5):4– 5, September/October 2004. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL

                                        http://csdl.computer.org/
                                        comp/mags/cg/2004/05/g5004.
                                        pdf;   http://csdl.computer.
                                        org/dl/mags/cg/2004/05/
                                        g5004.htm.

Singh:2005:DAR

[Sin05a] Gary Singh. Digital art revolution. IEEE Computer Graphics and Applications, 25(2):4–5, March/ April 2005. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL http://csdl.
                                        computer.org/comp/mags/
                                        cg/2005/02/g2004.pdf.

Singh:2005:FR

[Sin05b] Gary Singh. Filters rock. IEEE Computer Graphics and Applications, 25(4):4–5, July/August 2005. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cg/2005/04/g4004.pdf.

Singh:2005:FF

[Sin05c] Gary Singh. Fractal fracas.

IEEE Computer Graphics and Applications, 25(3):4–5, May/June 2005. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cg/2005/03/g3004.pdf.

Singh:2005:A

[Sin05d]  Gary Singh.  Into the

abstract. IEEE Computer Graphics and Applications, 25(6):4–5, November/ December 2005. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cg/2005/06/g6004.pdf.

Singh:2005:SS

[Sin05e] Gary Singh. Stone secrets.

IEEE Computer Graphics and Applications, 25(1):4– 5, January/February 2005. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL

http://csdl.computer.org/
comp/mags/cg/2005/01/g1004.
pdf.

Singh:2005:SF

[Sin05f]  Gary Singh.  Stringing

the fractals. IEEE Computer Graphics and Applications, 25(5):4–5, September/October 2005. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL http://csdl.
                                        computer.org/comp/mags/
                                        cg/2005/05/g5004.pdf.

Singh:2006:BCU

[Sin06a] Gary Singh. Branching into cosmic unity. IEEE Computer Graphics and Applications, 26(4):4–5, July/ August 2006. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Singh:2006:FV

[Sin06b] Gary Singh. A fractal voyage. IEEE Computer Graphics and Applications, 26(3): 4–5, May/June 2006. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

                                        csdl.computer.org/comp/
                                        mags/cg/2006/03/g3004.
                                        pdf.

Singh:2006:IP

[Sin06c] Gary Singh. Infinite possibilities. IEEE Computer Graphics and Applications, 26(6):4–5, November/ December 2006. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cg/2006/06/g6004.pdf.

Singh:2006:MC

[Sin06d]  Gary Singh.  Machine

creation. IEEE Computer Graphics and Applications, 26(5):4–5, September/October 2006. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cg/2006/05/g5004.pdf.

Singh:2006:SD

[Sin06e]  Gary Singh.  A space

to dream. IEEE Computer Graphics and Applications, 26(1):4–5, January/ February 2006. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Singh:2006:TBB

[Sin06f] Gary Singh. Things between and behind. IEEE Computer Graphics and Applications, 26(2):4–5, March/ April 2006. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cg/2006/02/g2004.pdf.

Singh:2007:BC

[Sin07a]  Gary Singh.  Beauty

in chaos. IEEE Computer Graphics and Applications, 27(5):4–5, September/October 2007. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Singh:2007:BI

[Sin07b]  Gary Singh.  The blitz

of infinity. IEEE Computer Graphics and Applications, 27(3):4–5, May/ June 2007. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL http://csdl.
                                        computer.org/comp/mags/
                                        cg/2007/03/g3004.pdf.

Singh:2007:DC

[Sin07c] Gary Singh. The dynamics of change. IEEE Computer Graphics and Applications, 27(4):4–5, July/ August 2007. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Singh:2007:UE

[Sin07d]  Gary Singh.  Untitled

evolution. IEEE Computer Graphics and Applications, 27(2):4–5, March/ April 2007. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL http://csdl.
                                        computer.org/comp/mags/
                                        cg/2007/02/g2004.pdf.

Singh:2007:VU

[Sin07e] Gary Singh. Vamping on ultrafractal. IEEE Computer Graphics and Applications, 27(1):4–5, January/ February 2007. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cg/2007/01/g1004.pdf.

Singh:2008:ACM

[Sin08a] Gary Singh. About the cover: Mad science in art. IEEE Computer Graphics and Applications, 28(6):4–5, November/December 2008. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Singh:2008:AF

[Sin08b] Gary Singh. Abstracts and fragments. IEEE Computer Graphics and Applications, 28(3):4–5, May/June 2008. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Singh:2008:AE

[Sin08c]  Gary Singh.  Artificial

evolution. IEEE Computer Graphics and Applications, 28(2):4–5, March/ April 2008. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cg/2008/02/mcg2008020004.
pdf.

Singh:2008:BDC

[Sin08d] Gary Singh. Back to digital crayons. IEEE Computer Graphics and Applications, 28(5):4–5, September/October 2008. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Singh:2008:ETA

[Sin08e] Gary Singh. Expression through art. IEEE Computer Graphics and Applications, 28(4):4–5, July/ August 2008. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Singh:2008:CC

[Sin08f] Gary Singh. From chemicals to creation. IEEE Computer Graphics and Applications, 28(1):4–5, January/ February 2008. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Singh:2009:ACDa

[Sin09a] Gary Singh. About the cover: Digital master. IEEE Computer Graphics and Applications, 29(3):4–5, May/ June 2009. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Singh:2009:ACDb

[Sin09b] Gary Singh. About the cover: Dressing it up. IEEE Computer Graphics and Applications, 29(4):4–5, July/ August 2009. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Singh:2009:ACF

[Sin09c] Gary Singh. About the cover: Fun with fractal art. IEEE Computer Graphics and Applications, 29(1):4– 5, January/February 2009. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Singh:2009:ACS

[Sin09d] Gary Singh. About the cover: The spiral effect. IEEE Computer Graphics and Applications, 29(2):4–5, March/April 2009. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Singh:2009:RV

[Sin09e] Gary Singh. Revisualizing visualization. IEEE Computer Graphics and Applications, 29(5):4–5, September/October 2009. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Singh:2009:SW

[Sin09f] Gary Singh. Soft as a whisper. IEEE Computer Graphics and Applications, 29(6): 4–5, November/December 2009. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Singh:2010:AC

[Sin10a]  Gary Singh.  An ab-

stract career. IEEE Computer Graphics and Applications, 30(1):4–5, January/ February 2010. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Singh:2010:AD

[Sin10b] Gary Singh. Alive and digital. IEEE Computer Graphics and Applications, 30(3):4–5, May/June 2010. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Singh:2010:EM

[Sin10c]  Gary Singh.  Emotions

in motion. IEEE Computer Graphics and Applications, 30(4):4–5, July/ August 2010. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Singh:2010:SC

[Sin10d] Gary Singh. Serial crossover. IEEE Computer Graphics and Applications, 30(6):8–9, November/December 2010. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Singh:2010:VTW

[Sin10e] Gary Singh. A vehicle through which to travel. IEEE Computer Graphics and Applications, 30(5):4– 5, September/October 2010. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Singh:2010:WFM

[Sin10f] Gary Singh. Working in a fast medium. IEEE Computer Graphics and Applications, 30(2):4–5, March/ April 2010. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Singh:2011:AG

[Sin11a]  Gary Singh.  Against

the grain. IEEE Computer Graphics and Applications, 31(1):3–4, January/ February 2011. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Singh:2011:DF

[Sin11b] Gary Singh. Diving into the flow. IEEE Computer Graphics and Applications, 31(4):3–4, July/ August 2011. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Singh:2011:HOV

[Sin11c] Gary Singh. Her own virtual gallery of art. IEEE Computer Graphics and Applications, 31(6):4–5, November/ December 2011. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Singh:2011:IBT

[Sin11d] Gary Singh. An infinite bag of tricks. IEEE Computer Graphics and Applications, 31(2):4–5, March/ April 2011. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Singh:2011:PT

[Sin11e] Gary Singh. The power of tutorials. IEEE Computer Graphics and Applications, 31(5):4–5, September/October 2011. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Singh:2011:WWB

[Sin11f] Gary Singh. When work becomes bliss. IEEE Computer Graphics and Applications, 31(3):4–5, May/June 2011. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Singh:2012:BBR

[Sin12a] Gary Singh. The best of both realities. IEEE Computer Graphics and Applications, 32(2):4–5, March/ April 2012. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Singh:2012:RBL

[Sin12b] Gary Singh. From right brain to left brain. IEEE Computer Graphics and Applications, 32(4):4–5, July/ August 2012. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Singh:2012:HWA

[Sin12c] Gary Singh. Her work is all the buzz. IEEE Computer Graphics and Applications, 32(1):4–5, January/ February 2012. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Singh:2012:JPD

[Sin12d] Gary Singh. Jerry Palmer:

Digital painter. IEEE Computer Graphics and Applications, 32(6):4–5, November/ December 2012. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Singh:2012:RPH

[Sin12e] Gary Singh. Reflecting the prism of his life. IEEE Computer Graphics and Applications, 32(3):4–5, May/ June 2012. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Singh:2012:SFL

[Sin12f] Gary Singh. A saturation of firsts: Lillian F. Schwartz.

IEEE Computer Graphics and Applications, 32(5):4– 5, September/October 2012. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Singh:2013:CDA

[Sin13a] Gary Singh. Creation and deconstruction [about the cover]. IEEE Computer Graphics and Applications, 33(3):3–4, May/June 2013. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Singh:2013:DTS

[Sin13b] Gary Singh. Dreams, tea, and science. IEEE Computer Graphics and Applications, 33(5):6–8, September/October 2013. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Singh:2013:CA

[Sin13c] Gary Singh. From code to art. IEEE Computer Graphics and Applications, 33(6): 4–5, November/December 2013. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Singh:2013:JED

[Sin13d] Gary Singh. The joy of exploration and discovery. IEEE Computer Graphics and Applications, 33(2):4–5, March/April 2013. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Singh:2013:MSC

[Sin13e] Gary Singh. The mind that sees and creates. IEEE Computer Graphics and Applications, 33(4):4–5, July/ August 2013. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Singh:2013:SN

[Sin13f] Gary Singh. Something out of nothing. IEEE Computer Graphics and Applications, 33(1):4–5, January/ February 2013. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Singh:2014:CDC

[Sin14a] Gary Singh. CultLab3D:

Digitizing cultural heritage. IEEE Computer Graphics and Applications, 34(3):4–5, May/June 2014. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Singh:2014:ETA

[Sin14b] Gary Singh. The eventual triumph of art. IEEE Computer Graphics and Applications, 34(5):4–5, September/October 2014. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://csdl.
computer.org/csdl/mags/

                                        cg/2014/05/mcg2014050004.
                                        html.

Singh:2014:FW

[Sin14c]  Gary Singh.  Filtering

the world. IEEE Computer Graphics and Applications, 34(6):4–5, November/ December 2014. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL http://csdl.
                                        computer.org/csdl/mags/
                                        cg/2014/06/mcg2014060004.
                                        html.

Singh:2014:PC

[Sin14d] Gary Singh. The poetry of colors. IEEE Computer Graphics and Applications, 34(2):4–5, March/ April 2014. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Singh:2014:RL

[Sin14e] Gary Singh. The repurpose of life. IEEE Computer Graphics and Applications, 34(4):4–5, July/ August 2014. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Singh:2014:TF

[Sin14f] Gary Singh. Transforming fractals. IEEE Computer Graphics and Applications, 34(1):4–5, January/ February 2014. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Singh:2015:ACP

[Sin15a] Gary Singh. Abstract creative practice. IEEE Computer Graphics and Applications, 35(3):4–5, May/ June 2015. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
cg/2015/03/mcg2015030004.
html.

Singh:2015:EDA

[Sin15b]  Gary Singh.  Electrify-

ing digital abstract art. IEEE Computer Graphics and Applications, 35(2):4–5, March/April 2015. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
cg/2015/02/mcg2015020004.
html.

Singh:2015:EFT

[Sin15c] Gary Singh. Exploring the fractal terrain. IEEE Computer Graphics and Applications, 35(1):4–5, January/ February 2015. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://csdl.
computer.org/csdl/mags/
cg/2015/01/mcg2015010004-
abs.html.

Singh:2015:IFL

[Sin15d] Gary Singh. Infinite flexibility with layers. IEEE Computer Graphics and Applications, 35(4):4–5, July/ August 2015. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL http://www.
                                        computer.org/csdl/mags/
                                        cg/2015/04/mcg2015040004.
                                        html.

Singh:2015:PP

[Sin15e] Gary Singh. Pass the parcel. IEEE Computer Graphics and Applications, 35(6):2–3, November/December 2015. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL

                                        http://www.computer.org/
                                        csdl/mags/cg/2015/06/mcg2015060002.
                                        html.

Singh:2015:WBS

[Sin15f] Gary Singh. The wonder and beauty of science. IEEE Computer Graphics and Applications, 35(5):2– 3, September/October 2015. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL

                                        http://www.computer.org/
                                        csdl/mags/cg/2015/05/mcg2015050002.
                                        html.

Singh:2016:GFC

[Sin16a] Gary Singh. A genealogy of fractal code [about the cover]. IEEE Computer Graphics and Applications, 36(2):3–4, March/ April 2016. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2016/02/mcg2016020003.
html.

Singh:2016:ITD

[Sin16b] Gary Singh. Improvising in three dimensions. IEEE Computer Graphics and Applications, 36(4):4–5, July/ August 2016. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2016/04/mcg2016040004-
abs.html.

Singh:2016:LDB

[Sin16c] Gary Singh. Landscapes of the digital baroque. IEEE Computer Graphics and Applications, 36(3):4–5, May/ June 2016. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2016/03/mcg2016030004-
abs.html.

Singh:2016:ODA

[Sin16d] Gary Singh. Outstanding design [about the cover]. IEEE Computer Graphics and Applications, 36(5):4– 6, September/October 2016. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL

                                        https://www.computer.org/
                                        csdl/mags/cg/2016/05/mcg2016050004.
                                        html.

Singh:2016:PA

[Sin16e]  Gary Singh.  Pure ab-

straction. IEEE Computer Graphics and Applications, 36(6):4–5, November/ December 2016. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        cg/2016/06/mcg2016060004-
                                        abs.html.

Singh:2016:SHE

[Sin16f] Gary Singh. Seeds of human experience. IEEE Computer Graphics and Applications, 36(1):2–3, January/ February 2016. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL http://www.
                                        computer.org/csdl/mags/
                                        cg/2016/01/mcg2016010002-
                                        abs.html.

Singh:2017:CRC

[Sin17a] Gary Singh. Control and release in complex relationships. IEEE Computer Graphics and Applications, 37(6):3–4, November/ December 2017. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        cg/2017/06/mcg2017060003.
                                        html.

Singh:2017:IN

[Sin17b]  Gary Singh.  Illuminat-

ing nature. IEEE Computer Graphics and Applications, 37(5):3–4, September/October 2017. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2017/05/mcg2017050003.
html.

Singh:2017:SD

[Sin17c] Gary Singh. Stories in the data. IEEE Computer Graphics and Applications, 37(3):4–5, May/ June 2017. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2017/03/mcg2017030004.
html.

Singh:2017:TF

[Sin17d] Gary Singh. Taking flight.

IEEE Computer Graphics and Applications, 37(1):4– 5, January/February 2017. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL

https://www.computer.org/
csdl/mags/cg/2017/01/mcg2017010004-
abs.html.

Singh:2017:VS

[Sin17e]  Gary Singh.  Visualiz-

ing science. IEEE Computer Graphics and Applications, 37(4):4–5, July/ August 2017. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2017/04/mcg2017040004.
html.

Singh:2017:P

[Sin17f] Gary Singh. Without a plan. IEEE Computer Graphics and Applications, 37(2):4–5, March/April 2017. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2017/02/mcg2017020004-
abs.html.

Singh:2018:AT

[Sin18a] Gary Singh. Alchemical transformation. IEEE Computer Graphics and Applications, 38(3):5–7, May/ June 2018. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2018/03/mcg2018030005.
html.

Singh:2018:CLA

[Sin18b] Gary Singh. From command lines to the Arctic. IEEE Computer Graphics and Applications, 38(6): 5–8, ???? 2018. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL https:

//ieeexplore.ieee.org/
document/8617731/.

Singh:2018:IVC

[Sin18c] Gary Singh. Improving visual communication. IEEE Computer Graphics and Applications, 38(1):8–10, January/February 2018. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL https:/

/www.computer.org/csdl/
mags/cg/2018/01/mcg2018010008.
html.

Singh:2018:NC

[Sin18d] Gary Singh. The nature of collaboration. IEEE Computer Graphics and Applications, 38(2):4–6, March/ April 2018. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2018/02/mcg2018020004.
html.

Singh:2018:RM

[Sin18e] Gary Singh. The right medium. IEEE Computer Graphics and Applications, 38(5):7–10, September/October 2018. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2018/05/mcg2018050007.
html.

Singh:2018:WMH

[Sin18f] Gary Singh. Wearing multiple hats. IEEE Computer Graphics and Applications, 38(4):6–8, July/ August 2018. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        cg/2018/04/mcg2018040006.
                                        html.

Singh:2019:HD

[Sin19a] G. Singh. Hungry for data. IEEE Computer Graphics and Applications, 39(5):5– 7, September/October 2019. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Singh:2019:TD

[Sin19b] G. Singh. In time with data. IEEE Computer Graphics and Applications, 39(6):5–7, November/December 2019. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Singh:2019:MIV

[Sin19c] G. Singh. Making the invisible visible. IEEE Computer Graphics and Applications, 39(4):5–7, July/ August 2019. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Singh:2019:RAA

[Sin19d] G. Singh. Resident artist [about the cover]. IEEE Computer Graphics and Applications, 39(3):5–7, May/ June 2019. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Singh:2019:EE

[Sin19e] Gary Singh. Economies of exascale. IEEE Computer Graphics and Applications, 39(1):9–11, January/February 2019. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL https:

//ieeexplore.ieee.org/
document/8664665/.

Singh:2019:GP

[Sin19f] Gary Singh. Going permanent. IEEE Computer Graphics and Applications, 39(2):5–7, March/ April 2019. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL https:

//ieeexplore.ieee.org/
document/8673008/.

Singh:2020:KR

[Sin20a] G. Singh. Killing rainbows.

IEEE Computer Graphics and Applications, 40(2):5–7, March/April 2020. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Singh:2020:MPD

[Sin20b] G. Singh. Machine, perchance to dream. IEEE Computer Graphics and Applications, 40(5):5–7, September/October 2020. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Singh:2020:PT

[Sin20c] G. Singh. The power of transformation. IEEE Computer Graphics and Applications, 40(3):5–8, May/ June 2020. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Singh:2020:M

[Sin20d] G. Singh. To the moon. IEEE Computer Graphics and Applications, 40(1):5– 7, January/February 2020. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Singh:2020:VV

[Sin20e]  G. Singh.  Valley viz.

IEEE Computer Graphics and Applications, 40(6):5–7, November/December 2020. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Singh:2020:ZMM

[Sin20f] G. Singh. Zen mind, machine mind. IEEE Computer Graphics and Applications, 40(4):5–7, July/ August 2020. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Singh:2021:DR

[Sin21a] G. Singh. Data redesign.

IEEE Computer Graphics and Applications, 41(2):5–7, March/April 2021. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Singh:2021:HH

[Sin21b]  G. Singh.  From hu-

man hands. IEEE Computer Graphics and Applications, 41(1):5–7, January/ February 2021. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Singh:2021:BW

[Sin21c] Gary Singh. Brad’s world.

IEEE Computer Graphics and Applications, 41(3):5–7, May/June 2021. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Singh:2021:DG

[Sin21d] Gary Singh. The dharma of GANs. IEEE Computer Graphics and Applications, 41(5):5–6, September/October 2021. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Singh:2021:RA

[Sin21e]  Gary Singh.  Radical

alliances. IEEE Computer Graphics and Applications, 41(4):5–6, July/ August 2021. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Singh:2021:UAM

[Sin21f] Gary Singh. Under the algorithmic microscope [about the cover]. IEEE Computer Graphics and Applications, 41(6):5–6, November/ December 2021. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Singh:2022:AF

[Sin22a] Gary Singh. About face. IEEE Computer Graphics and Applications, 42(3):5–6, May/June 2022. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Singh:2022:BC

[Sin22b] Gary Singh. Build this city. IEEE Computer Graphics and Applications, 42(2):5–7, March/April 2022. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Singh:2022:GS

[Sin22c] Gary Singh. The game of slime. IEEE Computer Graphics and Applications, 42(6):5–6, November/ December 2022. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Singh:2022:IA

[Sin22d]  Gary Singh.  Intelligent

accidents. IEEE Computer Graphics and Applications, 42(5):5–6, September/October 2022. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Singh:2022:MT

[Sin22e] Gary Singh. The Medusa touch. IEEE Computer Graphics and Applications, 42(4):5–7, July/August 2022. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Singh:2022:WYD

[Sin22f]  Gary Singh.  Wake up

your data. IEEE Computer Graphics and Applications, 42(1):5–6, January/ February 2022. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Singh:2023:NSC

[Sin23a] Gary Singh. And now for someone completely different. IEEE Computer Graphics and Applications, 43 (2):5–7, March/April 2023. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Singh:2023:BWA

[Sin23b] Gary Singh. Bacteria for the win [about the cover].

IEEE Computer Graphics and Applications, 43(1):5– 6, January/February 2023. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Singh:2023:CQR

[Sin23c] Gary Singh. Causal quartets to the rescue. IEEE Computer Graphics and Applications, 43(5):5–7, September/October 2023. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Singh:2023:DD

[Sin23d] Gary Singh. Driven by data. IEEE Computer Graphics and Applications, 43(3):5–7, May/June 2023. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Singh:2023:LLS

[Sin23e] Gary Singh. Long live SIGGRAPH! IEEE Computer Graphics and Applications, 43(4):5–9, July/ August 2023. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Singh:2023:OD

[Sin23f] Gary Singh. Open density. IEEE Computer Graphics and Applications, 43(6):5–6, November/December 2023. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Singh:2024:ADJ

[Sin24a]  Gary Singh.  All that

data jazz. IEEE Computer Graphics and Applications, 44(1):5–7, January/ February 2024. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Singh:2024:HLA

[Sin24b] Gary Singh. Holon loosely [about the cover]. IEEE Computer Graphics and Applications, 44(2):5–7, March/ April 2024. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Stein:2016:DCA

[SJB^{+}16] Manuel Stein, Halldor Janetzko, Thorsten Breitkreutz, Daniel Seebacher, Tobias Schreck, Michael Grossniklaus, Iain D. Couzin, and Daniel A. Keim. Director’s cut: Analysis and annotation of soccer matches. IEEE Computer Graphics and Applications, 36(5):50–60, September/October 2016. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2016/05/mcg2016050050-
abs.html.

Stein:2019:TSS

[SJSK19] M. Stein, H. Janetzko, T. Schreck, and D. A. Keim. Tackling similarity search for soccer match analysis: Multimodal distance measure and interactive query definition. IEEE Computer Graphics and Applications, 39(5):60–71, September/October 2019. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Staudhammer:1986:DMM

[SK86] John Staudhammer and Anoop S. Khurana. Display of molecular models with interactive graphics. IEEE Computer Graphics and Applications, 6(1):26– 31, January/February 1986. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Shinagawa:1991:CRG

[SK91] Yoshihisa Shinagawa and Tosiyasu L. Kunii. Constructing a Reeb graph automatically from cross sections. IEEE Computer Graphics and Applications, 11(6):44–51, November/ December 1991. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Shirman:1997:FAT

[SK97] Leon Shirman and Yakov Kamen. Fast and accurate texture placement. IEEE Computer Graphics and Applications, 17(1):60– 66, January/February 1997. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Sheppard:2002:CSD

[SK02] Laurel Sheppard and Mike Kohorst. CAD speeds up dinnerware designs. IEEE Computer Graphics and Applications, 22(5):16–20, September/October 2002. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL

http://dlib.computer.org/
cg/books/cg2002/pdf/g5016.
pdf.

Seitz:2003:PVM

[SK03] Steven M. Seitz and Jiwon Kim. Projects in VR: Multiperspective imaging. IEEE Computer Graphics and Applications, 23(6):16–19, November/ December 2003. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http:

//csdl.computer.org/dl/
mags/cg/2003/06/g6016.
pdf.

Suzuki:2005:AJD

[SK05] Yuriko Suzuki and Minoru Kobayashi. Air jet driven force feedback in virtual reality. IEEE Computer Graphics and Applications, 25(1):44–47, January/ February 2005. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Schell:2013:BVW

[SK13] Jesse Schell and Chris Klug. Building virtual worlds carrying on the legacy of Randy Pausch’s “Head fake”. IEEE Computer Graphics and Applications, 33(3):12–15, May/June 2013. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Skall:1986:NRC

[Ska86] Mark W. Skall. NBS’s role in computer graphics standards. IEEE Computer Graphics and Applications, 6(8):66–70, August 1986. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Schwenk:2012:PNR

[SKBF12] Karsten Schwenk, Arjan Kuijper, Johannes Behr, and Dieter W. Fellner. Practical noise reduction for progressive stochastic ray tracing with perceptual control. IEEE Computer Graphics and Applications, 32(6): 46–55, November/December 2012. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Skinner:1981:IWS

[Ski81] Frank D. Skinner. The interactive wiring system. IEEE Computer Graphics and Applications, 1(2):38–42, 44– 46, 48–51, April 1981. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Schroeder:2013:CWA

[SKJ^{+}13] David Schroeder, Fedor Korsakov, Joseph Jolton, Francis J. Keefe, Alex Haley, and Daniel F. Keefe. Creating widely accessible spatial interfaces: Mobile VR for managing persistent pain. IEEE Computer Graphics and Applications, 33(3):82– 88, May/June 2013. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Shinagawa:1991:SCB

[SKK91] Yoshihisa Shinagawa, Tosiyasu L.

Kunii, and Yannick L. Kergosien. Surface coding based on Morse theory. IEEE Computer Graphics and Applications, 11(5):66–78, September 1991. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Slater:2009:VRE

[SKMY09] Mel Slater, Pankaj Khanna, Jesper Mortensen, and Insu Yu. Visual realism enhances realistic response in an immersive virtual environment. IEEE Computer Graphics and Applications, 29(3):76–84, May/ June 2009. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Sugimoto:2005:TFV

[SKN^{+}05] Maki Sugimoto, Georges Kagotani, Hideaki Nii, Naoji Shiroma, Masahiko Inami, and Fumitoshi Matsuno. Time follower’s vision: a teleoperation interface with past images. IEEE Computer Graphics and Applications, 25(1):54–63, January/ February 2005. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Sagong:2015:SDS

[SKN^{+}15] Donghoon Sagong, Nahyup Kang, Junyong Noh, Xiaogang Jin, and Joseph S. Shin. Simulating drops settling in a still liquid. IEEE Computer Graphics and Applications, 35(1):46–55, January/February 2015. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

csdl.computer.org/csdl/
mags/cg/2015/01/mcg2015010046-
abs.html.

Shah:2009:ISS

[SKP09] Musawir A. Shah, Jaakko Konttinen, and Sumanta Pattanaik. Image-space subsurface scattering for interactive rendering of deformable translucent objects. IEEE Computer Graphics and Applications, 29(1):66–78, January/February 2009. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Samsel:2018:CRT

[SKR18] Francesca Samsel, Sebastian Klaassen, and David H. Rogers. ColorMoves: Realtime interactive colormap construction for scientific visualization. IEEE Computer Graphics and Applications, 38(1):20–29, January/ February 2018. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2018/01/mcg2018010020-
abs.html.

Schilling:1996:TSM

[SKS96a] Andreas Schilling, Günter Knittel, and Wolfgang Strasser. Texram: a smart memory for texturing. IEEE Computer Graphics and Applications, 16(3):32–41, May/June 1996. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Sclabassi:1996:NCI

[SKS^{+}96b] Robert J. Sclabassi, Donald Krieger, Robert Simon, Ray Lofink, Greg Gross, and Donald M. De-Lauder. NeuroNet: Collaborative intraoperative guidance and control. IEEE Computer Graphics and Applications, 16(1):39–45, January/February 1996. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Scharnowski:2013:IVC

[SKS^{+}13] Katrin Scharnowski, Michael Krone, Filip Sadlo, Philipp Beck, Johannes Roth, Hans-Rainer Trebin, and Thomas Ertl. 2012 IEEE Visualization Contest winner: Visualizing polarization domains in barium titanate. IEEE Computer Graphics and Applications, 33(5): 9–17, September/October 2013. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

L:2010:VAO

[SKUT^{+}10] László Szirmay-Kalos, Tamás Umenhoffer, Balázs Tóth, László Szécsi, and Mateu Sbert. Volumetric ambient occlusion for realtime rendering and games. IEEE Computer Graphics and Applications, 30(1):70– 79, January/February 2010. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Schroeder:2012:EVS

[SKW^{+}12] David Schroeder, Timothy Kowalewski, Lee White, John Carlis, Erlan Santos, Robert Sweet, Thomas S. Lendvay, Daniel F. Keefe, and Troy Reihsen. Exploratory visualization of surgical training databases for improving skill acquisition. IEEE Computer Graphics and Applications, 32(6):71–81, November/ December 2012. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Stephenson:2000:WSW

[SL00] Peter Stephenson and Bruce Litow. Why step when you can run? iterative line digitization algorithms based on hierarchies of runs. IEEE Computer Graphics and Applications, 20(6):76–84, November/December 2000. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

computer.org/cga/cg2000/
g6076abs.htm;   http:/
/dlib.computer.org/cg/
books/cg2000/pdf/g6076.
pdf.

Shapiro:2011:PCP

[SL11] Ari Shapiro and Sung-Hee Lee. Practical character physics for animators. IEEE Computer Graphics and Applications, 31(4):45–55, July/August 2011. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Shen:2023:IVC

[SLC^{+}23] Siqi Shen, Xingui Lai, Siming Chen, Qinghong Wang, and Junting Gao. IEEE VAST Challenge 2021 winner: Visual analytics for spatial temporal situation awareness. IEEE Computer Graphics and Applications, 43(5):12–22, September/October 2023. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Shen:2024:NFF

[SLC24] I-Chao Shen, Hao-Kang Liu, and Bing-Yu Chen. NeRF-In: Free-form inpainting for pretrained NeRF with RGB-D priors. IEEE Computer Graphics and Applications, 44(2):100–109, March/April 2024. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Setlur:2007:RIV

[SLNG07] Vidya Setlur, Tom Lechner, Marc Nienhaus, and Bruce Gooch. Retargeting images and video for preserving information saliency. IEEE Computer Graphics and Applications, 27(5): 80–88, September/October 2007. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Sun:2022:KBD

[SLSP22] Shantong Sun, Rongke Liu, Shuqiao Sun, and Unsang Park. Keypoint-based disentangled pose network for category-level 6-D object pose tracking. IEEE Computer Graphics and Applications, 42(5):28–36, September/October 2022. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Sears:1984:SVM

[SM84] Ken H. Sears and Alan E.

Middleditch. Set-theoretic volume model evaluation and picture-plane coherence. IEEE Computer Graphics and Applications, 4(3):41– 46, March/April 1984. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Schwartz:1986:CND

[SM86] Eric L. Schwartz and Bjorn Merker. Computer-aided neuroanatomy: Differential geometry of cortical surfaces and an optimal flattening algorithm. IEEE Computer Graphics and Applications, 6(3):36–44, March/ April 1986. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Sandor:1991:SD

[SM91] Ellen Sandor and Stephan Meyers. Science in depth.

IEEE Computer Graphics and Applications, 11(2): 10–12, March/April 1991. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Sheng:1995:GTS

[SM95] Xuejun Sheng and Ingo R. Meier. Generating topological structures for surface models. IEEE Computer Graphics and Applications, 15(6):35–41, November/December 1995. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Shon:2004:EDS

[SM04] Youngung Shon and Sara McMains. Evaluation of drawing on 3D surfaces with haptics. IEEE Computer Graphics and Applications, 24(6):40–50, November/December 2004. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http:

//csdl.computer.org/dl/
mags/cg/2004/06/g6040.
htm;   http://csdl.computer.
org/dl/mags/cg/2004/06/
g6040.pdf.

Satilmis:2022:DSC

[SMDBR22] Pinar Satilmis, Demetris Marnerides, Kurt Debattista, and Thomas Bashford-Rogers. Deep synthesis of cloud lighting. IEEE Computer Graphics and Applications, 42(5):8–18, September/October 2022. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Schatz:2021:ISV

[SMG^{+}21] Karsten Schatz, Christoph Müller, Patrick Gralka, Moritz Heinemann, Alexander Straub, Christoph Schulz, Matthias Braun, Tobias Rau, Michael Becher, Steffen Frey, Guido Reina, Michael Sedlmair, Daniel Weiskopf, Thomas Ertl, Patrick Diehl, Dominic Marcello, Juhan Frank, and Thomas Müller. 2019 IEEE scientific visualization contest winner: Visual analysis of structure formation in cosmic evolution. IEEE Computer Graphics and Applications, 41(6):101–110, November/December 2021. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Smith:1983:IKC

[Smi83] Bradford M. Smith. IGES: a key to CAD/CAM systems integration. IEEE Computer Graphics and Applications, 3(8):78–83, November/December 1983. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Smith:1984:HFI

[Smi84a] Michael J. Smith. Human factors issues in VDT use: Environmental and workstation design considerations. IEEE Computer Graphics and Applications, 4(11): 56–63, November/December 1984. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Smith:1984:PME

[Smi84b] Michael J. Smith. Physical, mental, and emotional stress effects of VDT work. IEEE Computer Graphics and Applications, 4(4):23– 27, April 1984. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Smith:1987:EAR

[Smi87] Randall B. Smith. Experiences with the alternate reality kit: an example of the tension between literalism and magic. IEEE Computer Graphics and Applications, 7(9):42–50, September 1987. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Smuc:2009:SST

[SML^{+}09] Michael Smuc, Eva Mayr, Tim Lammarsch, Wolfgang Aigner, Silvia Miksch, and Johannes Gärtner. To score or not to score? Tripling insights for participatory design. IEEE Computer Graphics and Applications, 29(3):29–38, May/ June 2009. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Sochting:2024:LIV

[SMMS24] Maximilian Söchting, Miguel D.

Mahecha, David Montero, and Gerik Scheuermann. Lexcube: Interactive visualization of large earth system data cubes. IEEE Computer Graphics and Applications, 44(1):25–37, January/ February 2024. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Seebacher:2019:VAV

[SMP^{+}19] D. Seebacher, M. Miller, T. Polk, J. Fuchs, and D. A. Keim. Visual analytics of volunteered geographic information: Detection and investigation of urban heat islands. IEEE Computer Graphics and Applications, 39(5):83–95, September/ October 2019. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Shoman:2022:RSF

[SMP^{+}22] Sota Shoman, Tomohiro Mashita, Alexander Plopski, Photchara Ratsamee, and Yuki Uranishi. Realto-synthetic feature transform for illumination invariant camera localization. IEEE Computer Graphics and Applications, 42(1):47– 55, January/February 2022. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Schwartz:1988:ACG

[SMWS88] Eric L. Schwartz, Bjorn Merker, Estarose Wolfson, and Alan Shaw. Applications of computer graphics and image processing to 2D and 3D modeling of the functional architecture of visual cortex. IEEE Computer Graphics and Applications, 8(4):13–23, July/ August 1988. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Stern:1986:YIB

[SN86] Richard H. Stern and Margaret Neal. Are your images being ripped off? IEEE Computer Graphics and Applications, 6(3):45– 51, March/April 1986. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Smets:1995:TBR

[SO95] Gerda J. F. Smets and Kees J. Overbeeke. Trade off between resolution and interactivity in spatial task performance. IEEE Computer Graphics and Applications, 15(5):46–51, September 1995. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Swarz:2010:CSB

[SOM^{+}10] Jeff Swarz, Anita Ousley, Adriane Magro, Marie Rienzo, David Burns, A. M. Lindsey, Ben Wilburn, and Susan Bolcar. CancerSpace: a simulationbased game for improving cancer-screening rates. IEEE Computer Graphics and Applications, 30(1):90– 94, January/February 2010. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Suntinger:2008:ETE

[SOSG08] Martin Suntinger, Hannes Obweger, Josef Schiefer, and M. Eduard Gröller. Event tunnel: Exploring eventdriven business processes. IEEE Computer Graphics and Applications, 28(5): 46–55, September/October 2008. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Sourin:2004:NTU

[Sou04] Alexei Sourin. Nanyang Technological University Virtual Campus. IEEE Computer Graphics and Applications, 24(6):6–8, November/December 2004. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http:

//csdl.computer.org/dl/

                                        mags/cg/2004/06/g6006.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/cg/2004/06/
                                        g6006.pdf.

Sowizral:2000:SGN

[Sow00]  Henry Sowizral.  Scene

graphs in the new millennium. IEEE Computer Graphics and Applications, 20(1):56–57, January/ February 2000. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL http://dlib.
                                        computer.org/cg/books/
                                        cg2000/pdf/g1056.pdf.

Solachidis:2004:WPL

[SP04a] Vassilios Solachidis and Ioannis Pitas. Watermarking polygonal lines using Fourier descriptors. IEEE Computer Graphics and Applications, 24(3):44–51, May/June 2004. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL http://csdl.
                                        computer.org/comp/mags/
                                        cg/2004/03/g3044abs.htm;
                                         http://csdl.computer.
                                        org/dl/mags/cg/2004/03/
                                        g3044.htm;   http://csdl.
                                        computer.org/dl/mags/cg/
                                        2004/03/g3044.pdf.

Strickon:2004:GEI

[SP04b] Joshua Strickon and Joseph A. Paradiso. Guest Editors’ introduction: Emerging technologies at Siggraph 2003. IEEE Computer Graphics and Applications, 24(1):24–25, January/ February 2004. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cg/2004/01/g1024.pdf.

Stephenson:2006:SBU

[SP06] Peter D. Stephenson and Joan Peckham. Seeing is believing: Using computer graphics to enthuse students. IEEE Computer Graphics and Applications, 26(6):87–91, November/December 2006. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Shneiderman:2015:SAF

[SP15] Ben Shneiderman and Catherine Plaisant. Sharpening analytic focus to cope with big data volume and variety. IEEE Computer Graphics and Applications, 35(3):10–14, May/ June 2015. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
cg/2015/03/mcg2015030010-
abs.html.

Sikudova:2016:GMF

[SPA^{+}16a] Elena Sikudova, Tania Pouli, Alessandro Artusi, Ahmet Oguz Akyuz, Francesco Banterle, Zeynep Miray Mazlumoglu, and Erik Reinhard. A gamutmapping framework for color-accurate reproduction of HDR images. IEEE Computer Graphics and Applications, 36(4):78–90, July/ August 2016. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2016/04/mcg2016040078-
abs.html.

Spagnuolo:2016:SSM

[Spa16b] Michela Spagnuolo. Shape 4.0: 3D shape modeling and processing using semantics. IEEE Computer Graphics and Applications, 36(1):92–96, January/ February 2016. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
cg/2016/01/mcg2016010092-
abs.html.

Spiers:1986:RAI

[Spi86] Ray G. Spiers. The realization and application of an intelligent GKS workstation. IEEE Computer Graphics and Applications, 6(5):58–65, May/June 1986. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Sacks:1999:VCS

[SPJ99] Elisha Sacks, Charles Pisula, and Leo Joskowicz. Visualizing 3D configuration spaces for mechanical design. IEEE Computer Graphics and Applications, 19(5):50–53, September/ October 1999. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

computer.org/cga/cg1999/
g5050abs.htm;   http:/
/dlib.computer.org/cg/
books/cg1999/pdf/g5050.
pdf.

Spring:2010:UMR

[SPM10] Adam P. Spring, Caradoc Peters, and Tom Minns. Using mid-range laser scanners to digitize culturalheritage sites. IEEE Computer Graphics and Applications, 30(3):15–19, May/ June 2010. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Sato:2014:CID

[SPMK14] Toshiki Sato, Jefferson Pardomuan, Yasushi Matoba, and Hideki Koike. Claytric-Surface: An interactive deformable display with dynamic stiffness control. IEEE Computer Graphics and Applications, 34(3):59– 67, May/June 2014. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Sainz:2004:SAP

[SPMS04] Miguel Sainz, Renato Pajarola, Albert Mercade, and Antonio Susin. A simple approach for pointbased object capturing and rendering. IEEE Computer Graphics and Applications, 24(4):24–33, July/ August 2004. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http:/

                                        /csdl.computer.org/dl/
                                        mags/cg/2004/04/g4024.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/cg/2004/04/
                                        g4024.pdf.

Slater:1999:PVP

[SPS99] Mel Slater, David-Paul Pertaub, and Anthony Steed. Projects in VR: Public speaking in virtual reality: Facing an audience of avatars. IEEE Computer Graphics and Applications, 19(2):6–9, March/ April 1999. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL http://dlib.
                                        computer.org/cg/books/
                                        cg1999/pdf/g2006.pdf.

Santos:2015:TSU

[SPT^{+}15] Marc Ericson C. Santos, Jarkko Polvi, Takafumi Taketomi, Goshiro Yamamoto, Christian Sandor, and Hirokazu Kato. Toward standard usability questionnaires for handheld augmented reality. IEEE Computer Graphics and Applications, 35(5):66–75, September/October 2015. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
cg/2015/05/mcg2015050066-
abs.html.

Strogatz:1984:ESC

[SPW84] Steven H. Strogatz, Melvin L.

Prueitt, and Arthur T. Winfree. Exotic shapes in chemistry and biology. IEEE Computer Graphics and Applications, 4(1):66– 69, January/February 1984. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Stromer:2005:JHJ

[SQG^{+}05] Julie N. Stromer, Gerald T.

Quon, Paul M. K. Gordon, Andrei L. Turinsky, and Christoph W. Sensen. Jabiru: Harnessing Java 3D behaviors for device and display portability. IEEE Computer Graphics and Applications, 25(2):70–80, March/ April 2005. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Satterfield:1985:PGC

[SR85] Steven G. Satterfield and David F. Rogers. A procedure for generating contour lines from a B-spline surface.

IEEE Computer Graphics and Applications, 5(4):71– 75, April 1985. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Smith:1993:AMC

[SR93] Brian C. Smith and Lawrence A. Rowe. Algorithms for manipulating compressed images. IEEE Computer Graphics and Applications, 13(5):34–42, September 1993. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Shapira:1995:SBU

[SR95] Michal Shapira and Ari Rappoport. Shape blending using the star-skeleton representation. IEEE Computer Graphics and Applications, 15(2):44–50, March/ April 1995. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Sanchez-Reyes:1997:STN

[SR97] Javier Sánchez-Reyes. A simple technique for NURBS shape modification. IEEE Computer Graphics and Applications, 17(1):52–59, January/February 1997. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Stapleton:2010:MRI

[SR10] Christopher Stapleton and Jannick Rolland. Mixing realities at Ismar 2009: Scary and wondrous. IEEE Computer Graphics and Applications, 30(3):89–95, May/ June 2010. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Schmidt:2013:DFM

[SR13] Ryan Schmidt and Matt Ratto. Design-to-fabricate: Maker hardware requires Maker software. IEEE Computer Graphics and Applications, 33(6):26–34, November/December 2013. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Schmidt:2016:GBD

[SR16] Jens Schmidt and Stephan Rudolph. Graph-based design languages: a lingua franca for product design including abstract geometry. IEEE Computer Graphics and Applications, 36(5):88–93, September/ October 2016. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2016/05/mcg2016050088-
abs.html.

Sra:2023:ESP

[Sra23] Misha Sra. Enhancing the sense of presence in virtual reality. IEEE Computer Graphics and Applications, 43(4):90–96, July/ August 2023. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Shlyakhter:2001:RTM

[SRDT01] Ilya Shlyakhter, Max Rozenoer, Julie Dorsey, and Seth Teller. Reconstructing 3D tree models from instrumented photographs. IEEE Computer Graphics and Applications, 21(3):53–61, May/June 2001. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

computer.org/cga/cg2001/
g3053abs.htm;   http:/
/dlib.computer.org/cg/
books/cg2001/pdf/g3053.
pdf.

Schulz:1998:AES

[SRE98] Martin Schulz, Thomas Reuding, and Thomas Ertl. Analyzing engineering simulations in a virtual environment. IEEE Computer Graphics and Applications, 18(6):46–52, November/December 1998. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

computer.org/cga/cg1998/
g6046abs.htm;   http:/

/dlib.computer.org/cg/
books/cg1998/pdf/g6046.
pdf.

Shen:2006:IDD

[SRF^{+}06] Chia Shen, Kathy Ryall, Clifton Forlines, Alan Esenther, Frédéric D. Vernier, Katherine Everitt, Mike Wu, Daniel Wigdor, Meredith Ringel Morris, Mark Hancock, and Edward Tse. Informing the design of direct-touch tabletops. IEEE Computer Graphics and Applications, 26(5):36–46, September/October 2006. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Segall:1999:VVU

[SRJ99] Matthew D. Segall, Dan Z.

Reinstein, and Neil F. Johnson. Visualizing VHF ultrasound of the human cornea. IEEE Computer Graphics and Applications, 19(4):74– 82, July/August 1999. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

computer.org/cga/cg1999/
g4074abs.htm;   http:/
/dlib.computer.org/cg/
books/cg1999/pdf/g4074.
pdf.

Sung:1990:CEP

[SRK90] Hsien Ching Kelvin Sung, Greg Rogers, and William Kubitz. A critical evaluation of PEX. IEEE Computer Graphics and Applications, 10(6):65–75, November/December 1990. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Shamir:2006:GCI

[SRL06] Ariel Shamir, Michael Rubinstein, and Tomer Levinboim. Generating comics from 3D interactive computer graphics. IEEE Computer Graphics and Applications, 26(3):53–61, May/ June 2006. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Stevens:1983:USG

[SRS83] Albert Stevens, Bruce Roberts, and Larry Stead. The use of a sophisticated graphics interface in computer-assisted instruction. IEEE Computer Graphics and Applications, 3(2):25–28, 30–31, March/April 1983. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Sachs:1991:ATD

[SRS91] Emanuel Sachs, Andrew Roberts, and David Stoops. 3-Draw: a tool for designing 3D shapes. IEEE Computer Graphics and Applications, 11(6):18–26, November/December 1991. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Segal:1988:PPO

[SS88] Mark Segal and Carlo H. Sequin. Partitioning polyhedral objects into nonintersecting parts. IEEE Computer Graphics and Applications, 8(1):53–67, January/ February 1988. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Salisbury:1997:PVP

[SS97] J. Kenneth Salisbury and Mandayam A. Srinivasan. Projects in VR: Phantombased haptic interaction with virtual objects. IEEE Computer Graphics and Applications, 17(5):6–10, September/October 1997. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL

http://dlib.computer.org/
cg/books/cg1997/pdf/g5006.
pdf.

Schell:2001:PVD

[SS01] Jesse Schell and Joe Shochet.

Projects in VR: Designing interactive theme park rides. IEEE Computer Graphics and Applications, 21(4):11– 13, July/August 2001. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http:

//dlib.computer.org/cg/

books/cg2001/pdf/g4011.
pdf.

Steinicke:2015:SI

[SS15] Frank Steinicke and Wolfgang Stuerzlinger. Spatial interfaces. IEEE Computer Graphics and Applications, 35(4):6–7, July/ August 2015. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
cg/2015/04/mcg2015040006.
html.

Sahillioglu:2017:SBA

[SS17] Yusuf Sahillioglu and Metin Sezgin. Sketch-based articulated 3D shape retrieval. IEEE Computer Graphics and Applications, 37(6):88–101, November/ December 2017. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2017/06/mcg2017060088-
abs.html.

Schmalstieg:2019:UPR

[SS19] Dieter Schmalstieg and Andre Stork. Unified patterns for realtime interactive simulation in games and digital storytelling. IEEE Computer Graphics and Applications, 39(1):100–106, January/February 2019. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL https:

//ieeexplore.ieee.org/
document/8663513/.

Srinivasan:2020:HAW

[SS20] A. Srinivasan and J. Stasko.

How to ask what to say?: Strategies for evaluating natural language interfaces for data visualization. IEEE Computer Graphics and Applications, 40(4):96–103, July/August 2020. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Sedef:2006:RTF

[SSB06] Mert Sedef, Evren Samur, and Cagatay Basdogan. Real-time finite-element simulation of linear viscoelastic tissue behavior based on experimental data. IEEE Computer Graphics and Applications, 26(6): 58–68, November/December 2006. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Sheng:2011:CUR

[SSC^{+}11] Bin Sheng, Hanqiu Sun, Shunbin Chen, Xuehui Liu, and Enhua Wu. Colorization using the rotation-invariant feature space. IEEE Computer Graphics and Applications, 31(2):24–35, March/ April 2011. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Sakas:1995:PVR

[SSG95] Georgios Sakas, Lars-Arne Schreyer, and Marcus Grimm. Preprocessing and volume rendering of 3D ultrasonic data. IEEE Computer Graphics and Applications, 15(4):47–54, July/August 1995. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Slusallek:1998:CLS

[SSH^{+}98] Philipp Slusallek, Marc Stamminger, Wolfgang Heidrich, Jan-Christian Popp, and Hans-Peter Seidel. Composite lighting simulations with lighting network. IEEE Computer Graphics and Applications, 18(2):22– 31, March/April 1998. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http:

                                        //dlib.computer.org/cg/
                                        books/cg1998/pdf/g2022.
                                        pdf.

Sherman:2007:OST

[SSM^{+}07] William R. Sherman, Simon Su, Philip A. McDonald, Yi Mu, and Frederick Harris, Jr. Open-source tools for immersive environmental visualization. IEEE Computer Graphics and Applications, 27(2):88–91, March/ April 2007. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Suzuki:2003:VSI

[SSMH03] Yoshio Suzuki, Kazunori Sai, Nobuko Matsumoto, and Osamu Hazama. Visualization systems on the information-technology-based laboratory. IEEE Computer Graphics and Applications, 23(2):32–39, March/ April 2003. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://dlib.
computer.org/cg/books/
cg2003/pdf/g2032.pdf;
http://www.computer.org/
cga/cg2003/g2032abs.htm.

Steed:2012:BAT

[SSO^{+}12] Anthony Steed, William Steptoe, Wole Oyekoya, Fabrizio Pece, Tim Weyrich, Jan Kautz, Doron Friedman, Angelika Peer, Massimiliano Solazzi, Franco Tecchia, Massimo Bergamasco, and Mel Slater. Beaming: An asymmetric telepresence system. IEEE Computer Graphics and Applications, 32(6):10–17, November/December 2012. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Smallman:2001:IAD

[SSOC01] Harvey S. Smallman, Mark St. John, Heather M. Oonk, and Michael B. Cowen. Information availability in 2D and 3D displays. IEEE Computer Graphics and Applications, 21(5):51–57, September/ October 2001. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL http://dlib.
                                        computer.org/cg/books/
                                        cg2001/g5051abs.htm;
                                        http://dlib.computer.org/
                                        cg/books/cg2001/pdf/g5051.
                                        pdf.

Stansfield:1995:VBM

[SSRH95] S. Stansfield, D. Shawver, D. Rogers, and R. Hightower. VR blackboard: Mission visualization for planning and training. IEEE Computer Graphics and Applications, 15(5):12–14, September 1995. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Sourin:2000:PVV

[SSS00] Alexei Sourin, Olga Sourina, and Howe Tet Sen. Projects in VR: Virtual orthopedic surgery training. IEEE Computer Graphics and Applications, 20(3):6–9, May/June 2000. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL http://dlib.
                                        computer.org/cg/books/
                                        cg2000/pdf/g3006.pdf.

Santoni:2018:LCG

[SSTP18] Christian Santoni, Gabriele Salvati, Valentina Tibaldo, and Fabio Pellacini. LevelMerge: Collaborative game level editing by merging labeled graphs. IEEE Computer Graphics and Applications, 38(4):71–83, July/August 2018. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2018/04/mcg2018040071-
abs.html.

Schmalstieg:2007:MCA

[SSW^{+}07] Dieter Schmalstieg, Gerhard Schall, Daniel Wagner, István Barakonyi, Gerhard Reitmayr, Joseph Newman, and Florian Ledermann. Managing complex augmented reality models. IEEE Computer Graphics and Applications, 27(4): 48–57, July/August 2007. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Szekely:2015:GBM

[SSZP15] Eniko Szekely, Arnaud Sallaberry, Faraz Zaidi, and Pascal Poncelet. A graphbased method for detecting rare events: Identifying pathologic cells. IEEE Computer Graphics and Applications, 35(3):65–73, May/ June 2015. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
cg/2015/03/mcg2015030065-
abs.html.

Szafir:2021:ECV

[SSZS21] D. A. Szafir, F. Samsel, S. Zeller, and R. Saltus. Enabling crosscutting visualization for geoscience. IEEE Computer Graphics and Applications, 41(1):49– 57, January/February 2021. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Sprague:2009:VVM

[ST09] David W. Sprague and Melanie Tory. Visualization viewpoints: Motivation and procrastination: Methods for evaluating pragmatic casual information visualizations. IEEE Computer Graphics and Applications, 29(4):86–91, July/ August 2009. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Stone:2010:GH

[ST10] Maureen Stone and Gabriel Taubin. Goodbye, hello. IEEE Computer Graphics and Applications, 30(1):6– 7, January/February 2010. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Staudhammer:1991:CGH

[Sta91a] John Staudhammer. Computer graphics hardware. IEEE Computer Graphics and Applications, 11(1):42– 44, January/February 1991.

CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Staudhammer:1991:GEI

[Sta91b] John Staudhammer. Guest Editor’s introduction: computer graphics — toward the next millennium. IEEE Computer Graphics and Applications, 11(1):21–22, January/February 1991. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Stallmann:2017:AAG

[Sta17] Matthias F. Stallmann. Algorithm animation with Galant. IEEE Computer Graphics and Applications, 37(1):8–14, January/ February 2017. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2017/01/mcg2017010008-
abs.html.

Steinberg:1984:SSB

[Ste84] Herbert A. Steinberg. A smooth surface based on biquadratic patches. IEEE Computer Graphics and Applications, 4(9):20–23, September 1984. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Stewart:1991:VBM

[Ste91] W. Kenneth Stewart. Visualization blackboard — multisensor visualization for underwater archaeology. IEEE Computer Graphics and Applications, 11(2):13–18, March/April 1991. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Stevens:2007:VCA

[Ste07] Jeffrey A. Stevens. Visualization of complex automotive data: a tutorial. IEEE Computer Graphics and Applications, 27(6): 80–86, November/December 2007. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Steinberger:2018:DSG

[Ste18] Markus Steinberger. On dynamic scheduling for the GPU and its applications in computer graphics and beyond. IEEE Computer Graphics and Applications, 38(3):119–130, May/ June 2018. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2018/03/mcg2018030119-
abs.html.

State:2020:DUW

[STJ^{+}20] A. State, H. Towles, T. Johnson, R. Schubert, B. Walters, G. Welch, and H. Fuchs. The A-Desk: a unified workspace of the future. IEEE Computer Graphics and Applications, 40(1):56– 71, January/February 2020. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Stone:1999:GEI

[Sto99] Maureen Stone. Guest Editor’s introduction: Virtual Reality Modeling Language. IEEE Computer Graphics and Applications, 19(2):17, March/April 1999. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://dlib.
computer.org/cg/books/
cg1999/pdf/g2017.pdf.

Stone:2000:GW

[Sto00] Maureen C. Stone. The graphic Web. IEEE Computer Graphics and Applications, 20(1):58–59, January/ February 2000. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://dlib.
computer.org/cg/books/
cg2000/pdf/g1058.pdf.

Stone:2001:CBA

[Sto01] Maureen C. Stone. Color and brightness appearance issues in tiled displays. IEEE Computer Graphics and Applications, 21(5):58–66, September/ October 2001. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://dlib.
computer.org/cg/books/

                                        cg2001/g5058abs.htm;
                                        http://dlib.computer.org/
                                        cg/books/cg2001/pdf/g5058.
                                        pdf.

Stone:2005:RCT

[Sto05] Maureen C. Stone. Representing colors as three numbers. IEEE Computer Graphics and Applications, 25(4):78–85, July/August 2005. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Stone:2007:IIC

[Sto07] Maureen C. Stone. Introducing the IEEE CG&A Editorial Board. IEEE Computer Graphics and Applications, 27(4):6–7, July/ August 2007. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL http://csdl.
                                        computer.org/comp/mags/
                                        cg/2007/04/g4006.pdf.

Stone:2008:WB

[Sto08] Maureen Stone. Welcome, and beyond. IEEE Computer Graphics and Applications, 28(1):10–11, January/ February 2008. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL http://csdl.
                                        computer.org/comp/mags/
                                        cg/2008/01/mcg2008010010.
                                        pdf.

Stone:2012:CPS

[Sto12] Maureen Stone. In color perception, size matters. IEEE Computer Graphics and Applications, 32(2): 8–13, March/April 2012. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Stork:2015:VCC

[Sto15] Andre Stork. Visual computing challenges of advanced manufacturing and Industrie 4.0 [Guest Editors’ introduction]. IEEE Computer Graphics and Applications, 35(2):21–25, March/ April 2015. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
cg/2015/02/mcg2015020021.
html.

Stork:2021:CGR

[Sto21] André Stork. Computer graphics: From research to commodity to new highs [graphically speaking]. IEEE Computer Graphics and Applications, 41(4): 16–19, July/August 2021. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Stork:2023:EEa

[Sto23a] André Stork. EIC’s editorial. IEEE Computer Graphics and Applications, 43 (2):8–10, March/April 2023.

CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Stork:2023:EEb

[Sto23b] André Stork. EIC’s editorial. IEEE Computer Graphics and Applications, 43(3):8–9, May/June 2023. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Stork:2023:EEc

[Sto23c] André Stork. EIC’s editorial. IEEE Computer Graphics and Applications, 43(4): 10–11, July/August 2023. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Stork:2023:EEd

[Sto23d] André Stork. EIC’s editorial. IEEE Computer Graphics and Applications, 43(5): 8–11, September/October 2023. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Stork:2023:EEe

[Sto23e] André Stork. EIC’s editorial. IEEE Computer Graphics and Applications, 43(6): 7–8, November/December 2023. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Stork:2023:GNE

[Sto23f] André Stork. Greetings from the new EIC. IEEE Computer Graphics and Applications, 43(1):7–9, January/ February 2023. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Stork:2024:EEa

[Sto24a] André Stork. EIC’s editorial. IEEE Computer Graphics and Applications, 44 (1):8–9, January/February 2024. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Stork:2024:EEb

[Sto24b] André Stork. EIC’s editorial. IEEE Computer Graphics and Applications, 44 (2):8–9, March/April 2024. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Strauss:1990:RLM

[Str90] Paul S. Strauss. A realistic lighting model for computer animators. IEEE Computer Graphics and Applications, 10(6):56–64, November/December 1990. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Strickon:2004:IET

[Str04]  Joshua Strickon.  Inter-

acting with emerging technologies. IEEE Computer Graphics and Applications, 24(1):C2, January/ February 2004. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cg/2004/01/g10c2.pdf.

Sturman:1998:CP

[Stu98] David J. Sturman. Computer puppetry. IEEE Computer Graphics and Applications, 18(1):38–45, January/ February 1998. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

computer.org/cga/cg1998/
g1038abs.htm;   http:/
/dlib.computer.org/cg/
books/cg1998/pdf/g1038.
pdf.

Shi:2008:VTS

[STW^{+}08] Kuangyu Shi, Holger Theisel, Tino Weinkauf, Hans-Christian Hege, and Hans-Peter Seidel. Visualizing transport structures of timedependent flow fields. IEEE Computer Graphics and Applications, 28(5):24–36, September/October 2008. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Stytz:1996:DVE

[Sty96] Martin R. Stytz. Distributed virtual environments. IEEE Computer Graphics and Applications, 16(3):19–31, May/June 1996. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Sugihara:1992:IAB

[Sug92] Kokichi Sugihara. An intersection algorithm based on Delaunay triangulation. IEEE Computer Graphics and Applications, 12(2): 59–67, March/April 1992. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Sun:2021:LHV

[Sun21] Qi Sun. Leveraging human visual perception for an optimized virtual reality experience. IEEE Computer Graphics and Applications, 41(6):164–170, November/ December 2021. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Stytz:1997:SSM

[SVB97] Martin R. Stytz, John Vanderburgh, and Sheila B. Banks. The solar system modeler. IEEE Computer Graphics and Applications, 17(5):47–57, September/October 1997. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://dlib.
computer.org/cg/books/
cg1997/pdf/g5047.pdf.

Salzman:1989:GSL

[SvNG89] David Salzman, John von Neumann, and Jack Grimes.

Graphics superworkstations and the last hurrah. IEEE Computer Graphics and Applications, 9(4):27–29, July/ August 1989. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Sherstyuk:2010:TNS

[SVT10] Andrei Sherstyuk, Dale Vincent, and Anton Treskunov. Toward natural selection in virtual reality. IEEE Computer Graphics and Applications, 30(3):93–96,C3, May/ June 2010. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Shrinivasan:2009:SEA

[SvW09] Yedendra Babu Shrinivasan and Jarke van Wijk. Supporting exploration awareness in information visualization. IEEE Computer Graphics and Applications, 29(5):34–43, September/October 2009. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Sabella:1983:TFC

[SW83] Paolo Sabella and Michael J. Wozny. Toward fast colorshaded images of CAD/ CAM geometry. IEEE Computer Graphics and Applications, 3(8):60–71, November/December 1983. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Sequin:1985:VFR

[SW85] Carlo H. Sequin and Paul R.

Wensley. Visible feature return at object resolution. IEEE Computer Graphics and Applications, 5(5):37– 50, May/June 1985. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Samet:1988:HDS

[SW88] Hanan Samet and Robert E.

Webber. Hierarchical data structures and algorithms for computer graphics — II: Applications. IEEE Computer Graphics and Applications, 8(4):59–75, July/ August 1988. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Sun:2022:DVD

[SW22] Cheng Sun and Ko-Chih Wang. DLA-VPS: Deeplearning-assisted visual parameter space analysis of cosmological simulations. IEEE Computer Graphics and Applications, 42(3):41– 52, May/June 2022. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Smith:1998:EFS

[SWD^{+}98] Joshua Smith, Tom White, Christopher Dodge, Joseph Paradiso, Neil Gershenfeld, and David Allport. Electric field sensing for graphical interfaces. IEEE Computer Graphics and Applications, 18(3):54–60, May/ June 1998. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

computer.org/cga/cg1998/
g3054abs.htm;   http:/
/dlib.computer.org/cg/
books/cg1998/pdf/g3054.
pdf.

Schubert:2016:HDS

[SWDR16] Ryan Schubert, Greg Welch, Salam Daher, and Andrew Raij. HuSIS: A dedicated space for studying human interactions. IEEE Computer Graphics and Applications, 36(6):26–36, November/December 2016. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL https:/

/www.computer.org/csdl/
mags/cg/2016/06/mcg2016060026-
abs.html.

Sweeney:2001:AWW

[Swe01] Tony Sweeney. Applications: Wired worlds: Exploring the digital frontier: a pioneering gallery of digital media discovery. IEEE Computer Graphics and Applications, 21(1):72–77, January/ February 2001. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://dlib.

computer.org/cg/books/
cg2001/pdf/g1072.pdf.

Swing:2010:PAA

[Swi10] Edward Swing. Prajna:

Adding automated reasoning to the visual-analysis process. IEEE Computer Graphics and Applications, 30(1):50–58, January/ February 2010. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Schonhut:1991:CPD

[SWS91] Jurgen Schönhut, Hans-Peter Wiedling, and Veronika Samara. Constructive page description. IEEE Computer Graphics and Applications, 11(2):71–78, March/ April 1991. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Seeling:2007:GBI

[SWS07] Chris Seeling, Greg Watson, and Kaiwei Sun. GPUbased interactive, stereoscopic visualization of automotive crash simulations. IEEE Computer Graphics and Applications, 27(6): 6–11, November/December 2007. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Scheepens:2012:IDM

[SWvdWvW12] Roeland Scheepens, Niels Willems, Huub van de Wetering, and Jarke J. van Wijk. Interactive density maps for moving objects. IEEE Computer Graphics and Applications, 32(1):56– 66, January/February 2012. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Sturman:1994:SGI

[SZ94] David J. Sturman and David Zeltzer. A survey of glovebased input. IEEE Computer Graphics and Applications, 14(1):30–39, January/ February 1994. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Szeliski:1996:VMV

[Sze96]  Richard Szeliski.  Video

mosaics for virtual environments. IEEE Computer Graphics and Applications, 16(2):22–30, March/ April 1996. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Shaverdian:2012:GAS

[SZMJ12] Anna A. Shaverdian, Hao Zhou, George Michailidis, and Hosagrahar V. Jagadish. A graph algebra for scalable visual analytics. IEEE Computer Graphics and Applications, 32(4): 26–33, July/August 2012. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Sharma:2000:SGI

[SZP^{+}00] Rajeev Sharma, Michael Zeller, Vladimir I. Pavlovic, Thomas S. Huang, Zion Lo, Stephen Chu, Yunxin Zhao, James C. Phillips, and Klaus Schulten. Speech/ gesture interface to a visualcomputing environment. IEEE Computer Graphics and Applications, 20(2):29– 37, March/April 2000. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

computer.org/cga/cg2000/
g2029abs.htm;   http:/
/dlib.computer.org/cg/
books/cg2000/pdf/g2029.
pdf.

Shi:2018:NRV

[SZZ^{+}18] Yang Shi, Ying Zhao, Fangfang Zhou, Ronghua Shi, and Yaoxue Zhang. A novel radial visualization of intrusion detection alerts. IEEE Computer Graphics and Applications, 38(6):83– 95, ???? 2018. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL https:

//ieeexplore.ieee.org/
document/8617765/.

Toga:1985:IAB

[TA85] Arthur W. Toga and Tamara L.

Arnicar. Image analysis of brain physiology. IEEE Computer Graphics and Applications, 5(12):20–25, December 1985. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Tachi:1998:PVR

[Tac98] Susumu Tachi. Projects in VR: Real-time remote robotics — toward networked telexistence. IEEE Computer Graphics and Applications, 18(6):6–9, November/December 1998. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http:

//dlib.computer.org/cg/
books/cg1998/pdf/g6006.
pdf.

Tachi:2016:TEH

[Tac16] Susumu Tachi. Telexistence: Enabling humans to be virtually ubiquitous. IEEE Computer Graphics and Applications, 36(1):8–14, January/February 2016. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http:/

/www.computer.org/csdl/
mags/cg/2016/01/mcg2016010008-
abs.html.

Takala:1985:UIM

[Tak85] Tapio Takala. User interface management system with geometric modeling capability: A CAD system’s framework. IEEE Computer Graphics and Applications, 5(4):42–50, April 1985. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Takacs:2005:SER

[Tak05] Barnabas Takacs. Special education and rehabilitation: Teaching and healing with interactive graphics. IEEE Computer Graphics and Applications, 25(5): 40–48, September/October 2005. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Talbert:1997:THS

[Tal97] Nancy Talbert. Toward human-centered systems. IEEE Computer Graphics and Applications, 17(4): 21–28, July/August 1997. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Taubin:1994:RAC

[Tau94] Gabriel Taubin. Rasterizing algebraic curves and surfaces. IEEE Computer Graphics and Applications, 14(2):14–23, March/April 1994. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Taubin:2010:CNA

[Tau10a] Gabriel Taubin. CG&A’s new Associate Editor in Chief. IEEE Computer Graphics and Applications, 30(4):6–7, July/August 2010. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Taubin:2010:EBA

[Tau10b]  Gabriel Taubin.  Edito-

rial Board: Additions and subtractions. IEEE Computer Graphics and Applications, 30(5):6, September/ October 2010. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Taubin:2010:ODC

[Tau10c] Gabriel Taubin. Our departments: Charting the path of computer graphics. IEEE Computer Graphics and Applications, 30(3):6–7, May/ June 2010. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Taubin:2011:R

[Tau11a] Gabriel Taubin. 3D rotations. IEEE Computer Graphics and Applications, 31(6):84–89, November/ December 2011. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Taubin:2011:CGC

[Tau11b] Gabriel Taubin. Comings, goings, and a contest. IEEE Computer Graphics and Applications, 31(5):3, September/October 2011. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Taubin:2011:NDP

[Tau11c] Gabriel Taubin. A new department and a panel discussion. IEEE Computer Graphics and Applications, 31(6):3, November/ December 2011. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Taubin:2011:NED

[Tau11d] Gabriel Taubin. The new Education Department and Editorial Board members. IEEE Computer Graphics and Applications, 31(1):5– 6, January/February 2011. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Taubin:2012:FLR

[Tau12a] Gabriel Taubin. Farewell to Larry Rosenblum. IEEE Computer Graphics and Applications, 32(3):3, May/ June 2012. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Taubin:2012:IGP

[Tau12b] Gabriel Taubin. Introduction to geometric processing through optimization. IEEE Computer Graphics and Applications, 32(4): 88–94, July/August 2012. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Taubin:2012:NDE

[Tau12c]  Gabriel Taubin.  New

department, editors, and awards. IEEE Computer Graphics and Applications, 32(4):6–7, July/ August 2012. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Taubin:2012:NY

[Tau12d] Gabriel Taubin. A new year. IEEE Computer Graphics and Applications, 32(1): 6, January/February 2012. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Taubin:2012:TCO

[Tau12e] Gabriel Taubin. Thanks to Carol O’Sullivan. IEEE Computer Graphics and Applications, 32(6):3, November/December 2012. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Taubin:2013:INA

[Tau13a] Gabriel Taubin. Introducing a new Associate Editor in Chief. IEEE Computer Graphics and Applications, 33(2):3, March/April 2013. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Taubin:2013:INE

[Tau13b] Gabriel Taubin. Introducing the new Editor in Chief. IEEE Computer Graphics and Applications, 33(5):4– 5, September/October 2013. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Taylor:2002:VVV

[Tay02] Russell Taylor. Visualization viewpoints: Visualizing multiple fields on the same surface. IEEE Computer Graphics and Applications, 22(3):6–10, May/ June 2002. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

computer.org/cga/cg2002/
g3006abs.htm;   http:/
/dlib.computer.org/cg/
books/cg2002/pdf/g3006.
pdf.

Tominski:2023:VUS

[TBB^{+}23] Christian Tominski, Michael Behrisch, Susanne Bleisch, Sara Irina Fabrikant, Eva Mayr, Silvia Miksch, and Helen Purchase. Visualizing uncertainty in sets. IEEE Computer Graphics and Applications, 43(5): 49–61, September/October 2023. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Tory:2023:FTD

[TBFGC23] Melanie Tory, Lyn Bartram, Brittany Fiore-Gartland, and Anamaria Crisan. Finding their data voice: Practices and challenges of dashboard users. IEEE Computer Graphics and Applications, 43(1):22–36, January/ February 2023. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Turton:2020:EDE

[TBO^{+}20] T. L. Turton, D. Banesh, T. Overmyer, B. H. Sims, and D. H. Rogers. Enabling domain expertise in scientific visualization with CinemaScience. IEEE Computer Graphics and Applications, 40(1):90–98, January/ February 2020. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Tiddeman:2001:PTF

[TBP01] Bernard Tiddeman, Michael Burt, and David Perrett. Prototyping and transforming facial textures for perception research. IEEE Computer Graphics and Applications, 21(5):42–50, September/ October 2001. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://dlib.
computer.org/cg/books/
cg2001/g5042abs.htm;

http://dlib.computer.org/
cg/books/cg2001/pdf/g5042.
pdf.

Treinish:1996:VBW

[TC96] L. A. Treinish and Z. D.

Christidis. Visualization blackboard: Weather forecasting for the 1996 Olympics. IEEE Computer Graphics and Applications, 16(4):10–13, July/ August 1996. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Thomas:2006:VAA

[TC06] James J. Thomas and Kristin A. Cook. A visual analytics agenda. IEEE Computer Graphics and Applications, 26(1):10–13, January/February 2006. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Tory:2015:PVP

[TC15] Melanie Tory and Sheelagh Carpendale. Personal visualization and personal visual analytics [Guest Editors’ introduction]. IEEE Computer Graphics and Applications, 35(4):26–27, July/ August 2015. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
cg/2015/04/mcg2015040026.
html.

Teller:2003:PPA

[TCB03] Seth Teller, Jiawen Chen, and Hari Balakrishnan. Pervasive pose-aware applications and infrastructure. IEEE Computer Graphics and Applications, 23(4):14– 18, July/August 2003. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http:

//csdl.computer.org/dl/
mags/cg/2003/04/g4014.
htm;   http://csdl.computer.
org/dl/mags/cg/2003/04/
g4014.pdf.

Tsingos:2002:VAS

[TCE^{+}02] Nicolas Tsingos, Ingrid Carlbom, Gary Elbo, Robert Kubli, and Thomas Funkhouser. Validating acoustical simulations in Bell Labs Box. IEEE Computer Graphics and Applications, 22(4):28– 37, July/August 2002. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

computer.org/cga/cg2002/
g4028abs.htm;   http:/
/dlib.computer.org/cg/
books/cg2002/pdf/g4028.
pdf.

Taylor:1996:IRC

[TCH^{+}96] Valerie E. Taylor, Jian Chen, Milana Huang, Thomas Canfield, and Rick Stevens. Identifying and reducing critical lag in finite element simulations. IEEE Computer Graphics and Applications, 16(4):67–71, July/ August 1996. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Theisel:1997:CCC

[TF97] Holger Theisel and Gerald Farin. The curvature of characteristic curves on surfaces. IEEE Computer Graphics and Applications, 17(6):88–96, November/December 1997. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

computer.org/cga/cg1997/
g6088abs.htm.

Trumblin:2001:GEI

[TF01] Jack Trumblin and James A.

Ferwerda. Guest Editors’ introduction: Applied perception. IEEE Computer Graphics and Applications, 21(5):20–21, September/October 2001. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://dlib.
computer.org/cg/books/
cg2001/g5020abs.htm;
http://dlib.computer.org/
cg/books/cg2001/pdf/g5020.
pdf.

Thatte:2021:RWV

[TG21] Jayant Thatte and Bernd Girod. Real-world virtual reality with head-motion parallax. IEEE Computer Graphics and Applications, 41(4):29–39, July/August 2021. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Tiller:1984:OTP

[TH84] Wayne Tiller and Eric G. Hanson. Offsets of twodimensional profiles. IEEE Computer Graphics and Applications, 4(9):36–46, September 1984. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Titsworth:2001:ACE

[TH01] Tammi Titsworth and Susana Maria Halpine. About the cover: Explaining the brave new world of information. IEEE Computer Graphics and Applications, 21(4):4–5, July/ August 2001. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL http://dlib.
                                        computer.org/cg/books/
                                        cg2001/pdf/g4004.pdf.

Taylor:2014:RPE

[TH14] Russell M. Taylor II and Jonathan Harter. Random per-element luminance modulation for improved visual tracking. IEEE Computer Graphics and Applications, 34(6):83–87, November/ December 2014. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL http://csdl.

computer.org/csdl/mags/
cg/2014/06/mcg2014060083-
abs.html.

Traore:2017:IES

[TH17] Michael Traore and Christophe Hurter. Interactive exploration of 3D scanned baggage. IEEE Computer Graphics and Applications, 37(1):27–33, January/ February 2017. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2017/01/mcg2017010027-
abs.html.

Tiede:1990:IMA

[THB^{+}90] Ulf Tiede, Karl Heinz Hoehne, Michael Bomans, Andreas Pommert, Martin Riemer, and Gunnar Wiebecke. Investigation of medical 3D-rendering algorithms. IEEE Computer Graphics and Applications, 10(2):41–53, March/April 1990. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Thompson:1984:MAC

[Tho84] Andrew V. Thompson.

Micro-Cap: an analog circuit design system for personal computers. IEEE Computer Graphics and Applications, 4(4):30–32, April 1984. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Thomas:2000:ICG

[Tho00] James J. Thomas. The impact of computer graphics. IEEE Computer Graphics and Applications, 20(1): 17, January/February 2000. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL

                                        http://dlib.computer.org/
                                        cg/books/cg2000/pdf/g1017.
                                        pdf.

Thomas:2001:ECM

[Tho01] James J. Thomas. Editor in Chief’s message. IEEE Computer Graphics and Applications, 21(3):0, May/ June 2001. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL http://dlib.
                                        computer.org/cg/books/
                                        cg2001/pdf/g3c2.pdf.

Thomas:2002:ECM

[Tho02a] James J. Thomas. Editorin-Chief’s message: Welcome aboard. IEEE Computer Graphics and Applications, 22(3):C2, May/ June 2002. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

                                        computer.org/cga/cg2002/
                                        g30c2abs.htm;   http:/
                                        /dlib.computer.org/cg/
                                        books/cg2002/pdf/g30c2.
                                        pdf.

Thomas:2002:EMC

[Tho02b] James J. Thomas. EIC’s message: Changing of the guard. IEEE Computer Graphics and Applications, 22(6):c2, November/ December 2002. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://dlib.
computer.org/cg/books/
cg2002/pdf/g60c2.pdf.

Tilove:1981:LPC

[Til81] Robert B. Tilove. Line/ polygon classification: a study of the complexity of geometric computation. IEEE Computer Graphics and Applications, 1(2):75– 78, 80–82, 84, April 1981. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Tiller:1983:RBC

[Til83a] Wayne Tiller. Rational Bsplines for curve and surface representation. IEEE Computer Graphics and Applications, 3(6):61–64, 66–69, September 1983. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Tilove:1983:ESM

[Til83b] Robert Bruce Tilove. Extending solid modeling systems for mechanism design and kinematic simulation. IEEE Computer Graphics and Applications, 3(3):9–19, May/June 1983. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Titsworth:2007:VSS

[Tit07]  Frances Titsworth.  Vi-

sual symphonies: Setting intricate images to music. IEEE Computer Graphics and Applications, 27(6):4–5, November/December 2007. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Takacs:2003:VHI

[TK03] Barnabás Takács and Bernadette Kiss. The virtual human interface: a photorealistic digital human. IEEE Computer Graphics and Applications, 23(5):38–45, September/October 2003. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http:

//csdl.computer.org/dl/
mags/cg/2003/05/g5038.
pdf;   http://csdl.computer.
org/dl/mags/cg/2003/05/
g5038abs.htm.

Tecchia:2002:IBC

[TLC02] Franco Tecchia, Céline Loscos, and Yiorgos Chrysanthou. Image-based crowd rendering. IEEE Computer Graphics and Applications, 22(2):36–43, March/ April 2002. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

computer.org/cga/cg2002/
g2036abs.htm;   http:/
/dlib.computer.org/cg/
books/cg2002/pdf/g2036.
pdf.

Thudt:2017:ERM

[TLCC17] Alice Thudt, Bongshin Lee, Eun Kyoung Choe, and Sheelagh Carpendale. Expanding research methods for a realistic understanding of personal visualization. IEEE Computer Graphics and Applications, 37(2):12–18, March/ April 2017. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2017/02/mcg2017020012-
abs.html.

Tebo:1996:ODF

[TLL^{+}96] Scot A. Tebo, Donald A.

Leopold, Donlin M. Long, S. James Zinreich, and David W. Kennedy. An optical 3D digitizer for frameless stereotactic surgery. IEEE Computer Graphics and Applications, 16(1):55– 64, January/February 1996. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Turkay:2014:CCS

[TLS^{+}14] Cagatay Turkay, Alexander Lex, Marc Streit, Hanspeter Pfister, and Helwig Hauser. Characterizing cancer subtypes using dual analysis in Caleydo StratomeX. IEEE Computer Graphics and Applications, 34(2): 38–47, March/April 2014. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Tan:2011:SPD

[TLT11] Jie Tan, Karen Liu, and Greg Turk. Stable proportionalderivative controllers. IEEE Computer Graphics and Applications, 31(4):34–44, July/August 2011. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Thatch:1988:PBG

[TM88] B. R. Thatch and A. Myklebust. A PHIGS-based graphics input interface for spatial-mechanism design. IEEE Computer Graphics and Applications, 8(2):26– 38, March/April 1988. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http:/

                                        /www.computer.org/csdl/
                                        mags/cg/02/00000501-abs.
                                        html;   http://www.computer.
                                        org/csdl/mags/cg/02/mcg1988020026-
                                        abs.html.

Tory:2005:EVD

[TM05] Melanie Tory and Torsten Moller. Evaluating visualizations: Do expert reviews work? IEEE Computer Graphics and Applications, 25(5):8–11, September/October 2005. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Tena:2013:FFP

[TMB^{+}13] J. Rafael Tena, Moshe Mahler, Thabo Beeler, Max Grosse, Hengchin Yeh, and Iain Matthews. Fabricating 3D figurines with personalized faces. IEEE Computer Graphics and Applications, 33(6):36–46, November/December 2013. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Teoh:2004:DFI

[TMWJK04] Soon Tee Teoh, Kwan-Liu Ma, Soon Felix Wu, and T. J. Jankun-Kelly. Detecting flaws and intruders with visual data analysis. IEEE Computer Graphics and Applications, 24(5):27–35, September/ October 2004. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http:

//csdl.computer.org/dl/
mags/cg/2004/05/g5027.
htm;   http://csdl.computer.
org/dl/mags/cg/2004/05/
g5027.pdf.

Tsetserukou:2010:IAE

[TN10] Dzmitry Tsetserukou and Alena Neviarouskaya. iFeel IM!: Augmenting emotions during online communication. IEEE Computer Graphics and Applications, 30(5): 72–80, September/October 2010. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Tamura:2007:SBA

[TNTM07] Nobuhiko Tamura, Toshiya Nakaguchi, Norimichi Tsumura, and Yoichi Miyake. Springbead animation of viscoelastic materials. IEEE Computer Graphics and Applications, 27(6):87–93, November/December 2007. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Tudhope:1983:HRS

[TO83] Douglas S. Tudhope and John V. Oldfield. A high-level recognizer for schematic diagrams. IEEE Computer Graphics and Applications, 3(3):33–38, 40, May/June 1983. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Tonti:2024:LDC

[TPA24] Claudia Melis Tonti, Lorenzo Papa, and Irene Amerini. Lightweight 3-D convolutional occupancy networks for virtual object reconstruction. IEEE Computer Graphics and Applications, 44(2):23–36, March/ April 2024. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Tovanich:2022:VCC

[TPR^{+}22] Natkamon Tovanich, Alexis Pister, Gaëlle Richer, Paola Valdivia, Christophe Prieur, Jean-Daniel Fekete, and Petra Isenberg. VAST 2020 contest challenge: Graph-MatchMaker visual analytics for graph comparison and matching. IEEE Computer Graphics and Applications, 42(4):89–102, July/ August 2022. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Teter:1983:PVK

[TR83] James Teter and Donna Roberts. Production volume: the key to successful slide-making systems. IEEE Computer Graphics and Applications, 3(4):27–30, July/ August 1983. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Tumblin:1993:TRR

[TR93] Jack Tumblin and Holly E.

Rushmeier. Tone reproduction for realistic images. IEEE Computer Graphics and Applications, 13(6): 42–48, November/December 1993. CODEN ICGADZ.

ISSN 0272-1716 (print), 1558-1756 (electronic). also appeared as Tech. Report GIT-GVU-91-13, Graphics, Visualization & Usability Center, Coll. of Computing, Georgia Institute of Tech.

Turner:1994:GEI

[TR94] Joshua Turner and Jarek Rossignac. Guest Editor’s introduction: Solid modeling. IEEE Computer Graphics and Applications, 14 (2):13, March/April 1994. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Treinish:1995:VSM

[Tre95] Lloyd A. Treinish. Visualization of scattered meteorological data. IEEE Computer Graphics and Applications, 15(4):20–26, July/ August 1995. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Treinish:1998:VBF

[Tre98] Lloyd Treinish. Visualization blackboard: Flodar: Flow visualization of network traffic. IEEE Computer Graphics and Applications, 18(5):6–8, September/October 1998. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL http://dlib.
                                        computer.org/cg/books/
                                        cg1998/pdf/g5006.pdf.

Treinish:1999:TSV

[Tre99] Lloyd A. Treinish. Taskspecific visualization design. IEEE Computer Graphics and Applications, 19(5):72–77, September/ October 1999. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

computer.org/cga/cg1999/
g5072abs.htm;   http:/
/dlib.computer.org/cg/
books/cg1999/pdf/g5072.
pdf.

Treinish:1997:VBW

[TS97] Lloyd Treinish and Deborah Silver. Visualization blackboard: Worm plots. IEEE Computer Graphics and Applications, 17(6):17–20, November/ December 1997. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://dlib.
computer.org/cg/books/
cg1997/pdf/g6017.pdf.

Tumen:2013:DTS

[TS13] R. Sinan Tumen and T. Metin Sezgin. DPFrag: Trainable stroke fragmentation based on dynamic programming. IEEE Computer Graphics and Applications, 33(5):59–67, September/October 2013. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Tumen:2022:SRO

[TS22] Recep Sinan Tümen and Metin Sezgin. Segmentation and recognition of offline sketch scenes using dynamic programming. IEEE Computer Graphics and Applications, 42(1):56–72, January/ February 2022. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Tromp:1998:SGB

[TSF^{+}98] Jolanda Tromp, Anthony Steed, Emmanuel Frécon, Adrian Bullock, Amela Sadagic, and Mel Slater. Small group behavior experiments in the Coven Project. IEEE Computer Graphics and Applications, 18(6):53–63, November/ December 1998. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

computer.org/cga/cg1998/
g6053abs.htm;   http:/
/dlib.computer.org/cg/
books/cg1998/pdf/g6053.
pdf.

Tiede:1996:VBV

[TSH96] Ulf Tiede, Thomas Schiemann, and Karl Heinz Hohne. Visualization blackboard: visualizing the visible human. IEEE Computer Graphics and Applications, 16(1):7–9, January/ February 1996. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Tate:1997:VRU

[TSK97] David L. Tate, Linda Sibert, and Tony King. Virtual reality: Using virtual environments to train firefighters. IEEE Computer Graphics and Applications, 17(6):23–29, November/December 1997. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

computer.org/cga/cg1997/
g6023abs.htm.

Tan:2012:IIC

[TSLX12] Li Tan, Yangqiu Song, Shixia Liu, and Lexing Xie. ImageHive: Interactive content-aware image summarization. IEEE Computer Graphics and Applications, 32(1):46–55, January/ February 2012. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Toriya:1986:URO

[TSUC86] Hiroshi Toriya, Toshiaki Satoh, Kenji Ueda, and Hiroaki Chiyokura. Undo and redo operations for solid modeling. IEEE Computer Graphics and Applications, 6(4):35–42, April 1986. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Tuy:1984:DDO

[TT84] H. K. Tuy and L. T. Tuy. Direct 2-D display of 3-D objects. IEEE Computer Graphics and Applications, 4(10):29–33, October 1984. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Tominaga:2000:ERP

[TT00] Shoji Tominaga and Norihiro Tanaka. Estimating reflection parameters from a single color image. IEEE Computer Graphics and Applications, 20(5):58–66, September/ October 2000. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

                                        computer.org/cga/cg2000/
                                        g5058abs.htm;   http:/
                                        /dlib.computer.org/cg/
                                        books/cg2000/pdf/g5058.
                                        pdf.

Toennies:1990:ROS

[TUH^{+}90] Klaus D. Toennies, Jayaram K. Udupa, Gabor T. Herman, Isaac L. Wornom III, and Steven R. Buchman. Registration of 3D objects and surfaces. IEEE Computer Graphics and Applications, 10(3):52–62, May/ June 1990. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Turner:1988:ASM

[Tur88] Joshua U. Turner. Accurate solid modeling using polyhedral approximations. IEEE Computer Graphics and Applications, 8(3):14–28, May/ June 1988. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Turnage:2002:RAN

[Tur02a]  Anna Turnage.  Reduc-

ing aircraft noise with computer graphics. IEEE Computer Graphics and Applications, 22(3):16–21, May/ June 2002. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

computer.org/cga/cg2002/
g3016abs.htm;   http:/
/dlib.computer.org/cg/
books/cg2002/pdf/g3016.
pdf.

Turner:2002:PCD

[Tur02b] Alan Turner. The premier CG&A digital media issue. IEEE Computer Graphics and Applications, 22(4):c2, July/August 2002. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

computer.org/cga/cg2002/
g40c2abs.htm;   http:/
/dlib.computer.org/cg/
books/cg2002/pdf/g4c2.
pdf.

Turnage:2003:MSB

[Tur03] Anna Turnage. Modeling supernovae: Braving a bold new frontier. IEEE Computer Graphics and Applications, 23(6):6–11, November/December 2003. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http:

                                        //csdl.computer.org/dl/
                                        mags/cg/2003/06/g6006.
                                        pdf.

Terzopoulos:1988:PBM

[TW88] Demetri Terzopoulos and Andrew Witkin. Physically based models with rigid and deformable components. IEEE Computer Graphics and Applications, 8(6):41–51, November/December 1988. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Turquin:2007:SBI

[TWB^{+}07] Emmanuel Turquin, Jamie Wither, Laurence Boissieux, Marie-Paule Cani, and John F. Hughes. A sketchbased interface for clothing virtual characters. IEEE Computer Graphics and Applications, 27(1):72–81, January/February 2007. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Tu:2023:IKL

[TWQ^{+}23] Yamei Tu, Xiaoqi Wang, Rui Qiu, Han-Wei Shen, Michelle Miller, Jinmeng Rao, Song Gao, Patrick R. Huber, Allan D. Hollander, Matthew Lange, Christian R. Garcia, and Joe Stubbs. An interactive knowledge and learning environment in smart foodsheds. IEEE Computer Graphics and Applications, 43(3):36–47, May/ June 2023. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Tamura:2001:MRF

[TYK01] Hideyuki Tamura, Hiroyuki Yamamoto, and Akihiro Katayama. Mixed reality: Future dreams seen at the border between real and virtual worlds. IEEE Computer Graphics and Applications, 21(6):64–70, November/ December 2001. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://dlib.
computer.org/cg/books/
cg2001/g6064abs.htm;
http://dlib.computer.org/
cg/books/cg2001/pdf/g6064.
pdf.

Urschler:2012:FCA

[UBS^{+}12] Martin Urschler, Alexander Bornik, Eva Scheurer, Kathrin Yen, Horst Bischof, and Dieter Schmalstieg.

Forensic-case analysis: From 3D imaging to interactive visualization. IEEE Computer Graphics and Applications, 32(4):79–87, July/ August 2012. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Uyttendaele:2004:IBI

[UCK^{+}04] Matthew Uyttendaele, Antonio Criminisi, Sing Bing Kang, Simon Winder, Richard Szeliski, and Richard Hartley. Image-based interactive exploration of real-world environments. IEEE Computer Graphics and Applications, 24(3):52–63, May/ June 2004. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL http://csdl.
                                        computer.org/comp/mags/
                                        cg/2004/03/g3052abs.htm;
                                         http://csdl.computer.
                                        org/dl/mags/cg/2004/03/
                                        g3052.htm;   http://csdl.
                                        computer.org/dl/mags/cg/
                                        2004/03/g3052.pdf.

Ullrich:2011:LAS

[UF11] Torsten Ullrich and Dieter W. Fellner. Linear algorithms in sublinear time — a tutorial on statistical estimation. IEEE Computer Graphics and Applications, 31(2):58–66, March/ April 2011. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Upson:1989:AVS

[UFK^{+}89] Craig Upson, Thomas A.

Faulhaber, Jr., David Kamins, David Laidlaw, David Schlegel, Jeffrey Vroom, Robert Gurwitz, and Andries van Dam. The Application Visualization System: a computational environment for scientific visualization. IEEE Computer Graphics and Applications, 9(4):30–42, July/ August 1989. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Urness:2006:SVM

[UIL^{+}06] Timothy Urness, Victoria Interrante, Ellen Longmire, Ivan Marusic, Sean O’Neill, and Thomas W. Jones. Strategies for the visualization of multiple 2D vector fields. IEEE Computer Graphics and Applications, 26(4):74–82, July/ August 2006. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Ulfsby:1982:TDB

[UMO82] Stig Ulfsby, Steinar Meen, and Jorn Oian. Tornado: a data-base management system for graphics applications. IEEE Computer Graphics and Applications, 2(3):71–76, 78–79, May/ June 1982. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Udupa:1991:FVM

[UO91] Jayaram K. Udupa and Dewey Odhner. Fast visualization, manipulation, and analysis of binary volumetric objects. IEEE Computer Graphics and Applications, 11(6):53–62, November/December 1991. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Udupa:1993:SR

[UO93] Jayaram K. Udupa and Dewey Odhner. Shell rendering. IEEE Computer Graphics and Applications, 13(6):58–67, November/December 1993. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Umetani:2017:SSM

[US17] Nobuyuki Umetani and Ryan Schmidt. SurfCuit: Surface-mounted circuits on 3D prints. IEEE Computer Graphics and Applications, 37(3):52–60, May/ June 2017. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        cg/2017/03/mcg2017030052-
                                        abs.html.

Uselton:1993:CRD

[Use93] S. P. Uselton. Comments, with reply, on “Delaunay triangulation using a uniform grid” by T. P. Fang and L. A. Piegl. IEEE Computer Graphics and Applications, 13(6):4, November/ December 1993. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Uselton:1998:EWT

[Use98] Samuel P. Uselton. ExVis and wind tunnel experiment data. IEEE Computer Graphics and Applications, 18(4):75–77, July/ August 1998. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

computer.org/cga/cg1998/
g4075abs.htm.

Usui:2017:CGA

[USW17] Yoko Usui, Katsumi Sato, and Shinichi Watabe. Computer graphics animation for objective self-evaluation. IEEE Computer Graphics and Applications, 37(6):5–9, November/December 2017. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL

https://www.computer.org/
csdl/mags/cg/2017/06/mcg2017060005-
abs.html.

Umetani:2011:RFE

[UTMI11] Nobuyuki Umetani, Kenshi Takayama, Jun Mitani, and Takeo Igarashi. A responsive finite element method to aid interactive geometric modeling. IEEE Computer Graphics and Applications, 31(5):43–53, September/October 2011. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

UUIG:1995:VBA

[UVa95] UVa User Interface Group. VR blackboard: Alice: Rapid prototyping for virtual reality. IEEE Computer Graphics and Applications, 15(3):8–11, May/June 1995. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Usman:2023:MGD

[UZM^{+}23] Muhammad Usman, Honglu Zhou, Seonghyeon Moon, Xun Zhang, Petros Faloutsos, and Mubbasir Kapadia. A multiscale geospatial dataset and an interactive visualization dashboard for computational epidemiology and open scientific research. IEEE Computer Graphics and Applications, 43(1):39– 52, January/February 2023. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

VanAken:1984:EEA

[Van84] Jerry R. Van Aken. An efficient ellipse-drawing algorithm. IEEE Computer Graphics and Applications, 4(9):24–35, September 1984. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

vanWijk:1993:FVS

[van93] Jarke J. van Wijk. Flow visualization with surface particles. IEEE Computer Graphics and Applications, 13(4):18–24, July/August 1993. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

vandePanne:1996:PGS

[van96] Michiel van de Panne. Parameterized gait synthesis. IEEE Computer Graphics and Applications, 16(2): 40–49, March/April 1996. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

VanLoocke:2010:PBF

[Van10] Philip Van Loocke. Polygonbased fractals from compressed iterated function systems. IEEE Computer Graphics and Applications, 30(2):34–44, March/April 2010. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Varsa:2024:RBA

[VB24] Petri M. Varsa and Gladimir V. G. Baranoski. Rendering the bluish appearance of snow: When light transmission matters. IEEE Computer Graphics and Applications, 44(1):50–61, January/ February 2024. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

vanBasten:2012:MTT

[vBE12] Ben van Basten and Arjan Egges. Motion transplantation techniques: a survey. IEEE Computer Graphics and Applications, 32 (3):16–23, May/June 2012. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Vanek:2011:LSP

[VBHS11] Juraj Vanek, Bedrich Benes, Adam Herout, and Ondrej Stava. Large-scale physics-based terrain editing. IEEE Computer Graphics and Applications, 31(6): 35–44, November/December 2011. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Varshney:1994:CSM

[VBW94] Amitabh Varshney, Frederick P. Brooks, Jr., and William V. Wright. Computing smooth molecular surfaces. IEEE Computer Graphics and Applications, 14(5):19–25, September 1994. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Vannier:1997:VPF

[VCBS97] Michael W. Vannier, Paul K.

Commean, Barry S. Brunsden, and Kirk E. Smith. Visualization of prosthesis fit in lower-limb amputees. IEEE Computer Graphics and Applications, 17(5):16–29, September/ October 1997. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://dlib.
computer.org/cg/books/
cg1997/pdf/g5016.pdf.

VanSintJan:1998:VBV

[VCR98] Serge L. Van Sint Jan, Gordon J. Clapworthy, and Marcel Rooze. Visualization blackboard: Visualization of combined motions in human joints. IEEE Computer Graphics and Applications, 18(6):10–14, November/December 1998. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http:

//dlib.computer.org/cg/
books/cg1998/pdf/g6010.
pdf.

Jan:2000:MBD

[VCR00] Serge L. Van Sint Jan, Gordon J. Clapworthy, and Marcel Rooze. Morphologybased data elimination from medical image data. IEEE Computer Graphics and Applications, 20(2):46–52, March/April 2000. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

                                        computer.org/cga/cg2000/
                                        g2046abs.htm;   http:/
                                        /dlib.computer.org/cg/
                                        books/cg2000/pdf/g2046.
                                        pdf.

Vesna:2019:VVI

[VCS19] V. Vesna, B. Campbell, and F. Samsel. Victoria Vesna: Inviting meaningful organic art science collaboration. IEEE Computer Graphics and Applications, 39(4):8–13, July/ August 2019. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

vanDam:2000:BW

[vD00]  Andy van Dam.  Be-

yond WIMP. IEEE Computer Graphics and Applications, 20(1):50–51, January/ February 2000. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL http://dlib.
                                        computer.org/cg/books/
                                        cg2000/pdf/g1050.pdf.

vanDam:2005:VRP

[vD05] Andries van Dam. Visualization research problems in next-generation educational software. IEEE Computer Graphics and Applications, 25(5):88–92, September/October 2005. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

vanDam:2012:CM

[vD12] Andy van Dam. Carl Machover, 1927–2012. IEEE Computer Graphics and Applications, 32(4):20–22, July/August 2012. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

vandenElzen:2023:FTV

[vdEAA^{+}23] Stef van den Elzen, Gennady Andrienko, Natalia Andrienko, Brian D. Fisher, Rafael M. Martins, Jaakko Peltonen, Alexandru C. Telea, and Michel Verleysen. The flow of trust: a visualization framework to externalize, explore, and explain trust in ML applications. IEEE Computer Graphics and Applications, 43(2):78–88, March/ April 2023. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

vanDam:2000:IVS

[vDFL^{+}00] Andries van Dam, Andrew S. Forsberg, David H. Laidlaw, Joseph J. LaViola, Jr., and Rosemary M.

Simpson. Immersive VR for scientific visualization: a progress report. IEEE Computer Graphics and Applications, 20(6):26–52, November/December 2000. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

computer.org/cga/cg2000/
g6026abs.htm;   http:/
/dlib.computer.org/cg/
books/cg2000/pdf/g6026.
pdf.

Vote:2002:DPA

[VFLJ02] Eileen Vote, Daniel Acevedo Feliz, David H. Laidlaw, and Martha Sharp Joukowsky. Discovering Petra: Archaeological analysis in VR. IEEE Computer Graphics and Applications, 22(5):38–50, September/October 2002. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://dlib.
computer.org/cg/books/
cg2002/pdf/g5038.pdf;
http://www.computer.org/
cga/cg2002/g5038abs.htm.

Vinberg:1981:CGB

[VG81] Anders Vinberg and James E. George. Computer graphics and the business executive — the new management team. IEEE Computer Graphics and Applications, 1(1):57–58, 61–68, 70– 71, January/February 1981. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Verbeck:1984:CLD

[VG84] Channing P. Verbeck and Donald P. Greenberg. A comprehensive light-source description for computer graphics. IEEE Computer Graphics and Applications, 4(7):66–75, July/ August 1984. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

vanWijk:1994:TWS

[vHdP94] Jarke J. van Wijk, Andrea J. S. Hin, Willem C. de Leeuw, and Frits H. Post. Three ways to show 3D fluid flow. IEEE Computer Graphics and Applications, 14(5):33–39, September 1994. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

vanHam:2009:GEI

[vHV09] Frank van Ham and Fernanda B. Viégas. Guest Editors’ introduction: Collaborative visualization. IEEE Computer Graphics and Applications, 29(5):20–21, September/October 2009. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Vlahakis:2002:AAR

[VIK^{+}02] Vassilios Vlahakis, Nikolaos Ioannidis, John Karigiannis, Manolis Tsotros, Michael Gounaris, Didier Stricker, Tim Gleue, Patrick Daehne, and Luís Almeida. Archeoguide: An augmented reality guide for archaeological sites. IEEE Computer Graphics and Applications, 22(5):52–60, September/October 2002. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL

http://dlib.computer.org/
cg/books/cg2002/pdf/g5052.
pdf;   http://www.computer.
org/cga/cg2002/g5052abs.
htm.

Vasylevska:2017:CVF

[VK17] Khrystyna Vasylevska and Hannes Kaufmann. Compressing VR: Fitting large virtual environments within limited physical space. IEEE Computer Graphics and Applications, 37(5):85–91, September/October 2017. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL

https://www.computer.org/
csdl/mags/cg/2017/05/mcg2017050085-
abs.html.

vanLent:2006:GEI

[vL06] Michael van Lent. Guest Editor’s introduction: Interactive narrative. IEEE Computer Graphics and Applications, 26(3):20–21, May/ June 2006. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cg/2006/03/g3020.pdf.

vonLukas:2016:UVC

[vL16] Uwe Freiherr von Lukas. Underwater visual computing: The grand challenge just around the corner. IEEE Computer Graphics and Applications, 36(2):10–15, March/April 2016. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2016/02/mcg2016020010-
abs.html.

vonLandesberger:2018:IVC

[vL18] Tatiana von Landesberger.

Insights by visual comparison: The state and challenges. IEEE Computer Graphics and Applications, 38(3):140–148, May/ June 2018. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2018/03/mcg2018030140-
abs.html.

vonLandesberger:2017:TUS

[vLBW17] Tatiana von Landesberger, Sebastian Bremm, and Marcel Wunderlich. Typology of uncertainty in static geolocated graphs for visualization. IEEE Computer Graphics and Applications, 37(5):18–27, September/October 2017. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2017/05/mcg2017050018-
abs.html.

Valbom:2007:IMI

[VM07] Leonel Valbom and Adérito Marcos. An immersive musical instrument prototype. IEEE Computer Graphics and Applications, 27(4): 14–19, July/August 2007. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Volino:1996:ESS

[VMJT96] Pascal Volino, Nadia Magnenat Thalmann, Shen Jianhua, and Daniel Thalmann. An evolving system for simulating clothes on virtual actors. IEEE Computer Graphics and Applications, 16(5):42–51, September 1996. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Voorhies:1989:RG

[Voo89] Douglas Voorhies. Reducedcomplexity graphics. IEEE Computer Graphics and Applications, 9(4):63–70, July/ August 1989. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Vossler:1985:SCU

[Vos85] Donald L. Vossler. Sweepto-CSG conversion using pattern recognition techniques. IEEE Computer Graphics and Applications, 5(8):61–68, August 1985. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

vonOmpteda:2015:TQW

[vOW15] Karin von Ompteda and Kevin Walker. Translating the quantum world to human scale: An art–science collaboration. IEEE Computer Graphics and Applications, 35(3):74–81, May/ June 2015. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
cg/2015/03/mcg2015030074-
abs.html.

Voyatzis:1999:PDI

[VP99] George Voyatzis and Ioannis Pitas. Protecting digital-image copyrights: a framework. IEEE Computer Graphics and Applications, 19(1):18–24, January/ February 1999. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

computer.org/cga/cg1999/
g1018abs.htm;   http:/

                                        /dlib.computer.org/cg/
                                        books/cg1999/pdf/g1018.
                                        pdf.

Vuillemot:2016:STP

[VP16] Romain Vuillemot and Charles Perin. Sports tournament predictions using direct manipulation. IEEE Computer Graphics and Applications, 36(5):62–71, September/October 2016. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL

                                        https://www.computer.org/
                                        csdl/mags/cg/2016/05/mcg2016050062-
                                        abs.html.

Vannier:1991:FSS

[VPB^{+}91] Michael W. Vannier, Tom Pilgram, Gulab Bhatia, Barry Brunsden, and Paul Commean. Facial surface scanner. IEEE Computer Graphics and Applications, 11(6):72–80, November/December 1991. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Vasilakis:2021:RTA

[VPVG21] Andreas A. Vasilakis, G. Papaioannou, N. Vitsas, and A. Gkaravelis. Remote teaching advanced rendering topics using the rayground platform. IEEE Computer Graphics and Applications, 41(5):99–103, September/ October 2021. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Varela:2007:VED

[VS07] José M. Varela and C. Guedes Soares. A virtual environment for decision support in ship damage control. IEEE Computer Graphics and Applications, 27(4): 58–69, July/August 2007. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Varela:2014:RDW

[VS14] Jose Miguel Varela and Carlos Guedes Soares. Ring discretization of the wave spectrum for sea surface simulation. IEEE Computer Graphics and Applications, 34(2):58–71, March/ April 2014. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Victor:2022:VPS

[VSLG22] Viny Saajan Victor, Andre Schmeißer, Heike Leitte, and Simone Gramsch. Visual parameter space analysis for optimizing the quality of industrial nonwovens. IEEE Computer Graphics and Applications, 42(2): 56–67, March/April 2022. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Vartiainen:2024:HTI

[VT24] Henriikka Vartiainen and Matti Tedre. How text-to-image generative AI is transforming mediated action. IEEE Computer Graphics and Applications, 44(2): 12–22, March/April 2024. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Vilbrandt:2011:DIT

[VVP^{+}11] Turlif Vilbrandt, Carl Vilbrandt, Galina Ivanovna Pasko, Cherie Stamm, and Alexander Pasko. Digitally interpreting traditional folk crafts. IEEE Computer Graphics and Applications, 31(4):12–18, July/ August 2011. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

vanWijk:2006:BG

[vW06] Jarke J. van Wijk. Bridging the gaps. IEEE Computer Graphics and Applications, 26(6):6–9, November/ December 2006. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

vanWijk:2010:GEI

[vWNS10] Jarke van Wijk, Stephen North, and Han-Wei Shen. Guest Editors’ introduction: Highlights from IEEE Pacific Visualization. IEEE Computer Graphics and Applications, 30(6):14–15, November/December 2010. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Wallace:2005:TAL

[WAB^{+}05] Grant Wallace, Otto J. Anshus, Peng Bi, Han Chen, Yuqun Chen, Douglas Clark, Perry Cook, Adam Finkelstein, Thomas Funkhouser, Anoop Gupta, Matthew Hibbs, Kai Li, Zhiyan Liu, Rudrajit Samanta, Rahul Sukthankar, and Olga Troyanskaya. Tools and applications for large-scale display walls. IEEE Computer Graphics and Applications, 25(4):24–33, July/ August 2005. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Wrzesien:2013:TLT

[WAB^{+}13] Maja Wrzesien, Mariano Alcañiz, Cristina Botella, Jean-Marie Burkhardt, Juana Bretón-López, Mario Ortega, and Daniel Beneito Brotons. The therapeutic lamp: Treating smallanimal phobias. IEEE Computer Graphics and Applications, 33(1):80–86, January/ February 2013. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Walton:2021:DDV

[WAH^{+}21] J. Walton, S. Adams, W. Hayek, P. Florek, and H. Dyson. Dynamic 3-D visualization of climate model development and results. IEEE Computer Graphics and Applications, 41(1):17– 25, January/February 2021. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Wallin:1991:CIC

[Wal91] Ake Wallin. Constructing isosurfaces from CT data. IEEE Computer Graphics and Applications, 11(6): 28–33, November/December 1991. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Wang:1993:PIA

[Wan93] Fangju Wang. A parallel intersection algorithm for vector polygon overlay. IEEE Computer Graphics and Applications, 13(2): 74–81, March/April 1993. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Wang:2010:DPF

[Wan10] Charlie Wang. From designing products to fabricating them from planar materials. IEEE Computer Graphics and Applications, 30(6): 74–85, November/December 2010. CODEN ICGADZ.

ISSN 0272-1716 (print), 1558-1756 (electronic).

Wang:2022:VTT

[Wan22] Chaoli Wang. VisVisual:

a toolkit for teaching and learning data visualization. IEEE Computer Graphics and Applications, 42(4): 20–26, July/August 2022. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Ware:2006:VUB

[WAPW06] Colin Ware, Roland Arsenault, Matthew Plumlee, and David Wiley. Visualizing the underwater behavior of humpback whales. IEEE Computer Graphics and Applications, 26(4): 14–18, July/August 2006. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Warner:1981:PDI

[War81a]  J. R. Warner.  Princi-

ples of device independent computer graphics software. IEEE Computer Graphics and Applications, 1(10):85– 100, October 1981. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Warner:1981:PDC

[War81b] James R. Warner. Principles of device-independent computer graphics software. IEEE Computer Graphics and Applications, 1(4):85– ??, October 1981. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Warner:1985:SGS

[War85] James R. Warner. Standard graphics software for highperformance applications. IEEE Computer Graphics and Applications, 5(3):74– 79, March/April 1985. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Ware:1988:CSU

[War88]  Colin Ware.  Color se-

quences for univariate maps: Theory, experiments, and principles. IEEE Computer Graphics and Applications, 8(5):41–49, September/October 1988. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Warnock:2000:CLG

[War00] John E. Warnock. The changing landscape of graphic arts. IEEE Computer Graphics and Applications, 20(1):32–33, January/February 2000. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL

http://dlib.computer.org/
cg/books/cg2000/pdf/g1032.
pdf.

Ward:2004:FNL

[War04] Matthew Ward. Finding needles in large-scale multivariate data haystacks. IEEE Computer Graphics and Applications, 24(5): 16–19, September/October 2004. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL

http://csdl.computer.org/
dl/mags/cg/2004/05/g5016.
htm;   http://csdl.computer.
org/dl/mags/cg/2004/05/
g5016.pdf.

Ware:2008:TPT

[War08] Colin Ware. Toward a perceptual theory of flow visualization. IEEE Computer Graphics and Applications, 28(2):6–11, March/ April 2008. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Watson:2006:BOC

[Wat06] Benjamin Watson. Broadening our collaboration with design. IEEE Computer Graphics and Applications, 26(5):18–21, September/ October 2006. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Ward:1985:IRH

[WB85] Jean Renard Ward and Barry Blesser. Interactive recognition of handprinted characters for computer input. IEEE Computer Graphics and Applications, 5(9):24–37, September 1985. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Want:2000:GEI

[WB00a] Roy Want and Gaetano Borriello. Guest Editors’ introduction: Information appliances. IEEE Computer Graphics and Applications, 20(3):21–23, May/ June 2000. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://dlib.
computer.org/cg/books/
cg2000/pdf/g3021.pdf.

Want:2000:SIA

[WB00b] Roy Want and Gaetano Borriello. Survey on information appliances. IEEE Computer Graphics and Applications, 20(3):24–31, May/ June 2000. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

computer.org/cga/cg2000/
g3024abs.htm;   http:/
/dlib.computer.org/cg/
books/cg2000/pdf/g3024.
pdf.

Wilson:2014:IAB

[WB14] Andy Wilson and Hrvoje Benko. Interacting above and beyond the display [guest editorial]. IEEE Computer Graphics and Applications, 34(3):20–21, May/ June 2014. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Wood:1996:WSC

[WBA^{+}96] Frank Wood, Daniel Brown, Robert A. Amidon, Jonathan Alferness, Brian Joseph, Richard E. Gillilan, and Carlos Faerman. WorkSpace and the study of Chagas’ disease. IEEE Computer Graphics and Applications, 16(4):72–78, July/ August 1996. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). WorkSpace is a 3-D extension of the 2-D graphics model provided by the Java Abstract Window Toolkit.

Weiler:2005:HAF

[WBS^{+}05] Manfred Weiler, Ralf Botchen, Simon Stegmaier, Thomas Ertl, Jingshu Huang, Yun Jang, David S. Ebert, and Kelly P. Gaither. Hardwareassisted feature analysis and visualization of procedurally encoded multifield volumetric data. IEEE Computer Graphics and Applications, 25(5):72–81, September/October 2005. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Wilson:1991:EIS

[WC91] Michael Wilson and Anthony Conway. Enhanced interaction styles for user interfaces. IEEE Computer Graphics and Applications, 11(2):79–90, March/ April 1991. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Wei:2011:IIH

[WC11] Xiaolin K. Wei and Jinxiang Chai. Intuitive interactive human-character posing with millions of example poses. IEEE Computer Graphics and Applications, 31(4):78–88, July/ August 2011. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Weber:2017:ARA

[WCE^{+}17] Gunther H. Weber, Sheelagh Carpendale, David Ebert, Brian Fisher, Hans Hagen, Ben Shneiderman, and Anders Ynnerman. Apply or die: On the role and assessment of application papers in visualization. IEEE Computer Graphics and Applications, 37(3):96–104, May/ June 2017. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        cg/2017/03/mcg2017030096.
                                        html.

Wong:2011:GAL

[WCG^{+}11] Pak Chung Wong, Chaomei Chen, Carsten Gorg, Ben Shneiderman, John Stasko, and Jim Thomas. Graph analytics — lessons learned and challenges ahead. IEEE Computer Graphics and Applications, 31(5):18–29, September/October 2011. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Weber:2018:SWM

[WCS18] Sally Weber, Bruce Campbell, and Francesca Samsel. Sally Weber: Making art from light. IEEE Computer Graphics and Applications, 38(3):8–12, May/ June 2018. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2018/03/mcg2018030008.
html.

Wei:2015:PML

[WD15a] Li Wei and Zhigang Deng.

A practical model for live speech-driven lip-sync. IEEE Computer Graphics and Applications, 35(2):70– 78, March/April 2015. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http:/

/www.computer.org/csdl/
mags/cg/2015/02/mcg2015020070-
abs.html.

Wiens:2015:CESa

[WD15b] Alex Wiens and Gitta Domik. In-class exercise for shadow mapping algorithms. IEEE Computer Graphics and Applications, 35(3):15–19, May/ June 2015. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
cg/2015/03/mcg2015030015-
abs.html.

Wiens:2015:CESb

[WD15c] Alex Wiens and Gitta Domik. In-class exercises for shadow volume algorithms. IEEE Computer Graphics and Applications, 35(4):108–112, July/ August 2015. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
cg/2015/04/mcg2015040108-
abs.html.

Wu:2023:GCV

[WDC^{+}23] Aoyu Wu, Dazhen Deng, Min Chen, Shixia Liu, Daniel Keim, Ross Maciejewski, Silvia Miksch, Hendrik Strobelt, Fernanda Viégas, and Martin Wattenberg. Grand challenges in visual analytics applications. IEEE Computer Graphics and Applications, 43(5): 83–90, September/October 2023. CODEN ICGADZ.

ISSN 0272-1716 (print), 1558-1756 (electronic).

Wnuk:2017:CTV

[WDK17] Agnieszka Wnuk, Konrad J. Debski, and Marcin Kozak. CI thermometer: Visualizing confidence intervals in correlation analysis. IEEE Computer Graphics and Applications, 37(6):103–108, November/ December 2017. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2017/06/mcg2017060103-
abs.html.

Weatherford:1998:ACB

[Wea98a] Margaret Weatherford. About the cover: Beyond the virtual salon — software games for girls. IEEE Computer Graphics and Applications, 18(1):4–6, January/ February 1998. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

computer.org/cga/cg1998/
g1004abs.htm;   http:/
/dlib.computer.org/cg/
books/cg1998/pdf/g1004.
pdf.

Weatherford:1998:ACK

[Wea98b] Margaret Weatherford. About the cover: Keyframing to fame. IEEE Computer Graphics and Applications, 18(4):4–5, July/ August 1998. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Weatherford:1999:ACAb

[Wea99a] Margaret Weatherford. About the cover: Animal, vegetable, digital. IEEE Computer Graphics and Applications, 19(5):4–5, September/October 1999. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL http://dlib.
                                        computer.org/cg/books/
                                        cg1999/pdf/g5004.pdf.

Weatherford:1999:ACAa

[Wea99b] Margaret Weatherford. About the cover: Apparatus for digesting 3D thoughts. IEEE Computer Graphics and Applications, 19(3):4–5, May/ June 1999. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL http://dlib.
                                        computer.org/cg/books/
                                        cg1999/pdf/g3004.pdf.

Weber:2008:FSR

[Web08] Jason P. Weber. Fast simulation of realistic trees. IEEE Computer Graphics and Applications, 28(3):67– 75, May/June 2008. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Weeks:2002:RTR

[Wee02] Jeff Weeks. Real-time rendering in curved spaces. IEEE Computer Graphics and Applications, 22(6): 90–99, November/December 2002. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL

http://dlib.computer.org/
cg/books/cg2002/pdf/g6090.
pdf;   http://www.computer.
org/cga/cg2002/g6090abs.
htm.

Weisberg:1984:INS

[Wei84] David E. Weisberg. Impact of network system architecture on CAD/CAM productivity. IEEE Computer Graphics and Applications, 4(8):36–40, August 1984. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Weiler:1985:EDS

[Wei85] Kevin Weiler. Edge-based data structures for solid modeling in curved-surface environments. IEEE Computer Graphics and Applications, 5(1):21–40, January/ February 1985. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Wein:2023:HHC

[Wei23] Marceli Wein. From holocaust hidden child to computer animation laboratory. IEEE Computer Graphics and Applications, 43(1): 103–109, January/February 2023. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Waller:2000:ECM

[WEWL00] Marcus D. Waller, Jon P. Ewins, Martin White, and Paul F. Lister. Efficient coverage mask generation for antialiasing. IEEE Computer Graphics and Applications, 20(6):86–93, November/December 2000. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

computer.org/cga/cg2000/
g6086abs.htm;   http:/
/dlib.computer.org/cg/
books/cg2000/pdf/g6086.
pdf.

Wood:1993:AIA

[WF93] Joseph D. Wood and Peter F. Fisher. Assessing interpolation accuracy in elevation models. IEEE Computer Graphics and Applications, 13(2):48–56, March/ April 1993. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Welch:2002:MTN

[WF02] Greg Welch and Eric Foxlin. Motion tracking: No silver bullet, but a respectable arsenal. IEEE Computer Graphics and Applications, 22(6):24–38, November/ December 2002. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://dlib.
computer.org/cg/books/
cg2002/pdf/g6024.pdf;
http://www.computer.org/
cga/cg2002/g6024abs.htm.

Wanger:1992:PSR

[WFG92] Leonard R. Wanger, James A.

Ferwerda, and Donald P. Greenberg. Perceiving spatial relationships in computer-generated images. IEEE Computer Graphics and Applications, 12(3):44– 58, May/June 1992. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Wong:2000:DSV

[WFL^{+}00] Pak Chung Wong, Harlan Foote, Ruby Leung, Dan Adams, and Jim Thomas. Data signatures and visualization of scientific data sets. IEEE Computer Graphics and Applications, 20(2):12– 15, March/April 2000. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http:

//dlib.computer.org/cg/
books/cg2000/pdf/g2012.
pdf.

Wilson:1985:IDT

[WFOP85] P. R. Wilson, I. D. Faux, M. C. Ostrowski, and K. G. Pasquill. Interfaces for data transfer between solid modeling systems. IEEE Computer Graphics and Applications, 5(1):41–51, January/ February 1985. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Welch:2000:PIY

[WFR^{+}00] Greg Welch, Henry Fuchs, Ramesh Raskar, Herman Towles, and Michael S. Brown. Projected imagery in your “Office of the future”. IEEE Computer Graphics and Applications, 20(4):62–67, July/ August 2000. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

                                        computer.org/cga/cg2000/
                                        g4062abs.htm;   http:/
                                        /dlib.computer.org/cg/
                                        books/cg2000/pdf/g4062.
                                        pdf.

Wu:2021:PVD

[WFZB21] Yingcai Wu, Siwei Fu, Jian Zhao, and Chris Bryan. Powering visualization with deep learning. IEEE Computer Graphics and Applications, 41(5):16–17, September/October 2021. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Wee:1995:ERPa

[WG95a] Chionh Eng Wee and Ronald N. Goldman. Elimination and resultants: Part 1 — elimination and bivariate resultants. IEEE Computer Graphics and Applications, 15(1):69–77, January/ February 1995. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Wee:1995:ERPb

[WG95b] Chionh Eng Wee and Ronald N. Goldman. Elimination and resultants: Part 2 — multivariate resultants. IEEE Computer Graphics and Applications, 15(2): 60–69, March/April 1995. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Willett:2015:UDN

[WGI15] Wesley Willett, Pascal Goffin, and Petra Isenberg. Understanding digital notetaking practice for visualization. IEEE Computer Graphics and Applications, 35(4):38–51, July/ August 2015. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
cg/2015/04/mcg2015040038-
abs.html.

Wegenkittl:1997:VDB

[WGP97] Rainer Wegenkittl, Eduard Gröller, and Werner Purgathofer. Visualizing the dynamical behavior of Wonderland. IEEE Computer Graphics and Applications, 17(6):71–79, November/December 1997. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

                                        computer.org/cga/cg1997/
                                        g6071abs.htm.

Wang:2018:OQA

[WGZ^{+}18] Shiqi Wang, Ke Gu, Kai Zeng, Zhou Wang, and Weisi Lin. Objective quality assessment and perceptual compression of screen content images. IEEE Computer Graphics and Applications, 38(1):47–58, January/ February 2018. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        cg/2018/01/mcg2018010047-
                                        abs.html.

Wiley:1997:ISA

[WH97] Douglas J. Wiley and James K. Hahn. Interpolation synthesis of articulated figure motion. IEEE Computer Graphics and Applications, 17(6):39–45, November/December 1997. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

                                        computer.org/cga/cg1997/
                                        g6039abs.htm.

Wong:1999:DTF

[WH99] Stephen T. C. Wong and Kent Soo Hoo, Jr. Digital teaching files in diagnostic imaging. IEEE Computer Graphics and Applications, 19(3):56–65, May/ June 1999. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

computer.org/cga/cg1999/
g3056abs.htm;   http:/
/dlib.computer.org/cg/
books/cg1999/pdf/g3056.
pdf.

Whitman:1994:RDP

[WHC94] Scott Whitman, Charles D.

Hansen, and Thomas W. Crockett. Recent developments in parallel rendering. IEEE Computer Graphics and Applications, 14(4): 21–22, July/August 1994. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Wong:2014:VAP

[WHC^{+}14] Pak Chung Wong, Zhenyu Huang, Yousu Chen, Patrick Mackey, and Shuangshuang Jin. Visual analytics for power grid contingency analysis. IEEE Computer Graphics and Applications, 34(1):42–51, January/ February 2014. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Wong:2001:IRP

[WHF01] Tien-Tsin Wong, Pheng-Ann Heng, and Chi-Wing Fu. Interactive relighting of panoramas. IEEE Computer Graphics and Applications, 21(2):32–41, March/ April 2001. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

                                        computer.org/cga/cg2001/
                                        extras/g2032x1.mpg;   http:
                                        //computer.org/cga/cg2001/
                                        extras/g2032x2.mpg;   http:
                                        //computer.org/cga/cg2001/
                                        extras/g2032x3.mpg;   http:
                                        //computer.org/cga/cg2001/
                                        g2032abs.htm;   http:/
                                        /dlib.computer.org/cg/
                                        books/cg2001/pdf/g2032.
                                        pdf.

Whitton:1984:MDR

[Whi84] Mary C. Whitton. Memory design for raster graphics displays. IEEE Computer Graphics and Applications, 4(3):48–65, March/ April 1984. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Whitney:1987:GSS

[Whi87] John Whitney, Jr. The graphics soapbox — the second generation. IEEE Computer Graphics and Applications, 7(6):62–63, June 1987. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Whitehouse:1994:MMM

[Whi94a] Karen Whitehouse. Mike Miller’s many hats. IEEE Computer Graphics and Applications, 14(1):4–6, January/February 1994. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Whitman:1994:DLB

[Whi94b] Scott Whitman. Dynamic load balancing for parallel polygon rendering. IEEE Computer Graphics and Applications, 14(4):41–48, July/August 1994. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Whitehouse:1995:KEE

[Whi95a] K. Whitehouse. Keeping an electronic eye on the road. IEEE Computer Graphics and Applications, 15(1):16– 17, January/February 1995. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Whitehouse:1995:ACA

[Whi95b] Karen Whitehouse. About the cover: Art without pretense. IEEE Computer Graphics and Applications, 15(6):4–5, November/ December 1995. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Whitehouse:1995:ACP

[Whi95c] Karen Whitehouse. About the cover: Pictures worth more than a thousand words. IEEE Computer Graphics and Applications, 15(5):4–6, July/August 1995. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Whitehouse:1995:NLI

[Whi95d] Karen Whitehouse. In the news: Licensing issues could affect MPEG’s future; Blinn moves to Microsoft; US Supreme Court to decide software copyright issues; compromise carries the day in videodisc war; the Visible Human steps out; intellectual property news. IEEE Computer Graphics and Applications, 15(6):94–95, November/December 1995. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Whitehouse:1996:AVA

[Whi96a] K. Whitehouse. Applications: VRML adds a new dimension to Web browsing. IEEE Computer Graphics and Applications, 16(4):7–9, July/August 1996. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Whitehouse:1996:ACB

[Whi96b] Karen Whitehouse. About the cover: Building cyberspace. IEEE Computer Graphics and Applications, 16(4):4–5, July/ August 1996. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Whitehouse:1996:ACM

[Whi96c] Karen Whitehouse. About the cover: Making connections. IEEE Computer Graphics and Applications, 16(3):4–5, May/June 1996. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Whitehouse:1996:AWW

[Whi96d] Karen Whitehouse. Applications: Weather without the weatherman [on the World Wide Web]. IEEE Computer Graphics and Applications, 16(2):12–15, March/ April 1996. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). Discusses use of Java for browsing of weather data on the Web, with samples available at http:/

/cirrus.sprl.umich.edu/
javaweather/.

Whitehouse:1997:CGE

[Whi97] Karen Whitehouse. Computer graphics education issue on the Web. IEEE Computer Graphics and Applications, 17(1):20–21, January/ February 1997. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Whitted:1999:VVD

[Whi99] Turner Whitted. Visualization viewpoints: Draw on the wall. IEEE Computer Graphics and Applications, 19(4):6–9, July/ August 1999. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL http://dlib.
                                        computer.org/cg/books/
                                        cg1999/pdf/g4006.pdf.

Whitaker:2004:MDS

[Whi04]  Ross Whitaker.  Model-

ing deformable surfaces with level sets. IEEE Computer Graphics and Applications, 24(5):6–9, September/October 2004. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http:

                                        //csdl.computer.org/dl/
                                        mags/cg/2004/05/g5006.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/cg/2004/05/
                                        g5006.pdf.

Whitted:2020:OGI

[Whi20] T. Whitted. Origins of global illumination. IEEE Computer Graphics and Applications, 40(1):20–27, January/February 2020. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Whitton:2023:RFP

[Whi23] Mary Whitton. Remembering Frederick P. Brooks, Jr.

(1931–2022). IEEE Computer Graphics and Applications, 43(1):109, January/ February 2023. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Wakefield:2013:SIM

[WHKM^{+}13] Graham Wakefield, Tobias Hollerer, JoAnn Kuchera-Morin, Charles Roberts, and Matthew Wright. Spatial interaction in a multiuser immersive instrument. IEEE Computer Graphics and Applications, 33(6):14–20, November/December 2013. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Wan:2019:ICV

[WHR19] Yong Wan, Charles Hansen, and Theresa-Marie Rhyne. Independent and collaborative visualization tool development. IEEE Computer Graphics and Applications, 39(1):44–52, January/February 2019. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL https:

//ieeexplore.ieee.org/
document/8663651/.

Wang:2017:NPT

[WHSM17] Feng Wang, Brett Hansen, Ryan Simmons, and Ross Maciejewski. Name profiler toolkit. IEEE Computer Graphics and Applications, 37(5):61–71, September/October 2017. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        cg/2017/05/mcg2017050061-
                                        abs.html.

Willmert:1982:VHB

[Wil82] K. D. Willmert. Visualizing human body motion simulations. IEEE Computer Graphics and Applications, 2(9):35–38, November/December 1982. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Williams:1984:GIE

[Wil84] Thomas Lee Williams. A graphical interface to an economist’s workstation. IEEE Computer Graphics and Applications, 4(8):42– 47, August 1984. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Wilson:1985:EFG

[Wil85] P. R. Wilson. Euler formulas and geometric modeling. IEEE Computer Graphics and Applications, 5(8):24– 36, August 1985. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Williams:1986:AUI

[Wil86] A. Williams. An architecture for user interface research and development. IEEE Computer Graphics and Applications, 6(7):39– 50, July/August 1986. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Wilhelms:1987:TAM

[Wil87a] Jane Wilhelms. Toward automatic motion control. IEEE Computer Graphics and Applications, 7(4):11– 22, April 1987. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Wilhelms:1987:UDA

[Wil87b] Jane Wilhelms. Using dynamic analysis for realistic animation of articulated bodies. IEEE Computer Graphics and Applications, 7(6):12–27, June 1987. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Wilson:1987:CRS

[Wil87c] Peter R. Wilson. Conic representations for shape description. IEEE Computer Graphics and Applications, 7(4):23–30, April 1987. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Wilson:1991:SPT

[Wil91] Peter R. Wilson. Standards: Past tense and future perfect? IEEE Computer Graphics and Applications, 11(1):44–48, January/ February 1991. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Wilhelms:1997:AA

[Wil97] Jane Wilhelms. Animals with anatomy. IEEE Computer Graphics and Applications, 17(3):22–30, May/ June 1997. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Luk:2013:ASR

[Wil13] Lukáš Hošek Alexander Wilkie. Adding a solarradiance function to the Hošek–Wilkie skylight model. IEEE Computer Graphics and Applications, 33(3):44– 52, May/June 2013. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Wixson:1983:FPT

[Wix83] Steven E. Wixson. Fourdimensional processing tools for cardiovascular data. IEEE Computer Graphics and Applications, 3(5):53– 59, August 1983. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Ware:1988:UBS

[WJ88] Colin Ware and Danny R. Jessome. Using the BAT: a six-dimensional mouse for object placement. IEEE Computer Graphics and Applications, 8(6):65–70, November/December 1988. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Wallace:1993:APC

[WJ93] David R. Wallace and Mark J. Jakiela. Automated product concept design: Unifying aesthetics and engineering. IEEE Computer Graphics and Applications, 13(4):66–75, July/ August 1993. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Wasee:2022:PSR

[WJP22] Farhan Rahman Wasee, Alen Joy, and Charalambos Poullis. Predicting surface reflectance properties of outdoor scenes under unknown natural illumination. IEEE Computer Graphics and Applications, 42(5):19–27, September/October 2022. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Wells:1985:ALM

[WK85] Charles H. Wells and Vladimir Keene. Application of a large-image memory graphics workstation. IEEE Computer Graphics and Applications, 5(2):58– 70, February 1985. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Wolff:1990:RTP

[WK90] Lawrence B. Wolff and David J. Kurlander. Ray tracing with polarization parameters. IEEE Computer Graphics and Applications, 10(6):44–55, November/December 1990. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Wang:1994:VM

[WK94] Sidney W. Wang and Arie E. Kaufman. Volume-sampled 3D modeling. IEEE Computer Graphics and Applications, 14(5):26–32, September 1994. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Ward:2011:DDS

[WKB11] Ben Ward, Sing Bing Kang, and Eric P. Bennett. Depth director: a system for adding depth to movies. IEEE Computer Graphics and Applications, 31(1):36– 48, January/February 2011. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Wong:1996:MIF

[WKHL96] Stephen T. C. Wong, Robert C. Knowlton, Randy A. Hawkins, and Kenneth D. Laxer. Multimodal image fusion for noninvasive epilepsy surgery planning. IEEE Computer Graphics and Applications, 16(1):30– 38, January/February 1996. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Watson:2004:STS

[WKM^{+}04] Benjamin Watson, Janet Kim, Tim McEneany, Tom Moher, Claudia Hindo, Louis Gomez, and Stephen Fransen. StorySpace: Technology supporting reflection, expression, and discourse in classroom narrative. IEEE Computer Graphics and Applications, 24(2):13–15, March/April 2004. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http:

//csdl.computer.org/dl/
mags/cg/2004/02/g2013.
htm;   http://csdl.computer.
org/dl/mags/cg/2004/02/
g2013.pdf.

Wyvill:1986:SDR

[WKS86] Geoff Wyvill, Tosiyasu L.

Kunli, and Yasuto Shirai. Space division for ray tracing in CSG (constructive solid geometry). IEEE Computer Graphics and Applications, 6(4):28–34, April 1986. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Wu:2023:ENA

[WKY23] Hsiang-Yun Wu, Karsten Klein, and Da Yan. Effective network analytics: Network visualization and graph data management. IEEE Computer Graphics and Applications, 43(3):10–11, May/ June 2023. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Wolfe:1988:IVS

[WL88] R. H. Wolfe, Jr. and C. N. Liu. Interactive visualization of 3D seismic data: a volumetric method. IEEE Computer Graphics and Applications, 8(4):24–30, July/ August 1988. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Wan:2012:PWM

[WLC^{+}12] Yong Wan, A. Kelsey Lewis, Mary Colasanto, Mark van Langeveld, Gabrielle Kardon, and Charles Hansen. A practical workflow for making anatomical atlases for biological research. IEEE Computer Graphics and Applications, 32(5):70–80, September/October 2012. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Wetzstein:2012:CLF

[WLH^{+}12] Gordon Wetzstein, Douglas Lanman, Matthew Hirsch, Wolfgang Heidrich, and Ramesh Raskar. Compressive light field displays. IEEE Computer Graphics and Applications, 32(5): 6–11, September/October 2012. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Wong:2009:DCV

[WLL^{+}09] Pak Chung Wong, L. Ruby Leung, Ning Lu, Michael J. Scott, Patrick Mackey, Harlan Foote, James Correia, Jr., Z. Todd Taylor, Jianhua Xu, Stephen D. Unwin, and Antonio Sanfilippo. Designing a collaborative visual analytics tool for social and technological change prediction. IEEE Computer Graphics and Applications, 29(5):58–68, September/October 2009. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Wang:2016:MVS

[WLL^{+}16] Xi Wang, David Lindlbauer, Christian Lessig, Marianne Maertens, and Marc Alexa. Measuring the visual salience of 3D printed objects. IEEE Computer Graphics and Applications, 36(4):46–55, July/ August 2016. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        cg/2016/04/mcg2016040046-
                                        abs.html.

Wheless:1996:VCB

[WLVL^{+}96] Glen H. Wheless, Cathy M. Lascara, Arnoldo Valle-Levinson, Donald P. Brutzman, William Sherman, William L. Hibbard, and Brian E. Paul. Virtual Chesapeake Bay: Interacting with a coupled physical/biological model. IEEE Computer Graphics and Applications, 16(4):52–57, July/August 1996. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Welty:1985:TEC

[WMB85] Jeffrey J. Welty, John W. Moser, Jr., and Michael J. Bailey. TIMCOGS: An educational, computer graphics timber-marking simulator. IEEE Computer Graphics and Applications, 5(1):61– 67, January/February 1985. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Wong:2013:VMC

[WMFM13] Pak Chung Wong, Patrick Mackey, Harlan Foote, and Richard May. Visual matrix clustering of social networks. IEEE Computer Graphics and Applications, 33(4):

88–96, July/August 2013. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Wyvill:1997:MF

[WMG97] Geoff Wyvill, Dean McRobie, and Michael Gigante. Modeling with features. IEEE Computer Graphics and Applications, 17(5): 40–46, September/October 1997. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL

http://dlib.computer.org/
cg/books/cg1997/pdf/g5040.
pdf.

Wittenbrink:2000:LEI

[WMG00] Craig Wittenbrink, Tom Malzbender, and Mike Goss. Letter to the editor: Interpolation for volume rendering. IEEE Computer Graphics and Applications, 20(5):6, September/October 2000. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http:

//dlib.computer.org/cg/
books/cg2000/pdf/g5006.

pdf. See [Lev88, Lev00].

Wang:1992:PAG

[WMS92] Yulun Wang, Amante Mangaser, and Partha Srinivasan. A processor architecture for 3D graphics. IEEE Computer Graphics and Applications, 12 (5):96–??, September 1992.

CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Watson:2008:PUM

[WMV^{+}08] Benjamin Watson, Pascal Müller, Oleg Veryovka, Andy Fuller, Peter Wonka, and Chris Sexton. Procedural urban modeling in practice. IEEE Computer Graphics and Applications, 28(3):18–26, May/ June 2008. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Witten:1982:UPC

[WN82] Ian H. Witten and Radford M. Neal. Using Peano curves for bilevel display of continuous-tone images. IEEE Computer Graphics and Applications, 2(3):47– 48, 50–52, May/June 1982. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Wolfe:1992:VRS

[WNAJ92] Robert H. Wolfe, Mark Needels, Tomas Arias, and John D. Joannopoulos. Visual revelations from silicon ab intio calculations. IEEE Computer Graphics and Applications, 12(4): 45–??, July/August 1992. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Wang:2013:BMA

[WO13] Xinchang Wang and Peichang Ouyang. Beautiful math — aesthetic patterns based on logarithmic spirals. IEEE Computer Graphics and Applications, 33(6): 21–23, November/December 2013. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Woerdenweber:1983:STP

[Woe83]  Burkard  Woerdenweber.

Surface triangulation for picture production. IEEE Computer Graphics and Applications, 3(8):45–51, November/December 1983. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Wahle:1996:HRB

[WOF96] Andreas Wahle, Helmut Oswald, and Eckart Fleck. 3D heart-vessel reconstruction from biplane angiograms. IEEE Computer Graphics and Applications, 16(1):65– 73, January/February 1996. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Wong:1999:GEI

[Won99] Pak Chung Wong. Guest Editor’s introduction: Visual data mining. IEEE Computer Graphics and Applications, 19(5):20–21, September/October 1999.

CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL

http://dlib.computer.org/
cg/books/cg1999/pdf/g5020.
pdf.

Wong:2008:VVA

[Won08] Pak Chung Wong. Visualization viewpoints: Adding a visualization feature to Web search engines: It’s time. IEEE Computer Graphics and Applications, 28(6):11–13, November/ December 2008. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Woo:1985:CAB

[Woo85] Tony C. Woo. A combinatorial analysis of boundary data structure schemata. IEEE Computer Graphics and Applications, 5(3):19– 27, March/April 1985. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Woodwark:1988:ERP

[Woo88] J. R. Woodwark. Eliminating redundant primitives from set-theoretic solid models by a consideration of constituents. IEEE Computer Graphics and Applications, 8(3):38–47, May/ June 1988. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Woodwark:1991:RHC

[Woo91a] John Woodwark. Reconstructing history with computer graphics. IEEE Computer Graphics and Applications, 11(1):18–20, January/ February 1991. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Woolsey:1991:MS

[Woo91b] Kristina Hooper Woolsey.

Multimedia scouting. IEEE Computer Graphics and Applications, 11(4):26–38, July/August 1991. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Woo:1993:ESC

[Woo93]  Andrew Woo.  Efficient

shadow computations in ray tracing. IEEE Computer Graphics and Applications, 13(5):78–83, September 1993. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Wood:2015:VPP

[Woo15]  Jo Wood.  Visualizing

personal progress in participatory sports cycling events. IEEE Computer Graphics and Applications, 35(4):73–81, July/ August 2015. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL http://www.
                                        computer.org/csdl/mags/
                                        cg/2015/04/mcg2015040073-
                                        abs.html.

Wordenweber:1983:STP

[Wor83]  B. Wordenweber.  Sur-

face triangulation for picture production. IEEE Computer Graphics and Applications, 3(11):45–51, November/December 1983. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

World:1993:OUM

[Wor93] Linda World. Opera uses master computer score to light the way. IEEE Computer Graphics and Applications, 13(5):84–85, September 1993. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

World:1994:AGF

[Wor94] L. World. Applications: GIS feels the Earth move in Los Angeles. IEEE Computer Graphics and Applications, 14(3):16, May/June 1994. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

World:1995:ACA

[Wor95] Linda World. About the cover: Advancing traditions. IEEE Computer Graphics and Applications, 15(3):5–7, May/June 1995. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

World:1996:ASE

[Wor96a] Linda World. Aesthetic selection: The evolutionary art of Steven Rooke. IEEE Computer Graphics and Applications, 16(1):4– 5, January/February 1996. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

World:1996:NIT

[Wor96b] Linda World. In the news:

Intel and TI will enter 3D graphics market; NRC releases report on cryptography. IEEE Computer Graphics and Applications, 16(4):92, July/August 1996. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Wozny:2021:TK

[Woz21] M. J. Wozny. Tosiyasu L.

Kunii, 1938–2020. IEEE Computer Graphics and Applications, 41(2):114–119, March/April 2021. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Ward:1987:DTP

[WP87] Jean Ward and Mark J.

Phillips. Digitizer technology: Performance characteristics and the effects on the user interface. IEEE Computer Graphics and Applications, 7(4):31–44, April 1987. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Woo:1990:SSA

[WPF90] Andrew Woo, Pierre Poulin, and Alain Fournier. A survey of shadow algorithms. IEEE Computer Graphics and Applications, 10(6): 13–32, November/December 1990. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Westberg:2022:UCV

[WPL22] Jimmy Johansson Westberg, Karljohan Lundin Palmerius, and Jonas Lundberg. UTM city visualization of unmanned aerial vehicles. IEEE Computer Graphics and Applications, 42(5):84–89, September/ October 2022. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Wylie:2001:SRP

[WPLM01] Brian Wylie, Constantine Pavlakos, Vasily Lewis, and Ken Moreland. Scalable rendering on PC clusters. IEEE Computer Graphics and Applications, 21(4):62– 70, July/August 2001. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

computer.org/cga/cg2001/

g4062abs.htm;   http:/
/dlib.computer.org/cg/
books/cg2001/pdf/g4062.
pdf.

Woo:1996:IRR

[WPO96] Andrew Woo, Andrew Pearce, and Marc Ouellette. It’s really not a rendering bug, you see. IEEE Computer Graphics and Applications, 16(5):21–25, September 1996. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Wu:1989:DGE

[WR89] Xiaolin Wu and Jon G.

Rokne. Double-step generation of ellipses. IEEE Computer Graphics and Applications, 9(3):56–69, May/ June 1989. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Wright:1983:NLF

[Wri83] Joan A. Wright. Notes from left field: Corporate slide presentations. IEEE Computer Graphics and Applications, 3(4):39–40, 42, 44, July/August 1983. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Wright:1990:PBL

[Wri90] William E. Wright. Parallelization of Bresenham’s line and circle algorithms.

IEEE Computer Graphics and Applications, 10(5):60– 67, September 1990. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Wright:1997:BVA

[Wri97] William Wright. Business visualization applications. IEEE Computer Graphics and Applications, 17(4): 66–70, July/August 1997. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Wright:1998:SNB

[Wri98]  William Wright.  Short

note: Business visualization adds value. IEEE Computer Graphics and Applications, 18(4):39, July/ August 1998. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

computer.org/cga/cg1998/
g4039abs.htm.

Wang:1990:VDP

[WS90a] Sue-Ling Chen Wang and John Staudhammer. Visibility determination on projected grid surfaces. IEEE Computer Graphics and Applications, 10(4):36–43, July/August 1990. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Wilhelms:1990:IBA

[WS90b] Jane Wilhelms and Robert Skinner. A “Notion” for interactive behavioral animation control. IEEE Computer Graphics and Applications, 10(3):14–22, May/ June 1990. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Watanabe:1992:TPM

[WS92] Yasuhiko Watanabe and Yasuhito Suenaga. A trigonal prism-based method for hair image generation. IEEE Computer Graphics and Applications, 12(1):47– 53, January/February 1992. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Warren:2004:FAS

[WS04] Joe Warren and Scott Schaefer. A factored approach to subdivision surfaces. IEEE Computer Graphics and Applications, 24(3):74–81, May/June 2004. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cg/2004/03/g3074abs.htm;
 http://csdl.computer.
org/dl/mags/cg/2004/03/
g3074.htm;   http://csdl.
computer.org/dl/mags/cg/
2004/03/g3074.pdf.

Wagner:2009:PVMa

[WS09a] Daniel Wagner and Dieter Schmalstieg. Projects in VR: Making augmented reality practical on mobile phones, Part 1. IEEE Computer Graphics and Applications, 29(3):12–15, May/ June 2009. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Wagner:2009:PVMb

[WS09b] Daniel Wagner and Dieter Schmalstieg. Projects in VR: Making augmented reality practical on mobile phones, Part 2. IEEE Computer Graphics and Applications, 29(4):6–9, July/ August 2009. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Wong:2012:TCE

[WSJ^{+}12] Pak Chung Wong, Han-Wei Shen, Christopher R. Johnson, Chaomei Chen, and Robert B. Ross. The top 10 challenges in extreme-scale visual analytics. IEEE Computer Graphics and Applications, 32(4):63–67, July/ August 2012. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Wei:2000:VRL

[WSK^{+}00] Bin Wei, Claudio Silva, Eleftherios Koutsofios, Shankar Krishnan, and Stephen North. Visualization research with large displays. IEEE Computer Graphics and Applications, 20(4):50– 54, July/August 2000. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

computer.org/cga/cg2000/
g4050abs.htm;   http:/
/dlib.computer.org/cg/
books/cg2000/pdf/g4050.
pdf.

Wang:2019:ERV

[WSK^{+}19] Y. Wang, A. Segal, R. Klatzky, D. F. Keefe, P. Isenberg, J. Hurtienne, E. Hornecker, T. Dwyer, and S. Barrass. An emotional response to the value of visualization. IEEE Computer Graphics and Applications, 39(5):8–17, September/October 2019. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Woo:2009:LPM

[WSKY09] Jeong-Ho Woo, Ju-Ho Sohn, Hyejung Kim, and Hoi-Jun Yoo. A low-power multimedia SoC with fully programmable 3D graphics for mobile devices. IEEE Computer Graphics and Applications, 29(5):82–90, September/October 2009. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Windhager:2020:MVE

[WSL^{+}20] F. Windhager, S. Salisu, R. A. Leite, V. Filipov, S. Miksch, G. Schreder, and E. Mayr. Many views are not enough: Designing for synoptic insights in cultural collections. IEEE Computer Graphics and Applications, 40(3):58–71, May/ June 2020. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Wong:2012:ESV

[WSP12] Pak Chung Wong, Han-Wei Shen, and Valerio Pascucci. Extreme-scale visual analytics. IEEE Computer Graphics and Applications, 32(4): 23–25, July/August 2012. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Walny:2020:PSP

[WSP^{+}20] J. Walny, S. Storteboom, R. Pusch, S. M. Hwang, S. Knudsen, S. Carpendale, and W. Willett. Pixel-Clipper: Supporting public engagement and conversation about visualizations. IEEE Computer Graphics and Applications, 40(2): 57–70, March/April 2020. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Ware:2023:RCA

[WSS23] Colin Ware, Maureen Stone, and Danielle Albers Szafir. Rainbow colormaps are not all bad. IEEE Computer Graphics and Applications, 43(3):88–93, May/ June 2023. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Wischgoll:2023:MTV

[WSSS23] Thomas Wischgoll, André Stork, Herbert Schilling, and Gerik Scheuermann. Metaverse: Technologies for virtual worlds. IEEE Computer Graphics and Applications, 43(2):11–12, March/ April 2023. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Weidlich:2008:AUV

[WSW08] Dieter Weidlich, Sandra Scherer, and Markus Wabner. Analyses using VR/AR visualization. IEEE Computer Graphics and Applications, 28(5):84–86, September/October 2008. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Wong:2004:VA

[WT04] Pak Chung Wong and Jim Thomas. Visual analytics. IEEE Computer Graphics and Applications, 24(5):20–21, September/ October 2004. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cg/2004/05/g5020.pdf;
http://csdl.computer.org/
dl/mags/cg/2004/05/g5020.
htm.

Wang:2015:DEP

[WTLM15] Shimin Wang, Yuzuru Tanahashi, Nick Leaf, and Kwan-Liu Ma. Design and effects of personal visualizations. IEEE Computer Graphics and Applications, 35(4):82–93, July/ August 2015. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
cg/2015/04/mcg2015040082-
abs.html.

Wolter:2009:SIT

[WTPHK09] Marc Wolter, Irene Tedjo-Palczynski, Bernd Hentschel, and Torsten Kuhlen. Spatial input for temporal navigation in scientific visualizations. IEEE Computer Graphics and Applications, 29(6):54–64, November/December 2009. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Wu:2011:NCP

[WTR11] Xiaomao Wu, Maxime Tournier, and Lionel Reveret.

Natural character posing from a large motion database. IEEE Computer Graphics and Applications, 31(3):69–77, May/ June 2011. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Wu:2018:TMD

[WVL^{+}18] Shin-Ting Wu, Raphael Voltoline, Wallace Loos, Jose Angel Ivan Rubianes Silva, Lionis Watanabe, Barbara Amorim, Ana Coan, Fernando Cendes, and Clarissa L. Yasuda. Toward a multimodal diagnostic exploratory visualization of focal cortical dysplasia. IEEE Computer Graphics and Applications, 38(3):73–89, May/ June 2018. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2018/03/mcg2018030073-
abs.html.

Wang:1986:GMS

[WW86] W. P. Wang and K. K.

Wang. Geometric modeling for swept volume of moving solids. IEEE Computer Graphics and Applications, 6(12):8–17, December 1986. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

World:1995:NRC

[WW95] Linda World and Karen Whitehouse. In the news: The reality of cybersickness; political, technological changes prompt changes in US patent law; Encarnação wins Coons Award; ETOMs: Cheap, fast optical storage. IEEE Computer Graphics and Applications, 15 (5):95–96, September 1995. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Wiley:2007:RIS

[WW07] Keith Wiley and Lance R. Williams. Representing interwoven surfaces in 2-1/2D drawings. IEEE Computer Graphics and Applications, 27(4):70–83, July/ August 2007. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Watson:2008:PMU

[WW08] Benjamin Watson and Peter Wonka. Procedural methods for urban modeling. IEEE Computer Graphics and Applications, 28(3):16–17, May/June 2008. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL http://csdl.
                                        computer.org/comp/mags/
                                        cg/2008/03/mcg2008030016.
                                        pdf.

Wu:2024:PRT

[WW24] Jian Wu and Lili Wang.

Panoramic ray tracing for interactive mixed reality rendering based on 360^{∘} RGBD video. IEEE Computer Graphics and Applications, 44(1):62–75, January/ February 2024. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Wells:1991:DTG

[WWC91] Stuart C. Wells, Grant J.

Williamson, and Susan E. Carrie. Dithering for 12-bit true-color graphics. IEEE Computer Graphics and Applications, 11(5):18–29, September 1991. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Wang:2000:AEY

[WWF00] Lifeng Wang, Christine Wang, and Alain Fournier. Applications: Envisioning Yuan Ming Yuan (Garden of Perfect Brightness). IEEE Computer Graphics and Applications, 20(1):10–14, January/February 2000. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http:

//dlib.computer.org/cg/
books/cg2000/pdf/g1010.
pdf.

Wesche:1997:PVV

[WWG97] Gerold Wesche, Jürgen Wind, and Martin Göbel. Projects in VR: Visualization on the Responsive Workbench. IEEE Computer Graphics and Applications, 17(4):10–12, July/ August 1997. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Wu:2021:SBP

[WWL^{+}21] Zhongke Wu, Xingce Wang, Shaolong Liu, Quan Chen, Hock Soon Seah, and Feng Tian. Skeleton-based parametric 2-D region representation: Disk B-spline curves. IEEE Computer Graphics and Applications, 41(3):59– 70, May/June 2021. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Wyvill:1987:STS

[WWM87] Geoff Wyvill, Brian Wyvill, and Craig McPheeters. Solid texturing of soft objects. IEEE Computer Graphics and Applications, 7(12):20– 26, December 1987. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Wingrave:2010:WBS

[WWV^{+}10] Chadwick A. Wingrave, Brian Williamson, Paul D. Varcholik, Jeremy Rose, Andrew Miller, Emiko Charbonneau, Jared Bott, and Joseph J. LaViola¡ Jr. The Wiimote and beyond: Spatially convenient devices for 3D user interfaces. IEEE Computer Graphics and Applications, 30(2):71–85, March/April 2010. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Wen:2023:DDB

[WWW^{+}23] Xiaolin Wen, Yong Wang, Meixuan Wu, Fengjie Wang, Xuanwu Yue, Qiaomu Shen, Yuxin Ma, and Min Zhu. DiffSeer: Difference-based dynamic weighted graph visualization. IEEE Computer Graphics and Applications, 43(3):12–23, May/ June 2023. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Wall:2022:PVG

[WXK22] Emily Wall, Cindy Xiong, and Yea-Seul Kim. VisHikers — guide to evaluation: Competing considerations in study design. IEEE Computer Graphics and Applications, 42(3):29–38, May/ June 2022. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Wei:2012:VAP

[WYG^{+}12] Jishang Wei, Hongfeng Yu, Ray W. Grout, Jacqueline H. Chen, and Kwan-Liu Ma. Visual analysis of particle behaviors to understand combustion simulations. IEEE Computer Graphics and Applications, 32(1):22–33, January/ February 2012. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Wang:2021:DDL

[WYHS21] Xiaoqi Wang, Kevin Yen, Yifan Hu, and Han-Wei Shen. DeepGD: a deep learning framework for graph drawing using GNN. IEEE Computer Graphics and Applications, 41(5): 32–44, September/October 2021. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Wang:2010:ADC

[WYM10] Chaoli Wang, Hongfeng Yu, and Kwan-Liu Ma. Application-driven compression for visualizing largescale time-varying data. IEEE Computer Graphics and Applications, 30(1):59– 69, January/February 2010. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Wang:2021:KKS

[WZL^{+}21] Jin Wang, Xinwei Zhang, Guodong Lu, Jianhui Fu, Jinsong Xiao, and Xusheng Zhang. Kineticist: Kinetic sculpture design using multilevel skeletons. IEEE Computer Graphics and Applications, 41(5):113–123, September/October 2021. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Wang:2011:CBB

[WZM^{+}11] Yanhui Wang, Liqiang Zhang, Jingtao Ma, Liu Liu, Dongqin You, and Lixin Zhang. Combining building and behavior models for evacuation planning. IEEE Computer Graphics and Applications, 31(3):42–55, May/June 2011. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Wu:2019:ADS

[WZQ^{+}19] Zizhao Wu, Ming Zeng, Feiwei Qin, Yigang Wang, and Jiří Kosinka. Active 3-D shape cosegmentation with graph convolutional networks. IEEE Computer Graphics and Applications, 39(2):77–88, March/ April 2019. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL https:

//ieeexplore.ieee.org/
document/8611124/.

Xu:2015:APS

[XAJK^{+}15] Kai Xu, Simon Attfield, T. J. Jankun-Kelly, Ashley Wheat, Phong H. Nguyen, and Nallini Selvaraj. Analytic provenance for sensemaking: A research agenda. IEEE Computer Graphics and Applications, 35(3):56–64, May/ June 2015. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
cg/2015/03/mcg2015030056-
abs.html.

Xie:2023:UCG

[XDJ23] Charles Xie, Xiaotong Ding, and Rundong Jiang. Using computer graphics to make science visible in engineering education. IEEE Computer Graphics and Applications, 43(5):99–106, September/October 2023. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Xia:2017:VRT

[XHC^{+}17] Jing Xia, Yumeng Hou, Yingjie Victor Chen, Zhenyu Cheryl Qian, David S. Ebert, and Wei Chen. Visualizing rank time series of Wikipedia topviewed pages. IEEE Computer Graphics and Applications, 37(2):42–53, March/ April 2017. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2017/02/mcg2017020042-
abs.html.

Xiang:1994:CIQ

[XJ94] Zhigang Xiang and Gregory Joy. Color image quantization by agglomerative clustering. IEEE Computer Graphics and Applications, 14(3):44–48, May/ June 1994. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Xian:2018:ECW

[XJW18] Chuhua Xian, Shuo Jin, and Charlie C. L. Wang. Efficient C2-weighting for image warping. IEEE Computer Graphics and Applications, 38(1):59–76, January/ February 2018. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2018/01/mcg2018010059-
abs.html.

Xia:2021:VCF

[XLJ^{+}21] Jiazhi Xia, Weixing Lin, Guang Jiang, Yunhai Wang, Wei Chen, and Tobias Schreck. Visual clustering factors in scatterplots. IEEE Computer Graphics and Applications, 41(5): 79–89, September/October 2021. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Xu:2008:NM

[XMJ08] Jiayi Xu, Xiaoyang Mao, and Xiaogang Jin. Nondissipative marbling. IEEE Computer Graphics and Applications, 28(2):35–43, March/April 2008. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Xie:2019:MBC

[XNSW19] W. Xie, Y. Nie, Z. Song, and C. C. L. Wang. Meshbased computation for solving photometric stereo with near point lighting. IEEE Computer Graphics and Applications, 39(3):73–85, May/June 2019. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Xu:2005:NMC

[XP05] Ruifeng Xu and Sumanta N. Pattanaik. A novel Monte Carlo noise reduction operator. IEEE Computer Graphics and Applications, 25(2): 31–35, March/April 2005. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Xu:2005:HDR

[XPH05] Ruifeng Xu, Sumanta N. Pattanaik, and Charles E. Hughes. High-dynamic-range still-image encoding in JPEG 2000. IEEE Computer Graphics and Applications, 25(6):57–64, November/December 2005. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Xu:2020:TTS

[XSN^{+}20] K. Xu, S. Salisu, P. H.

Nguyen, R. Walker, B. L. W. Wong, A. Wagstaff, G. Phillips, and M. Biggs. Time-Sets: Temporal sensemaking in intelligence analysis. IEEE Computer Graphics and Applications, 40 (3):83–93, May/June 2020. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Xu:2001:VHI

[XY01] Zhan Xu and Xue Dong Yang. V-HairStudio: An interactive tool for hair design. IEEE Computer Graphics and Applications, 21(3):36– 43, May/June 2001. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

computer.org/cga/cg2001/
g3036abs.htm;   http:/
/dlib.computer.org/cg/
books/cg2001/pdf/g3036.
pdf.

Xing:2014:HPA

[XZWH14] Lianping Xing, Xiaoting Zhang, Charlie C. L. Wang, and Kin-Chuen Hui. Highly parallel algorithms for visual-perception-guided surface remeshing. IEEE Computer Graphics and Applications, 34(1):52–64, January/February 2014. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Xu:2019:IPR

[XZZL19] Cheng Xu, Cheng Zhang, Xiaochen Zhou, and Biao Leng. Improved panoramic representation via bidirectional recurrent view aggregation for three-dimensional model retrieval. IEEE Computer Graphics and Applications, 39(2):65–76, March/ April 2019. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL https:

                                        //ieeexplore.ieee.org/
                                        document/8565889/.

Yan:1985:ACI

[Yan85] Johnson K. Yan. Advances in computer-generated imagery for flight simulation. IEEE Computer Graphics and Applications, 5(8):37– 51, August 1985. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Yan:2001:IAD

[Yan01] Hong Yan. Image analysis for digital media applications. IEEE Computer Graphics and Applications, 21(1):18–26, January/ February 2001. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

                                        computer.org/cga/cg2001/

g1018abs.htm;   http:/
/dlib.computer.org/cg/
books/cg2001/pdf/g1018.
pdf.

Yan:2021:RRD

[Yan21] Ling-Qi Yan. Realistic rendering in details. IEEE Computer Graphics and Applications, 41(4):20–26, July/August 2021. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Yoo:2012:VCH

[YBL^{+}12] Terry S. Yoo, Donald Bliss, Bradley C. Lowekamp, David T. Chen, Gavin E. Murphy, Kedar Narayan, Lisa M. Hartnell, Thao Do, and Sriram Subramaniam. Visualizing cells and humans in 3D: Biomedical image analysis at nanometer and meter scales. IEEE Computer Graphics and Applications, 32(5):39–49, September/October 2012. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Yang:2005:NPP

[YCB05] Yonggao Yang, Jim X.

Chen, and Mohsen Beheshti. Nonlinear perspective projections and magic lenses: 3D view deformation. IEEE Computer Graphics and Applications, 25(1):76–84, January/February 2005. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Yuan:2022:SVA

[YCB^{+}22] Jun Yuan, Gromit Yeuk-Yin Chan, Brian Barr, Kyle Overton, Kim Rees, Luis Gustavo Nonato, Enrico Bertini, and Claudio T. Silva. SUBPLEX: a visual analytics approach to understand local model explanations at the subpopulation level. IEEE Computer Graphics and Applications, 42(6):24–36, November/December 2022. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Yagel:1992:DRT

[YCK92] Roni Yagel, Daniel Cohen, and Arie Kaufman. Discrete ray tracing. IEEE Computer Graphics and Applications, 12(5):19–??, September 1992. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Yuan:2015:HCD

[YCKF15] Xiaoru Yuan, Baoquan Chen, Koji Koyamada, and Issei Fujishiro. Humancentered data visualization [Guest Editors’ introduction]. IEEE Computer Graphics and Applications, 35(6):18–19, November/ December 2015. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
cg/2015/06/mcg2015060018.
html.

Yeary:1997:ACM

[Yea97a] Robin Yeary. About the cover: Making the unseen visible. IEEE Computer Graphics and Applications, 17(4):4–5, July/ August 1997. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Yeary:1997:AS

[Yea97b]  Robin Yeary.  The art

of science. IEEE Computer Graphics and Applications, 17(2):4–5, March/ April 1997. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Yilmaz:2009:VMP

[YIÇ09] Erdal Yilmaz, Veysi Isler, and Yasemin Yardimci Çetin. The virtual marathon: Parallel computing supports crowd simulations. IEEE Computer Graphics and Applications, 29(4), July/ August 2009. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Yamaguchi:1985:DFL

[YIK85] Kazunori Yamaguchi, Naota Inamoto, and Tosiyasu Kunii. A data flow language for controlling multiple interactive devices. IEEE Computer Graphics and Applications, 5(3):48–60, March/ April 1985. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Yamaguchi:1995:NTB

[YK95] Yasushi Yamaguchi and Fumihiko Kimura. Nonmanifold topology based on coupling entities. IEEE Computer Graphics and Applications, 15(1):42–50, January/ February 1995. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Yamaguchi:1984:ODS

[YKFT84] K. Yamaguchi, T. L. Kunii, K. Fujimura, and H. Toriya. Octree-related data structures and algorithms. IEEE Computer Graphics and Applications, 4(1):53–59, January/February 1984. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Yamaguchi:1984:CMS

[YKR^{+}84] K. Yamaguchi, T. L. Kunii, David F. Rogers, Steven G. Satterfield, and Francisco A. Rodriguez. Computerintegrated manufacturing of surfaces using octree encoding. IEEE Computer Graphics and Applications, 4(1):60–62, 64–65, January/ February 1984. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Yoo:2009:CMF

[YL09] Min-Joon Yoo and In-Kwon Lee. Creating musicalfountain shows. IEEE Computer Graphics and Applications, 29(5):6–13, September/October 2009. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Yoo:2011:AEB

[YL11] Min-Joon Yoo and In-Kwon Lee. Affecticon: Emotion-based icons for music retrieval. IEEE Computer Graphics and Applications, 31(3):89–95, May/ June 2011. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Yang:2010:NIV

[YLL10] Jing Yang, Dongning Luo, and Yujie Liu. Newdle: Interactive visual exploration of large online news collections. IEEE Computer Graphics and Applications, 30(5):32–41, September/October 2010. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Yoon:2017:BFD

[YLR17] Seung-Hyun Yoon, John Lewis, and Taehyun Rhee. Blending face details: Synthesizing a face using multiscale face models. IEEE Computer Graphics and Applications, 37(6):65–75, November/December 2017. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL

https://www.computer.org/
csdl/mags/cg/2017/06/mcg2017060065-
abs.html.

Ynnerman:2018:ENS

[YLT18] Anders Ynnerman, Jonas Lowgren, and Lena Tibell. Exploranation: A new science communication paradigm. IEEE Computer Graphics and Applications, 38(3):13– 20, May/June 2018. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL https:/

/www.computer.org/csdl/
mags/cg/2018/03/mcg2018030013-
abs.html.

Yamrom:1995:VFA

[YM95] Boris Yamrom and Kenneth M. Martin. Vector field animation with texture maps. IEEE Computer Graphics and Applications, 15(2):22–24, March/ April 1995. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Yoshida:1993:CPS

[YMDO93] Ryo Yoshida, Tatsuo Miyazawa, Akio Doi, and Taisuke Otsuki. Clinical Planning Support System — CliPSS. IEEE Computer Graphics and Applications, 13(6): 76–84, November/December 1993. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Yu:2012:VRE

[YMK^{+}12] Insu Yu, Jesper Mortensen, Pankaj Khanna, Bernhard Spanlang, and Mel Slater. Visual realism enhances realistic response in an immersive virtual environment. Part 2. IEEE Computer Graphics and Applications, 32(6):36–45, November/December 2012. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Yu:2001:IRU

[YMS01] Xiaohua Yu, Bryan S.

Morse, and Thomas W. Sederberg. Image reconstruction using datadependent triangulation. IEEE Computer Graphics and Applications, 21(3):62– 69, May/June 2001. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

computer.org/cga/cg2001/
g3062abs.htm;   http:/
/dlib.computer.org/cg/

                                        books/cg2001/pdf/g3062.
                                        pdf.

You:1999:OTO

[YNA99] Suya You, Ulrich Neumann, and Ronald Azuma. Orientation tracking for outdoor augmented reality registration. IEEE Computer Graphics and Applications, 19(6):36–42, November/December 1999. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

                                        computer.org/cga/cg1999/
                                        g6036abs.htm;   http:/
                                        /dlib.computer.org/cg/
                                        books/cg1999/pdf/g6036.
                                        pdf.

Yoo:1992:DVV

[YNF^{+}92] Terry S. Yoo, Ulrich Neumann, Henry Fuchs, Stephen M. Pizer, and Tim Cullip. Direct visualization of volume data. IEEE Computer Graphics and Applications, 12(4):63–71, July/ August 1992. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Yasuda:2004:TKH

[YNH04] Kazutaka Yasuda, Takeshi Naemura, and Hiroshi Harashima. Thermo-Key: Human region segmentation from video. IEEE Computer Graphics and Applications, 24(1):26–30, January/ February 2004. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cg/2004/01/g1026abs.htm;
http://csdl.computer.
org/dl/mags/cg/2004/01/
g1026.pdf.

Yang:2008:AMT

[YP08] Chuan-Kai Yang and Li-Kai Peng. Automatic moodtransferring between color images. IEEE Computer Graphics and Applications, 28(2):52–61, March/April 2008. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Yerry:1983:MQA

[YS83] Mark A. Yerry and Mark S.

Shephard. A modified quadtree approach to finite element mesh generation. IEEE Computer Graphics and Applications, 3(1):39– 46, January/February 1983. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Yoshida:2016:RDE

[YSS^{+}16] Shigeo Yoshida, Takumi Shirokura, Yuta Sugiura, Daisuke Sakamoto, Tetsuo Ono, Masahiko Inami, and Takeo Igarashi. RoboJockey: Designing an entertainment experience with robots. IEEE Computer Graphics and Applications, 36(1):62–69, January/ February 2016. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
cg/2016/01/mcg2016010062-
abs.html.

Yen:1991:PBB

[YSSP91] Jonathan Yen, Susan Spach, Mark T. Smith, and Ronald W. Pulleyblank. Parallel boxing in B-spline intersection. IEEE Computer Graphics and Applications, 11(1):72– 79, January/February 1991. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Yamada:1984:LDS

[YST84] Jiro Yamada, N. Saito, and A. Tamura. A low-cost drafting system based on a personal computer. IEEE Computer Graphics and Applications, 4(5):61–65, May/ June 1984. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Yuan:2023:SSN

[YSW^{+}23] Jiang Yuan, Chen Shanxiong, Gao Weize, Peng Maling, and Jiang Lihua. SFF-Siam: a new oracle bone rejoining method based on siamese network. IEEE Computer Graphics and Applications, 43(6):22–32, November/December 2023. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Yang:2008:ISR

[YSWW08] Gang Yang, Hanqiu Sun, Enhua Wu, and Lifeng Wang. Interactive fur shaping and rendering using nonuniform-layered textures. IEEE Computer Graphics and Applications, 28(4):85–93, July/August 2008. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Yamaguchi:1984:UAB

[YT84] Fujio Yamaguchi and Toshiya Tokieda. A unified algorithm for Boolean shape operations. IEEE Computer Graphics and Applications, 4(6):24–37, June 1984. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Yamaguchi:1985:SMD

[YT85] Fujio Yamaguchi and Toshiya Tokieda. A solid modeler with a 4 × 4 determinant processor. IEEE Computer Graphics and Applications, 5(4):51–59, April 1985. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Yu:2010:SVL

[YWG^{+}10] Hongfeng Yu, Chaoli Wang, Ray W. Grout, Jacqueline H. Chen, and Kwan-Liu Ma. In situ visualization for large-scale combustion simulations. IEEE Computer Graphics and Applications, 30(3):45–57, May/ June 2010. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Yin:2009:TGB

[YWR09] Xuetao Yin, Peter Wonka, and Anshuman Razdan. Tutorial: Generating 3D building models from architectural drawings: a survey. IEEE Computer Graphics and Applications, 29(1):20– 30, January/February 2009. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Yeo:1999:WOV

[YY99] Boon-Lock Yeo and Minerva M. Yeung. Watermarking 3D objects for verification. IEEE Computer Graphics and Applications, 19(1):36–45, January/ February 1999. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

computer.org/cga/cg1999/
g1036abs.htm;   http:/
/dlib.computer.org/cg/
books/cg1999/pdf/g1036.
pdf.

Yanaka:2016:IDC

[YY16] Kazuhisa Yanaka and Toshiaki Yamanouchi. 3D image display courses for information media students. IEEE Computer Graphics and Applications, 36(2):68–73, March/April 2016. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2016/02/mcg2016020068-
abs.html.

Yasuda:1992:SMC

[YYiTI92] Takami Yasuda, Shigeki Yokoi, Jun ichiro Toriwaki, and Katsuhiko Inagaki. A shading model for cloth objects. IEEE Computer Graphics and Applications, 12(6):15–24, November/December 1992. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Yasuda:1992:VAE

[YYOiT92] Takami Yasuda, Shigeki Yokoi, Hiroshi Ohshita, and Jun ichiro Toriwaki. 3D visualization of an ancient Egyptian mummy. IEEE Computer Graphics and Applications, 12(3):13–17, May/June 1992. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Yang:2011:ESL

[YZM^{+}11] Ling Yang, Liqiang Zhang, Jingtao Ma, Zhizhong Kang, Lixin Zhang, and Jonathan Li. Efficient simplification of large vector maps rendered onto 3D landscapes.

IEEE Computer Graphics and Applications, 31(2): 14–23, March/April 2011. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Zachrisen:1984:ASB

[Zac84] Morten Zachrisen. Adding structure to bit-map displays. IEEE Computer Graphics and Applications, 4(7):47–51, July/August 1984. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Zheng:2014:UGI

[ZCBM14] Lin Zheng, Abhijit J. Chaudhari, Ramsey D. Badawi, and Kwan-Liu Ma. Using global illumination in volume visualization of rheumatoid arthritis CT data. IEEE Computer Graphics and Applications, 34(6):16–23, November/ December 2014. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL http://csdl.
                                        computer.org/csdl/mags/
                                        cg/2014/06/mcg2014060016-
                                        abs.html.

Zhang:2006:VSH

[ZCS^{+}06] Keqi Zhang, Shu-Ching Chen, Peter Singh, Khalid Saleem, and Na Zhao. A 3D visualization system for hurricane storm-surge flooding. IEEE Computer Graphics and Applications, 26(1):18– 25, January/February 2006.

CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Zyda:1998:PVN

[ZD98] Michael Zyda and Rudolph P.

Darken. Projects in VR: The Naval Postgraduate School’s Moves curriculum. IEEE Computer Graphics and Applications, 18(3):8–11, May/ June 1998. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://dlib.
computer.org/cg/books/
cg1998/pdf/g3008.pdf.

Zhao:2023:SWT

[ZE23] Zhenpeng Zhao and Niklas Elmqvist. The stories we tell about data: Surveying datadriven storytelling using visualization. IEEE Computer Graphics and Applications, 43(4):97–110, July/ August 2023. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Zielke:2009:SGI

[ZED^{+}09] Marjorie A. Zielke, Monica J. Evans, Frank Dufour, Timothy V. Christopher, Jumanne K. Donahue, Phillip Johnson, Erin B. Jennings, Brent S. Friedman, Phonesury L. Ounekeo, and Ricardo Flores. Serious games for immersive cultural training: Creating a living world.

IEEE Computer Graphics and Applications, 29(2): 49–60, March/April 2009. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Zeltzer:1982:MCT

[Zel82] David Zeltzer. Motor control techniques for figure animation. IEEE Computer Graphics and Applications, 2(9):53–59, November/December 1982. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Zacharatos:2014:AER

[ZGC14] Haris Zacharatos, Christos Gatzoulis, and Yiorgos L. Chrysanthou. Automatic emotion recognition based on body movement analysis: A survey. IEEE Computer Graphics and Applications, 34(6):35–45, November/ December 2014. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://csdl.
computer.org/csdl/mags/
cg/2014/06/mcg2014060035-
abs.html.

Zhang:2017:VTS

[ZGL^{+}17] Qian Zhang, Yu Guo, Pierre-Yves Laffont, Tobias Martin, and Markus Gross. A virtual tryon system for prescription eyeglasses. IEEE Computer Graphics and Applications, 37(4):84–93, July/ August 2017. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2017/04/mcg2017040084-
abs.html.

Zhu:2013:EDP

[ZGP^{+}13] Yufeng Zhu, Pradeep Garigipati, Pieter Peers, Paul Debevec, and Abhijeet Ghosh. Estimating diffusion parameters from polarized spherical-gradient illumination. IEEE Computer Graphics and Applications, 33(3):34–43, May/ June 2013. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Zhang:1996:VBV

[Zha96a]  Ping Zhang.  Visualiza-

tion blackboard: Visualizing production planning data. IEEE Computer Graphics and Applications, 16(5):7–10, September 1996. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Zhang:1996:FAD

[Zha96b] Yuefeng Zhang. A fuzzy approach to digital image warping. IEEE Computer Graphics and Applications, 16(4):34–41, July/August 1996. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Zhang:2007:API

[Zha07] Kang Zhang. From abstract painting to information visualization. IEEE Computer Graphics and Applications, 27(3):12–16, May/ June 2007. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Zhang:2021:DVE

[Zha21] R. Zhang. Design and visualization for exploring realworld data. IEEE Computer Graphics and Applications, 41(2):23–24, March/ April 2021. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Zhang:2022:BPA

[Zha22a] Richard Zhang. Best paper awards 2021. IEEE Computer Graphics and Applications, 42(6):7–8, November/ December 2022. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Zhang:2022:CGV

[Zha22b] Richard Zhang. Computer graphics and VR/AR: From clouds, to cartoons, to a 3-D virtual museum. IEEE Computer Graphics and Applications, 42(5):7, September/October 2022. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Zhang:2022:NTT

[Zha22c] Richard Zhang. The new test of time award. IEEE Computer Graphics and Applications, 42(2):89, March/ April 2022. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Zhang:2022:VAP

[Zha22d] Richard Zhang. Visual analysis and processing of diverse data. IEEE Computer Graphics and Applications, 42(4):27, July/August 2022. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Zyda:2003:ERD

[ZHM^{+}03] Michael Zyda, John Hiles, Alex Mayberry, Casey Wardynski, Michael Capps, Brian Osborn, Russell Shilling, Martin Robaszewski, and Margaret Davis. Entertainment R&D for defense. IEEE Computer Graphics and Applications, 23(1):28– 36, January/February 2003. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL

http://dlib.computer.org/
cg/books/cg2003/pdf/g1028.
pdf;   http://www.computer.
org/cga/cg2003/g1028abs.
htm.

Zhu:2012:IGC

[Zhu12]  Ying Zhu.  Introduc-

ing Google Chart Tools and Google Maps API in data visualization courses. IEEE Computer Graphics and Applications, 32(6):6–9, November/December 2012. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Zhu:2019:LSM

[Zhu19] Jun-Yan Zhu. Learning to synthesize and manipulate natural images. IEEE Computer Graphics and Applications, 39(2):14–23, March/ April 2019. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL https:

//ieeexplore.ieee.org/
document/8672983/.

Zhou:2015:EVA

[ZHZ^{+}15] Fangfang Zhou, Wei Huang, Ying Zhao, Yang Shi, Xing Liang, and Xiaoping Fan. ENTVis: A visual analytic tool for entropy-based network traffic anomaly detection. IEEE Computer Graphics and Applications, 35(6):42–50, November/ December 2015. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
cg/2015/06/mcg2015060042-
abs.html.

Zintel:2008:TPa

[Zin08a]  Ed Zintel.  Tools and

products. IEEE Computer Graphics and Applications, 28(3):104, 103, May/ June 2008. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://bell.
computer.org/dlcomments/
;   http://csdl.computer.
org/comp/mags/cg/2008/
03/mcg2008030104.pdf.

Zintel:2008:TPb

[Zin08b]  Ed Zintel.  Tools and

products. IEEE Computer Graphics and Applications, 28(4):103–104, July/ August 2008. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://bell.
computer.org/dlcomments/

.

Zintel:2008:TPc

[Zin08c]  Ed Zintel.  Tools and

products. IEEE Computer Graphics and Applications, 28(5):87–88, September/October 2008. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://bell.
computer.org/dlcomments/
;   http://www.adobe.com/
products/acrobat/readstep2.
html.

Zintel:2008:TPd

[Zin08d]  Ed Zintel.  Tools and

products. IEEE Computer Graphics and Applications, 28(6):14–17, November/December 2008. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Zhao:2018:PRP

[ZJH^{+}18] Haiming Zhao, Xiaogang Jin, Xiaojian Huang, Menglei Chai, and Kun Zhou. Parametric reshaping of portrait images for weightchange. IEEE Computer Graphics and Applications, 38(1):77–90, January/ February 2018. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        cg/2018/01/mcg2018010077-
                                        abs.html.

Ziemkiewicz:2010:BBS

[ZK10] Caroline Ziemkiewicz and Robert Kosara. Beyond Bertin: Seeing the forest despite the trees. IEEE Computer Graphics and Applications, 30(5):7–11, September/October 2010. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Zarsky:2022:ETC

[ŽLK22a] Jiří Žárský, Gaetan Lopez, and Tomáš Kliegr. Explainability of text clustering visualizations Twitter disinformation case study. IEEE Computer Graphics and Applications, 42(4):8–19, July/ August 2022. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Zotos:2022:DWC

[ZLK^{+}22b] Savvas Zotos, Marilena Lemonari, Michael Konstantinou, Anastasios Yiannakidis, Georgios Pappas, Panayiotis Kyriakou, Ioannis N. Vogiatzakis, and Andreas Aristidou. Digitizing wildlife: The case of a reptile 3-D virtual museum. IEEE Computer Graphics and Applications, 42(5): 51–65, September/October 2022. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Zhang:2019:ACS

[ZLMY19] Wei Zhang, Guanbin Li, Haoyu Ma, and Yizhou Yu. Automatic color sketch generation using deep style transfer. IEEE Computer Graphics and Applications, 39(2):26–37, March/ April 2019. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL https:

//ieeexplore.ieee.org/
document/8640099/.

Zhang:2020:BB

[ZM20] R. Zhang and T. Möller.

Best of the best. IEEE Computer Graphics and Applications, 40(6):8–11, November/December 2020. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Zyda:1988:FSU

[ZMR^{+}88] Michael J. Zyda, Robert B. McGhee, Ron S. Ross, Douglas B. Smith, and Dale G. Streyle. Flight simulators for under $100,000. IEEE Computer Graphics and Applications, 8(1):19– 27, January/February 1988. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Zhang:2015:TIA

[ZNMT15] Kang Zhang, Quang Vinh Nguyen, Zhao-Peng Meng, and Wen-Yuan Tao. Teaching information aesthetics as a research class in China. IEEE Computer Graphics and Applications, 35(1):56–64, January/ February 2015. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://csdl.
computer.org/csdl/mags/
cg/2015/01/mcg2015010056-
abs.html.

Ziemkiewicz:2012:UVU

[ZOC^{+}12] Caroline Ziemkiewicz, Alvitta Ottley, R. Jordan Crouser, Krysta Chauncey, Sara L. Su, and Remco Chang. Understanding visualization by understanding individual users. IEEE Computer Graphics and Applications, 32(6):88–94, November/December 2012. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Zollikofer:1995:TRP

[ZP95] Christoph P. E. Zollikofer and Marcia S. Ponce de Leon. Tools for rapid prototyping in the biosciences. IEEE Computer Graphics and Applications, 15(6): 48–55, November/December 1995. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Zeltzer:1995:VBT

[ZPA95] D. Zeltzer, N. J. Pioch, and W. A. Aviles. VR blackboard: Training the officer of the deck. IEEE Computer Graphics and Applications, 15(6):6–9, November/ December 1995. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Zhu:2021:IVH

[ZPG^{+}21] Feiyu Zhu, Yu Pan, Tian Gao, Harkamal Walia, and Hongfeng Yu. Interactive visualization of hyperspectral images based on neural networks. IEEE Computer Graphics and Applications, 41(5):57–66, September/October 2021. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Zikas:2023:MAW

[ZPL^{+}23] Paul Zikas, Antonis Protopsaltis, Nick Lydatakis, Mike Kentros, Stratos Geronikolakis, Steve Kateros, Manos Kamarianakis, Giannis Evangelou, Achilleas Filippidis, Eleni Grigoriou, Dimitris Angelis, Michail Tamiolakis, Michael Dodis, George Kokiadis, John Petropoulos, Maria Pateraki, and George Papagiannakis. MAGES 4.0: Accelerating the world s transition to VR training and democratizing the authoring of the medical metaverse. IEEE Computer Graphics and Applications, 43(2):43–56, March/ April 2023. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Zielasko:2024:SSV

[ZR24] Daniel Zielasko and Bernhard E. Riecke. Sitting or standing in VR: About comfort, conflicts, and hazards. IEEE Computer Graphics and Applications, 44(2): 81–88, March/April 2024. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Zingsheim:2021:CVB

[ZSK^{+}21] Domenic Zingsheim, Patrick Stotko, Stefan Krumpen, Michael Weinmann, and Reinhard Klein. Collaborative VR-based 3D labeling of live-captured scenes by remote users. IEEE Computer Graphics and Applications, 41(4):90–98, July/ August 2021. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Zyda:2007:ENG

[ZTJ^{+}07] Michael Zyda, Dhruv Thukral, Sumeet Jakatdar, Jonathan Engelsma, James Ferrans, Mat Hans, Larry Shi, Fred Kitson, and Venu Vasudevan. Educating the next generation of mobile game developers. IEEE Computer Graphics and Applications, 27(2):96, 92–95, March/April 2007. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://bell.
computer.org/dlcomments/

.

Zhou:1999:ERC

[ZTT99] Yong Zhou, Paul M. Thompson, and Arthur W. Toga. Extracting and representing the cortical sulci. IEEE Computer Graphics and Applications, 19(3):49–55, May/June 1999. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

computer.org/cga/cg1999/
g3049abs.htm;   http:/
/dlib.computer.org/cg/
books/cg1999/pdf/g3049.
pdf.

Zuckerman:1984:IDT

[Zuc84] Matthew M. Zuckerman. Innovative display technologies — why a flat panel when you can have a CRT? IEEE Computer Graphics and Applications, 4(4):9–15, April 1984. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Zhu:2019:FSI

[ZWWS19] L. Zhu, Z. Wang, Y. Wang, and A. Song. Follow the smoke: Immersive display of motion data with synthesized smoke. IEEE Computer Graphics and Applications, 39(4):86–94, July/ August 2019. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Zeng:2018:VVA

[ZY18] Wei Zeng and Yu Ye. VitalVizor: A visual analytics system for studying urban vitality. IEEE Computer Graphics and Applications, 38(5):38–53, September/October 2018. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        cg/2018/05/mcg2018050038-
                                        abs.html.

Zhou:2021:SNS

[ZYJ^{+}21] Zeyang Zhou, Zhiyong Yang, Shan Jiang, Xiaodong Ma, Fujun Zhang, and Huzheng Yan. Surgical navigation system for low-dose-rate brachytherapy based on mixed reality. IEEE Computer Graphics and Applications, 41(3):113–123, May/ June 2021. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Zhou:2017:VAS

[ZYL^{+}17] Zhiguang Zhou, Zhifei Ye, Yanan Liu, Fang Liu, Yubo Tao, and Weihua Su. Visual analytics for spatial clusters of air-quality data. IEEE Computer Graphics and Applications, 37(5):98–105, September/ October 2017. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2017/05/mcg2017050098-
abs.html.

Zheng:1999:AVR

[ZZ99] Jiang Yu Zheng and Zhong Li Zhang. Applications: Virtual recovery of excavated relics. IEEE Computer Graphics and Applications, 19(3):6–11, May/ June 1999. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL http://dlib.
computer.org/cg/books/
cg1999/pdf/g3006.pdf.

Zhang:2005:DLT

[ZZ05] Hua Zhang and Changqian Zhu. Deriving linear transformations in 3D using quaternion algebra. IEEE Computer Graphics and Applications, 25(2):93–95, March/April 2005. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).